A Complete Bibliography of Publications in Isis, 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.11

Title word cross-reference

-Ikos [2867]. -Ismos [2867]. -Logos [2867]. -theoretische [3715].

1 [354, 747, 651]. 1000 [2156]. 1200 [870]. 125 [756]. 1250-1270 [2729]. 1270 [2729]. 14 [501]. 1450 [3555]. 1520 [1359]. 1540/2 [2541]. 1542 [616]. 1543 [355]. 1550 [3232]. 16 [2607]. 1600 [586, 1173, 3389, 3729]. 1601 [703]. 1605-2005 [2745]. 1616-1703 [2731]. 1620 [1700]. 1625 [985]. 1630 [3313]. 1640 [2048, 2722, 786]. 1640-1940 [2722]. 1642 [1832]. 1659 [2731]. 1665 [1092]. 1666 [2739]. 1668 [2731]. 1670 [2341, 1245]. 1676 [1927]. 1680 [2491]. 1685 [3299]. 1687 [2040]. 1687-1727 [2742]. 17 [3402]. 1700 [3739, 987, 2538, 1889]. 1705 [3741]. 1715 [885]. 1716 [1921]. 1720 [1692, 1698]. 1727 [1707]. 1735- [3659]. 1736 [484]. 1740 [709]. 1745 [270]. 1750 [3496, 2746, 2273, 159]. 1750-90 [2746]. 1751 [3131]. 1759 [1787]. 1760 [486]. 1770 [3242]. 1772 [2257]. 1773 [446]. 1777 [2744, 1827]. 1777-1809 [2744]. 1780 [1102]. 1793 [2164]. 18 [3227, 2485, 3260]. 1800 [3072, 2268, 3662, 3317, 3155]. 1803 [797]. 1804 [2419]. 1805 [1512]. 1809 [2744]. 1819 [3050]. 1820 [989, 3241, 1831, 2482]. 1823 [1815]. 1824 [165]. 1827 [761, 275]. 1830 [2196, 1710, 2172, 3051, 3151, 2837, 1829, 3548, 3592]. 1835 [2688, 3584]. 1839-1939 [2752]. 1842 [1935, 804, 1208]. 1848 [2809, 1943]. 1850 [3598, 348, 853, 970, 500, 1388, 273, 3293, 2066, 2291]. 1851 [2980]. 1852 [2183]. 1853 [372]. 1854 [184]. 1855-1934 [2753]. 1856 [2898, 3674]. 1857 [912]. 1858 [3411]. 1860 [1009, 2851, 1610, 1125]. 1861 [1119]. 1863 [805]. 1867 [3169]. 1868 [1078]. 1870 [3303, 3745, 3382, 1678]. 1872 [1717]. 1875 [47, 1287, 1252]. 1877 [3759]. 1878 [720]. 1879 [2281]. 1880 [134]. 1884 [2769]. 1884-1920 [2769]. 1885 [3332, 3245]. 1886 [1718, 1722]. 1890 [2755]. 1890-2000 [2755]. 1890s [2390]. 1895 [632]. 19 [3101, 3500, 3679]. 1900 [905, 906, 636, 2990, 858, 278, 3075, 550, 3715, 2655, 1521, 3223]. 1900s [2478]. 1901 [2692]. 1902 [3250]. 1903 [373, 2680]. 1909 [1387]. 1910 [2832, 3754]. 1911 [181]. 1912 [899, 2069, 296, 1482, 1721]. 1914 [845, 131, 3748, 1124, 442, 337, 716, 2373, 814, 1553, 1381]. 1915 [2201, 1780, 3177]. 1917 [1354]. 1918 [1714, 400, 68]. 1919 [1847]. 1920 [3410, 2370, 2379, 766, 2769, 1285]. 1920s [2318]. 1921 [2751]. 1924 [914]. 1925 [291, 3416]. 1926 [1116, 178, 1728]. 1927 [2763, 212]. 1927-1945 [2763]. 1929 [3509]. 1930 [2434, 511, 1855, 1833, 504, 2836, 2780, 1942, 286, 2056, 387, 1008, 3664, 3172]. 1930-1970 [2780]. 1931 [1006, 3320]. 1933 [1615, 1716]. 1934 [2054, 2907, 2495]. 1937 [1544]. 1938 [1629, 1529, 3263]. 1939 [183, 2175, 2845, 341, 536, 820, 1526]. 1940 [855, 2722, 393, 1031, 1963, 2682, 851, 1936, 1383, 3170, 1611, 2165, 2625, 279, 1937, 1322, 3342, 2195, 1271, 3216]. 1940s [576]. 1942 [1605]. 1945 [176, 2763, 2695, 2770, 1026, 313, 2838, 2372, 1027, 777, 2423, 540, 860, 909, 2121, 2431, 1278, 1759, 1739, 1549, 3766]. 1945-1999 [2770]. 1946 [2782]. 1946-1957 [2782]. 1948 [74]. 1950 [852, 2998, 2918, 199, 2553, 1123, 200, 1348, 1860, 2079, 1533, 2694]. 1950s [1649, 2218, 1894]. 1951 [221]. 1952 [891, 1885]. 1953 [194, 93]. 1955 [2305]. 1957 [2782]. 1958 [2301, 3168]. 1959 [2492, 379]. 1960 [1380, 918, 563, 1386, 198, 1195, 3675]. 1960s [303, 3666, 662]. 1960s-1990s [662]. 1961 [731, 3609]. 1962 [2920]. 1963 [3185]. 1965 [2853, 1396, 1997, 2504, 2366]. 1968 [2908]. 1969 [320]. 1970 [2854, 219, 2858, 2780]. 1972 [280, 1051]. 1974 [2378, 1954]. 1975 [678, 1966, 2444, 962, 3426, 3522]. 1977 [3422]. 1979 [2683]. 1980 [1135, 2932, 1758]. 1981 [1148, 1680]. 1987 [70, 1146]. 1990 [554, 2308, 927, 3077, 1884, 1757, 3519]. 1990s [2478, 2218, 662]. 1992 [315, 2475, 325]. 1993 [2568]. 1995 [1697, 740, 1412, 546, 445, 2506]. 1997 [578, 2022, 1766]. 1998 [20, 3100, 1542, 2452]. 1999 [1178, 837, 2098, 2770, 790, 1360].

2 [651]. 20 [3761, 1279]. 20-kh-nachale [1267]. 2000 [579, 435, 239, 1560, 2129, 1188, 1654, 240, 958, 1072, 754, 2208, 1951, 2755, 1425, 2176, 3081, 3031, 1323, 1861, 3427]. 2001 [1071, 2007, 1342, 1561]. 2002 [2472, 1799, 1340, 1454, 1655, 1455, 2008, 1341, 1081, 2367]. 2003 [1896, 2240, 1798, 1559, 1653, 1800, 1562, 1895, 2009, 1652]. 2004 [2130, 2393, 2324, 2464, 2466, 2395, 2467, 3093]. 2005 [2588, 2961, 2465, 2396, 2397, 2855, 3424, 3191]. 2006 [2873, 3124, 2960, 3125, 2806, 2915]. 2007 [3458, 3026, 3371, 3253]. 2008 [3702, 3627, 3457, 3543]. 2009 [3703]. 2nd [680].

30-kh [1267]. 3032 [370]. 360/970 [264].

4300a [755].

5 [3605].

62 [3778]. 65 [1746].

750 [3107].

8 [1178]. 82 [1830].

90 [2746]. 90e [1310].

Âge [3159, 1691, 614, 2887, 790, 3739, 3355, 2823, 472]. Être [982]. êtres [3635]. îles [2894].

A-G [1872]. A. [3046, 2330, 1655, 1364, 1824]. A.A [1527]. A.D [1081, 3379]. A.D. [2334]. Aaboe [3026]. Aant [2992]. Aaron [2400, 240]. Aaserud [3465]. Abade [3327]. Ábbasid [1365]. Abbate [523]. Abbeys [872]. Abbott [2220, 937]. ABC [1629]. Abe [799]. Abel [855, 2198]. Abele [1956]. Abgrund [2837]. Abhandlungen [2997, 3171]. Abir [680]. Abir-Am [680]. Aboard [1010]. Aboriginal [1278]. Abortion [1423]. Abraham [1858, 79, 1493, 534, 2123]. Abrams [2009]. abri [3267]. Abriendo [2133]. Abroad [2124, 1245, 2201]. Absence [1550, 2179]. Absenkung [309]. Absolute [1207]. Abstraction [698, 230]. Abt [596]. Abu [2082, 264]. Abundance [698]. Abuse [1292]. Academia [3083, 1202]. Academic [942, 3099, 643, 559, 1250, 1202, 3375, 3653]. académie [1372, 2164]. Academy [3100, 2472, 1362, 4, 1057, 421]. Accelerating [3300]. Accelerator [734]. Access [954]. accipitrum [472]. Account [1747, 2391, 857, 202]. Accounts [2266, 2787]. Accuracy [2310]. Achievements [224, 1463]. Achilladelis [553]. Achim [2117]. Achinstein [1429]. achttiende [499]. Acid [1180, 1038, 2767]. Acker [2168]. Ackland [810]. Aclimatación [3056]. Aconcio [3102]. Acot [506]. Acoustic [794]. Acoustics [1615, 340]. acoustique [976]. acque [2338]. Acquiring [1250]. Acres [220]. across [1357, 589, 967, 1609, 661]. Acrostic [3730]. Actes [361].

Action [2767, 3154, 2384]. actividad [1242]. Activism [222, 2399]. Activist [2293]. activité [3361]. Activity [2238]. Acutorum [352]. Ad [1936, 1634]. Adam [202, 119, 3709, 1826, 3593, 589]. Adams [986, 3758, 215, 1007, 236]. Adaptation [2581]. adaptées [2091]. Addiction [2168]. Adelaide [1797]. Adele [87]. Administration [2172, 316, 3089]. Adolescent [322]. Adolf [170, 2117]. Adolph [2560]. Adolphe [2073]. Adoption [752]. Adriaan [385]. Adrian [1607, 109, 2937]. Adrienne [2333]. Adudaddaula [264]. Advanced [3559, 3345, 1306]. Advancement [595, 3117, 1290, 259, 2745]. Advancements [3552]. Advantage [815]. Adventurer [1338]. Adventures [3269, 558, 557, 2779, 2065, 2856, 2384, 1410, 2102, 1074, 3433, 1618, 1621, 2201]. Adversity [3050]. Advice [1746, 3185, 1053]. Advisor [2169, 3178]. Advocacy [49, 926]. Advocate [482]. Aebersold [407]. Aerial [2030]. Aerodynamics [690]. Aeronautical [1057]. Aesthetic [2535, 166]. Aesthetics [2283, 2160]. Aether [478]. Aetna [3725]. Afar [1621]. Affair [1838, 1878, 306]. affaire [2738]. Affairs [1069]. Affection [1622]. affectis [141]. Affects [3523, 3700]. Affiliation [1513]. Affinity [2349]. Afield [454]. Africa [505, 3081, 1713, 974, 1024, 776, 218, 1422, 3752, 3334]. African [2161, 1404, 1154, 858, 1104, 342, 1953]. African-American [1953]. Africanizing [1024]. afrikanischer [909]. After [1747, 3204, 2952, 2940, 533, 3163, 1821, 1421, 2652, 3784, 1043]. Afterlife [1871, 3514]. Aftermath [3453]. Afterword [3701]. Again [2991]. Against [1158, 1075, 2987, 745, 7, 1302, 1282, 3469, 3597]. Agar [2438]. Agassiz [1725]. Agatha [1058]. Age [2697, 1326, 1779, 2638, 1081, 853, 3329, 966, 911, 184, 1660, 2449, 2085, 970, 2736, 2559, 1707, 531, 2332, 3760, 1618, 1008, 3252, 3596, 2821, 2672, 145, 3051, 2049, 1709, 2030, 1834, 1887, 3469, 2486, 445, 3675, 2714, 3076, 2743, 466, 77, 1712, 2766, 3673, 403, 3249, 3709, 43, 2005, 3603, 700, 1824, 2989, 3090, 1137, 2199, 726, 2247, 261, 63, 1395, 2018, 961, 3713, 3683, 3030, 1535, 3244, 3770]. Age-Old [2697]. Agency [2700, 3359, 1146, 1724]. Agents [3020]. Ages [2021, 1821, 2243, 1812, 1490, 3488, 338, 2881, 787, 151, 615, 3147, 1440, 1491, 1911, 3562]. aggregationis [3483]. Aging [2945]. Agnesi [3492, 751]. Agnoletti [1198]. Agnotology [3614]. Agostoni [2832]. Agrarian [3349]. Agricultural [2907, 766, 739]. Agriculture [740, 2492, 546, 2634, 2165, 913, 660, 2181]. Agronomy [2367]. Agusti [134, 2938, 1952]. Ägypten [1089]. Ahead [1418]. Ahmed [1875]. Ahmes [584]. AI [3429, 3735, 1465]. Aid [3769]. AIDS [2502, 549, 222, 218, 2105]. aiguës [1686]. Aiken [203]. Aileen [2550]. Aimé [1648]. Aiming [1137]. Ainley [3702]. Air [390, 2134, 3337, 294, 914, 3774, 1057, 971]. Aira [120]. Aircraft [1544, 2084]. Airship [542]. akademicheskoi [2191]. Akademie [2997]. Akademii [1354]. akademischen [67]. Akademisches [3477]. Akera.S [3344]. Akteure [2368]. Al [2245, 1497, 2022, 146, 1425, 3486, 18, 1364, 264, 1399, 3486, 1364, 1497]. al-Amid [264]. al-Andalus [1497]. al-Aqsà [1497]. al-Asbahi [3486]. Al-Birjandi [2245]. al-Farisi [3486]. al-Haytham [1364]. al-Magrib [1497]. al-Manāzir [1364]. Alain [3232, 503, 1014]. Alan [359, 304, 2878, 1935, 780, 623, 1121, 1974, 1859, 2771, 2045, 2073, 996, 2934]. Alaska [2212, 1424]. Albanese [29]. Albert [161, 712, 2683, 592, 396, 3410, 2751, 34, 68, 3028, 3483, 91, 343, 2886, 2751, 958, 68]. Alberti [873, 3467]. Alberto [156, 58, 2299, 777]. Albertus [2407, 3019]. Albillos [3056]. Albrecht [1827, 1103]. Albree [421]. Alchandreana [3388]. Alchemical [114, 2043]. Alchemie [3227, 617, 617]. Alchemiebibliothek [95]. Alchemist [3042]. Alchemy [2185, 1315, 1358, 611, 2983, 2490, 3055, 3304, 3567, 2892, 603, 1591, 2599, 3550, 1731, 3484, 2043, 3392, 3493]. Alchimia [2405]. alchimie [1310, 3311]. alchimique [361]. Alchymy [874]. Alcock [1037]. Alcohol [1393]. Alcoholism [2693]. Alcón [1820]. Alder [2666]. Aldrich [804]. Aldrovandi [2726]. Alec [2369]. Alejandro [714]. Aleksandra [3761]. Aleph [1345]. Alessandro [2418, 3238, 3034]. Aletta [2676]. Alex [3668, 2568, 2287, 2286, 3097]. Alexander [446, 290, 95, 3331, 2074, 765, 309, 1589, 553, 1147, 755, 590, 2695, 2204, 1454, 2847, 1320]. Alexandra [2849]. Alexandria [862, 3308, 3718]. Alexei [2779]. Alexis [1718]. Alfonsine [2404]. Alfred [1025, 1848, 2435, 3005, 555, 898, 2994, 53, 1118, 2280, 897, 1392, 2617, 2024, 925, 1603, 2633]. Alfredo [701]. Algebra [1496, 3583, 3257, 3306]. Algebraic [1258]. Algebralehrbücher [1920]. algébrique [1907]. Alhacen [1364, 1364]. alia [2386]. Alice [731, 117, 1755, 297, 1715]. Alien [193, 1867, 2259, 193, 3771]. alimentorum [1688]. Alisa [2178]. Alison [1837, 2442, 984]. Alistair [1811]. Alister [2311]. Aliveness [3579]. Alix [3049]. alla [994, 324, 3732]. Allan [1630, 1032, 2848]. Allegany [564]. Allemagne [1735]. Allen [160, 3638, 2447, 220, 176, 148, 656, 2930, 856, 1919, 1930, 2629]. Allergies [3220]. Allergy [2928]. Allianzen [2838]. Allin [505]. Allison [1135]. Alm [1038]. Almagest [462]. Almeder [2577]. Almost [3673, 3414, 2566]. Almroth [904]. along [736]. Alpert [1411]. Alphabet [3690]. Alquibla [1497]. Alsted [1093]. Alta [190]. altägyptischen [867]. alten [1089]. Alter [2613, 516]. Alternative [2111, 1197, 128, 1883, 1605]. Alternativen [171]. Alternatives [1186, 2936]. Altertum [2144]. Altitude [452]. Alton [1852]. Altruism [2878]. Always [3700]. Alwyn [1075]. Alzheimer [2899]. Am [680, 1470, 1279, 67, 2372, 3103]. AMA [1273]. Amadae [2083]. Amaldi [526]. Amalia [493]. Amanda [771]. Amateur [1909, 1865, 2713]. Amazing [2978, 3385]. Amazonia [2035]. Ambassador [2042]. Ambassadors [2152]. Ambiguity [3471]. Ambition [2859]. Ambitions [2556, 987, 2599, 2713, 2248]. Ambonese [881]. Ambrose [2132]. Amchitka [2212]. America [3132, 2105, 3111, 740, 731, 2195, 595, 2693, 1299, 659, 643, 181, 1761, 2827, 3085, 3079, 1627, 1143, 372, 2899, 1197, 1271, 1759, 1615, 1854, 3131, 916, 563, 3139, 1130, 1846, 2594, 3501, 320, 1680, 3354, 1721, 1479, 344, 3140, 2363, 1932, 2508, 124, 2079, 198, 2804, 2582, 3283, 957, 2234, 3693, 2714, 576, 2071, 2678, 2346, 2097, 2985, 2077, 1423, 2924, 557, 1401, 2849, 539, 2978, 456, 3091, 2914, 1127, 2188, 1883, 2348, 2631, 2371, 2778, 661, 566, 2190, 2660, 180, 2302, 206, 27, 2047, 1820, 2111, 2885, 3380, 3518, 3162, 3336, 1237, 3274, 1256, 3474]. America [3588, 3737, 1253, 2193, 113, 3199, 3603]. American [1632, 855, 2622, 637, 3776, 2168, 2614, 1945, 191, 596, 2364, 2434, 2885, 182, 632, 595, 2834, 511, 818, 2081, 3048, 1009, 345, 3476, 3341, 117, 977, 727, 3548, 2093, 396, 1844, 1946, 3339, 1144, 655, 2546, 2755, 696, 1166, 757, 1273, 3522, 3596, 1382, 3187, 1732, 1290, 2478, 1935, 660, 2858, 3093, 1155, 1675, 89, 858, 1118, 899, 251, 537, 2305, 278, 1835, 891, 88, 2036, 178, 1834, 2379, 3663, 185, 321, 1019, 1130, 2497, 1954, 990, 1431, 3168, 373, 3294, 1147, 3509, 2576, 1287, 2034, 87, 1413, 3332, 3750, 91, 1391, 817, 3587, 1208, 962]. American [2069, 2421, 291, 3417, 2181, 391, 2201, 1480, 1927, 3266, 2773, 633, 1032, 2529, 2186, 1750, 3270, 3216, 342, 3019, 2436, 3535, 7, 2462, 2554, 2740, 2303, 3456, 674, 3541, 1108, 3212, 1949, 1181, 948, 1056, 1600, 1461, 598, 1513, 3189, 571, 77, 81, 279, 1953]. Americana [1830]. americanas [768]. Americanist [1176, 90]. americanistas [3174]. Americans [3059, 851, 858, 1020, 992]. Americas [86, 868]. Amerika [525]. Amerindian [1404]. Ames [588]. Ames-Lewis [588]. Amherst [2472]. Ami [797]. Amid [264]. Amir [1589]. amis [1310]. Among [288, 186, 119]. Amoretti [2977]. Ampère [3071, 1596]. Amsterdam [3175, 940]. Amsterdamse [1936]. Amundson [2630]. Amy [1309, 247, 2434, 1951, 2060]. Anaconda [914]. Anales [2366]. Analog [1762]. Analyse [3715, 2739]. Analysis [1180, 3640, 889, 3759, 2785, 1850, 1989, 19]. Analytical [296]. Analytics [2575]. anatomia [58]. Anatomical [2334, 2040]. Anatomice [2040]. Anatomie [3747, 2905, 2485]. Anatomies [368]. Anatomista [1999]. anatomo [610]. anatomo-patologica [610]. Anatomy [3163, 2303, 1253, 1287, 1451, 594, 3585]. Anaximander [2402]. Ancestors [1312, 3593, 1621]. Ancestral [118]. Ancestry [2552]. ancienne [2822, 826]. anciens [3388]. Ancient [3040, 465, 1996, 2333, 2038, 1214, 3231, 609, 468, 980, 244, 1487, 3725, 1538, 1217, 938, 463, 2095, 2602, 3724, 1490, 3480, 2336, 2248, 3723, 438, 1893, 3301, 3297, 2207, 2533, 262, 977, 2150, 978, 465, 1443, 374, 461]. Andalus [1497]. Andean [2039]. Andere [126, 2182, 2724]. Anders [341]. Andersen [3467]. Anderson [2748, 2053, 1198, 690, 343]. Andre [2834, 483, 2100, 1825]. Andrea [88, 2266, 2401, 3039]. Andreas [436, 3685, 2607, 337, 1856, 2258, 1826, 2910, 3567]. Andree [381]. Andrei [2772, 1536]. Andres [141]. Andrew [1262, 608, 482, 2884, 3501, 2964, 536, 3190, 296, 532, 2220, 2491, 2571, 1117, 937, 2409, 1879, 1884, 1839]. Andrews [2364, 2054, 3653, 299]. Andy [561]. Anecdotal [1542]. Anemia [2446]. Anfang [67]. Anfänge [1856, 587]. Anfängen [3227, 2818]. Ángel [2742, 639]. Angela [1396]. Angèle [743]. Angeles [2845]. Angelis [2347]. Angelo [2996]. Angels [3730, 611, 2658, 954]. Angles [20]. Angleterre [494]. Anglo [2081, 393, 851, 891, 1927, 2292]. Anglo-American [2081, 891, 1927]. Anglo-Egyptian [393]. Anglo-European [2292]. Angst [663]. Angus [2970]. anhand [3406]. Anhang [719]. Anima [607]. Animal [1063, 1881, 412, 3676, 2552, 2329, 2029, 1945, 1909, 49, 2574].

Animal-Human [2552]. animale [2352]. animales [3056, 1820, 2386]. animali [3161]. animalium [3483]. Animals [1063, 1665, 2885, 2086, 1105, 1223, 2188, 2870, 2305, 412, 1511, 3156, 2574, 2029, 2979]. animaux [2979]. Anita [713, 3659, 177, 2617, 2029]. Anjing [3275]. Anker [2423]. Ann [805, 68, 564, 514, 1430, 2157, 920]. Anna [3496, 845, 501, 686, 693, 2798]. Annals [1756, 2420]. Anne [1175, 626, 1948, 3252, 3154, 1087, 788, 710, 3031, 1150, 1125, 1829]. Annelies [95]. Annelore [500]. Annet [554]. Annette [556]. Annie [1147]. anniversaire [1310]. Anniversary [3196]. Anno [866]. Annotated [1600, 1683, 1166, 1249, 1167, 1608, 1693]. Annual [579]. anonimo [35]. anonyme [2337]. Anstaltspsychiatrie [3101]. Ant [1068, 2631]. Antarctic [2680, 1251, 211, 2680]. Antarctica [3076]. Antecedents [280]. Anthony [305, 3739, 355, 612, 624, 873, 2146, 2253, 992, 2018]. Anthrax [3281, 1961]. Anthropological [177, 3771]. Anthropologie [2818, 2763, 3179]. Anthropologist [202, 1426, 2769, 2260]. Anthropologists [119]. anthropologues [2549]. Anthropology [1779, 1024, 2771, 2242, 2358, 1211, 3439, 1117, 2834, 119, 2546, 2295, 1959, 1312, 3756, 1732, 1247, 1953]. Anthropomorphism [412, 2574]. Anti [1304]. Anti-Cancer [1304]. Antibiotics [2914]. Anticommunism [77]. Antidepressant [318]. Antiguas [2148]. Antihumanism [1117]. antiken [1489]. antimonistische [2205]. Antimony [857, 857]. Antinuclear [532]. Antipodes [3729]. Antique [1996]. Antiquité [790]. Antiquities [1851]. Antiquity [3107, 3136, 2881, 2728, 3485, 1167, 2030, 3488, 3712, 2018, 1911, 2033, 862, 3480, 3481]. Antisemitismus [1109]. Antoine [1509, 502, 1376, 2620]. Antoine-Augustin [502]. Antoine-Laurent [1376]. Antonella [613]. Antonia [3052]. Antonio [44, 768, 3054, 3735, 498, 3728, 2829, 2988, 879, 3053]. Antony [12]. Anwendung [171]. Anxiety [77, 403]. anys [2430]. Anyway [3207]. Apache [1248]. Aparicio [141]. Apart [2565]. Ape [2552, 2160]. Apes [2552]. Aphasia [813]. aphasies [2905]. Aphorisms [3312]. apocalisse [493]. Apocalypse [3341]. Apocalyptic [3055]. Apollo [3611, 2244, 2378, 1294, 2160, 1767]. Apollonius [1684, 3647]. Apotheker [274, 1574]. Apotheker/Arzt [274]. Apotheosis [3307]. Apparatus [1509]. Apparent [971]. Appeal [3682]. Appel [962]. Appl [195]. Application [1634]. Applications [734, 1318]. Applied [3641]. Applied-Science [3641]. apport [1352]. Appreciated [2124]. Appreciation [2919, 3315]. Apprenticeship [3118]. Approach [2362, 3094, 1043, 195, 2749, 2057]. Approaches [1318, 1059, 2521, 2965, 296, 1198, 1663]. Appropriate [2957]. Appropriating [2562]. Appropriation [536]. April [3543]. apropriación [884]. aproximació [1462]. Aqsà [1497]. Aqua [1374]. Aquarium [800]. Arabatzis [2749]. arabe [2879, 3388]. Arabern [264]. Arabic [3646, 3312, 1365, 705, 3365, 2245]. Arabic-English [3312]. Arabischer [34, 607]. Arago [1557]. Aragón [3056]. Aratos [32]. Arbeit [3134]. Arbeiten [3757]. Arbitrary [1670]. Arch [3640]. Archaeoastronomy [979]. Archaeological [3760, 456, 2082]. Archaeologist [1965]. Archaeologists [3038, 308]. Archaeology [977, 2669, 3760, 542, 765, 3675, 117, 90, 774, 2623, 1176, 1046]. Archaic [3644]. archeologie [176]. Archeology [1389]. Archetype [2301]. Archimedes [1633, 3485, 2821]. Architect [1408]. Architectonica [3727]. Architects [205]. Architectural [1615]. Architecture [158, 2825, 3155, 509, 1733, 581, 2332, 2336, 1160, 1634]. Architectures [1029, 1467]. Architektur [1470]. Archiv [3727, 3260]. Archival [2992]. Archive [2525]. Archives [1979, 2472, 2277, 96, 1108, 939, 2088]. Archivgespräche [1643]. arco [498]. Arcos [2662]. Arctic [1325, 3339, 3076, 245, 389, 342, 1424, 1209]. Ardenne [361]. Area [1487]. Arend [704]. Areteo [610]. Argand [797]. argent [2706]. Argentina [3182, 2505]. Argentine [3601]. Argonaut [730]. Argumenten [3715]. aria [2842]. Ariadnefaden [360]. Arianna [3727]. Ariel [136]. Arien [412]. Ariew [154, 157]. Aristides [679]. Aristote [3649]. Aristotele [607]. Aristotelian [1704, 145, 865]. Aristotelianism [1704]. Aristotelis [2729]. Aristotle [2575, 3379, 467, 2479, 2147, 3378, 864, 2880, 3039, 467]. Arithmetic [667]. Arithmetik [1920]. Arithmetik- [1920]. arithmétique [2408]. Arjo [568]. Ark [269]. Arleen [3006]. Armageddon [1629]. Armando [2099]. armée [1710]. Armstrong [2984, 2544, 47, 2774]. Army [313, 130]. Arnal [200]. Arnaud [2532]. Arney [421]. Arnold [2898, 812, 3044, 1714, 250, 1595, 752, 812]. Arnoud [3646]. Arnzen [607]. Arquitectos [692]. Arrangement [2373]. Arrizabalaga [1697]. Arrows [2333, 3713]. Ars [1697]. Art [1438, 247, 3467, 464, 2881, 2542, 1592, 993, 2026, 1509, 3323, 2241, 1598, 3590, 2821, 2341, 1591, 2296, 3232, 2020, 3000, 3022, 78, 1667, 1582, 2659, 1915, 1085, 1811, 2523, 434, 1687, 2532, 111, 2315, 1730]. arte [147, 2338]. Artefakte [2368]. Arthur [1870, 765, 2762, 3773, 2563, 75, 1159, 2659, 2782, 2989, 630]. Artibus [3248]. Articella [1697]. Article [1974]. articles [3348, 1995]. Artifacts [1965]. Artificial [3546, 3490, 1426, 3700]. Artigas [2050]. artigiani [3735]. artillerie [364]. Artis [3175]. Artisan [2542]. Artisans [159]. Artist [1408]. artistas [1602]. Artists [594]. Arts [1632, 446, 2195, 1495, 1817, 3313, 2033]. Arun [3045]. Arwen [293]. Aryans [120]. Arznei [163]. Arznei- [163]. Arzt [274, 164]. Ärztinnen [1254]. Asbahi [3486]. Asbestos [1894]. Asger [3026]. Ash [519, 2604, 405]. Ashley [3154]. Asia [1787, 3678, 3351, 3223, 653, 1984, 1194, 1193]. Asian [3351, 979]. Asif [2378]. Asimov [1409]. Asklepios [3679]. aspectibus [1364]. Aspects [1059, 3727, 361]. aspetti [610]. Assault [801]. Assembling [1046]. Assessment [224]. Assimilation [1449]. Assistants [703]. Assmann [3297]. Associated [2993]. Associates [2782]. Association [221, 595, 3117, 1290, 1722, 3351, 3268, 1355]. Assumptions [3411]. Assuring [545]. Asteroids [1347]. Asthetik [256]. Ästhetische [2831]. Astore [2985]. Astral [1485]. Astrida [2167]. Astro [3300].

Astro-Physicists [3300]. Astrobiology [2925]. Astrofuturism [2032]. Astrolabe [3727, 2154, 705]. Astrolabes [1998, 1498, 3730]. astrolabii [1222]. Astrologer [3042, 2253]. astrologi [1685]. Astrología [2148]. Astrological [995]. Astrologie [3482]. astrologiques [1691, 3388]. astrologues [354]. Astrology [1315, 1782, 2819, 2481, 1474]. Astrometry [20]. Astronautics [3100, 1645]. astronome [3572]. Astronomers [966, 727, 1363]. Astronomes [968]. astronomi [583]. astronomia [2022]. astronómica [625]. Astronomical [3235, 191, 596, 2883, 894, 2158, 2859, 979, 385, 3644, 755]. Astronomie [3500, 3486, 2738]. astronomii [2262]. Astronomy [464, 2022, 244, 3408, 889, 3557, 123, 583, 1033, 1383, 236, 2273, 461, 2245, 3504, 245, 2335, 3135, 279, 835, 3427, 3512, 3386, 1349, 2149, 485, 2141, 2531, 708, 2140, 2283, 705, 262, 1925, 1232, 3143, 3037, 3724, 967, 1493, 1503]. Astrophysical [596]. Astrophysics [3009, 123]. Asúa [2474]. AT&T [431]. Atatürk’s [3181]. Athanasius [1823, 3406]. Atheism [2358]. Athena [1682]. Ätherwindexperiments [2183]. Atkins [1805, 528]. Atkinson [678]. Atlantic [2740, 3733, 1239, 2540]. Atlantis [590]. Atlantropa [309]. Atlantropa-Projekt [309]. Atlas [2263]. Atmosphere [1064]. Atmospheric [1293, 3041]. Atom [2658, 2301]. atome [2915]. Atomi [877]. Atomic [3776, 924, 1139, 1618, 534, 891, 2365, 528, 2519, 961, 647]. atomique [3355]. Atomism [2888, 879]. Atomisme [1936]. Atomismus [171]. Atomista [1701]. Atomistics [631]. atomizzazione [25]. Atoms [2269, 2892, 1840, 277]. Atomzeitalters [1956]. atribuidos [1500]. Atsushi [3344]. Attained [3231]. Attending [794]. Attention [521]. Attis [365]. Attitudes [438]. Attractions [1945]. attribué [3483]. Atwater [1166]. Aube [3159, 2896]. Aubin [3364]. Aubin-Louis [3364]. Aucante [2973, 490]. Audience [3117]. Audiences [341, 3744, 3413]. Audion [1864]. Audoin [1139]. Audrey [3124]. Aufbruch [1320]. Aufklärung [2291, 2831]. Aufklärung. [2897]. Aufnahme [2406]. Auftrag [2143]. Augsburg [619]. August [3124, 42, 578]. Augustanus [619]. Auguste [518, 2452]. Augustin [502, 2547, 2477]. Auk [1115]. Aull [747]. aura [1593]. aura-t-elle [1593]. Aurelian [1679]. Aurora [3633]. Auschwitz [2116]. Ausgang [2831]. ausgehend [2183]. Aussage [1574]. Ausstellungen [2368]. Austen [3591]. Austin [2169]. Australia [3100, 2058, 2768, 3351, 1588, 2053, 132]. Australian [286, 342]. Auswahlbibliographie [2205]. Authoritative [1441]. Authority [3619, 3493, 3677, 2135, 321]. Authors [3162, 3520]. Authorship [1000, 2033]. Autobiographical [921]. Autobiographies [2062]. Autobiography [808, 1960, 1565, 711, 1876]. autodétermination [2737]. Autonomy [1434, 30, 2451]. Autoproduction [2737]. Autour [2738, 761]. autre [3639]. autres [2979]. Aux [1610, 3635, 2738, 3494, 2088]. avalanche [3495]. avant [3739, 3348]. avant-propos [3348]. Avenging [639]. Aveni [2146, 2018]. aventure [183, 335]. Averaged [3187]. Averroès [2879]. Aviation [2781, 3337, 3168]. Avicenna [2480]. AVISTA [3232]. Awards [3275]. Away [550]. Aymerich [176]. Ayres [2833]. Ayval [2488].

Azarkh [531].

B [1175, 2943, 2938, 3158, 161, 1243, 447, 884, 2077, 452, 3736, 2741, 153, 1020, 3642, 50, 2779, 123, 2951, 720, 1400, 369, 699, 104, 357, 2190, 1303, 506, 275, 2442, 353, 2979, 780, 3523, 1811, 215, 2369, 2096, 597, 711, 821, 843, 843]. B. [104]. B.C [1081, 3379, 606]. Babbage [890]. Babies [3469]. Babini [2986, 3141]. Baby [1744]. Babylon [910, 3038, 2017, 2140]. Babylonian [3385]. babylonischen [10]. Bacchic [1393]. Bachelard [2222]. Bachman [632]. Bachmann [617]. Back [3, 100, 233, 328, 424, 429, 573, 671, 749, 829, 831, 1334, 1793, 2231, 2516, 3044, 774, 2805, 2652, 1022, 818, 62]. Backbone [1673]. Background [154, 416]. Backstage [407]. Bacon [621, 2890, 3490, 2745, 36, 1097, 3366, 793, 622, 623, 3660]. Bad [2816, 1479]. Bada [930]. Badischen [95]. Baffin [288]. Baghdad [1365]. Bahn [2308]. bai [3200]. baijia [307]. Bailey [446, 3118, 571]. Baird [2448, 53, 1145]. Bake [1130]. Baker [2447, 2812, 2003]. Bala [3045]. Balance [1032, 868]. Bald [628]. Baldi [38]. Bali [1526]. Ball [2020, 2019, 2018]. Ballenger [2899]. Ballester [1654]. Ballistic [742]. Balloons [797]. Ballstadt [3731]. Baltas [679]. Baltimore [1418]. Ban [404, 3185]. Banks [2545, 3241, 3151, 43, 3147, 2482, 771]. Barbara [1162, 3679, 1270, 563, 285, 2911, 2824, 882, 995, 1123, 2935]. Barbarians [3752]. Barber [2102]. Barberá [1820]. Barbo [798]. Barcelona [498]. Barcelonesa [2430]. Bard [2200]. Bare [1795]. Barger [821]. Baring [103]. Barkan [728]. Barnaby [2409]. Barnes [2987, 2809, 605, 3688]. Barnhart [2546]. Baroque [1592, 1823]. Barr [2207, 288]. Barras [627]. Barrel [643]. Barriers [3099]. Barron [736]. Barrow [1424]. Barry [193, 3688, 314, 1935, 3270]. Bart [2811, 1871, 2765]. Barth [2834]. Barthélemy [2408]. Bartholomew [193]. Barton [3742]. Bartram [1263, 990, 3323]. Basalt [712]. base [498]. Based [3641, 463, 2296, 2704]. Basic [833]. Basil [553]. Basilica [753]. Basis [2420, 1893]. Bassett [2559]. Bastien [86]. Bates [3046]. Bateson [1526, 433]. Bath [1829]. Batten [1651]. Battimelli [526]. Battista [873, 2726]. Battle [221, 1200, 2073, 281]. Bauer [823]. Baumann [1004]. Baur [1964]. Baur-Fischer-Lenz [1964]. Bauten [1052]. Bayle [3578]. Bayung [1526]. Be [935, 2845]. Beagle [2065]. Beam [2777]. Bearing [3235]. Beasts [1534, 767]. Beatrice [1004]. Beatrix [2789, 2999, 1127]. Beauchamp [1863]. Beaufort [885]. Beautiful [166, 2074, 1623, 3015]. Beauty [951, 441, 3172]. Became [2627, 3079, 323, 1569, 3537]. Beche [420]. Bechtel [2852]. Beck [247, 117]. Beckham [2055]. Beckman [812]. Beckons [3652]. Beckwith [2293]. Becomes [1753]. Becoming [1256, 2798, 2309, 3690]. Bedeviled [460]. Bedini [1929, 477, 159]. Beeckman [1701]. Beeley [2731, 162]. Beeman [913]. Been [3286, 1409]. Before [2952, 1767, 3709, 2940, 2015, 2914, 1812, 1677, 3729, 273, 1286, 3677, 682, 3164]. Began [3408]. Begets [900]. Beginn [1279, 1385]. Beginning [1797, 816]. Beginnings [858, 1124, 2606, 1703, 3555, 2251, 2071]. Begriff [1219]. Begun [1030, 1030]. Behavior [2615, 604]. Behavioral [223, 2647, 2554]. Behind [2472, 1212, 3272, 2696, 3237, 2992, 2018, 3150]. Behm [210]. Behnegar [1849]. Behrens [194]. Beidleman [2900]. Being [2882, 3623, 4, 698, 3690]. Beisner [2789]. Beispiel [1470, 2616]. Beisswanger [1912]. Beitrag [1402, 2846]. Beitrage [52, 199, 150, 3227, 1178, 1319, 476, 1324, 2143]. Belehrung [500]. België [3766]. Belief [1281, 700, 3150]. Beliefs [1307, 127]. Believe [1045]. Believing [2524, 2714]. Bell [393, 152, 3775, 1282]. Bella [2173]. Bellamy [907, 3077]. Beller [530]. Bellos [685]. Bellows [561]. Belmar [3054]. Below [1424, 3559]. Beltran [44, 3053]. Bemuhungen [274]. Ben [2974, 2505, 1136, 2219]. Ben-Chaim [2974]. Bender [2614]. Bending [3610]. Bends [294]. Benedict [882, 1134, 712, 2886, 1397]. Benedikt [2182]. Benefit [1463, 450]. Benefits [1035]. Benison [821]. Benjamin [3166, 3048, 3057, 3500, 695, 1841, 807, 2655, 2963]. Bennett [2159, 269, 2069]. Benoît [2906]. Benoit [1006]. Bensaude [825, 826, 341]. Bensaude-Vincent [825, 826, 341]. Bensman [816]. Benson [806]. Bent [2932]. Benton [130]. Benvenuto [2132]. Benz [1371]. Benzer [2856]. Beobachtung [10]. Berechnung [10]. Beretta [2810, 779, 1376, 3238, 25]. Berg [561]. Bergbau [3227]. Berge [514]. Berger [807]. Bergerac [2258]. Beriberi [597]. Berichte [1178, 1324]. Berkel [592]. Berkeley [559]. Berkner [1032]. Berkowitz [2732]. Berlin [3103, 2372, 3410, 2751, 68, 3414, 2910, 3254]. Berliner [3515, 3672]. Berlinghoff [2021]. Berlins [556]. Bernabé [3140]. Bernadette [825, 826, 341, 2844, 2360]. Bernard [1405, 161, 92, 565, 818, 3449, 295, 762, 1139, 203, 2404, 228, 501, 3011, 501]. Bernardino [38, 2260]. Bernd [2838]. Bernhard [170, 3063, 799]. Bernoulli [483, 42, 1923]. Bernoullis [2814]. Bernstein [1268]. Bernus [95]. Berridge [820]. Bert [1180, 2614, 906]. Berthold [153]. Bertomeu [3054, 2938]. Bertomeu-Sánchez [2938]. Bertram [719]. Bertrand [1629]. Bertsch [189]. Berube [2696]. Beside [3690]. Bess [2208]. Bessarion [3730]. Besserat [242]. Best [395]. Bestimmung [3486]. Beth [399]. Bethe [649]. Bethell [69]. Bethlehem [350]. Better [3069, 1744, 2849, 2567, 397, 925]. Bettina [909]. Bettini [493]. Betty [1141]. Bettyann [2566]. Between [2790, 6, 3512, 418, 1738, 2612, 3654, 1235, 1881, 2727, 333, 935, 770, 2904, 2907, 1231, 2996, 2685, 1597, 3381, 944, 545, 568, 1677]. Bevalac [1337]. Beverley [541, 554]. bewahre [48]. Bewegung [1734]. Bewick [3165]. Beyer [340]. Beyond [3732, 2140, 444, 301, 2771, 535, 2865, 193, 3136, 1200, 3429, 1194, 225, 1570, 3720]. bez [1291]. Bhopal [2872]. Biagioli [2826, 107]. Bias [2445, 2598]. Bibiliothèque [3321]. Biblical [1095, 104, 269]. Bibliografia [1553]. Bibliographia [586]. Bibliographic [1165, 1966]. Bibliographie [2202, 790]. Bibliography [2, 427, 2517, 2796, 3116, 3447, 3392, 2055, 1631, 830, 1333, 1792, 2230, 1167, 1600]. Biblioteca [3141, 2988, 2977, 1696]. bibliothek [48]. Bibliothèque [3157, 3355, 3152]. Bicycle [3473]. Biederman [78, 78]. Bier [1402]. Biermann [1799]. Big [2502, 216, 2839, 3262, 2658, 1149, 1337]. Bild [1052]. Bildung [2846, 337]. Bilharz [564]. Bill [517, 1938, 1527]. Bilson [811]. Bilstein [2084]. Binary [2039]. Bindman [2160]. bing [3517]. Bingen [2972]. Biocapital [3195]. biochemische [2838]. Biochemist [3414]. Biodiversity [2998]. Bioethics [1299]. Biogeochemistry [915]. biografia [3327]. biogràfica [1462]. Biographia [2607]. Biographical [2472, 1848, 938, 2749, 571, 1431, 231]. Biographie [1826, 2753]. Biographien [2607]. Biographies [3245, 782]. Biography [1908, 45, 757, 3153, 3505, 2294, 2587, 1051, 2585, 1715, 2483, 3772, 287, 2210, 507, 3226, 1603, 2765, 3215, 1461, 1536]. biológica [449]. Biological [2333, 2661, 1888, 2051, 1746, 1627, 2087, 652, 962]. biológico [1602]. Biologie [2276, 3013, 3355]. biologii [1314, 1267]. Biologische [2838]. biologischen [2818]. biologischer [2661]. Biologist [1715]. Biologists [1118]. Biology [1314, 3286, 931, 3303, 3186, 413, 2479, 676, 1421, 770, 2092, 1210, 172, 2853, 2451, 962, 1623, 2573, 1269, 3442, 198, 2447, 2375, 897, 2852, 80, 558, 2578, 197, 3005, 2067, 2813, 3097, 2246, 9, 3636, 770, 1572]. Biomedical [1434, 2557, 1417, 2305, 3680, 332, 2299]. biomédicas [2444]. Biomedicine [1737, 3440, 2570, 1197, 3680]. Biometry [2901]. Biopolis [1604]. Bioprospecting [2540]. Biosocial [3450]. Biosphere [915, 1741]. Biot [1557]. Biotech [3193, 1149]. Biotechnologie [3193]. Biotechnology [676, 2656, 407]. Biow [2152]. Biowarrior [2087]. Biowissenschaften [2116]. Bipolar [3608]. Bird [1909]. Birjandi [2245]. Birkhead [1909]. Birth [1065, 1093, 283, 966, 254, 1950, 1423, 2974, 257, 3052, 655, 1947, 1534, 2725, 1094, 3155, 1247, 1104, 385, 3045, 1860, 771]. Birthday [1153]. Birthing [2747, 2668]. Birthplace [917]. Births [3046]. bis [1815, 3227, 2308, 2258, 3035, 2607, 347, 2553, 3477, 1017, 2056, 2144, 3260, 587, 1386, 2818, 2831]. Bitbol [72]. Bizarre [1852]. Bjelić [2221]. Blaauw [385]. Black [51, 2098, 2532, 3732, 3093, 916, 1864, 3425, 14, 2161, 733]. Black-Box [1864]. Blackett [2529]. Blackmore [276]. Blackwell [1179, 266]. Blaise [2673]. Blanchard [2733]. Blanckaert [2675]. Blanes [2995]. Blanes-Nadal [2995]. Blass [3078]. Blay [357]. Bleker [67]. Blessing [694]. Blessings [1874]. Blick [121]. Blind [7, 3168]. Bliss [2750, 515]. Blood [2933, 3342, 1747, 1069, 549, 1481]. Bloom [1367]. Blu [2631]. Blue [1598]. Blueprints [1290]. Blues [2446]. Blum [1622]. Bluma [186]. Blyman [3124]. Bo [3261]. Boas [288, 1918]. Bob [501]. Bodies [3697, 1781, 2108, 295, 2869, 1241, 1280, 529, 3160, 323, 1253, 767, 1667, 2040, 1464, 1728, 487]. Bodily [1785]. Bodleian [1054]. Body [1785, 436, 468, 3055, 2542, 1486, 2221, 512, 1150, 166, 1949, 901, 2049, 3532, 1667, 3560, 2891, 2785, 1927, 3620, 1169, 436, 2596]. Bodysnatchers [1770]. Boeckmann [899]. Boerhaave [2487, 1233]. Boethium [473]. Boethius [3171, 3145]. Boewe [908, 1851]. Böhler [3695]. Bohm [78, 1619, 78]. Bohm-Biederman [78]. Bohr [1967, 301, 3008, 3508, 3465, 2904, 2458]. Bois [3254, 3178]. Bois-Reymond [3254, 3178]. Boke [481]. Bolea [1399]. Bolin [1180]. Bologna [3319, 2022, 2726, 693]. Bologne [2257]. Bölsche [1847]. Bolshevik [2054]. Bolsheviks [3282]. Bolt [686, 1841]. Bolton [1844]. Boltzmann [277, 276]. Bolyai [2618]. Bolzano [501]. Bomb [650, 3776, 2212, 534, 2658, 649, 3119, 2365, 1158, 3264]. Bombe [1260]. Bombers [313]. Bombing [390]. Bombs [2804, 2333, 3274, 3262]. Bon [1006]. Bonaparte [513, 335]. Bondeson [167, 1828]. Bone [1621, 3620, 1262]. Bones [436, 1451, 3514]. Bonn [1532]. Bonnefoi [44]. Bonner [1858, 1800]. Bonnie [110]. Bonoli [382]. Book [1029, 852, 994, 2173, 736, 2989, 3601, 650, 973, 3647, 1025, 183, 1576, 3194, 908, 2353, 2678, 3344, 3631, 2697, 691, 873, 1160, 3002, 3236, 2882, 3132, 2823, 1255, 3040, 3034, 761, 3572, 2405, 2852, 465, 489, 1254, 867, 1870, 2498, 2742, 2987, 602, 1263, 2898, 3269, 3755, 1484, 2185, 2933, 974, 2208, 2493, 2789, 942, 1174, 1178, 1179, 1180, 1181, 1308, 1309, 1310, 1311, 1312, 1313, 1314, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1323, 1324, 1325, 1326, 1327, 1328, 943, 944, 945, 946, 947, 948, 949, 950, 951, 1058]. Book [1059, 1060, 1061, 1062, 1063, 1064, 1065, 1169, 1170, 1171, 1172, 1173, 1175, 1176, 1177, 1432, 1640, 1644, 1645, 1778, 1779, 1780, 1781, 1782, 1783, 1784, 1785, 1786, 1787, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1441, 1442, 1443, 1444, 1445, 1446, 1632, 1633, 1634, 1635, 1636, 1637, 1638, 1639, 1641, 1642, 1643, 1889, 1998, 2105, 2106, 2107, 2108, 2109, 2110, 2111, 2112, 2113, 2114, 2115, 2116, 2117, 2119, 2120, 2121, 1890, 1975, 1976, 1977, 1978, 1979, 1980, 1981, 1982, 1983, 1984]. Book [1985, 1986, 1987, 1988, 1989, 1990, 1991, 1992, 1993, 1994, 1995, 1996, 1997, 1999, 3098, 3099, 3100, 3101, 3102, 3103, 3104, 3105, 3106, 3107, 3108, 3109, 3110, 3111, 2944, 2943, 2940, 2945, 2939, 2534, 3158, 2407, 2425, 3215, 855, 2622, 1501, 2536, 1952, 3553, 1856, 2284, 490, 1702, 1832, 3157, 3426, 348, 789, 1201, 3719, 1357, 738, 1843, 2695, 2735, 3342, 890, 2556, 1406, 800, 2431, 2307, 2811, 243, 637, 3063, 3436, 1753, 2076, 1158, 1840, 2175, 3776, 3071, 280, 2075, 1204, 812, 167, 849, 960, 2168, 2502, 2333, 3319, 3471, 402, 921, 44]. Book [768, 2101, 1553, 3235, 390, 1690, 621, 702, 1093, 1024, 2063, 2638, 2575, 3495, 3778, 1851, 73, 418, 242, 684, 1908, 247, 513, 2998, 924, 1121, 1752, 2215, 549, 268, 3314, 1044, 1849, 1405, 1974, 52, 1056, 1088, 2986, 3420, 1415, 2025, 2346, 862, 274, 1923, 156, 3467, 2614, 2555, 3164, 3600, 1030, 72, 1945, 45, 158, 1577, 3324, 1358, 2419, 51, 175, 562, 1018, 1021, 1665, 1939, 2661, 3240, 740, 191, 596, 2338, 3044, 2134, 2429, 2364, 1362, 2883, 1488, 2480, 2744, 2352, 2912, 2969, 2189, 794, 3753, 80, 1109, 1762, 2681, 3070, 3515]. Book [1685, 2038, 2434, 2767, 216, 743, 1704, 1736, 629, 1967, 940, 910, 2545, 446, 1504, 2761, 1951, 2739, 2260, 2885, 3243, 182, 731, 1349, 2060, 2137, 3180, 3496, 1371, 644, 907, 1716, 3220, 1218, 3730, 1206, 2146, 1807, 3142, 1122, 1264, 3640, 3222, 2704, 161, 712, 969, 1243, 3358, 1701, 2662, 3613, 2915, 2223, 2906, 1888, 2328, 3412, 3670, 2195, 2911, 3254, 3749, 1165, 1931, 783, 404, 3585, 2027, 2259, 632, 3166, 283, 843, 1689, 3646, 447, 2474, 1081, 1725, 464, 2936, 905, 1968, 2900, 853, 1815, 605, 1719, 3276, 3511, 1961, 2054, 2098, 3234, 1214, 1492, 210]. Book [221, 301, 569, 745, 1467, 1957, 49, 1523, 401, 595, 3250, 1664, 3549, 3380, 92, 398, 565, 2433, 3069, 2350, 3565, 3329, 2844, 2510, 202, 2834, 3756, 966, 2312, 911, 3507, 3175, 558, 1420, 2255, 2254, 2927, 665, 551, 2693, 3606, 1082, 985, 1494, 1828, 3308, 2698, 3335, 1084, 184, 2292, 2343, 1052, 845, 3394, 931, 2021, 3573, 3136, 1600, 3346, 1848, 611, 647, 2097, 2532, 679, 1660, 2627, 3177, 449, 884, 1394, 1713, 2051, 2133, 2264, 3056, 1626, 3754, 85, 2683, 324, 2680, 1299, 436, 1744, 3054, 1688, 1859, 2171, 2985, 1904]. Book [3711, 160, 2840, 3293, 2337, 223, 2086, 511, 1260, 1012, 1519, 2078, 1157, 2888, 1200, 375, 2058, 417, 2791, 3785, 254, 658, 771, 2270, 1265, 2449, 2881, 631, 205, 1772, 3521, 548, 1585, 742, 1616, 2088, 3389, 287, 634, 1950, 2085, 2288, 1497, 2894, 1300, 2620, 3267, 119, 2436, 2663, 31, 3561, 3635, 15, 818, 964, 2272, 3184, 2202, 3357, 1607, 1215, 3090, 3328, 179, 659, 1046, 485, 970, 3275, 545, 3261, 1000, 2630, 3303, 3783, 2267, 1496, 58, 25, 1141, 643, 654, 1760, 2732, 3083, 3072, 3046, 3325, 3338, 1090, 1502, 140, 3231, 3393, 2983]. Book [3713, 2285, 1506, 2753, 3130, 3376, 9, 359, 1808, 2667, 3662, 2081, 1962, 2068, 471, 30, 929, 3516, 2500, 3048, 181, 1009, 1578, 2557, 131, 1010, 3611, 1940, 592, 2048, 3247, 345, 2077, 164, 207, 1761, 3476, 3764, 460, 1359, 2244, 2827, 642, 888, 609, 2728, 900, 1105, 3685, 2981, 2381, 497, 3518, 3326, 378, 2770, 346, 468, 1666, 1727, 2220, 2528, 3341, 2784, 368, 1747, 2771, 3767, 734, 763, 3400, 2615, 975, 61, 1266, 3295, 648, 2164, 144, 982, 2132, 470, 2427, 193, 1307, 861, 894, 1867, 2621, 1423, 2720, 2895, 3391, 2196, 1855, 1038, 255, 452]. Book [1737, 690, 1412, 1833, 14, 3144, 668, 2829, 3221, 714, 2351, 3349, 2257, 533, 3735, 1746, 2490, 3055, 3710, 1971, 2578, 117, 519, 1278, 1397, 2613, 3059, 2212, 108, 707, 3029, 500, 1268, 871, 2896, 3390, 3085, 1522, 2045, 295, 2691, 2967, 2929, 522, 1137, 1738, 2209, 2700, 3591, 3405, 2385, 2828, 1483, 2141, 1873, 2378, 3009, 1347, 541, 776, 66, 290, 510, 1262, 847, 1388, 1941, 2754, 157, 2666, 2349, 62, 638, 3067, 980, 194, 773, 2418, 3736, 2282, 2416, 453, 366, 2610, 364, 1593, 2022, 896, 3183, 244, 2306, 3683, 2334, 1471, 2023, 499, 1925, 3316]. Book [3715, 554, 1216, 2491, 1613, 304, 1380, 2217, 2723, 3353, 3359, 1226, 3732, 844, 1842, 977, 1617, 277, 723, 1163, 2330, 3419, 2974, 57, 3379, 26, 2161, 2887, 521, 3686, 2091, 1103, 1827, 2265, 2671, 3161, 3248, 2737, 3237, 1301, 1277, 2042, 2733, 3525, 2924, 593, 3682, 299, 3079, 3608, 3227, 3304, 3645, 2542, 3574, 1679, 146, 2041, 2362, 1592, 2162, 3162, 3656, 2968, 3133, 1528, 1756, 1853, 3678, 2824, 406, 3398, 787, 1507, 2339, 3251, 869, 1026, 727, 3548, 109, 3336, 111, 69, 1612, 1668, 2093, 1858, 3340, 870, 3256, 504, 2741, 153, 2736, 3438, 3010, 3265]. Book [2559, 3429, 222, 3192, 396, 1295, 1674, 2403, 2061, 300, 1820, 1602, 365, 451, 687, 1487, 1683, 2277, 2531, 3300, 3725, 993, 2492, 106, 2632, 1627, 1006, 90, 639, 3551, 881, 3186, 3343, 494, 989, 1707, 2664, 3036, 1844, 557, 1115, 1809, 1909, 2810, 197, 1486, 1538, 1152, 3008, 3057, 2400, 3415, 3559, 1770, 3345, 125, 168, 3716, 271, 2612, 2184, 1508, 2722, 3583, 1392, 2435, 3745, 1129, 38, 133, 903, 1401, 1876, 2166, 2035, 2361, 2561, 3578, 8, 495, 2044, 2402, 3500, 1532, 413, 999, 877, 1587, 2099, 2221, 2050, 2609, 2738, 3053, 3395, 559]. Book [762, 918, 3201, 3242, 82, 531, 1411, 134, 892, 601, 850, 1297, 1020, 1946, 2210, 382, 3485, 2083, 2921, 2332, 1139, 3508, 3567, 2026, 1419, 2849, 2781, 3687, 257, 619, 906, 1284, 1466, 1710, 3271, 3633, 1509, 483, 3315, 3654, 1003, 1568, 3638, 2445, 151, 863, 2571, 2565, 475, 1583, 1142, 1261, 1463, 3339, 1944, 2886, 3642, 2275, 138, 1050, 2104, 3530, 3784, 1194, 2809, 1168, 1699, 1162, 2159, 1235, 1393, 393, 1389, 517, 467, 2430, 683, 717, 2854, 2892, 3546, 1482, 2170, 3241, 3502, 3005, 1551, 454, 2999, 266, 555, 491, 3052, 3576, 1190, 3671, 2052]. Book [298, 1116, 620, 1028, 1144, 3038, 1354, 2204, 409, 2836, 790, 1768, 173, 1164, 655, 701, 2703, 626, 83, 256, 1275, 2178, 267, 1217, 1474, 2406, 16, 1552, 2877, 3361, 636, 539, 616, 774, 1366, 1078, 2619, 825, 50, 803, 627, 1023, 1728, 759, 1527, 1143, 1257, 1363, 107, 3061, 2978, 3475, 1133, 512, 1077, 2089, 2748, 3255, 3399, 3506, 756, 2448, 2918, 3356, 2039, 2669, 2546, 3760, 526, 1465, 3141, 2262, 2772, 3014, 3513, 456, 608, 628, 3752, 2304, 758, 1742, 356, 1228, 1819, 29, 3408, 3757, 2308, 407, 1618, 976, 3233, 1475, 1381, 1921]. Book [380, 1751, 3058, 1022, 1404, 1670, 1414, 147, 2995, 419, 1469, 2426, 3657, 1907, 2568, 3087, 1922, 225, 1730, 70, 377, 1355, 1407, 1554, 2031, 2224, 1929, 2734, 3520, 3762, 507, 3094, 3679, 523, 1043, 2879, 1629, 938, 1008, 1814, 3323, 226, 678, 1232, 2242, 1947, 1193, 1579, 1049, 498, 1241, 1100, 1881, 3163, 1111, 2087, 95, 769, 927, 2850, 349, 663, 313, 646, 1031, 568, 1223, 2015, 2755, 2600, 3481, 1470, 2484, 1743, 1267, 1289, 915, 1741, 1620, 186, 696, 488, 1495, 1821, 2779, 3091, 1166, 2548, 1564, 1624, 2727, 757, 1529, 2473, 2535, 2295, 889, 2241]. Book [141, 2479, 2815, 213, 520, 1128, 1221, 1534, 1042, 3589, 2914, 3607, 1273, 2253, 354, 1293, 2169, 3522, 2764, 2477, 2298, 3580, 882, 995, 1095, 1948, 3252, 3660, 1240, 219, 2172, 2358, 3566, 199, 934, 2370, 2838, 653, 1230, 3153, 859, 1305, 1871, 1718, 1837, 1861, 2453, 3042, 1120, 3095, 2268, 96, 1590, 1822, 2922, 3043, 968, 1229, 872, 737, 3596, 1852, 703, 458, 2206, 2503, 2413, 2893, 3226, 3505, 3734, 3655, 1382, 1598, 2855, 3259, 3331, 3590, 3609, 1127, 1192, 372, 2065, 2984, 898, 2639, 2899, 3187, 808, 2562, 195, 797, 1279, 2725, 381]. Book [542, 2197, 1237, 263, 74, 2314, 2149, 3577, 374, 1732, 3311, 1140, 1584, 2258, 1135, 3727, 1304, 2821, 2287, 2437, 1290, 2991, 315, 2839, 3681, 2768, 1197, 67, 1402, 3035, 2841, 2372, 3707, 1055, 3557, 2225, 2188, 981, 1581, 2341, 128, 819, 1153, 2971, 1117, 2158, 2963, 2478, 1669, 2073, 3479, 1148, 1417, 2990, 3274, 2572, 2878, 1935, 3593, 1883, 2994, 169, 370, 1154, 1865, 2752, 3498, 253, 1258, 2670, 546, 21, 40, 1280, 3721, 2174, 218, 706, 2672, 2902, 310, 3641, 3643, 660, 1863, 3748, 384, 469, 936, 288, 1271, 1302, 3714, 496, 987, 2163, 2876, 3401, 2344, 1697, 2243]. Book [1076, 123, 1955, 1812, 1134, 53, 971, 2858, 603, 1775, 318, 1625, 1588, 37, 1958, 2607, 1533, 145, 613, 12, 305, 525, 728, 747, 1677, 2147, 2412, 2438, 3093, 2604, 1276, 2439, 394, 785, 1155, 1692, 1722, 1830, 2145, 1917, 2348, 2673, 2890, 3490, 998, 319, 2668, 2919, 3086, 3337, 3217, 1477, 347, 2103, 317, 1868, 3312, 3317, 3392, 79, 527, 809, 1159, 1520, 2070, 2635, 1365, 136, 935, 1249, 1167, 1270, 2817, 778, 3741, 666, 1963, 2682, 2692, 1754, 3690, 786, 2631, 39, 1591, 1150, 1675, 2443, 118, 2780, 1256, 166, 1461, 89, 203, 421, 1196, 1244]. Book [1131, 851, 1614, 1874, 1966, 3718, 1734, 1949, 630, 695, 2889, 1970, 2371, 3474, 824, 127, 583, 1195, 858, 3588, 1118, 2067, 3636, 899, 1138, 2599, 3482, 2505, 155, 1102, 3001, 2017, 1236, 1717, 3728, 251, 537, 720, 1027, 1759, 1615, 2778, 3777, 1854, 2624, 901, 1089, 2593, 3413, 3550, 2410, 2820, 2746, 459, 1838, 11, 1421, 2398, 2305, 2686, 278, 661, 1835, 2861, 3742, 1536, 3769, 2345, 149, 1777, 1473, 1250, 294, 383, 530, 2916, 2279, 2560, 2191, 2640, 2404, 2997, 3375, 1033, 3431, 2634, 1132, 2554, 2928, 3229, 3680, 729, 1468, 1972, 1936, 3181, 400, 1211]. Book [534, 719, 3068, 3330, 3765, 3466, 1124, 3149, 1400, 2982, 3639, 3472, 3410, 625, 2182, 2608, 505, 983, 2046, 3378, 1544, 2913, 2757, 474, 864, 1047, 2880, 3051, 1545, 891, 2905, 810, 502, 724, 1037, 1726, 1826, 3302, 2777, 2192, 478, 24, 770, 2553, 2904, 1913, 713, 904, 1831, 2074, 2049, 2596, 2688, 3066, 2993, 3296, 1709, 2297, 2740, 2030, 412, 2926, 932, 1662, 588, 2825, 3299, 3402, 463, 3143, 3309, 1857, 2151, 515, 473, 1224, 865, 369, 2037, 2294, 2504, 110, 1428, 2296, 1813, 2976, 699, 2199, 2447, 2903, 1281, 1841, 1973, 922, 1872, 3262, 1126]. Book [3219, 1546, 715, 208, 220, 2399, 2616, 1695, 2606, 2843, 2932, 1282, 2790, 681, 1491, 1817, 3649, 3532, 3726, 618, 229, 1912, 2420, 88, 269, 704, 933, 1094, 3131, 2444, 48, 1212, 566, 2303, 1885, 2699, 2092, 201, 414, 1960, 2177, 2300, 351, 3077, 775, 804, 916, 2036, 598, 2094, 482, 2428, 1099, 2547, 3159, 3075, 516, 1199, 1918, 2266, 3440, 3528, 1806, 2980, 2550, 805, 3360, 2907, 228, 529, 744, 2758, 1343, 1586, 3733, 1245, 178, 782, 1234, 3154, 1294, 1766, 3272, 316, 1036, 1834, 2082, 1379, 1887, 1210, 1418, 721, 764, 3334, 2187, 2856, 3033, 3469, 735]. Book [1087, 19, 357, 586, 1231, 3403, 192, 2658, 3434, 3318, 1711, 2917, 411, 97, 176, 765, 1731, 3397, 20, 1112, 2859, 3257, 1700, 416, 455, 1123, 308, 563, 2095, 1383, 2488, 1758, 2374, 3073, 3610, 187, 2641, 2486, 3387, 2356, 2756, 3139, 708, 1698, 1106, 1773, 2222, 2379, 2150, 3768, 1671, 3170, 139, 1296, 923, 2140, 188, 3037, 1048, 2367, 1425, 766, 1714, 1604, 3096, 1882, 3637, 450, 826, 1285, 2665, 3717, 2602, 172, 1575, 2446, 3422, 3663, 1015, 2996, 441, 1959, 1903, 518, 1460, 1774, 2884, 3152, 3652, 3737, 3477, 3307, 2816, 200, 1114, 3182]. Book [442, 2696, 3684, 1880, 2327, 3050, 1739, 3155, 607, 59, 64, 3224, 2783, 2601, 1259, 1476, 1014, 726, 480, 1083, 445, 1705, 2674, 1549, 1351, 2705, 3751, 3160, 779, 1110, 1524, 227, 3347, 42, 358, 795, 876, 1703, 2411, 1283, 3430, 1513, 1521, 322, 1764, 1203, 1755, 2084, 2441, 3030, 185, 2567, 114, 547, 2311, 3151, 3599, 1298, 321, 1019, 1498, 2909, 1213, 3484, 3517, 2401, 2679, 323, 1011, 1569, 1547, 3012, 1306, 1130, 1222, 1378, 3545, 1361, 248, 3724, 259, 282, 722, 1119, 1720, 1847, 1906, 3232, 2637, 1942, 3759, 1776, 1233, 2152, 1846, 2830]. Book [196, 3198, 2910, 2497, 2594, 265, 866, 550, 3174, 3501, 1225, 1113, 2156, 651, 1145, 991, 2190, 2851, 3592, 47, 657, 2476, 60, 371, 1274, 3558, 65, 2347, 3047, 1510, 170, 914, 3676, 1303, 2214, 606, 309, 1757, 2660, 2684, 152, 2053, 2708, 2595, 3439, 2846, 1426, 440, 2762, 1377, 3351, 320, 2835, 2845, 3468, 2415, 2043, 1916, 3065, 1589, 1954, 2432, 3189, 1723, 2283, 408, 1146, 3519, 93, 3424, 669, 1075, 1205, 1571, 1877, 2837, 990, 2180, 3078, 784, 941, 1535, 2749, 506, 1512, 2020, 3000, 3604, 2286, 3665, 250, 1680, 806, 822, 204, 2966]. Book [285, 1920, 3195, 1543, 3146, 3729, 481, 1219, 798, 2747, 2657, 3185, 3354, 3773, 337, 509, 1910, 2544, 1691, 2603, 937, 1253, 1040, 1810, 3584, 902, 2605, 71, 2617, 3011, 698, 1818, 2833, 472, 571, 1431, 1490, 1926, 2355, 2726, 3218, 1765, 1034, 3675, 2495, 3428, 3327, 3480, 3569, 28, 397, 567, 589, 887, 1247, 1943, 78, 928, 610, 1505, 2570, 716, 148, 580, 1227, 878, 2731, 3575, 3739, 1209, 1611, 813, 2776, 1207, 1239, 3489, 2558, 1348, 2263, 880, 1462, 2685, 1098, 2373, 1823, 3138, 1721, 1869, 3774, 524, 1238, 967, 2016, 2842, 3527, 2812, 3674, 3667, 553]. Book [1248, 2623, 1051, 297, 767, 1667, 1745, 3650, 1136, 992, 2247, 2383, 3167, 3668, 3594, 2193, 1608, 1107, 17, 2299, 2626, 2988, 2813, 692, 448, 788, 2040, 2203, 2628, 3744, 1531, 1481, 2750, 3039, 1573, 32, 3779, 875, 1696, 3015, 652, 739, 1540, 86, 143, 2301, 1619, 2496, 399, 986, 570, 479, 2143, 3708, 1621, 2219, 3435, 3503, 1242, 2165, 2252, 777, 1599, 1582, 3168, 1427, 2964, 3441, 3661, 2563, 1609, 2340, 1004, 1376, 405, 1479, 2908, 3560, 137, 273, 275, 379, 746, 997, 1104, 2310, 2643, 3614, 180, 2423, 1493, 484, 2269, 2167, 2485, 2384, 224, 1500, 150]. Book [361, 874, 1374, 174, 373, 249, 585, 2028, 3313, 656, 3294, 302, 1387, 1356, 2751, 3264, 2386, 1035, 1550, 1938, 2860, 3669, 1503, 2819, 1345, 2499, 2689, 3529, 1969, 3555, 645, 540, 1539, 1570, 3076, 2930, 1595, 1017, 2659, 3298, 3488, 2359, 3598, 162, 360, 1092, 2822, 3659, 493, 2743, 3497, 3740, 1676, 2442, 1054, 2422, 1147, 376, 1403, 303, 33, 1499, 1360, 2256, 3509, 3651, 2853, 353, 2826, 2072, 3514, 3003, 3746, 3478, 1408, 2409, 979, 3225, 260, 2935, 292, 344, 2569, 3411, 3672, 367, 3193, 113, 2057, 552, 2618, 2891, 3492, 2576, 3278, 3526, 725]. Book [1353, 2205, 2276, 3006, 3586, 3743, 34, 466, 1580, 2249, 1079, 2019, 1915, 694, 1630, 796, 640, 477, 3486, 2273, 2759, 3244, 3322, 1933, 1367, 2290, 2760, 3499, 461, 2336, 2250, 3568, 755, 171, 341, 1001, 2357, 2832, 3140, 2543, 2139, 2365, 926, 2024, 2090, 177, 2577, 120, 2326, 2501, 75, 18, 2625, 2454, 732, 2366, 23, 1085, 326, 1706, 3731, 1464, 1724, 94, 385, 677, 893, 1566, 2176, 2923, 3758, 1511, 2925, 1605, 1429, 68, 799, 1013, 2194, 2376, 3605, 1763, 2687, 2342, 13, 604, 1572, 2451, 420, 3253, 130, 1287, 1565, 2034, 1286, 289, 689]. Book [1525, 2302, 2642, 3032, 3556, 3712, 2775, 312, 2066, 2920, 1292, 3409, 3775, 1693, 1220, 1914, 802, 116, 680, 1364, 87, 978, 772, 2245, 2315, 1537, 1594, 2363, 1410, 2636, 1430, 1729, 2677, 543, 1748, 710, 885, 886, 1375, 3570, 2745, 2293, 2218, 1369, 536, 2360, 126, 264, 3470, 1694, 286, 1478, 1686, 3722, 3581, 2979, 3060, 3238, 3571, 2730, 3148, 284, 3178, 3747, 984, 352, 617, 3612, 501, 2375, 3200, 261, 2675, 3666, 2055, 2970, 1878, 307, 1373, 3582, 1932, 2056, 1396, 2144, 535, 1045, 2597, 77, 1517, 912, 1514, 1934, 1041, 846]. Book [2965, 3761, 2787, 63, 3348, 3437, 281, 780, 591, 1631, 3165, 919, 3082, 2977, 528, 649, 1352, 1769, 1610, 2702, 3273, 3563, 3564, 1518, 3191, 1091, 3770, 3305, 1672, 3199, 760, 1370, 1413, 2281, 2157, 2701, 3089, 3384, 154, 2489, 2785, 3156, 437, 3523, 258, 879, 3493, 476, 1924, 2248, 2574, 641, 355, 457, 492, 1811, 3491, 1712, 2475, 3239, 3404, 996, 3332, 3750, 3423, 3504, 3333, 3432, 2251, 503, 3228, 1681, 486, 3190, 115, 963, 2540, 3579, 3677, 2862, 2179, 2774, 36, 3723, 1080, 3028, 2198, 2278, 387, 590, 1002, 386, 3524, 3720, 1530, 2629]. Book [1086, 1057, 2508, 3483, 91, 415, 2102, 2788, 3688, 3064, 3223, 1391, 306, 1733, 2709, 817, 3587, 1541, 3355, 816, 444, 1489, 2380, 81, 206, 215, 581, 1208, 1108, 1682, 1879, 1884, 2135, 2032, 2369, 2417, 2566, 2766, 3041, 2973, 3510, 3634, 1074, 1395, 1845, 2901, 801, 1422, 2154, 3081, 792, 1708, 2487, 3245, 350, 338, 2018, 3045, 2847, 622, 1161, 962, 3088, 2552, 3179, 3673, 3320, 245, 2059, 2080, 3473, 2611, 3396, 403, 842, 2440, 3763, 3134, 2096, 2814, 2069, 339, 2676, 3766, 132, 231, 3173, 2549, 3321, 1350, 10, 709, 1485, 2335, 741, 1542, 3615]. Book [2160, 2897, 3260, 27, 1096, 2706, 2421, 291, 3263, 3388, 988, 2271, 2721, 3249, 3417, 587, 2690, 438, 600, 3416, 3487, 1886, 538, 1016, 1875, 2064, 1623, 3080, 1149, 965, 2181, 2769, 811, 930, 1398, 3771, 1386, 2100, 2313, 1125, 1372, 3554, 1390, 2573, 246, 3135, 296, 2377, 3176, 3407, 1673, 2213, 1911, 1516, 682, 2408, 2261, 3246, 815, 2931, 2047, 1825, 84, 214, 279, 1767, 142, 1905, 3780, 422, 3494, 3709, 3188, 314, 410, 925, 1053, 1866, 3092, 2183, 2564, 1399, 584, 3007, 340, 392, 3383, 2694, 615, 391, 667, 662, 913, 1715, 293, 1663, 2201]. Book [43, 54, 814, 3031, 3664, 2331, 3171, 1480, 1771, 3258, 3427, 3632, 2148, 3782, 3648, 1816, 2155, 325, 895, 3547, 1956, 2511, 3230, 1603, 3562, 22, 135, 2598, 2494, 1953, 3531, 159, 1007, 1661, 1601, 1927, 2274, 2765, 3301, 3382, 272, 2633, 1684, 1191, 1472, 121, 1384, 1385, 705, 230, 3145, 3385, 564, 395, 2136, 1151, 1678, 1515, 1005, 1416, 2289, 3266, 3603, 3421, 3602, 3781, 124, 854, 2481, 3406, 3137, 129, 2773, 793, 1039, 2483, 1409, 1272, 1291, 920, 1269, 1597, 3297, 163, 209, 860, 2291, 3049, 917, 2818, 2707, 2424, 3442, 2972, 46, 633, 807, 2071]. Book [3097, 388, 1805, 791, 2538, 3147, 3377, 76, 311, 1032, 1749, 1548, 2216, 2529, 3197, 3268, 122, 594, 686, 1344, 2062, 2414, 343, 857, 3062, 189, 820, 848, 2848, 3013, 612, 2541, 1622, 2507, 2309, 3772, 1864, 55, 2246, 56, 635, 2186, 3172, 3418, 335, 1202, 939, 2506, 389, 730, 2452, 1740, 3512, 1606, 1964, 2763, 1735, 3350, 2368, 700, 823, 1288, 961, 1836, 2207, 3084, 3433, 212, 514, 3414, 561, 2509, 2551, 582, 544, 3352, 35, 614, 2280, 3597, 868, 1198, 2033, 1567, 1687, 2992, 3150, 3738, 3425, 697, 733, 883, 1596, 1965, 2200, 2329, 3595, 252]. Book [487, 41, 112, 336, 508, 624, 664, 688, 718, 1246, 856, 1101, 1829, 1839, 1919, 2975, 3653, 556, 363, 2029, 443, 2153, 2539, 1862, 532, 270, 599, 1860, 897, 1750, 3658, 2079, 2482, 2729, 2211, 2138, 597, 1937, 3074, 3310, 2533, 1930, 2934, 3306, 3644, 1526, 1850, 1928, 3004, 211, 560, 711, 1251, 972, 1346, 2782, 362, 1628, 1824, 821, 2786, 3270, 2325, 623, 1097, 2382, 2537, 2831, 439, 685, 1252, 2450, 3216, 3277, 1368, 3689, 909, 217, 781, 3386, 2724, 2354, 198, 2857, 190, 276, 693, 165, 2142, 3381, 3552, 262, 3196, 342, 1156, 1424, 2530]. Book [3169, 1574, 462, 466]. Books [2472, 2865, 1095, 1196, 2390, 3651, 1364, 1631, 3381, 1578, 3389, 963, 3147]. Bookseller [2319]. Boom [916, 1744]. Boomgaard [1348]. Border [2346, 2067, 661, 1208]. Borderlands [3352]. Borderline [1892]. Borders [2060, 1038, 767, 3720]. Borges [2629]. Boring [1715]. Boris [3320]. Born [2681, 2243]. Borrelli [3727]. Borsay [1829]. Bos [3312, 2410]. Bosch [2676, 1016]. Bose [645]. Bosons [675]. Boss [304]. Bossolo [83]. Boston [1112, 3512]. botanica [2401]. Botanical [1009, 3219, 1083, 3509, 1878, 1825]. botánicas [768]. Botanik [347]. Botanist [1721]. Botanophilia [1099]. Botany [2610, 2275, 1490, 3509, 834]. Botey [586]. Both [519]. Bothered [2778]. Boudet [354]. Boudia [1111].

Bougainville [3489]. Bought [1972]. Boulaine [1084]. Boulnois [982]. Boundaries [1038, 1624, 11]. Boundary [1892, 912, 2552]. Boundless [478]. Bounds [2678]. Bounty [3144]. Bourdelais [2811]. Bousé [729]. Bouveresse [2820]. Bouwsma [786]. bóvedas [2662]. Bowden [2944]. Bowen [2206]. Bowerbank [2667]. Bowers [481]. Bowker [781]. Bowler [2720]. Bows [3713]. Box [1865, 1926, 3193, 1864, 3425]. Boy [1828, 2866]. Boyd [2669]. Boyle [3158, 960, 2974, 1590, 3158, 1502, 580]. Boys [1533, 1759]. Bozeman [314]. Brace [757]. Brackenridge [1895]. Bradley [3315]. Brady [764]. Bragg [1797, 2298]. Brahe [3235, 703, 39]. Brahmin [1112]. Brain [973, 401, 2334, 1116, 377, 1076, 1212, 2213, 3658, 3384]. Brains [3716]. Brand [1909]. Brand-New [1909]. Brandon [618]. Brandy [797]. Brantlinger [1833]. Brasil [2495]. Brauer [888]. Braun [1415, 3510, 1516]. Bray [3634]. Brazil [2361]. Brazilian [903]. Break [3091]. Breaking [2848]. Breast [2193, 735]. Breath [1938]. Breathing [3220]. Breeding [2849, 1395, 3302, 397]. Breeds [2225]. Breidbach [3047, 55]. Bremen [909]. Brenda [113, 3067]. Bresadola [2352]. Brésil [1736]. Bret [364]. Brett [368, 2684, 2821, 3007]. Brian [45, 2812, 1957, 970, 1592, 3038, 2169, 916, 2289]. Bridge [1811, 1609]. Bridges [1776]. Bridging [2787]. Brief [164, 3057, 1726, 3618, 1346]. Briefe [48, 264]. Brieftagebuch [170]. Briefwechsel [74, 630, 1847, 1714]. Briggs [1070]. Brigham [2932]. Bright [2020, 187]. Brigitte [770, 561]. Brill [3154]. Brilliant [3178]. Bringing [774]. Brinkley [1852]. Briskorn [909]. Brissoni [2099]. Bristol [757]. Britain [905, 184, 3754, 2985, 3072, 3162, 1627, 2854, 3760, 2149, 1722, 2909, 3584, 3675, 379, 1125, 272, 820, 1101, 2866, 3540, 717, 1837, 2550, 2679, 2891, 641, 2424, 1636, 1746, 1288]. Britanniques [2894]. British [3040, 348, 3778, 2834, 2292, 3349, 1388, 844, 1679, 3576, 278, 2740, 1882, 1114, 3151, 2286, 3668, 180, 2423, 376, 3411, 43, 1678, 1121, 1336, 2236, 3117, 2412, 3399]. Britons [2668]. Britta [171]. Britton [3509]. Broadcast [918, 816]. Broadcasting [651]. Broadening [2261]. Broader [1494]. Brock [1967, 3665, 656, 591, 2937, 3607]. Brockman [738]. Broecke [2481]. Broelmann [2056]. Broken [1597, 3090]. Broker [3353, 3679]. Broks [3010]. Bronfen [253]. Bronn [3670]. Brook [3424]. Brooke [600, 1071]. Brookhaven [1051]. Brooks [1165, 3074]. Broom [3619]. Broome [1131]. Brosche [20]. Brose [131]. Brotons [146, 586]. Brown [2076, 2923, 1161, 15, 222, 646, 46]. Browne [2670, 1720]. Bruce [794, 824, 669, 686, 890, 2983, 2294, 1895]. Bruch [3761]. Brückner [2827, 850]. Bruno [138, 266, 356, 39, 33, 3355, 2823, 1933, 3148]. Brush [737]. Brutal [3156]. Brute [2427]. Brutes [1105]. Bruyn [2171]. Bryson [1834]. Bub [73]. Buchanan [656]. Buchwald [3410, 2751]. Buck [3775]. Buckyball [1358]. Bud [295, 133, 762]. Bueno [768, 2829]. Buffon [44, 1100, 44, 3400]. Bugs [2520]. Buhs [2631]. Builder [873, 872]. Building [2307, 2434, 3222, 2461, 496, 2931, 2551, 1744, 252]. Built [1949]. Buisseret [2153]. Bulmer [2901]. Bunsen [48]. Burbank [674]. Burd [436]. Burden [919]. Burdens [259]. Burdman [2617]. Burdon [1853]. Bureau [280, 2234]. Bureaucracy [606]. Bureaucratic [942]. Burgerliche [337]. bürgerlichen [635]. Burgio [777]. Burian [2572]. Buridan [1817]. Burkhardt [2615, 3458, 805, 3169]. Burnett [1114, 621]. Burney [3328]. Burnham [125]. Burns [1408]. Burnside [888]. Burrow [2170]. Burrows [326]. Bursting [2743]. Burton [2970]. Busard [3145, 147]. Bush [3089]. Bushnell [2256]. Business [3338, 1380, 413, 1812, 2624]. Butenandt [2117]. Butler [3597, 541, 215]. Butterflies [1679, 3752]. Butterfly [1881]. Butters [353]. Buttons [3147]. Buzz [2696]. Byers [2939]. Bynum [3031]. Byrd [858, 212].

C [852, 2937, 1851, 2545, 1081, 2292, 511, 375, 1607, 471, 2330, 1627, 197, 125, 3315, 50, 1921, 96, 898, 2437, 1883, 747, 421, 2297, 2980, 2856, 2665, 3224, 2830, 371, 6, 204, 2623, 992, 137, 3529, 3555, 376, 312, 772, 415, 206, 403, 741, 988, 1416, 2529, 2507, 2309, 487, 1919, 1628, 781, 1820, 379, 2979, 352, 486, 1829, 1166, 1921, 1697, 3598]. C. [921, 2098, 2531, 922, 2301, 2551]. Côme [701]. ca [2156, 2729]. Cabala [611, 1819]. Cabinet [3396, 881]. Cabinets [3044]. Cabot [2548]. Caccamo [1022]. Cadavers [1810]. Cadeddu [2988]. Caffey [2835]. Cahan [757]. Cahiers [3311]. Caiazzo [2041]. Cajal [2444]. cajas [2133]. Calculating [3344]. Calculation [2883]. cálculo [2662]. Calculus [442]. Calder [3005]. Caldwell [2294]. Calendar [243, 140, 980]. Calendars [2146, 980, 3724, 464]. California [1940, 559, 775, 2900, 2089, 1751]. Californian [286]. Californian-Australian [286]. Call [2883]. Called [1565]. Callendar [3255]. Calling [1594]. Calvert [176]. Calvinist [1093, 2273, 2487]. cambio [2430]. Cambridge [2113, 1987, 3109, 2945, 2472, 1664, 3785, 1814, 3589, 3252, 2876, 1697, 2593, 358, 3668, 3003, 3722, 3190, 1708, 1472, 2483, 1839, 1937, 3196, 1353, 177]. Cambrosio [2299]. Came [853]. Camel [3473]. Camerota [2411]. Camille [503, 2721]. Camillo [58]. Camp [2186]. Campaign [2692, 1942]. Campaigns [3498]. Campanus [3145]. Campbell [375, 203, 1061]. campo [1705]. Campus [3523, 3606]. Campuses [431]. Can [2805, 3210, 1130, 897, 935]. Canada [2307, 280, 2058, 2929, 2492, 2682, 3077, 2442, 1860, 1252, 1301]. Canaday [2365]. Canadian [3108, 1278, 3345, 774, 1208]. Canadian-American [1208]. Canal [3020]. Canals [179, 273]. Cancer [1577, 1756, 1304, 2682, 1123, 1185, 2193, 3618, 548]. Candid [1417]. Candolle [2547]. Canestrini [1599]. Canham [114, 114]. Canine [546]. Canning [2214]. Cannon [821]. ’Cannot [299]. Canonical [1368]. Canonis [462]. Canterbury [2154]. cantieri [691]. Cantor [1515, 600, 584]. Canyon [3525]. Cao [3261]. Capabilities [396]. Cape [2556, 45]. Capelotti [542]. Capitalism [27, 3523, 2933]. Capitalist [2083]. Caplan [655]. Cappadocia [610]. Capria [2940]. Caprina [2257]. Capshew [320]. Captain [1743]. Carbon [3255]. Cardano [1499, 2253, 2403]. Cardiac [237, 1155]. Cardiff [2714]. Cardinal [3730]. Cardiology [2081]. Care [1433, 182, 2844, 3325, 1123, 2218, 391, 1771, 2704, 1155, 258]. Career [238, 3118, 2124, 1376, 807]. Careers [740, 1852, 2327, 279]. Caribbean [2165]. Caring [855]. Carl [170, 1016, 163, 3418, 3414, 2089, 1153, 1416, 811]. Carl-Henry [2089]. Carla [509]. Carlat [720]. Carles [2995]. Carlo [277, 226]. Carlos [190]. Carlson [488]. Carman [436]. Carmen [2632, 732]. Carnap [667, 3589]. Carnegie [1262, 1077]. Carneiro [2242, 2561]. Carney [2161]. Carnival [17]. Carol [2190, 47, 91, 2544]. Carolina [343]. Caroline [111, 892, 3680, 514, 2168, 2062, 2062]. Carolingian [3386]. Carolyn [1940, 2129, 1949, 1425, 2034]. Carone [2969]. Carozzi [161, 2886, 712]. Carpenter [83, 597]. Carrier [390]. Carriers [2080]. Carrington [3408]. Carroll [3744]. Carson [771]. Carsten [2916]. carte [2277]. Carter [511]. Cartesian [3781]. cartesiane [156]. Cartilages [436]. Cartographers [933]. Cartographic [2244, 149, 2457, 1843]. Cartography [2244, 3569, 2965, 1096, 965, 2539, 3569, 342]. Cartoons [3623]. Cartwright [2812]. cas [2988]. casa [2044]. Casaubon [706]. Case [1059, 2582, 470, 7, 1302, 1970, 675, 3469, 2603, 2685, 2002, 3499, 2069, 22, 2598]. Cases [3358, 2023]. Cash [3166]. Cashin [990]. caso [2986]. Caspar [2543]. Casper [837]. Cassandra [2577]. Cassidy [1002, 2529]. Cassini [626, 968]. Cassirer [324, 229]. Castejón [1399]. Castellani [529]. Casti [1349]. Castonguay [2492]. Castro [2279]. catalans [1462]. Catalogue [940, 1807, 1166, 97, 2016, 2040, 2414]. Catalogues [1054]. Catastrophe [1041]. Catastrophic [1289, 2760]. Categories [3540]. Catesby [247]. Cathedral [1719]. Cathedrals [336, 872, 3667]. Catherine [2544, 2745, 443, 2933]. Cathode [192]. Catholic [2844, 751, 2605]. Catholicism [3638, 3019, 3638, 270]. Cathy [1808, 899]. católica [3053]. Caton [694]. Caucasus [772]. causa [2737, 3563]. Causality [2597]. Cause [1946, 266, 3669, 597]. causes [1686]. Cavalli [2641]. Cavalli-Sforza [2641]. Cavanilles [2829]. Cavazza [2726]. Cavendish [2889, 3001, 710, 1375]. Cay [2372]. Cay-Rüdiger [2372]. Cayley [2989]. CD [2296]. CD-ROM [2296]. Ceccarelli [2023]. Cecilia [1061, 2729, 2729]. Cédric [2619]. Ceglia [994]. Celebrate [3196]. Celebrating [1625, 1978]. Celebration [2648]. célébrations [3739]. celebres [354]. Celebrity [1262, 3192, 2123]. Celeste [1880]. Celestial [461, 1344]. Celestiall [2415]. Cell [2852, 2933, 2120, 2446, 3198, 257]. Cells [2300]. Cellucci [226]. Censorship [945]. Census [2472, 1693, 1252]. Centenary [2022, 192, 177]. Centennial [596, 381, 1933]. Center [3773, 1884, 312, 214, 3350]. Central [974, 3342, 280, 1024, 1202]. centrali [58]. Centuries [193, 3735, 1365, 2726, 591, 3473, 1451, 123, 770, 588, 2557]. Centuries/Dagli [3735]. Century [3194, 2987, 1433, 1632, 1633, 2111, 1990, 2943, 2939, 2472, 1974, 596, 2429, 910, 1122, 3412, 3276, 3511, 1626, 1744, 1506, 929, 3516, 1266, 193, 871, 2209, 3009, 2334, 2924, 299, 3162, 396, 3415, 271, 903, 1509, 1003, 1583, 1463, 683, 1077, 2669, 526, 3141, 1404, 1922, 1743, 2548, 1624, 2473, 1128, 2914, 1948, 3505, 3259, 3590, 1237, 3727, 2994, 1865, 2174, 706, 310, 3748, 1697, 2668, 2631, 3001, 2554, 3680, 2982, 2608, 1817, 2699, 1099, 2266, 2356, 1513, 1755, 1378, 3759, 1512, 1253, 1926, 1239, 2623, 3650, 2193, 2299, 3744, 3661, 2123, 2409, 367]. Century [3322, 3499, 3731, 385, 1511, 2687, 284, 2375, 3491, 3504, 206, 3417, 1825, 391, 2633, 2424, 1548, 2529, 3062, 3772, 1288, 1246, 2539, 2729, 362, 1424, 2718, 1068, 1186, 2954, 2319, 3693, 3668, 2717, 834, 1449, 673, 191, 818, 3083, 3059, 2929, 69, 147, 2197, 37, 809, 2428, 516, 114, 3773, 928, 1136, 739, 2219, 886, 1932, 879, 816, 2160, 913, 272, 2529, 944, 1784, 1446, 2108, 1978, 1501, 3235, 3175, 1738, 3265, 976, 938, 1280, 2243, 3765, 3440, 1130, 1227, 1676, 3246, 895, 3301, 1749]. Ceramic [1676]. Cercignani [277]. Cerebralis [24]. CERN [675]. Cerruti [2173]. Certain [2681]. cervello [2250]. Cesarani [1159]. Ceteris [1981]. Ceylon [2909]. cha [3186]. Chabás [2404]. Chabrán [884]. chacun [825]. Chadarevian [1421]. Chadwick [184]. Chaga [86]. Chaim [2974]. Chains [3623]. Chair [2593]. Chairs [771]. Chakrabarti [2677]. Challenge [3251, 3368, 2294, 3526, 2378]. Challenged [3499]. Challenges [2110, 2680, 2741]. Chamberlain [884]. Chambers [3241, 3151]. Champagne [361]. Champions [65]. Chan [1921]. Chance [268, 2673, 1773, 656, 2664, 137]. Chandak [638, 2918]. Chandler [2633]. Chandra [1140, 645]. Chang [2734, 311]. Change [1180, 3144, 850, 927, 2371, 766, 355, 206, 3020, 3268, 2789, 2705, 248, 1772, 1297, 3255, 1741, 2593]. Changed [569, 125, 2443, 764, 1416]. Changing [1637, 2630, 1211, 114, 753, 3266, 1064]. Chaomei [1955]. Chaos [633, 2220]. Chaoui [2560]. chapitres [2822]. Chaplin [3048, 1927]. Chapter [2245]. Chapters [1683]. Char [2095]. Character [899]. Characters [554]. Charakterisierung [2183]. Charap [2085]. Charcot [2178]. Chardin [2072]. Chareix [3316]. Charges [2583]. Charis [2656]. Charisma [3375]. Charity [1829]. Charlatanism [3574]. Charles [3158, 1851, 513, 2912, 2171, 1012, 3591, 366, 3745, 808, 805, 1106, 78, 275, 2057, 1149, 2274, 560, 711, 3169, 908, 3002, 2947, 890, 1908, 2081, 1359, 888, 2754, 2484, 2682, 2825, 1608, 1410, 3600, 724, 1720]. Charles-Lucien [513]. Charleston [632]. Charlotte [267, 3069, 1092]. Chart [122]. Charting [359, 1700]. Chasing [2476]. Chaucer [2154]. Cheaper [2567]. Cheated [1529]. Checklist [1697]. Chemical [2333, 2027, 631, 2349, 844, 3567, 305, 3317, 1591, 330, 1343, 547, 3000, 2269, 3298, 434, 895, 2633, 1384, 2373]. Chemicals [1028]. Chemie [3227, 1017, 3227, 1508]. Chemikers [194]. cheminement [3639]. Chemische [3715]. Chemischer [171]. Chemist [547, 2380, 2487, 991]. Chemistry [1358, 2027, 631, 2983, 9, 1747, 333, 2593, 3550, 770, 2428, 1521, 114, 2659, 2581, 341, 1001, 879, 325, 2938, 2472, 2744, 75, 1808, 2073, 2916, 306, 2136, 844, 3304, 3580, 458, 1233, 3667, 2064]. Chemla [2822]. Chen [1955, 2276]. Chene [865, 878]. Cheng [589, 3262, 3200]. Cheng-Chung [589]. Chernobyl [1041, 1041]. Cherry [1130, 911]. Cheyne [713]. chez [502]. Chiara [3147]. Chicago [937, 2391, 2690]. Chicxulub [1213]. Chikara [2489]. Child [1323, 1229, 752, 321, 2656, 2214]. Childbed [773]. Childbirth [2052, 2243, 694]. Childhood [1758, 1942]. Childrearing [3518]. Children [471, 1861, 317, 2866, 3468, 367, 599]. Chilvers [3320]. chimica [2173, 2267]. Chimie [1352]. Chimique [2312]. Chimisso [2222, 3666]. China [1783, 3553, 85, 2683, 3561, 3275, 3261, 2655, 2809, 2158, 2963, 2412, 3219, 3430, 1521, 3517, 1092, 979, 2249, 543, 1369, 3200, 3634, 3268, 3306, 3277, 2683, 2953, 2655, 2954, 2955, 6, 2957, 2956, 2027, 695, 2375, 2248, 1096, 854, 1484, 3479, 1715]. Chinatown [2377]. Chine [2822]. Chinese [2259, 31, 1486, 1404, 2158, 1813, 3333, 2533]. Chinnici [2277]. Chipley [1058]. Chiro [2562]. Chirurgien [627]. Chisholm [3607]. Chloroform [694]. Cho [653]. Choice [2083, 1276, 3469, 656, 3782]. Chompalov [3524]. Chooses [3548]. Choosing [753]. Chor [2561]. Choreography [2656]. Chr [264]. chrétienne [3649]. Chrimes [273]. Chris [373, 2701]. Christ [1240, 3730]. Christiaan [3316]. Christian [737, 935, 1826, 478, 1887, 3012, 866, 716, 3581, 2511, 2077, 1907, 2665, 2965]. Christiane [2881]. Christianity [1992, 2328, 3295]. Christianson [601, 703]. Christine [1105, 896, 1965]. Christl [1912]. christlichen [1574]. Christoph [1815, 3035, 3477, 2731, 2751, 126, 363]. Christophe [2780]. Christopher [184, 1166, 3320, 3031, 2076, 847, 2548, 2686, 2816, 997, 930]. Christus [1574]. Chromosome [2775]. Chronicle [3432]. Chronicles [1297, 3170]. chroniques [1686]. Chronological [1106]. chrysopée [2337]. Chrysophiles [1894]. Chrysophobes [1894]. Chu [870]. Chung [589]. Chunglin [2724]. Church [2609, 336, 2612, 945]. Churella [396]. Chyldren [481]. Chymistry [3496, 3304, 2892]. Chymists [3304]. Ciência [2561]. ciências [2495]. CIA [2217]. cicala [3138]. Cidadão [3327]. ciel [2277]. Ciencia [2742, 2264, 714, 1492, 300, 2386, 190, 2474, 2829, 1462]. Ciencias [3056, 2444, 1976, 732]. Científica [3732, 2742, 449, 3174, 3604, 3327, 1242]. Científica/Beyond [3732]. científicos [2133]. Cincinnati [1605]. Cinema [2816]. Cinematic [2913]. Cinematographic [1186]. CINP [404]. circa [2952]. Circle [632, 847, 1231, 440, 3452]. Circulation [3223]. Circumcision [577, 2679]. Citations [2394, 2527, 3625, 1896]. Cities [1143, 1821, 3773, 2690, 296, 3472]. Citizens [3187, 2845]. Citizenship [222]. City [1487, 1604, 2446, 1942, 550, 344, 2832, 2385, 2844, 1132, 654, 3596, 2398, 640]. Civic [3336]. Civil [1309, 943, 182, 2081, 3283, 695, 2825, 273]. Civilian [84]. Civilisation [2027, 1689]. Civilisations [3159]. Civilization [2883, 3455, 2883, 140, 1172]. Civilizational [570]. Civilizations [3559, 261, 3045, 3221]. Civilized [1201]. Civilizing [3132, 2626, 3182]. Clagett [2588, 465]. Clairaut [683]. Claire [626, 2670, 883]. Clancey [2836]. Clark [2172, 2345, 796, 680, 2055]. Clarke [41, 204, 87]. Class [1197, 1759, 1108, 785]. Classic [2168, 3307]. Classical [2728, 3144, 977, 3481, 53, 529, 3724, 2459, 3512]. Classics [1890]. Classification [781]. Classifications [1311]. Classifying [167]. Classique [3739, 2822, 582, 2896]. Classroom [352, 1750]. Claude [2532, 2086, 1139, 2675]. Claudia [346, 3679, 121, 3457, 2832]. Claudine [142]. Claudio [779]. Claus [617]. Clausberg [3500]. Clavis [1489]. Clavius [2971]. Clay [375]. Clayton [858]. clé [2971]. Clean [345]. Clearance [2216]. Clearcutting [1036]. Cleavage [735]. Clements [1129]. Clericuzio [3728, 879]. Clerk [169]. Cleveland [1083]. Cliff [296]. Clifford [1851, 345, 2132, 821, 435, 3133]. Climate [1180, 970, 850, 1297, 3255, 3399, 248, 3061]. Climates [348]. Clinic [3606]. clinica [165]. Clinical [2704, 620, 237, 3554, 2494]. clinique [3321]. Clio [2875, 2808]. Clive [2149]. Clock [1499, 1139]. Clocks [2146, 878, 2496, 3673]. Clonal [555]. Cloning [2300]. Close [3018]. Closed [357, 2992]. Clouds [26]. Clow [1123]. Club [547]. Clues [972]. Co [605, 3359, 2459]. Co-Creation [2459]. Co-Evolution [605]. Co-Production [3359]. Coal [3072]. Coates [438]. Cobb [1808]. Cobbett [3062]. Cochrane [817]. Code [1273, 2848, 2926]. Codex [3485]. Códice [1820]. Coding [2039]. Cody [2668]. Coerced [3302]. Coercion [3435]. Cofounder [1527]. Cogito [3153]. Cognition [615, 871]. Cognitive [1362, 3376, 3356, 1298, 1661]. Cografya [1370]. Cohen [203, 113, 3449, 203, 113, 142]. Cohesion [1672]. Cohn [1747]. Coil [3637]. Coin [2360]. Coining [2867]. Cold [3344, 1021, 2698, 2097, 2791, 1737, 1756, 82, 2083, 3607, 1871, 3431, 3272, 652, 77, 76, 1032, 544, 3352, 532, 1750, 2582, 3535, 3426, 1207]. Coldest [3339, 1251]. Cole [1023, 2295]. Colección [2133]. Colfax [2835]. Colgrove [2924]. Colin [1950]. Collaboration [3524]. Collaborative [2445]. Collected [301, 2362, 1133, 934, 3465, 3410, 1872, 2751, 3497, 75, 68, 3028, 1133, 1872, 2213]. Collecting [3151, 3744, 510, 563]. Collection [1995, 1600, 1166, 2315, 1057, 1196, 2016]. Collections [98, 232, 327, 423, 572, 670, 748, 2944, 2051, 2810, 1591]. Collective [2110, 680]. Collectors [2998, 3764, 1631, 1679]. College [2945, 3271, 128, 2752, 588, 1605, 433]. Colleges [2907]. Colliding [1034]. Collier [890, 216]. Collins [2439, 2788, 2155]. Collip [2442]. Colloid [3501]. colloque [361]. Colón [599]. Colonel [51, 1403, 1057]. Colonial [45, 1082, 2237, 3261, 57, 3502, 1947, 3655, 1797, 2740, 2626, 2908, 1595, 344, 1729, 2677, 2540, 1096, 343, 1795, 714, 2610, 3334, 3663, 132]. Colonialism [1928, 3349]. Colonienses [2041]. colonisation [183]. Colony [2909]. Color [1271, 457, 3491, 2995, 2026, 2020, 457]. Colorado [1939]. Colored [2957]. colorimetría [2995]. Colors [187]. Colour [1742]. Colours [1597]. Columbia [2034]. Columbus [3565]. com [2729]. Come [1398]. Comentarios [146]. Comets [145]. Comfort [403]. Coming [742]. Comitato [2842]. Command [1382]. Commemorate [2745]. Commemorative [36, 680]. Commentaire [2879, 2879, 2879]. commentaires [2822]. Commentaries [1542]. commentarii [616]. Commentariorum [1489]. Commentarius [1103]. Commentary [2697, 1364, 3041, 3512, 3320, 1583, 1690]. Comments [2606]. Commerce [2610, 3030, 2342, 2539, 2782, 2092, 1481]. Commercial [2714]. commerciale [2408]. Commercialization [676, 3665]. Commission [2478].

Commissioners [299]. Committee [2164, 544]. Commodity [901]. Common [1636, 1019, 674]. Commoner [3270]. Commonwealth [3081]. Communicate [2784]. Communicating [1280, 1974, 341]. Communication [1440, 2341, 1027, 3222, 1616]. Communicative [2554]. communicazione [3161]. Communist [1202]. Communities [179]. Community [2431, 3235, 3283, 279, 549, 628, 547]. Companies [2559, 1401, 305]. Companion [1987, 3109, 3785, 2031, 3589, 358, 3722, 1708, 113]. Company [2782, 3247]. Comparative [1180, 2953, 517, 1404, 933]. Comparison [2521]. Compass [1498]. Compendium [2945, 2243, 3650]. Compendy [2408]. Compensation [1185]. Competence [2307]. Competing [1556]. Competition [1951, 333]. Complaint [3476]. Complementarity [301]. Complete [3312]. complètes [1228]. compositio [1222]. Composition [2269]. Compound [874]. Compounding [2415]. Compounds [857]. Comprendre [1400]. Compressed [294]. Computational [3640]. Computer [2307, 205, 406, 1042, 203, 204, 524, 2782, 2438, 203, 2776, 685]. Computers [1029, 3344, 1043, 3428, 2782]. Computing [2568, 1677, 1966, 2187, 2563, 1628, 1441]. Comte [518, 2452]. comunidad [3174]. comunità [2267, 1381]. Conan [115]. Concealment [1323]. Conceived [3096]. Conceiving [3263]. Concentration [2186]. Concept [1235, 2600, 2410, 2746, 229, 3545, 2534, 2192]. Conception [1660, 2668, 799]. Conceptions [3046, 3300, 865, 392, 518]. Concepts [1059, 3106, 418, 845, 2815, 681, 339, 2261, 3136, 3783, 1669]. Conceptual [2705, 228, 1537]. Concerns [40]. Conchis [473]. Concise [1809]. Concrete [2434]. Condit [1880]. Conditions [1147, 2891]. Condon [2234]. Conevery [1844]. Conference [1546]. Conferences [1832]. Confessions [1227]. Confinement [736, 1997]. Conflict [1405, 783, 2461, 1881, 3401, 2356, 593]. Conflicts [2048]. conflit [2738]. Confluence [2554]. Conforti [3728]. Confrontation [2612]. Confronting [3431, 1025]. Confronts [3250, 1772]. Confucian [854]. Confutation [1583]. cong [3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277]. Congo [1947]. Congress [643, 7]. Conica [3647]. Conjecture [2015]. Conley [149]. Conn [3059, 178]. connaissance [2921]. connaissances [1552]. Connecticut [445]. Connecting [1776]. Connection [2126]. Connectionism [227]. Conner [3133]. Connor [1355, 1575]. conquérants [2915]. Conquest [1126, 1207, 2920, 1344, 698]. Conrad [3736]. Conscience [3429]. Conscientious [2319]. Consciousness [3059, 780]. Consensus [93]. Consent [2926]. Consequences [1423, 850, 307, 781, 3194]. Conservation [1420, 1129, 1550, 723, 3525]. Conservationist [1565]. conservée [2277]. Conserving [2929, 2350]. Considerations [2451, 2454]. Conspirator [1745]. Constantin [1034]. Constantine [2421]. Constants [606]. Constellating [31]. Constitution [2632, 613, 3195, 1350, 2820]. Constitutions [348]. construcción [3174]. Constructing [14, 409, 252, 927, 401, 514]. Construction [1061, 413, 173, 1775, 1150, 3223, 3020, 2492, 2410].

Constructivism [2100]. Consuegra [71]. Consultations [620]. Consumerism [1129]. Consumers [2371]. Consumption [293]. Contact [193, 1085]. Contagion [2913, 736]. Contagious [2377]. Contardi [2044]. contemporanea [3319]. Contemporaries [3654]. Contemporary [2536, 1406, 3561, 460, 3710, 2723, 3094, 333, 1885, 3080, 3531, 1039, 3439]. Contempt [3302]. Contend [307]. Contents [1331, 1790, 2228, 2514, 2794, 3114, 3445, 3788, 1697]. Contest [785]. Contested [2318, 3760]. contesto [1517]. Context [3647, 119, 3325, 66, 2402, 678, 3490, 2003, 3131, 3541, 143, 3555, 1915, 2970, 3082, 1384, 2541, 1100, 1587]. Contexts [2109, 1138, 1661, 1224]. Continental [1413, 1034]. Continents [79, 1034]. contingentia [144]. Continuity [2951, 355]. Continuous [2261]. Contraceptives [219]. Contracting [693]. Contradictions [460]. Contraption [2664]. Contraries [1405]. Contre [613]. contribución [449]. Contribution [1316, 2417]. Contributions [2939, 631, 2557, 1404, 770, 2355, 399, 3158, 2076]. Contributors [952, 1066, 1182, 1329, 1447, 1555, 1646, 1788, 1891, 2000, 2122, 2226, 2316, 2387, 2455, 2512, 2792, 2579, 2644, 2710, 2863, 2948, 3016, 3112, 3202, 3279, 3362, 3443, 3533, 3616, 3691, 3786]. Contro [1685]. Control [1025, 402, 2164, 1423, 1677, 3422, 1942, 3119, 198, 2168, 3631, 1171, 223, 1270]. Controlling [736, 662, 854, 88, 408]. Controls [1616]. Controriforma [2411, 3395]. controverse [2257]. Controverses [3361]. controvérsia [1619]. Controversies [1068, 679, 3654, 237]. Controversy [1583, 1148, 3001, 2746, 735, 3694, 1550, 281, 1894, 3425, 1648, 560, 3503]. Convent [710]. Convention [3095]. Converging [1034]. Conversations [1870, 1417, 611]. Conway [2770]. Conwy [3760]. Cook [359]. Cooke [1938]. Cookson [1950]. Cookstove [1130]. Cooley [2735]. Cooper [2159, 3049, 2975]. Cooperation [2783]. Copenhagen [2904]. Copernican [2536, 1095]. Copernicana [2607]. Copernicanischen [2406]. Copernicans [2273]. Copernicus [2967, 1822, 2607, 2536, 3379, 2476, 1693]. Copernicus-Biographien [2607]. Copies [1697]. Coping [975]. Copper [914]. Copy [13]. coraje [2174]. cordée [969]. Cordis [1826, 1103]. Core [871]. Coriell [1626, 1626]. Corinna [2558, 2336]. Cornelia [1414]. Cornelius [1697]. Corner [1620]. Corning [1157, 2369]. Cornish [984]. Coronelli [495]. Corones [355]. Corporate [431]. Corporation [3077, 1157]. corporis [2040]. corps [3635, 3568, 2549]. Corpse [1040]. Corpuscles [879]. Corpuscular [1366]. Correia [3327]. Correlations [1567]. Correspondence [959, 2277, 3241, 370, 2070, 3410, 1872, 805, 78, 2731, 68, 3169, 745, 491, 41]. Corrupt [3610]. Corruzione [2270]. Corsi [3060]. cose [1694]. Cosmic [1873, 3009, 1033, 1476, 2801, 2848, 2141, 3632]. cosmique [2896, 582]. Cosmo [493]. Cosmogram [2802]. Cosmograms [2801]. cosmologia [146, 3395]. cosmologico [39]. cosmologie [2824, 469, 3615]. Cosmology [2969, 475, 1095, 3504, 3547, 3386, 2621, 3009, 3300, 3557]. Cosmopolitanism [2830, 3456]. Cosmos [268, 3300, 2253, 478, 3230, 582, 1206, 1215, 3557]. Cost [2567]. Costa [1420, 2922, 2625, 2351]. costas [2263]. Cotte [1610]. Cotter [2367]. Cottingham [154]. Cottrell [1941]. Couch [2094]. Coulston [2484]. Count [2490]. Countenance [871]. Countercultural [3193]. Counterfactuals [3288]. Counting [242]. Countries [2811]. Country [2735, 1844, 2371, 2686]. Counts [2596]. County [2835]. Couple [2384]. Couprie [2402]. cour [701]. Cournot [502]. Cours [3361]. Court [2435, 2798, 3499, 3775]. Courts [2869, 914]. Cowan [251]. Cowboys [3205]. Cowen [2873]. Coworkers [2362]. Coxeter [3087]. Craddock [608, 640]. Craft [1281, 2369, 1586]. Crafting [3554]. Craig [546, 2813, 1932]. Craige [1141]. Craik [3668]. Crary [521]. Crasta [998]. Crater [1213]. Cratylus [1670]. Crawford [2472, 986]. creación [2430]. Creager [1396]. Crease [1051, 384]. Create [2167]. Created [1909]. Creating [3429, 1256, 1854, 2168, 3783, 2557, 76]. Creation [2852, 2078, 1397, 2691, 1648, 475, 653, 1846, 1543, 373, 893, 281, 3491, 3677, 2459, 2071, 1000]. Creation/Evolution [281]. Creationism [3224, 3712, 3084, 1970]. Creationists [3084]. Creative [3535, 137, 543]. Creativity [2332, 1730, 78, 689]. Creator [3090]. Creatures [2998, 2885, 3559]. Credibility [11]. Credit [2826]. Creed [715]. Creese [278, 2355, 278]. Crelinsten [2902]. Cremer [1385]. Crenner [2548]. Crettaz [161]. Crick [1339]. Crilly [2989]. Crime [2938, 2179]. Crimes [2926]. Criminal [1023, 2179, 860, 2498]. Criminology [860]. Crippled [1047]. Crises [3598]. Crisis [3019, 3769, 3759, 2702, 2322, 2481, 3085, 1958, 2145, 1015]. Crisóstomo [3620]. Crist [412]. Cristina [3666, 2222]. Critchlow [1423]. Critic [2222]. critica [1238]. Critical [378, 117, 1583, 507, 2242, 3435, 2028, 1364, 3612, 3041, 1542, 3597, 1771]. Criticism [2865]. critico [1701]. Critics [186]. critique [2879, 160, 3712]. Crocker [2483]. Croesu [608]. Croissant [2562]. Croker [1391]. Crone [457]. Crookes [3665]. Croone [795]. Cropper [1079]. Crosbie [717]. Crosby [2024]. Crosland [2976]. Cross [1369, 2868, 143]. Cross-Cultural [1369]. Cross-Examination [2868]. Crosscurrents [508]. Crossing [1038, 1038, 1152]. Crossroads [3303]. Crotty [1418]. Crouch [1137]. Crow [314]. Crown [3077]. Crowther [2678, 3252]. Crowther-Heyck [2678]. Crucible [3339]. Crusade [236]. Crusader [1859]. Crystal [2018]. crystalline [3200]. Crystallization [2206]. Crystallization-Differentiation [2206]. Cuddington [2789]. Cueto [60]. cuisine [33]. culinaire [2823]. Culmann [719, 719]. Culmann-Texten [719]. Culpa [2174]. culpabilizacion [255]. Cult [3079, 3172]. Cultivation [2981, 2053, 2161]. cultura [2411, 324, 1483]. Cultural [2, 427, 1325, 2517, 2796, 3116, 3447, 1024, 2195, 3069, 1299, 3328, 3365, 989, 3716, 717, 2836, 2521, 2242, 882, 2963, 2127, 2412, 166, 695, 127, 2641, 3584, 2964, 3003, 116, 1369, 2597, 281, 2785, 81, 1879, 1422, 2585, 662, 230, 830, 1333, 1792, 2230, 11, 392]. Culture [3601, 1316, 1321, 945, 2118, 1201, 3324, 1665, 3222, 1931, 3646, 565, 3175, 2288, 3303, 3072, 3338, 2771, 510, 1336, 106, 134, 655, 539, 3520, 2473, 3226, 3748, 1692, 3337, 1365, 2505, 3001, 1027, 1615, 901, 3765, 3299, 3131, 2699, 3397, 1112, 3759, 3501, 3592, 2286, 3665, 3480, 524, 3744, 1427, 405, 3659, 13, 3666, 1712, 2523, 2154, 988, 3258, 209, 2972, 3772, 2033, 1795, 701, 613, 1868, 1044, 3070, 460, 1867, 521, 3079, 1128, 3681, 2422, 1108, 2335, 1060, 1062, 1785, 1444, 1632, 1467, 202, 1760, 3341, 3036, 3339, 1579, 3590, 3532, 1085, 3266]. culturelle [3495]. Cultures [1357, 446, 3222, 3454, 2321, 3567, 195, 2740, 2092, 2428, 967, 846, 1191, 2492, 2146, 1034, 1171]. Culure [1853]. Cumberland [492]. Cuomo [862, 3480, 978]. cúpulas [2662]. Cure [3535, 1756, 1948, 237, 1711, 2193, 3435, 693, 597]. Cures [1883]. Curie [2915, 1111, 64, 798]. Curing [577]. curiosità [3394]. Curiosities [3040]. Curiosity [881, 575, 2740, 2248, 882, 3218]. Curious [3044, 3559, 2991, 1810, 3514]. Current [2, 427, 2517, 2796, 3116, 3447, 2680, 3417, 830, 1333, 1792, 2230, 1193]. Cursing [1076]. Curt [2764]. Curtin [628]. Curtis [888, 123, 89, 1416, 1347]. Curve [1418]. Cusanus [1371]. Cusanus-Forschung [1371]. Cushing [2750, 418, 2063]. Customs [396]. Cut [577]. Cutler [2045]. Cutter [3085]. Cutting [411]. Cyanotype [1598]. Cybernetics [323, 1677, 1769]. Cyberspace [602, 2187]. Cyberspeak [1769]. Cyborg [1753]. Cycles [345, 140]. Cymbalum [1238]. Cynicism [2790]. Cynthia [884, 195, 2214]. Cyrano [2258]. Cyril [2775]. Cystic [3198].

D [1309, 3324, 632, 3507, 558, 375, 2272, 643, 30, 690, 1137, 2700, 3133, 365, 1261, 3005, 454, 83, 1921, 2821, 2188, 2858, 203, 3434, 2486, 1882, 2441, 2497, 813, 1823, 3668, 1054, 2581, 2357, 94, 2315, 2633, 343, 270, 211, 711, 754, 1851, 1815, 897]. D-Day [1261]. D. [1114, 2684, 875, 1039]. Dadswell [2162]. Dag [363]. Dagli [3735]. d’agronomes [1084]. Dahl [2175, 192, 383]. Dahlemer [1643]. dai [3561, 3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277]. Daiber [264]. daishu [3306]. Dal [1465, 2022, 469]. d’al-Kindî. [469]. Dale [3005, 1078]. d’Alexandre [2202]. Dalibor [2332]. Dalitz [792]. Dalla [324]. Dallas [2882]. Dall’onnipotenza [25]. Dalrymple [66]. Dalton [2309]. Daly [2704]. Dam [564]. Dame [1715]. Damerow [230]. Dams [2803]. Dan [656, 3262, 3200]. Dances [3749]. Dangerous [1908, 1709, 2816, 2196]. Danian [2683]. Daniel [489, 1923, 283, 1157, 519, 2184, 2855, 40, 656, 2751, 3523, 410, 853, 2092, 3075, 2689, 2290, 1149, 1866]. Daniela [3605]. Daniels [656]. Dann [661, 187]. dans [2534, 3157, 2620, 3248, 2277, 2619, 2426, 3482, 1225, 3740]. Dansgaard [2420]. Danson [2346]. Dante [602, 984, 869, 3615]. D’Antonio [3315]. dao [3200]. Daoism [3484]. Daqing [3517]. Darás [586]. Darby [2679]. d’archives [1610]. Dardo [2472]. Darian [2223, 2359]. Dario [3735]. d’Aristote [2879]. d’Aristotile [35]. Dark [2502, 2911, 2968, 1833]. Darker [2193]. Darkness [1539]. Darling [3407]. Darlington [2775]. DARPA [2568]. Darrel [2819]. Darrigol [2814, 1596]. Dartmouth [2557]. d’Arturo [382]. Darwin [3164, 3749, 3250, 2698, 1848, 2449, 378, 3295, 3591, 2754, 2435, 3745, 2065, 724, 805, 3558, 3669, 2501, 2642, 1074, 930, 1005, 3425, 3169, 3286, 3622, 1336, 2984, 1119, 2689, 2648, 3624, 1601, 3698, 1971, 2967, 847, 1720, 3169, 1987, 3600, 1719, 3183, 172, 2674, 3669, 3409]. Darwinian [935, 689, 3504, 1395, 3097]. Darwinism [3670, 1523, 1268, 722, 1390, 3012, 2614, 3251, 516]. Darwinismus [2553, 1317]. Darwinist [1599]. Darwins [3579, 3743, 3766]. Darwinsche [2553]. Dasgupta [645]. Dasmann [1565]. Dassow [1931]. Daston [782]. Dating [911, 456]. Dauben [1167]. Daughter [647, 1387, 1601]. Daum [337]. d’autourserie [472]. Davenport [3154, 3516]. David [1061, 280, 621, 191, 1362, 2883, 161, 1165, 404, 301, 545, 3144, 366, 723, 727, 539, 50, 313, 757, 263, 374, 603, 318, 319, 421, 699, 1418, 93, 204, 3569, 78, 585, 2751, 3555, 979, 1724, 3523, 206, 741, 587, 3297, 2529, 544, 688, 3074, 711, 685, 342, 462, 1025, 2987, 2898, 2130, 2912, 1888, 2097, 2873, 3083, 1962, 1010, 2141, 2723, 2664, 2184, 2361, 3508, 2565, 2854, 2669, 2855, 2990, 1159, 3001, 2345, 1033, 933, 1562, 1112, 923, 2140, 2696, 1351, 2835, 3065, 2966, 1611, 1619, 2961, 1595, 2687, 919, 2160, 2071, 2153, 2450, 1424, 1919]. Davis [3158, 1896, 374, 53, 3124, 192, 2573, 1157, 2448, 2206, 1877]. Davy [991]. Dawkins [3032]. Dawn [3683, 3596, 726, 3770, 3579]. Day [3055, 892, 1261, 231, 709, 84, 139]. Days [1762, 995]. Dead [1122, 1253, 376, 594]. Deadly [3426, 1961, 1942, 2660, 2047]. Deaf [1282]. Deahl [174]. Deal [316]. Dean [727, 805, 689, 2212]. Deanna [594]. Dear [3475, 987]. Death [2945, 254, 3163, 1743, 737, 2653, 2651, 1040, 2650, 86, 2833, 738]. debat [16]. Debate [1200, 2700, 2904, 1233, 3676, 1680, 3185, 484, 1288, 3004, 39, 193, 1038, 2634]. Debates [1523, 554, 1880, 401, 848, 1292]. Debating [3586, 1074]. Débats [2091, 761, 3740]. DeBevoise [357]. Deborah [611, 2771, 3340, 133, 371, 1881, 2181, 1622]. Debs [3095]. Debus [148]. Decade [1767, 1430]. Decades [3098, 3449]. décalages [3639]. December [239, 3340, 1697, 3410, 371, 240, 958, 3371]. Deception [926]. Deceptions [2714]. Dechend [1342]. decimal [498]. Deciphering [296]. Declaration [3685]. Declercq [866]. Decline [1295, 145, 3501, 3598, 376]. deCODE [3767]. d’Éduard [2620]. Deduction [1362]. deduktiver [2347]. Dee [3393]. Deep [2761, 505, 1766, 6, 1539, 477, 3384, 444, 3076]. Deep-Sea [1539]. Deeper [2449]. Defects [1169, 788]. Defenders [1200]. Defending [2790]. Defense [926, 3739, 3272]. Defibrillator [1155]. Deficiency [379]. Defining [303, 2622, 2700]. Definition [3347]. Definitive [1041]. Deforesting [2145]. Deformations [2330]. Deformed [3046]. Degaev [1838]. degli [58]. DeGrandpre [3079]. Degree [1250]. Degrees [3065]. DeGregori [2634]. deiatel’nosti [1354]. Deirdre [1152]. Dekanatsbuch [1815]. del [2173, 768, 3394, 884, 2264, 3056, 2270, 58, 25, 255, 1483, 146, 1820, 498, 102, 625, 3174, 60, 610, 2263, 2252, 493, 732, 35, 18]. Delano [443]. Delano-Smith [443]. Delaporte [298]. Delbourgo [2978]. Delbrück [2458]. Delft [2486]. Delgarno [2608]. della [691, 3730, 324, 25, 3326, 3395, 1465, 226, 998, 2504, 2411, 2401, 610, 777, 493, 2977, 3615]. Delta [2569]. Deluge [2341]. Delusion [40, 3032]. Delusions [3523]. Democracy [2318, 2083, 957, 379, 222]. Democratic [225]. Democritus [235]. démon [2343]. Demonization [2679]. Demonstration [2575, 1509, 568, 3390]. Demonstrativo [3390]. Denery [2882]. d’enginyeria [2430]. Denis [2905, 1356]. Denise [242, 29]. Denmark [3760]. Dennis [1415, 2437, 865, 878]. Denominative [2867]. Dental [1576]. Dentist [1007]. Dentiste [486]. Dentistry [1007]. Denton [1629]. Deomocracy [2510]. Departing [2302]. Department [315, 2813, 2930, 2923, 937]. Departure [2932]. DePauli [1349]. Depicting [2153]. Depletion [1180]. Depreciated [2124]. Depression [713, 1621]. Derangement [2313]. Deregulation [537]. Derek [1896, 729, 1850]. Desaguliers [1556]. desarrollo [768]. Descarte [2410, 2416, 154]. Descartes [2981, 2733, 155, 227, 3781, 2737, 491, 1098, 2489, 1904, 878, 2973, 490, 1702, 1492, 157, 491, 3153, 3154]. Describing [3391]. Descripcion [2263]. Description [1697, 2420, 3396]. Descriptive [2016, 2414]. desde [2742]. Desert [696]. Design [3132, 268, 1971, 3370, 3469, 524, 3619, 3712, 314, 3425, 3084, 3295]. Designing [3413, 2673, 1933]. Designs [3110, 1421]. Desire [2256, 2958]. Desires [1480]. Desor [2620]. d’Espagnat [3011, 228]. Despoix [2887]. Desrosières [503]. Destined [2274, 390]. Destinies [408]. Destiny [2053]. Detailed [1747]. Detection [3328]. Detective [2068]. Detectors [3294]. Determinants [305]. Determination [2951]. Detlef [799]. Detroit [893]. deutsche [337]. Deutschen [2368, 1402, 3193, 2831]. deutscher [2840, 525, 34]. Deutschland [2308, 199, 1402, 347, 3330, 2553, 126, 3260, 2818]. deutschsprachigen [171]. deuxième [1735]. Devant [2341]. Devastation [2440]. Develop [317]. Development [2844, 1747, 734, 3349, 2546, 1404, 2572, 21, 2505, 2356, 442, 3497, 576, 2442, 2925, 543, 3470, 338, 2136, 389, 1337, 17, 1980]. Developmental [3507]. Developments [631]. Deventer [2016]. Deviance [640, 2302]. Device [2582]. Devices [3110, 1480]. Devil [6]. Devo [3507, 2630]. DeVorkin [191, 727]. Dewey [2309]. DeWitt [757, 1591]. Dewsbury [2853]. DGGTB [1178]. d’Herelle [197]. d’Histoire [3157, 3495, 3018, 3355, 3321]. d’Història [2430]. d’Humanisme [3311]. Diablo [3525]. Diagrams [2565]. Dialectics [208]. Diálogo [479]. Dialogue [1917, 530, 3045]. Diamond [388]. Diamonds [249]. Diana [728, 3410, 2751, 2768, 2187, 1426, 1422]. diao [3186]. Diaries [1248]. Diary [2279, 2688, 212, 922]. Dias [2549]. Diaspora [2058]. dibattito [2977]. Dibner [3507, 2994]. Dick [2985, 193, 20, 2176, 2925, 2200]. Dickens [2068, 900]. Dickenson [247]. Dictation [1567]. Dictatorship [3519, 3337]. Dictionaries [2422]. Dictionary [938, 1814, 3726, 1431, 2643, 231, 2817, 571, 3215]. Dictionnaire [364]. Did [1650, 362]. Diderot [3731]. Didi [2178]. Didi-Huberman [2178]. Didier [1718, 3311]. Dienel [636]. Diesel [396]. Diet [1580]. Dieter [769, 3035].

Diethard [609]. Dietrich [790, 34, 2380]. Dieu [2896]. Difference [3346, 3207, 1541, 1451]. Differences [3654, 889]. Different [3710, 1351]. Differentialgeometrie [483]. Differentiation [2206]. difficile [1381]. Diffusion [772, 641, 613]. dificil [200]. Digby [1054]. Digital [2559, 2808, 3611]. Dijck [2785]. Dikes [2803]. Dilemma [858, 2907]. Dilemmas [2309]. diluvio [18]. Dimension [2241, 3605, 3501]. Dimensions [1957, 2969, 1249, 1661]. Dimitri [1365]. Dinosaur [1868, 116, 560, 1262]. Dinosaurs [847, 1538, 445, 1213, 1522]. d’interprétation [2706]. Diodati [3157]. Dioskurides [34]. Dioxide [3255]. Dipak [2762]. Diphtheria [1942]. Direction [585]. Directions [2953]. Directorate [2217]. Dirk [525, 2402, 1072]. Dirt [3221]. d’Isaac [2979]. Disability [1047, 1085, 1573]. Disappearance [2497]. Disappointment [3468]. Disaster [1305, 3073, 1430]. Disasters [1035]. Disciples [2698]. Discipline [1327, 2834, 937, 2551, 2451]. Disciplines [108, 2952, 3581, 612, 2220, 3710]. Disciplining [87]. Discontents [2707, 253]. Discord [1630]. Discorsi [2099]. discours [2733]. Discourse [1833, 3055, 678, 2913, 2283, 2735]. Discourses [536]. Discover [1621, 535]. Discovered [2045, 847, 993]. Discoveries [1501, 2141, 1842, 2065, 2345, 973, 189, 1496, 1212]. Discovering [2852, 3001]. Discovers [1286]. Discovery [2761, 3180, 2784, 1522, 2224, 675, 699, 192, 2658, 3592, 1589, 214, 1384, 2786, 1339, 2486, 997, 1384]. Discrete [2261]. Discrimination [7]. Discursive [201]. discurso [2351]. Disease [3755, 946, 1577, 905, 1278, 2385, 554, 1241, 2815, 2268, 2899, 3588, 956, 1123, 2214, 2660, 86, 640, 296, 3176, 2047, 1687, 856, 597, 2658, 3347, 2728, 1126, 217, 2105, 1814, 3198, 1729]. Diseases [2086, 218, 3517]. Disguise [1874]. Diskontinuitäten [3761]. Diskussii [1267]. Disorder [3608]. Disorders [2171]. Dispatches [186]. Dispersion [2565, 807]. Display [1945, 3351, 509]. Disposal [1768]. Dispositifs [2887]. Dispute [1892, 2784]. Disputes [455]. Disrupting [3522]. Dissection [3066, 789, 3149]. Disseminating [722]. Dissent [2121, 85, 3619]. Dissenter [482, 3258]. Dissertation [1250]. Dissertations [42]. Distance [585]. Distilling [2983]. Distinguished [2274]. Distorted [3025]. Distortion [3779, 3089]. Distributed [3690]. distribution [3248]. District [2966]. Distrust [675]. Disturbed [3286]. Disziplin [1017]. Disziplinbildung [1385]. Disziplinen [2211]. d’Italia [777]. Dittrich [500]. Divagazioni [3138]. Dive [390]. Divergence [1486]. Diversis [3248]. Diversity [555]. Divide [2771]. Divided [2332, 3609, 1272, 3401]. Divides [2377]. Dividing [1482]. divin [2337]. Divination [2335, 3482, 461]. Divine [871, 2737]. Diving [294]. Divining [432]. divulgazione [2282]. Dixon [2346, 3151]. Djerassi [1153]. Dmitrii [2357]. DNA [1148, 414, 1623, 676, 3305, 3558]. Do [2197, 152, 2279, 3327]. Doborah [2941]. Dobzhansky [2023]. Doctor [3777, 1250, 322, 1696, 2272, 2828]. Doctoring [1246]. Doctors [1009, 2728, 1154, 2193, 2625, 1937, 2152]. Doctrine [871]. Doctrines [3349]. Documentary [213, 3423]. Documents [1309, 943, 1738, 62, 2904, 84, 1310, 1726]. Dodge [51, 51, 1403]. Dodgen [854]. Dodgson [2057]. Dodonaeus [1377]. Doers [1137]. Does [1971, 3208, 3207]. Dohner [2225]. Doing [1575]. Doktor [556]. D’Olivier [1084]. Dollars [676]. Dolly [3033]. Dolman [2086]. Dolomites [2281]. Domain [2867]. Domaradskij [2087]. Domestic [1130]. Domesticating [3748]. Dominance [3132]. Domini [866]. Dominion [280]. Domurat [297]. Don [1401, 3638, 739]. Donahue [708]. Donald [1061, 108, 774, 803, 3087, 186, 1846, 694, 1423, 1389, 1148, 914, 2853]. Dong [653]. Dong-sung [653]. Donna [932, 932]. Dooley [731]. Doolittle [1012]. Doom [3055]. Doomsday [3434, 1289]. Doors [2992]. Dora [884, 1400]. Dorado [1114]. d’origine [3388]. Doris [1089, 65]. Dorland [2821]. Dorman [3340, 371]. Dorn [131, 437]. Dorofeeva [3634]. Dorofeeva-Lichtmann [3634]. Dorothy [1653, 386]. Dörries [2904]. Dorrit [1874]. Dose [928]. dottrina [2099]. Double [1204, 3204, 1339]. Doug [2055]. Douglas [1175, 175, 1747, 2210, 660, 2032, 972, 1620, 1126, 2152]. Dow [2055]. Dowe [2642]. Down [1451]. Downlink [1766]. Downs [1030]. Doyle [281, 2068]. Dr [884, 2548, 2503, 40, 3434, 3432, 884, 3166]. Dr. [731, 1401, 2788]. Draelants [3483]. Draft [3561, 3306, 2420]. Dragon [847, 854]. drahtlosen [814]. Drama [1053]. Dramatize [971]. Draper [1870]. Drawing [2346, 2565]. Drawings [2420]. Drawn [247]. Drayton [1101]. Dream [2098, 2433, 3516, 2349, 2656, 3434, 3198, 204, 3273, 1628, 3335, 546]. Dreamer [3510]. Dreamers [1137]. Dreaming [1751]. Dreams [1753, 2770, 3489, 535, 3601]. Dreger [297]. dreizehnten [474]. Drift [1034, 3076, 1413]. Dritten [2117]. Driver [2736, 846]. Dropping [1295]. Dross [2616]. Drucke [95]. Drug [1025, 3079, 2860, 1304]. Drugoi [2499]. Drugs [3111, 3347, 2094]. Dryden [214]. DSIR [315]. Dubois [1005]. Dubow [3081, 776]. Ducal [353]. Duchamp [66]. Duchess [885]. Ducks [3536]. Ducos [614]. due [3141]. Duffin [2815, 451]. Dugatkin [2878]. Duke [885]. Dumbarton [1600]. Dumont [1084]. d’un [3157, 1006, 3639, 1350, 3494]. Dunar [1884]. Dunbar [3318]. Duncan [805, 1156, 2382]. d’une [613, 2278, 3355, 582]. Dunhill [788]. Dunkerque [498]. Dunkle [3715]. Dunlap [2058]. Dunmore [3652]. Dunn [1188]. Dunnill [904]. Duns [144, 982, 143]. DuPont [3274]. Dupree [3252]. Dupuit [369]. Durana [1175]. durch [3578, 3068]. Durchschaut [1284]. During [1747, 1370, 3344, 2614, 3753, 182, 31, 3261, 2741, 2174, 583, 2486, 376, 91, 3020]. Durkheim [2453]. durkheimienne [2921]. Dusan [2221]. Dusek [112]. Düsseldorf [2616]. Dutch [554, 1245, 3030, 1721, 2273, 3244, 2850, 1795]. Duverney [1647]. Dwayne [84]. Dwight [2613, 678]. Dyck [3171]. Dyer [1157]. Dyestuffs [134]. Dying [2757, 2446, 1723]. Dynamics [1741]. Dynamiste [2312]. Dynamo [1292]. Dynasty [3306]. Dyson [977, 763].

E. [2943, 2951, 1569]. Eagle [628]. Ear [3721]. Earl [958]. Earliest [2330]. Early [1315, 945, 1060, 1782, 1785, 1889, 2109, 1978, 1982, 1991, 3106, 3109, 2735, 3044, 794, 2004, 2885, 3234, 3046, 3325, 2667, 3662, 2048, 2827, 238, 2895, 108, 3029, 2869, 2610, 3304, 3574, 3265, 3052, 2052, 1366, 1363, 2978, 3475, 4, 976, 1404, 1232, 1241, 3163, 2548, 575, 882, 1095, 1229, 2893, 2341, 2670, 3401, 1697, 1365, 3368, 2410, 149, 699, 269, 3131, 235, 308, 3160, 2830, 65, 784, 2605, 2373, 3774, 767, 2650, 3650, 448, 3744, 652, 3497, 2256, 367, 2891, 433, 2249, 2342, 1594, 3570, 2745, 834, 3156, 2523, 3245, 709, 1096, 2261, 3547, 3049]. Early [857, 3062, 1288, 883, 487, 3700, 3658, 1097, 693, 2718, 2392, 1451, 2414, 2109]. Early-Modern [2052, 1097]. Earnest [4]. Earth [2697, 3496, 1081, 2500, 2244, 304, 3559, 683, 83, 2850, 1741, 2477, 2145, 2156, 2020, 3674, 3497, 2326, 3253, 760, 1413, 2271, 538, 1016, 422, 2972, 688, 972, 2966, 2773, 1824, 3034, 644, 911, 2045, 535, 2134, 1296, 3684]. Earthmoving [3273]. Earthquake [409, 2836]. Earthquakes [2089]. Earthrise [3684]. East [1787, 3247, 1194, 3752, 1193, 927, 653, 3351, 3519, 772, 1202, 1154, 979]. Eastern [1026, 3724]. Eating [2799]. Ebeling [3297]. Eberhard [707, 56]. Ebing [1120]. eccellent [875]. Eccentricity [2865, 959, 3744]. Eckert [1714, 2782]. Eckert-Mauchly [2782]. Eclectic [1605]. Eclipse [2614, 1363, 2287, 1338]. Ecofeminism [2652]. Ecological [2789]. Ecologies [2667, 3176]. Ecologist [1141, 1941]. Ecology [2208, 2789, 740, 907, 2755, 2651, 506, 2423, 2689, 331, 913, 633, 1834, 1040, 1391]. Economic [1727, 1069, 93, 260, 1039]. Economics [1753, 1972, 3360, 1569, 2917, 2756, 2037]. économie [2921, 502, 502]. Economies [2933]. économiques [3320]. Economist [2076]. Economy [2520, 3252, 3360, 2379, 2851, 1682]. Ecosystem [1141]. écrites [3739]. Écrits [490]. Écriture [2738, 2820]. Ecstatic [1823]. Eddington [1650, 3419, 1338]. Eddleman [1591]. Ede [3501]. Edelson [1349]. Eden [2440]. Edens [3773]. Edge [1784, 2943, 1307, 1305, 1562, 1795]. Edgerton [2854, 2990, 2296, 2296]. Edinburgh [2686, 2688, 3653]. Edison [720, 721, 1608, 3677]. Edited [3331, 1081, 3465, 2751]. Edition [1583, 1917, 1364, 3747, 3041, 1106, 2879, 3312, 3612, 2041]. Editor [2472, 1656, 1801, 1803, 1457, 2277, 2609, 1658, 1901, 1458, 840, 1898, 1459, 1899, 2301, 1563, 1900, 1456, 838, 2216, 2874, 2962, 2014, 2654, 2011, 3290, 3706, 105, 3630, 3544, 3449, 3289, 1189, 3704, 3291, 3372, 3461, 3214, 2719, 3464, 241, 3027, 3292, 2470, 334, 3459, 3462, 3373, 2590, 2591, 3629, 3628, 3127, 1073, 3128, 2468, 3213, 2013, 3460, 3463, 3705]. Editorial [2807, 3374, 2001, 2389]. Editors [2994, 2657, 2731, 1198, 3448]. Edizione [3735, 1238]. Edmond [359]. Edmund [1028]. Edoardo [2132, 526]. Edsall [1340]. Edson [351]. Eduard [850]. Education [1316, 1896, 1666, 3271, 3208, 1290, 1835, 2993, 2356, 1739, 2391, 391, 1796, 1678, 3306, 30, 606, 1751, 1750]. Educational [1739]. Edward [3158, 402, 1349, 404, 511, 378, 3379, 3315, 1166, 822, 652, 849, 2503, 2234, 2884, 990, 3015, 1548, 89]. Edwardian [285]. Edwin [374, 2346, 1747]. eens [906]. Eero [431]. eeuw [499]. Effect [2097, 3255, 2640]. Effects [1205]. Effervescence [42]. Efficiency [3631, 395]. Effland [1175]. Effluent [1143]. Egerton [1941]. Eggert [628]. Eglash [2562]. Ego [3452]. Egypt [3040, 3553, 577, 980, 2336, 3502]. Egyptian [465, 393, 1389]. Eibesfeldt [55]. Eibl [55]. Eibl-Eibesfeldt [55]. Eigen [2179]. Eight [2420]. Eighteenth [944, 1433, 3735, 3162, 271, 1509, 683, 2669, 1237, 2243, 123, 2668, 2982, 770, 1099, 2266, 1512, 1926, 1239, 2717, 3322, 3731, 1511, 3491, 2160, 1825, 272, 1246, 2539, 362, 2123, 1449]. Eighteenth-Century [3162, 271, 1509, 683, 2669, 1237, 2668, 2982, 1099, 2266, 1512, 1926, 1239, 3322, 1511, 3491, 1825, 1246, 2539, 1449]. Eighth [1365]. Eijk [2728]. Eileen [738, 785, 412]. einem [1319]. Einfuhrung [346, 3689]. Einschneidende [56]. Einstein [3136, 2683, 522, 2531, 3008, 3500, 3508, 2991, 3410, 1387, 2751, 68, 3173, 3512, 3070, 2126, 2204, 2902, 2127, 1869, 2496, 2751, 2096, 2992, 2940, 3410, 1596, 648, 3513, 3590, 3378, 302, 1387, 3264, 3470, 3028, 1740]. Eintritt [67]. Eisele [2129]. Eisenhart [210]. Eisenhower [544, 3185]. Eisenstaedt [2991]. eisernen [1985]. Eisner [2154, 2059]. Ekelund [369]. Ekkehard [1178]. Ekmeleddin [583]. el-Haj [2082]. Elder [863]. Eldredge [3224]. Eleanor [606]. Election [555]. Electric [2418, 1949, 537, 1841, 919, 2875, 2808, 215, 1781]. Electrical [1950, 2310]. Electricity [2978, 3748, 2658, 367, 3677, 2213, 538, 1781]. Electroconvulsive [3686]. Electrodynamics [1596]. Electromagnetic [112]. Electron [1065, 198, 192, 192]. Electronic [3090, 1055, 2247, 326]. Electronics [762]. Electrons [2749]. Electropolis [2910]. Electroshock [2190]. Elegance [951]. Elegant [1957]. Elektrische [3330]. Elektrizitat [54]. Elektrodynamik [3071]. Elektroenzephalographie [3058]. Element [2682]. Elementarlehre [2846]. Elemente [2194]. Elements [2545, 2269, 1384, 879, 1014, 2147, 2019, 3145]. Elena [529]. Elephant [477, 1511]. elettricità [2352]. Eleventh [3727]. Eleventh-Century [3727]. élèves [1310]. Elhanan [497]. Eli [687, 819]. Eliane [1825]. Elichirigoity [406]. Elie [1269, 3157]. Elin [821]. Elinor [2102]. Elio [38]. Eliot [1360]. Elisabeth [2472, 253, 2472, 2598]. Elitegehirnforschung [2530]. Elitism [225]. Elixir [3199]. Elizabeth [210, 519, 219, 1027, 442, 59, 2009, 802, 811, 1945, 1590, 1830, 2891, 2069]. Elizabethan [3233, 3042, 2604, 791, 2730]. elle [1593]. Ellen [1939, 2124]. Elliot [2784]. Elliott [26, 604, 680]. Ellis [1012]. Elman [2963, 2655, 695]. Eloge [3702, 2130, 435, 3627, 3703, 239, 2873, 837, 1799, 2324, 2240, 1071, 1560, 1340, 1798, 2129, 1454, 2588, 3457, 2464, 2466, 3458, 1559, 3026, 1188, 2395, 3124, 2467, 2007, 1653, 1342, 1800, 1562, 1561, 3543, 1654, 578, 1655, 2960, 1895, 2961, 1455, 2009, 240, 1652, 2008, 958, 1072, 3125, 1341, 2806, 2465, 3371, 2396, 2397, 754, 826]. Eloges [3626]. Eloquence [3758]. Elsa [556]. Elsasser [80]. Elusive [1666, 2349, 2280]. Elzinga [2992]. Emancipatory [2302]. Emanuel [998]. Embankment [1283]. Embedded [2950]. Emblematic [3046]. Embodied [1253]. Embrace [285]. Embryo [2630]. Embryological [2583]. Embryology [3507, 470, 2930]. Embryos [2300, 1530]. Emergence [3335, 378, 3029, 406, 1669, 2505, 3299, 2835, 2326, 289, 1795]. Emerson [268, 46, 3653, 1931]. Emigranten [525]. Emigration [642]. Emil [3254, 3178]. Emile [2315, 1558]. Emilie [3043, 772]. Emily [497, 616, 581, 855, 1615]. Eminent [2106, 3276]. Emmanuel [2312, 1904]. Emmart [616]. Emory [50]. Emotional [3716]. Emotions [3697, 3699, 3698]. Emperor [513, 3351]. Empire [2885, 3243, 3565, 1082, 2292, 2436, 2236, 1240, 2412, 2980, 3733, 1700, 1070, 1943, 2239, 2423, 3509, 3003, 2540, 3020, 43, 2287, 2556, 247, 49, 2736, 1821, 846, 1779, 3576, 3493]. Empires [2496, 2146]. Empirical [1181, 948, 2974, 2454]. Empiricism [2893, 3785, 1551]. Empiricist [3758]. Empiricists [2888]. Empirico [1685]. Empirismus [2347]. Emptoz [2088]. Enchantment [2286]. Encounter [2511, 2412]. Encounters [3328, 936, 3468]. Encyclopaedic [2422]. Encyclopedia [2225, 2689, 528, 3723, 422, 2472, 133, 1168, 1822, 669, 1431, 1926, 2709]. Encyclopédie [364, 2258]. End [3107, 2943, 2681, 1504, 611, 861, 2358, 1213, 1535, 63, 3427]. Endangered [471, 2225]. Ende [3227]. Endicott [652]. Ending [3503]. Endings [3180]. Endless [2725, 291]. Endocrinology [3282]. Enduring [3699, 1479]. Energy [1438, 717, 1143, 2519, 2139, 383, 528]. enfermedad [1399, 255]. enfermedades [141]. Engagement [959, 957, 600]. Engels [199]. Engendering [1169]. Engineer [1950, 3510, 2106]. Engineered [1909]. Engineering [3099, 216, 2350, 179, 1129, 1043, 2825, 201, 3519, 273, 1862, 2782, 3469, 1950, 294, 524, 3643]. Engineers [3344, 292, 854, 369]. Engines [890]. England [2121, 794, 3496, 3389, 2196, 1707, 3242, 1241, 488, 1692, 2348, 3490, 3741, 2606, 2266, 1698, 492, 3062, 487, 575, 3325, 2667, 2669, 1922, 2604, 715, 3651, 1594, 2261, 2539, 3163, 2356, 3160, 1723, 3570, 3156, 1246]. English [1177, 3044, 1165, 2058, 3247, 510, 2491, 1627, 989, 151, 1728, 1055, 1917, 3312, 3392, 1711, 1700, 2966, 3313, 2256, 2409, 1364, 3041, 3150, 270, 1326, 443]. English-Speaking [1165]. Englishwomen [49]. Engraver [3165]. Engs [345]. Engstrom [552, 2494]. Enigma [3272, 395]. Enigmas [3730]. enjeux [761]. Enlightened [2419, 796]. Enlightenment [2350, 2272, 2418, 3162, 2275, 2978, 1240, 1588, 713, 432, 3397, 3737, 887, 3497, 2422, 1932, 3249, 3653, 751, 366, 2741, 2821, 1812, 1099, 2658, 2884, 2611, 2271, 1439, 2109, 3399, 1709, 1379, 3218, 1712]. Enough [1466]. Enriching [1016]. Enrico [2252, 2519]. Enrique [146]. Ensaios [2561]. Ensayos [190]. enseignement [3639]. Ensenyat [3604]. Enstehung [3578]. Entanglement [3437]. Entdeckung [1944, 1826]. Enterprise [1786, 1988, 2622, 1096, 2442, 2290, 1082]. Enterprises [8]. Entertainment [2658]. Enthusiastical [40]. Entities [2749]. Entomological [3020]. entomologie [2492]. Entomology [740, 2520]. Entre [3054, 1736, 969, 1492, 2343, 2738, 3053, 3075, 479, 3320, 2444, 3482]. Entrepreneur [1401, 1002]. Entrepreneurial [1752]. Entropie [3177]. Entropy [2362]. Entstehung [3671, 199, 10]. Entstehungsgeschichte [188].

Entwicklung [2837]. Envelope [2090]. Environment [348, 2364, 2058, 844, 1834, 2665, 1283, 2637, 2165, 3176, 1143, 1019, 801, 2134, 2306]. Environmental [1180, 949, 950, 1064, 1177, 1984, 2429, 2364, 1716, 3380, 287, 1940, 3336, 696, 1958, 1549, 2441, 2423, 2569, 2034, 1292, 286, 331, 1879, 2690, 3020, 2931, 2773, 1577]. Environmentalism [3417, 3270]. Environmentalist [1141]. Enzymes [9]. Eolith [1648]. Ephemera [1578]. Ephraim [374, 2546]. Epic [2612, 1481]. epidemias [60]. Epidemic [1830]. Epidemics [2174, 296, 2377, 2811, 1026]. Epidemiological [1627]. Epigrams [3730]. Epilepsy [1124]. Epilogue [3560]. Episodes [1781, 3257]. Epistemes [2313]. Epistemologie [3689, 3013]. Epistemologies [1239]. Epistemology [1308, 947, 1061, 1669, 2572, 1723, 2128, 2577, 1572]. Epistula [1222]. epocas [44]. époque [982, 1225]. Époques [3494]. Equation [2878]. Equations [3167]. equilibrio [2662]. Equilibrium [3360]. Era [2168, 2318, 859, 2094, 1834, 2066, 742, 318, 721, 866, 391]. Eradicate [1962, 2692]. Eradication [3426, 3422, 2214]. Erasmus [378, 2674, 1336]. Erblehre [2763]. Erbpathologie [2695]. Erdos [380]. Eresus [3309]. Erfinder [3605]. Erganzung [34]. Ergo [3153]. Eri [2362, 2362]. Eric [131, 655, 1618, 552, 633, 1144, 3014, 2604, 2494]. Erica [767]. Erickson [2055]. Erie [662]. Erik [709, 2545, 2770]. Erika [764]. Erin [901]. Erinnerungen [194]. Erkenntnis [704]. Erkenntnis- [704]. Erläuterungen [3486]. Ernan [2609]. Ernest [3067, 440, 1840]. Ernst [994, 1639, 3063, 3670, 3707, 305, 3728, 2903, 1731, 282, 3178, 55, 2545, 2583, 2451, 2397, 2831, 324, 229, 1110, 1847]. Erosion [3221, 1866]. Errata [3126]. Erratum [1183, 3129]. Errol [1138]. Error [2666, 3531]. Erste [163]. ersten [2276]. erstes [556]. Erudition [2893]. Eruptions [290, 760, 1205]. Escuela [2430, 692]. Esotericism [2817]. éspace [1350]. Espada [1394]. España [3732, 146, 2742, 768, 625, 2444, 2263, 2366, 300, 18, 1361]. español [768]. española [1399, 2366]. españolas [732]. españoles [1713]. Espionage [272]. essai [3248, 2706]. Essay [2472, 3465, 2606, 2198]. Essays [942, 1312, 1058, 1443, 167, 1888, 92, 565, 2244, 453, 138, 266, 491, 526, 568, 2725, 2341, 1972, 2904, 139, 887, 17, 2819, 177, 1220, 2745, 230, 3781, 2707, 3512, 3552, 934, 2572, 71, 2055, 36, 135, 1982, 3105, 2132, 1258, 1234]. Essence [808]. Essential [1986, 3734]. est [2877]. Established [3255]. Establishment [1025, 595, 2870]. Esther [224, 211, 3000]. Esther-Mirjam [224]. Estrogen [3199]. estructuras [2662]. Estudios [1492]. Etablierung [1964]. Etat [939, 518]. Eternal [1539]. Eternity [2903]. ETH [2855, 387]. Ethan [3086]. Ethical [2969, 408, 1866]. Ethics [1434, 663, 1273, 3469, 1739, 2501, 3747, 3523, 1597, 1606, 3685, 2120]. Ethik [2407]. Ethiopia [1621]. Ethnicity [2137, 1197, 3198]. Ethnographic [2288]. Ethnographica [909]. Ethnography [1935, 1915, 1096, 764]. Ethnology [985, 2288]. Ethology [3536, 2615]. Étienne [2425]. etnográfico [2351]. Étude [2041, 3388, 2343]. Euclid [3145, 1469, 3037]. Euclide [2971]. Euclidean [2618]. euclidienne [3075]. Eudemus [1913]. Eudoxus [2531]. Eugene [1668, 757, 75, 1141, 75, 1005]. Eugenic [2849]. Eugenics [3342, 1744, 2003, 566, 2303, 397, 379, 2501, 3775, 2768, 1151, 1104, 3516]. Eugenie [3224]. Eugenik [2763]. eugénisme [2091]. Euler [3573, 3315]. Europa [3728, 3479]. Europa-Islam-China [3479]. europäischer [2897]. Europe [3157, 3342, 2195, 845, 3662, 1026, 134, 3649, 201, 1343, 442, 886, 1352, 2278, 3064, 3223, 3321, 709, 3401, 2523, 3234, 3046, 108, 2893, 3727, 699, 269, 2830, 3189, 448, 796, 2157, 3049, 612, 883, 2153, 1315, 1321, 946, 1782, 1433, 1437, 3511, 3391, 1495, 2342, 3491, 1202]. European [3236, 1060, 1785, 1632, 789, 985, 1404, 987, 1588, 305, 2746, 3472, 1146, 506, 2355, 3569, 3555, 2576, 2354, 2292, 1795]. Europeanization [2158]. Europeans [2885]. europeo [495]. Euthanasia [1882, 1739]. Eutrophication [662]. Eva [3672, 126, 3433]. Evaluation [2914]. Evangelical [2550, 3003]. Evangelicalism [2985]. Evans [1420, 765, 262]. Eve [199, 176]. Eve-Marie [199]. Evelyn [702, 1912, 351, 929]. Evelynn [1942]. Evemarie [587]. Evenden [2441]. Event [1467]. Events [1397, 422, 1289]. Ever [455]. Everett [2399]. Everhardus [881]. Every [709, 2866, 2181]. Everyday [3211, 3556, 209, 3655]. Everything [3673]. Evidence [2068, 2445, 1728, 2015, 2454, 1893, 2233, 2704, 3015, 1429, 2992]. Evidence-Based [2704]. Evo [3507, 2630]. Evo-Devo [3507, 2630]. evolutietheorie [3766]. Evolution [1432, 2431, 390, 1665, 1888, 605, 1719, 3250, 3295, 1971, 1268, 2691, 2330, 1944, 173, 2755, 1741, 2914, 1293, 2922, 2206, 2583, 2572, 305, 2103, 1970, 3545, 2648, 604, 3556, 281, 457, 2633, 230, 848, 2246, 3433, 2280, 3425, 849, 2614, 3639, 3507, 2449, 2689, 2598, 1601, 1416, 2848, 3613, 3745, 1773, 3224, 3669, 1541]. evolutionaren [256]. Evolutionary [1061, 2630, 1336, 3707, 172, 1547, 2072, 2501, 1525, 1269]. Evolutionism [1187, 2242]. Evolutionist [104, 1941]. évolutionnisme [2619]. Evolutionsbiologie [199, 2103]. Evolutionsdiagramme [3743]. Evolutionsforschung [896]. Evolving [3546, 224, 3556]. Ewan [1188]. Ewen [965]. Exact [1443, 1689]. Exactness [2410]. exakten [3331]. Exaltation [3393]. Examen [3477]. Examination [2577, 2868, 2162]. Examinations [695]. Excavating [3421]. Excavation [1262]. Excavations [608]. Exceeding [2936]. Excellence [2472, 3345, 788, 812, 1156]. Exchange [3281, 3030]. Excluded [1312]. Exclusion [2758]. Exemplary [3358]. Exhibition [367, 1933, 1566]. Existence [2074]. Exobiology [955]. Exotic [1511, 3644]. Expanding [2090]. Expansion [1787, 3283]. Expectancy [1674]. Expectations [224]. expedición [2351]. expediciones [768]. Expedition [2680, 1935, 2055, 1338, 381, 2279, 174, 177, 1251]. Expeditions [2716, 2287]. Experience [2098, 1010, 207, 2542, 3281, 2409, 2576, 508, 1725]. Experiences [3412]. Experiencing [2885]. Experiment [1983, 3317, 3741, 3768, 367, 2290, 1623, 1386, 1630, 1681]. Experimental [2578, 2892, 3282, 1210, 2327, 991, 1375, 1187, 1396, 2974]. Experimentalisierung [2695]. Experimentalism [2798]. Experimentation [1063, 3685, 3368, 3366, 3592, 2862, 2064, 2445]. Experimente [3071]. Expérimenter [3635]. Experimenting [2029, 1436].

Experiments [1501, 2673, 2746, 1709, 220, 1294, 3661, 2428]. Expert [1064, 1892, 3349, 1746, 2348, 2798, 1053, 3536]. Expertise [2869, 2604, 2699, 3201, 3528]. Experts [1043]. Explain [217]. Explained [1086]. Explaining [3725, 2085, 3230]. Explanation [1318]. Explanations [2509]. explorador [1394]. Exploration [974, 2761, 511, 3339, 2457, 1834, 1114, 1070, 1040, 1209, 1611, 1104, 846, 3417, 2887, 542, 1539]. Explorations [3243, 3484]. Explorative [3071]. Explore [1884]. Explorer [3489]. Explorers [3652]. Exploring [1935, 2067, 2084, 2296, 1309, 943, 634, 84]. Explosion [1840]. Explosives [3377]. Exposing [762]. Exposition [3756]. Expression [1128]. Expressiveness [1486]. Expunging [3422]. exstaticum [3406]. Extension [1514]. External [84]. Extinction [1264, 1833, 560]. Extinctions [972, 1213]. Extracts [2442]. extraordinaire [2343]. Extraordinary [1000, 847, 3170]. Extraterrestrial [193, 861, 139]. Extraterrestrials [193, 1867]. Extravehicular [1306]. Extremes [3269]. Extremities [1785]. exzentrische [121]. Eye [3754, 2244, 2950, 330, 1703, 2765, 3382]. Eyes [985, 2646, 851, 3588]. Eyewitness [1293]. Ezio [41].

F [1025, 2175, 2215, 2449, 2285, 2081, 1867, 863, 83, 1414, 2899, 537, 11, 383, 369, 2294, 1561, 192, 2641, 669, 992, 2577, 1565, 1041, 2251, 3031, 1597, 860, 211, 2382, 685]. F. [2639]. Fa [2412]. Fa-ti [2412]. Fabian [974]. Fabiani [586]. Fabien [3316]. Fabric [436]. fábrica [479, 2662]. fabrique [3568, 3355]. Fabrizio [382]. Fabulous [1209, 2760]. face [1322, 1471]. Faced [145]. Faces [26]. Facets [2646]. Facing [63]. Facsimiles [13]. Fact [1692, 28, 3423]. Factor [2374, 794]. Factory [2290, 2181]. Facts [1045, 2082]. facultatibus [1688]. Faculty [2622]. Fadl [264]. Fagan [970, 3038]. Fahnestock [666]. Faigman [603]. Failure [1185]. Fair [3756]. Fairchild [2060]. Fairfield [2289]. Fairy [2052]. Faith [3401, 3595, 2903, 2207, 3167, 3425]. Fakultät [1815]. Falchetta [3146]. Falcon [3039]. Falconer [192]. Fall [3090, 103, 1295, 2050, 683, 1854, 2297]. Fallout [3525]. Fallows [45]. Falls [2944]. Falsely [187]. Fame [902]. Families [752, 321]. Family [3118, 822, 275, 817, 2069, 3263, 711, 2646, 2069, 718]. Famous [2346, 1417]. Fan [2412]. fang [3219]. Fangerau [3101, 1964]. Fantastic [2911]. Fantastical [3559]. Far [818, 1205]. Far-Reaching [1205]. Fara [2658]. Faraday [3071, 370]. Farben [3753]. Farber [2061]. Fardon [2210]. Faria [2279]. Farisi [3486]. Farm [3118, 316, 2853, 2181, 817]. farmacia [129]. farmaco [25]. Farmer [1600]. Farming [175, 1048]. Farrell [2381]. Farther [454]. Fascist [283]. Fashioned [546]. Fashioning [2082]. Fast [313]. Faster [2567]. Faszinationsgeschichte [3177]. Fatal [1251]. Fate [2101, 3337, 1841, 766, 3652, 3456, 992, 3664]. fated [2275]. Fateful [1850]. Father [1248, 3536]. Father-Son [1248]. Fathoming [2761]. Faulk [50]. Fault [661]. Faults [2849]. Fausto [1150]. Fausto-Sterling [1150]. Favorable [421]. Favrholdt [301]. Faye [1904, 151]. Fe [2835]. Fear [1348, 2193]. Fearful [441]. Fearon [45]. February [240]. Febvre [3348]. Federal [1423, 1770]. Fedunkiw [2993]. Feedback [1677]. Feejee [167]. Feelings [3701]. Feferman [2617]. Feingold [1633]. Felicity [686]. Felipe [449, 1602]. Felix [165, 2736, 846, 197]. Fell [1017]. Female [1255, 468, 2479, 577, 2598]. Feminism [569, 459]. Feminisms [2652]. Feminist [2861, 2577]. femmes [3639]. fen [3277]. Fenn [1830]. Fennell [70]. fenomenen [499]. Ferdinand [290, 1002, 1516]. Ferenczi [2075]. Ferguson [49, 122]. Fermat [922]. Fermentation [42, 3298]. Fermi [2519]. Fernand [2595]. Fernanda [2495]. Fernando [2021, 406, 352, 2351]. Ferrari [324]. Ferry [386]. Fertility [3777]. Festgabe [769]. Festschrift [1975, 1668]. Festungsbau [3402]. Feuds [455]. Fever [1627, 488, 773]. Fevers [3426]. Feyerabend [745, 2381, 698]. Feynman [2565, 527]. Feza [3074]. Fi [3433]. Fiber [654]. Fibrosis [3198]. Fichman [2280]. Ficin [2534]. Fiction [2068, 899, 1138, 1107, 1286, 561, 2629, 458]. Fictions [1843, 2052, 142, 1241]. Field [1725, 558, 818, 2169, 1270, 2067, 1537, 3512]. Fields [1616, 2385, 658, 75]. Fieldwork [1024, 400, 2186, 1526]. fièvre [1736]. Fifteenth [37]. Fifties [1138]. Fiftieth [1153]. Fifty [931, 306, 214, 1424]. Fifty-Year [306]. Figala [617]. Fight [1962]. Fighting [1946, 1028, 3596]. Figments [167]. Figueirôa [2495]. Figures [666, 2919]. Figuring [3036]. File [894]. Film [3104, 2526, 659, 1138]. Filmography [1608]. Films [1739, 3423, 561, 729, 3423]. Films/Reich [1739]. filosofare [1705]. Filosofia [3395, 324, 998, 3615, 190, 1492, 3148]. filosofii [1267]. Filth [2987, 3242]. fin [3311, 582]. fina [58]. Finanzmathematik [707]. Finches [2065]. Find [304]. Finding [585, 2061, 176, 3675]. Findings [3768]. Finds [2944]. Finger [2828, 1805]. Fingerprinting [1023]. Finish [2638]. Finite [889]. Finkel [210]. Finn [295, 762, 3465, 1965, 512]. Finns [120]. Finocchiaro [2475, 225]. Finucane [471]. fiorentino [35]. Fire [2333, 2134, 1068, 1295, 2631, 2966, 2024, 2440, 1346, 3567, 590]. Fireballs [2944]. Firenze [2044]. Fireplace [1130]. Fireworks [1581]. First [1029, 1784, 191, 2260, 1218, 931, 3048, 2557, 3085, 1909, 676, 2692, 2030, 916, 3684, 2623, 992, 2365, 385, 1364, 2508, 3150, 1424, 3516, 3590, 3440, 2774, 2552]. Fischer [2697, 609, 1964, 274, 790]. Fischer-Mauch [274]. Fischler [842]. Fish [2418, 2441]. Fisher [2888, 1591, 3503, 731]. Fisheries [1958]. Fishes [2689, 2979]. Fishing [1716]. Fishman [319, 2180]. Fishy [800, 413]. fisica [2504, 714, 2366, 732, 2366]. fisici [2363]. Fisico [2044, 1483]. fisiologia [994]. Fission [2175]. Fit [2845]. Fitness [2768]. Fits [3498]. Fitting [103]. Fitzgerald [2181]. Fitzpatrick’s [908]. Five [205]. Fix [3023]. Fizeaus [2183]. Flagellation [3730]. Flamsteed [2345]. Flash [2097, 192]. Flatland [759, 1249]. Flats [810]. Flattened [2271]. Flatterland [759]. Flaugergues [3572]. Flaws [2657]. Fleeming [1950]. Fleischhacker [631]. Fleming [248]. Flesche [806]. Flesh [1452, 2185, 2049]. Flexner [1858]. Flies [196]. Flight [213, 2030, 2090, 1057, 1884, 214, 2781, 65, 1990, 557]. Flights [214]. Floodplain [2569]. Florence [1745, 2623, 353, 639]. Florian [3297]. Florida [2853, 1825]. Flourishing [1092]. Flow [2814]. Flowers [2544, 1018].

Flowing [1414]. Flucht [642]. Flügel [2837]. Fluid [3488]. Fluids [1842]. Fluoroscope [103]. Flux [3424]. Flying [690]. Föger [2182]. foi [2620]. Folgen [3260, 282]. Foligno [1368]. Folk [1831]. Folkerts [1291]. Folklore [464]. folletos [586]. Following [1641]. fonds [1610, 939]. Food [3144, 1529, 1048, 1580, 1016, 1149]. Foodstuffs [1688]. Foolishness [1973]. Foot [3473]. Footnotes [2359]. Footprints [1924]. Footsteps [1641]. Forbes [1391, 331]. Forbidden [575, 37]. Force [2427, 1235, 1057, 2060]. Forced [564]. forces [3248, 2304, 1672]. Ford [90, 2766]. Förderung [1052]. Fordist [3370]. Forecast [1261]. Forecasting [1764]. Forensic [1040]. Forensics [1040, 2871]. Foreskin [2679]. Forest [2905, 1036, 1198]. foresterie [2492]. Foreword [3324, 3151, 3297]. Forged [93]. Forgotten [2417, 3414]. Form [865, 2843, 512]. Forman [3042, 2668, 995]. Formation [1669, 706, 1735, 1252, 2543, 2278]. formatione [1103, 2543]. Formative [375, 1011, 1387, 1861]. formazione [610, 2282]. forme [3326]. Forrester [186]. Forschung [2661, 1371, 2838, 163]. Forschung/Places [2661]. Forschungen [476]. Forschungspraxis [2183]. Forsythe [1426, 517]. Fortis [2977]. Fortress [3131]. Fortune [623]. Forty [1791]. Foss [1412]. Foss-Mollan [1412]. Fossil [1218, 2865, 444, 1706]. Fossilists [847]. Fossils [3674, 289, 3603, 3205, 2065]. Foster [907, 3536]. Foucault [30, 2500]. Fouché [2562]. Fouke [40]. Foul [2268]. Found [2991, 1005]. Foundation [2993, 753, 962, 434, 1591]. Foundations [1975, 2890, 1473, 2049, 463, 228, 2415, 75, 1537, 792, 1515]. Founding [45, 2615, 541, 331]. Fountain [2751]. Fouque [290]. Four [3098, 2834, 2420, 2726, 3473]. Fourfold [235]. Fournier [2414]. Fourteenth [1583, 147, 1697, 1817]. Fourteenth-Century [1583, 1697, 1817]. Fourth [1365]. Fourth/Eighth [1365]. Fowler [1362]. Fox [2941, 845, 929, 134, 2752, 791]. Foxes [2259]. Foy [693]. fra [2173, 3146]. Fractals [1890]. Fractionation [1747]. Fracture [2218, 3695]. Fragile [1586]. Fragmented [2584]. Fragments [647]. Frameworks [305]. francais [614]. francaise [3348, 2879]. France [2208, 2744, 3293, 533, 1710, 2484, 2172, 3577, 3311, 2243, 3639, 1099, 1512, 716, 3598, 3666, 1610, 272, 1735, 2539, 1186, 2521, 432, 673, 2052, 1111, 2358, 149, 2982, 2266, 3749, 968, 3595]. Frances [3196]. Francesca [3730, 3634, 998, 2335]. Francesco [620, 994]. Francis [621, 1946, 3660, 2890, 588, 2606, 1104, 793, 623, 3366, 2315, 2745, 3125, 2901, 1097, 1156, 591, 622]. Francisco [884, 71, 18, 2377, 640]. Franck [436]. Franco [2444, 332]. Francois [1310, 298, 16, 196, 2202, 2922, 502, 3652, 3074]. francophone [2619]. Frängsmyr [3633]. Frank [2063, 446, 182, 3083, 2068, 3340, 1142, 3657, 370, 176, 371, 361, 292, 2941, 2196, 1941, 2240, 2835, 2194, 2920, 2564]. Frankenstein [1568, 3592, 367]. Frankland [2297]. Franklin [3048, 2828, 1841, 3033, 1630, 897]. Frankreich [2258, 1017]. Franz [288, 288]. Fraser [2441, 1811, 783, 809]. Frau [1505]. Fraud [2583, 1973, 1878]. Frauen [67, 1912]. Frauenberufs [1912]. Fräulein [556]. Fraunhofer [1281]. Frawley [2424]. Frayling [2816]. Frayn [2904]. Fred [784, 2636, 2713, 2551]. Freddy [725, 1109]. Frederic [419, 2856, 2327, 1652, 1623]. Frederick [616, 421, 805, 802, 395, 3169, 3458]. Frederika [1564]. Fredrickson [1148]. Fredrik [2834]. Free [3278]. Freedom [1620, 2772]. Freeland [355]. Freeman [2472, 763, 2960, 1850]. Freemon [182]. Freeze [3076]. Frege [52]. Frei [387]. Freiheit [3418]. Freiherr [3418]. Freire [1619]. French [2350, 2834, 1393, 626, 3577, 2073, 1697, 1812, 3051, 3334, 1460, 1943, 3216, 1368, 789, 2275, 2239, 2717, 2799, 414]. frente [1399]. Frenza [2977]. Frequency [1139]. Frercks [2183]. Fresh [3720]. Freud [2075, 62, 186, 1734, 3067, 2505, 2359, 186, 377, 813]. Frewer [3685]. Freyhofer [2926]. Friedewald [1042, 814]. Friedlander [1873, 1124]. Friedman [2472, 2994, 94, 94, 686]. Friedmann [3331]. Friedrich [1279, 3035, 2841, 2543, 590, 3071]. Friends [1703, 1057]. Fritz [48, 302, 1016, 2361, 934, 2616, 2380, 1574]. Frobenius [888]. Frogs [447]. Front [1, 101, 234, 329, 426, 430, 574, 672, 750, 827, 832, 1335, 1794, 2232, 2518, 1939]. Frontier [2900, 757, 2725, 1927, 175, 990]. frontières [2278]. Frontiers [2120, 3341, 1419, 2849, 384, 1955, 319, 3360, 2570, 1348, 393]. Frontispiece [953, 1067, 1184, 1332, 1448, 2229, 2317, 2388, 2456, 2515, 2795, 2580, 2645, 2711, 2864, 2949, 3017, 3115, 3203, 3280, 3363, 3446, 3534, 3617, 3692, 3789]. Frost [1405]. Frozen [2420]. Frugoni [3147]. Frühaufklärung [2347]. frühe [3227]. fruhen [347, 150, 56, 2840, 769, 3193, 3743, 814, 3035]. Früher [3645]. Fruits [3474]. Fruton [2136, 3298, 3626]. Fry [2326]. Fuchs [616, 474]. Fudge [767]. Fuegians [2065]. Fugitive [2040]. Fulgham [59]. Full [2735, 1422, 1715]. Fuller [664]. Fullmer [239, 991]. Function [973]. Functional [1180]. Functions [2329]. Fund [1870]. Fundamental [3106, 596, 2582]. Fundamentals [770, 1524]. fundamentos [1015]. Funding [1870, 2993, 643]. Fungal [2833]. Funken [814]. Funktionen [3405]. Furth [1092]. Fury [1075]. Fusion [1871]. Füßl [2753]. futura [144]. Future [644, 2037, 215, 3263, 1149, 3425, 1351]. Fyfe [2550].

G [2882, 2941, 243, 1974, 2189, 3670, 2900, 2098, 634, 2613, 2967, 903, 616, 628, 3757, 927, 3095, 2893, 372, 374, 53, 305, 394, 478, 110, 1872, 516, 2374, 766, 2884, 445, 204, 148, 399, 2252, 405, 1054, 3486, 680, 1002, 2934, 342, 462]. G. [2944, 1091]. Gabler [3083]. Gabriel [52, 3060]. Gabriela [2969]. Gabriele [1912]. Gabrielle [533]. Gadgets [2804]. Gaetana [751, 3492]. Gage [2026, 902]. Gágyor [3747]. Gail [2180]. Gaither [544]. Gal [2415]. Galan [1952, 2938, 134]. Galápagos [849]. Galaxies [718]. Galbreath [315]. Galdeano [146]. Gale [601]. Galen [141, 1219, 2029, 2334, 2950, 1580, 1638, 3722, 3648]. Galenism [1638]. Galeno [2250]. Galilée [3157, 2738]. Galilei [1319, 2099, 2411, 360, 3239]. galileiane [3138]. Galileiani [877]. Galileo [3326, 2612, 2099, 3053, 1234, 1703, 3661, 1079, 2475, 3150, 2612, 2221, 2050, 2411, 2797, 2826, 2642, 1805, 2609, 358, 1587, 3395, 3734, 3577, 3404, 682, 3738].

Galison [2496, 581, 111]. Gall [309]. Gallagher [397]. Galliher [1025]. Galton [1946, 1104, 2901]. Galvani [1170, 2352, 447]. Game [911, 2157]. Games [655, 2673]. Gangrene [182]. gao [3561, 3343, 3277, 3186]. Garber [1309, 442]. Garcia [3157, 3054, 1654, 625]. Garden [3018, 1600, 2306, 269, 1460, 3509, 115]. Gardens [1600, 286, 2256]. Garfield [1668]. Garfinkel [205]. Garland [2930, 2447]. Garniss [1416]. Garrett [718]. Garrigós [2995]. Garrigós-Oltra [2995]. Gary [2939, 2039, 1019, 2787]. Gascoigne [1588, 43]. Gases [1590]. Gaskell [1115]. Gasman [283]. Gassendi [2888, 3052, 3739, 2534]. Gasser [2091]. Gaston [2222]. Gates [602, 285]. Gathorne [898]. Gathorne-Hardy [898]. Gatti [691, 356]. Gaugier [2881]. Gaukroger [2416, 1097, 155]. Gausemeier [2838]. Gaußens [2997]. Gauvin [446]. Gavin [1865]. Gaw [641]. Gayle [731]. Gayon [3060]. Gazalé [584]. Gaze [2898]. ge [3343, 3200]. Gebarens [126]. Gebetszeiten [3486]. Geburt [126]. Geburtstag [769]. Geddes [1604, 3450]. Gedé [1526]. Gee [444]. Gegenstand [1584]. Gegenwart [3035, 2144]. Geheimnisse [617]. Gehirn [24]. Gehirne [2530]. Geigerzähler [1956]. Geison [2007]. Geist [3134]. Geistesgeschichte [54]. geistigen [1042]. Gelehrtenbriefwechsel [2897]. Gelehrtenrepublik [67]. Gelman [551]. geloof [2676]. Gema [917]. Geminos [3143]. Gender [2941, 3234, 565, 1744, 3036, 559, 2172, 3748, 936, 785, 1150, 3149, 110, 2094, 2650, 3650, 1107, 1092, 2002, 2577, 3504, 1108, 293, 1480, 1796, 1007, 209, 2799, 1590, 722, 1197]. Gendered [3160, 3643]. Gene [2856, 1880, 3558, 929, 1419, 1754]. généalogie [982]. Genealogy [2244, 2349, 3033, 1277]. General [1404, 706, 747, 1575, 692, 1013, 2391, 1187, 3512, 1829]. générale [3396, 2879]. Generali [3735]. Generalist [2323, 2320]. Generall [706]. Generals [709]. Generating [2478, 3160]. Generation [3345, 555, 3149, 1398, 2343, 1254, 1523]. Generations [1157, 1022, 3775, 1407, 564]. generazione [2270]. Generosity [2123]. Genes [9, 82, 289, 2293, 2573, 1886]. Genese [2291, 3740, 582]. Genesis [3760, 1355, 1970, 2558, 1057, 3512, 2246, 768]. Genetic [1270, 201, 3469, 2641, 3198, 2857, 2771]. Genetically [1909, 1540]. Genetics [1870, 3767, 2572, 1970, 1649, 656, 433, 307, 1542, 1349, 2856, 2685]. Genetiker [2695]. genetiki [2191]. Geneva [59]. Geneviève [2091]. Geniale [2530]. Genius [3090, 2045, 3485, 1730, 1865, 2163, 2674, 3065, 3594, 689, 1086, 3048, 2050, 3513]. Gennady [2772]. Genocide [1026]. Genome [408, 1541]. Genomes [2573]. Genomics [3767, 2638, 2632]. Genova [691]. Genres [899, 3540]. Gentile [1368]. Gentle [2021]. Gentleman [1007]. Gentlemen [3646, 1937, 292]. Gentzen [3594]. Genuine [3287, 61]. Genuth [3524]. Geodesy [2697]. Geodetic [2716]. Geoffrey [114, 2822, 2154, 600, 781, 2663, 2249, 2248]. Geoffroy [2425]. Geographer [2436, 2922]. Geographers [2457]. Geographic [2827, 626]. Geographical [2663, 2420, 2156, 1370, 2079]. Geographically [3249]. Geographies [1102, 2760, 2450]. Geography [1683, 2922, 1533, 1700, 1114, 2912, 1359, 846, 1327, 366, 626]. Geohistorik [2837]. Geohistory [2743, 3709]. Geologic [445]. Geological [3642, 1554, 2966, 1706, 3253, 3333, 3473, 1930]. geológicas [2495]. Geologist [3602, 2754]. Geologists [1416]. geologo [1999]. Geology [510, 1388, 372, 3584, 2726, 3675, 1287, 3333, 3642, 2535, 3497, 688, 3582]. Geometría [2662]. Geometric [2843]. Geometrical [147, 2410, 1589]. Geometrie [1470, 1226, 1907, 3159, 3075]. geometrische [2331]. Geometry [1634, 3467, 1469, 889, 872, 2043, 3313, 2618, 3087, 1201]. Geophysical [2713, 3076, 2923]. Geopolitics [82]. Georg [994, 1815, 758, 3331, 363, 2284]. George [650, 2735, 44, 2472, 193, 2546, 713, 1655, 249, 2171, 2791, 3454, 915, 1382, 3453, 2036, 3456, 874, 3452, 3074, 1641, 723]. George-Louis [44]. Georges [3634, 685, 2178, 866]. Georgia [1717]. Georgian [1829]. Georgina [2995, 386]. Georgius [881]. Gerald [628, 259, 3008, 1946, 2007, 2660, 2047]. Gerard [2402, 2170, 2600, 2730, 2088]. Gerbert [2735]. Gerhard [704, 3594, 3070, 2846, 2144]. Gerhardt [3761]. Gerhild [2895]. Gerlinde [163]. Germ [2658]. German [1314, 1779, 3670, 2834, 3130, 1268, 1401, 520, 195, 1237, 1477, 482, 2910, 2558, 405, 302, 2380, 2201, 895, 860, 917, 3414, 1611, 2233]. German-Jewish [3414]. German-Speaking [2201]. Germana [3728]. Germany [2895, 927, 3609, 1276, 2907, 1739, 3519, 2494, 3172, 1606, 1995, 1450, 2288, 3259, 1117, 2605, 2501, 1581, 3397]. Germs [637, 905, 955, 2987]. Gerrit [3312]. Gertrude [1567]. Gesammelte [3762]. gescheiterten [309]. Geschichte [1178, 1324, 1815, 3685, 3578, 3657, 199, 2370, 1734, 2553, 1912, 309, 1505, 2143, 150, 2144, 814, 3171, 1956, 3013, 2530, 1574, 2818, 1944, 1815, 3035, 1478, 2368]. Geschichtsphilosophie [3578]. geschiedenis [2724]. Geschlechterfrage [67]. Geschossfotografien [1639]. Geschwind [2089]. Gesell [752]. Gesellschaft [2117, 1052, 3477, 3761, 56, 3515]. Gesellschaftsdeutungen [635]. Gesprach [121]. Gespräche [1260]. Gessner [3736]. gest [264]. Gestalt [2684, 405]. Get [2805]. Getting [3023]. Getz [151]. gg [1267]. Ghamari [2635]. Ghamari-Tabrizi [2635]. Ghiselin [23]. Ghosts [2809, 3690]. Gian [2726]. Giancarlo [877]. Gianna [693, 2893]. Giant [2859, 2714]. Giants [2486]. Giberson [1846]. Giberti [1933]. Gideon [1522]. Gidley [89]. Giere [3529, 417]. Gieryn [11]. Giessen [274]. Gift [2732, 3196]. Gifts [1682]. Gilbert [2162, 767]. Gilded [2714]. Giles [3091]. Gilgen [91]. Gilles [2091]. Gillham [1104]. Gillian [1950]. Gillispie [2484]. Gillmor [2551]. Gilman [166]. Gilson [869]. Gimbel [1057]. Gimelli [2745]. Gingerich [2476]. Gingras [2721]. Gingrich [2834]. Ginzberg [819]. Gioia [495]. Giordano [266, 1049, 3148, 266, 356]. Giorgio [610, 2028]. Giovanna [2562]. Giovanni [526, 1599]. Gipper [2258]. Giraffe [505]. Girl [2866, 1709, 398]. Girls [1835]. Girolamo [1499]. Giselher [377]. Giselle [196]. Gissis [665]. Gitelman [721]. Giuseppe [1917, 1049]. Givens [3232]. Giving [1132, 1860]. Glands [2918]. Glasgow [3653]. Glasmeier [1951].

Glasner [1583]. Glass [2532, 66]. Glasses [2957]. Glassmaking [1586]. Glassworks [3404]. Glassy [808]. glave [1354]. Gleditsch [2124]. Glen [1767]. Glenda [562]. Glenn [2288, 1618]. Gliboff [3670]. Glicksman [3703]. Glimpse [2992]. Glitz [2930]. Global [650, 1180, 949, 950, 1984, 1951, 1157, 1962, 406, 2953, 1674, 601, 1297, 3061, 1948, 2990, 3498, 2145, 570, 3190, 296, 2786, 2478, 2214]. Globalization [2913, 2436]. Glory [390, 182]. Glosae [2041, 473]. GM [431]. Gnosis [2817]. God [849, 3164, 2985, 871, 3768, 3167, 3032, 700, 2520, 1095, 2437, 1491, 3492]. Godel [71, 522, 1349, 522, 1515]. Godin [2906]. Godless [398]. Godlewska [626]. Godly [3003]. Gods [464, 2809, 286]. Goerler [212]. Goes [3756, 3192]. Goethe [2272, 507, 2167, 1597, 1660]. Goethes [3047, 629]. Goggin [738, 738]. Gogol [567]. Gohau [3060]. Going [1405]. Golan [665, 2348]. Golas [2027]. Gold [1940, 1876, 608, 775]. Goldberg [1142]. Goldberger [1859]. Golden [3030, 3244, 3603]. Goldie [900]. Goldsborough [3005]. Goldstein [2404]. Golem [2788]. Golgi [58]. Golinski [796]. Gómez [3604]. Gondhalekar [1080]. gong [3262, 3200]. Gontier [2737]. González [768, 2829, 1048]. Good [45, 2981, 1129, 1946, 1575, 152, 3332, 1208, 1409, 1077, 1395, 422]. Gooday [2752, 3748, 2310]. Goodchild [2503]. Goodeve [932]. Goodman [2771]. Goodness [2878]. Goodstein [3170]. Goodwin [1132]. Goon [1622]. Goonatilake [570]. Gorbachëv [1620]. Gordin [2357]. Gordon [809, 1616, 926, 1886]. Gorelik [2772]. Gorilla [3295]. Göring [1274]. Gorn [2090]. Gors [171]. Gospel [637]. Gossin [3504]. Gosztonyi [3702]. Got [3713]. Gothic [3232]. Gottfried [3103, 52, 707]. Gotthard [2480]. Göttingen [2997]. Gottweis [201]. gou [3261]. Gouguenheim [3649]. Gouk [1922]. Gould [1455]. Gouvêa [2021]. Gove [734]. Governance [2638, 1064]. Governing [201]. Government [2697, 1423, 2438, 2234, 576, 925, 1101]. Govoni [3319, 2282]. Gowland [2970]. Gozza [778]. grabado [1602]. Graceful [2712]. Gracefully [1854]. Graeco [1365, 577]. Graeco-Arabic [1365]. Graeco-Roman [577]. Graells [3056]. Graeme [2752, 3748, 2298, 2310]. Graf [707]. Graffi [2028]. Grafton [873, 2253, 612]. Graham [378, 1114, 2369, 1886, 1603, 207]. Grain [1571, 3597]. Gramsci [225]. Gran [912, 176]. Gran-Aymerich [176]. Granada [39]. Grand [2285, 2664, 1337, 3483]. Grandin [2657]. Grange [518]. granic [1291]. Grant [1580, 799, 3379]. Graphical [2712]. Graphics [524, 3634]. graphische [719]. Grasp [2936]. Grattan [2722]. Grattan-Guinness [2722]. Gravitation [3512, 1080, 1236]. Gravitational [2439, 3173]. Gravity [2991, 1236, 1080, 2439]. Gray [1044, 2345, 3257, 2618]. Great [2987, 1636, 1716, 3231, 2093, 1115, 3502, 616, 2779, 1830, 2692, 661, 1087, 3073, 3652, 174, 3598, 842, 1805, 697, 560, 1079, 1893, 455]. Greatbatch [1625]. Greatest [3485, 450]. Greco [772]. Greco-Roman [772]. grecques [3649, 1488]. Greece [3725, 981, 2248, 468, 2249, 2038]. Greek [1362, 1218, 1486, 2402, 2600, 3480, 3723, 3385, 2333, 1365, 835, 466].

Greeks [1217]. Green [2208, 1420, 82, 397, 2891, 3750, 1879, 913, 1940, 2256]. Greenberg [2464, 2659, 3523, 1866, 683, 813, 410]. Greene [731, 1957]. Greenhouse [601]. Greening [3130]. Greenland [2420]. Greenspan [2681]. Greenwich [1807, 1807, 541]. Greenwood [3638, 2497]. Gregg [659]. Gregor [1349]. Gregorio [2903]. Gregory [595, 203, 422, 1526, 2836, 3494]. Greifswald [1815]. Greiner [361]. Grene [3703]. Grenzen [1291]. Grenzüberschreitungen [2763]. Greppin [772]. Gresham [588]. Grey [316]. Gribbin [966, 527, 527]. Griffith [375]. Griffiths [12]. Grimaldus [472]. Grimoult [2619]. Gringauz [83]. Grip [1080]. Grit [308]. Grit-Tempered [308]. Grmek [754]. Grob [2942, 2660, 2047]. Groddeck [2075]. Grooves [721]. Grosholz [497]. Gross [1974, 1157, 2934]. große [3406]. Ground [2520, 2082]. Groundbreaking [1501]. Grounds [3273]. Group [1180, 949, 950]. Groups [1258, 1258]. Growing [3085, 2509]. Growth [2555, 406, 90, 915, 2780, 1343]. Grubrich [62]. Grubrich-Simitis [62]. Grunden [2839]. Grundlage [163]. Grundlagen [274]. Grundzuge [256]. Guagnini [845]. Guangqi [1369]. Guardians [1355]. Gudrun [1912]. Guedj [1356]. Guenter [2741]. Gueriguian [772]. guérir [1735]. Guérout. [3320]. guerra [2363]. Guerraggio [2996]. guerre [1593]. Guerrini [2029, 713]. Guesnerie [16]. Gugerli [2184, 2855]. Guicciardini [484]. Guidance [321]. Guide [1179, 1165, 2088, 96, 2314, 1966, 445, 2034, 1161, 3531, 848, 941, 338]. Guided [3714, 463, 2019]. Guides [3638, 152]. Guido [382]. Guidobaldo [102]. Guihan [3219]. Guilaume [1310]. Guild [3019]. Guillelmus [473]. Guillemin [1961]. Guillermo [2430, 692]. Guinea [1770]. Guinness [2722]. Guinot [1139]. Gulgul [34]. Gunergun [3074]. Gunnison [2076]. Gunpowder [1581, 3377]. Gunther [387, 2819]. Guo [2822, 3261]. Gustav [1383]. Gustave [1006]. Guston [545]. Gutachten [1519]. Gutachten- [1519]. Gutas [1365]. Gutenberg [3387]. Guthrie [26]. Guttmann [377]. Guy [3255, 355]. Guyer [160]. Gwyneth [1018]. Gyan [246]. Gynaecology [3650]. Gypsy [2692]. Gyung [2349].

H [736, 973, 2944, 402, 2472, 1358, 191, 1896, 161, 85, 545, 140, 1578, 2771, 2542, 727, 2632, 82, 892, 555, 50, 147, 1629, 349, 2914, 458, 3458, 2725, 374, 2604, 2951, 2345, 891, 2926, 708, 923, 1575, 1083, 371, 3665, 3773, 2819, 3529, 3651, 113, 1079, 2090, 420, 289, 1410, 1396, 591, 792, 2313, 2573, 815, 3145, 1416, 1597, 2424, 2396, 823, 1243, 3670, 3757, 1872]. H-Z [1872]. H. [2288]. H.-B [161]. H.M.S. [724]. hôpitaux [939]. ha-Penini [1583]. Haack [2790, 135]. Hård [3472, 2964, 536]. Haber [2380, 1016]. Habilitationen [1519]. Habitat [1476]. Hächler [2897]. Hacker [1760]. Hackett [36]. Hacks [2215, 2319]. Hacohen [1011]. Haeckel [3670, 3707, 2903, 282, 1847, 3178, 55, 2583, 283, 3179]. Haeckel-Wilhelm [1847]. Haeckels [3063]. Hagerman [652]. Haggett [296]. Hagner [2530, 24]. Hahn [2402, 1912]. Hahnemann [128]. Hail [563]. Haim [1011]. Haiyan [3262]. Haj [2082]. Hakfoort [837]. Hal [455, 204]. Halbwachs [3348]. Half [1784, 1626, 3431, 385, 1424, 3431]. Half-Century [385]. Half-Lives [3431]. Half-Truths [3431]. Halfte [171, 2276]. Halfway [3437]. Halifax [2993]. Hall [1419, 859, 1918, 2773, 271]. Halle [2143]. Halle-Wittenberg [2143]. Haller [181, 1009, 1103, 2746, 2347, 1605]. Hallers [1827, 2897]. Halley [359]. Hallgrímsson [2200]. Halliday [3699]. Hallion [2030]. Halliwell [508]. Hallyn [2595]. Haltung [525]. Ham [3681]. Hamblin [2698]. Hamilton [2063, 889]. Hamlin [184]. Hamm [506]. Hammonds [1942]. Han [31, 3261]. Hanafi [880]. Hanauer [230]. Hancocks [1351]. Hand [3475, 2725, 2997, 3239]. Handbook [3144, 3530, 147, 747]. Handbuch [867]. Hands [519, 1757, 1039, 883]. Handschriften [95]. Handwerk [619]. Hanegraaff [2817]. Hankinson [459]. Hannaway [3680, 2806, 514]. Hannover [3103]. Hanns [2611]. Hanover [3506]. Hans [2695, 1815, 636, 2560, 480, 126, 264, 3260, 2942, 2284, 3096, 2763, 3013, 3581]. Hans-Christoph [126]. Hans-Georg [2284]. Hans-Jörg [3013]. Hans-Liudger [636]. Hans-Uwe [3260]. Hans-Walter [2763]. Hansen [2640, 2774]. Hanson [1945]. Hantzsch [630]. Happened [2811, 1410]. Haraldur [669, 760]. Harari [2575]. Haraway [932, 932]. Harbor [1021]. Hard [1412]. Harding [2861, 685]. Hardy [1948, 898, 799, 3504, 1125, 3031]. Hargittai [2472, 1417]. Haring [3681]. Harkness [611]. Harley [2564]. Harlow [1622, 249]. Harm [305]. Harman [2775, 169]. Harmon [1974]. Harmonious [3259]. Harold [2296, 437, 2296, 2690]. Harper [1003]. Harriet [3158, 167]. Harriot [791]. Harris [446, 3587, 1398, 257, 272, 252]. Harrison [348]. Harro [2756]. Harry [734, 1622, 2076, 1393, 1120, 2439, 1559, 2833, 2788, 3064, 917]. Hart [2942, 711, 93]. Hartley [1128, 1919]. Hartmann [2276, 1498]. Hartog [16]. Hartouni [392]. Hartvig [3026]. Hartz [2178]. Harvard [69]. Harvest [2367, 1149]. Harvey [268, 2750]. Harwell [2169]. Harwood [2907]. Hashagen [2563, 3171]. Häsler [1004]. Hasok [2734]. Hass [1405]. Hasse [363]. Hatred [3669]. Hatten [2923]. Haught [2449]. Haunting [1040]. Hauptportal [3515]. Hauptschriften [707]. haut [472]. haute [790]. Hawking [2113, 2642, 1079]. Hawkins [2865]. Haycock [2669]. Hayden [174, 1002]. Hayek [2294, 2294]. Hayes [1027]. Hayles [2776, 323]. Haynes [1126]. Hayter [2682]. Haytham [1364]. Hazard [2089]. Hazel [669]. Hazen [1896, 1077, 1932, 388, 697]. Head [1221, 3772]. Headed [1828]. Headrick [853]. Heads [1665]. Heal [641, 1338, 391]. Healer [1594]. Healers [2625, 693]. Healing [2111, 2077, 2809, 1197, 823, 258, 3648]. Health [946, 3108, 348, 221, 184, 1859, 345, 2728, 2385, 1528, 1844, 1166, 3607, 1861, 1127, 1135, 2174, 1155, 317, 1963, 858, 3680, 2993, 2266, 316, 2374, 3610, 1757, 2053, 2845, 3598, 376, 2832, 1673, 124, 3172, 3231, 2909, 1745, 1125, 2285, 1948, 2913, 2446, 553, 1433, 3383, 820]. Healthcare [1224]. Healthy [1778, 3551]. Healy [404, 863, 318]. Hearing [3478, 2216]. Heart [2081, 394, 505, 477, 1468]. Hearts [3697, 394, 1155, 855]. Heath [2771]. Heather [644, 393, 322]. Heaven [3034, 2881, 2566]. Heavenly [2335].

Heavens [3235, 2883, 359, 2477, 478, 3386]. Heavy [313, 675, 383]. Heazle [2927]. Hebert [369]. Hecht [2358, 2777, 654, 533]. Hedin [677]. Hegel [2312]. Hegemony [3189]. Heidi [520]. Heiko [2205]. Heilbron [1840, 567]. Heinecke [153]. Heiner [3101, 1964]. Heinrich [1093, 2843, 176, 2144, 195, 53]. Heinz [150]. Heirs [2821, 3723]. Heisenberg [1519, 3008, 3508, 2904, 1260]. Helander [480]. Held [674]. Helden [592]. Helen [1896, 468, 3650, 3529, 2101, 2761, 1624, 2147, 3066]. Helena [209]. Hélène [1558]. Helfand [1578]. Helge [2621, 1407]. Helix [3204, 1339]. Hellebust [2185]. Hellenistic [3143, 3037, 466]. Heller [2091, 616]. Hellman [455]. Hellyer [2605]. Helm [2485]. Helmholtz [3068, 3178, 3068]. Helmholtzian [1967]. Helmont [153, 1233]. Helms.S [1175]. Helmut [596, 194, 1401, 2837]. Helmuth [295, 762]. Helsham [365]. Helsinki [3685]. Helwig [1815]. hemelse [499]. Hemisphere [1673]. Hemophilia [549]. Hempstead [1950]. Henderson [66, 524]. Heng [2276]. Heng-an [2276]. Henk [2410]. Henle [3178, 3747]. Hennessey [861]. Henri [2922, 48, 200, 1004]. Henry [1641, 175, 1896, 2254, 3340, 257, 40, 2889, 3001, 371, 2283, 444, 2766, 84, 3772, 823, 2089, 1797, 2234, 1070, 2302, 1398, 2765, 2289, 2483, 3340, 1918, 727, 1612, 1724]. Henslow [1119]. Hentschel [2840, 2997, 170, 68, 1386]. Hepatitis [1303]. Hepler [1135]. Her [1558, 3144, 2362, 3599, 1147, 1451, 1850]. Herbal [616]. Herbalism [1490]. Herbals [2155]. herbaria [480]. herbarum [3483]. Herbert [3420, 2754, 201, 997, 487, 2678, 1336, 2112, 3256]. Herbicides [576]. Herbs [2809, 1490]. Herder [1247]. Here [3623, 3713, 2903]. Hereditarian [1946]. Heredity [3516, 900, 2901, 1880, 3303, 3469]. Hereford [1690]. Heresy [186]. Heretic [2435]. Heretics [3636]. Hereward [2490]. Heringman [2535]. Heritage [1889, 541, 1591, 1546]. Herkunft [256]. Herle [177]. Herman [2635, 368, 2487]. Hermann [3726, 309, 3178, 587, 1013, 387]. Hermaphrodites [3234, 297]. Hermeneutics [1174]. herméneutique [3566]. Hermes [3297]. Hermeticism [3297]. hermetischen [617]. Hernández [884]. Herodotus [1915]. Heroic [466]. Herreman [1014]. Herring [312]. Herrnstein [2935]. Herrschaft [48]. Herschel [2556, 2062, 2712]. Herschels [3506]. Hershey [897]. Hertha [1342]. Hertler [896]. Hertz [2843, 53]. Herz [842]. Herz-Fischler [842]. Herzenberg [892]. Hese [1494]. Hess [3638]. Hessen [274, 3320]. Hester [1807]. Heterodoxies [823]. Heuck [176]. Heuvel [2825]. Hewett [1941]. Heyck [2678]. Heyes [1061]. Heyman [3105]. Hibner [920]. Hickman [1427]. Hidden [1349, 1957, 2666, 3618]. Hide [1968]. Hielkunde [1089]. High [452, 546, 2780, 2770, 452]. High-Altitude [452]. High-Speed [2770]. High-Stakes [546]. High-Tech [546]. Higher [3271]. Highest [634]. Highland [1526]. Higton [1807]. Hilaire [2425]. Hilary [356, 1134]. Hilbert [1473]. Hildegard [2972]. Hilgartner [1053]. Hill [775]. Hillel [13]. Hills [51, 3050]. Him [3700]. Himmel [48]. Himrod [1165]. Hindle [1071, 1077]. Hine [1921]. Hinter [1985]. Hintereingang [3515]. Hintergrund [1219]. Hintergründe [725]. Hippocrate [468].

Hippocrates [1313, 2602, 1914]. Hippokrates [3260]. Hippolyte [2183]. Hirabayashi [2186]. Hirai [2534]. Hirnforschung [3134]. Hirnströme [3058]. Hiro [2534]. Hiroshima [734, 208]. Hirschmann [2751]. Hirsh [537]. Hiscock [872]. Hispana [1713]. Hispanic [3104]. hispanica [586]. Hispanu [352]. Hissarlik [176]. Histoire [969, 3400, 2823, 761, 3495, 2738, 2619, 976, 1806, 1014, 2988, 2619, 2905, 298, 3396]. Histology [3620]. Historia [884, 146, 3053, 2174, 190, 1976, 3054, 616, 498, 2495, 2893, 2142, 1361, 2040, 129]. Historian [1772, 2525]. Historians [3697, 3540]. Historic [861, 2225, 122]. histórica [1696]. Historical [942, 1323, 1980, 1994, 3108, 3110, 418, 3542, 679, 2680, 378, 975, 2720, 3710, 3059, 3539, 2700, 2306, 2723, 2362, 3186, 133, 850, 3687, 774, 526, 1814, 2314, 1669, 1835, 2904, 2596, 110, 2420, 416, 2996, 408, 1345, 2128, 979, 2717, 3322, 2326, 680, 2787, 2702, 3228, 1682, 1542, 1893, 22, 3531, 124, 2509, 1850, 3381, 2472, 75, 1059, 1434, 1979, 248, 528, 3782, 3531]. Historicity [2946]. histórico [2166, 2474]. histórico-naturais [2166]. Histories [2111, 3328, 2983, 2963, 3003, 3554, 3066, 2760, 115, 1065, 218]. Historiographical [3782]. Historiographie [2619, 360]. Historiographie-Mathematik-Wirkung [360]. Historiographies [1239]. Historiography [3288, 1857, 2650, 3080, 2943, 3466]. Historique [3649]. historiques [1907]. Historische [1681, 3134, 2194, 3689]. History [1029, 3002, 3034, 602, 1263, 3755, 2, 427, 1180, 1309, 1312, 1316, 1325, 1328, 943, 1175, 1432, 1781, 1435, 1441, 1633, 2105, 2113, 2114, 1980, 1989, 1996, 2517, 2937, 3100, 3105, 2944, 2796, 3116, 3447, 3448, 1201, 1357, 800, 2811, 280, 3319, 242, 684, 2215, 3467, 2463, 1358, 562, 1665, 2429, 2364, 1362, 2767, 1896, 3243, 3496, 1716, 1264, 3640, 3670, 1689, 2936, 221, 1664, 3549, 3380, 3069, 3507, 1420, 2021, 3754, 2171, 3711, 223, 2078, 2058, 254, 119, 2663, 3561, 970, 3261, 1000, 2520, 3325, 3376, 1808, 2557, 592, 2244]. History [1666, 2132, 2720, 3391, 452, 690, 14, 3144, 117, 3449, 2045, 1738, 3009, 2392, 3454, 3736, 2875, 844, 1163, 2330, 3701, 3686, 3608, 3304, 3133, 3548, 69, 2953, 2531, 3300, 3343, 989, 1809, 3057, 3559, 3345, 8, 559, 3271, 3633, 3123, 1168, 3122, 517, 717, 2170, 1023, 3061, 3475, 3208, 3369, 3760, 456, 608, 1404, 1414, 1554, 2031, 1008, 1193, 1100, 646, 568, 3481, 1743, 186, 1564, 3209, 2477, 882, 219, 934, 2808, 96, 737, 3226, 3505, 1598, 3259, 2991, 315, 3557, 2963, 2990, 1883, 2994, 1154, 1258, 1863, 2876, 1076, 1958, 12, 747, 2438, 2348, 2323, 1167, 2692]. History [2443, 166, 421, 695, 127, 583, 858, 251, 3367, 2462, 294, 2928, 1124, 2526, 1047, 1726, 3302, 770, 2993, 2740, 588, 2037, 110, 956, 3219, 1546, 2399, 3131, 1212, 1885, 804, 598, 1234, 3154, 1766, 2187, 3210, 3207, 3455, 1087, 19, 192, 3257, 1383, 1758, 2374, 3139, 3170, 1296, 2140, 1425, 1882, 3637, 450, 1460, 3307, 1739, 1549, 779, 322, 1764, 2084, 2441, 185, 3599, 3517, 3232, 2594, 2190, 2660, 440, 3189, 1146, 1205, 3456, 3146, 2657, 2587, 3675, 3022, 3480, 3569, 28, 928, 1611, 1239, 2373, 967, 2383, 1481, 1573, 3708, 2952, 3435, 2964, 1609]. History [746, 2269, 3024, 3294, 2819, 1539, 1570, 3298, 1092, 2853, 3618, 2569, 552, 3120, 433, 3322, 1933, 1367, 1464, 2302, 312, 1220, 2360, 3238, 3747, 3666, 3582, 2597, 1041, 1631, 3082, 1769, 1091, 1672, 3199, 760, 1370, 3384, 437, 258, 2322, 457, 3750, 3423, 503, 590, 3720, 1057, 444, 215, 1884, 403, 2814, 339, 231, 2064, 1623, 965, 3135, 3176, 1673, 2047, 84, 3188, 2564, 2585, 3632, 2494, 3301, 3382, 1472, 3137, 2773, 3297, 860, 3049, 2707, 2972, 3212, 343, 857, 3172, 3350, 830, 1333, 1792, 2230, 2329, 112, 664, 688, 1829, 1862, 599, 2138, 3306]. History [3644, 2325, 685, 262, 3196, 579, 3625, 3059, 3638, 1755, 651, 2762, 2563, 243, 167, 44, 551, 1848, 1940, 1347, 1627, 125, 2035, 2242, 2815, 213, 2899, 88, 2036, 3140, 2024, 2034, 2965, 919, 81, 686, 443, 1346, 1645, 1787, 3100, 3111, 2946, 1265, 3357, 3130, 3685, 108, 1674, 451, 1474, 1742, 70, 1192, 1276, 1477, 2017, 1211, 1210, 1703, 651, 3354, 3560, 1969, 3478, 385, 963, 3487, 1663, 1771, 3377, 2506, 1198, 3658]. Hitler [1260, 642, 2501]. Hitoshi [1406]. HIV [218]. HIV/AIDS [218]. Hoax [1841]. Hoaxing [1850]. Hobbe [1704]. Hobbes [3398, 1231]. Hochschild [1401]. Hochschule [194]. Hockey [682]. Hodgkin [386]. Hoechst [3753]. Hofer [2284]. Hoff [552]. Hoffleit [1874]. Hoffman [1127, 380]. Hofmeier [617]. Hogendijk [843, 3646]. Holbein [2042]. Holism [80, 405]. Holistic [112]. Holland [2811]. Hollow [3559]. Holly [2052, 2283]. Holmberg [1383]. Holmes [2017, 3410, 2856, 2327, 1652, 1623, 287]. Holmesburg [220]. Holocaust [2926]. Holographic [3188]. Holthuis [2979]. Holton [3008, 259, 259]. Holtzmann [2566]. Holy [3493]. Holyoke [2622]. Homburg [305]. Home [2124, 1107]. Homeland [3342]. Homeless [1159]. homme [1904]. Homo [24]. Homoeopathy [1435]. Homosexual [2302]. Homosexuality [1675]. Honest [3353]. Hong [1864]. Honor [1058, 1975, 1668, 3633, 3244]. Honorat [1820]. Honoré [3572]. Honour [3105, 3646, 3552]. Hooijmaijers [2942]. Hook [2187]. Hooke [2415, 2975, 2159]. Hooker [3746]. Hoolihan [1166]. Hoover [1129]. Hope [45, 138, 2682, 3469, 2193]. Hopes [2458]. Hopkin [3668]. Hoppe [770]. Horace [973, 605, 712, 2886, 2554]. Horace-Bénédict [2886]. Horary [1807]. Horgan [63]. Horizons [1494, 1033, 122]. Hormone [3199]. Hormones [2918, 3282, 576]. Horn [382, 3051]. Hornblum [220]. Horney [92, 565]. Horoscope [753]. Horoscopes [2819]. Horoscopy [2335]. Horowitz [293]. hors [2278]. Horst [2926]. Horticultural [3474]. Hösle [3012]. Hospice [97]. Hospital [128, 1854, 2932, 1575]. Hospitals [2844, 2268, 2197, 819, 2688, 550, 1464]. Hoßfeld [1944, 2818]. Hostage [623]. Hostetler [1096]. Hot [2778]. Hotson [1093]. Houck [2778]. Houghton [669]. House [1122, 3233]. Houzel [1907]. Hovel [163]. Howard [658, 193, 995, 203, 264, 1093, 1076, 2567, 2766, 203]. Howarth [2944]. Howe [1086]. Howell [1095]. Howes [892]. Höxtermann [1178]. Hoyle [1018, 2636]. HP [939]. Hsi [870]. HSS [1896, 2394, 2527, 2393]. Hu [2683, 3186]. Hua [3219, 3343]. Huang [3726, 2533]. Hubble [1033]. Hubbub [3242]. Huberman [2178]. Hubert [1371, 2746]. Hubris [2964]. Huerta [2662, 2486]. Hugh [1978, 639, 1388, 2035, 918, 3602]. Hughes [1058, 53, 2658, 1058]. Hugo [1721]. Huguenot [3131]. hui [3517, 3333]. Huibert [499]. Huisman [2564]. Hull [1061, 1888]. Hulse [519]. humain [2549, 3568]. humaine [2671, 2737]. humaines [2537]. humains [3635]. Human [973, 1785, 783, 436, 3231, 3611, 3685, 975, 2632, 1419, 542, 3593, 1677, 3690, 118, 3149, 220, 1880, 3468, 408, 1205, 1810, 2570, 553, 767, 1667, 1621, 746, 3669, 3612, 1541, 2552, 3664, 1601, 1416, 1919, 3066, 2935, 3428]. humana [2386]. humani [2040]. Humanism [2152, 2344, 2114]. humanismo [146]. Humanist [2432]. Humanistic [2134]. humanité [518]. Humanities [3532, 326]. Humanity [3156, 450, 3721]. Humankind [3586, 1463]. Humans [2108, 1028, 412, 2029]. Humble [2045]. Humboldt [1320, 2847, 3417, 626]. Humboldt-Aufbruch [1320]. Hume [624]. Humphry [991]. Hundisburg [3103]. Hundred [812, 2472, 601, 3315, 547, 2966, 399, 2028, 307, 340, 937]. Hunger [2989, 168, 3257, 1422]. Hunt [1947, 1303, 1054]. Hunter [3158, 960, 2159, 2298, 2975, 2678]. Hunters [1218]. Hunting [894]. Huntington [2259]. Hupp [1920]. Hutchison [1018]. Huygens [3316, 3237]. Huysbou [2825]. Huzel [3062]. Hybrids [2964]. Hydraulic [1414]. Hydro [1178]. Hydro- [1178]. Hydrodynamics [2814]. Hydrological [673]. Hygiene [1948, 3663]. Hypatia [3308]. Hype [2696]. Hyperprofessionalism [2322]. Hyperspace [1469]. Hypnose [1856]. Hypnose-Labor [1856]. Hysteria [2178, 253]. Hysterical [1855].

I. [3449]. Ian [3328, 759, 286, 231, 22, 857, 685, 2285, 2603, 1603]. I’AP [939]. I’AP-HP [939]. IAU [385]. Iberian [2885, 3243, 1493, 1390, 2237]. IBM [431]. Ibn [2879, 34, 3486, 1364, 264, 1226]. Ibrahim [843, 1226]. Ice [970, 1297, 2420, 2966, 389]. Iceland [3767, 2200]. Icelandic [2420]. Iceman [734]. ICES [1624]. Ichthyology [2689]. Ichuanxue [307]. Icon [3622, 2127, 116]. Iconicity [2867]. Iconoclast [1858]. Iconography [2178, 118, 1376]. Icy [2791]. Ida [1896]. Idea [1727, 125, 903, 3742, 180, 2160, 3102, 1908]. Ideal [1130, 2181]. Idealismus [2831]. Ideals [2647, 3278, 1032]. Ideas [631, 3029, 3633, 3490, 3581, 2970, 760, 1805, 2541, 1564, 2663, 1783]. ideengeschichtlichen [162]. idées [2738]. Identification [1023]. Identities [1204, 2319, 1023, 1239]. Identity [2638, 1359, 533, 1389, 1256, 320, 1253, 2424, 1843, 2827, 1120]. ideologia [877]. Ideologie [1017, 1317]. Ideology [2392, 3285, 310, 1649, 283, 2204]. Ier [701]. If [3284, 2862]. Ifrah [685]. IG [3753]. Iglesia [3053]. Igneous [1877]. Ignited [2681]. Ignorance [2872, 825, 3614]. Igor [2087]. Ihde [240]. ihre [3515, 629, 3405, 1944]. ihrem [162]. Ihsanoglu [583]. II [2882, 1787, 3390, 1043, 646, 2839, 1417, 2858, 1256, 2355, 2245, 2315, 1747, 533, 1131, 1421, 91]. III [2164, 3390, 1958, 1416]. III-II [3390]. Ikos [2867]. Ilana [150]. Ildikó [3747, 1912]. Ileana [2277]. Ilhsanoglu [583]. Iliaster [769]. Iliffe [1154]. Ill [2275, 1479]. Ill-fated [2275]. Illich [2110]. Illies [3012]. Illness [3192, 3686, 1854]. Illuminating [1021]. Illumination [1255, 871]. Illusions [3278]. Illustraciones [2386]. Illustrated [69, 3557, 1210, 2084, 185, 264]. Illustration [3739, 2043, 3549].

Illustrations [2583, 2051]. Illustrative [2155]. Illy [3410, 2751, 68]. Ilse [62]. ilustración [1976, 449, 1602, 2264, 1242]. Ilustrada [768]. ilustrado [2351]. Im [630, 704, 2143, 2117, 3101, 1856, 1052, 164, 346, 2184, 758, 3679, 3479, 1089, 2997, 3068, 474, 3402, 188, 337, 2485, 360, 3486, 171, 126, 264, 476, 1386, 56, 635, 3418, 1964, 2211, 909]. Image [1440, 1637, 516, 1379, 3232, 1145, 3146, 355, 106]. Image-Making [3232]. imagen [2742, 449]. Imagery [1730, 3632, 2533]. Images [1991, 910, 1867, 14, 118, 2843, 174, 2826, 3427, 3632, 2773, 686, 494, 3566, 1691, 2824, 946, 412]. Imaginaire [3521]. Imaginary [439]. Imagination [3601, 2556, 2054, 568, 1245, 1286, 246, 270, 955, 2079, 167, 259, 3328, 2757, 2222, 2244]. Imagined [763]. Imaging [1376, 2785]. Imagining [3559, 2570, 2256, 1232]. Imbeciles [3775]. Imitations [2025]. immagini [877, 1483]. Immanuel [160]. Immigrant [2060]. Immunity [2924, 3554]. Immunology [555, 3554]. Impact [650, 1762, 690, 3387, 2665, 2708, 2812, 3225, 1367, 3074, 923]. Impediments [3366]. Imperatives [3132, 646]. Imperatorskoi [1354]. Imperfect [868]. Imperial [3002, 1484, 2259, 2288, 638, 2655, 3271, 1117, 695, 1521, 2494, 854, 1101, 1126, 2423, 3746]. Imperiale [2044]. Imperialism [348, 3261, 3678, 3074]. imperio [884]. Implantable [1155]. Implications [2657]. Importation [4]. Imprese [2842]. Impressions [2544]. Improbable [1131]. Improve [3231]. Improvement [1101]. Impure [2092]. Imre [745]. Inc. [434]. inchiesta [3161]. Inchofer [3150]. Incident [3085]. Including [745, 3150]. Inclusion [3346]. Inclusive [1312]. Incognita [3729]. inconnu [1190]. Inconvenient [1045]. Incorporated [89]. Incorporating [45]. incubato [1103]. Independence [453]. Index [1330, 1789, 2227, 2513, 2793, 3113, 3444, 3787, 2472]. India [2898, 3553, 348, 985, 3247, 3678, 3422, 477, 2677, 2292, 453, 2626, 1595, 246, 650]. Indian [1851, 89, 534, 806, 645, 1403, 1091, 3644, 3247]. Indians [851, 992]. Indies [3565]. Indigenous [933, 3049]. individual [255]. individualisme [1936]. Individualität [1371]. Indo [1404, 2979]. Indo-European [1404]. Indo-Ouest [2979]. indochinoise [183]. Indomitable [1959]. Indonesia [1795]. Induced [1185]. Induction [3029]. Inductive [2707]. Industrial [3324, 2060, 845, 179, 134, 2430, 315, 1343, 2356, 2891, 1733, 741, 988, 2181, 2005, 2633, 431, 2430, 272]. Industrialisation [22, 3475]. Industrialisierungsgeschichte [1402]. Industrialization [3051, 291]. Industrialized [2427]. Industrializing [131]. Industrie [2117, 1111, 1352]. industrielle [2088, 1610, 2721]. industrielles [1352]. Industries [2766, 741, 2633]. Industry [1951, 1143, 653, 2314, 305, 2980, 1343, 2657, 576, 3618, 2218, 895, 396, 537, 3000, 844, 1545, 3193]. inédite [2277]. inédites [2979]. Inequality [2861, 3216]. Inerrantist [104]. Inescapable [3176]. Inevitability [3285]. Infanticide [1323]. Inference [3029]. infinie [2896]. Infinite [3087, 357, 2896]. Infinitesimal [3654]. infinito [3564]. Infinity [2729]. Infirmary [1829]. Influence [2881, 2458, 3584, 2481, 2541, 2601]. Influences [2, 427, 2517, 2796, 3116, 3447, 1266, 830, 1333, 1792, 2230].

Informatics [323]. Information [425, 428, 828, 1438, 853, 205, 1546]. Informed [2926]. Infrastructure [344]. Inge [619, 377]. Ingenieros [692]. Ingenieure [636]. ingenios [1500]. Ingenious [496]. ’inghilterra [493]. inglesi [2729]. Ingredients [2019]. Ingrid [1920, 1823]. Inheriting [566]. Initiative [317]. Initiatives [1737]. Ink [3247]. Inka [2039]. Inkster [22]. Innes [447]. Innocence [1245]. Innocents [471]. Innovation [313, 2780, 2567, 1547, 3198, 3773, 553, 585, 206, 132, 314, 2369, 3098, 1301, 1813]. Innovations [3194, 3643, 1676]. Innovators [1903]. Inquest [1728]. Inquietudes [2430]. Inquiry [3475, 882, 1886]. Inquisition [3738]. Insane [1997, 1854]. Insanity [1203, 517]. insbesondere [347]. Inscription [2457]. Insects [1028, 2059]. Insider [1296]. Insights [1730]. insignes [616]. Inspection [2060]. Inspector [2997]. Inspirations [112]. Inspired [2458]. institucional [2495]. Institut [2763, 1052]. Institute [3507, 1626, 3186, 3345, 756, 1414, 2994, 3641, 1739, 1069, 3455, 2002, 931, 1274]. Instituten [2838, 2116]. Institutes [3261, 1745]. Institution [1636, 1012, 1077, 2478]. Institutional [3555, 1795]. Institutionalization [3277]. Institutions [3261, 2647, 8, 517, 2478, 3321]. Instituto [732]. Instruire [1735]. Instrument [2221, 3259, 3659, 2730, 1386, 1681]. Instrument-Experiment [1681]. Instrumental [1795, 2727, 1259]. Instrumentation [2883]. Instrumentenbauer [2997]. Instrumentos [2995, 2133]. Instruments [2025, 3324, 2883, 2727, 981, 2158, 2866, 1259, 3441, 2826, 2581, 159, 686, 3196, 2942, 2448, 1446, 133, 1196, 2064, 3644]. insulin [3200]. Insurance [1127]. Insurrection [1282]. Integrating [770, 3532]. Integration [2536]. Integrity [545]. Intellect [402, 996, 3197]. Intellectual [2112, 2210, 3281, 2870, 588, 2294, 178, 3154, 1369, 536, 2765, 2541, 3119, 1025, 520]. intellectuel [1006]. intellectuelles [1691]. intellettuale [495]. Intelligence [1870, 3716, 2568, 3609, 3700, 3216, 1426]. Intelligent [3295, 3712, 3084, 3425]. Intelligibility [3217]. Intended [1423]. Inter [2478]. Inter-American [2478]. Interaction [2665, 2708, 2812, 3225, 3074]. Interactions [3381]. Interest [1148, 598]. Interests [3610]. Interfaces [3222]. Interfere [299]. Interferon [2860]. Intermolecular [1672]. internacionais [2561]. Internacional [2148, 3174]. Internalization [1649]. International [1327, 1170, 1778, 1997, 3100, 2431, 1737, 317, 651, 1431, 3452, 3076, 1933, 385, 22, 361, 2713, 3456, 2685]. internationale [150]. Internationalism [833, 3694]. Internationalität [1385]. Internationalization [1738]. Internationalizing [2937]. Internet [2118, 3521, 602, 523, 3718]. Interplay [9, 2642]. Interpretation [1215, 488, 1095, 2290]. Interpretations [3196]. Interpretationstendenzen [1371]. interpretazione [1483]. Interpretazioni [610]. Interpreting [2095, 73, 529]. Interrelations [770]. Interrogating [1732]. Interview [932]. Interwar [2866, 1011, 2002]. intestinorum [2543]. intestinorum/La [2543]. intestins [2543]. Intimate [3061, 2656]. introduccion [165, 2742]. Introduction [2800, 2715, 2868, 2649, 378, 1583, 2125, 3021, 3094, 2879, 12, 3143, 3451, 3284, 2584, 1013, 2235, 3621, 3041, 1663, 3538, 3512, 3696, 3224, 264, 451, 437, 2520, 1552, 411, 3135]. Introduzione [994]. Intuition [2056]. Intuitive [2635]. Inuit [288, 245]. Invalidism [2424]. Invariance [3095]. Invasion [284, 815]. Invented [2775, 3773]. Inventible [1312]. Inventing [1532, 2030, 3072, 2734, 860, 3049, 1437, 1761, 523, 1680]. Invention [702, 1762, 3256, 1050, 3475, 3699, 1865, 3065, 2020, 3773, 297, 685, 1710, 3310, 2178, 1625]. Inventions [1501, 657, 3147, 1142]. Inventor [1865, 546]. inventores [1952]. Investigación [2148]. Investigation [1961]. Investigations [2068, 343]. Investigative [2327]. invisible [494, 1419]. Inviting [1305]. Involvement [2697]. Inwardness [487]. Ioannis [188]. Ionian [3230]. Iowa [1414]. Ira [2632]. iranien [1225]. Ireland [3162, 3760, 2149, 97, 1928, 1743]. Irenaus [55]. Irene [2697, 2041]. Iris [3410, 311, 2326]. Irish [1903]. Iron [628]. Ironies [735]. Ironworks [628]. Irony [2310]. Irregular [2048]. Irresistible [1567]. Irritability [2746]. Irritable [2081]. Irritating [2746]. Irvine [773]. Irving [51, 1403]. Isaac [1952, 1507, 2163, 2606, 2979, 1924, 3245, 1373, 1409, 1701, 271, 876]. Isabel [255, 479, 3074]. Isabelle [3157, 3483, 1593, 1050]. Isagoge [352]. ISIS [2793, 2137, 1331, 1330, 1789, 2227, 3113, 2513, 2807, 3126, 3129, 3444, 3788, 3787, 3449, 3454, 3453, 954, 1791]. Islam [1786, 1988, 3553, 3479, 264]. Islamic [1172, 2883, 1689, 3646, 3043, 585, 1969, 3486, 1367, 2315, 3236]. islamique [1225]. islamischen [3486]. Island [703, 288, 2702]. Islands [849, 2984]. Ismos [2867]. Isobel [1018]. Israel [720, 1649]. Israeli [2082]. Issitt [3258]. issledovanija [2943]. Issue [596]. Issues [1197, 1791, 2861, 2447, 2563]. istochniki [2943, 2262]. istokov [2191]. istorii [2262, 2499]. Istorija [1291]. istoriografija [2943]. István [2472, 1417]. Italia [3319, 2267, 1381, 2977]. Italian [873, 1173, 702, 1600, 3141, 1695, 2996]. italiana [1553, 2267, 2504, 2986]. italiani [2363]. Italiano [3728, 2842]. Italians [152]. italo [2401]. italo-ungheresi [2401]. Italy [2920, 3238, 2152, 945, 3574]. itinéraire [2426]. Itineraries [1446]. Itinerários [2166]. Itinerarium [3406]. Itinerarius [1494]. Itinerary [2063]. Itself [3440]. Itty [534]. Itzkoff [229]. IUPAC [70]. Iuskevich [1291]. Ivan [2110, 3524]. Ivanovich [915]. Ivor [2722]. Ivory [1776]. Iwan [2720, 367, 2627]. IX [1054].

J [1309, 1175, 2118, 2944, 2938, 849, 2912, 446, 783, 2027, 1081, 92, 565, 2097, 2985, 2086, 742, 287, 2663, 1359, 2728, 193, 2720, 3735, 707, 510, 366, 1380, 153, 396, 90, 3638, 266, 83, 1078, 1527, 1821, 797, 542, 374, 981, 2073, 370, 305, 3093, 2817, 2692, 786, 2410, 1936, 1124, 3472, 932, 922, 3077, 3543, 192, 3257, 416, 1048, 2762, 1954, 990, 571, 2731, 2373, 2310, 373, 2751, 2743, 2409, 552, 2618, 240, 2139, 2176, 2925, 68, 420, 116, 1594, 591, 1086, 3355, 1884, 2494, 2274, 272, 2216, 730, 443, 597, 3004, 1346, 1628, 439]. J [462, 1747, 2529, 3751]. J. [1840, 1162, 192, 1575, 928, 2442]. J.-Matthias [707]. Jaap [2608]. Jacalyn [451, 2815]. Jack [458, 459, 2555, 2624, 1817]. Jackson [175, 3259, 2928, 2708, 1609, 2908, 3473, 179, 554, 258]. Jacob [31, 2980, 2830, 196, 2965, 988, 2698]. Jacobs [2676]. Jacobson [735].

Jacopo [3102]. Jacqueline [211]. Jacques [3105, 44, 2091, 2671, 2820, 2850]. Jacyna [1116, 3031, 3772]. Jagadis [645]. Jagdish [3678, 75]. Jahncke [769]. Jahncke. [3035]. Jahrbuch [1644]. Jahre [619]. Jahren [3715, 48]. Jahrestagung [1178]. Jahrhundert [3101, 2997, 3068, 337, 126, 264, 3679, 3402, 2485, 2184, 758, 347, 474, 3477]. Jahrhundert-Legitimation [3101]. Jahrhunderts [164, 3227, 3500, 2607, 3761, 3260, 1385, 1279, 67, 1920, 2276, 171, 56]. Jake [560]. Jakob [483, 3178, 3747]. James [2989, 738, 890, 418, 731, 1349, 634, 485, 643, 30, 378, 50, 1077, 349, 1095, 169, 3641, 3001, 3050, 248, 651, 320, 376, 466, 2925, 1932, 437, 801, 913, 508, 262, 2054, 1000, 959, 2164, 2924, 2978, 3699, 1883, 370, 2774, 1002, 1395, 1161, 3088, 325, 3062, 1836, 61, 1263, 90, 168, 3050, 1272]. Jamie [1921]. Jamison [2964, 1879, 536]. Jammer [648]. Jan [167, 843, 3646, 3402, 796, 1072, 1091, 132, 3297, 1021, 1828, 2183, 2246]. Janaab [3307]. Jane [3507, 774, 993, 2300, 2930, 1360, 3591]. Janes [785]. Janet [1921, 523, 1720, 2225]. Janine [2560]. Janis [2350]. Janković [989]. János [2618]. Jansen [2370]. Janssen [2751, 68]. January [239, 837, 3340, 3457, 370, 371, 2806, 754]. Janus [145]. Japan [2767, 2839, 1885, 1377, 979, 815, 1406, 310, 3007, 1579, 2699, 3351]. Japanese [3261, 2836]. Jarcho [620]. Jardine [2159, 623, 496]. Jason [1132]. Jaume [1820]. jaune [1736]. Java [1416, 1795]. Javier [1015]. Jay [3083, 319, 1455, 2576, 2764]. Jean [761, 2168, 1084, 1141, 133, 1509, 354, 469, 3232, 32, 3060, 561, 2915, 1084, 2202, 1607, 2896, 2671, 2733, 2991, 3075, 1087, 3652]. Jean-Daniel [3075]. Jean-Francois [2202]. Jean-Marc [2671]. Jean-Patrice [354]. Jean-Paul [1084]. Jean-Pierre [1344, 2915]. Jean-Vincent [2733]. Jeanmonod [2091]. Jeanne [666, 3750, 2704, 1961]. Jedaiah [1583]. Jeff [654, 2658, 1100, 3051, 2777]. Jefferson [1929, 992]. Jeffrey [1155, 289, 81, 930, 2217, 2902, 1966, 2379]. Jeffries [3340, 371]. Jena [52, 1934]. Jena-Weimar [1934]. Jenkin [1950]. Jennifer [557, 2562, 1355, 2358]. Jenny [2638]. Jensen [1870]. Jepsen [852]. Jérôme [3572]. Jeremiah [36, 3258]. Jeremy [3325, 14, 2187, 3257, 2055, 1115, 2618]. Jerez [449, 1602]. Jerne [1960]. Jeronimo [146]. Jerry [641, 2432]. Jesper [2843]. Jesse [2899, 3659]. Jessica [77, 211, 2344, 1040]. Jesuit [1633, 2477, 2605, 2238, 1640]. Jésuite [613]. jésuites [2733]. Jesuitische [758]. Jesuits [446]. Jesús [2444]. Jevons [2756]. Jew [3551, 2380]. Jewel [3233]. Jewish [2989, 738, 1401, 475, 520, 3414, 7, 3251, 699]. Jews [787, 1477]. Jha [1428]. ji [3561, 3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277, 3261]. Jian [3268, 3277]. jianbian [3306]. Jiang [3343, 979, 3275]. Jianping [3261]. jiaoyu [3306]. Jie [3200]. Jifeng [3262]. Jim [269, 2159]. Jiménez [1394]. jin [3561, 3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277, 3219]. jindai [3306]. jing [3726, 3200, 2533]. Jingsheng [3186]. jiu [3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277].

Jiuchen [3333]. Jo [254, 2529]. Joachim [1178, 144, 95, 769, 630]. Joan [2656, 985]. Joan-Pau [985]. Joanne [2443]. Joao [129]. Jobs [1680]. Jobst [2056]. Joe [467, 421]. Joel [3524, 3090, 2179]. Joëlle [614]. Johach [3672]. Johan [567, 1795]. Johann [1093, 619, 42]. Johanna [67, 1842]. Johannes [1494, 144, 2895, 147, 2486, 708, 2979, 587, 974, 1349, 1956, 1995, 485, 147, 3747]. John [1025, 1263, 1309, 943, 1975, 800, 446, 1349, 632, 221, 436, 223, 2270, 287, 179, 3083, 3338, 181, 452, 690, 2418, 1853, 69, 125, 82, 683, 616, 803, 756, 3757, 507, 186, 653, 96, 1852, 703, 797, 2437, 527, 2003, 3777, 2345, 294, 2658, 3652, 227, 185, 722, 1119, 265, 1145, 657, 371, 3189, 669, 2731, 399, 3758, 312, 772, 63, 154, 245, 600, 538, 84, 279, 2564, 43, 3258, 2309, 730, 700, 1246, 276, 2556, 2307, 2429, 907, 966, 1626, 2449, 2085, 1009, 1867, 1163, 2042, 2968, 2026, 863, 1340, 1164, 1363, 2464, 2712, 1154]. John [1588, 1970, 2624, 1037, 1817, 2497, 2708, 2812, 2203, 986, 236, 2581, 2759, 2365, 1605, 2096, 2313, 2511, 1556, 1851, 3393, 3403]. Johns [109]. Johnson [2880, 2941, 313, 306]. Johnston [1083]. Joined [2176]. Joint [3448]. Jole [2541]. Jolivet [469]. Jon [349, 1697, 492, 2438, 1621, 2293]. Jónas [2200]. Jonathan [521, 299, 805, 2744, 898, 2094, 2907, 1367]. Jones [924, 1105, 2981, 3067, 111, 2188, 2909, 321, 755]. Jonkers [1824]. Joost [2881]. Joost-Gaugier [2881]. Jordan [2858]. Jordanova [3324, 594]. Jörg [3013]. Jorge [165]. Jörn [2407]. Jose [300, 1820, 1868, 3140, 2938, 449, 3054, 2829, 2404, 2386, 2785, 2142]. Josef [631]. Joseph [2989, 1974, 1030, 2419, 1896, 1612, 3340, 168, 1194, 517, 3241, 1281, 765, 416, 371, 86, 1938, 3298, 3746, 43, 2482, 2946, 1958, 1167, 1188, 2367, 2705, 3626, 2136, 3151]. Josephson [2427, 2637]. Joshua [2631, 955]. Jost [74]. Journal [596, 1689, 3740, 1717]. Journals [3448, 51, 288, 1614, 174, 1403, 779]. Journey [2420, 2783, 1823, 477, 3228, 1548, 1825, 1248]. Journeys [2180]. jours [1352]. Joy [2714, 564, 1027]. Joyce [1935, 3258, 3048, 1927]. Jozsef [68, 2751, 3410]. JPL [3093]. Jr [1009, 2615, 3057, 1083, 2708, 1619, 2923, 1605, 1811, 2465, 2633]. Jr. [181]. Juan [714, 1500, 3140, 2264]. Juanelo [1500]. Judaism [3551, 1507, 1345]. Judgement [3782]. Judith [3170, 1745, 3305, 124, 1548, 2161, 2778, 2747]. Judy [3750]. juge [3572]. Jules [48]. Julian [1026, 2926]. Julie [1959, 2247, 2745]. Juliette [518]. Julius [2334, 2725]. Jullien [3159]. July [3125]. Julyan [903]. June [239, 991, 687]. Jung [2301, 2433]. jungen [162]. Jungles [1795]. Junker [1944, 199, 2553]. Junkie [2168]. Jupiter [682]. Jürgen [2485, 3512]. Jurine [627]. Jury [2902]. Juschkewitsch [1291]. Justice [1641, 184, 1302, 1047, 656].

Kaasch [1178]. kabbale [1310]. Kadletz [476]. Kaeser [2620]. Kahn [3311, 2635]. Kaibab [1550]. Kain [443]. Kaiser [2117, 2723, 2565, 2838, 2763, 2116, 3515]. Kaiser-Wilhelm-Gesellschaft [3515]. Kaiser-Wilhelm-Gesellschaft-Wissenschaft [2117]. Kaiser-Wilhelm-Institut [2763]. Kaiser-Wilhelm-Instituten [2838, 2116]. Kaiserreich [1254, 1017]. Kaji [1384]. Kalb [1621]. Kalendarium [265]. Kaluza [3605]. Kamal [1089]. Kamerlingh [1842]. Kamil [3131]. Kampfplatz [3177]. Kanigel [395]. Kant [944, 160, 1473, 667, 1597, 2537, 3578, 1247]. Kantian [2994]. Kaplan [684]. Karen [2989, 3101, 210, 92, 565, 168, 1140, 3257, 3232, 399, 1843, 2305]. Karen-Beth [399]. Kargon [3773]. Karin [2485, 617]. Karine [2822]. Karl [2189, 401, 3500, 2070, 1714, 2657, 476, 2199, 1011, 1846, 1878, 1045, 719]. Karlsruhe [95]. Karsten [170]. kartografische [2184]. Kary [281]. Kassell [3042]. Kassler [1921]. Kastner [2189]. Katalog [95]. Kate [1412]. Katharine [2748]. Katherine [323, 2776, 2261]. Kathleen [3234, 3340, 321, 371, 1832, 1131]. Kathryn [524]. Kaufman [562]. Kavita [2626]. Kay [2700, 1560, 1831]. Kealey [260]. Kean [1382]. Keating [2299, 2721]. Kedi [3200]. Keel [3321]. Keen [2841]. Keep [3683]. Kehoe [117]. Keil [2563, 619]. Keil-Slawik [2563]. Keith [720, 2072, 689, 2446, 2139]. Kelleher [57]. Keller [929, 3736]. Kellert [3529]. Kelley [108]. Kelly [365]. Kelvin [3065]. Kemilainen [120]. Kemp [1882]. Ken [3044, 2666, 1863]. Kendrick [1129]. Kennedy [822, 705, 3350, 2382, 446]. Kennefick [2751]. Kenneth [3529, 3497, 391, 597, 2767, 1095]. Kent [2445]. Kenyon [1105]. Kenyon-Jones [1105]. Kepler [1995, 1349, 485, 708, 2595, 267, 1217, 1503]. Keplero [3395]. Kern [2597]. Kernenergie [525]. Kerstin [480]. Keuren [2130]. Kevin [541, 661, 187, 2268]. Kevles [2566]. Key [2944, 3123, 3480, 3632]. Keynes [1601, 724]. Keys [1025]. kh [1267]. Khalili [2315]. Khipu [2039]. Kieckhefer [37]. Kienzler [52]. Kilgore [2032]. Kill [3768]. Killed [1115]. Killen [2910]. Killer [1303]. Killing [810]. Kim [2349, 870, 2287, 3002, 2789, 1835]. Kime [51]. Kin [855, 667]. Kind [2259, 766, 902]. Kindî. [469]. Kindly [181]. Kindred [1105, 285]. Kinds [2705]. King [2883, 2684, 3650, 468, 3087, 585, 486, 2627, 608, 394]. Kingdom [1529, 1240, 1209, 1678]. Kings [3408]. Kingsland [2755]. Kingsley [1825]. Kino [3500]. Kinsey [898, 1118]. Kinship [1134]. Kinzelbach [1178]. Kinzua [564]. Kirby [2520]. Kircher [1823, 3406]. Kiriwina [400]. Kirk [1155]. Kirsanov [3371]. Kirsch [919]. Kirschke [2189]. Kirschner [2607]. Kirsten [2255, 2913]. Kirsti [3467]. Kiss [86]. Kistemaker [31]. Kitāb [1364]. Kitchen [559, 1130]. Kitcher [2510]. Kite [1841]. Kitty [122]. Klaas [592]. Klangmaschinen [3330]. Klassifizieren [3715]. Klassiker [3331]. Klaus [609, 790, 256, 2182, 170, 68, 54, 2840, 2997, 1386]. Klaus-Dietrich [790]. Klaus-Metrich [609]. Kleber [335]. Klee [411]. Kleeberg [3063]. Klein [2428, 3322]. kleinkinderen [3766]. Kleinman [2092]. Kline [2371]. Klio [3260]. Kneeland [2190]. Knell [1081, 510]. Knew [731]. Knight [3158]. Knobloch [707]. Knoeff [2487]. Knorr [578, 1443]. Knots [1671]. Knotted [2039, 253]. Knotted-String [2039]. Know [1723]. Knowing [2895, 1163, 2950, 3527, 2409, 1724]. Knowledg [2735]. Knowledge [942, 1311, 1064, 1024, 2767, 3222, 853, 2983, 3662, 1450, 108, 3539, 871, 3390, 2334, 109, 1668, 3365, 2461, 3502, 3283, 2448, 3752, 1355, 3094, 2727, 575, 934, 102, 3655, 2341, 2478, 987, 1955, 2872, 2604, 3718, 229, 269, 2486, 1755, 2156, 28, 674, 2935, 3244, 331, 63, 834, 2131, 996, 3190, 3223, 1682, 1879, 3634, 3081, 3473, 2440, 709, 3531, 1191, 3049, 2507, 1836, 2033, 883, 2138, 2533, 2575, 2491, 2101, 85, 933, 187, 570, 415, 2450, 1662, 552, 3688, 3176, 3780]. Knowledges [2952]. Known [3231, 3229]. Knows [1624]. Knox [2559]. Koblitz [920]. Kockelmans [416]. Kocku [2819]. Koertge [2577]. Koestler [1159]. Kohler [2998, 2067]. Kohlhoff [2212]. Kohlstedt [595]. Kojevnikov [2779]. Kokowski [2536]. Kolchinsky [2943]. Kolleginnen [3515]. Kollerstrom [2417]. Kolmogorov [1617]. Kolta [1089]. Koltsova [91]. kommentierte [3747]. kommentierter [34]. Kommerz [500]. Kommunikationssystems [814]. Konashev [2943]. konce [2943]. Konflikt [3418]. Königsberg [1230]. Konjunkturen [2855]. Konkreten [3013]. Konrad [2615, 2182, 3536]. Konsonanzwertung [1584]. Konstgenoten [499]. Konstrukts [2370]. Kontext [704, 162]. Kontingenz [144]. Kontinuitat [162]. Kontinuitäten [3761]. Kontroverse [3406]. Konzepte [188]. Koplow [1962]. Korallen [3743]. Korea [652, 979]. Kormos [728, 3410, 2751]. Korrespondenz [1827, 1291, 164]. Kosmologie [1470]. kosmologische [3406]. Kox [2751, 68]. Koyré [2202]. Krafft [1120, 1574, 48]. Krafft-Ebing [1120]. Kragh [2621, 1407]. Krankenhaus [2616]. Krankheit [635]. Kranksheitsvorstellung [1089]. Krau [1279]. Kraut [1859]. Krebs [941, 338]. Krebszelle [3672]. Krech [563]. Kreiseltechnik [2056]. Krementsov [2685]. Kremer [743]. Kremer-Marietti [743]. Kren [3457]. Krieg [2284]. Krige [3189]. Krise [636]. Krisenbewältigung [2284]. Kristall [3645]. Kristen [3681]. Kritik [3101]. Kroge [917]. Kruk [3646]. Kues [1371]. Kühlmann [769]. Kuhn [3687, 664, 2742, 1049, 190, 1635, 1199]. Kühne [2607]. Kukla [2100]. Kuletz [696]. Kultur [3436, 337, 126]. kultureller [3177]. Kulturgeschichte [3058, 1910, 126]. Kulturkritik [635]. Kulturvergleich [3479]. Kulturwissenschaft [3672]. Kunitzsch [705]. Kunkel [347]. Kunstformen [55]. Künstler [3239]. Kunstreligion [2831]. Kuper [202, 119]. Kupper [2855]. Kurian [2472]. Kurt [15, 71, 3731, 476, 1291, 1799, 71]. Kurt-Reinhard [1799]. Kuru [3764]. Kusch [2138]. Kushner [1076]. Kwa [2724]. Kyle [2936].

L [1061, 1840, 2075, 44, 2364, 1349, 2060, 1888, 1081, 884, 1012, 658, 2881, 205, 2873, 2981, 3144, 977, 2402, 892, 3638, 3642, 2275, 2809, 683, 83, 616, 50, 756, 147, 2464, 2850, 696, 2172, 859, 3699, 3095, 2206, 2562, 195, 1135, 370, 603, 1963, 166, 851, 294, 478, 473, 1973, 2932, 804, 1099, 586, 2641, 3050, 722, 65, 2762, 2835, 397, 2563, 3497, 2759, 2577, 2376, 2302, 772, 641, 1811, 590, 81, 2690, 279, 3007, 2201, 1007, 3145, 1548, 3772, 3084, 544, 3653, 2782, 821, 264]. l-Fadl [264]. L. [1116]. Lab [2067, 3367, 1001, 3178]. Labeling [1943]. labiles [3495]. Labor [2060, 1135, 1533, 1512, 3193, 1856]. laboratoire [1111]. Laboratoires [716]. laboratori [1465]. Laboratorien [3254].

Laboratories [2752, 2853, 314, 845, 3595]. Laboratory [205, 2632, 3502, 3370, 3369, 3368, 2686, 278, 1378, 2355, 1051, 2290, 2694, 1424, 2764, 2923]. Labour [2694]. Labouvie [126]. Labs [2559, 1954]. Labscapes [2067]. l’Académie [3152]. Lachieze [1344]. Lachieze-Rey [1344]. Ladies [2355, 278]. LaDuke [3750]. Lady [3514]. L’affaire [3267]. Lagier [2537]. Lagrange [1235]. Lai [589]. Laid [2811]. Laidler [2139]. Laird [1917]. Lakatos [745]. Lakatos-Feyerabend [745]. Lake [2966, 331, 543, 730, 662]. Lakes [1716]. Lakoff [3190]. Lalande [3572]. L’alchimie [701]. l’alimentation [1607]. Lamarck [3060]. Lambright [84]. Lammel [3260]. Lamotius [2979]. Lamotius/Isaac [2979]. Lamp [591]. Lamphere [209]. Lamps [797]. Lana [127, 2257]. Lance [231]. Land [117, 1834, 2569, 2760, 913, 3433, 1844]. Landau [553]. Landes [2295]. Landesbibliothek [95]. Landings [3168]. Landmark [2722]. Landmarks [961]. Lands [3559]. Landscape [2898, 3336, 916, 373, 1287, 868]. Landscapes [1437, 1589, 3584, 2067, 3220]. Landschaften [2184]. Landschaftsveraenderung [2308]. Lane [214, 2186]. Lang [2147]. Langermann [787]. Langins [2350]. Langley [2217]. Language [2945, 2223, 3623, 1116, 3392, 2976, 516, 442, 813, 1836, 2613, 1436, 1670, 3490, 3676]. Languages [3490, 2608]. langue [2278]. Lankford [279]. l’Antiquité [2721]. lapidum [3483]. Laplace [2426]. LaPorte [2705]. Lapsley [1134]. Larabee [1430]. Large [66, 2330, 2712, 503]. Large-Wavelength [2330]. Largo [1705]. Larry [2980, 1427]. Larson [849, 2911]. Larsson [677]. L’art [2337]. Laser [1410, 2777]. Lashley [401]. Last [1504, 1264, 157, 2214, 275, 1938, 116]. Lastanosa [1500]. Late [1484, 2933, 1996, 3235, 862, 1704, 5, 2259, 818, 3415, 3475, 695, 2457, 865, 1521, 1378, 2043, 2299, 3708, 2310, 1369, 1449, 1684, 854, 807, 3306, 1068, 2655, 1366]. Late-Twentieth-Century [2299]. lateinischen [363]. Later [2882, 615, 897, 276]. Latin [2105, 1982, 3380, 2156, 3140, 1364]. Latins [3388, 790]. Latour [138, 2313]. L’aube [1488]. Laubichler [3507]. Laue [74]. Laugwitz [799]. Launius [1309, 84]. Laura [284, 1931, 1096, 3004]. Laureate [2989, 2380]. Laureates [2472, 325]. Lauren [3042]. Laurence [183, 210, 2375]. Laurent [1376]. Laurie [488]. Laurioux [2823, 33]. Lauxtermann [1597]. L’avènement [3321]. Lavoisier [2267, 1376, 364, 1508]. Law [1484, 2109, 3098, 975, 2869, 1590, 7, 1773, 802, 1384, 693, 665, 603]. Lawless [254]. Lawrence [3158, 2061, 3315, 2298, 21, 2686, 2856, 2327, 1652, 1623, 3031, 2068, 1591]. Laws [3358, 2348, 941, 260, 1080, 2348, 417, 1981]. Laylin [325]. Lazardzig [3402]. Le-May [3225]. Lead [737, 1758, 3073, 2581]. Lead-Paint [1758]. Leader [859, 806, 2359]. Leaders [1056]. Leadership [741]. Leading [3659, 1079]. Leaf [932]. Leah [2023]. Leaps [2968]. Lear [2999]. Learned [207, 971, 2603, 1795]. Learning [1180, 1311, 949, 950, 210, 2752, 706, 886, 2745, 591, 3552, 1858, 1180, 949, 950, 3487]. Leary [389]. Leavitt [124]. Leben [1004, 3605]. Lebenden [2372]. Lebensläufe [1254].

Lebenswissenschaften [2846]. Lebenswissenschaftlers [1284]. Lebenswunder [282]. Lebrun [2687]. Leclerc [44]. L’école [2921]. Lectures [2114, 745, 2041, 92, 365, 1257]. Lécuyer [2780]. Led [3085]. Lederberg [955]. Lederer [1568]. Lee [2078, 375, 90, 1852, 2092, 1811, 46, 2878]. Lefèvre [2538]. Lefèvre. [3322]. Left [225]. Legacies [3349, 3431, 1606]. Legacy [402, 1725, 1679, 1946, 3590, 2994, 3078, 204, 2636, 3581, 350, 3603, 730, 2112]. Legal [2732, 603]. Legend [3732, 1275, 1527, 590, 245]. Legends [1690]. legibus [492]. Legierungen [3763]. Legitimation [3101]. Legros [1084]. Legs [3069, 1763]. L’Egypte [335]. Lehner [2751]. Lehrach [525]. Lehre [144, 3068]. Lehrer [2189]. Leibe [346]. Leibniz [944, 3103, 2732, 2981, 707, 3654, 162, 2671, 2265, 2608, 497, 618, 41]. Leibnizsche [3657]. Leidensia [3634]. Leigh [781]. Leisured [1107]. Lemaire [83]. Lemelson [3773]. Lemmerich [74]. Lemuria [2760]. Len [810]. l’Encyclopédie [3348, 3314]. l’Enginyería [2430]. Lens [2646]. L’enseignement [3639]. Lenz [1964]. Leo [1849]. Leon [873, 2922, 1696, 2500]. Leonard [372, 2421]. Leonardo [2159, 1667, 2118]. Leonhard [3573]. Leonhart [616]. Leopard [3018]. Leopold [170, 170]. Lerner [1855, 582, 736, 582]. Lesassier [2074]. Leshem [2488]. Leslie [3000, 1038]. L’espistemologia [1517]. l’esprit [1400]. Lessons [1939, 2241, 1305]. Lester [632, 3005]. Lestienne [137]. Let [307]. L’État [1710]. Letter [3290, 3289, 1189, 3291, 3214, 3292, 334, 3127, 1073, 3128, 3213]. Letters [3364, 2874, 2962, 3706, 105, 3630, 3544, 3704, 3372, 3745, 3461, 2719, 3464, 241, 3027, 370, 3459, 3462, 3373, 1513, 2301, 3629, 174, 2057, 3628, 3460, 3463, 2482, 3705, 921, 168, 288, 1640, 1614, 1916]. lettore [1701]. lettori [2022]. Lettres [3572]. Letture [156]. Leu [3736]. leur [3248]. L’Europe [1350]. Levelt [1842]. Levere [1358]. Leverington [1033, 2140]. Leviathan [3499]. Levin [2622, 520]. Levins [817]. Levit [915]. Levitt [460]. Levy [2141, 186, 923]. Lewenstein [595]. Lewin [1416]. Lewis [911, 981, 218, 3490, 1118, 2055, 856, 1081, 756, 588, 342]. Lexicon [1947]. Lexikon [617]. l’expression [2905]. Leyden [1250]. Leyenda [3732]. L’herbier [3571]. L’histoire [2192]. L’homme [142, 3240]. Li [2442, 3432, 3275, 3261, 3277]. Liaison [3260]. Liang [3261, 3561, 3262]. Liba [3196, 2150]. Libavius [3567]. Libellus [480]. Liber [472, 3483, 2729]. Liberal [3664, 3604]. Liberalism [3329, 2285, 2083]. Librandi [35]. Libraries [1631]. Library [940, 3496, 3736, 1355, 2187, 47, 3298, 1054, 1057, 3483, 3736]. libro [146, 1465]. Libros [586, 1696, 1500]. licheng [3306]. Licht [3715]. Lichtbildarchive [3500]. Lichtenberg [363]. Lichter [1577]. Lichtmann [3634]. Licklider [1628]. Lie [1258, 3294]. Liebersohn [3064]. Liebigs [2189]. Lienhard [2897]. Lieserl [1387]. lieu [1593]. Life [2697, 1778, 1445, 1636, 1978, 3719, 44, 247, 2419, 1018, 1021, 2681, 1349, 644, 3142, 1931, 3166, 1523, 3018, 1848, 1660, 884, 1859, 1157, 3231, 2500, 2981, 61, 193, 452, 1858, 1674, 168, 2435, 2159, 2999, 803, 1527, 3255, 2918, 507, 2295, 520, 2169, 2764, 2298, 3153, 898, 2639, 2841, 3707, 384, 79, 527, 1614, 1118, 1421, 713, 904, 2074, 515, 2296, 2199, 3131, 1960, 2300, 3440, 178, 1418, 2641, 3170, 139, 1575, 1959, 3652, 2458, 2601, 3751, 3078, 2617, 1765, 3594, 2750, 3708, 1104, 2586, 2384, 3411, 3006, 1079, 2326, 2775, 802, 116, 87, 1396]. Life [535, 3165, 3190, 2774, 3587, 1875, 930, 1398, 2213, 3442, 3414, 623, 276, 865, 2882, 637, 2254, 861, 1392, 1129, 1411, 2752, 1874, 1604, 386, 444, 392, 1409, 209, 644, 3329, 558, 193, 3256, 1571, 3195, 3556, 3780]. Lifelong [1005]. Lifesaving [1625]. Ligaments [436]. Light [2208, 1967, 3754, 654, 871, 869, 1742, 1382, 1250, 2458, 1259, 2043, 2648, 2298]. Light-Green [2208]. Lightning [2169, 1518]. Lights [457]. Like [900, 759, 932, 900]. Likenesses [13]. Lila [1003]. Lillian [275, 711]. Lily [1560]. Liminal [2570]. Limited [314]. Limiting [1670]. Limits [406, 2781, 833, 1732, 321, 2743, 63, 3782, 2481, 3360]. Limoges [2721]. Lincoln [702]. Linda [210, 61, 66, 858, 2809, 2999, 2172, 2641]. Lindberg.d [3555]. Lindee [2771, 2857]. Lindemann [448]. Lindesmith [1025]. Lindley [3508, 3065]. Lindner [1279]. Lindqvist [677]. Line [2346, 524, 2552, 11]. Linea [912]. Linear [2657]. Lines [2933, 1469, 3137]. linguaggi [3615]. Linguerri [2842]. Linguistic [1404, 1670, 516]. Linguistics [1404, 1404]. Linhard [3657, 2194]. Lining [1035]. Link [1005, 2221]. Linking [136]. Links [1314, 3623]. Linnaeus [2061, 1510, 3238]. Linne [265, 163]. l’INPI [2088]. Linton [2531]. Linus [1748]. Lion [1924]. Lipke [2979]. líquido [2995]. Lisa [2159, 496, 208, 721, 2409, 130, 623, 2668, 2074]. Lise [3515, 1948, 74, 726]. Lissa [3475]. Listening [1615]. Lister [2623]. Listerism [641]. Lists [724]. Literacy [2827]. Literarische [2258]. Literarischen [3757]. Literary [3646, 512, 1100, 2919, 1678]. Literate [2409]. Literatur [769]. literatura [2474, 2264]. Literature [2185, 3104, 1600, 3261, 1078, 3163, 2341, 3714, 583, 1834, 445, 323, 2851, 2365, 3409, 3570, 284, 2055, 1370, 270, 2812, 1165, 454, 195, 2344, 1083, 1869, 1091, 1177, 1438, 1204, 1285, 1203, 3484]. literatürü [583, 1370]. Lithosphere [2330]. Litten [1109, 725]. littéraire [3260]. Little [970, 1620, 2197, 1044, 942]. Littleton [3158]. Littlewood [541]. Liu [3262]. Liudger [636]. Liveliest [455]. Livers [2815]. Lives [731, 644, 3220, 3276, 938, 3252, 3652, 1810, 2570, 292, 2584, 209, 189, 252, 3431]. Livestock [2225]. Living [345, 3637, 152, 285, 2925, 3425]. Livingston [658]. Livingstone [2912, 366, 2450]. livres [33, 3568]. Liwei [3261]. Llasera [494]. Lloyd [1591, 2822, 2249, 2248, 2598, 1032]. Lluís [2995]. Local [3061, 331, 3049]. Locality [1797]. localizacion [141]. Locating [2564]. locis [141]. Locke [2974, 3490, 1326]. Locomotive [396, 1295]. Lodewijk [592]. Lodi [568]. Logic [106, 784, 1582, 2603, 2791, 2617, 1349, 3594]. logica [226]. Logicae [3390]. Logical [3785]. Logics [1515]. Logie [1145]. Logique [1226]. Logos [2867]. Logsdon [1309, 84]. Logsdon.S [943]. Loïc [798]. lokale [2616]. LoLordo [3052]. Lombardo [3775]. Londa [569, 2540]. London [489, 2048, 3271, 3233, 678, 995, 1948, 3042, 2268, 2372, 1102, 588, 3659, 3253, 2730, 2159, 910, 96, 367, 2179, 1506, 1229, 1283]. Londres [3320]. Long [3234, 3559, 3676, 2033, 1687, 3639]. Longair [3009]. Longer [1582, 3637]. Longing [3700]. Longino [2101, 3529]. Longitude [3023, 2417]. Longo [2360]. Longrigg [466]. Loo [2929]. Look [3192, 2612, 2279, 618]. Looking [3044, 3622, 304]. López [1820, 449, 2386, 2142]. L’opinion [825]. L’optique [2733]. Lorch [705, 705]. Lord [3659, 2339]. L’ordre [3310]. Lords [2114, 1149]. Lore [245]. Loren [207]. Lorenz [2615, 2182, 3695, 3536]. Lorenzo [1015, 3604]. Loretta [2725]. L’oroscopo [2252]. Lorraine [782]. Lorton [3297]. Loschmidt [631]. Loss [537]. Lost [2789, 2063, 1496, 3764, 2991, 3022, 3594, 2760, 261, 3007, 1116]. Lothar [500]. Lou [3433]. Loudon [773]. Louis [3364, 44, 646, 859, 720, 3489, 1609, 1349, 1607, 2923, 627, 2177]. Louise [209, 973, 1349]. Louisiana [3756]. Louvain [2481]. Love [306, 2059, 3097, 1622]. Loved [2557, 380]. Loveland [1100]. Lovers [2815]. Loves [1968]. Low [1406, 2811, 3324, 3133, 2567, 3351, 3168]. Low-Cost [2567]. Low-Visibility [3168]. Lowell [1112]. Lowen [747, 76]. Lowenthal [723]. Lowman [558]. Lozovsky [2156]. LSD [3606]. Lubbock [3338]. Luc [3222, 2897]. Luca [2641]. Lucasian [2113]. Lucca [266]. Lucia [2977]. Luciana [691, 2736]. Lucien [513, 3348]. Lucille [44]. Luckey [843]. Luckhurst [999]. Lucky [94]. Lucretius [2382]. Lucrezia [788]. Lucrezio [1701]. Lucy [1128]. Ludger [288]. Ludmerer [391]. Ludmilla [3324, 594]. Ludwig [277, 276]. Ludy [3057]. Lugt [2343, 501]. Luigi [2173, 1170]. Luis [255, 1868, 1654]. Luiz [2279]. Luke [1566]. lumière [1702]. Lumières [3159, 3572, 3314]. Lumières. [2887]. Luminaries [1698]. Luminet [1344]. luna [2115]. Lunacy [299]. Lunar [1294, 2417, 965]. Lunbeck [2069]. Lund [709]. Lundgren [341]. L’univers [2620]. L’università [1553]. l’Université [2257]. Luo [3219]. Lure [1307]. Lurquin [2641]. Lusa [2430, 692]. l’Utérus [2257]. Luther [2143, 1291, 674]. Lutwidge [2057]. Lützen [2843]. Lydia [158]. Lydon [97]. Lyell [372]. Lyman [85, 90]. Lyn [1024]. Lynda [910]. Lynette [2908]. Lynn [459, 2982, 2007, 1151, 980, 2440]. Lynne [308]. Lynx [1703]. Lyon [389]. Lyons [218]. Lysenko [2640].

M [1309, 943, 2943, 3158, 158, 80, 2761, 401, 595, 1859, 2171, 2085, 2770, 1747, 2771, 2330, 1853, 2531, 2083, 3638, 475, 3038, 1363, 757, 882, 3252, 2478, 2572, 169, 310, 1591, 2410, 278, 1126, 715, 220, 805, 357, 1604, 2696, 1119, 3232, 152, 2708, 93, 540, 3598, 303, 552, 1367, 461, 68, 130, 2920, 339, 1825, 84, 391, 3031, 1291, 388, 389, 2207, 697, 3074]. M. [2945, 981, 1628, 1198]. M.D [806]. M.D. [3006]. Ma [1282]. Maaike [2343]. Maas [1936, 2756]. Maat [2608]. MacDonald [3066, 245]. MacGillivray [765]. Mach [2545, 1110]. Machamer [679, 358]. Machine [3376, 3611, 2306, 2568, 1677, 2438, 3139, 878, 880, 2247, 326, 3695, 2766, 1628, 2087, 1753, 3429]. Machinenstadt [3254]. Machinery [2344, 3472]. Machines [3559, 1031, 102, 1280, 1949, 1759, 721, 2538, 1155, 690, 295]. Machs [1639]. Mack [312, 216, 412]. Mackowski [2781]. MacLachlan [890]. Maclean [2603]. MacLeod [715, 540]. MacMillan [914, 902]. Macray [1054]. Macrobe [2041]. Macrobium [2041]. Mad [1246, 2816, 1479]. Mad-Doctoring [1246]. Madame [3572]. Maddison [2315, 3125]. Maddox [3067]. Made [970, 752, 149, 735, 923, 3667, 1628, 3091]. Madness [974, 3325, 517, 3334, 1203]. Madrid [3056, 692]. Madrigal [3140]. Magallón [732]. Magazine [2473]. Magdalena [2762]. Magi [350]. Magia [2148]. Magic [1996, 266, 1922, 3042, 1865, 1378, 880, 2315, 2018, 1996]. Magical [3393, 1285]. Magician [1731]. Magick [1808]. Magie [1089, 3482, 2151]. magiques [1691]. Magma [1877]. magnete [1222]. Magnetic [2581, 823, 1030, 3682]. Magnetism [2923, 1824]. Magnificent [65]. Magnus [3019, 2407]. Magnusson [1821]. Magoc [373]. Magoun [973, 2554]. Magrib [1497]. Mahanti [453]. Mahoney [3627, 1745]. Maienschein [3507, 2300, 2930]. Maier [2490]. Maillard [1607]. Maillet [2532]. Mailloux [2762]. Maimonides [3312, 1818]. Maintaining [2951]. Maio [2561]. Majesty [3599]. Major [676, 1205]. Make [2777, 3207]. Maker [802, 3659]. Makers [3735, 3259, 2730, 388, 1998]. Makes [3217, 2451]. Makin’ [203]. Making [3236, 3755, 1024, 740, 3622, 2433, 3247, 62, 3353, 2968, 1120, 3505, 3187, 1140, 315, 2341, 128, 676, 1302, 2163, 1625, 1759, 530, 534, 2756, 751, 1498, 3232, 991, 3614, 1001, 2293, 486, 1879, 988, 3620, 2522, 2539, 1832, 2720, 1853, 1411, 1958, 2780, 2305, 2656, 1051, 2293, 109, 2730, 3324, 3662, 628, 166, 810, 1662, 1531]. mal [2091]. Malachi [1011]. maladies [1686]. Malady [2928]. Malaria [2105, 3426, 2292, 2920, 3755, 1528]. Malaspina [714, 2351]. Malay [1348]. Malcolm [1389, 2465, 342, 3009, 864, 902]. Male [577]. Malin [1968]. Malinowski [2769, 400]. Malmstrom [140]. Malquias [3074]. Malthus [3062]. Mammals [1538]. Mamone [2940]. Man [1415, 2038, 911, 2086, 1215, 3338, 2500, 277, 3237, 1401, 380, 3087, 1153, 2074, 1709, 2297, 735, 923, 3684, 1075, 3078, 2072, 2775, 2774, 2271, 1005, 1416, 791, 1648, 229, 2289]. Man-Made [735]. Manāzir [1364]. Manage [1180, 949, 950]. Managed [391]. Management [406, 1043, 310, 1512, 2360, 2782]. Managerial [396]. Managing [2364, 2364]. Manchester [2690]. Manchukuo [3261]. Mandelbrote [1924, 269]. Manfred [3507]. Manhattan [892, 2658, 3352]. Mania [3608]. Manifesting [295]. Manifesto [135]. Mankind [450]. Mann [925]. Manned [2781]. Mannheim [2070]. Mannix [2725]. Manson [1126, 956]. Mantell [1522]. Mantissa [908]. Mantra [3631]. Manual [37, 1498]. Manuel [300, 2240, 625, 617]. Manufacturing [1951, 310, 215]. Manuscript [1493, 3186]. Manuscripts [3158, 1697, 1054, 712]. manuscrit [364]. Many [2646, 1249, 3723]. Manzoni [2250]. Mao [85]. Maor [687]. Map [1690, 2954, 971, 149, 2965, 3137, 2255, 3146]. Mapmaker [2153]. Mapmakers [351]. Mappae [3566]. Mapping [3729, 912, 3034, 1955, 965, 243, 1520, 351]. Maps [1843, 2244, 2827, 557, 2420, 2496, 443, 2539, 2255, 14, 585]. Maquet [42, 795]. máquina [1952]. máquinas [1500]. Mara [530]. Marbles [977]. Marburg [48]. Marburgo [324]. Marc [2472, 3735, 3340, 627, 371, 2620, 2671, 1344]. Marc-Antoine [2620]. Marcel [2578]. Marcello [679]. March [837, 3274, 578, 320, 958, 996, 1251, 754]. Marchal [2088]. Marché [2858]. Marcia [3006]. Marco [2352, 25, 19, 779, 3138, 3238, 2940, 2352, 2739, 2418, 2810, 1376]. Marconi [1865, 1864]. Marcos [60, 2561, 1394]. Marcus [1959, 2605, 2773]. mare [2842]. Maresi [559]. Margaret [602, 247, 210, 558, 494, 681, 744, 710, 1850, 2048, 1077, 1134, 2980, 2917, 2909, 2830, 1375, 2508, 2369, 988, 1526]. Margins [210]. Margolis [2946]. Margrit [629]. Marguerite [3252]. Mari [44, 3053]. Maria [449, 495, 998, 3728, 3492, 3074, 1820, 751, 2424, 479, 2386, 2142, 186, 2444, 1614]. Marian [2414]. Mariana [2505]. Marianne [3702, 2993]. Mariano [3056, 2050]. Mariculture [6]. Marie [3496, 626, 199, 308, 2930, 3006, 2915, 1111, 1918, 1796, 64]. Marielle [1385]. Marietti [743]. Marijane [2154]. Marika [3702, 677]. Marina [382, 561]. Marine [2698, 2306, 3005, 1624, 2979]. Marinell [519]. Marinella [788]. marins [2979]. Mario [107, 2903, 2826]. Marion [185]. Maritime [1807, 541]. Maritimes [2694]. Marjorie [3703]. Mark [348, 3338, 463, 1364, 27, 971, 2928, 2297, 1580, 1937, 685]. Märker. [1714]. Market [2379]. Marketing [2539]. Marketplace [3412]. Markets [3735, 305]. Marking [3651, 3711, 3439]. Markley [2757]. Marlene [114, 174]. Maroto [479]. Marpeau [1006]. Marriage [1321, 925]. Marrinan [308]. Marrone [871]. Mars [2457, 2757]. Marsh [723]. Marshal [2169]. Marshall [973, 465, 2833, 1884, 30, 2588]. Marsile [2534]. Marta [2270]. Martens [2736]. Martha [571]. Martial [1495]. Martin [2663, 483, 1279, 3659, 344, 2745, 3092, 508, 761, 2189, 2827, 2385, 1143, 1943, 32, 2143, 2743, 2576, 2897, 2280, 2138]. Martineau [3062]. Martínez [3620]. Martyr [3308]. Marvelous [3559, 880]. Marvels [2342, 1828]. Marvin [816, 686]. Marx [1268, 3320, 907]. Mary [2941, 647, 1921, 527, 278, 775, 744, 28, 448, 2529, 964, 2210, 1810, 2355, 2271, 2539, 1255]. Maryinez [218]. Maryland [1360]. Marz [1003]. Más [3732]. Masaki [561]. Masanori [1384]. Maschinen [2997]. Maschinen-Inspector [2997]. Masculine [1759]. Masculinity [1759]. Masochism [657]. Mason [2346, 1868]. Masons [933]. Mass [2883, 734, 1867, 1475, 3187, 1213, 972]. Massachusetts [2692]. Masse [3642]. Masses [2054]. Massimi [2758]. Massimo [324]. Massnahmen [56]. Master [873, 872, 1518, 1817]. Masters [3241, 3088, 1114, 1839]. Mastery [657]. Mastodon [3603]. Masturbation [1087]. matematica [1015]. Matematicas [146]. matematiche [3326]. matematici [38]. matematik [583]. matematiki [1291]. Materia [34]. Material [1321, 1851, 106, 901, 3131, 3651]. materiale [1465]. Materialien [3657]. Materialism [907, 2805, 3712]. Materials [2330, 421, 3430, 2659, 154, 3322]. Materiaux [826]. Materie [2846, 1013]. Maternal [1963]. Mateu [2938]. Math [2021]. Mathematica [501, 1684, 586]. Mathematical [1783, 1787, 2431, 3467, 2531, 3583, 1699, 683, 380, 2017, 1250, 463, 681, 357, 1569, 440, 1589, 484, 2057, 2489, 2198, 835, 1515, 1839, 2325, 889]. Mathematician [2989, 3308, 3170, 3492, 2759]. Mathematicians [1256, 2376]. Mathematics [2718, 1980, 1982, 3553, 3471, 862, 3467, 1362, 1738, 2722, 1751, 2521, 1290, 1167, 3741, 421, 583, 19, 3257, 2488, 1671, 3170, 751, 2996, 606, 1070, 3167, 3668, 2717, 2336, 3750, 2716, 1515, 3512, 3306, 465, 15, 267, 799, 978, 2113, 1977, 3099, 1357, 2431, 1473, 1887, 2383, 3385]. Mathematik [758, 360, 387, 3171]. Mathematikegeschichte [1291]. Mathematiker [1920, 642]. mathématique [613, 2822, 2820]. mathématiques [1488, 1014]. mathematischen [1373]. Mathesis [3316, 2971]. Mathew [379]. Mathews [653]. matière [2534]. Mats [480]. Matteo [745]. Matter [3697, 3, 100, 233, 328, 424, 1, 101, 234, 329, 426, 429, 573, 671, 749, 829, 430, 574, 672, 750, 827, 831, 832, 1334, 1335, 1793, 1794, 2231, 2232, 2516, 2518, 1358, 447, 1366, 1777, 2980, 3210, 3545, 3437, 1927, 2216, 362, 2621]. Matters [1063, 475, 3209, 3030]. Matthew [1527, 296, 2981, 3077, 2441]. Matthias [707, 552, 3068, 2904]. Mattmüller [483]. Mauch [274]. Mauchly [2782]. Maudslay [1603]. Maupertuis [2271]. Maura [2781]. Maurer [719]. Maurice [225, 3348, 2976, 2475]. Mauricio [884, 1401]. Mauritius [57]. Maurits [1795]. Mauro [2472, 3146]. Maverick [394]. Mavericks [1808, 3636]. Max [1849, 2681, 1052, 648, 74, 2458, 170, 2276, 2684, 1727, 3671, 1731]. Max-Planck-Gesellschaft [1052]. Maxine [1942]. Maxwell [169, 2595]. Maxwell’s [2919]. May [3340, 3641, 3410, 3225, 3371]. May-December [3410]. Maya [464, 3307, 1603, 2095]. Mayaud [2738]. Mayer [1856]. Mayerne [1592]. Mayhem [1808]. Mayhew [2479]. Maynard [1142]. Mayor [2333, 3005]. Mayr [2451, 2397]. Mazlish [824]. McBirney [290, 590]. McCabe [1938]. McCall [2944]. McCallum [857]. McCartney [420]. McCauley [2844]. McClain [885]. McClellan [2164, 3641, 437]. McClintock [1270]. McCloskey [1152]. McConnell [3659]. McCook [2165]. McCormmach [2889]. McCosker [800]. McCray [1586, 2859]. McCuaig [3147]. McCurdy [2567]. McDonald [445]. McDougall [561]. McElroy [398]. McGill [2993]. McGovern [2114]. McGowan [847]. McGrath [1380, 2311, 2960]. McGrayne [1001, 189]. McGucken [662]. McGuire [1162]. McKnight [2890]. McMahon [402]. McMullen [268]. McMullin [2609]. McNeely [2285]. McNeil [231]. McNeill [2429]. McNutt [669]. McPherson [185]. Me [3700]. Mead [1134, 1526, 1850]. Meadows [2555, 591]. Meanest [2415]. Meaning [2532, 2026, 3513, 1128, 40, 2198, 250, 3437]. Meanings [1880, 2564]. Means [2587, 1164]. Measure [1206, 2666, 898, 3721, 1356, 1208, 3673, 3216]. Measured [966]. Measurement [1139, 2734, 2310, 2906]. Measurements [1742]. Measures [2354]. Measuring [2881, 1475, 3737, 122]. mecaniche [2046]. Mecca [585]. Mechanical [1235, 3765]. Mechanicks [3133]. Mechanics [2132, 1235, 102, 1917, 42, 3246, 3427, 3512, 75, 72, 631, 1917, 228, 3640, 3378, 2843, 3547]. Mechanicus [2997]. Mechanik [704, 1109]. Mechanisation [1704]. Mechanisms [2852, 2329]. Mechanismus [162]. Mechanistic [2843]. Mechanization [1298]. Medal [1896]. médecine [1607, 1400, 3321, 183]. Media [1780, 2473, 532]. Mediators [744]. medica [34, 449]. Medical [3002, 1063, 1781, 1642, 2108, 2114, 3110, 2945, 3215, 182, 2060, 905, 221, 1828, 3346, 1626, 2557, 3685, 2869, 1592, 2741, 1627, 125, 1355, 2548, 1273, 1055, 128, 3312, 858, 2596, 2688, 2993, 2926, 1758, 2374, 450, 1739, 1233, 1490, 297, 1531, 3560, 2798, 1092, 2442, 694, 2625, 1605, 1091, 2785, 1911, 2564, 2694, 391, 3031, 1796, 857, 1606, 2533, 2048, 3574, 3252, 3312, 1739, 352]. médicale [2257, 2973]. médicales [1735]. Medicalism [181]. Medicalization [1947]. Medicamenta [1490]. medicamentorum [163]. Medicating [3111, 551]. médicaux [790, 490]. medicina [2405, 2742, 1696, 1399, 2142]. Medicine [3697, 1314, 1059, 1782, 1439, 2111, 2114, 2938, 3448, 3496, 2704, 2195, 3166, 3325, 2983, 2285, 181, 3476, 609, 975, 1278, 295, 2385, 3736, 2491, 2875, 1853, 451, 903, 2781, 1194, 1393, 393, 2170, 1728, 1166, 1948, 1237, 1135, 2963, 1883, 2243, 1812, 1675, 1195, 2778, 2686, 1831, 2993, 2926, 1126, 598, 316, 3663, 3182, 1203, 3030, 3484, 3232, 3198, 880, 1481, 2952, 1479, 1595, 3298, 466, 2249, 2625, 1605, 1220, 772, 1594, 3747, 3487, 3080, 1125, 2972, 2541, 2799, 1937, 1368, 3321, 1697, 2472, 2728, 3042, 1812, 1360, 1829, 1122, 771, 452, 2828, 1163, 903, 3043]. Medicine [515, 59, 322, 2603, 2299, 2788, 514, 2857, 2343, 1576, 1173, 1979, 3192, 3551, 1486, 3271, 151, 1477, 1813, 2602, 448, 1499, 3383, 3772, 1606]. Medicines [2659]. Médicis [701]. Medico [2196, 1602]. Medico-Moral [2196]. Medieval [2882, 1437, 1443, 1634, 1889, 1982, 2945, 2883, 1494, 471, 3485, 475, 1366, 872, 1581, 2243, 478, 1224, 351, 3232, 1490, 3488, 1364, 264, 2157, 3547, 2972, 3482, 869, 3043, 2155, 2115, 470, 1223, 2151]. médiévale [2824]. médiévales [2343, 2041]. médiévaux [33]. Medikalisierung [126]. medioevo [2405]. Medir [498]. Meditations [154, 1424]. meditsine [1314]. Medium [1042]. Medizin [3685, 3679, 2372, 867, 164]. Medizinalwesens [2291]. Medizinhistorische [1826]. Medizinische [3418]. Medizinischen [1815, 2485, 3672, 1489]. Meeresbiologie [1178]. Meet [2651, 1992, 2328, 3438]. Meeting [579, 1697, 2904, 3437]. Meets [522]. megaterio [18]. Mehra [75]. Meier [2301]. Meikle [81]. Meiosis [2648]. Meitner [3515, 74, 726]. Meitner-Max [74]. Melancholy [2970, 3325]. Melbourne [3100]. Melchior [3150]. Melling [517]. Melon [705]. Melon-Shaped [705]. Melosi [1143, 344]. Melting [760]. Meltzer [374]. Melville [368, 368]. Members [3175]. Membranes [284]. Memoir [1148, 1152, 531]. Mémoire [3739]. Mémoires [2547]. Memoirs [1030, 1382, 2966, 3509, 1548, 94, 2432]. memoria [190, 3054]. Memorial [3186]. Memories [2432]. Memory [1443, 2132, 227, 680, 3082, 883, 208, 3711]. Men [637, 1044, 2255, 1855, 1743, 2348, 1759, 110, 2699, 3434, 3668, 2747, 196, 788]. Mendel [1349, 3503, 2648]. Mendel-Fisher [3503]. Mendeleev [2357, 1384]. Mendelians [433]. Mendelism [2951]. Mendelsohn [2399]. Mending [1464]. Mendonca [2495]. Menlo [720]. Menopause [2778]. Menou [335]. Menschen [2695]. Menschenversuche [2116]. menschliche [2763]. mensurandi [147]. Mental [3686, 1854, 822, 379, 2179, 3108].

Mentalität [2840]. Mentally [1479]. menti [2729]. Mentor [1119]. Menzhulin [2499]. meraviglie [3394]. mercati [3735]. Mercator [3137]. Merchant [1940, 2034, 2055]. Merchants [2342]. Merciful [1882]. Meric [706]. Meridian [3025]. meridiano [498]. Méridien [761]. Méridienne [761]. Merit [3216]. Meriwether [1118]. Mermaid [167]. Merrill [2570, 174]. Merton [1798, 2102]. Mesa [2623]. Meselson [1623]. Mesmerism [3595]. Mesmerized [1837]. Mesoamerican [140]. Mesopotamia [3553, 3038, 1485]. Mesopotamian [2335, 606]. Message [20]. Messenger [2298, 1533]. Messianism [1060]. Messzirkel [2560]. mesure [2549, 2887]. Metabiography [2847]. Métailié. [3634]. Metal [2185]. Metamorphosenlehre [3047]. Metaphor [1869]. Metaphorik [3672]. Metaphors [269, 284]. Metaphysics [1889, 1235, 3131, 1773, 3527, 2552, 467, 1308, 947, 23]. métaphysique [982, 469]. metaulx [875]. metaura [35]. Metchnikoff [1269]. Meteorite [2944]. Meteoritics [2944]. météorologie [614]. Meteorologists [3520, 971]. Meteorology [2748, 2150, 1261]. Method [745, 2712, 1229, 956, 1413]. Methode [2347]. Methoden [896]. méthodique [3314, 364]. Methodology [784, 1039]. Methods [2916, 2677, 1198, 2136, 339]. Methuen [267]. Metrich [609]. métrico [498]. metro [498]. Metrology [3561]. Metropolitan [2677]. Mettenleiter [1826]. Metzger [1558]. Metzl [2094]. Meulenbeld [1091]. Mexican [884]. Mexico [3426, 1482, 1027, 2367, 2835, 2832, 912, 1214, 3352]. Meyer [1020, 616, 1567]. Meyerowitz [2443]. Meyers [247]. Meyerson [1558]. Meyerstein [2997]. Mezzo [1337]. Mi [2349]. Mice [2305, 196]. Michael [1178, 3158, 1974, 605, 595, 742, 2490, 222, 90, 531, 2159, 512, 2358, 3095, 370, 858, 400, 505, 474, 24, 515, 1428, 316, 19, 192, 1714, 172, 3151, 2684, 148, 792, 350, 314, 667, 814, 807, 487, 2975, 960, 905, 3627, 2927, 1848, 1971, 2691, 1873, 2974, 1809, 1042, 2994, 2145, 935, 2982, 2904, 904, 1834, 1869, 2750, 2357, 2090, 1086, 2901, 2769, 3092, 848, 2975, 2530]. Michał [2536]. Michaux [1825]. Michel [72, 3649, 2094, 357, 2751, 68, 115, 582, 2208, 2820, 1610, 30]. Michel-Pierre [582]. Michela [2758]. Michele [804, 2411]. Michelle [1880, 1108, 2931]. Mick [89]. Mickey [403]. Mickulas [3509]. Mico [165]. Microbe [637]. Microbes [1349]. Microbial [3771, 2573]. Microbiologist [2086]. Microcinematography [2526]. Microcosm [331]. Microeconomics [369]. Micrologu [3483]. Micrologus [2115]. Microphysics [1065, 106]. Microscope [3382, 198]. Microscopes [3735, 2414]. Microscopi [3735]. Microscopy [3427]. Microwave [1761]. Mid [938, 114, 2319, 2123]. Mid-Eighteenth [2123]. Mid-Nineteenth [2319]. Mid-Twentieth [938, 114]. Middle [1440, 2881, 787, 151, 1821, 1055, 2243, 1812, 1491, 1490, 3488, 772, 1911, 615, 3562, 3147]. Middleton [1715]. Middletown [1022]. Midgley [964]. Midhat [584]. Midway [390]. Midwifery [993, 1791, 3650]. Midwives [771, 1506, 3133, 993]. MiG [1544]. Might [3286]. Miguel [39, 2474]. Mikael [3472, 536, 2964]. Mike [273, 1598]. Mikko [2569]. Mikroskop [1284]. Mikroskopie [1856].

Mikulás [1402]. Milbrath [464]. Mile [2702]. milenio [2386]. Miles [1896, 1102]. Miletus [2251]. Milgram [3078]. Militant [846]. Military [2350, 646, 1031, 3522, 421]. Mill [1144, 1512]. Millán [2995]. Millán-Verdú [2995]. Millenarianism [1093, 1060]. Millennia [2787]. millennio [493]. Millennium [244]. Miller [1978, 2753, 3001, 886, 3602, 85, 394, 275, 711]. Milligan [50]. Millikan [3340, 371]. Millin [3364]. Mills [921, 223]. Milton [218, 399, 94, 487]. Milwaukee [1412]. min [3261]. Mind [2195, 401, 3711, 1558, 378, 1837, 3490, 412, 1298, 1377, 3560, 2066, 3666, 2507, 1246, 2871, 2725, 1877, 1121, 3338, 1159, 1255, 3376, 655, 3442]. Mind-Body [3560]. Mind-Brain [401]. Mindell [1010]. Mindful [3475]. Minds [3697, 974, 3441, 1212, 1353]. Mineke [2676]. Mineral [673]. minerale [2977]. Mineralogy [688]. Miners [3133]. Ming [2158, 1369, 3268]. Minguo [3333]. miniere [2977]. Minima [1826]. Mining [1401, 570, 2027]. Minna [2849]. Minnesota [3529]. Minoan [765]. Minor [812, 403]. Minotaur [765]. Minta [2155]. Minton [2302]. minus [439]. Miquel [1820, 165]. Miracle [2860]. Miracles [471]. Miriam [2622]. Mirjam [224]. Mirko [754]. Mirror [2532, 1499, 3404]. Misa [3472, 2118]. Miscellanea [501]. Miscellaneous [889]. Misconceptions [338]. Misfortunes [1874]. Miss [1959]. Missile [742, 3272, 926, 543]. Missing [3623, 1005]. Mission [3003, 3132, 2700]. Missionaries [3752]. Missionary [3002]. Missions [2238]. Mississippi [374, 1609, 2569]. Mistake [2612]. Mistreatment [1479]. Misunderstood [1754]. Misuse [603]. Mitchell [2933, 1614, 405, 926, 1628, 774, 116]. Mitchinson [1860]. Mitigation [2089]. Mitman [659]. Mitteilungen [3103]. Mittelalter [3486, 476]. mittelalterlichen [1584]. Mittelmeers [309]. Mittelpunkt [3103]. mixte [826]. Mixtures [3033]. Mladjenovic [303]. Mobility [1947]. Mociño [2051]. Modal [1889]. Model [3358, 3262, 2909, 2657, 1396, 415]. Modeling [3716]. Models [2241, 936, 2428, 1530, 2573, 744]. Moderate [135]. Modern [3601, 2678, 2852, 1315, 945, 1060, 1169, 1782, 1785, 1441, 1889, 2105, 2109, 1991, 3106, 3109, 3111, 2735, 794, 2885, 2060, 3234, 2433, 631, 3561, 3275, 2732, 3046, 3325, 3535, 2667, 3662, 2048, 1761, 3685, 2528, 1867, 2621, 1423, 2720, 2895, 1737, 108, 2869, 2392, 2610, 521, 3304, 3574, 2953, 3365, 3256, 3551, 557, 3057, 2849, 3638, 1261, 1050, 3052, 2052, 1366, 3408, 1670, 2031, 1008, 1232, 1579, 1241, 3163, 1534, 575, 882, 1095, 2893, 3590, 752, 2899, 2188, 2341, 2963, 3274, 1154, 2670, 546, 3401, 53, 728, 3337, 1675, 1949, 2889, 3368, 1759, 2778, 2410]. Modern [149, 2916, 1132, 2928, 1124, 3472, 1047, 2049, 369, 699, 2606, 269, 1094, 235, 529, 3257, 2756, 1296, 172, 441, 1774, 3182, 3160, 1703, 3517, 3545, 2830, 2835, 6, 784, 2747, 2605, 28, 3575, 3667, 767, 2650, 3650, 448, 1531, 3708, 2256, 2576, 1001, 677, 2342, 2677, 3570, 2745, 3156, 2958, 3223, 2523, 1422, 3045, 2459, 709, 1096, 246, 2261, 3547, 2633, 3297, 3049, 2956, 3268, 883, 3595, 487, 3306, 1097, 693, 2207, 2286, 2558, 3583]. moderna [2742, 3319]. Moderne [3101, 636, 982, 2733, 176, 2278, 3321, 1320, 3645]. modernen [3671, 3013]. modernidad [1492]. Modernism [2195, 3772, 415]. Modernist [1467, 1869, 3452, 2629, 2004]. modernità [1381]. Modernitätskritik [2284]. modernité [1736]. Modernity [2208, 3749, 1078, 2358, 1102, 1615, 2462, 3299, 2626, 3598, 87, 536, 3664, 2910]. modernizacion [200]. Modernization [2434, 660, 3181, 2832]. Modernizing [2507]. Modest [1337]. Modified [1540]. Moffitt [1867]. Mogren [1144]. Mohr [1795]. Mohum [293]. Moira [49]. Moisants [65]. moitié [3740, 1735]. Mojsisovics [2281]. moléculaire [3355]. Molecular [197, 555, 676, 1421, 2573, 3097, 2329, 897]. molécule [2192]. Molecules [201, 1408]. Molella [3773]. Moleski [3092]. Moletti [1917]. Molina [2845]. Moliner [1602, 449]. Mollan [1412]. Molland [1655]. Molly [885]. Molnar [350]. moment [2738]. Momentous [189]. Moments [1782, 2857]. Monarquía [714]. Monasteries [1821]. Mond [3239]. Mond-Die [3239]. monde [2887, 2619, 2426, 1225, 582]. mondi [3564]. mondi/Über [3564]. mondial [2363]. mondo [1483, 884]. Mondphasen [10]. Mondtheorie [10]. Money [2686, 3244, 3523, 1149, 1866]. Monforte [2430, 692]. Monge [2351, 3467]. Mongolia [1538]. Monism [283]. Monistische [2205]. Monkey [2853, 3295]. Monmonier [971]. Monnais [183]. Monnais-Rousselot [183]. Monro [1246]. Monsieur [2877]. Monster [1709, 880]. Monsters [3046, 2947]. Mont [3649]. Mont-Saint-Michel [3649]. montagne [969]. Montague [1147]. Montana [1538, 914]. montañés [479]. Monte [102, 2880]. Montgolfier [1512]. Montgomery [2169]. Montiel [255]. Montreal [2993]. Monturiol [1952]. Monuments [3038, 374, 2832]. Moog [1762]. Mooij [554]. Moon [140, 980, 3725, 1294, 1698, 965, 1415]. Mooney [2701, 265]. Moonwatch [3683]. Moore [3098, 1970, 2759, 1395, 634]. Moral [2407, 3325, 2196, 1276, 3708, 1724, 649, 2574, 2135, 1399]. Morality [996]. Morals [2310]. Moran [1616, 2983]. Morandi [1504]. Mordechai [1633]. Morgagni [610]. Morgan [744, 2951]. Morgentaler [900]. Moritz [1401, 2997]. Morphologische [896]. Morrell [2624]. Morris [3351, 1406, 305, 211]. Morrison [3324, 681, 744]. Morrison-Low [3324]. Mort [2196]. Mortal [3637]. Mortality [376]. Morton [539]. Morus [2627, 2720, 367]. MOS [2559]. Mosaic [399, 1396, 5]. Mosca [225]. Moscow [3184]. Moser [3040, 118]. Moses [2751]. Mosini [3360]. Moss [1831]. Most [2346, 3079, 2978, 2918, 1865, 2074, 2420, 1623]. Moth [2692]. Mother [2776, 3536]. Motherhood [3518, 3536]. Mothers [1135]. Motion [447, 3365, 1608, 1080, 2729]. Motions [1583, 2415]. Motor [3596]. Mott [21]. Motterlini [745]. Mould [1041]. Mount [2622, 2848]. Mountain [1939, 397]. Mountains [2420]. Mountaintop [2045]. Mouse [1761]. moustiques [1736]. Movable [3415]. Movement [1009, 1365, 2390, 1524, 42, 795, 1757, 2391, 186, 1882]. Movements [345, 3522, 1191, 532]. Mowery [206, 741]. Moy [1031]. Moyen [2823, 3159, 1691, 472, 614, 790, 2879]. Moyer [958, 2157]. Mozzochi [922]. Mr [1527, 3668]. MRI [3682]. MS [2945]. Much [731, 2197]. Muezzin [2883]. Mühlberger [476]. Muir [287, 2437]. mujeres [732]. Muller [288, 2407, 3747, 1109, 500, 2361, 769, 3035, 2846, 2831]. Müller-Strahl [2846]. Muller-Wille [288]. Müller’s [3178]. Multhauf [2324]. Multiculturalism [1732]. Munchen [194, 3171]. Münchener [725]. mundi [3566, 1238]. Mundo [1820, 2264, 3327, 2252]. Munoz [146]. Munro [322]. Muraskin [317]. Murder [1323, 2597, 3163]. Muris [147]. Murky [2958]. Murphy [1611, 2931, 254, 454]. Murphy-Lawless [254]. Murri [2986]. Muscles [436, 2213, 42, 795]. Muse [2365]. Museo [3056, 2044]. Museum [2460, 1807, 541, 2461, 96, 1425, 3151, 420, 46, 3196, 1566, 909, 3018, 3040, 1905]. Museums [2463, 2288, 1389, 2462, 509, 2368, 3044, 2810, 1665, 178, 677]. Music [976, 1921, 2606, 1107, 1472, 1344, 1922, 3741]. Musical [778, 3624]. musicale [976]. Musicians [3259, 1107]. Musick [2606]. Musicus [3068]. Musik [3068]. Musiktheorie [1584]. Musselman [2891]. Musser [1608]. Mutation [2648]. Mutel [1414]. Mutilation [577]. My [2697, 1411, 1874, 2213, 1816, 852, 2776]. Myers [660, 812]. Myles [3259]. Mynttinen [2751]. myopathies [298]. Myrmecology [3069]. Mysteries [140, 173]. Mystery [2126, 193, 173]. Mystic [3419]. Mysticism [3317]. Mystik [153]. Myth [3682, 2664, 2127, 1244, 661, 1811, 2078, 765, 3642, 1731]. Mythe [161]. Mythen [1985]. Mythology [1868]. Myths [2255, 3300, 2659].

N [1995, 738, 2307, 2364, 2912, 417, 548, 3090, 179, 1941, 366, 3678, 2170, 2660, 2310, 3529, 461, 886, 1396, 91, 3473, 2047, 3074, 2450, 264]. N. [2206, 1882, 2776]. N.Y [2472]. NAACP [7]. NACA [216, 2084, 2090]. nach [2116, 1734, 3179, 163]. nachale [1267]. Nachkriegszeit [2840, 2818]. Nachtsheim [2695]. Nacional [732]. Nadal [2995]. Naddaf [2402, 2600]. Nadia [2082]. Nadine [401, 2708]. Nahin [439]. Naish [503]. Naissance [3157, 3494, 176]. Nakayama [1406]. Names [2867]. Naming [965]. Nance [1592]. Nancy [637, 210, 2893, 308, 397, 612, 2681, 1947, 2303]. Nano [2696]. Nano-Hype [2696]. Nanotechnology [2696]. Naomi [128, 64, 1413]. Napoleon [1393, 1389]. Napoleonic [2484]. Narcotic [2168]. Narender [453]. Narrating [1325]. Narrative [2400, 2457, 1723, 2259]. Narratives [2718, 2614, 1003, 1116, 2716, 2071, 3358, 1494]. Narrators [554]. NASA [216, 2770, 2700, 3424, 2925, 214, 2084, 2090, 1545]. nascita [2401, 2252]. Nash [456]. Nasser [2315, 1849]. Nastasi [2996]. Nat [3433]. Natalia [2156, 2845]. Natalie [347, 2560]. Nathan [2249, 206, 3074, 2395]. Nation [3243, 49, 2184, 2849, 2836, 2668, 1027, 3694, 1017, 3598, 1510]. National [2307, 1807, 2078, 1359, 2827, 533, 541, 1301, 1532, 1389, 2768, 1461, 1954, 3541, 1051, 1745, 420, 130, 3333, 962, 314, 2234, 562]. Nationalism [3342, 1027, 3264, 270, 2303]. Nationalizing [2073]. Nationalsozialismus [2182]. Nations [131, 589, 589]. Native [3059, 851, 563]. Natur [55]. Natura [2115, 3326, 998, 3615, 1483]. Natura-Cultura [1483]. Naturae [629, 492]. naturais [2166]. Natural [3344, 1263, 1632, 2109, 1983, 167, 44, 684, 1665, 1704, 2545, 5, 3496, 1689, 3389, 1000, 1266, 2427, 1450, 3391, 2416, 3398, 870, 365, 2810, 2035, 1699, 3546, 976, 1922, 1100, 575, 96, 2922, 3655, 169, 2889, 1709, 2740, 956, 804, 516, 744, 2917, 2486, 1698, 1460, 876, 1703, 3232, 2605, 767, 1035, 1147, 3731, 3332, 3723, 590, 3562, 3301, 2773, 3049, 3433, 3381, 884, 146, 2705, 134, 3576, 343, 612]. Naturale [2044]. Naturales [3056]. naturaleza [44]. naturaliser [2278]. Naturalism [1670, 2203, 1777]. Naturalist [2162, 2061, 2361, 3118, 661, 2309, 2425, 3587]. naturalistas [1713]. naturaliste [627]. naturalisti [3735]. Naturalistic [1061]. Naturalists [2998, 632, 3591, 2412, 2180, 2900]. Naturbegriff [2347]. Nature [1315, 247, 513, 1945, 562, 2885, 2911, 1968, 1719, 1068, 884, 818, 659, 3231, 2667, 2068, 2244, 2728, 2771, 2427, 1450, 2929, 3183, 109, 2061, 3725, 557, 8, 454, 1028, 3757, 1620, 1120, 3655, 1304, 2147, 2348, 3490, 3217, 2631, 3413, 2390, 411, 1351, 2653, 2311, 2651, 1510, 3000, 285, 2603, 1040, 148, 2650, 3039, 2165, 249, 1147, 3746, 2618, 3586, 2454, 2342, 1287, 3612, 3332, 3587, 2611, 27, 600, 2652, 612, 2533, 3248, 2492, 2675, 1968, 2058, 1883, 2390, 3003, 3020, 1101, 907, 611, 413, 1568, 2999, 2600, 2437, 2599, 2544, 2167, 1919, 2382, 743, 3060]. Nature [1984, 3164, 3243, 1082, 2473, 3468, 509, 3499, 3165, 2135, 3673, 438]. Nature-Oriented [454]. Nature-Study [2390]. Nature/Culture [2771]. naturelle [1702, 2343, 3400, 3316, 502, 3159, 875, 3396, 3018]. Natures [2626]. Naturethik [2103]. Naturforschers [1004]. Naturforschung [3418]. Naturganzen [3063]. Naturgeschichte [3743]. Naturkunde [769]. Natürliche [2838, 2407]. Naturmodell [629]. Naturschutz [2103]. Naturvetenskapen [2080]. Naturwissenschaft [264]. Naturwissenschaften [3171, 3479, 1934]. naturwissenschaftliche [337]. naturwissenschaftlichen [2189]. naturwissenschaftlicher [2258]. Naturzerstörung [1505]. Natuurkunde [1936]. natuuronderzoekers [906]. Nauenberg [792]. nauk [1354]. Nauka [1536, 2943]. Nauta [568, 473]. Naval [3155, 1019, 2176, 1424]. Navarro [146, 586, 165]. Navegación [625]. Navi [691]. Navigation [1070]. Navigator [2254]. navíos [479]. Navy [1131, 833, 3599]. Nayan [2377]. Nazi [548, 1276, 2926, 1739, 1606]. Nazionale [3735]. Nazis [2376]. Nazism [3181]. Nead [910]. Neal [2261]. Near [3724]. Nearest [667]. nebst [34]. Nebular [1383]. nébuleuse [716]. Neck [1087]. Necromancer [37]. Nederlandse [499, 906]. Need [3208]. Needell [1032]. Needham [2270, 1194, 2214]. Needles [2809]. Needs [2942]. Neglect [2224]. Neglected [3501]. Negotiating [2552, 1123, 3250]. Negra [3732]. negras [2133]. Neher [137]. nei [3735, 1483, 3726, 2533]. neiges [3495]. Neil [593, 3241, 3131, 3151, 2774, 252, 2436]. Neither [3231]. Nel [1517, 2405, 2338, 3394, 777]. Nélia [2549]. Nelkin [1653]. nell [493]. nella [2401, 777]. nelle [1483]. nell’età [3319, 2411, 3395]. nell’Italia [3394, 2282, 2842]. Nelly [3679]. Nelson [2197, 459, 459, 741]. nemetskie [1314]. Nenci [38]. Neoplatonism [1684]. Neotropen [1004]. Neptune [894, 2306]. Nerad [559]. Nernst [3580, 728]. Nerve [2329, 3180]. Nervenklinik [188]. Nervenschwäche [2284]. Nerves [2784, 2910, 1136, 2219, 2213]. nervoso [58]. Nervous [2891]. Nesta [1188]. Netherlands [940, 592, 3475, 2066, 145]. Network [1237, 1766].

Networking [2187, 1280]. Networks [1024, 2996]. Netz [1362, 2897]. Neuberg [3414]. Neubestimmung [1279]. Neudobnye [2499]. neue [126]. neuf [2822]. Neufeld [742]. Neumann [556]. Neural [2554]. Neurology [2171]. Neuroscience [973, 2554]. Neurosciences [377]. Neuroses [2799]. Neurotransmitters [2784]. Neutrons [2519]. Neuzeit [3645, 769, 3035, 150]. Neville [1858, 1800]. Nevskaia [1995]. Newbold [644]. Newengland [1600]. Newman [2892, 2599]. Newnham [433]. Newport.S [1309]. News [579, 2959, 99]. Newspeak [1769]. Newton [944, 2113, 271, 1235, 683, 3505, 2606, 2980, 876, 2347, 2415, 484, 3470, 1373, 1924, 3245, 3320, 2739, 2345, 2488, 2417, 2974, 3326, 1507, 1822, 2413, 2672, 2163, 1708]. Newtonian [3505, 792, 1556]. Newtonianism [2413]. Newtons [3657]. Next [1629]. Nexus [2657]. nian [3432]. Niccolò [704, 484, 1999]. Nice [2313]. Nicholas [186, 445, 2265, 2932, 1104, 2862, 2417, 2274, 1928]. Nichols [932]. nichtreduktive [3763]. Nickel [609]. Nicola [999]. Nicolaas [2847]. Nicolas [198, 1806, 1696]. Nicolaus [2476]. Nicole [238, 3041]. Nicolle [3002]. niebla [2351]. Niels [301, 2904, 1960, 2458, 1967, 301, 3465]. Nier [720]. Nieto [1952, 884, 2938, 134]. Nieto-Galan [1952, 2938, 134]. Nietzsche [1275]. Nigel [3759, 872]. Nightingale [639]. Nightwork [2215]. nigromance [3482]. NIH [1185]. Niko [2615]. Nikolai [2685]. Nikolaus [1371, 363]. Niles [3224]. Nils [2640]. Nina [2657, 2939]. Nineteenth [2718, 2987, 1433, 1978, 910, 1122, 3412, 3511, 3175, 1266, 3059, 299, 3415, 903, 1003, 1404, 1743, 2473, 1128, 3505, 3259, 2197, 2994, 1865, 1280, 3748, 123, 3001, 2428, 516, 2356, 1513, 3759, 1253, 3668, 3744, 367, 3499, 2687, 284, 1932, 3504, 3417, 3301, 2424, 3062, 3772, 1186, 2319, 3693, 834]. Nineteenth-Century [2987, 910, 1122, 3412, 3511, 1266, 299, 3415, 903, 1003, 1404, 2473, 1128, 3505, 3259, 2994, 3001, 2356, 1513, 1253, 3744, 367, 3499, 2687, 284, 3417, 2424, 3062, 1186, 3693, 834]. niu [3200]. No [3144, 1624, 3006, 3775, 2495, 550]. Noah [2535]. Nobel [1979, 2472, 2681, 2380, 1875, 2992, 2472, 325, 189]. Nobility [788]. Noble [102, 1244, 1745]. Nobler [2415]. Nobody [2476]. Nocturnals [1807]. Noel [2072, 2738]. nog [906]. Noise [3242, 3765]. Nolla [1820]. Nollet [1509]. Nolte [3101]. nombre [3348]. nombres [2408]. nome [777]. Nomenclature [965]. Nominators [2472]. Nominees [2472]. Non [1357, 967, 1374, 2618, 2091, 3075]. Non-Euclidean [2618]. Non-Western [1357, 967]. Nonnoi [877]. Noosphere [915]. nor [3231]. Nora [735]. Norberg [2563, 2782]. Nordenskjöld’s [2680]. Nordic [1325]. Nordmann [2994, 53]. Noretta [2577]. Normal [2498, 683, 2299, 518]. Norman [460, 2187, 656, 1825]. Normative [3288]. Normativity [3359]. Norms [3523]. Norris [727, 50]. Norsk [2493]. Norte [2279]. North [1632, 2111, 740, 511, 2042, 456, 1921, 381, 542, 1237, 89, 2606, 3334, 2783, 3501, 180, 2201, 2462]. Northern [1433, 1048]. Northwest [1958, 1070]. Norton [746, 1424]. Norway [2124]. nos [1352]. Nostrums [1578]. Notable [1431]. Note [3374, 2006]. Notebook [3512, 212]. Notes [952, 1066, 1182, 1329, 1447, 1555, 1646, 1788, 1891, 2000, 2122, 2226, 2316, 2387, 2455, 2512, 2792, 2579, 2644, 2710, 2863, 2948, 3016, 3112, 3202, 3279, 3362, 3443, 3533, 3616, 3691, 3786, 724, 2069, 3381, 2472]. Nothing [684]. Notions [3266, 3106]. Notorious [995]. Nouhuys [145]. nouveau [826, 716]. nouvelle [1300, 2738, 3355]. Nova [1222]. Novara [3145]. Novecento [2173]. Novel [3504, 1356, 1711]. Novelists [3591]. Novellierung [274]. Novels [2597]. Novelty [291]. November [579, 3124]. novembre [361]. novus [480]. Nowhere [2830]. Noyes [657]. Nozal [768]. Nuclear [650, 2175, 3341, 533, 2212, 3525, 531, 2304, 2169, 2802, 957, 383, 810, 726, 3424, 3185, 303, 2581, 2365, 2702, 3273, 3770, 2440, 3119, 3352, 891, 1768]. nuestra [2742]. nueva [2430]. Nuevo [884]. Number [1440, 3126, 3129, 2664, 1257, 2261, 778, 584]. Numbers [380, 203, 3347, 722, 503, 1663, 3084, 685]. Numerical [3335]. Nuncius [3141, 2988]. Nunes [1496]. Nuns [2197]. Nuremberg [2926]. Nurses [1276]. Nursing [2197, 1771, 1480]. Nurturing [599]. Nut [906]. Nutrition [1673]. Nutritional [2171]. Nutzpflanzen [163]. Nye [2071, 2529]. Nylon [3274].

O [812, 2061, 2297, 2751, 2033, 794, 3646]. O-Factor [794]. Oaks [1600]. Oaxaca [1048]. Obesity [713]. Object [2241]. Objectifying [901]. Objectivity [2445, 3698, 3095, 405, 3466]. Objects [2460, 3662, 2950, 529, 1187, 3246, 2288, 782]. Objektivierung [3068]. objetividad [1976]. Oblivion [2274]. O’Boyle [1697]. O’Brien [90, 1047]. Obrist [2824]. Observateur [2894]. Observation [3121, 2714, 1983, 3232]. Observationes [363]. Observations [979, 1363, 1375]. Observatoire [2277]. observatoires [968]. Observatories [2477, 336]. Observatory [45, 3343, 2176, 2848, 1140, 893]. Observed [69, 3096, 3150]. Observed/The [3096]. Observers [2950, 3332]. Observing [682, 2985]. Obsession [1675, 3294]. Obsolescence [3091]. Obstetric [254]. Obstetrics [3650]. ocaso [768]. occasion [1310]. Occident [3482]. Occult [1378, 2018]. occulte [1322]. Occultism [3393, 2286, 2558, 3595]. Occupation [3261, 1885]. Occupational [1135]. Ocean [2761, 3599, 1019, 2176, 3771, 2544, 2787]. Oceanographer [1763]. Oceanographers [2698]. Oceanographic [1570]. Oceanography [3269]. Oceans [3770]. Ocherki [1354]. Ochoa [2444, 332]. Ochre [3514]. Ochs [2329]. Ockham [3390]. O’Connor [901]. O’Connor.S [1755]. October [3457, 370, 2806]. Odd [902]. Odie [50]. Odilon [2911]. O’Donnell [1626]. Odum [1141]. Odyssey [2666, 2641, 2769]. Oeconomicus [163]. Oeconomies [2672]. Oestmann [2819]. œuvre [2620]. Oeuvres [1228, 469]. Ofer [2415]. Off [539]. Offentlichkeit [337]. Offers [3532]. offerta [3735]. Office [852, 280, 2548, 1739, 2697]. Officer [560]. Officers [1019]. Official [96]. Offit [3085]. Ogborn [1102]. O’Grady [2251]. ohne [1291]. Oil [916]. Ökologie [1324, 2103]. Olaf [55, 3047]. Olarte [884]. Old [2697, 519, 1055, 2990, 546, 2686, 432, 3318, 3675, 1460, 1758]. Old-Fashioned [546]. Oldenburg [1918]. Oldroyd [2966, 688]. O’Leary [3638]. Oleinik [91]. Oliner [2762]. Olival [1619]. Oliver [20, 508]. Oliverio [58]. Olivier [982, 2814, 1596]. Olmsted [802]. Olson [2967]. Olsson [480]. Oltra [2995]. Omaha [806]. O’Malley [446]. Omega [2072]. One [812, 2472, 911, 3262, 547, 550, 937, 399, 395, 2834, 439, 3498]. Online [2875, 954]. Only [3175, 759, 380]. Onnes [1842]. Ontario [1007]. ontdekking [2724]. Ontological [2656]. Ontology [1162, 3322]. onwrikbaar [2676]. O’Odham [1397]. Oosterhoff [1564]. Oosterhuis [1120]. Opdycke [550]. Open [3784, 2703, 394]. Open-Heart [394]. Opening [1768, 3425]. Openness [2033]. Opera [1222]. Operation [3683]. Operations [3168]. Opere [3735]. Opinion [2982, 432]. Opinionated [1161]. Opinions [1816]. Opioids [1994]. Opium [2292]. Oppenheimer [3513, 649, 2216, 2529, 3197]. Opposite [1583]. Opposites [1670]. Opposition [532]. Optical [1925, 708, 1259]. Optics [2882, 2582, 869, 463, 2043, 3041, 654, 1925, 1281, 708]. Opticus [619]. Optimismus [636]. optisches [619]. Opuscule [875]. Oral [219]. Orange [2853]. Orbis [3146]. Order [2061, 976, 2147, 2908, 2423, 3499, 3688]. Ordered [2357]. Ordering [3386]. Orders [2272]. Ordinary [2865]. Ordinatio [3390]. Ordnung [2184]. Ordnungen [2838]. Ordover [2303]. Oren [2775]. Orent [2087]. Oreskes [1413]. Oresme [238, 3041]. Orfila [2938]. organi [58]. Organic [631, 2634, 2428, 2581, 306]. organica [2977]. Organicisme [2675]. organischer [2846]. Organism [878]. Organisms [80, 1187]. Organismus [1279]. Organization [3607]. Organizational [396]. Organizations [97, 2440]. Organotherapy [3282]. Organs [2933]. Orgasm [2598]. Oriented [454]. origenes [18]. Origin [712, 3481, 2926, 2758, 23, 1845, 3385]. originale [969]. Originality [2536]. Originals [2025]. origine [2270]. origines [2739, 1610, 142]. Origins [1176, 3670, 2097, 1299, 2081, 2161, 1336, 197, 2402, 2083, 2781, 2892, 3576, 1732, 2878, 1588, 537, 2390, 3375, 3149, 369, 1711, 2917, 1774, 3366, 1298, 1306, 1846, 866, 506, 3575, 2203, 3076, 3193, 289, 2730, 1682, 2289, 1202, 1606, 897, 2634, 2187, 118, 3593, 136, 689]. Orle [2167]. Orna [2575]. Ornithology [1809]. Ørsted [3581]. Orthodox [2076]. Orthopaedic [2233]. Orville [215]. Osborn [2154, 2289]. Osiris [680, 2137]. Oskar [2753]. Osler [515]. Osmanli [583, 583, 1370]. Osservazione [1694]. Österreich [3227, 3330]. österreichischen [2284]. Ostherr [2913]. Ostwald [630]. Ostwalds [3331]. Otechestvennaja [2943]. Other [167, 158, 1828, 1616, 1496, 193, 2984, 2341, 155, 2017, 412, 148, 1706, 1220, 710, 2979, 3147, 823]. Otherness [554]. Othmar [3321]. Otis [1132, 284]. Otte [19]. Otter [368, 180]. Otto [2680, 638]. Ottoman [583, 1370]. oubliés [1310]. Ouest [2979]. Our [974, 2114, 3220, 2936, 2844, 1155, 2805, 1282, 2156, 1757, 408, 2552, 340, 1416, 122, 664, 3476, 1476]. Ourselves [2364, 3690, 1680]. Outbreak [1961]. Outer [1873]. Outpost [1795]. Outposts [3283, 3076]. Outrageous [127]. Overseas [2238]. Overthrow [333]. Overview [2326, 3597]. ovo [1103]. Owen [1562, 2286, 2476, 2806]. Own [1558, 2655, 1282, 322, 1757, 1147, 893, 1816]. Owns [1445]. Owsei [2008]. Oxford [2472, 3760, 2031, 2752, 2729, 2643]. Oxford/New [2472]. Oxy [755]. Oxyrhynchus [755]. Ozone [1180].

P [2986, 843, 3646, 3234, 2021, 2381, 871, 222, 365, 3141, 2850, 2268, 542, 169, 384, 2345, 1936, 2993, 2030, 308, 139, 2708, 506, 3773, 1051, 2290, 755, 1884, 2766, 705, 1291, 3062, 443, 3277]. P. [2936, 3434]. Paarlberg [739, 739]. Paavo [78]. Pacemaker [1625, 1155]. Pacey [250]. paciente [255]. Pacific [1036, 540, 1570, 3003, 2979, 342, 2783, 3064]. Pacífico [2351]. Pacifique [2979]. Padovan [1160]. paese [3394]. Page [560]. Pahta [470]. Paige [1309]. Pain [1994, 694, 658]. Painlevé [561]. Paint [1758]. Painting [1287, 3427]. Paintings [2979]. Pais [79]. paisaje [2351]. països [1462]. Päivi [470]. Pakal [3307]. Pala [156]. Palaeolithic [3675]. Pale [2865]. Palenque [3307]. Paleoanthropology [3586]. paleontologia [18]. Paleontology [1218, 3549]. Palevsky [647]. Palladino [740]. Palle [522]. Palm [592]. Palma [2762]. Palmer [2625]. Pálsson [2420]. Pamela [216, 774, 3504, 2542, 2033]. Pampas [2505]. Pampered [1511]. Pamphlets [2057]. Panama [3020]. Panda [3425]. Pandemic [3107]. Pandora [138]. Pang [2287]. Panisperna [2363]. Pankenier [979]. Panorama [2308]. Pantin [3157]. Panza [2739, 19]. Paola [3319, 2282]. Paolo [994, 740, 778, 3552, 1094]. Paoloni [526]. papel [255]. Paper [1557, 2428, 524, 2040, 1367]. Papermaking [1512]. Papers [3158, 1021, 596, 2362, 1612, 3340, 889, 1697, 720, 3410, 371, 275, 2751, 75, 68, 420, 1269, 711, 3641, 3028]. Pappus [862]. Papyri [755]. Para [1976, 768, 884]. Paracelsian [2541, 1782]. paracelsisme [3311]. Paracelsismus [150]. Paracelsus [150]. Paradigm [1616, 333]. Paradigmawechsel [3679]. paradigmi [1049]. Paradigms [2789]. Paradise [3034, 2180, 1093]. Paradox [1393]. Paradoxical [1351]. Paradoxien [3402]. Paralipomena [1925, 708]. Parallax [1206]. Parallel [1469, 3312, 2486]. Parapegmata [3724]. Paraphrase [607]. Parasites [1770, 3407]. Parasitic [956]. Parcours [1006]. Pardo [3604]. Parents [2656]. Paribus [1981]. Paris [761, 2987, 1783, 2164, 683, 2688, 1372, 939, 514, 92, 565, 2277, 505, 1511]. Park [2664, 720, 1973, 2853, 1622, 2078, 130, 802]. Parke [1058]. Parker [2759, 193]. Parkin [2834, 492]. Parks [562]. Parrett [2400]. Parrish [2740]. Parrots [1511]. Parshall [2989, 3257, 168]. Parson [2735]. Parsons [718]. Part [2027, 1925, 123, 708, 75]. Particle [809, 957, 3427, 3545]. Particles [1873, 75]. particularis [1222]. Particulars [612]. particulière [3396]. Partington [3757]. Partnership [2062]. Parts [3620]. Pasachoff [64]. Pascal [2981, 2733, 3361, 506, 2673, 2015]. pasión [2829, 3604]. Pasnau [615]. Passage [1070]. Passages [252]. Passing [2170]. Passion [2502, 1876]. Passionate [3758, 135]. Passions [1307, 1466, 1711, 1740, 761, 186, 3585]. Past [1046, 3287, 774, 3252, 3478, 1836, 856, 14]. Pasteur [3002, 1349, 2988, 2177]. Pastrologla [2252]. Pasts [1008]. Pat [1005]. Patent [676, 2519]. Patenting [3119]. Patents [2870, 434]. Paternalism [1434]. Path [1641, 2772, 128, 3051, 2541]. Pathological [2299]. Pathologie [1279, 2372]. Pathologies [3663]. Pathologische [3747]. Pathology [375, 901, 3407]. Paths [206]. Pathways [2327]. Patients [3192, 1273, 2188, 693, 1435]. patologica [610]. Patrice [364, 298, 354, 2811]. Patricia [1175, 513, 2658, 726, 209, 2813, 893, 2251]. Patrick [1025, 454, 2855, 1126, 1586, 2859, 1604, 3473, 1833, 1380, 2984, 956, 3450]. Patrie [3267]. Patriotism [434]. Patriots [49]. Patronage [2048, 3653]. Patrons [2647, 2884, 159]. Patterns [1270, 2327, 2615]. Patton [3338]. Pau [985]. Paul [843, 745, 1084, 160, 375, 609, 3638, 1393, 608, 407, 380, 186, 720, 891, 2926, 3726, 414, 97, 2641, 42, 795, 371, 698, 113, 552, 420, 3775, 2055, 476, 3763, 439, 217, 2292, 2078, 2427, 1855, 3085, 1026, 2061, 1857, 2637, 1597, 2533]. Paula [912]. Pauli [2301, 2758]. Pauli/Jung [2301]. Pauling [1748]. Paulo [2561]. Pauly [1118, 3543, 2689]. Pauper [3062]. Pauwels [3222]. Paved [847]. Pavel [2304]. Paviland [3514]. Pavlov [2290]. Pawliczek [3761]. Pay [2519]. Paz [3056]. Peace [3264, 576, 384]. Peacemaker [1748]. Peale [275, 711, 711]. Pearce [44]. Peard [903]. Pearly [602]. Pearson [2199]. Pedagogical [957]. Pedagogy [2723]. Pederf [2423]. Pediatrics [599]. Pedley [2539]. Pedro [2263, 1500, 1496]. Peebles [1347]. Peenemunde [742]. Peer [1616]. Pegasus [1819]. Peirce [1106, 808, 808, 1773]. Peiresc [3364, 886]. Pelayo [18]. peligro [1399]. Pelis [3002]. Pelling [2048]. Pelouze [48]. Peña [1949]. pendant [1735]. Pendulum [2221]. Penelope [1922]. Penetrating [1568]. Penicillin [3296]. Penini [1583]. Peninsula [1493]. Penn [3488]. Penny [2288]. pensamiento [714]. pensant [2257]. Pensée [3739]. Penser [2896]. pensiero [3102]. People [910, 1009, 1278, 3133, 1770, 1425, 3430, 1943, 1348, 94, 422]. Pera [679]. Peral [1952]. Perception [521, 1147, 1364, 1524]. Percezioni [3615]. Percival [1112]. perelome [2943]. Perez [622, 165]. Perfect [2599, 2656, 3695, 3518]. Perfection [3469, 1904, 890]. Perfectly [2932]. Performing [3744]. Perga [3647, 1684]. Perifano [701]. Perils [3341, 3523]. Period [3261, 1105, 583, 1885, 1370, 767, 1377, 3547]. Periodic [3014, 2357, 1384]. Periodical [3415, 2473]. Periodicals [2473]. Perkins [723, 82]. Perkovich [650]. Perlin [538]. permanences [3639]. Permanent [913]. Permissible [928]. Pérouse [2275, 3652]. persischer [607]. Person [1279]. Personal [2697, 30, 1539, 1767, 1628]. Personalcomputers [1042]. Personalities [3774, 3667]. Perspective [1993, 3467, 3542, 2004, 378, 2420, 3541, 2128, 1968, 517, 1835, 1778, 1994, 3110, 1617, 3716, 2596, 1887, 2702, 2354]. Perspectives [1323, 1176, 1434, 1997, 2680, 3710, 2306, 774, 3094, 1193, 1564, 1236, 744, 248, 408, 552, 689, 680, 2574, 3720, 1541, 2509, 1633, 679, 2723, 271, 22, 1786, 1988, 3108, 3776, 975, 3265, 1211, 115, 1542]. Perspectivism [3529]. Perspektiv [504]. Persuasion [1915]. Peru [2174, 60, 2174]. pervenstvuiushchego [1354]. Pesic [2198]. Pesticide [1068]. Peter [1639, 2027, 679, 111, 3638, 3475, 3760, 305, 2932, 20, 358, 753, 3509, 2179, 581, 2721, 587, 438, 296, 230, 2075, 2254, 3072, 2720, 3010, 1627, 1392, 1629, 3043, 2503, 987, 3093, 1954, 2833, 1348, 1098, 2373, 2299, 2496, 1429, 886, 2198, 2611, 1805, 961, 856, 1930, 972]. Peterburge [2191]. Peterburgskoi [1354]. Petersburg [4]. Peterson [2215, 375]. Petherbridge [594]. Petra [3486]. Petrarca [1492]. Petrolia [916]. Petrology [1877]. Petrus [352, 2541]. Petterssen [1261]. Petty [2894]. Pfizer [1896]. Pflock [1045]. Ph.D.’s [3750]. Phaenomena [3037]. Phaer [481]. Phage [2856]. Phantastik [3757]. Pharaohs [1389]. Phares [354]. Pharmaceutical [1237, 2314, 2633, 553, 3190]. Pharmacology [3079]. Pharmacy [2744]. Pharmazie [347, 1912, 3035, 1478]. Pharmaziegeschichtsschreibung [587]. Phenomena [3143, 823, 422]. Phenomenalism [1110]. Phenomenes [32]. Phenomenology [3442]. Phenomenon [3023]. Phil [213, 2642]. Philadelphia [128, 1933, 1240]. Philanthropy [1737, 3407]. Philip [489, 3166, 62, 218, 1868, 3001, 2626, 739, 162, 291, 2510, 2728, 1118, 1561, 3543, 2020, 2731, 2019]. Philippe [1400, 2887, 2921]. Philippines [3663]. Phillip [2324, 408]. Phillips [2781, 2624, 294]. philosophe [3060]. Philosophen [3679]. Philosopher [2070, 2157, 2309, 53, 3731, 3092, 276]. Philosophers [3231, 2728, 186, 2486]. philosophes [3571]. Philosophic [1583]. Philosophical [1434, 679, 3029, 3288, 1162, 491, 678, 186, 1382, 1777, 956, 2606, 408, 71, 3555, 3740, 75, 2454, 3470, 1886, 793, 2541, 664, 2608, 1427, 418, 3441]. Philosophie [3063, 3685, 3657, 162, 3763, 2343, 3316, 826, 875, 2973, 2312, 1904, 743, 1190, 3159, 3400]. Philosophien [2831]. Philosophies [667]. philosophiques [469]. philosophische [2407]. Philosophy [1160, 1179, 1181, 1316, 1326, 944, 948, 1633, 1204, 418, 3420, 72, 1967, 5, 3496, 1888, 3646, 1660, 2888, 2791, 3389, 3325, 1502, 3231, 2728, 1266, 2720, 668, 2385, 2974, 870, 3201, 3687, 3123, 3122, 2448, 575, 169, 3317, 2889, 459, 2606, 1234, 411, 1773, 1110, 1011, 2605, 767, 2383, 180, 484, 2167, 1503, 3529, 3120, 3486, 3666, 2574, 3416, 2100, 3301, 3512, 1556, 3004, 276, 2578, 2545, 2416, 365, 2402, 475, 3052, 568, 1428, 19, 3012, 3011, 1375, 2251, 2862, 2138, 1097, 3109, 2947, 1704, 3398, 1699, 416, 227, 876, 1969, 3230, 3781]. Phineas [902]. Phoenix [2490]. Phone [1282]. Phonesthemic [2867]. Phonograph [764]. Photogenic [1186]. Photograph [47]. Photographic [1265, 2178, 1598, 1557]. Photographs [1526]. Photography [400, 3351, 3427, 682]. Phrenology [2203]. Phyllis [1107]. Physic [986]. Physical [3467, 631, 1808, 756, 3095, 728, 1250, 2916, 681, 2420, 440, 75]. Physician [1592, 1273, 995, 3042, 2487]. Physicians [1122, 2048]. Physicist [2681, 2683, 2772, 53, 79]. Physicists [3300, 3259, 1079, 2779, 593]. physico [586]. physico-mathematica [586]. Physicorum [2729]. Physics [1975, 1204, 418, 2472, 1967, 5, 1968, 2582, 301, 2627, 648, 1509, 717, 526, 3580, 1382, 333, 2147, 957, 770, 442, 2605, 3011, 1051, 143, 303, 2365, 3470, 3437, 2459, 3416, 3427, 2529, 3512, 2207, 112, 2729, 2940, 2472, 2752, 75, 2939, 1967, 2085, 668, 2565, 756, 529, 441, 1839, 3545]. Physik [1373, 3671]. Physiker [2840, 1516]. Physio [181]. Physio-Medicalism [181]. Physiognomik [346, 121]. Physiognomy [1128]. Physiological [2334]. physiologique [3248]. physiologiques [490]. Physiology [452, 2290, 1187, 487]. physique [1471]. Physis [1219]. Picatrix [2151]. Piccardi [3642]. Piccolino [2352, 2418, 3138]. PICES [2783]. Pickering [3241]. Pickled [1665]. Pickstone [1163]. Pico [2481]. Picotte [806]. Pictura [3566]. Picture [3550, 1926, 2525, 198]. Pictures [2583, 1608, 834, 2801]. Picturing [1867, 2538, 111]. Pie [1130]. Pieces [2057]. Piero [3730, 3146]. Pierre [761, 2534, 2426, 1344, 582, 2915, 2888, 2738, 3052, 798]. Pierre-Noël [2738]. Pierry [3572]. Pieters [2860]. Pietro [2996, 779, 3060]. Pietsch [2979]. Piety [1450]. Pilavas [188]. Pill [219, 1153, 1153]. Piltdown [2072]. Pimentel [2264, 714]. Pinch [2788]. Pinel [1400]. Pinell [298]. Piñero [1820, 2142, 449, 2386]. Pingree [2961, 462]. Pink [1959]. Pinnick [2577]. Pinpoint [1040]. Pioneer [1870, 1030, 2162, 3005, 3580, 3318, 2901, 203, 1145]. Pioneered [394]. Pioneering [938, 2180, 631, 3750]. Pioneers [894, 2850, 213, 1212, 59, 1953, 888]. Pioneras [732]. Pionierzeit [3330]. pionnier [2886]. Pious [5, 1237]. Pipe [3073]. Piper [1843]. Pipes [1838]. Piszkiewicz [1415]. Pita [129]. Pittier [2922, 200, 1004]. Place [1437, 378, 903, 2461, 2147, 722, 2286, 3586, 1796, 2729, 2934, 2450, 1720]. Places [2661]. Placing [3369, 3249]. Plague [3107, 940]. Plagues [1241, 640]. Plains [1278, 3474]. planches [2979]. Planck [1052, 3671, 170, 1517]. Planet [894, 2757, 3228, 682, 2457, 406]. Planeta [2263]. Planetaria [2858]. Planetary [2140, 244, 1217, 123, 535]. Planets [2415]. Planned [815]. Planning [2825]. Plans [872]. Plant [375, 1941, 2813, 576, 3579]. plantarum [163]. Plantas [1820, 768]. Plants [2870, 3579, 2540]. Plasmasphere [83]. Plastic [81]. Plastik [3436]. Plate [2330, 834, 1296]. Plateau [1550]. Plates [1557]. Platforms [2299]. Plato [1362, 1670, 904, 2969, 1090]. Platonic [872]. Platonist [2483]. Platt [2690]. Platypus [167]. Plays [1566, 710]. Please [2805]. Pleasure [710, 834]. Plinio [146]. Pliny [863]. plis [161]. Plitzner [54]. Plot [3768]. Plotkin [2505]. Plowshare [3273]. Plum [3424]. Pluralism [3529]. Plurality [1113]. plus [354, 3388]. Pluto [3228]. Pneumonia [2914]. Pnina [680]. Pockenschutzimpfung [56]. poder [3053]. Podolsky [555, 2914]. Podvig [2304]. Poe [2068]. Poems [15, 1532]. Poet [2200]. Poetics [1263, 1232, 3226, 3674, 3313, 3305]. poétiques [494]. Poetry [818, 2385, 964]. Pogliano [779]. Poincaré [2126, 2877]. Poincaré’s [2496]. Point [2072, 3348]. Points [799, 1085]. Poirier [2915]. poiskakh [1267]. Poison [2333, 3328, 3770]. Poisoned [1529]. Poisoning [1758, 737]. poisson [3267]. poissons [2979]. Polanyi [1428, 3092]. Polar [1611]. polarforskning [504]. polarhistorie [2493]. Polarity [3546]. Pole [381, 542, 212]. Polemical [3567, 1100]. Polenberg [2216]. Poles [3269]. Policey [2291]. Policies [3680]. Policy [946, 1746, 3353, 1144, 2631, 1047, 1755, 2909, 93, 2657, 379, 307, 3212, 2364, 319, 3082]. Polio [3085]. Politic [2891]. política [714, 2561]. Political [1577, 221, 3754, 2520, 3784, 305, 957, 2699, 3360, 2851, 2867, 2057, 75, 1866]. políticas [2174]. Politics [1778, 2111, 2502, 2472, 2767, 2803, 3613, 2927, 3346, 85, 3293, 545, 3303, 643, 3338, 2196, 1855, 1412, 2610, 3359, 2924, 1756, 3398, 3365, 1707, 1532, 1069, 2836, 1768, 1728, 2089, 3522, 1127, 1304, 1148, 3593, 1722, 317, 1963, 1150, 2640, 201, 3440, 2446, 1549, 1011, 2651, 6, 3779, 3264, 2002, 926, 1566, 680, 331, 492, 503, 1879, 2931, 410, 1866, 2529, 2186, 3150, 1252, 2632, 1849, 3680, 284, 1548, 1504, 1299, 3353, 1770, 1241]. Politik [2117, 2184, 2616, 3761, 3418]. politique [2620, 502, 2706, 701, 1736]. Politiques [3321]. politische [3436, 2838]. politischen [2370, 3672]. Polity [2484]. Polkinghorne [2511, 700]. Pollock [3086]. Pollution [3072, 1343, 914]. Polly [562]. Polonium [3618]. Poluhowich [730]. Polygraphs [2804]. Polymer [3077, 3277]. Pomar [1820]. Pomata [2893, 693]. Ponceau [1404]. ponctualité [3310]. Poor [1433, 2268, 1755, 819]. Poovey [28]. Pope [477]. Popkin [2396]. Popper [1011, 419]. Popplestone [185]. Popular [1044, 2054, 2985, 3539, 3010, 1166, 1868, 1027, 432, 2550, 3540, 3541, 3674, 2625, 1932, 3266, 2982]. Popularisation [453]. Popularization [3062]. Popularize [1556]. Popularizers [3413, 330, 3774]. Population [2472, 3252, 1649, 2266, 1252, 1487]. Populism [2625]. por [2829]. Pork [643]. Pormann [3043]. Porphyry [353]. Porras [255]. Porter [2049, 2199, 805, 1341, 593, 450]. porteurs [2721]. Portolés [732]. Portrait [203, 1817, 1339]. Portraits [3622, 1084, 2093]. Portraiture [1592]. Portrayal [247]. Portugal [1787]. Portuguese [1787]. Poser [2171]. Positive [1475]. positivisme [3314]. positivistes [2452]. Possibilities [3287]. Possibility [1098]. Possible [2663]. Post [946, 85, 3293, 2875, 3370, 3283, 195, 1256, 130, 3504, 415, 3352]. Post-1750 [2875]. Post-Cold [3352]. Post-Darwinian [3504]. Post-Fordist [3370]. Post-Mao [85]. Post-modernism [415]. Post-Revolutionary [3293]. Post-War [195, 946]. Post-World [1256]. Postcolonial [3502, 2861, 534]. Postel [1310]. Posterior [2575]. Posters [1578]. Postgenomic [3195]. Posthuman [323]. PostHumanities [3305]. Postmodern [1887]. Postponed [1093]. Postpositivism [2313]. Postrevolutionary [3364]. Postscript [2463, 3320]. Poststructuralist [2652]. Postwar [2647, 2565, 957, 2234, 3189, 3354]. Poteat [859]. potente [2173]. Potter [2999, 1590, 609, 667]. Potthast [2103]. poudres [364]. Poultry [2225]. Pour [3361, 364, 2971, 1014]. Poverty [1528, 640, 1755]. Powell [803]. Power [1058, 1991, 1616, 533, 57, 2461, 2169, 3440, 1123, 2441, 1720, 137, 552, 2625, 1566, 834, 279, 2773, 2604, 1884, 390, 2562, 281, 537, 3677]. Powerful [3393, 2699]. Powerless [7]. Powers [1837]. Pox [1830]. pp [2472]. Prabhakar [1080]. Practical [1634, 3281, 3313, 3523, 3677, 46, 2980, 3419]. Practice [905, 1388, 2491, 2723, 2548, 3401, 577, 901, 1557, 2907, 2082, 320, 1360, 3450, 3526, 352, 262, 110, 1589, 2310, 2494, 3531]. Practices [1787, 3222, 845, 3662, 3735, 1777, 3680, 2952, 3746, 680, 1606, 2677, 942, 1325, 3581]. Practika [1498]. Practitioner [2798, 2463]. Practitioners [2048]. Praed [904]. Praetorius [2895]. Prag [758]. Pragmatism [3336, 1427]. Praises [3624]. Prakash [246]. Praktische [1470]. Prandtl [2814]. Pranks [2215]. pratiche [3735]. praticque [2408]. Pratik [2677]. Pratique [2988, 1735]. Pratiques [2091, 1691]. Pratt [1078]. Praxis [67, 1089]. Prayer [1996, 3485]. Pre [453, 2808, 2066, 3750]. Pre-1750 [2808]. Pre-1940 [3750]. Pre-Independence [453]. Pre-Statistical [2066]. Précédé [3348]. Preceptum [462]. Precision [1446, 1281, 2730]. Precolumbian [868]. Predators [1550]. Predecessors [1425]. Predict [971]. Predicting [2748]. Prediction [3335, 3520]. Predictions [1363]. Preface [3157]. Pregnant [2052]. prehistoire [142, 3267]. Prehistoric [1215, 2149, 1648]. préhistorien [2620]. Prehistory [973, 117, 3760, 2145, 3555, 961, 685]. Prelude [2436]. Preludes [2621]. Prematurity [2224]. Premier [790, 1854]. Premodern [1321, 2972]. Prendcourt [1921]. Preoccupations [1785]. Preparatoria [692]. Prescott [322]. Prescribing [2094, 3347]. Prescription [3111]. Present [3476, 1167, 571, 1974, 2906, 344, 694, 343, 857, 1836, 856, 2118, 3712]. presented [3641]. Preservation [1556]. Preserving [1147]. Presidio [130]. Press [2472, 3415, 1697, 3062]. Presso [2022]. Pressure [3526]. Preston [2670]. prevencion [255]. Prevention [1948]. prezidentov [1354]. Price [1896, 3523, 557]. Price/Rod [1896]. Priesner [617]. Priestley [2419]. Primary [2808, 421]. Primate [2853, 936]. primatologische [2560]. Primeval [930]. primitifs [3388]. Primitive [1833]. primo [2252]. Prince [353, 2254]. Principe [1591, 3158]. Principia [1699, 484, 3320]. principio [3563]. Principle [266, 2758, 2758, 3237]. Principles [1699, 941, 879, 697, 2843]. Pringle [2751]. Print [940, 3247, 109, 1196, 1493]. Printing [1598, 3387]. Printmaker [702]. Printmaking [12]. Prints [1578, 12]. Prinzip [3563]. Prinzipien [1373]. Prinzipienphysik [2194]. Prison [220]. Pristina [1490]. Pritchard [913, 247]. Privacy [3588]. Private [2556, 3329, 3736, 2810, 3264, 2548]. Privileges [2870]. Prix [1404]. Prize [2472, 1896, 189, 2992, 2394, 2527, 3625, 1875]. Prizes [1979, 2472, 2472]. probabilité [1471]. Probability [3029, 2015, 889, 1613, 2707]. Problem [2614, 2936, 683, 858, 379, 2931, 3641]. Problems [3122, 3765, 28, 1630, 87]. Proceedings [3100, 1546]. Process [92, 2447, 3298, 1547]. Processing [3144]. Procreate [2243]. Procreation [1321]. Proctor [548]. Prodigies [1707]. prodigieux [1718]. Produced [3303]. Producing [111, 901]. Production [3359, 1036, 3634, 3624, 291, 2332, 1016]. Productive [3471]. Productivity [545]. Profession [1327, 2959, 125, 2594, 1017, 1360, 2555, 1154]. Professional [2319, 2172, 1256, 3693, 320, 3523, 1032]. Professionalization [1046, 1343]. Professionals [1273]. Professions [2152, 1466]. Professorenkarriere [2189]. Professors [2113]. Profile [1605]. Profiles [2353, 1175]. Profiting [3077]. progetto [3403]. Prognostica [352]. Program [2393, 3093, 3424, 3450, 84, 1309, 943, 1767, 2519, 2567, 2508, 1294]. Programs [333, 316]. Progress [2832, 2734]. Progressive [2318, 859, 1127].

Progressive-Era [859]. Prohibition [2693]. Project [216, 1697, 3262, 3726, 2658, 397, 3273, 3352, 3200, 892, 408, 1541]. Projectile [2024]. Projection [3137]. Projects [1984]. Projekt [309]. Prolegomena [1684]. Proliferation [650]. Prologue [3390]. prolongación [498]. Promethean [2599]. Prometheans [1001]. Prometheus [2472, 3519, 460]. Prominent [2420]. Promise [3317, 1118, 2859, 3512]. Promising [3767]. Promoting [834, 409]. Promotionen [1519, 3477]. Pronatalism [3263]. Proof [2198]. Properties [1688]. Property [2870, 3119]. Prophecy [1504]. Prophet [723]. Prophylaxis [2913]. Proportion [1160]. propos [3348]. propres [518]. propriété [2088]. Proskouriakoff [2095]. Prospect [2449]. Prostitute [2498]. Protection [1620, 928, 2492, 1881]. Protections [2870]. Protein [1747, 3200]. Proteins [9, 1475]. Protest [3525]. Protestant [3055, 270]. Protestantism [27]. Proteus [3206, 2687]. Proto [355]. Proto-Scientific [355]. Protogaea [3656]. Proust [2629, 512]. Proust’s [999]. Prove [1005]. Proved [2500]. Providence [1707]. Provinces [2016]. Provincial [1336]. Provincialism [1779]. Proving [3273]. proyecto [768, 2430]. Prozac [2094]. Prüfungsprotokolle [1519]. Prüll [2372]. Prussian [1598]. prussienne [3694]. Pseudoscience [823]. psikhiatrii [2499]. Psoriasis [2144]. Psyche [1856, 3293, 258]. Psychedelic [3606]. Psychiatric [321, 2494]. Psychiatrie [2284, 188]. Psychiatrische [635]. Psychiatrists [1136, 2219]. Psychiatry [3606, 1855, 1008, 1120, 3334, 2190, 2908, 3560, 2494, 552, 3190, 3435]. Psychic [823]. Psychoanalyse [1856]. Psychoanalysis [738, 2075, 186, 2069, 62, 186, 565, 3067, 1857, 813]. Psychoanalytic [186, 2505]. psychoanalytischen [1734]. Psychological [258, 2507, 2138]. Psychologists [320]. Psychology [2195, 401, 2433, 1848, 378, 1266, 40, 2594, 405, 3560, 604, 1187, 91, 223, 3057, 1946, 185, 2497, 2937, 3710, 2594]. Psychopathology [2799]. Psychopharmacology [404, 1543]. Psychophysiological [1187]. Psychosocial [3699]. Psychotherapy [655, 1274]. Ptolemy [462, 836, 2950, 1683, 463, 462]. Ptolomei [462]. pu [3432]. pubblico [2282]. Public [1025, 946, 2556, 402, 3166, 2054, 1859, 2285, 3539, 1528, 3192, 2810, 1948, 2172, 1148, 319, 2631, 2003, 3765, 2982, 598, 432, 3610, 1880, 550, 2845, 3264, 2819, 2832, 2625, 1053, 955, 2374, 184, 3187]. Publication [1000]. Publications [940, 712, 2164, 3642, 3473, 1372]. publique [1710, 825]. Published [1993, 596]. Publishers [3162]. Publishing [2550, 591]. puertos [2263]. Pupils [3141]. Puppet [3261]. Purchase [3756]. Pure [160, 3167, 410]. Purgatory [414]. Purity [1582, 2488]. Purplemountain [3343]. Purpose [268, 1971]. Purposes [1636]. Pursglove [531]. Pursuing [2725]. Pursuit [3048, 1307, 833, 1304, 3558, 2193, 46]. Pursuits [496, 421]. Pushbutton [2190]. Pushing [1236]. Put [3499]. Putting [378, 1427, 2450]. Pylkkanen [78]. Pyne [1346]. Pyramid [842]. Pyramus [2547]. Pythagoras [2881, 2601].

Q [2021]. qi [3261]. Qibla [3486]. Qing [2158, 2412, 3306, 1096]. Qingdao [307]. Qu [3275]. Quack [1578, 1578]. Quaderns [2430]. Quadrants [1807]. Quadratum [1634]. Quaker [1338]. Qualifikationswesen [3477]. Qualitates [163]. quan [3200]. quanta [1619]. Quantification [2596]. Quantifying [2266]. Quantitative [1663]. Quantum [73, 72, 2681, 1967, 1968, 2582, 2571, 228, 529, 75, 1537, 3437, 3427, 1407, 530]. Quark [2445, 2571]. Quarters [3018]. Quarto [1054]. Quaternary [3675]. Queen [398]. Queen’s [2993]. Queen’s/Associated [2993]. Queer [2303]. Quellen [642]. Quenet [3494]. Quest [2697, 1957, 2490, 1297, 2568, 1948, 1955, 2599, 405, 535, 1080, 2417, 403, 3173, 1005, 122, 2153]. Question [2332, 899]. Questioni [2729]. Questions [352, 186, 647]. Quetchenbach [818]. Quick [594]. química [3054, 2366, 732]. Quincy [236, 3758]. Quine [2313]. Quinlan [3598]. Quintessence [2918]. Quinton [624]. Quotations [2643].

R [2944, 2945, 2622, 247, 51, 2429, 794, 182, 1349, 853, 2058, 485, 207, 2427, 1038, 108, 290, 3340, 90, 2050, 918, 2892, 3005, 512, 3014, 507, 2850, 1620, 1852, 53, 747, 1917, 2673, 2003, 1966, 2371, 2599, 278, 2404, 2634, 2982, 805, 316, 20, 3170, 172, 2637, 265, 371, 1589, 408, 669, 2966, 2355, 3015, 2751, 1035, 360, 926, 312, 264, 2774, 590, 816, 215, 350, 741, 272, 705, 599, 1628, 1824, 2786, 3277, 3196]. R&D [314]. R. [1852, 1054, 2759, 857]. rêve [798]. Rabinow [414]. Rabusseau [3495]. Raby [1392]. Race [1870, 348, 1843, 2137, 1206, 632, 1744, 2378, 304, 1528, 1419, 1197, 3593, 2902, 1271, 2127, 2003, 858, 383, 2777, 2303, 516, 3272, 3694, 2446, 3663, 722, 2214, 2845, 1943, 2626, 1621, 120, 1108, 2032, 2160, 2377, 2289, 2638, 903, 2708, 1195]. Races [1833, 2537]. Racial [3342, 3341, 2003, 2053, 3778]. racines [3649, 3320]. Racism [1870, 899, 566, 2708, 2501, 1422, 120]. Radar [646, 917]. Radder [3096]. Rader [2305]. Radiance [533]. Radiation [928, 1185, 3618, 731]. Radiation-Induced [1185]. Radical [222, 997, 3258]. Radio [3681, 3774, 918, 1027, 2297]. Radioactive [1768, 3431, 3770]. radioactivité [1111]. Radioactivity [2002, 64]. Radiochemist [2124]. Radioepidemiologic [1185]. Radium [2682, 2002]. Rado [186]. Rafael [884, 2742]. Rafaelli [2252]. Raffael [3736]. Raffaella [3141, 2267]. Raffles [2035]. Rafinesque [1404, 2421, 1851, 2353]. ragioni [226]. Ragnar [1178]. Ragone [209]. Ragouet [3361]. Rahel [520]. Rahiq [3486]. Rails [1609]. Railway [180]. Railways [273, 180]. Raimund [212]. Rain [1180, 2767, 1038, 1873, 1036, 1180]. Rainbow [1811]. Rainbows [1811]. Rainer [3477]. Rainforest [2119]. raison [2733]. raisonnements [2671]. Rajala [1036]. Ralph [553]. Ramage [608]. Ramaswamy [2760]. Ramberg [2373]. Ramón [2938, 3054, 1399]. Ramsden [3659]. Rand [186, 2782]. Randal [859, 913, 1601]. Randall [854]. Randles [478]. Rane [2352]. Range [1939]. Rania [2259]. Rank [2075]. Ransome [2880]. Raphael [663, 1527, 419, 3495, 2537]. Rapporti [2401]. rapports [2343]. Rash [801].

Rashed [469]. Rashid [319]. Rasmussen [1425, 198]. rassenhygienischen [1964]. Ratcliff [3735, 627]. Ratio [1089, 3727]. Rational [2083, 224, 497]. Rationalism [1938]. Rationality [224, 3156, 2381]. Rationalization [2479]. Rationalized [3695]. Rationalizing [2083]. Rauch [996]. Raúl [768]. Raum [3331, 171, 1013, 3657]. Raw [901]. Ray [2418, 1140]. Raymond [270, 927, 1565, 1811]. Rayner [114, 114]. Rayner-Canham [114, 114]. Raynor [296]. Rays [1475, 192, 2233]. Rayvon [2562]. razza [777]. razzismo [777]. re [480, 2162]. Re-examination [2162]. Reaching [1205, 1383]. Reaction [2384]. Reactions [1768, 3577]. Read [2476]. Reader [3757, 107]. Readers [3651]. Readership [2322]. Reading [468, 2473, 2341, 47, 2075, 989, 984, 254, 148, 484]. Readings [2137, 1192, 983, 124, 1986]. Reagan [1309]. Real [2503, 810, 3434, 1164]. Reale [2044]. Realigning [2299]. Realism [3527]. Realitaten [1985]. Reality [1968, 61, 138, 2382, 204, 2511]. Reaper [546]. Reardon [2638]. Rearranging [2916]. Reason [2678, 974, 853, 2479, 2049, 2790, 1491, 795, 2642, 3190, 246, 160, 2104, 1240, 3249, 667]. Reasonable [2699]. Reasoning [440, 3411, 3156, 503, 357]. Reassessment [1151]. Rebecca [1374, 76, 2256, 1594]. Rebel [1938]. rebelles [2091]. Rebels [3636]. Rebert [912]. Rebound [3206]. Recasting [2766]. Reception [2683, 1000, 3760, 3757, 2955, 2273, 1390]. Receptions [589]. recherche [1702, 494, 1710, 1735, 1190]. Recherches [1907]. rechtlichen [274]. rechtvaardigheid [2676]. récit [2988]. Reclaiming [1777, 2651]. Reclassification [108]. Recollections [526]. Recombinant [1148]. Reconceiving [2856]. Reconciling [3164, 1288]. Reconfiguring [415]. Reconsidered [835]. Reconsidering [1428, 1990]. Reconstructing [3083, 600]. Reconstruction [2693, 2825, 3189, 2743, 3709, 1750, 1362]. Record [511, 539, 1041, 444]. Recording [1030, 539]. Records [97, 979, 296, 2039, 1554, 2314]. Recreating [3505]. recueil [354]. Red [1892, 2457, 3514, 3519]. Redefining [2410]. Redhead [3095]. Rediscovering [3661]. Redon [2911]. Redondi [779]. Redrawing [2954]. Reductionism [3097]. Reed [378]. Reeds [3232]. Reefs [2281]. Reel [659]. Reenacting [3661]. Rees [3659]. Reexamining [2648]. Reference [941, 338, 592, 1055]. Refiguring [2745]. Refining [608]. Reflect [644]. Reflection [3701, 1250, 1039, 2110]. Reflections [1725, 2098, 3516, 3539, 1946, 1153, 3367, 3540, 75, 2957, 1767, 3382, 3781, 1022, 80]. Reform [299, 3336, 1290, 1722, 957, 3668, 286, 2690, 1166, 345, 2773, 3769, 3709]. reformas [2430]. Reformation [1450, 575, 3055]. réformatrice [716]. Réforme [613]. Reformer [806]. Reformers [1963]. Reforming [1751, 3004]. Reforms [980]. Refounding [2718]. Refraction [1250, 3041]. Refracts [2629]. Refugees [3181]. Regal [2289]. regard [1243]. Regime [432, 1460, 1202]. Regimen [352]. Regimes [2637]. Regiomontanu [3730]. régisseurs [364]. Regnault [48]. Regourd [3074]. regreso [60]. Regulation [918, 1529, 3779, 816, 1036]. Rehabilitated [662]. Rehbock [1561]. Reice [1035]. Reich [2117, 742, 1739, 1274, 738, 3753]. Reid [1389].

Reidy [1974]. Reign [3416]. Reigning [333]. Reill [2611]. Reimagining [851]. Reims [361]. Reims-Champagne-Ardenne [361]. Reinforced [2434]. Reingold [2395]. Reinhard [642, 2563, 1799]. Reinhardt [2916]. reinos [2263]. Reinterpretation [887]. Reinventing [1313, 3718]. Reisch [2791]. reivindicación [3148]. Rejection [1413]. Rejetées [2091]. Rekonstruktionsversuche [3406]. Rektoratskommission [2143]. relaciones [3053]. relacões [2561]. Related [1851, 1397, 66, 97, 3724, 2057]. relatif [2877]. Relation [418, 1631, 3560]. Relations [785, 891, 969]. Relationship [935]. Relationships [1273, 84]. Relativistic [1975]. Relativity [2991, 1869, 997, 3416, 2955, 2128, 2902, 3512]. relato [2474]. Release [1882]. Relief [1433, 1994]. Religion [1165, 632, 1719, 2449, 3325, 648, 2967, 3419, 3379, 3638, 1168, 2669, 1629, 3593, 3307, 722, 3619, 1292, 1932, 492, 1288, 1836, 270, 2921, 1492, 26, 935, 2642, 600, 3276]. Religiosität [2831]. Religious [2882, 605, 2077, 1307, 2621, 1450, 40, 3401, 2890, 482, 3555, 1227]. Relocating [3223]. Relocation [564]. Remains [3066]. Remak [3178]. Remaking [128, 3033, 392, 2507, 3270]. Remarkable [1865, 2297, 2633]. Remarks [577]. Remedios [884]. Remington [2782]. Reminiscences [384]. Remke [3646]. Remmert [360]. Removing [3099]. Remy [137]. ren [3200]. Renaissance [2118, 2534, 702, 1504, 3234, 985, 1904, 3391, 3475, 356, 976, 1495, 2253, 3311, 2341, 2344, 123, 613, 786, 1586, 2152, 152, 2043, 1691, 2603, 3218, 3313, 3488, 1499, 3651, 2970, 2157, 3487, 2481, 612, 2153, 3364, 5, 3075, 3569, 873, 789, 2042, 3488, 2033, 3236, 1311, 1998, 983, 1695, 1490, 3386]. Renate [170, 230]. Renaudot [1832]. Renault [2312]. rencontre [3320]. Rene [123, 155, 1243, 1702, 1084, 491, 627, 3153, 2467]. Renewal [1369]. Renn [3512]. Rennie [2077]. renovación [2742]. Renovating [3664]. Renowned [644]. Repair [2773]. Repertorio [2729]. Repertorium [1827]. Replacement [3199]. Replacing [3369, 222]. Replication [1623]. Replikation [2183]. Reply [2010, 1897, 841, 2592, 2589, 2469, 2012, 839, 1657, 2471, 1804, 1802, 1902, 1659]. Report [2393, 1862]. Reportorio [1585]. Representation [1438, 888, 2332, 2473, 1520, 1557, 3146, 2342, 3427, 230, 982, 582, 3471]. Representational [3222]. Representations [1632, 785, 2457, 524, 2247, 1107, 494]. Representing [721, 2749]. repressii [1267]. Reproduction [88, 3160, 87, 3263]. Reproductive [3548, 1963, 3777, 2656, 392, 1963]. repubblica [691]. Republic [1640, 1420, 97, 3430, 2273, 3364]. Republican [2701]. Republics [3216]. Reputation [1556]. Reputations [2968]. Require [3751]. Rereading [2951]. Rescher [2265, 2862]. Rescue [471]. Rescuers [2275]. Rescuing [3779]. Research [2697, 2120, 2431, 2168, 216, 845, 1626, 3261, 2557, 2418, 2559, 3271, 1404, 377, 1881, 833, 2727, 2808, 2725, 333, 315, 2478, 1797, 1520, 2305, 3375, 3680, 1739, 2966, 2442, 2002, 2090, 1187, 962, 214, 2186, 2782, 1424, 3344, 2420, 2002, 2661, 545, 1414, 2355, 260, 1850, 1063, 3346, 1666, 3345, 3610, 3717, 3779].

Researches [2943, 3261]. réseau [3157]. réseaux [716]. Resistance [2224]. Resnik [3523, 549]. Resonance [2581]. Resources [2875, 1966, 2626, 2637]. Respectability [3409]. responsabilidad [255]. Response [3448, 1651, 2682, 645, 694]. Responses [1180]. Responsibility [649, 2179, 663]. Rest [1062]. Restoration [492, 2773]. Restoring [1145]. Restructuring [537]. Résumé [3568]. Resurrecting [2652]. Retained [97]. Retardation [822]. rete [3161]. Rethinking [3222, 3455, 3688, 1337, 2453, 2382, 3528]. Retina [330]. Retooling [1772]. Retrying [2475]. Return [1432, 2174, 2104, 3038]. Reungoat [2894]. Revealed [510, 263]. Revealing [3485, 1140]. Revels [1203]. Reverent [330]. Reviel [1362]. Review [1029, 852, 994, 2173, 736, 2989, 3601, 650, 973, 3647, 1025, 183, 1576, 3194, 908, 2353, 2678, 3344, 3631, 2697, 691, 873, 1160, 3002, 3236, 2882, 3132, 2823, 1255, 3040, 3034, 761, 3572, 2405, 2852, 465, 489, 1254, 867, 1870, 2498, 2742, 2987, 602, 1263, 2898, 3269, 3755, 1484, 2185, 2933, 974, 2208, 2493, 2789, 942, 1174, 1178, 1179, 1180, 1181, 1308, 1309, 1310, 1311, 1312, 1313, 1314, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1323, 1324, 1325, 1326, 1327, 1328, 943, 944, 945, 946, 947, 948, 949, 950, 951, 1058]. Review [1059, 1060, 1061, 1062, 1063, 1064, 1065, 1169, 1170, 1171, 1172, 1173, 1175, 1176, 1177, 1432, 1640, 1644, 1645, 1778, 1779, 1780, 1781, 1782, 1783, 1784, 1785, 1786, 1787, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1441, 1442, 1443, 1444, 1445, 1446, 1632, 1633, 1634, 1635, 1636, 1637, 1638, 1639, 1641, 1642, 1643, 1889, 1998, 2105, 2106, 2107, 2108, 2109, 2110, 2111, 2112, 2113, 2114, 2115, 2116, 2117, 2119, 2120, 2121, 1890, 1975, 1976, 1977, 1978, 1979, 1980, 1981, 1982, 1983, 1984]. Review [1985, 1986, 1987, 1988, 1989, 1990, 1991, 1992, 1993, 1994, 1995, 1996, 1997, 1999, 3098, 3099, 3100, 3101, 3102, 3103, 3104, 3105, 3106, 3107, 3108, 3109, 3110, 3111, 2944, 2943, 2940, 2945, 2939, 2534, 3158, 2407, 2425, 3215, 855, 2622, 1501, 2536, 1952, 3553, 1856, 2284, 490, 1702, 1832, 3157, 3426, 348, 789, 1201, 3719, 1357, 738, 1843, 2695, 2735, 3342, 890, 2556, 1406, 800, 2431, 2307, 2811, 243, 637, 3063, 3436, 1753, 2076, 1158, 1840, 2175, 3776, 3071, 280, 2075, 1204, 812, 167, 849, 960, 2168, 2502, 2333, 3319, 3471, 402, 921, 44]. Review [768, 2101, 1553, 3235, 390, 1690, 621, 702, 1093, 1024, 2063, 2638, 2575, 3495, 3778, 1851, 73, 418, 242, 684, 1908, 247, 513, 2998, 2472, 924, 1121, 1752, 2215, 549, 268, 3314, 1044, 1849, 1405, 1974, 52, 1056, 1088, 2986, 3420, 1415, 2025, 2346, 862, 274, 1923, 156, 3467, 2614, 2555, 3164, 3600, 1030, 72, 1945, 45, 158, 1577, 3324, 1358, 2419, 51, 175, 562, 1018, 1021, 1665, 1939, 2661, 3240, 740, 191, 596, 2338, 3044, 2134, 2429, 2364, 1362, 2883, 1488, 2480, 2744, 2352, 2912, 2969, 2189, 794, 3753, 80, 1109, 1762, 2681]. Review [3070, 3515, 1685, 2038, 2434, 2767, 216, 743, 1704, 1736, 629, 1967, 940, 910, 2545, 446, 1504, 2761, 1951, 2739, 2260, 2885, 3243, 182, 731, 1349, 2060, 2137, 3180, 3496, 1371, 644, 907, 1716, 3220, 1218, 3730, 1206, 2146, 1807, 3142, 1122, 1264, 3640, 3222, 2704, 161, 712, 969, 1243, 3358, 1701, 2662, 3613, 2915, 2223, 2906, 1888, 2328, 3412, 3670, 2195, 2911, 3254, 3749, 1165, 1931, 783, 404, 3585, 2027, 2259, 632, 3166, 283, 843, 1689, 3646, 447, 2474, 1081, 1725, 464, 2936, 905, 1968, 2900, 853, 1815, 605, 1719, 3276, 3511, 1961, 2054, 2098, 3234, 1214]. Review [1492, 210, 221, 301, 569, 745, 1467, 1957, 49, 1523, 401, 595, 3250, 1664, 3549, 3380, 92, 398, 565, 2433, 3069, 2350, 3565, 3329, 2844, 2510, 202, 2834, 3756, 966, 2312, 911, 3507, 3175, 558, 1420, 2255, 2254, 2927, 665, 551, 2693, 3606, 1082, 985, 1494, 1828, 3308, 2698, 3335, 1084, 184, 2292, 2343, 1052, 845, 3394, 931, 2021, 3573, 3136, 1600, 3346, 1848, 611, 647, 2097, 2532, 679, 1660, 2627, 3177, 449, 884, 1394, 1713, 2051, 2133, 2264, 3056, 1626, 3754, 85, 2683, 324, 2680, 1299, 436, 1744, 3054, 1688, 1859, 2171]. Review [2985, 1904, 3711, 160, 2840, 3293, 2337, 223, 2086, 511, 1260, 1012, 1519, 2078, 1157, 2888, 1200, 375, 2058, 417, 2791, 3785, 254, 658, 771, 2270, 1265, 2449, 2881, 631, 205, 1772, 3521, 548, 1585, 742, 1616, 2088, 3389, 287, 634, 1950, 2085, 2288, 1497, 2894, 1300, 2620, 3267, 119, 2436, 2663, 31, 3561, 3635, 15, 818, 964, 2272, 3184, 2202, 3357, 1607, 1215, 3090, 3328, 179, 659, 1046, 485, 970, 3275, 545, 3261, 1000, 2630, 3303, 3783, 2267, 1496, 58, 25, 1141, 643, 654, 1760, 2732, 3083, 3072, 3046, 3325, 3338, 1090, 1502, 140]. Review [3231, 3393, 2983, 3713, 2285, 1506, 2753, 3130, 3376, 9, 359, 1808, 2667, 3662, 2081, 1962, 2068, 471, 30, 929, 3516, 2500, 3048, 181, 1009, 1578, 2557, 131, 1010, 3611, 1940, 592, 2048, 3247, 345, 2077, 164, 207, 1761, 3476, 3764, 460, 1359, 2244, 2827, 642, 888, 609, 2728, 900, 1105, 3685, 2981, 2381, 497, 3518, 3326, 378, 2770, 346, 468, 1666, 1727, 2220, 2528, 3341, 2784, 368, 1747, 2771, 3767, 734, 763, 3400, 2615, 975, 61, 1266, 3295, 648, 2164, 144, 982, 2132, 470, 2427, 193, 1307, 861, 894, 1867, 2621, 1423, 2720, 2895, 3391, 2196, 1855]. Review [1038, 255, 452, 1737, 690, 1412, 1833, 14, 3144, 668, 2829, 3221, 714, 2351, 3349, 2257, 533, 3735, 1746, 2490, 3055, 3710, 1971, 2578, 117, 519, 1278, 1397, 2613, 3059, 2212, 108, 707, 3029, 500, 1268, 871, 2896, 3390, 3085, 1522, 2045, 295, 2691, 2967, 2929, 522, 1137, 1738, 2209, 2700, 3591, 3405, 2385, 2828, 1483, 2141, 1873, 2378, 3009, 1347, 541, 776, 66, 290, 510, 1262, 847, 1388, 1941, 2754, 157, 2666, 2349, 62, 638, 3067, 980, 194, 773, 2418, 3736, 2282, 2416, 453, 366, 2610, 364, 1593, 2022, 896, 3183, 244, 2306, 3683, 2334, 1471, 2023]. Review [499, 1925, 3316, 3715, 554, 1216, 2491, 1613, 304, 1380, 2217, 2723, 3353, 3359, 1226, 3732, 844, 1842, 977, 1617, 277, 723, 1163, 2330, 3419, 2974, 57, 3379, 26, 2161, 2887, 521, 3686, 2091, 1103, 1827, 2265, 2671, 3161, 3248, 2737, 3237, 1301, 1277, 2042, 2733, 3525, 2924, 593, 3682, 299, 3079, 3608, 3227, 3304, 3645, 2542, 3574, 1679, 146, 2041, 2362, 1592, 2162, 3162, 3656, 2968, 3133, 1528, 1756, 1853, 3678, 2824, 406, 3398, 787, 1507, 2339, 3251, 869, 1026, 727, 3548, 109, 3336, 111, 69, 1612, 1668, 2093, 1858, 3340, 870, 3256, 504, 2741, 153]. Review [2736, 3438, 3010, 3265, 2559, 3429, 222, 3192, 396, 1295, 1674, 2403, 2061, 300, 1820, 1602, 365, 451, 687, 1487, 1683, 2277, 2531, 3300, 3725, 993, 2492, 106, 2632, 1627, 1006, 90, 639, 3551, 881, 3186, 3343, 494, 989, 1707, 2664, 3036, 1844, 557, 1115, 1809, 1909, 2810, 197, 1486, 1538, 1152, 3008, 3057, 2400, 3415, 3559, 1770, 3345, 125, 168, 3716, 271, 2612, 2184, 1508, 2722, 3583, 1392, 2435, 3745, 1129, 38, 133, 903, 1401, 1876, 2166, 2035, 2361, 2561, 3578, 8, 495, 2044, 2402, 3500, 1532, 413, 999, 877, 1587, 2099, 2221, 2050, 2609]. Review [2738, 3053, 3395, 559, 762, 918, 3201, 3242, 82, 531, 1411, 134, 892, 601, 850, 1297, 1020, 1946, 2210, 382, 3485, 2083, 2921, 2332, 1139, 3508, 3567, 2026, 1419, 2849, 2781, 3687, 257, 619, 906, 1284, 1466, 1710, 3271, 3633, 1509, 483, 3315, 3654, 1003, 1568, 3638, 2445, 151, 863, 2571, 2565, 475, 1583, 1142, 1261, 1463, 3339, 1944, 2886, 3642, 2275, 138, 1050, 2104, 3530, 3784, 1194, 2809, 1168, 1699, 1162, 2159, 1235, 1393, 393, 1389, 517, 467, 2430, 683, 717, 2854, 2892, 3546, 1482, 2170, 3241, 3502, 3005, 1551, 454, 2999, 266, 555, 491]. Review [3052, 3576, 1190, 3671, 2052, 298, 1116, 620, 1028, 1144, 3038, 1354, 2204, 409, 2836, 790, 1768, 173, 1164, 655, 701, 2703, 626, 83, 256, 1275, 2178, 267, 1217, 1474, 2406, 16, 1552, 2877, 3361, 636, 539, 616, 774, 1366, 1078, 2619, 825, 50, 803, 627, 1023, 1728, 759, 1527, 1143, 1257, 1363, 107, 3061, 2978, 3475, 1133, 512, 1077, 2089, 2748, 3255, 3399, 3506, 756, 2448, 2918, 3356, 2039, 2669, 2546, 3760, 526, 1465, 3141, 2262, 2772, 3014, 3513, 456, 608, 628, 3752, 2304, 758, 1742, 356, 1228, 1819, 29, 3408, 3757, 2308, 407]. Review [1618, 976, 3233, 1475, 1381, 1921, 380, 1751, 3058, 1022, 1404, 1670, 1414, 147, 2995, 419, 1469, 2426, 3657, 1907, 2568, 3087, 1922, 225, 1730, 70, 377, 1355, 1407, 1554, 2031, 2224, 1929, 2734, 3520, 3762, 507, 3094, 3679, 523, 1043, 2879, 1629, 938, 1008, 1814, 3323, 226, 678, 1232, 2242, 1947, 1193, 1579, 1049, 498, 1241, 1100, 1881, 3163, 1111, 2087, 95, 769, 927, 2850, 349, 663, 313, 646, 1031, 568, 1223, 2015, 2755, 2600, 3481, 1470, 2484, 1743, 1267, 1289, 915, 1741, 1620, 186, 696, 488, 1495, 1821, 2779, 3091, 1166, 2548, 1564, 1624, 2727, 757]. Review [1529, 2473, 2535, 2295, 889, 2241, 141, 2479, 2815, 213, 520, 1128, 1221, 1534, 1042, 3589, 2914, 3607, 1273, 2253, 354, 1293, 2169, 3522, 2764, 2477, 2298, 3580, 882, 995, 1095, 1948, 3252, 3660, 1240, 219, 2172, 2358, 3566, 199, 934, 2370, 2838, 653, 1230, 3153, 859, 1305, 1871, 1718, 1837, 1861, 2453, 3042, 1120, 3095, 2268, 96, 1590, 1822, 2922, 3043, 968, 1229, 872, 737, 3596, 1852, 703, 458, 2206, 2503, 2413, 2893, 3226, 3505, 3734, 3655, 1382, 1598, 2855, 3259, 3331, 3590, 3609, 1127, 1192, 372, 2065, 2984, 898, 2639, 2899, 3187, 808, 2562]. Review [195, 797, 1279, 2725, 381, 542, 2197, 1237, 263, 74, 2314, 2149, 3577, 374, 1732, 3311, 1140, 1584, 2258, 1135, 3727, 1304, 2821, 2287, 2437, 1290, 2991, 315, 2839, 3681, 2768, 1197, 67, 1402, 3035, 2841, 2372, 3707, 1055, 3557, 2225, 2188, 981, 1581, 2341, 128, 819, 1153, 2971, 1117, 2158, 2963, 2478, 1669, 2073, 3479, 1148, 1417, 2990, 3274, 2572, 2878, 1935, 3593, 1883, 2994, 169, 370, 1154, 1865, 2752, 3498, 253, 1258, 2670, 546, 21, 40, 1280, 3721, 2174, 218, 706, 2672, 2902, 310, 3641, 3643, 660, 1863, 3748, 384, 469, 936, 288, 1271, 1302, 3714, 496, 987]. Review [2163, 2876, 3401, 2344, 1697, 2243, 1076, 123, 1955, 1812, 1134, 53, 971, 2858, 603, 1775, 318, 1625, 1588, 37, 1958, 2607, 1533, 145, 613, 12, 305, 525, 728, 747, 1677, 2147, 2412, 2438, 3093, 2604, 1276, 2439, 394, 785, 1155, 1692, 1722, 1830, 2145, 1917, 2348, 2673, 2890, 3490, 998, 319, 2668, 2919, 3086, 3337, 3217, 1477, 347, 2103, 317, 1868, 3312, 3317, 3392, 79, 527, 809, 1159, 1520, 2070, 2635, 1365, 136, 935, 1249, 1167, 1270, 2817, 778, 3741, 666, 1963, 2682, 2692, 1754, 3690, 786, 2631, 39, 1591, 1150, 1675, 2443, 118, 2780, 1256, 166]. Review [1461, 89, 203, 421, 1196, 1244, 1131, 851, 1614, 1874, 1966, 3718, 1734, 1949, 630, 695, 2889, 1970, 2371, 3474, 824, 127, 583, 1195, 858, 3588, 1118, 2067, 3636, 899, 1138, 2599, 3482, 2505, 155, 1102, 3001, 2017, 1236, 1717, 3728, 251, 537, 720, 1027, 1759, 1615, 2778, 3777, 1854, 2624, 901, 1089, 2593, 3413, 3550, 2410, 2820, 2746, 459, 1838, 11, 1421, 2398, 2305, 2686, 278, 661, 1835, 2861, 3742, 1536, 3769, 2345, 149, 1777, 1473, 1250, 294, 383, 530, 2916, 2279, 2560, 2191, 2640, 2404, 2997, 3375, 1033, 3431, 2634, 1132, 2554, 2928, 3229, 3680]. Review [729, 1468, 1972, 1936, 3465, 3181, 400, 1211, 534, 719, 3068, 3330, 3765, 3466, 1124, 3149, 1400, 2982, 3639, 3472, 3410, 625, 2182, 2608, 505, 983, 2046, 3378, 1544, 2913, 2757, 474, 864, 1047, 2880, 3051, 1545, 891, 2905, 810, 502, 724, 1037, 1726, 1826, 3302, 2777, 2192, 478, 24, 770, 2553, 2904, 1913, 713, 904, 1831, 2074, 2049, 2596, 2688, 3066, 2993, 3296, 1709, 2297, 2740, 2030, 412, 2926, 932, 1662, 588, 2825, 3299, 3402, 463, 3143, 3309, 1857, 2151, 515, 473, 1224, 865, 369, 2037, 2294, 2504, 110, 1428, 2296, 1813, 2976, 699, 2199]. Review [2447, 2903, 1281, 1841, 1973, 922, 1872, 3262, 1126, 3219, 1546, 715, 208, 220, 2399, 2616, 1695, 2606, 2843, 2932, 1282, 2790, 681, 1491, 1817, 3649, 3532, 3726, 618, 229, 1912, 2420, 88, 269, 704, 933, 1094, 3131, 2444, 48, 1212, 566, 2303, 1885, 2699, 2092, 201, 414, 1960, 2177, 2300, 351, 3077, 775, 804, 916, 2036, 598, 2094, 482, 2428, 1099, 2547, 3159, 3075, 516, 1199, 1918, 2266, 3440, 3528, 1806, 2980, 2550, 805, 3360, 2907, 228, 529, 744, 2758, 1343, 1586, 3733, 1245, 178, 782, 1234, 3154, 1294, 1766, 3272, 316, 1036, 1834, 2082, 1379, 1887, 1210, 1418]. Review [721, 764, 3334, 2187, 2856, 3033, 3469, 735, 1087, 19, 357, 586, 1231, 3403, 192, 2658, 3434, 3318, 1711, 2917, 411, 97, 176, 765, 1731, 3397, 20, 1112, 2859, 3257, 1700, 416, 455, 1123, 308, 563, 2095, 1383, 2488, 1758, 2374, 3073, 3610, 187, 2641, 2486, 3387, 2356, 2756, 3139, 708, 1698, 1106, 1773, 2222, 2379, 2150, 3768, 1671, 3170, 139, 1296, 923, 2140, 188, 3037, 1048, 2367, 1425, 766, 1714, 1604, 3096, 1882, 3637, 450, 826, 1285, 2665, 3717, 2602, 172, 1575, 2446, 3422, 3663, 1015, 2996, 441, 1959, 1903, 518, 1460, 1774, 2884, 3152]. Review [3652, 3737, 3477, 3307, 2816, 200, 1114, 3182, 442, 2696, 3684, 1880, 2327, 3050, 1739, 3155, 607, 59, 64, 3224, 2783, 2601, 1259, 1476, 1014, 726, 480, 1083, 445, 1705, 2674, 1549, 1351, 2705, 3751, 3160, 779, 1110, 1524, 227, 3347, 42, 358, 795, 876, 1703, 2411, 1283, 3430, 1513, 1521, 322, 1764, 1203, 1755, 2084, 2441, 3030, 185, 2567, 114, 547, 2311, 3151, 3599, 1298, 321, 1019, 1498, 2909, 1213, 3484, 3517, 2401, 2679, 323, 1011, 1569, 1547, 3012, 1306, 1130, 1222, 1378, 3545, 1361, 248, 3724, 259, 282, 722, 1119, 1720, 1847, 1906]. Review [3232, 2637, 1942, 3759, 1776, 1233, 2152, 1846, 2830, 196, 3198, 2910, 2497, 2594, 265, 866, 550, 3174, 3501, 1225, 1113, 2156, 651, 1145, 991, 2190, 2851, 3592, 47, 657, 2476, 60, 371, 1274, 3558, 65, 2347, 3047, 1510, 170, 914, 3676, 1303, 2214, 606, 309, 1757, 2660, 2684, 152, 2053, 2708, 2595, 3439, 2846, 1426, 440, 2762, 1377, 3351, 320, 2835, 2845, 3468, 2415, 2043, 1916, 3065, 1589, 1954, 2432, 3189, 1723, 2283, 408, 1146, 3519, 93, 3424, 669, 1075, 1205, 1571, 1877, 2837, 990, 2180, 3078, 784, 941, 1535, 2749, 506, 1512, 2020, 3000, 3604]. Review [2286, 3665, 250, 1680, 806, 822, 204, 2966, 285, 1920, 3195, 1543, 3146, 3729, 481, 1219, 798, 2747, 2657, 3185, 3354, 3773, 337, 509, 1910, 2544, 1691, 2603, 937, 1253, 1040, 1810, 3584, 902, 2605, 71, 2617, 3011, 698, 1818, 2833, 472, 571, 1431, 1490, 1926, 2355, 2726, 3218, 1765, 1034, 3675, 2495, 3428, 3327, 3480, 3569, 28, 397, 567, 589, 887, 1247, 1943, 78, 928, 610, 1505, 2570, 716, 148, 580, 1227, 878, 2731, 3575, 3739, 1209, 1611, 813, 2776, 1207, 1239, 3489, 2558, 1348, 2263, 880, 1462, 2685, 1098, 2373, 1823, 3138, 1721, 1869, 3774, 524, 1238, 967]. Review [2016, 2842, 3527, 2812, 3674, 3667, 553, 1248, 2623, 1051, 297, 767, 1667, 1745, 3650, 1136, 992, 2247, 2383, 3167, 3668, 3594, 2193, 1608, 1107, 17, 2299, 2626, 2988, 2813, 692, 448, 788, 2040, 2203, 2628, 3744, 1531, 1481, 2750, 3039, 1573, 32, 3779, 875, 1696, 3015, 652, 739, 1540, 86, 143, 2301, 1619, 2496, 399, 986, 570, 479, 2143, 3708, 1621, 2219, 3435, 3503, 1242, 2165, 2252, 777, 1599, 1582, 3168, 1427, 2964, 3441, 3661, 2563, 1609, 2340, 1004, 1376, 405, 1479, 2908, 3560, 137, 273, 275, 379, 746, 997, 1104, 2310, 2643, 3614, 180, 2423, 1493]. Review [484, 2269, 2167, 2485, 2384, 224, 1500, 150, 361, 874, 1374, 174, 373, 249, 585, 2028, 3313, 656, 3294, 302, 1387, 1356, 2751, 3264, 2386, 1035, 1550, 1938, 2860, 3669, 1503, 2819, 1345, 2499, 2689, 3529, 1969, 3555, 645, 540, 1539, 1570, 3076, 2930, 1595, 1017, 2659, 3298, 3488, 2359, 3598, 162, 360, 1092, 2822, 3659, 493, 2743, 3497, 3740, 1676, 2442, 1054, 2422, 1147, 376, 1403, 303, 33, 1499, 1360, 2256, 3509, 3651, 2853, 353, 2826, 2072, 3514, 3003, 3746, 3478, 1408, 2409, 979, 3225, 260, 2935, 292, 344, 2569, 3411, 3672, 367, 3193, 113, 2057, 552]. Review [2618, 2891, 3492, 2576, 3278, 3526, 725, 1353, 2205, 2276, 3006, 3586, 3743, 34, 466, 1580, 2249, 1079, 2019, 1915, 694, 1630, 796, 640, 477, 3486, 2273, 2759, 3244, 3322, 1933, 1367, 2290, 2760, 3499, 461, 2336, 2250, 3568, 755, 171, 341, 1001, 2357, 2832, 3140, 2543, 2139, 2365, 926, 2024, 2090, 177, 2577, 120, 2326, 2501, 75, 18, 2625, 2454, 732, 2366, 23, 1085, 326, 1706, 3731, 1464, 1724, 94, 385, 677, 893, 1566, 2176, 2923, 3758, 1511, 2925, 1605, 1429, 68, 799, 1013, 2194, 2376, 3605, 1763, 2687, 2342, 13, 604, 1572, 2451]. Review [420, 3253, 130, 1287, 1565, 2034, 1286, 289, 689, 1525, 2302, 2642, 3032, 3556, 3712, 2775, 312, 2066, 2920, 1292, 3409, 3775, 1693, 1220, 1914, 802, 116, 680, 1364, 87, 978, 772, 2245, 2315, 1537, 1594, 2363, 1410, 2636, 1430, 1729, 2677, 543, 1748, 710, 885, 886, 1375, 3570, 2745, 2293, 2218, 1369, 536, 2360, 126, 264, 3470, 1694, 286, 1478, 1686, 3722, 3581, 2979, 3060, 3238, 3571, 2730, 3148, 284, 3178, 3747, 984, 352, 617, 3612, 501, 2375, 3200, 261, 2675, 3666, 2055, 2970, 1878, 307, 1373, 3582, 1932, 2056, 1396, 2144, 535, 1045]. Review [2597, 77, 1517, 912, 1514, 1934, 1041, 846, 2965, 3761, 2787, 63, 3348, 3437, 281, 780, 591, 1631, 3165, 919, 3082, 2977, 528, 649, 1352, 1769, 1610, 2702, 3273, 3563, 3564, 1518, 3191, 1091, 3770, 3305, 1672, 3199, 760, 1370, 1413, 2281, 2157, 2701, 3089, 3384, 154, 2489, 2785, 3156, 437, 3523, 258, 879, 3493, 476, 1924, 2248, 2574, 641, 355, 457, 492, 1811, 3491, 1712, 2475, 3239, 3404, 996, 3332, 3750, 3423, 3504, 3333, 3432, 2251, 503, 3228, 1681, 486, 3190, 115, 963, 2540, 3579, 3677, 2862, 2179, 2774, 36, 3723, 1080, 3028, 2198]. Review [2278, 387, 590, 1002, 386, 3524, 3720, 1530, 2629, 1086, 1057, 2508, 3483, 91, 415, 2102, 2788, 3688, 3064, 3223, 1391, 306, 1733, 2709, 817, 3587, 1541, 3355, 816, 444, 1489, 2380, 81, 206, 215, 581, 1208, 1108, 1682, 1879, 1884, 2135, 2032, 2369, 2417, 2566, 2766, 3041, 2973, 3510, 3634, 1074, 1395, 1845, 2901, 801, 1422, 2154, 3081, 792, 1708, 2487, 3245, 350, 338, 2018, 3045, 2847, 622, 1161, 962, 3088, 2552, 3179, 3673, 3320, 245, 2059, 2080, 3473, 2611, 3396, 403, 842, 2440, 3763, 3134, 2096, 2814, 2069, 339, 2676, 3766, 132, 231, 3173, 2549]. Review [3321, 1350, 10, 709, 1485, 2335, 741, 1542, 3615, 2160, 2897, 3260, 27, 1096, 2706, 2421, 291, 3263, 3388, 988, 2271, 2721, 3249, 3417, 587, 2690, 438, 600, 3416, 3487, 1886, 538, 1016, 1875, 2064, 1623, 3080, 1149, 965, 2181, 2769, 811, 930, 1398, 3771, 1386, 2100, 2313, 1125, 1372, 3554, 1390, 2573, 246, 3135, 296, 2377, 3176, 3407, 1673, 2213, 1911, 1516, 682, 2408, 2261, 3246, 815, 2931, 2047, 1825, 84, 214, 279, 1767, 142, 1905, 3780, 422, 3494, 3709, 3188, 314, 410, 925, 1053, 1866, 3092, 2183, 2564, 1399, 584, 3007, 340, 392, 3383, 2694]. Review [615, 391, 667, 662, 913, 1715, 293, 1663, 2201, 43, 54, 814, 3031, 3664, 2331, 3171, 1480, 1771, 3258, 3427, 3632, 2148, 3782, 3648, 1816, 2155, 325, 895, 3547, 1956, 2511, 3230, 1603, 3562, 22, 135, 2598, 2494, 1953, 3531, 159, 1007, 1661, 1601, 1927, 2274, 2765, 3301, 3382, 272, 2633, 1684, 1191, 1472, 121, 1384, 1385, 705, 230, 3145, 3385, 564, 395, 2136, 1151, 1678, 1515, 1005, 1416, 2289, 3266, 3603, 3421, 3602, 3781, 124, 854, 2481, 3406, 3137, 129, 2773, 793, 1039, 2483, 1409, 1272, 1291, 920, 1269, 1597, 3297, 163, 209, 860, 2291, 3049, 917]. Review [2818, 2707, 2424, 3442, 2972, 46, 633, 807, 2071, 3097, 388, 1805, 791, 2538, 3147, 3377, 76, 311, 1032, 1749, 1548, 2216, 2529, 3197, 3268, 122, 594, 686, 1344, 2062, 2414, 343, 857, 3062, 189, 820, 848, 2848, 3013, 612, 2541, 1622, 2507, 2309, 3772, 1864, 55, 2246, 56, 635, 2186, 3172, 3418, 335, 1202, 939, 2506, 389, 730, 2452, 1740, 3512, 1606, 1964, 2763, 1735, 3350, 2368, 700, 823, 1288, 961, 1836, 2207, 3084, 3433, 212, 514, 3414, 561, 2509, 2551, 582, 544, 3352, 35, 614, 2280, 3597, 868, 1198, 2033, 1567, 1687, 2992, 3150, 3738, 3425, 697]. Review [733, 883, 1596, 1965, 2200, 2329, 3595, 252, 487, 41, 112, 336, 508, 624, 664, 688, 718, 1246, 856, 1101, 1829, 1839, 1919, 2975, 3653, 556, 363, 2029, 443, 2153, 2539, 1862, 532, 270, 599, 1860, 897, 1750, 3658, 2079, 2482, 2729, 2211, 2138, 597, 1937, 3074, 3310, 2533, 1930, 2934, 3306, 3644, 1526, 1850, 1928, 3004, 211, 560, 711, 1251, 972, 1346, 2782, 362, 1628, 1824, 821, 2786, 3270, 2325, 623, 1097, 2382, 2537, 2831, 439, 685, 1252, 2450, 3216, 3277, 1368, 3689, 909, 217, 781, 3386, 2724, 2354, 198, 2857, 190, 276, 693, 165, 2142]. Review [3381, 3552, 262, 3196, 342, 1156, 1424, 2530, 3169, 1574, 462]. Revised [1214]. Revisions [3368]. Revisited [1398, 3526, 3738]. Revisiting [2458, 1975]. Revista [2148]. revolución [3054]. Revolution [1986, 2307, 268, 2744, 2885, 853, 2983, 2827, 2981, 2647, 1508, 82, 3687, 1273, 1822, 3285, 333, 2951, 3051, 2553, 2367, 3155, 2653, 990, 1512, 3575, 739, 2128, 2218, 3238, 2375, 920, 1628, 3732, 2820, 1610, 2721, 2185, 3324, 2912, 2681, 2350, 2349, 2542, 2892, 3252, 1865, 496, 778, 530, 2374, 148, 880, 2743, 355, 43, 3405, 3732, 3233, 3777, 1460, 1259, 1505, 2331]. revolutionäre [1944]. Revolutionaries [3300]. Revolutionary [3293, 2484, 2438, 3598]. Revolutionizing [553, 987]. Revolutions [2385, 2658, 2476, 3356]. Rewards [3523]. rex [1868]. Rey [2263, 1344]. Reymond [3254, 3178]. Rezensionsliteratur [1964]. Rezeption [150]. Rezeptionsgeschichte [1826]. Rheinberger [3013]. Rhetoric [2292, 2023, 3160, 926, 3570, 3383, 3648, 2934, 1100]. Rhetorical [2097, 236, 666]. rhétorique [2733, 2879]. rhétoriques [2595]. Rheumatic [1627]. Rhodes [1913]. Rhodri [3490]. Rhumb [3137]. Rhys [2627, 2720, 367]. Ri [3261]. Ribbon [2193]. riben [3261]. Rica [2625, 1420, 2922]. Rice [2161, 597]. ricerca [1381]. Ricerche [3319]. Rich [2811, 65]. Richard [51, 2086, 62, 365, 266, 3408, 2548, 706, 37, 537, 724, 3050, 2214, 143, 1403, 292, 725, 261, 492, 1057, 817, 792, 741, 212, 362, 1160, 2364, 1165, 2900, 2615, 2967, 3079, 2210, 1069, 2572, 1838, 2030, 1036, 2422, 2501, 3032, 1041, 1151, 860, 2216, 2396, 1101, 1919, 527]. Richards [243, 378]. Richardson [1121, 3335, 436, 3759]. Riché [939]. Richelson [2217]. Richness [698]. Richter [2764, 883, 256]. Rick [481]. Rickey [279, 421]. Rickover [1156]. Rieger [3068]. Rieke [500]. Riemann [799]. Rienk [2273]. Rife [726]. rigenerazioni [3161]. Right [225, 2508, 1005]. Rights [975, 1963, 2302, 2574, 2029]. Rights-Historical [975]. Rik [413]. Riley [562, 376]. Rina [2487]. Rinascimento [3728, 2338]. Ringer [934]. Ringler [2200]. Ripley [874]. Ripoll [161]. Riquier [939]. Rise [3344, 2622, 1752, 1044, 404, 3090, 103, 2559, 3583, 2050, 1261, 683, 555, 2172, 123, 2889, 1854, 1132, 2297, 1210, 2679, 2830, 3501, 567, 1745, 3650, 1147, 1391, 1007, 1839]. Rising [1674]. Risk [409, 2089, 2374]. Risks [949, 950, 1180]. risque [3494]. Risse [2741]. Rita [1820, 35, 1022]. Rites [37, 252]. Rithmomachia [2157]. Ritual [1947, 577]. Ritvo [167]. Rivalries [3667]. Rivalry [894]. Rivals [891]. River [1892, 803, 854, 2441]. Rivers [1939, 972]. rivoluzione [2267]. rivoluzioni [1049]. Roach [1810]. Road [1642, 2472, 1973, 736]. Roazen [186, 1857]. Robbins [1349]. Robert [2933, 3158, 684, 2472, 1405, 1577, 2834, 548, 1502, 193, 295, 133, 2402, 762, 134, 3315, 2159, 266, 409, 2850, 703, 2639, 384, 203, 48, 411, 2762, 3773, 580, 2751, 2577, 68, 3178, 2970, 535, 457, 338, 340, 615, 395, 388, 791, 2216, 697, 960, 2998, 1405, 2419, 845, 2272, 2381, 2324, 1798, 2242, 2479, 2752, 2692, 2067, 2757, 1973, 2486, 2679, 941, 1051, 1609, 2269, 3619, 2102, 1391, 2483, 2529, 2975, 2106, 1408]. Roberta [1821]. Roberto [668, 1048]. Roberts [3475, 3087, 349, 2072, 2581]. Robin [2745, 2841]. Robinson [1745]. Robot [3719]. Robson [606]. Robyn [1963]. Rocca [2334]. Rochberg [2335]. Rochelle [308]. Rochester [3002]. Rock [3777]. Rocke [2073]. Rockefeller [2557, 2686, 2993, 1737, 1738]. Rocket [742, 1862]. Rocketry [1645, 3100]. Rockmore [2946]. Rocks [2535]. Rocky [810]. Rod [1896]. Rodger [248]. Rodman [2639]. Rodriguez [146, 2742, 71, 768]. Rodriguez-Consuegra [71]. Roelcke [635]. Roger [1309, 789, 44, 157, 16, 1921, 70, 1697, 2606, 746, 154, 486, 84, 293, 1416, 3653, 443, 2732, 861, 2275, 1812, 1099, 2084, 842, 1368, 1921, 36]. Rogers [3742, 128]. Rogge [3679]. Rohrbasser [2671]. Rojecki [532]. Rokitansky [3418]. Roland [628, 2568, 20]. Role [2630, 1038, 2922, 2725, 2003, 722, 550, 3582]. Roles [114]. Rolf [2185]. Roll [2640]. Roll-Hansen [2640]. ROM [2296]. Roman [1218, 2844, 3638, 1821, 577, 1034, 3480, 772, 3493, 3386, 586]. Romance [659, 1249, 2069]. romande [2091]. Romano [1853, 613]. Romans [2408]. Romantic [2718, 1660, 1105, 3592, 3581, 2535, 3584, 633, 508]. Romantic-Period [1105]. Romanticism [1121, 46, 2385, 1729]. romántico [3604]. Rome [3144, 1487, 2050, 753, 981, 1823, 477, 3725]. Romeril [3125]. Rommevaux [2971]. Ron [543, 2630, 300, 2562]. Ronald [417, 471, 144, 244, 722, 3529, 2735, 2371, 3084]. Roof [3305]. Roos [3496]. Roosevelt [2436]. Root [439]. Rooted [2972, 2972]. Roots [1349, 2630, 2392, 3417, 1515]. Rootselaar [501]. Rosalind [1915]. Rose [980, 1947, 94, 172, 2957]. Rose-Colored [2957]. Rosemarie [693]. Rosemary [598]. Rosen [1963]. Rosenberg [3097, 206]. Rosenkranz [3410]. Roshdi [469]. Rosicrucianism [2490]. Roslin [39]. Rosner [2074]. Ross [315, 2559, 2435, 402]. Rosselló [586]. Rosser [2941]. Rossi [1094, 2336, 3552]. Rossii [1354, 2262]. Rostock [1178]. Roswell [1044, 1045]. Rotates [2500]. Roth [136]. Rothblum [211]. Rothenberg [3340, 371]. Rothman [1577]. Rothmann [39]. Rothschild [2656]. Rothstein [2374]. Rotverschiebung [1386]. Rouse [177]. Rousselot [183]. Routes [2789]. Routledge [3360, 3351]. Routledge/Asian [3351]. Rowan [889]. Rowe [2751]. Rowland [2612, 1382, 1823, 2573]. Rowley [3760]. Rowley-Conwy [3760]. Roxana [2969]. Roy [715, 450, 1569, 540, 2049, 1341, 3396]. Royal [1636, 678, 1229, 1918, 362, 45]. Royale [2164, 3152, 1372]. Rozanskaja [1291]. Rozwadowski [2761, 1624]. Rube [1142]. Rubiés [985]. Rücktritt [725]. Rudacille [1881]. Rudavsky [475]. Rudd [757]. Rudder [21]. Ruderman [699]. Rudgley [261]. Rüdiger [3761, 2372, 607]. Rudnytsky [2075]. Rudolf [2285, 3227, 630, 3672, 3178]. Rudolph [3035, 152]. Rudwick [2743]. Ruellius [587]. Ruggles [2149]. Rui [129]. Ruin [696]. Rule [2292, 1498]. Rules [1161, 1039]. Rum [1878]. Rumm [371]. Rumphius [881]. Runge [170]. Running [2697, 803]. Rupert [271]. Rupke [2847]. Rural [2371, 316]. Rusd [2879]. Ruse [1971, 2691, 935, 848, 173]. Ruskin [2556]. Rusnock [2266]. Russel [2254, 1848, 1392, 2435, 2280]. Russell [727, 1028, 1629, 1515, 2889]. Russia [1995, 2943, 2054, 3769, 3664, 1202, 2521, 1838, 920, 3337]. Russian [1314, 207, 2772, 2087, 1620, 2506, 2304]. Russian-German [1314]. Russko [1314]. Russko-nemetskie [1314]. Russland [347]. Rust [46]. Ruston [2628]. Ruth [345, 1397, 1134, 251, 767, 892, 1583, 2295, 1047]. Rutherford [1840]. Rutkin [2819]. Rutot [1648]. Ruzsits [1428]. Ryan [2076]. Ryden [480]. Rymer [669]. Ryo [2186].

S [2742, 2987, 1851, 1165, 1081, 3627, 375, 181, 1009, 378, 2784, 193, 3009, 3419, 1116, 3757, 915, 381, 1148, 305, 2147, 89, 278, 744, 1106, 2355, 3480, 1943, 3298, 2743, 113, 75, 893, 2923, 1605, 649, 3523, 306, 815, 410, 1866, 913, 705, 2136, 76, 3772, 1198, 211, 3196, 2006, 753, 1106, 1745]. S. [1577, 2083, 1119]. S.J [3092]. Saalmann [2144]. Saarinen [431]. Sabbagh [1878]. Sabine [3679, 2894, 2971]. Sabri [1089]. Sachs [319, 3198, 1040, 467]. Sacks [508]. sacré [2337]. Sacred [349, 3669, 1687]. Sacrificial [3055]. Saga [549]. Sagan [811]. Sage [1236]. Saggi [3138, 877]. Saggio [2099]. Sahagún [2260]. Saharan [218]. Saikku [2569]. Sail [1824]. Sailed [3565]. Saint [2425, 3649]. Saint-Hilaire [2425]. sainte [2738]. Sainthood [2045]. Sakharov [2772, 1536]. Salavert [586]. Salazar [2762]. Salazar-Palma [2762]. Salchers [1639]. Saldaña [3140]. Sale [3523]. Salem [488]. Sally [595, 2295]. Salmon [413, 1958, 352]. Salomone [2044]. Salons [3535]. salpêtres [364]. Salpêtrière [2178]. Salt [3496, 1571]. Salud [141, 60]. salute [25]. Salvation [2550]. Salzes [3645]. Sameness [1541]. samhället [2080]. Sammlungsgeschichte [909]. Samoan [1850]. Samuel [1009, 368, 928, 2421, 3407, 3597]. San [640, 2377]. Sanchez [300, 2742, 2938, 3054]. Sandage [2848]. Sander [3670, 166]. Sanders [808, 343]. Sanderson [1853]. Sandifer [3315]. Sandor [186]. Sandra [3736, 550, 177, 2754, 2861, 2842, 2201]. Sands [1144]. Sanford [2376]. Sangwan [453]. Sanitary [344]. Sankt [1354, 2191]. Sankt-Peterburge [2191]. Sankt-Peterburgskoi [1354]. Sanskrit [2245]. Santa [2835]. Santé [3321]. Santesmases [2444]. Santiago [2662, 3056]. Santorini [590, 290]. Santos [190]. Santponc [1952]. Sanz [1868]. saperi [2403]. sapientiae [1238]. Sapp [2246]. Sara [2783]. Sarah [44, 3340, 805, 799, 2693, 2370, 3033]. Sardis [608]. Sargent [224]. Sarkar [2762]. SARS [3082]. Sarton [1896, 3454, 3453, 3455, 3456, 3452]. Sasaki [2489]. Sassower [663]. Satellite [3262]. Satpal [453]. Sauer [3410]. Saul [776, 620, 821, 2888, 3081]. Saunterer [807]. Saussure [712, 1243, 2886, 161]. Savage [3043, 772, 2315, 643, 3341, 1244]. Savage-Smith [3043, 772, 2315]. Savages [1534]. savant [2721]. savante [335]. savantes [1352]. savants [1322, 3740]. Saved [3087]. Saving [562, 1464]. Savoia [2401]. Savoir [3739, 2721]. Savoirs [701, 142, 3321]. Saw [3684]. Saxony [2798]. Say [1650, 299, 2197]. Scafi [3034]. Scaife [718]. Scala [629]. Scale [2285]. Scalpel [1881]. Scandalous [2074, 2935]. Scandalously [451].

Scandinavian [3265]. Scanning [1764]. Scarce [413]. Scarth [1075]. Scene [2978]. Scenes [3272, 3150, 47]. Scerri [3014]. Schabas [2917]. Schädlinge [2370]. Schaffer [3475, 796]. Schafhauser [1089]. Schall [1639]. Schatteninstrumente [843]. Schellenberg [3088]. Schiebinger [2540, 569]. Schieder [2117]. schillerndes [3645]. Schirmbeck [195]. Schizophrenia [551]. Schlange [2144]. Schlange-Schöningen [2144]. Schlich [2218]. Schliemann [176]. Schmandt [242]. Schmandt-Besserat [242]. Schmaus [2453]. Schmidl [3486]. Schmidt [3325, 3685, 587]. Schmitt [3248]. Schmitz [3035]. Schmolders [346, 121]. Schmuhl [2763]. Schnalke [164]. Schneider [2375]. Schnell [2409]. Schoenfeld [3340]. Schoenfeldt [487]. Schoepflin [2077]. Schofield [2419]. Scholar [706, 1074]. Scholarly [1616]. Scholars [1616, 1943]. Scholarship [2321, 3475, 792]. Scholasticism [1368]. Scholastics [157]. Scholthof [399]. Scholz [2895, 377]. Scholz-Strasser [377]. Schonen [256]. Schönfeld [631]. Schöningen [2144]. Schonland [2169]. School [2557, 1842, 2170, 1948, 1382, 3759, 1531]. Schools [307]. Schopenhauer [1597]. Schorn [244]. Schott [150]. Schouls [1098]. Schrift [843]. Schriften [363, 3762]. Schrödinger [2023, 72]. Schroter [305]. Schrotflinte [2560]. Schule [3418]. Schulenberg [707]. Schulkin [2764]. Schullery [2078]. Schulmann [2751, 68]. Schulten [2079]. Schultz [2560, 2765]. Schultze [642]. Schumaker [1024]. Schuman [1896]. Schuppener [758]. Schur [888]. Schwann [3178]. Schwartz [251, 13, 289]. Schwarz [20]. Schwarzmann [1089]. Schwarzmann-Schafhauser [1089]. Schweber [649]. Schweizer [2347, 1004]. Schwerin [2695]. Schwerpunkt [3715]. Schwinges [3477]. Sci [3433]. Science [1160, 3236, 1255, 3697, 465, 2898, 2, 427, 579, 1311, 944, 946, 1432, 1640, 1782, 1784, 1786, 1436, 1438, 1446, 1632, 1633, 1636, 2108, 2120, 1979, 1988, 1992, 2517, 3099, 2943, 2796, 3116, 3447, 3625, 3448, 1832, 738, 2556, 849, 3319, 3778, 247, 2472, 1121, 1974, 1056, 2463, 3324, 3542, 2681, 2038, 2767, 216, 3019, 1896, 3243, 3496, 3222, 2704, 2906, 2328, 3670, 2911, 1931, 3646, 2936, 3276, 2054, 210, 569, 595, 1068, 3329, 2312, 3507, 1082, 2292, 1848, 2097, 1660, 2237, 2985, 2888, 1200, 2791, 205, 2620, 15, 3275, 545, 1892]. Science [643, 2732, 3338, 2983, 2646, 3535, 2500, 2557, 131, 592, 2077, 207, 460, 1359, 497, 3518, 1666, 2771, 959, 2720, 2647, 3391, 1746, 2613, 3449, 3539, 871, 2045, 66, 510, 2392, 3454, 62, 638, 3736, 453, 2610, 2023, 2217, 2723, 3353, 2875, 1163, 3419, 2974, 3701, 3133, 111, 2953, 3365, 3265, 2655, 2954, 3036, 3008, 133, 2221, 1069, 2026, 3271, 3633, 3638, 863, 1463, 3123, 138, 3288, 3530, 1194, 1168, 2319, 3122, 683, 717, 3576, 1164, 3370, 626, 1078, 3283, 107, 3475, 1077, 2089, 2748, 3208, 3369, 2669, 3141, 1742, 29, 1922, 1730, 3094, 1629, 938, 3323, 1193, 568]. Science [2015, 3481, 1289, 2779, 833, 1624, 2727, 3209, 2473, 2241, 3522, 2298, 2236, 3660, 2808, 1871, 2922, 458, 3226, 1598, 3259, 3590, 3609, 752, 2562, 195, 3117, 2821, 1290, 315, 2341, 2963, 2073, 1148, 1417, 2994, 3641, 3643, 3748, 936, 7, 1812, 603, 2412, 785, 1722, 2673, 2668, 3086, 3217, 2323, 2635, 136, 935, 2631, 1675, 2003, 421, 1195, 1138, 3001, 3413, 459, 11, 2686, 278, 1835, 294, 534, 2982, 2526, 2757, 2234, 1662, 3299, 2296, 2976, 1973, 3262, 1546, 330, 715, 2790, 3532, 2303, 1885, 235, 3159, 432, 2980, 2550, 2907, 1294, 1036, 1834, 3210, 3207]. Science [3455, 2658, 411, 3397, 1112, 416, 455, 3610, 2222, 139, 1048, 2665, 2996, 1903, 3737, 3307, 3182, 1739, 3155, 64, 3751, 3430, 1521, 1755, 3030, 3599, 1378, 259, 3232, 3759, 1776, 2594, 3501, 2524, 2851, 3592, 2053, 2708, 320, 1954, 3189, 2283, 6, 93, 784, 3665, 3456, 2657, 2587, 3185, 509, 1040, 3540, 571, 1431, 2355, 2239, 2558, 2685, 674, 3211, 3541, 2812, 3674, 3667, 1185, 1667, 2650, 1107, 17, 3744, 3039, 3779, 570, 2952, 2165, 1376, 1479, 3537, 2867, 2860, 1345, 3529, 3555, 540, 3076, 3298, 3659, 576, 1147, 3003, 2581, 3225, 3450, 2935, 2891]. Science [3278, 3120, 3526, 2249, 1915, 3486, 3322, 926, 2391, 2302, 2642, 680, 2677, 2218, 2360, 3581, 3238, 284, 2957, 3666, 1932, 2597, 77, 63, 780, 3191, 3089, 3523, 2322, 1712, 3423, 2958, 2251, 2629, 415, 3223, 3587, 2032, 2523, 1395, 2154, 3081, 2018, 1161, 962, 2080, 132, 600, 3080, 2313, 1449, 246, 1905, 925, 2585, 43, 3258, 2598, 1796, 159, 1927, 1678, 3266, 1039, 1548, 2956, 3268, 3212, 189, 1622, 3772, 3512, 673, 1288, 961, 830, 1333, 1792, 2230, 1337, 3150, 508, 718, 1101, 1750, 3074, 2934, 3004, 821, 3270, 2450, 3277, 1795, 3381, 3196, 1736, 161, 969, 2896]. science [1710, 3361, 825, 3482, 1225, 2595, 2988, 3355, 1406, 2472, 2060, 1888, 1165, 2054, 2510, 964, 2318, 2967, 999, 1162, 2484, 3117, 2861, 2036, 2550, 1595, 3006, 2315, 264, 1053, 1866, 2511, 920, 1032, 1928, 2622, 960, 1405, 2223, 1888, 605, 595, 2433, 2698, 1626, 1853, 3010, 3638, 1050, 683, 356, 1929, 1095, 3095, 2839, 2670, 384, 728, 527, 666, 2889, 2624, 459, 2640, 864, 1094, 744, 1418, 779, 1569, 3665, 2747, 78, 1098, 2964, 585, 645, 2577, 677, 2293, 3238, 1413, 2701, 3523, 1811, 1002, 2102, 581, 3045, 3088, 2100, 410, 1805, 791, 700, 1567, 697, 1174, 1179]. Science [1316, 1172, 1439, 1444, 2121, 1991, 3104, 2946, 1753, 1752, 2137, 3358, 3613, 3412, 2027, 632, 3511, 1664, 3175, 665, 85, 417, 3338, 3376, 1266, 2209, 776, 57, 3379, 3678, 3415, 3508, 3687, 2031, 3094, 757, 2473, 2241, 3505, 1290, 2876, 1722, 2398, 3769, 1972, 2399, 1960, 3733, 1234, 1379, 1774, 1110, 1298, 3424, 1765, 3575, 3774, 2016, 3744, 3661, 540, 3076, 3740, 3746, 3140, 3758, 1286, 1220, 3570, 1631, 437, 492, 3579, 3720, 2709, 2706, 3080, 2213, 3188, 43, 3632, 1661, 1191, 3266, 823, 561, 2509, 2426]. Science-Industry [2657]. Science-Within [2790]. Sciences [1787, 1443, 3157, 3467, 446, 1689, 631, 2164, 3454, 3678, 787, 1111, 2473, 1192, 987, 1250, 1513, 28, 1227, 3497, 3411, 796, 87, 332, 2787, 3190, 963, 36, 3355, 1485, 2271, 3664, 3442, 743, 1593, 2886, 16, 1552, 2619, 3639, 3152, 2675, 1350, 1372, 4, 2850, 824, 746, 1806, 1328, 1637, 1983, 3471, 2528, 1293, 1818, 422, 688, 1300]. Scientific [1870, 1325, 1787, 1986, 2942, 2622, 1501, 418, 268, 1849, 1974, 3324, 740, 2885, 745, 401, 511, 3261, 3783, 2983, 2068, 3048, 2981, 2381, 1307, 2621, 3735, 3121, 3390, 3732, 2542, 2968, 1853, 1336, 3716, 8, 1003, 3287, 1583, 2892, 3241, 3283, 2448, 3356, 3233, 2224, 2734, 2169, 2238, 2358, 1822, 3226, 3285, 315, 2341, 310, 7, 3714, 496, 1955, 2348, 319, 778, 1966, 2951, 899, 2345, 1777, 770, 3299, 699, 2199, 2447, 2903, 681, 229, 933, 2758, 782, 3318, 3694, 1773, 3155, 2783, 2653, 2311, 259, 1377, 1723, 506, 2587, 148, 3575, 880, 3527, 2203, 3779]. Scientific [143, 3708, 997, 2643, 2586, 3529, 3555, 3659, 2422, 260, 113, 3120, 1353, 3619, 2326, 75, 2454, 1013, 2451, 2636, 63, 834, 1631, 1672, 355, 3332, 3677, 3524, 1893, 1449, 3230, 2274, 1384, 3084, 3698, 1687, 2450, 2927, 941, 988, 2025, 679, 678, 1055, 3529, 338]. scientifica [2267, 2282, 2411]. scientifiche [3735, 1381]. scientifico [3102]. scientifique [3157, 3639, 2278, 3361, 798, 1350]. scientifiques [494, 469, 1552]. Scientism [3511]. Scientist [2555, 1725, 1950, 2162, 3005, 3255, 2169, 2922, 2070, 2866, 2003, 3693, 2816, 2432, 2586, 1748, 3092, 50, 1410, 649, 2200, 3485, 1740]. Scientists [3344, 2472, 644, 1616, 3764, 1038, 1909, 531, 1466, 3522, 2878, 1302, 1131, 2390, 3210, 2713, 1019, 1431, 77, 3723, 3673, 3205, 733, 1937, 217, 2472, 1380, 1417, 2327, 2093, 1750]. Scientists-Their [1466]. Scienza [3394, 2977, 2173, 3319, 2338, 2282, 1465, 3138, 3141]. Scienze [3138, 2115, 3728]. scintille [2352]. Scope [3360, 2123]. Scopes [1265, 1726]. Scorpion [2333]. Scot [982]. Scotland [1978, 1359, 3050, 2354, 299, 1378]. Scott [555, 2740, 269, 3224, 3092, 1251, 213, 2914, 1924]. Scottish [1439, 299, 3162, 2741, 887, 3653]. Scotus [143, 144]. Scourge [1942, 1962]. Scranton [291]. Screening [3469, 1138]. Scriabine [553]. Scriba [2731]. scribe [2721, 2945]. Scripps [2478]. Script [3247]. Scripts [721]. scriptura [3566]. Scripture [136]. Scrupulosity [960]. Scrutinizing [2577]. Sculptor [353]. scuola [324]. Sea [1624, 6, 1539, 2958, 590, 2761, 1763]. Seaborg [1618]. Seafaring [3005]. Seams [411]. Sean [402, 3598]. Séance [2594]. Search [2119, 2502, 5, 2704, 911, 193, 861, 1538, 380, 1564, 2878, 2439, 1270, 1731, 187, 2486, 139, 441, 1846, 1387, 415, 444, 3473, 1515, 2289, 972]. Searchers [3088]. Searching [884, 2477, 3588]. Seas [359, 3771]. Seashell [2045]. Seasons [621]. Seaver [2255]. Secession [1328]. Secoli [691, 3735]. secolo [2022, 35]. Second [3100, 924, 3449, 2334, 3678, 1365, 1378, 1220, 2071]. seconda [2252, 2363]. secondaire [3639]. Secondary [1165, 2391]. seconde [3740]. secondo [2250]. Secord [1000]. Secrecy [534, 2033, 2826]. Secret [1310, 1000, 2042, 2918, 186, 369, 3297, 2839]. Secretaries [2152]. Secretary [1012]. Secrets [731, 884, 3389, 1450, 1568, 652, 2069, 1315, 3149]. Secular [349]. Secularization [3768]. Secure [801]. Security [316, 2216]. seducción [1952]. Seeds [807]. Seeing [1077, 3478, 3041, 2882, 2950, 2524, 1204]. Seekers [847]. Seelenorgan [24]. Seen [2882, 187]. Seers [3088]. Segal [2376, 2766]. Segre [693]. Segregation [1302]. Seguin [1610]. segundo [146]. Sehgal [453]. Seibert [1912]. seicento [493]. Seidel [126]. Seidengart [2896]. Seife [1908]. seine [3254, 150, 725]. seiner [164, 1219]. seines [2183]. Seismicity [2836]. Seismik [1385]. Seite [2182]. Sektionen [2485]. Selborne [2162]. Selbstverstandnis [67, 1371]. Selden [566]. Selected [1309, 943, 1021, 596, 3745, 3784, 2572, 1591, 275, 84, 711, 2070, 3430, 793]. Selection [1888, 526, 2482, 1061, 3433]. Selections [3390]. Selective [1167]. Selectivity [1630]. Self [3293, 638, 149, 2082, 1272, 3698, 508, 2899].

Self-Fashioning [2082]. Self-Made [149]. Selfhood [2379]. Seligardi [2267]. Selleck [3433]. Sellers [1578]. Sellés [625]. Selling [373, 2860]. selon [2537]. Selves [3160, 487]. Semantics [2102]. semence [2534]. Semi [1030, 1030]. Semiconductor [653]. sémiotique [1014]. semplici [25]. Semyonovich [3371]. senda [1394]. Senecas [564]. Sengers [1842]. Sengoopta [2918, 638]. Sengör [2330]. Sengupta [2762]. Senility [2899]. sens [2549]. Sensation [1000]. Sense [3353, 3707, 3217]. Senses [1437]. Sensibilities [8]. Sensibility [2746, 3081]. Sensing [3478]. Sent [224]. Sentiment [3698]. sentimentale [1806]. Separating [3567]. September [3543, 2731]. Septimus [891]. Ser [2997]. Serafina [862, 978]. Serber [384]. Serendipity [2102]. Serfati [2820]. Serge [3320]. Sergey [531]. Serialized [2473]. Series [3315, 3351, 3497, 680]. Serjeantson [706]. Sermons [3295]. Serpent [1893]. Serra [2279, 3327]. Serres [1084]. Servants [8]. Service [2108, 3166, 2980, 3758, 43]. Services [2993]. servir [1735]. ses [1310, 1014, 716, 2822, 3568]. Sessé [2051]. Sesto [1685]. Set [1515]. Setel [218]. Seth [1035, 3089]. Settecento [3394]. Setting [1704]. Settlers [1844]. Seven [3449, 2666, 2878, 784, 741]. Seven-Year [2666]. Seventeenth [1446, 1632, 1633, 1501, 1974, 1506, 3735, 1922, 706, 2608, 588, 1378, 3661, 2409, 886, 879, 1451, 3246]. Seventeenth-Century [1632, 1633, 1506, 1922, 706, 1378, 3661, 2409]. Seventh [2022]. Severinus [2541]. Severo [332]. Sewall [1360]. Sex [1321, 2614, 638, 2918, 2668, 2443, 2951, 1452, 3598, 3006, 87, 2508, 856, 2945, 898, 297, 1484]. Sexing [1150]. Sexual [1120, 1451, 1399]. Sexualitätstheorie [2276]. Sexualities [2196]. Sexuality [1744, 2221, 1669, 1150]. Sexualized [3716]. Sexually [218]. Seymour [229, 2856]. Sforza [2641]. Shackelford [2541]. Shadow [1848, 3059, 2332, 2839, 2439, 2357, 120, 649]. Shadows [2939, 1742]. Shah [2377]. Shaikh [319]. Shakers [3088]. Shakespeare [487, 986]. Shall [2844]. Shamdasani [2433]. Shame [566]. shan [3343]. Shane [1418]. Shanly [292]. Shape [2697, 3220, 3357, 683, 842, 3658]. Shaped [3669, 705]. Shaping [1088, 1362, 2060, 1935, 1662, 3299, 1918, 2023, 962, 2633]. Shapiro [535]. Shared [1204, 3411]. Sharon [189, 2755, 2635, 2628, 1001]. Sharp [993]. Sharpe [420]. Sharrer [766]. Shaver [710]. Shaw [399]. Shawn [1108]. Shea [2050, 2673]. Sheet [2420]. Sheets [2040]. Sheffield [3225]. Sheila [805, 567]. Sheldon [551]. Shelley [2628]. sheng [3186, 3219]. Shenitzer [799, 799]. Shepard [563]. Shephard [1136, 2219]. Sherlock [2017]. Sherman [3651, 883]. Shermer [1848]. Sherry [851]. shi [3561, 3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3306, 3277]. Shifting [2916]. Shifts [3083]. shigao [3306]. Shigeru [1406]. Shimon [1968]. Shinn [3361, 3433]. Shipman [1005]. Ships [541, 3155]. Shiv [17]. Shock [2990, 2910, 1749, 2690, 3686]. Shocked [3078]. Shockley [3090]. Shoe [103]. Shoe-Fitting [103]. Shoemaker [923]. Shoolery [1548]. Shores [977]. Short [3755, 940, 3608, 451, 3637, 3135, 3306]. Short-Title [940]. Shorter [1582, 404, 822]. Shot [2502]. Shots [2502]. Show [3674]. Shrum [3524]. shu [3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277].

Shuchun [2822]. Shuldiner [2369]. Shulman [3089]. Shurkin [3090]. shuxue [3306]. Siberian [815]. Sichtbar [1284]. Sick [2114, 2844, 2931, 376]. Sickle [2446, 3198]. Sickness [975, 1127]. Sida [2174]. Siddiqi [2378]. Side [2911, 148]. Sidewinder [543]. Sidgwick [2765]. Sidgwick-Eye [2765]. Sidney [711, 2329]. siècle [982, 2733, 494, 3566, 2971, 3482, 3639, 1691, 716, 361, 3740, 1352, 3388, 2408, 3157, 3495, 2091, 3159, 2675, 2549, 1226, 3152, 2278, 1735, 3310]. siècles [2738, 3494, 614, 939, 3635]. Siegel [219]. Siegfried [2269]. Siegmund [642]. Siegmund-Schultze [642]. Siena [2268]. Sierra [1048, 634]. Sight [1378]. siglo [146, 300, 2430, 2995, 625, 3174, 60]. siglos [449, 1696, 2386]. Sigmund [2154]. signes [1014, 1686]. Significance [621, 1404, 3014]. Signor [1865]. Signs [2603, 1078, 2039, 3393, 3309]. Sigrist [1243, 627]. Sigrist.-B [1243]. Sigurdsson [669, 760]. Sik [870]. Sikorskii [2499]. Silencing [1616, 532]. Silent [62, 1028]. Silicon [2780, 2551, 1965]. Silkworm [311]. Silver [398, 1035]. Silverman [2834]. Silverthorne [59]. Silvia [2495]. Silvio [477, 159, 1929]. Simian [3676]. Simili [3141]. Simitis [62]. Simon [2678, 2744, 884, 510, 3475, 2426, 354, 995, 1871, 3042, 2825, 796, 2069, 730, 61, 869, 1023]. Simone [2044, 2347]. Simonton [689]. Simpson [2919]. Simson [205]. Simulation [406]. Simultaneity [3136]. Sin [62, 856]. Sinan [843, 1226]. Since [857, 1311, 1327, 3072, 1359, 2196, 3548, 1078, 2172, 2990, 1544, 1549, 550, 3456, 438, 1125, 820]. Singer [3331, 2201]. Singing [3624]. Sinha [3678]. Sinica [3634]. Sinnbildes [1574]. Sinnlichkeit [1584]. Sinnstiftung [3578]. Sinuosity [2712]. Sioban [2197]. Siobhan [3087]. Sir [621, 3338, 519, 1507, 3241, 889, 765, 1104, 3772, 2482, 588]. Siraisi [2893, 612]. sistema [58, 498]. Sisters [852, 2844]. Site [1768]. Sites [3412, 845]. Situated [209]. Situating [1796]. Sivasundaram [3003]. sive [1238, 2040]. Sivin [2249]. Six [3069, 3427]. Sixteenth [3235, 588, 1227, 3650, 1451]. Sixteenth-Century [3650]. Sixtieth [3196]. Sixty [3191]. Size [2697]. SJ [3406]. Skeletons [1647]. Skelettestreit [1453]. Skid [736]. Skies [3683, 989, 1764]. Skilled [434]. Skills [2798]. Skin [489, 220]. Skinner [2055]. Sklansky [2379]. Sky [31, 245, 2972, 2176]. Skywatchers [1214, 1214]. Slade [3091]. Slaton [2434]. Slavery [3669]. Slaves [1511]. Slavings [279]. Slawik [2563]. Slaybaugh [660]. Sledge [3473]. Sleep [3717, 897, 3229]. Sleigh [3069]. Sloan [604, 408]. Slopes [2791]. Slotkin [62]. Slotten [2435, 918]. Slow [2519]. Small [403, 3620, 639]. Smallman [296]. Smallman-Raynor [296]. Smallpox [1830, 3422, 1962]. Smelling [3478]. Smets [472]. Smil [1016]. Smith [1044, 2086, 2436, 2542, 774, 3043, 851, 3434, 589, 2935, 1364, 772, 2315, 2465, 443, 794, 717, 463, 110, 746, 403]. Smithsonian [3340, 371, 1012]. Smoke [3072, 914]. Smoot [3721]. Smout [281]. Snait [665]. Snead [544]. Snow [1018]. Snowden [2920]. Snyder [1309, 1040, 3004]. Sober [604]. sobre [1492, 2561, 2174, 479, 2148, 141]. Sociability [3117]. Social [3631, 1180, 1328, 949, 950, 1061, 1983, 1832, 2735, 1088, 184, 3535, 2528, 1738, 3678, 413, 559, 1069, 3784, 517, 3502, 173, 3522, 219, 934, 737, 2562, 2478, 1302, 1775, 785, 1722, 2003, 2371, 251, 2861, 2447, 2399, 1885, 744, 3139, 2665, 1774, 1755, 3517, 2497, 2594, 2708, 1253, 567, 2812, 2908, 379, 3003, 3225, 3526, 694, 2293, 3688, 3088, 403, 1384, 3137, 3172, 1829, 2138, 3074, 2495, 2100, 402, 1551, 1302]. sociales [2675, 3320]. Socialism [1532, 927]. Socialist [221, 1268, 2957, 1268]. socialistischen [48]. Sociedad [3056, 2366, 60]. società [2115, 3319, 2173]. société [2675]. sociétés [1352]. Societies [342, 3724]. Society [1484, 2208, 579, 1636, 3108, 3625, 1406, 3776, 3778, 191, 596, 2911, 205, 3535, 2720, 776, 844, 3642, 3784, 517, 756, 3094, 678, 3252, 1229, 1382, 2174, 1365, 3229, 2993, 2379, 1283, 2966, 1943, 448, 3225, 3253, 3333, 3190, 2709, 3587, 1395, 2080, 3473, 1449, 3268, 820, 362, 579, 99, 3456, 31, 1918, 2082, 28, 2507, 821, 1719, 3359, 1273, 936, 1675, 1776, 3004]. Socio [75]. Socio-Political [75]. Sociobiology [1200, 1037]. Sociohistorical [1162, 678]. sociologia [190]. Sociological [1022, 3094, 3229, 2596, 2102]. Sociologie [2921, 1300, 1552, 3361]. Sociologist [2070]. Sociology [3094, 933, 937, 2102, 3688, 1727]. Soder [144]. Söderqvist. [3226]. Software [2563]. soiaeuza [1267]. soigner [1400]. Soil [1175]. Sokal [595]. Solar [304, 2287, 3228, 538, 336]. Sold [1415, 1972]. Soldier [2169, 3489, 50]. Soldier-Scientist [50]. Soldiers [2081, 1136, 2219]. Sole [103, 2115]. Solid [478, 75]. Solis [190]. Solitary [1003]. Solomon [2095, 2617, 2775, 2745, 1229]. Some [3539, 3368, 577, 3381]. Somer [265]. Somerville [1255]. Something [2429]. Sommerfeld [1714]. Son [1248, 1310, 2312, 825, 1111]. Sonar [917]. Song [3624]. Songs [2576]. sonnambuli [2977]. Sonne [3239]. Sonne-Die [3239]. Sonnenspektrum [1386]. Sonnert [3070]. Sonu [2433]. Soo [158]. Soojung [2287]. Soojung-Kim [2287]. Sophie [685, 939]. Sophisticated [833]. Soraya [1111, 1421]. Sorell [154]. Sorgel [309]. Sorrenson [362]. Sorting [781]. sosloviia [1354]. Sothis [980]. Souden [263]. Soukup [3227]. Soul [3325, 2728, 378, 3508, 2358, 3131, 2379, 2558, 2049, 865]. Souls [3764, 1464]. Sound [539, 778, 3765]. Sounds [1921, 340]. Soundscape [1615]. Soup [930]. Soups [2784]. Source [465, 1473, 463, 466, 154, 3512, 2703]. Source-Based [463]. Sourcebook [1576, 3553]. Sources [2943, 3467, 1940, 850, 813, 2198, 2088, 741]. South [1984, 632, 3565, 985, 1401, 3752, 3737, 2626, 3223, 1422, 3081, 343, 859]. South-East [3752]. Southeast [1984, 3342, 1194, 2180]. Southeastern [308]. Southern [776, 990, 1831]. Southwest [2036, 3318]. souvenirs [2547]. Sovereign [2965]. sovetskogo [1267]. Soviet [2185, 1784, 2054, 2378, 2204, 2087, 2779, 3086, 3768, 1769, 531, 91]. Soviet/Russian [2087]. Sowjetischen [1985]. Space [602, 1174, 1309, 943, 3106, 3220, 3713, 1137, 2209, 2378, 3683, 3415, 2781, 3087, 2858, 3093, 1033, 1545, 1662, 208, 351, 1766, 2084, 2567, 1306, 1146, 3146, 312, 3470, 2508, 1884, 3510, 84, 122, 3350, 955, 1873, 2618, 2032, 2566, 633, 1344, 538]. Space-Time [3470]. Spaceflight [3611]. Spaces [1102, 406]. Spain [3732, 332]. Spallanzani [2977, 3161]. Spanish [2885, 1078, 2165]. Spark [930]. Sparked [1865]. Sparks [2784, 1523]. Spas [673]. spätantike [607]. Spatial [2373, 3313]. Speak [62]. Speaking [1165, 2201, 2667]. Spear [2692]. specchio [382]. Special [3642, 3610, 3473]. Specialist [2164]. Specialization [598]. Specialty [291]. Species [3438, 1846, 23, 289, 3183, 1845]. Specimen [724]. Specimens [1706]. Spectacle [521]. Spectacular [1667]. Spectre [172]. Spectrometry [734]. Spectrum [3315, 1520, 2787, 1382, 1281]. Speculation [2334, 3767]. Spéculations [1691]. Speculative [482, 2889]. Speed [3173, 2770]. Speich [2184, 2855]. spektrale [1386]. spekulativ [3578]. spekulativ-universalistischer [3578]. Spence [2520]. Spencer [2112, 3256, 1336, 2786, 3420]. Spencerian [3450]. Spenser [487]. Spermate [470]. sphère [2896]. Spheres [2819, 1344, 582]. Spherical [3037]. Spiegel [164, 1964]. Spiegels [2182]. Spinning [2247]. Spiral [764]. Spirit [2621, 1099, 3619]. Spirited [1808]. Spiritism [3595]. Spirits [3657, 2329, 878]. Spiritual [2490, 2488]. Splitting [924]. Sponberg [2539]. Spontaneous [1523, 1398]. Spoof [3433]. Spot [3287]. Spread [3238]. Spreading [905]. Spreadsheets [2325]. Spring [1021, 1028]. Springer [2835]. Springford [192]. Spurensicherung [52]. Spurious [2360]. Sputnik [1749]. Spying [3609]. Square [439]. Squaring [1231]. Squier [2546, 374, 2570]. St [1609, 3653]. St. [4, 753]. staatliches [3477]. Stachel [2096, 1975]. Stadtische [164]. Stage [235, 3313, 1053]. Stages [2327, 605]. Stahl [1623, 897, 994]. Stakes [546]. Staley [2445]. Stalin [1620, 2779, 3086]. Stallings [409]. Stammbach [387]. Standage [894]. Standard [3368]. Standardizing [2305]. Standards [2234]. Standardwerkes [1964]. Stands [3264]. Stanford [2936, 2551, 76]. Stanley [1577, 3419, 2756, 3078, 258, 2828, 1083, 2297]. Star [3272, 350, 245, 781, 464]. Stargazer [368, 3142]. starnitsy [2499]. Starring [2934, 1868]. Stars [1996, 31, 1137, 541, 1466, 1383, 3041, 984]. Start [2559]. Start-Up [2559]. Starving [1422]. State [2054, 1380, 2854, 3588, 804, 3488, 75, 43, 3377, 1032, 2956, 1252, 2924, 534, 2637, 397, 3182]. Statecraft [1369]. States [2058, 918, 1554, 1861, 1135, 421, 1758, 1757, 93, 652, 3076, 912, 1057, 3263, 2929, 2165, 331, 375, 977, 2443, 201, 308, 925, 733, 1528, 3302, 1733]. Statesman [2432, 3489]. Statik [719]. Station [421, 3424, 3076, 312]. Statistical [280, 631, 3029, 2199, 2454, 2066, 503, 339]. Statistics [280, 2906, 1252, 1642, 339]. Stätten [2661]. Status [2318]. Staum [1943]. Steam [396, 1295]. Steel [1532]. Steele [1363, 2821]. Steen [1967]. Stefan [2607, 2897]. Stefani [2270]. Stefania [1428]. Stein [1567]. Steiner [2921]. Steinhart [800]. Steinke [2746]. Steinle [3071]. Stellar [1383]. stellarum [3041]. Stem [2300, 2120]. Stench [3242]. Stengers [1593, 1050, 1087]. Stenhouse [722]. Stennis [312]. Stenone [1999]. Stepchildren [1120]. Stéphane [3157, 3248, 3060, 2492]. Stephanie [118, 3040]. Stephen [1309, 456, 155, 2345, 516, 669, 652, 3529, 528, 215, 1884, 339, 3031, 2613, 2416, 977, 1419, 2104, 2466, 2890, 1455, 2925, 2055, 331, 2597, 1391, 1053, 2207, 1346, 1097]. Stephens [3005, 632]. Stephenson [2106, 686]. Stereochemistry [2373]. Sterilization [3302, 2091]. Sterling [1150, 1420].

Stern [2849, 302]. Sternen [3500]. sterrenkunde [499]. Steve [664]. Steven [446, 287, 193, 566, 178, 3488, 415, 2556, 2223, 3059, 871, 2703, 2727, 2625, 2176, 2481, 2507, 1567]. Stevens [3641, 1119, 598]. Stevenson [802]. Stevin [2825]. Stewart [731, 26, 3255, 2980, 3626, 2551, 221, 759, 623]. Stiff [1810]. Stigler [339]. Stimulus [267]. Stink [2987]. stirpium [616]. Stix [669]. Stocker [458]. Stockholm [2472]. Stockinger [95]. Stocklöv [630]. Stoffel [2202]. Stokes [927]. Stoll [27]. Stoltzenberg [2380]. Stomach [1422, 2799]. Stone [2641, 3675, 261]. Stonehenge [1215, 263]. Stonemason [3602]. Stop [3097]. Stopping [2296]. Storey [57]. stori [1517]. Storia [2044, 753, 2977, 2986, 2504, 777, 493, 3138, 3319, 1465]. Stories [644, 1411, 902, 2564, 3427, 3184, 1531]. Stork [204]. Storm [1261, 3520]. Storms [3489]. Story [3719, 924, 404, 2255, 654, 894, 2045, 3683, 1756, 601, 3014, 380, 1469, 1304, 546, 394, 1874, 505, 566, 1877, 822, 1926, 3178, 1878, 2566, 538, 2633, 439, 3200, 775, 1846, 3550]. Stow [1119]. Strahl [2846]. Strahlen [3715]. Strahlenschutzes [1956]. Strahlungssorten [3715]. Strait [177]. Straker [2466]. Strange [3559, 1949]. Strangelove [3434, 2503]. Strangers [2830]. straordinarie [1694]. Strasse [2308]. Strasser [3355, 377]. Strategic [2304, 2568, 926]. Strategien [1584, 2258]. Strategies [2109]. Strategy [2123]. Stratification [559]. Stratton [2725]. Strauss [1849, 1112]. Streams [2554]. Street [904]. Streets [910]. Strick [2925]. Striking [2932, 13]. String [2039]. Striving [2768]. Strohmaier [2480]. Stroud [513]. Structural [631]. Structure [973, 1311, 999, 3356, 2373, 1525]. Structures [3105, 3640, 2595, 681, 3583, 3524]. Struggle [2987, 1158, 2502, 3508, 3707, 317, 1040, 1208, 1156, 2691]. Struggles [3677, 189]. Struik [1072]. Stuart [2165]. Stuber [2897]. Stuckrad [2819]. Student [2688]. Studi [2988]. Studie [162]. Studien [642, 3657, 2347, 360, 3013, 1681]. Studies [3002, 1436, 1633, 1979, 3319, 2472, 3467, 2134, 2883, 3222, 3670, 3646, 3507, 1046, 609, 1738, 3304, 2402, 3633, 107, 3760, 3094, 3252, 1871, 3259, 2994, 901, 588, 3219, 933, 3154, 1383, 2996, 3232, 3351, 3189, 3773, 1345, 3529, 3488, 3497, 3486, 3322, 3731, 3238, 3747, 3190, 415, 741, 22, 2972, 3512, 688, 3644, 2536, 138, 784, 2975, 2934, 1059, 3530, 2413]. Studio [2022, 1530]. Study [1849, 2419, 1362, 2545, 3670, 1494, 1848, 3029, 980, 1162, 454, 3143, 3037, 1426, 3468, 307, 1631, 879, 3332, 3333, 2102, 2313, 3268, 41, 2782, 2390, 463]. Studying [1426]. Stuff [2508]. Stuffed [1665]. Stukeley [2669]. Stukenbrock [2485]. style [2733]. Styles [1513, 3739, 2714, 2136, 3411]. sua [3138]. Sub [218]. Sub-Saharan [218]. Subject [1327, 253, 1927]. subjective [518]. Subjectivity [3440, 780]. Subjects [3502]. Subjektivität [1371]. Submarine [730, 389]. submarinos [1952]. Submission [657]. Subodh [453]. Subrata [645]. Substantiale [618]. Subtext [3647]. Succeeding [210]. succès [3483]. Success [2703]. sud [968]. Sudan [393]. Sudden [289]. Sudhoffs [3727, 3260]. Sue [2941]. Suffering [658]. Suffixes [2867]. Suggestions [3541]. Suhayl [1689]. Suhrkamp [3013]. sui [2737].

Suisse [3495, 2091]. Suit [1306]. suivi [3348]. Sujit [3003]. sulla [3319, 58]. sull’elettrometria [2977]. Sullivan [308]. Sum [3153]. Sumathi [2760]. Sumer [2325]. Summa [3390]. Summers [2856, 197]. Sun [2429, 31, 140, 892, 3408, 2287, 1698, 2851, 336, 980]. Sundials [1807, 1498, 1807]. sung [653]. Sungook [1864]. suo [3186, 495, 1517]. super [2041, 473]. Supernatural [3562, 2107]. Superpower [3076]. Supersonic [2770]. Superstrings [1957]. Superstructures [2415]. Superweapon [3434]. Supplement [908, 790]. Supplementary [3410]. Supply [1412, 1487]. Suppressed [2345]. Suppressing [2086]. Suppression [3089]. Supreme [3775]. Surface [3559]. Surfacing [2908]. Surgenor [1747]. Surgeons [2932]. Surgery [2932, 2233, 2218, 394, 166, 2750, 489]. Surgical [3252, 2679]. Surveillance [3588]. Survey [592, 2720, 3143, 804, 2355, 3333, 1862, 1930, 1175, 2028, 1327, 1554]. Surveyed [2346, 1114]. Surveying [981, 511]. Surveys [3187]. Survival [549, 540, 3270]. Suryadasa [2340]. Suryapraka’sa [2340]. sus [1015, 2263]. Susan [549, 464, 2771, 176, 767, 640, 135, 781, 1122, 1568, 2188, 2740, 2790, 2570, 2009, 2079, 2857, 806]. Susantha [570]. Suspect [1023]. Suspensions [521]. Sustainability [2767]. Sustaining [801]. Sutton [227]. Suzanne [353, 3225]. Svante [677]. Sveinn [2420]. Sven [2657]. svensk [504]. svenska [2080]. Sverre [1261]. sviazi [1314]. Svoboda [1536]. Swedenborg [998]. Swedish [1383, 2080]. Sweeney [1946]. Sweet [2972]. Sweetnam [1382]. Swerdloff [186]. Swerdlow [461]. Swisher [1416]. Swiss [3752]. Sydel [2834]. syllepticus [3150]. Syllogismo [3390]. Sylvain [3649, 939]. Sylvester [2989, 168]. Sylvia [2667, 2960]. Sylvie [3739]. Symbiosis [3551]. symbolique [2820]. Symbolism [2026]. symbolum [1238]. Symmetry [3095, 441, 2707]. Symposion [1319]. Symposium [3100, 307]. Synaesthesia [187]. Synapse [3180]. Synchrony [2883]. Syndrome [2931, 1076]. Synoptic [807]. Syntax [2028]. Synthèse [518]. Syntheses [75]. Synthesis [19, 1369, 3200]. Synthesize [1419, 3200]. Synthesizer [1762]. Synthetic [3000]. Synthetischen [199, 2553]. Syphilis [489, 1275]. System [2416, 1301, 3760, 915, 1954, 314, 3228]. systématique [3248]. système [2426, 518]. Systems [3358, 1217, 3752, 1043, 1197, 2597, 304, 1043, 1546]. Szász [1912]. Szientismus [3763]. Szulakowska [3055, 2043].

T [2947, 418, 446, 731, 783, 2054, 2270, 631, 2983, 1423, 2217, 69, 3057, 475, 1618, 981, 305, 187, 1575, 285, 148, 23, 420, 116, 1045, 2369, 340, 1824, 1593, 1061, 1868]. T. [2215, 2951]. Tabit [843]. Tabitta [145]. Table [3014, 339, 2357]. Tables [2404, 1185, 2325]. Taboo [780]. Tabrizi [2635]. Tadhkira [2245]. tai [3343]. Tailings [1144]. Tainted [696]. Taints [2867]. Take [2844]. Taken [3051]. Takes [3658]. Taking [2006, 2030, 1294]. Tal [665, 2348]. Talal [3095]. Talassografico [2842]. Tale [2052, 3408, 79, 3065, 698, 439, 800, 21]. Talento [3053]. Talents [3216]. Tales [2017, 674, 1763, 2154]. Talk [2241, 3266]. Tamera [2821]. Taming [2418, 321]. Tangled [1270]. Tanks [313]. Tannenbaum [1594]. Tansey [3031]. Tantillo [2167]. Tapan [2762]. Tappan [893]. Taraboletti [693]. Taraboletti-Segre [693]. Tarihi [1370, 583]. Tarski [2617]. Tartaglia [704]. Taschenbuch [3013]. Taschwer [2182]. Tasmanian [1264]. tasselli [382]. Tasting [3478]. Tatiana [2095]. Taton [123, 2467]. Taub [2150, 3196]. Tauber [555]. Taught [2557]. Taussig [3739, 3092]. Tavoni [495]. Tavormina [2945]. Taxol [1304]. Taxonomy [3209]. Tay [3198]. Tay-Sachs [3198]. Taylor [1958, 3497, 591, 3407, 1825, 395]. Teacher [2759]. Teachers [2021]. Teaching [2022, 1509, 1739, 2601, 819, 1520]. Tech [546, 2780]. Technical [3601, 3194, 646, 3681, 2356, 606, 3537, 2360, 3634, 2033, 2976]. Technik [636, 3171]. Technikentwicklung [3193]. Technique [1201]. Techniques [1520, 3159, 3661, 16, 12]. Technische [194]. technischen [1042]. Techno [3773]. Techno-Cities [3773]. Technological [3132, 2208, 2307, 1772, 21, 1680, 1427, 206, 2801, 3377, 3119, 1547]. Technologies [3110, 853, 1031, 752, 305, 3488, 392, 2656, 1058]. Technology [1181, 948, 1438, 1633, 2118, 3099, 3448, 1406, 3324, 2906, 3670, 2027, 3507, 3275, 3261, 3338, 2983, 1010, 207, 1266, 2427, 3144, 66, 2306, 3359, 2875, 1163, 918, 3271, 3530, 1194, 539, 3094, 1579, 927, 1821, 3091, 653, 3226, 3259, 2562, 1290, 3681, 2963, 2990, 2994, 3641, 3748, 1533, 2371, 3001, 1759, 3765, 3472, 1885, 721, 735, 1285, 1283, 3430, 1306, 1130, 3232, 2637, 3501, 3189, 93, 3665, 1431, 17, 2952, 2964, 1595, 3659, 3322, 2024, 536, 3666, 437, 3524, 2709, 132, 3080, 1480, 3258, 1927, 272, 2071, 718, 2782, 2941, 2098, 2907, 2217, 2559, 863, 251, 2859, 139]. Technology [250, 231, 293, 1308, 947, 1062, 3719, 1088, 131, 1305, 2356, 3139, 1903, 3354, 3480, 1286, 22]. Technopolitics [2770]. Technoscience [2931]. Technoscientific [663]. tecnica [1462]. tecnología [1361]. Tectonics [1296, 2330]. Ted [2162, 2851]. Tegni [352]. Teich [1402]. Teilhard [2072]. Telecommunications [2187]. Telegraph [852, 1533]. Telegraphic [852]. Telegraphy [1863]. Teleology [2880]. Telescope [3404, 1033, 3142]. Telescopes [2859, 2826]. Television [918, 1145, 3774, 3423]. Telle [769, 95]. Teller [2503, 1548]. teme [2252]. Temkin [2008]. Temperature [2734]. Tempered [308]. Temple [269]. temps [161, 3310, 2312]. Temptation [2679, 3468]. Ten [3122, 455, 1805, 498]. tension [190]. Tenth [3727, 1365]. Tenth- [3727]. teologia [3395]. Teoría [2742]. Teorie [493]. Tercentennial [2975]. Terence [260]. Teresa [2945]. Terman [2551]. Terms [2655, 1147]. terra [493, 3729]. Terrains [453]. Terrall [2271]. Terrarum [3146]. terre [3494, 1243, 2886]. Terrestrial [2923]. Terrie [1853]. Territorial [1662, 2082]. Territories [3710]. Territory [1403]. Terror [1838, 1087]. Terry [3361, 766, 2546]. Tertius [2729]. Terzian [811]. Tesla [1518]. tesoro [25]. Tesoros [1696]. Tessenow [3726]. Test [2902, 3185]. Test-Ban [3185]. Testi [2988, 1483]. Testigos [2264]. Testimony [1892, 2348]. Testing [2212, 2781, 2084]. Tewarson [520]. Texas [59]. Texeira [2263]. Texera [200]. Text [3647, 901, 3146, 34, 326, 3634, 2934, 2533, 609]. Textbooks [341]. Texte [2607, 3486, 3171, 2879, 3483, 2406, 1489]. Texten [719].

Textes [761, 2824, 3566, 790]. Texts [2945, 3662, 62, 2808, 1138, 1224, 3724, 3488, 3651, 3486, 326]. Textualization [2382]. TH [3171]. Thackray [812, 96]. Thaddeus [3587]. Thagard [217]. Thales [2251]. Thames [1283]. Theatermaschine [3402]. Theaters [2858]. Thebesius [1826]. Theerman [371]. Their [3194, 1998, 644, 179, 1679, 3162, 2655, 1844, 1020, 1466, 2188, 1759, 1212, 351, 1351, 322, 114, 2355, 2193, 341, 3178, 2564, 189, 41, 892]. Theistic [1661]. Them [3287, 2420]. Theme [2576]. Themes [592, 3061, 315, 1662, 3480]. Themselves [1844]. Theobald [2086]. Theodor [3605, 3178]. Theodore [2979, 2199, 2749]. Theological [267, 408, 143]. théologie [2343, 3740]. Theology [2882, 3646, 1507, 330, 516, 2311, 3486, 648, 3660, 3381]. Theophile [48]. Theophilus [1556]. Théophraste [1832]. Theophrastus [3309]. Theophysis [3063]. Theoretical [2789, 1318, 1967, 1683, 915, 2889, 442, 2749, 2965]. Theoretische [2276, 3715, 10]. Theorie [1324, 1944, 199, 474, 1386, 2406]. Theorien [2846]. Theories [2109, 905, 869, 2565, 1670, 3577, 681, 1259, 941, 120, 457, 1515, 2534, 2343, 418, 3106, 1366, 615]. Theorist [482]. Theory [2789, 1061, 1889, 3105, 3467, 80, 3640, 1467, 2683, 888, 3685, 1613, 26, 555, 3255, 976, 2991, 1236, 1250, 1211, 864, 3154, 1296, 1524, 2797, 2072, 1364, 3747, 1413, 2417, 3624, 3512, 1839, 1957, 2206, 2684, 567, 1039, 112, 1432, 1257, 1525, 1537, 1514, 1472]. Therapeutic [92, 2914]. Therapeutics [237]. therapeutischen [1089]. Therapie [2144]. Therapies [1393]. Therapy [3686, 3199]. Thermodynamik [3177]. Thermonuclear [2635]. These [1743]. Theunissen [906]. They’re [299]. Thick [2803]. Thierry [2737]. Thiher [2629]. Thin [1873]. Thing [2357, 2448]. Things [2666, 2802, 898, 2462, 2801, 781, 1398, 2241]. Think [2658, 2788]. Thinker [621]. Thinking [1285, 2574, 3249, 254, 230, 519, 3246]. Third [738, 3753, 244, 2241, 317, 1274, 2652]. Third-Wave [2652]. Thirteenth [871, 2729, 1697]. Thirteenth- [1697]. Thirty [3100, 205]. Thirty-Five [205]. Thirty-Second [3100]. Thomann [520]. Thomas [2742, 1058, 390, 2472, 268, 1704, 731, 404, 2985, 164, 1049, 199, 3226, 1949, 720, 11, 278, 3472, 588, 481, 1611, 1004, 224, 1915, 617, 3165, 682, 190, 852, 2118, 2865, 2058, 1858, 1944, 2308, 2670, 2919, 2103, 2634, 2553, 1800, 3078, 2218, 2059, 2465, 2309, 1635, 3398, 3687, 1199, 481, 1374, 3504, 791, 664]. Thompson [1615, 2656, 127, 581]. Thomson [1009, 192, 379]. Thoreau [1724, 807]. Thorndike [2681]. Thornton [1631]. Thorsheim [3072]. Those [1426]. Thought [605, 2881, 2630, 1268, 2330, 3336, 3660, 2919, 1365, 699, 2379, 784, 997, 1220, 2745, 2862, 3173, 1911, 2272, 2890, 2597, 2489, 3707]. Thoughts [3264, 2958, 2462]. Thread [2833, 311]. Threat [3448, 409]. Three [3315, 3760, 1022, 1364, 3775, 2702]. throughout [2965]. Throwing [2024]. Thümmel.J. [1815]. Thurmond [3144]. Thy [871]. Thylacine [1264]. Thyrza [932]. ti [2412]. tian [3343]. Tidal [2797]. Tides [3599]. tiempos [1585]. Tier [1505]. Tieren [500]. Tietge [2097]. Tiger [1264, 653]. Tigers [1348]. Tightrope [2381]. Tightrope-Walking [2381]. Tileston [1340]. Tilman [3410]. Tilney [2932]. Tilton [2490]. Tim [1909]. Timaeus [1090, 1216]. Time [3106, 243, 1951, 2146, 1467, 966, 522, 3415, 1139, 475, 456, 1414, 3252, 3498, 2858, 2296, 208, 351, 3050, 3439, 1040, 3675, 880, 2623, 753, 2743, 3470, 641, 444, 2154, 1875, 972, 2729, 924, 2895, 2496, 122, 1442, 783, 3357, 1538, 475, 391, 1191]. Timekeeping [2883]. Times [1978, 3142, 1157, 244, 2330, 1363, 938, 2779, 713, 904, 3170, 114, 344, 1079, 116, 340, 343, 857, 2938, 438, 3297, 664, 1218, 850]. Timothy [531, 1031, 2190, 1566]. Tina [2929]. Tinbergen [2615]. Tints [2867]. Tirard [3060]. Tissue [2933]. Titel [3477]. Titelbilder [3405]. Title [940]. Titmuss [1069]. Tjossem [2783]. Tobacco [3618, 1396, 399]. Tobies [170]. Tobin [2272, 2500]. Toby [326, 1044, 962]. Today [621]. Todd [132]. Todes [692, 2290]. Toil [3050]. Toine [2860]. Tokugawa [1377]. Tokyo [2472]. Told [1397]. Toledo [2404]. Tolerance [482]. Toleration [3401]. tolleranza [3102]. Tolley [1835]. Tom [149, 420, 154, 2946, 894, 1137, 1841]. Tome [3348]. Tomes [637]. Tone [88]. Tonempfindungen [3068]. Tönenden [814]. Tong [806]. Tongue [3676]. Tongues [1436]. Tonso [210]. Tony [2989, 924]. Too [731]. Tools [942, 2428, 1427, 3441, 353, 2315]. Toomer [462]. Top [2445]. Topalov [716]. Topham [805]. Topic [3120]. Topical [2036]. Topografien [2184]. topologie [1014]. Topology [747]. Torches [2170]. Tore [3633]. Torealis [3633]. Tormented [3178]. Tornado [1764]. Torner [2051]. Torok [186]. Toronto [2993, 1575]. torpedini [2352]. Torrens [1388]. Torres [177]. Torretti [668]. Torti [620]. Tosi [3238]. total [3200]. Toten [2372]. Touching [3478]. Toulmin [2104]. Tour [2019, 3714]. Touret [2850]. Tourette [1076]. Tout [2877]. Touwaide [3232]. Tower [269]. Town [2825, 3759]. Townes [1410]. Toys [1759]. tra [3138, 1049]. Traces [208, 227, 3385, 968]. Tracey [1007]. Tracing [2552]. Tractatus [3150]. Tracts [158, 1921]. Tracy [2693]. Trade [3283, 1621, 2730, 1631, 299]. Traders [1237]. tradicional [2662]. Tradition [1177, 1967, 2557, 609, 3251, 2061, 1168, 585, 3555, 3743, 3723, 3031, 3230, 361, 2732, 2400, 2453, 1982, 1687]. Traditional [342]. traditionale [56]. Traditions [1312, 1443, 1259, 3487, 2155]. tradizione [2403]. Tradizioni [1381]. traduction [2879, 2879]. Traduire [2278]. Traffic [3596, 1253]. Tragedy [773, 3408, 3065, 3296]. Tragic [2718, 3707, 992, 3197]. Training [2323]. Traister [995]. traité [2151, 472]. traités [3388]. Trajectoires [2452]. Tranquilizers [403]. Transactions [678, 3740, 2606]. Transatlantic [372, 2773, 508]. Transcendental [187]. Transcontinental [180]. Transcription [1690]. Transfer [132, 272]. Transfers [1202]. Transformation [3670, 1266, 2427, 3067, 3567, 3474, 2593, 2410, 2916, 2234, 1589, 2690, 1016, 3246, 76, 1097, 198, 179, 1879]. Transformationen [629]. Transformations [3670, 3539, 3259, 3717, 3189, 3322, 1202, 868]. Transformed [2791, 2666, 1273, 2658]. Transforming [1031, 1358, 3194]. Transit [3281, 2131, 687]. Transition [728, 1493, 1930]. Translated [3312, 3140, 3320, 3297, 3147, 2178, 3331]. Translating [1521]. Translation [1690, 3670, 3390, 1683, 102, 1917, 1365, 3143, 3037, 1377, 1364, 1191, 3150, 2245, 3041]. Translations [3365, 589]. Translunar [2400]. Transmission [1783, 1521, 609]. Transmitted [218]. Transmutation [2175]. Transmutations [1591]. Transmuting [3024]. Transnational [3541]. Transparency [1557, 3682]. Transparent [2785]. Transportation [2770]. transposer [2278]. Transsexuality [2443]. transversaliste [3361]. Tras [1394]. Trauma [1855, 1008, 1277]. Traumatic [1008]. Traumdeutung [1734]. Träume [1734]. Travaux [3495, 3361]. Travel [1446, 1494, 522, 1003, 113, 3473, 985]. Traveler [1538, 3064]. Travelers [1003, 3038]. Traveling [2898, 3173]. Travels [1721, 2102, 1540]. través [2366]. Travis [305]. Treason [1838]. Treasures [1083, 113, 1905]. Treasury [884, 1344]. Treatise [1993, 1583, 706, 2825, 3037, 3612, 3041, 2154]. Treatises [3312, 1582]. Treatment [3686, 3695]. Treetops [558]. Trefil [1077]. Treitel [2558]. tremblements [3494]. Trembley [2123]. Trends [2383]. tres [875]. tres-eccellent [875]. Trevor [2788, 1358]. Triads [3259]. Trial [1265, 2077, 3577, 1726, 2926, 3499, 3150, 3738]. Trials [3295, 2872, 237, 2926, 488]. Triassic [2281]. Tribute [2399]. Tributes [1424]. Tricksters [933]. Trifogli [2729]. Trinity [2945, 2511]. triomphe [582]. Trischler [1912, 295, 762]. Trismegistus [3297]. Triumph [636, 1037, 3296, 3050, 353, 1422, 1866, 3349]. Troesken [3073]. trois [3348]. Trophy [216]. Tropical [3755, 2171, 1948, 2478, 1126, 3663, 3407, 2736]. Tropics [2898, 3565, 903]. Trotula [2243]. Trouble [836, 414]. Troubled [3079, 1960, 3198, 623, 2367]. Troublesome [2050, 321]. Troubling [1666]. Troweling [2623]. Troy [176]. True [3485, 394, 505, 1878, 286]. Trueblood [616]. Truesdell [435, 2132]. Trunk [2117]. Trust [1185, 2310]. Trusted [277]. Truth [2510, 1200, 2889, 2696, 2660, 2935, 3523, 415, 2047, 2857, 1931, 380, 1564, 951]. Truths [3431]. Trying [3499]. Tsarist [1838]. Tseng [6]. Tsouyopoulos [3679]. Tsukerman [531]. Tsutsui [310]. tuan [3268]. Tuberculosis [736, 1743]. Tübingen [267]. Tubinger [188]. Tuchańska [1162]. Tuchman [3006]. Tucker [2052, 1140, 1841]. Tufte [3015]. Tulipmania [3244]. Tullio [2250]. Tully [1631]. Tumbleson [270]. Tuna [2478]. Tunisia [3002]. Turbines [718]. Turbulence [633]. Turbulent [79]. turbulenter [194]. Turkey [3181]. Turn [1186, 1744, 3288, 3141, 391, 1181, 948]. Turnbull [933]. Turned [550]. Turner [489, 480, 3739, 2730, 1836, 349, 264]. Turning [3764, 799, 2391]. Turquet [1592]. tussen [1936]. Twentieth [3194, 1784, 1446, 2108, 2111, 2939, 2472, 2429, 3276, 1068, 818, 3083, 193, 2929, 1738, 2924, 69, 3265, 396, 3141, 976, 938, 1743, 2548, 2914, 1948, 2174, 310, 2631, 2554, 3680, 3765, 770, 2699, 1755, 114, 1130, 3773, 928, 1136, 2193, 2299, 739, 2219, 1676, 2375, 816, 206, 913, 895, 1548, 2529, 1288, 2954, 673, 526]. Twentieth-Century [2939, 2472, 2429, 2924, 2548, 2914, 1948, 2631, 2375, 1548, 2954]. Twenty [3516, 3590, 3440, 2658]. Twenty-first [3516, 3590, 3440]. Twilight [63, 3512]. Twist [1671]. Two [2804, 1828, 2557, 2771, 193, 1738, 3454, 1909, 2321, 601, 2521, 1095, 195, 145, 79, 3262, 2996, 2966, 1630, 591, 340, 564, 1416, 3381, 2028, 94]. Two-Culture [2771]. Two-Faced [145]. Two-Headed [1828]. Two-Hundred-Year [601]. Tycho [3235, 703]. Types [3415]. Typhoid [3759]. Typhus [3002]. Tyranny [2345]. Tyrrell [286]. Tyson [513]. Tzvi [787].

U [2234, 3726, 2533, 3344, 1309, 943, 1175, 1755, 2191]. U.K [97]. U.S [2081, 1038, 313, 1031, 1131, 2567, 2176, 84, 314]. U.S. [2519]. Überlieferung [607]. Übersee [909]. Übersee-Museum [909]. Ubersetzung [34]. Übersetzungen [3486, 2607]. Ubrizsy [2401]. uchenogo [1354]. Udías [2477]. Uexküll [3500]. UFOs [193, 1307]. Uiblein [476]. Ulf [3685, 677, 3171, 2563]. Ulisse [2726]. Ulrike [1017]. Ultimate [1957]. umbral [1492]. Umbruch [2211]. Umbruchs [2189]. umfangreichen [719]. Umstande [126]. Unbound [2431, 626]. Uncertain [824, 1351, 1185]. Uncertainty [3329, 2927, 3748, 2931, 3508]. Uncharted [453]. Unconceived [2936]. Unconscious [186, 1649, 2179, 377]. Unconventional [1059, 2673]. und/als [1317]. Undead [1871]. Underground [112]. Undermining [3089]. Undersea [833]. Understand [1297, 1080]. Understanding [3158, 2248, 1416, 2223, 3010, 3098, 3470]. Understood [1844]. Undertaker [1246]. Underwood [2851]. Unemployment [1680]. Unendliche [3564]. UNESCO [3191]. Uneven [2374]. Unexpected [3408, 2139]. Unfashionable [135]. Unfettered [1162]. Unfinished [1917, 2825]. ungebändigte [1319]. ungheresi [2401]. Ungheria [2401]. Unicellular [1187]. Uniformity [3039, 3498]. Unifying [681]. Unimaginable [1232]. Union [2378, 385]. Unique [2451]. Unit [2586]. Unita [2842]. United [375, 2058, 977, 1528, 918, 1554, 1529, 1861, 1135, 2443, 421, 3302, 201, 308, 1758, 1757, 93, 652, 3076, 912, 1057, 1733, 3263, 925, 733, 331]. Unity [266, 864, 3039, 266]. Univers [3649, 2896]. Universal [2472, 3338, 2576, 685]. universalistischer [3578]. Universality [3061]. Universe [268, 1957, 2085, 15, 2621, 522, 3300, 1140, 2925, 3437, 3504, 2765, 122, 686, 966, 357, 440, 811]. universel [518]. Universitat [2133, 3515, 1230, 2143, 476]. Universitats [48, 2997, 3477]. Universitäts- [3477]. Universitats-bibliothek [48]. Universitäts-Mechanicus [2997]. Universitätsgeschichte [1644]. Universitätsjubiläum [2143]. Universitätsskandal [725]. Université [361]. Universities [1173, 643, 1695, 1700, 2201, 3653, 2725]. University [2113, 2472, 940, 3083, 2557, 69, 559, 2092, 76, 2551, 3196, 349, 1250, 3375]. universo [495, 3564]. Universum [3564]. Unknown [1309, 943, 565, 84]. Unmaking [478, 3614]. Unmix [1842]. Unnatural [167, 3046]. uno/Über [3563]. Unobservable [3527]. Unpublished [71, 1993]. Unrealized [3273]. Unreasonable [13]. Unschuld [3726, 2533]. Unselfish [604]. Unsichtbar [1284]. Unsichtbar-Sichtbar-Durchschaut [1284]. unsichtbare [2997]. Unsichtbares [3715]. Unsolvability [2198]. Unternehmensgründung [3193]. Untersuchungen [3134]. Unterwegs [500]. Unto [604]. Unveiled [686]. uomo [2252]. Updated [1214]. Upheaval [1202]. Uplink [1766]. Uplink-Downlink [1766]. upon [1375]. Uppingham [3759]. Uppsala [3238]. Ups [1030]. Uranium [1144]. Urban [179, 504, 2268, 381, 819, 344, 3472]. Urs [3736, 387]. Ursache [3563]. Ursprung [1574]. Ursprünge [3071]. Ursula [3322, 2428]. Ursus [39]. Urszula [3055, 2043]. Urton [2039]. USA [434]. Use [1063, 268, 2532, 603, 1292, 2957]. Used [3651]. Useful [959, 996, 2108]. User [2581]. Users [3735]. Uses [1259, 3650, 43, 857]. USS [1010]. usus [163]. Uta [3761]. Ute [274]. útiles [2995, 3056]. Utility [833, 537, 2958, 2720]. Utopia [2650, 2032, 1460]. Utopianism [1678]. Uwe [1944, 3260, 2818].

V [2941, 161, 712, 595, 1163, 2886, 83, 1143, 2087, 2225, 21, 203, 421, 344, 1002, 1032, 1314, 2943, 1267, 2191, 434, 3775, 1267, 2262]. V2 [1033]. Vaccination [2318, 2924]. Vaccine [2502, 3085, 317, 2502]. Vaccines [317]. Vaclav [1016]. Vadim [2499]. Vai [2726]. Vailati [41]. Val [112]. Valencia [165, 2133]. Valenciana [449, 1242]. valencianos [1602]. Valenius [1844]. Valenstein [2784]. Valentin [482]. Valeria [3141, 3360, 2986]. Valerie [696, 813, 392, 2088]. Validation [2758]. Valley [2780, 445, 374, 2551, 1965]. Vallisneri [3735]. Value [1725, 3190, 3278]. Value-Free [3278]. Values [2381, 3526, 783]. Valuing [2188]. Vancouver [579]. Vanden [2481]. Vanderjagt [568]. Vanishings [1833]. Vapores [1952]. Vardalas [2307]. Varey [884]. Variability [850]. Variations [2576]. Varieties [3539]. Variola [3422]. Variorum [3497]. Varnhagen [520]. Vasa [1826]. Vatican [3250]. Vatin [502]. Vauban [2350]. Vaughan [1374]. vegetabilium [163]. Vehicle [919]. veintiún [1500]. veke [2943]. Veld [2169]. Venereal [554, 2268]. venereo [1399]. Venezuela [200]. Veniamin [531]. Venice [1586]. Ventures [3229]. Venus [1186, 687]. ver [2343]. Vera [91, 3634]. Veranderung [171]. Verbindung [2116]. verdanken [3500]. Verdean [2623]. Verdú [2995]. vergleichende [2695]. Verhaltnis [274, 2103, 474]. Verità [3326]. vérité [1702]. vérités [2988]. Verkehrswege [2308]. verlorene [607]. Vermeer [2486]. Vermij [2273]. Vermonters [397]. Vernacular [3749, 470, 2562, 2976, 674]. Vernadsky [915]. Vernon [3118]. Vernunft [1584]. Veröffentlichungen [3477]. Véron [2671]. Versailles [431]. Verschiebungseffekte [1386]. Verse [15]. Versicherungs [707]. Versicherungs- [707]. Version [1214, 1364]. Versuche [163]. versus [1940, 2441, 698, 3712, 1894]. Verwaltung [2291]. verwandte [1386]. Very [800, 3135]. Vesalius [436]. vescovo [1999]. Vesely [2332]. Vestiges [1000]. Vetenskapens [504]. Vetenskapsbärarna [2080]. Veterinarians [2188]. Vexation [3008]. vi [2472]. via [2363]. Viaggio [3394]. viajes [2264]. Vicente [586, 479]. Vices [788]. victima [255]. Victor [146, 48, 586]. Victoria [247, 1425, 3744, 2972, 586]. Victorian [2989, 2107, 2555, 1523, 3754, 2985, 1950, 3328, 1941, 1853, 717, 1837, 2287, 1722, 2624, 901, 715, 2550, 1283, 3759, 1723, 285, 509, 3167, 1107, 2203, 2310, 3746, 3409, 641, 2179, 801, 2280, 3597, 3004, 910, 1000, 2068, 2457, 330, 997, 3413]. Victorians [2748, 996, 3421]. Victory [3008]. vida [449, 2386]. vie [2620, 2671]. Viejo [1820]. Vienna [1110, 2002, 638, 1011, 3329]. Vierge [2343]. Vies [3739]. Vieth [1138]. View [145, 3368, 2300, 2005, 1597]. Viewed [351, 2062]. Views [3669]. Vigor [1715]. VIH [2174]. VIH/Sida [2174]. VII [3348]. viii [2472]. vikingatåg [504]. Village [1932, 2766]. vils [3635]. Vim [1715]. Vincent [490, 140, 627, 3159, 2733, 825, 826, 341, 2973]. Vincenzo [495]. Vinci [1667]. Vinculum [618]. Vínland [2255]. Violence [3131, 3366]. Virchow [2285, 3178]. Virchows [3672]. Vires [163]. Virgin [1292]. Virginia [3410, 766, 820, 2478, 2283]. Virology [399]. Virtual [323, 1939]. Virtually [3455]. Virtue [886, 2981]. virtutibus [3483]. Virus [1770, 399, 1396, 1736, 1303]. vis [2582]. vis-à-vis [2582]. Visibility [3168]. Visible [2522]. Vision [3754, 2437, 2323, 2320, 309, 2274, 247, 143, 2687, 807, 3181]. Visionary [2425]. visione [3041]. Visions [2736, 237, 2422, 2032, 3188, 2773, 1834]. Visser [2850]. Visual [3222, 3036, 1520, 330, 1379, 524, 1364, 1108, 2523, 3620, 3222]. Visualization [237, 1955]. Visualizations [1906]. Visualizing [3232]. Visvanathan [17]. vita [3403]. Vitae [1374]. Vital [3298, 2266]. vitales [3248]. Vitality [2628]. Vitalizing [2611]. Vitamin [597]. vite [38]. Vitis [1374]. Vito [2842]. Vittorio [3012]. vivant [1190]. vivente [2270]. Vivisection [789]. vizcayno [479]. Vladimir [915, 989, 3371]. Vlahakis [3074]. Vo [1354]. Vocation [1003, 3708]. voci [3141]. Voelke [3075]. Voelkel [1349, 485]. Vogel [1291]. Vogt [556]. Voice [562, 2421]. Vol [2115, 1995, 3788, 3787]. Volcanic [760]. Volcano [590, 1075]. Volcanoes [669, 1205]. Volgarizzamento [35]. Volker [360, 635, 1604]. Volkstümliche [3486]. Volney [1404]. volontaire [2091]. Volta [2352, 2418, 1712]. Voltaire [3731]. Volterra [3170, 2842]. Volume [465, 425, 428, 828, 1180, 1331, 1330, 1790, 1789, 2228, 2227, 2514, 2513, 2794, 2793, 3114, 3113, 3445, 3126, 3129, 3444, 2883, 2027, 301, 1052, 871, 3340, 889, 354, 3035, 370, 469, 747, 3410, 805, 586, 1714, 3050, 2311, 1720, 651, 371, 3569, 78, 2731, 275, 2751, 3486, 75, 2315, 84, 3512, 582, 711, 3169, 2879, 192]. Volumes [3241, 3465, 3145]. vom [347, 24, 3477, 1017, 2144, 3761, 3515]. Voodoo [1973]. vor [642, 48, 1219]. vorfallenden [3103]. Vorgänger [1219]. Vorhang [1985]. Vorlesungsmitschrift [3747]. Vorlesungsverzeichnisse [1230]. Vorurteil [346]. Voyage [2556, 1875]. Voyager [2140]. Voyages [2275, 1070, 1589, 3771]. voyant [1718]. vraye [875]. Vries [1721]. Vrolijk [3646]. vs [3224]. Vucinich [2204, 1454]. vue [3348]. Vulcan [353, 1075]. Vulgarisierung [2258].

W [973, 2735, 247, 2472, 2912, 2189, 446, 2693, 2171, 160, 631, 818, 3046, 1359, 888, 2615, 2212, 1873, 366, 3340, 2445, 3642, 1393, 1144, 1527, 2850, 3252, 3259, 218, 1167, 203, 400, 478, 229, 321, 1846, 2190, 371, 2837, 337, 1609, 1004, 656, 1054, 979, 2024, 1566, 116, 2979, 591, 258, 2369, 2769, 84, 718, 897, 1937, 1424]. W. [2700, 1917, 858, 1586, 2859, 3031]. Waals [1842]. Wachhund [1956].

Wade [2274, 1039, 2612]. Wages [1127, 856]. Wailoo [2446]. Waitley [175]. Wake [1869]. Walcott [1012]. Waldby [2933]. Waldo [389]. Waldrop [1628]. Wales [420]. Walker [928, 3007]. Walking [2381, 2688]. Walks [1278, 1875]. Wallace [1848, 1392, 2960, 992, 1140, 2234, 780, 214, 2435, 2280]. Wallach [2141]. Wallach-Levy [2141]. Wallen [2385]. Waller [2968, 2658]. Wallis [3403, 2731, 1231]. Walls [1931, 176]. Walsh [2418, 2478]. Walter [80, 292, 590, 2839, 1124, 1825, 2763, 821]. Walters [1809, 1119]. Walther [3580, 728, 3171]. Walton [2727, 3488, 148]. Walzer [124]. Wandel [24, 2846]. Wandering [127]. Wandermenagerien [500]. Wang [3200, 77]. Waning [786]. War [3344, 3426, 2582, 2698, 2097, 1859, 2791, 548, 2081, 2784, 1747, 1855, 1737, 533, 3678, 2083, 1389, 1028, 3283, 1043, 1629, 2087, 646, 2727, 195, 2821, 2839, 384, 2692, 1256, 1131, 1421, 2926, 3272, 1231, 2835, 1136, 652, 2219, 3264, 540, 576, 3695, 2701, 91, 1338, 76, 1032, 2529, 3352, 1750, 946, 3535, 1010, 1028, 1031, 182, 3678, 82, 2635, 2030, 77, 544, 532, 1977, 1737, 1756, 3607, 3431, 3510, 709]. Ward [2833, 2507, 711, 972, 415]. Wards [2268]. Ware [1598]. Warfare [2333, 1746, 833, 652, 709, 2854]. Warfield [104]. Waring [1884]. Warm [1144]. Wärme [3715]. Warming [601, 2786]. Warner [2564, 45, 133]. Warp [3634]. Warren [2190, 2421, 2453, 737, 733]. Warriors [1131]. Wars [1068, 1738, 1262, 2339, 186, 2631, 3272, 2996, 914, 2685, 2193, 3137, 848, 3086, 1161]. Wartime [383]. Warwick [1839, 2053]. Was [62, 3500, 1235, 2991, 550, 2776]. Washington [1077, 1618]. Waste [1768, 3770]. Waszkis [1401]. Watch [1951]. Watching [3683]. Water [2134, 1412, 1487, 3001, 383, 3073, 2966, 3488, 1821]. Waterhouse [3166]. Waters [1482, 3529, 2958, 673]. Waterwheel [3631]. Waterworks [1821]. Watkins [219]. Watson [1896, 1941, 2571, 1054, 1339]. Watt [3001, 3050, 3050]. Watts [1618]. Wave [3545, 2652]. Wavelength [2330]. Waves [1411, 2293, 2439, 3173]. Wax [1530]. Way [847, 778, 764, 2658, 2249, 395, 312]. Wayne [51, 394]. Ways [2834, 2895, 1163]. Wayward [2502]. Wealth [28, 589]. Weapons [2839, 2440, 3119]. Wear [2491]. Weart [2786]. Weather [3335, 989, 3061, 2748, 3399, 3309, 3724, 815, 1020, 971]. Weatherall [1937]. Weathering [1261]. Web [1668]. Webb [2036]. Weber [1849, 1727, 2578, 2703, 552, 2205, 610]. Webster [1896, 62]. Weeks [482]. Weft [3634]. wei [3261, 3268]. Weidman [2708, 401]. Weidmann [3736]. Weigel [482]. Weightman [1865]. Weights [2354]. Weikart [2501]. Weil [483]. Weimar [1532, 1934]. Weimin [3200]. Weindling [1026, 2926]. Weiner [884, 1620, 1400]. Weininger [638]. Weinstock [211]. Weintraub [1569]. Weir [1019]. Weire [365]. Weiss [196]. Weitekamp [2508]. weitere [1178, 1324]. Welch [203]. Welfare [1963]. Weliver [1107]. Well [2357, 3770]. Well-Ordered [2357]. Welland [179]. Wellman [1832]. Wells [1122, 3757, 2274]. Welt [3331]. Weltanschauung [2205]. Welten [3564]. Welter [1604]. Welterklärung [3405]. Weltratsel [282]. Welts [562]. wen [3261, 3343, 3726, 2533, 3726]. Wendee [2141]. Wendy [398, 2087, 1860].

Weng [3432, 3432]. Wenhao [3432]. Were [3428]. Werk [2560, 282, 1004, 3605]. Werke [1923, 483]. Werkzeug [1284, 1042]. Werner [1349, 3227, 2904, 3073, 1519]. Wernher [1415]. Wertheim [602]. Wertheimer [2684]. Wesley [803, 3524]. West [2361, 2809, 803, 1224, 2156, 2355, 174, 2979, 1449, 452, 562, 1144, 696, 810, 988, 1147]. Westdeutschland [3436]. Westendorf [867]. Western [1437, 1357, 1968, 2532, 2244, 2400, 2722, 1168, 1564, 3609, 2817, 3219, 1521, 967, 3555, 645, 2251, 215, 438, 1673, 3031, 1472, 961, 2677, 1911]. Westernization [3306]. Westeuropa [347]. Westrum [543]. Westwick [3093, 1954]. weten [2724]. wetenschap [2724]. Wetenschapsbeelden [906]. Wetzell [860]. Weyl [1013, 387]. Whale [3499]. Whaling [2927]. Whatever [3253]. Wheat [82]. Wheel [2247]. Where [2651, 3652]. Whewell [1113]. Which [3231]. Whipple [1196, 3196, 2713]. Whitaker [965]. White [2161, 2162, 308, 185, 3081, 1894, 597, 292]. Whitemen [3764]. Whiteness [2053]. Whitesell [893]. Whitfield [961]. Whitney [2613]. Whittlesey [2078]. Whitworth [1869]. Who [1415, 2681, 731, 2500, 2045, 277, 3255, 380, 3087, 394, 923, 1426, 3078, 2775, 2271, 1005, 2942, 2844, 1115, 1161, 1445, 3548]. Whole [993, 2440, 46]. Wholeness [3620]. Whorton [1883]. Whose [1865, 1389, 2300]. Whyte [1928]. Wickliffe [1870]. Widerstreit [188]. Widmalm [2657]. Widmung [3405]. Wiederaufbau [525]. Wiedervereinigung [2211]. Wielen [20]. Wien [476]. Wiener [3418]. Wiesel [619]. Wiesemann [3679]. Wiesner [2432]. Wightman [75]. Wigner [75, 75]. Wihelm [630]. Wikler [656]. Wilbur [1443, 578]. Wild [562, 1709, 3468]. wilden [500]. Wildenberg [390]. Wilderness [2421]. Wildlife [659, 2929, 729]. Wilds [2437]. Wilhelm [2116, 2117, 3103, 1109, 3515, 707, 769, 2838, 48, 1847, 2763, 2753]. wilhelminischer [1516]. Wilkening [2767]. Wilkes [1935]. Wilkins [2608]. Wilkinson [1948]. Will [3231, 2167, 1045]. Willard [817]. Wille [288]. Willi [2420]. William [436, 658, 3090, 378, 61, 3390, 57, 197, 3323, 889, 2298, 310, 288, 317, 786, 2856, 708, 480, 795, 3665, 2726, 86, 3651, 796, 306, 3587, 662, 2274, 1272, 3147, 343, 389, 508, 2852, 2021, 2985, 2894, 2500, 1578, 959, 2613, 2050, 1020, 2892, 2669, 889, 859, 1797, 2673, 2599, 2374, 2756, 1113, 990, 433, 1079, 3092, 268, 175, 50, 660, 3742, 515, 3318, 3665]. Williams [2939, 44, 2098, 2895, 2275, 2437, 2145, 1131, 1099, 447, 507]. Willmoth [3196]. Wills [930]. Willson [275, 711]. Willstätters [725]. Wilmot [805]. Wilson [519, 1527, 3457, 864, 2848, 489, 372, 123, 1625, 2951, 784, 604, 633, 2023, 2061]. Winchell [2554]. Wind [1565, 3488]. Window [1469]. Wine [1393]. Wini [733]. Winn [488]. Winners [216]. Winnicott [2639]. Winslow [395]. Winter [1837, 801, 3144]. Winther [2292]. Wireless [1864, 2762]. Wirkung [282, 360]. Wirtschaft [1402]. Wisconsin [1531]. Wisdom [855]. Wise [872, 1208]. Wiseman [767]. Wissen [144, 474]. Wissens [474]. Wissenschaft [2117, 1985, 2258, 1402, 2485, 2056, 3761, 3013, 3418, 617, 642, 1985, 3254, 194, 153, 2211]. Wissenschaften [3331, 2997, 3763, 1052].

Wissenschafter [474]. Wissenschaftlichen [3405, 2370, 67, 1505]. wissenschaftlicher [163, 1714]. Wissenschaftsaustausch [347]. Wissenschaftsforschung [3227]. Wissenschaftsgeschichte [3227, 3477, 3761]. wissenschaftshistorische [3715]. Wissenschaftskultur [1934, 3260]. Wissenschaftslabyrinth [360]. Wissenschaftslandschaften [3763]. Wissenschaftslegitimierung [3405]. Wissenschaftsorganisation [3171]. Wissenschaftspopularisierung [337]. Wissenschaftstheorie [704]. Wissensgeschichte [3645]. Wissensproduktion [3402]. Witch [488]. Witelo [1925, 708]. Withers [2912, 1359, 366]. Within [2790]. without [3358, 417, 861, 3039, 1039]. Witkowski [1021, 1806]. Witness [2895]. Witt [2032]. Witte [1494]. Wittenberg [2143]. Wizard [3067, 720]. Wizards [2217]. Wohl [1939]. Wöhler [2841]. Wolf [769, 3035, 587]. Wolf-Dieter [769]. Wolfe [2086, 1142, 2344, 821, 2947, 797]. Wolff [2543, 56]. Wolfgang [2117, 52, 20, 3322, 2538]. Wolfhart [867]. Wolves [3007]. Woman [2498, 562, 731, 558, 468, 1717, 1763, 1451, 2498]. Womb [127]. Women [852, 3099, 2939, 855, 2622, 637, 562, 1896, 1122, 3234, 2667, 892, 1003, 3141, 938, 2172, 1192, 3117, 2243, 1131, 1759, 2778, 278, 3149, 110, 308, 1513, 1130, 1757, 285, 2747, 571, 1431, 2355, 2650, 3650, 2193, 788, 433, 1594, 3582, 3332, 3750, 2508, 2566, 2931, 2201, 920, 189, 733, 2941, 1135, 2247, 1107, 1449, 1796, 1134, 127, 210, 1256, 1224, 59, 114, 3225, 124, 211]. Wonder [2094, 3218]. Wonderful [1450, 3433]. Wonders [2141, 1707, 2978]. Wondrous [3040]. Wood [651, 685, 160, 651]. Woods [1706]. Woodward [3569, 1408, 342]. Wooley [2081]. Woolf [1559, 2283]. Worboys [905]. Word [2726, 2249, 355, 1440]. Words [1116, 2699]. Work [2419, 2911, 1859, 3338, 1747, 2490, 168, 2159, 3255, 520, 315, 2841, 2296, 2641, 2851, 2409, 3414, 2639, 1427, 1924, 3441, 1013]. Worker [628]. Workers [3020, 2694, 2931]. Working [102, 2952, 292, 3554]. Workingmen [376]. Workmen [2319]. Workplace [1047]. Works [1995, 301, 884, 869, 1133, 2253, 995, 3312, 1591, 1872, 580, 75, 1733, 66, 3465, 793, 1133]. Workshop [849, 3641, 1170]. Workshops [845]. World [3344, 2882, 1632, 247, 794, 2681, 2545, 2885, 1349, 644, 1165, 3234, 2081, 1747, 3767, 1737, 533, 519, 1738, 2042, 3079, 3678, 2664, 1389, 1028, 2772, 1043, 646, 3607, 2253, 3655, 2839, 546, 145, 1256, 1131, 155, 1421, 2913, 2030, 3131, 2092, 1245, 357, 1698, 3096, 2996, 3737, 1083, 1351, 3151, 3724, 2830, 1426, 3078, 3146, 1926, 3022, 1611, 1239, 1348, 2685, 3211, 3650, 1356, 1408, 3492, 3556, 437, 2248, 3332, 91, 3064, 2440, 1016, 296, 3407, 3301, 1597, 2989, 3695, 2801, 2333, 73, 247, 513, 2429, 2427, 2666, 29, 2727, 317, 79, 2740, 351, 172, 285, 302]. World [1367, 1001, 3504, 2540, 1101, 1996, 3104, 3243, 2610, 3217, 3733, 1239, 585, 1390]. World-Maps [585]. World-Pictures [2801]. World-Renowned [644]. World-View [1597]. Worldly [1779]. Worlds [2663, 193, 2635, 3000, 2970, 2153, 861, 2814, 763, 1113, 3709]. Worldview [1336]. Worldwide [3083]. Wormbs [2657]. Worrying [3097]. Worse [925]. Worster [803]. Worton [280]. Wosk [2247]. Would [3700]. Wounds [1338, 1221]. Wouter [2817]. Wråkberg [504]. Wrakberg [381]. Wren [158]. Wright [921, 1104, 2787, 3074, 904]. Write [1239]. Writer [195, 3258, 3602]. Writers [2319, 2629]. Writing [1122, 125, 2670, 149, 721, 3744, 2335, 3080, 1567, 883, 1105, 2409, 2360, 1980, 3666]. Writings [884, 2722, 526, 1921, 1055, 3751, 2751, 2213, 921, 158, 155, 3784, 3392, 1106]. Wrong [62, 2508]. Wu [3261, 3343, 3186, 3219]. Wuensch [3605]. Wunderbare [2258]. Würdigung [1826]. Wurttemberg [56]. Wurtz [2073]. Würzburger [1920]. Wurzeln [1042]. Wyder [629]. Wyhe [2203]. Wylie [1018, 1422]. Wyndham [2768]. Wyse [3473].

X [1140, 3092, 2472, 2233]. X-Ray [1140]. X-rays [2233]. Xe [1226, 3388]. Xenotransplantation [319]. XI [2386, 3219]. xian [3561, 3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277]. Xiaochun [31]. Xiaoyuan [3343]. Xie [3262]. xihua [3306]. XII [2022]. XIIe [3482, 1691]. XIIe-XVe [3482, 1691]. XIIème [939]. XIIIe [982, 3566, 614]. XIIIe-XIVe [614]. XIIIe-XVIe [3566]. xin [3268, 3277, 3277]. xing [3262]. Xingsui [3261]. Xiong [3200]. XIV [35]. XIVe [982, 614]. XIX [2943, 449, 3174, 2995]. XIXe [3635, 3639, 2675, 1352, 2278, 2549, 1735]. XIXe-XXe [3639]. Xu [979, 1369]. xue [3275, 3261, 3186, 3343, 3219, 3517, 3200, 3333, 3432, 3268, 3277]. Xuetong [3432]. xv [2472]. XVe [3482, 1691, 2408]. XVI [449, 146, 1696, 691]. XVI-XVIII [691]. XVIe [2738, 3566, 2971, 3310]. XVII [1696, 2472]. XVIIe [3157, 2733, 494, 2738, 3159, 361, 3740, 3494]. XVIII [2262, 625, 691, 3735, 1696]. XVIIIe [3495, 3635, 3152, 3494]. XX [2943, 2022, 2386, 60]. XXe [2091, 3639]. XXème [939]. XXVII [3735].

Yacher [2922]. Yagi [2362]. Yakira [497]. Yakov [1411]. Yale [2170]. Yan [3343, 3275, 3261, 3186, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3277]. yanjiu [307]. Yanni [509]. Yanqiong [3262]. Yaotiao [979]. Yazoo [2569]. Yazoo-Mississippi [2569]. Ye [3646]. Year [1783, 2666, 601, 3076, 306, 1965, 2713]. Years [812, 2472, 595, 375, 205, 3340, 756, 1861, 3410, 2658, 3050, 547, 1011, 371, 2966, 399, 275, 2028, 1387, 2751, 303, 68, 3191, 214, 340, 1424, 3340, 2484, 931]. Yellow [854]. Yellowstone [2078, 1147, 174, 373]. Yeo [2422, 785]. Yerkes [2853]. Yervant [811]. Yerxa [1846]. yi [3261, 3262, 3200, 3268]. Yin [1092]. Yochelson [1012]. Yoder [2923, 815]. Yolanda [200]. Yoneyama [208]. York [2472, 2844, 804, 1942, 550, 3509, 3499, 3414, 3655]. Yoshioka [1406]. Yost [1966]. Young [287, 2206, 2665, 1877, 2769, 400, 991]. Yourgrau [522]. Youschkevich [1291]. Youshkevich [1291]. Youshkevich-K [1291]. yu [3261, 3262, 307, 3333, 3277]. yun [3268]. Yung [870]. Yuri [91]. Yves [2721].

Z [991, 2096, 1872]. Zacut [1493]. Zagorin [622]. zai [3261, 3268, 3277]. Zaitlin [399]. Zakiya [880]. Zakrzewska [3006, 1796]. Zammito [2313]. Zapotec [1048]. Zapovedniks [2506]. Zarafa [505]. Zavella [209]. Zealand [315, 2058]. Zebrowski [440]. Zecaire [875]. Zee [441]. Ze’ev [3410].

Zegher [2544]. Zeichen [474]. Zeichentheorie [474]. Zeit [194, 2837, 1013, 2897, 3331]. Zeitalter [635]. Zeiten [2189]. zeitgenössischen [704, 1964]. Zeitzeugen [1260]. Zell [2846]. Zell-Theorien [2846]. Zellenstaat [3672]. Zeller [2308]. Zellularpathologie [3672]. Zero [1207, 1424, 1908, 684]. Zewail [1875]. Zhang [3517, 3333, 3268, 3277]. zhengming [307]. Zhentao [979]. zhi [3275, 3261, 3200, 3333]. zhizni [1354]. zhong [3268]. Zhongguo [3275, 3261, 3186, 3343, 3262, 3219, 3517, 3200, 3333, 3432, 3268, 3306, 3277, 3561, 3275, 3517, 3306]. zi [3277]. Ziche [3763]. Ziegler [1530]. zielgerichteten [163]. Ziggelaar [42]. Zijin [3343]. Ziman [1164]. Zimbabwe [2908]. Zimmerman [239, 1117, 368]. Zinaida [531]. Zinguer [150]. Zinsser [2747]. Zionism [3264]. Zone [63, 1540]. Zonggang [3186]. Zoo [3175, 1534]. Zoological [3219]. zoológicas [2386]. Zoológico [3056]. Zoologist [1599]. Zoology [724]. Zoos [1351, 1945, 1910]. zoroastriana [2252]. Zuccoli [382]. zufolo [3138]. Zuidervaart [499]. Zukunftsmaschine [2855]. Zum [2103, 3515, 3227, 769, 347, 24, 3477, 2347, 2831, 1386, 274]. zuotanhui [307]. Zupko [1817]. Zur [3177, 3578, 67, 525, 188, 3672, 3486, 2291, 2530, 1178, 1324, 52, 1052, 642, 707, 3227, 2406, 3757, 3657, 199, 2258, 1402, 3035, 474, 2143, 150, 162, 34, 2144, 3134, 2897, 3171, 3013, 3689, 909, 1505, 10]. Zurich [387, 3512, 2855]. Zusammenfassungen [3486]. Zusammenspiel [1386]. zwanzigsten [2276]. zweite [2553]. Zweiten [1017, 171]. zwischen [3103, 500, 3645, 3578, 347, 2347, 170, 1291, 3500, 2291]. Zwischenkriegszeit [1017].

References

Anonymous:2000:FMa

[1] Anonymous. Front matter. Isis, 91(??):1–19, ???? 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/3080726.

Anonymous:2000:CBH

[2] Anonymous. Current bibliography of the history of science and its cultural influences, 2000. Isis, 91(??):20–200, ???? 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/3080727.

Anonymous:2000:BMa

[3] Anonymous. Back matter. Isis, 91(??):??, ???? 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/3080728.

Gordin:2000:IBE

[4] Michael D. Gordin. The importation of being earnest: The early St. Petersburg Academy of Sciences. Isis, 91(1):1–31, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237556.

Blair:2000:MPS

[5] Ann Blair. Mosaic physics and the search for a pious natural philosophy in the late Renaissance. Isis, 91(1):32–58, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237557.

Neushul:2000:BDD

[6] Peter Neushul and Zuoyue Wang. Between the Devil and the deep sea: C. k. tseng, mariculture, and the politics of science in modern China. Isis, 91(1):59–88, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237558.

Jackson:2000:BLP

[7] John P. Jackson, Jr. Blind law and powerless science: The American Jewish Congress, the NAACP, and the scientific case against discrimination, 1945–1950. Isis, 91(1):89–116, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237559.

Findlen:2000:BRB

[8] Paula Findlen. Book review: Servants of Nature: A History of Scientific Institutions, Enterprises, and Sensibilities. Isis, 91(1):117–120, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237560.

Cohen:2000:BRBa

[9] Seymour S. Cohen. Book review: Proteins, Enzymes, Genes: The Interplay of Chemistry and Biology. Isis, 91(1):120–124, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237561.

Steele:2000:BRB

[10] John M. Steele. Book review: Zur Entstehung der babylonischen Mondtheorie: Beobachtung und theoretische Berechnung von Mondphasen. Isis, 91(1):125–126, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237562.

Kohler:2000:BRB

[11] Robert E. Kohler. Book review: Cultural Boundaries of Science: Credibility on the Line by Thomas F. Gieryn. Isis, 91(1):127–128, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237563.

Johns:2000:BRB

[12] Adrian Johns. Book review: Prints and Printmaking: An Introduction to the History and Techniques by Antony Griffiths. Isis, 91(1):128–129, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237564.

Rudavsky:2000:BRB

[13] Shari Rudavsky. Book review: The Culture of the Copy: Striking Likenesses, Unreasonable Facsimiles by Hillel Schwartz. Isis, 91(1):129–130, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237565.

Curry:2000:BRB

[14] Michael Curry. Book review: Maps and History: Constructing Images of the past by Jeremy Black. Isis, 91(1):131, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237566.

Chianese:2000:BRB

[15] Robert Chianese. Book review: Verse and Universe: Poems about Science and Mathematics by Kurt Brown. Isis, 91(1):131–132, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237567.

Gingras:2000:BRB

[16] Yves Gingras. Book review: Des sciences et des techniques: Un debat by Roger Guesnerie; François Hartog. Isis, 91(1):132–133, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237568.

Paul:2000:BRB

[17] John J. Paul. Book review: A Carnival for Science: Essays on Science, Technology, and Development by Shiv Visvanathan. Isis, 91(1):133–134, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237569.

Romero:2000:BRB

[18] Aldemaro Romero. Book review: Del diluvio al megaterio: Los origenes de la paleontologia en España by Francisco Pelayo. Isis, 91(1):134–135, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237570.

Malet:2000:BRBa

[19] Antoni Malet. Book review: Analysis and Synthesis in Mathematics: History and Philosophy by Michael Otte; Marco Panza. Isis, 91(1):135– 136, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237571.

Marche:2000:BRB

[20] Jordan D. Marché, II. Book review: The Message of the Angles: Astrometry from 1798 to 1998 by Peter Brosche; Wolfgang R. Dick; Oliver Schwarz; Roland Wielen. Isis, 91(1):136–137, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237572.

Hutchinson:2000:BRB

[21] Gillian Hutchinson. Book review: The Development of the Rudder: A Technological Tale by Lawrence V. Mott. Isis, 91(1):137–138, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237573.

Tsutsui:2000:BRB

[22] William M. Tsutsui. Book review: Technology and Industrialisation: Historical Case Studies and International Perspectives by Ian Inkster. Isis, 91(1):138–139, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237574.

Rose:2000:BRB

[23] Michael R. Rose. Book review: Metaphysics and the Origin of Species by Michael T. Ghiselin. Isis, 91(1):139–140, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237575.

Laubichler:2000:BRB

[24] Manfred D. Laubichler. Book review: Homo Cerebralis: Der Wandel vom Seelenorgan zum Gehirn by Michael Hagner. Isis, 91(1):140–141, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237576.

Cingolani:2000:BRB

[25] Enrico Cingolani. Book review: Il tesoro della salute: Dall’onnipotenza dei semplici all’atomizzazione del farmaco by Marco Beretta. Isis, 91(1): 141–142, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237577.

Drees:2000:BRB

[26] Willem B. Drees. Book review: Faces in the Clouds: A New Theory of Religion by Stewart Elliott Guthrie. Isis, 91(1):142, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237578.

Steinberg:2000:BRB

[27] Ted Steinberg. Book review: Protestantism, Capitalism, and Nature in America by Mark Stoll. Isis, 91(1):142–143, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237579.

Olson:2000:BRBa

[28] Richard Olson. Book review: A History of the Modern Fact: Problems of Knowledge in the Sciences of Wealth and Society by Mary Poovey. Isis, 91(1):143–144, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237580.

Gossin:2000:BRB

[29] Pamela Gossin. Book review: New Science, New World by Denise Albanese. Isis, 91(1):144, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237581.

Colwell:2000:BRB

[30] C. Colwell. Book review: Michel Foucault: Personal Autonomy and Education by James D. Marshall. Isis, 91(1):145, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237582.

Chen:2000:BRB

[31] Cheng-Yih Chen. Book review: The Chinese Sky during the Han: Constellating Stars and Society by Xiaochun Sun; Jacob Kistemaker. Isis, 91(1):145–147, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237583.

Pendergraft:2000:BRB

[32] Mary L. B. Pendergraft. Book review: Phenomenes by Aratos; Jean Martin. Isis, 91(1):147, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237584.

Reeds:2000:BRB

[33] Karen Meier Reeds. Book review: Les livres de cuisine médiévaux by Bruno Laurioux. Isis, 91(1):147–148, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237585.

Richter-Bernburg:2000:BRB

[34] Lutz Richter-Bernburg. Book review: Die Erganzung Ibn Gulgul’s zur Materia medica des Dioskurides: Arabischer Text nebst kommentierter deutscher Ubersetzung by Albert Dietrich. Isis, 91(1):148–149, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237586.

White:2000:BRB

[35] Kevin White. Book review: La metaura d’Aristotile: Volgarizzamento fiorentino anonimo del XIV secolo by Rita Librandi. Isis, 91(1):149– 150, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237587.

Sidelko:2000:BRB

[36] Paul L. Sidelko. Book review: Roger Bacon and the Sciences: Commemorative Essays by Jeremiah Hackett. Isis, 91(1):150–151, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237588.

Jenkins:2000:BRB

[37] Jane E. Jenkins. Book review: Forbidden Rites: A Necromancer’s Manual of the Fifteenth Century by Richard Kieckhefer. Isis, 91(1):151–152, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237589.

Field:2000:BRBa

[38] J. V. Field. Book review: Le vite de’ matematici by Bernardino Baldi; Elio Nenci. Isis, 91(1):152–153, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237590.

Kelter:2000:BRB

[39] Irving A. Kelter. Book review: El debate cosmologico en 1588: Bruno, Brahe, Rothmann, Ursus, Roslin by Miguel A. Granada. Isis, 91(1):153– 154, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237591.

Hutton:2000:BRB

[40] Sarah Hutton. Book review: The Enthusiastical Concerns of Dr. Henry More: Religious Meaning and the Psychology of Delusion by Daniel Fouke. Isis, 91(1):154–155, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237592.

Wilson:2000:BRBa

[41] Catherine Wilson. Book review: Leibniz and Clarke: A Study of Their Correspondence by Ezio Vailati. Isis, 91(1):155–156, March 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237593.

Meli:2000:BRBa

[42] Domenico Bertoloni Meli. Book review: Dissertations: On the Mechanics of Effervescence and Fermentation and on the Mechanics of the Movement of the Muscles by Johann Bernoulli; Paul Maquet; August Ziggelaar. Isis, 91(1):156–157, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237594.

Tobin:2000:BRB

[43] Beth Fowkes Tobin. Book review: Science in the Service of Empire: Joseph Banks, the British State, and the Uses of Science in the Age of Revolution by John Gascoigne. Isis, 91(1):157–158, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237595.

Barahona:2000:BRB

[44] Ana Barahona. Book review: Buffon: A Life in Natural History by Jacques Roger; L. Pearce Williams; Sarah Lucille Bonnefoi; Las epocas de la naturaleza by George-Louis Leclerc Buffon; Antonio Beltran Mari. Isis, 91(1):158–159, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237596.

Bennett:2000:BRBa

[45] Jim Bennett. Book review: Royal Observatory, Cape of Good Hope, 1820–1831: The Founding of a Colonial Observatory: Incorporating a Biography of Fearon Fallows by Brian Warner. Isis, 91(1):159, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237597.

Walls:2000:BRB

[46] Laura Dassow Walls. Book review: The Emerson Museum: Practical Romanticism and the Pursuit of the Whole by Lee Rust Brown. Isis, 91(1):159–161, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237598.

Moser:2000:BRB

[47] Stephanie Moser. Book review: Scenes in a Library: Reading the Photograph in the Book, 1843–1875 by Carol Armstrong. Isis, 91(1):161–162, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237599.

Lotze:2000:BRB

[48] Siegfried Lotze. Book review: “… der Himmel bewahre Sie vor einer socialistischen Herrschaft!” Briefe von Robert Wilhelm Bunsen an Theophile Jules Pelouze und Henri Victor Regnault aus den Jahren 1841, 1848 und 1851 in der Universitats-bibliothek Marburg by Fritz Krafft. Isis, 91(1):162–163, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237600.

Browne:2000:BRB

[49] Rory Browne. Book review: Animal Advocacy and Englishwomen, 1780– 1900: Patriots, Nation, and Empire by Moira Ferguson. Isis, 91(1):163– 164, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237601.

Goetzmann:2000:BRB

[50] William H. Goetzmann. Book review: William H. Emory: Soldier-Scientist by David L. Norris; James C. Milligan; Odie B. Faulk. Isis, 91(1):164–165, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237602.

Benson:2000:BRBa

[51] Maxine Benson. Book review: The Black Hills Journals of Colonel Richard Irving Dodge by Richard Irving Dodge; Wayne R. Kime. Isis, 91(1):165, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237603.

Beaney:2000:BRB

[52] Michael Beaney. Book review: Frege in Jena: Beitrage zur Spurensicherung by Gottfried Gabriel; Wolfgang Kienzler. Isis, 91(1):165–166, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237604.

Janik:2000:BRB

[53] Allan Janik. Book review: Heinrich hertz: Classical Physicist, Modern Philosopher by Davis Baird; R. I. G. Hughes; Alfred Nordmann. Isis, 91(1):166–167, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237605.

Todd:2000:BRB

[54] Edmund N. Todd. Book review: Elektrizitat in der Geistesgeschichte by Klaus Plitzner. Isis, 91(1):167–168, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237606.

Weikart:2000:BRB

[55] Richard Weikart. Book review: Kunstformen der Natur by Ernst Haeckel; Olaf Breidbach; Irenaus Eibl-Eibesfeldt. Isis, 91(1):168–170, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237607.

Weindling:2000:BRB

[56] Paul Weindling. Book review: Einschneidende Massnahmen: Pockenschutzimpfung und traditionale Gesellschaft im Wurttemberg, des fruhen 19. Jahrhunderts by Eberhard Wolff. Isis, 91(1):170, March 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237608.

Drayton:2000:BRB

[57] Richard Drayton. Book review: Science and Power in Colonial Mauritius by William Kelleher Storey. Isis, 91(1):170–171, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237609.

Cimino:2000:BRB

[58] Guido Cimino. Book review: Sulla fina anatomia degli organi centrali del sistema nervoso by Camillo Golgi; Alberto Oliverio. Isis, 91(1):171– 173, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237610.

McGrath:2000:BRB

[59] Sylvia W. McGrath. Book review: Women Pioneers in Texas Medicine by Elizabeth Silverthorne; Geneva Fulgham. Isis, 91(1):173, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237611.

Moulin:2000:BRB

[60] Anne Marie Moulin. Book review: El regreso de las epidemias: Salud y sociedad en el Peru del siglo XX by Marcos Cueto. Isis, 91(1):173– 174, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237612.

Croce:2000:BRB

[61] Paul Jerome Croce. Book review: Genuine Reality: A Life of William James by Linda Simon. Isis, 91(1):174–175, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237613.

Decker:2000:BRB

[62] Hannah S. Decker. Book review: Back to Freud’s Texts: Making Silent Documents Speak by Ilse Grubrich-Simitis; Philip Slotkin; Why Freud Was Wrong: Sin, Science, and Psychoanalysis by Richard Webster. Isis, 91(1):175–177, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237614.

Schweber:2000:BRB

[63] S. S. Schweber. Book review: The End of Science: Facing the Limits of Knowledge in the Twilight Zone of the Scientific Age by John Horgan. Isis, 91(1):177–179, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237615.

McGrayne:2000:BRB

[64] Sharon Bertsch McGrayne. Book review: Marie Curie and the Science of Radioactivity by Naomi Pasachoff. Isis, 91(1):179–180, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237616.

Muenger:2000:BRB

[65] Elizabeth A. Muenger. Book review: The Magnificent Moisants: Champions of Early Flight by Doris L. Rich. Isis, 91(1):180, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237617.

Dean:2000:BRBa

[66] Dennis R. Dean. Book review: Duchamp in Context: Science and Technology in the “Large Glass” and Related Works by Linda Dalrymple Henderson. Isis, 91(1):180–182, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237618.

Hoppe:2000:BRB

[67] Brigitte Hoppe. Book review: Der Eintritt der Frauen in die Gelehrtenrepublik: Zur Geschlechterfrage in akademischen Selbstverstandnis und in der wissenschaftlichen Praxis am Anfang des 20. Jahrhunderts by Johanna Bleker. Isis, 91(1):182–183, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237619.

Rowe:2000:BRA

[68] David E. Rowe. Book review: Albert Einstein, Robert Schulmann, A. J. Kox, Michel Janssen, Jozsef Illy, Ann M. Hentschel, and Klaus Hentschel: The Collected Papers of Albert Einstein: The Berlin Years: Correspondence, 1914–1918. Isis, 91(1):183–185, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237620.

Elliott:2000:BRB

[69] Clark A. Elliott. Book review: Harvard Observed: An Illustrated History of the University in the Twentieth Century by John T. Bethell. Isis, 91(1):185–186, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237621.

Greenaway:2000:BRB

[70] Frank Greenaway. Book review: History of IUPAC, 1919–1987 by Roger Fennell. Isis, 91(1):186, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237622.

Oberdan:2000:BRB

[71] Thomas Oberdan. Book review: Kurt Godel: Unpublished Philosophical Essays by Kurt Godel; Francisco A. Rodriguez-Consuegra. Isis, 91(1): 186–187, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237623.

Ben-Menahem:2000:BRB

[72] Yemima Ben-Menahem. Book review: Schrodinger’s Philosophy of Quantum Mechanics by Michel Bitbol. Isis, 91(1):187–188, March 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237624.

Barrett:2000:BRBa

[73] Jeffrey Barrett. Book review: Interpreting the Quantum World by Jeffrey Bub. Isis, 91(1):188–189, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237625.

Hiebert:2000:BRL

[74] Erwin N. Hiebert. Book review: Lise Meitner-Max von Laue Briefwechsel, 1938–1948 by Jost Lemmerich. Isis, 91(1):189–190, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237626.

Roman:2000:BRE

[75] Paul Roman. Book review: Eugene Paul Wigner: The Collected Works of Eugene Paul Wigner. Arthur S. Wightman: The Scientific Papers. Volume 3, Part 1: Particles and Fields. The Scientific Papers. Volume 3, Part 2: Foundations of Quantum Mechanics. The Scientific Papers. Volume 4. Part 1: Physical Chemistry The Scientific Papers. Volume 4. Part 2: Solid State Physics. Jagdish Mehra: Historical, Philosophical, and Socio-Political Papers. Volume 6: Philosophical Reflections and Syntheses. Isis, 91(1):190–192, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237627.

Wang:2000:BRBa

[76] Jessica Wang. Book review: Creating the Cold War University: The Transformation of Stanford by Rebecca S. Lowen. Isis, 91(1):192–193, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237628.

Schrecker:2000:BRB

[77] Ellen Schrecker. Book review: American Science in an Age of Anxiety: Scientists, Anticommunism, and the Cold War by Jessica Wang. Isis, 91(1):193–194, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237629.

Olwell:2000:BRB

[78] Russell Olwell. Book review: Bohm-Biederman Correspondence. Volume 1: Creativity in Art and Science by David Bohm; Charles Biederman; Paavo Pylkkanen. Isis, 91(1):194–195, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237630.

Kaiser:2000:BRB

[79] David Kaiser. Book review: A Tale of Two Continents: A Physicist’s Life in a Turbulent World by Abraham Pais. Isis, 91(1):195–196, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237631.

Beyler:2000:BRB

[80] Richard H. Beyler. Book review: Reflections on a Theory of Organisms: Holism in Biology by Walter M. Elsasser. Isis, 91(1):196, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237632.

Smith:2000:BRBa

[81] John Kenly Smith, Jr. Book review: American Plastic: A Cultural History by Jeffrey L. Meikle. Isis, 91(1):196–198, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237633.

Fitzgerald:2000:BRB

[82] Deborah Fitzgerald. Book review: Geopolitics and the Green Revolution: Wheat, Genes, and the Cold War by John H. Perkins. Isis, 91(1):198– 199, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237634.

Gillmor:2000:BRB

[83] C. Stewart Gillmor. Book review: The Earth’s Plasmasphere by J. F. Lemaire; K. I. Gringauz; D. L. Carpenter; V. Bossolo. Isis, 91(1):199– 200, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237635.

Tatarewicz:2000:BRBa

[84] Joseph N. Tatarewicz. Book review: Exploring the Unknown: Selected Documents in the History of the U.S. Civilian Space Program. Volume 2: External Relationships by John M. Logsdon; Dwayne A. Day; W. Henry Lambright; Roger Launius. Isis, 91(1):200–201, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237636.

Cao:2000:BRB

[85] Cong Cao. Book review: Science and Dissent in Post-Mao China: The Politics of Knowledge by H. Lyman Miller. Isis, 91(1):201, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237637.

Perleth:2000:BRB

[86] Matthias Perleth. Book review: The Kiss of Death: Chagas’ Disease in the Americas by Joseph William Bastien. Isis, 91(1):201–202, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237638.

Saetnan:2000:BRB

[87] Ann Rudinow Saetnan. Book review: Disciplining Reproduction: Modernity, American Life Sciences, and “The Problems of Sex.” by Adele E. Clarke. Isis, 91(1):202–203, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237639.

Long:2000:BRBa

[88] Diana E. Long. Book review: Controlling Reproduction: An American History by Andrea Tone. Isis, 91(1):203–204, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237640.

Kidwell:2000:BRBa

[89] Clara Sue Kidwell. Book review: Edward S. Curtis and the North American Indian, Incorporated by Mick Gidley. Isis, 91(1):204–205, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237641.

Fagette:2000:BRB

[90] Paul Fagette. Book review: James A. Ford and the Growth of Americanist Archaeology by Michael J. O’Brien; R. Lee Lyman. Isis, 91(1):

205–206, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237642.

Sirotkina:2000:BRB

[91] Irina Sirotkina. Book review: Soviet and American Psychology during World War II by Albert Gilgen; Carol K. Gilgen; Vera A. Koltsova; Yuri N. Oleinik. Isis, 91(1):206, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237643.

Buhle:2000:BRBa

[92] Mari Jo Buhle. Book review: The Therapeutic Process: Essays and Lectures by Karen Horney; Bernard J. Paris. Isis, 91(1):206–207, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237644.

Neushul:2000:BRB

[93] Peter Neushul. Book review: Forged Consensus: Science, Technology, and Economic Policy in the United States, 1921–1953 by David M. Hart. Isis, 91(1):207–208, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237645.

Rotella:2000:BRB

[94] Elyce J. Rotella. Book review: Two Lucky People: Memoirs by Milton Friedman; Rose D. Friedman. Isis, 91(1):208–209, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237646.

Gunnoe:2000:BRB

[95] Charles D. Gunnoe, Jr. Book review: Die Alchemiebibliothek des Alexander von Bernus in der Badischen Landesbibliothek Karlsruhe: Katalog der Drucke und Handschriften by Annelies Stockinger; Joachim Telle. Isis, 91(1):209–210, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237647.

Hedlin:2000:BRB

[96] Edie Hedlin. Book review: A Guide to the Official Archives of the Natural History Museum, London by John C. Thackray. Isis, 91(1):210–211, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237648.

Maples:2000:BRB

[97] Philip G. Maples. Book review: A Catalogue of Records Retained by Hospice and Related Organizations in the U.K. and Republic of Ireland by Paul Lydon. Isis, 91(1):211, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237649.

Anonymous:2000:Ca

[98] Anonymous. Collections. Isis, 91(1):211–214, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237650.

Gregory:2000:NS

[99] Frederick Gregory, Edith Sylla, Michael H. Shank, and Keith R. Benson. News of the Society. Isis, 91(1):215–225, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237651.

Anonymous:2000:BMb

[100] Anonymous. Back matter. Isis, 91(1):226–232, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237652.

Anonymous:2000:FMb

[101] Anonymous. Front matter. Isis, 91(1):??, March 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237555.

Henninger-Voss:2000:WMN

[102] M. Henninger-Voss. Working machines and noble mechanics: Guidobaldo del Monte and the translation of knowledge. Isis, 91(2):233–259, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236915.

Duffin:2000:BSR

[103] Jacalyn Duffin and Charles R. R. Hayter. Baring the sole: The rise and fall of the shoe-fitting fluoroscope. Isis, 91(2):260–282, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236916.

Livingstone:2000:BBW

[104] David N. Livingstone and Mark A. Noll. B. B. Warfield (1851–1921): A biblical inerrantist as evolutionist. Isis, 91(2):283–304, June 2000.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236917.

Chen:2000:LE

[105] Joseph C. Y. Chen and Christopher Cullen. Letters to the Editor. Isis, 91(2):305–309, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236918.

Everitt:2000:BRB

[106] C. W. F. Everitt and Anna Muza. Book review: Image and Logic: A Material Culture of Microphysics. Isis, 91(2):310–313, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236919.

Golinski:2000:BRB

[107] Jan Golinski. Book review: The Science Studies Reader by Mario Biagioli. Isis, 91(2):314, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236920.

Daston:2000:BRB

[108] Lorraine Daston. Book review: History and the Disciplines: The Reclassification of Knowledge in Early Modern Europe by Donald R. Kelley. Isis, 91(2):314–315, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236921.

Eisenstein:2000:BRB

[109] Elizabeth L. Eisenstein. Book review: The Nature of the Book: Print and Knowledge in the Making by Adrian Johns. Isis, 91(2):316–317, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236922.

Lerman:2000:BRB

[110] Nina E. Lerman. Book review: The Gender of History: Men, Women, and Historical Practice by Bonnie G. Smith. Isis, 91(2):317–318, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236923.

Elkins:2000:BRB

[111] James Elkins. Book review: Picturing Science, Producing Art by Caroline A. Jones; Peter Galison. Isis, 91(2):318–319, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236924.

Wilson:2000:BRBb

[112] Andrew D. Wilson. Book review: The Holistic Inspirations of Physics: The Underground History of Electromagnetic Theory by Val Dusek. Isis, 91(2):319–320, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236925.

Rice:2000:BRB

[113] Richard E. Rice. Book review: America’s Scientific Treasures: A Travel Companion by Paul S. Cohen; Brenda H. Cohen. Isis, 91(2):320–321, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236926.

Miller:2000:BRB

[114] Jane A. Miller. Book review: Women in Chemistry: Their Changing Roles from Alchemical Times to the Mid-Twentieth Century by Marlene Rayner-Canham; Geoffrey Rayner-Canham. Isis, 91(2):321–322, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236927.

Shteir:2000:BRB

[115] Ann B. Shteir. Book review: Perspectives on Garden Histories by Michel Conan. Isis, 91(2):322–323, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236928.

Rydell:2000:BRB

[116] Robert W. Rydell. Book review: The Last Dinosaur Book: The Life and Times of a Cultural Icon by W. J. T. Mitchell. Isis, 91(2):323, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236929.

Darnell:2000:BRB

[117] Regna Darnell. Book review: The Land of Prehistory: A Critical History of American Archaeology by Alice Beck Kehoe. Isis, 91(2):323–324, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236930.

Kennedy:2000:BRB

[118] Kenneth A. R. Kennedy. Book review: Ancestral Images: The Iconography of Human Origins by Stephanie Moser. Isis, 91(2):324–326, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236931.

Cheater:2000:BRB

[119] Christine Cheater. Book review: Among the Anthropologists: History and Context in Anthropology by Adam Kuper. Isis, 91(2):326–327, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236932.

Roll-Hansen:2000:BRB

[120] Nils Roll-Hansen. Book review: Finns in the Shadow of the “Aryans”: Race Theories and Racism by Aira Kemilainen. Isis, 91(2):327–328, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236933.

Tytler:2000:BRB

[121] Graeme Tytler. Book review: Der exzentrische Blick: Gesprach über Physiognomik by Claudia Schmolders. Isis, 91(2):328–329, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236934.

Warner:2000:BRB

[122] Deborah Warner. Book review: Measuring the Universe: Our Historic Quest to Chart the Horizons of Space and Time by Kitty Ferguson. Isis, 91(2):329, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236935.

Jaki:2000:BRB

[123] Stanley L. Jaki. Book review: Planetary Astronomy from the Renaissance to the Rise of Astrophysics. Part B: The Eighteenth and Nineteenth Centuries by Rene Taton; Curtis Wilson. Isis, 91(2):329–331, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236936.

Verbrugge:2000:BRB

[124] Martha H. Verbrugge. Book review: Women and Health in America: Historical Readings by Judith Walzer Leavitt. Isis, 91(2):331–332, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236937.

Fee:2000:BRB

[125] Elizabeth Fee. Book review: How the Idea of Profession Changed the Writing of Medical History by John C. Burnham. Isis, 91(2):332–333, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236938.

Sauerteig:2000:BRB

[126] Lutz D. H. Sauerteig. Book review: Eine neue “Kultur des Gebarens”: Die Medikalisierung von Geburt im 18. und 19. Jahrhundert in Deutschland by Hans-Christoph Seidel; Andere Umstande: Eine Kulturgeschichte der Geburt by Eva Labouvie. Isis, 91(2):333–334, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236939.

King:2000:BRB

[127] Helen King. Book review: The Wandering Womb: A Cultural History of Outrageous Beliefs about Women by Lana Thompson. Isis, 91(2):334– 335, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236940.

Howell:2000:BRB

[128] Joel D. Howell. Book review: An Alternative Path: The Making and Remaking of Hahnemann Medical College and Hospital of Philadelphia by Naomi Rogers. Isis, 91(2):335, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236941.

Vernia:2000:BRB

[129] Pedro Vernia. Book review: Historia da farmacia by Joao Rui Pita. Isis, 91(2):335–336, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236942.

Runte:2000:BRB

[130] Alfred Runte. Book review: The Presidio: From Army Post to National Park by Lisa M. Benton. Isis, 91(2):336, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236943.

Constant:2000:BRB

[131] Ed Constant. Book review: Technology and Science in the Industrializing Nations, 1500–1914 by Eric Dorn Brose. Isis, 91(2):336–337, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236944.

Stanbury:2000:BRB

[132] Peter Stanbury. Book review: Colonial Technology: Science and the Transfer of Innovation to Australia by Jan Todd. Isis, 91(2):337, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236945.

Field:2000:BRBb

[133] J. V. Field. Book review: Instruments of Science: An Historical Encyclopedia by Robert Bud; Deborah Jean Warner. Isis, 91(2):338, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236946.

Fleck:2000:BRB

[134] George Fleck. Book review: Natural Dyestuffs and Industrial Culture in Europe, 1750–1880 by Robert Fox; Agusti Nieto-Galan. Isis, 91(2):338– 339, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236947.

Tuana:2000:BRB

[135] Nancy Tuana. Book review: Manifesto of a Passionate Moderate: Unfashionable Essays by Susan Haack. Isis, 91(2):339–340, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236948.

Kalthoff:2000:BRB

[136] Mark A. Kalthoff. Book review: Origins: Linking Science and Scripture by Ariel A. Roth. Isis, 91(2):340–341, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236949.

Porter:2000:BRBa

[137] Theodore M. Porter. Book review: The Creative Power of Chance by Remy Lestienne; E. C. Neher. Isis, 91(2):341, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236950.

Fuller:2000:BRB

[138] Steve Fuller. Book review: Pandora’s Hope: Essays on the Reality of Science Studies by Bruno Latour. Isis, 91(2):341–342, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236951.

Marvin:2000:BRB

[139] Ursula B. Marvin. Book review: The Search for Extraterrestrial Life: Essays on Science and Technology by P. Day. Isis, 91(2):342–343, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236952.

Closs:2000:BRB

[140] Michael Closs. Book review: Cycles of the Sun, Mysteries of the Moon: The Calendar in Mesoamerican Civilization by Vincent H. Malmstrom. Isis, 91(2):343–344, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236953.

Hankinson:2000:BRB

[141] R. J. Hankinson. Book review: Sobre la localizacion de las enfermedades: De locis affectis by Galen; Salud Andres Aparicio. Isis, 91(2):344–345, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236954.

Tattersall:2000:BRB

[142] Ian Tattersall. Book review: L’homme des origines: Savoirs et fictions en prehistoire by Claudine Cohen. Isis, 91(2):345–346, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236955.

Perreiah:2000:BRB

[143] Alan R. Perreiah. Book review: The Physics of Duns Scotus: The Scientific Context of a Theological Vision by Richard Cross. Isis, 91(2): 346–347, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236956.

Cross:2000:BRB

[144] Richard Cross. Book review: Kontingenz und Wissen: Die Lehre von den futura contingentia bei Johannes Duns Scotus by Joachim Ronald Soder. Isis, 91(2):347, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236957.

Jervis:2000:BRB

[145] Jane L. Jervis. Book review: The Age of Two-Faced Janus: The Comets of 1577 and 1618 and the Decline of the Aristotelian World View in the Netherlands by Tabitta Van Nouhuys. Isis, 91(2):347–348, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236958.

Eastwood:2000:BRB

[146] Bruce S. Eastwood. Book review: Matematicas, cosmologia, y humanismo en la España del siglo XVI: Los Comentarios al segundo libro de la Historia natural de Plinio de Jeronimo Munoz by Victor Navarro Brotons; Enrique Rodriguez Galdeano. Isis, 91(2):348–349, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236959.

Grant:2000:BRB

[147] Edward Grant. Book review: Johannes de Muris, De arte mensurandi: A Geometrical Handbook of the Fourteenth Century by Johannes de Muris; H. L. L. Busard. Isis, 91(2):349–350, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236960.

Osler:2000:BRB

[148] Margaret J. Osler. Book review: Reading the Book of Nature: The Other Side of the Scientific Revolution by Allen G. Debus; Michael T. Walton. Isis, 91(2):350–351, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236961.

Konvitz:2000:BRB

[149] Josef W. Konvitz. Book review: The Self-Made Map: Cartographic Writing in Early Modern France by Tom Conley. Isis, 91(2):351–352, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236962.

Priesner:2000:BRB

[150] Claus Priesner and Ulrich Neumann. Book review: Paracelsus und seine internationale Rezeption in der fruhen Neuzeit: Beitrage zur Geschichte des Paracelsismus by Heinz Schott; Ilana Zinguer. Isis, 91(2):352–353, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236963.

French:2000:BRB

[151] Roger French. Book review: Medicine in the English Middle Ages by Faye Getz. Isis, 91(2):353–354, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236964.

Musacchio:2000:BRB

[152] Jacqueline Marie Musacchio. Book review: How to Do It: Guides to Good Living for Renaissance Italians by Rudolph M. Bell. Isis, 91(2): 354–355, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236965.

Emerton:2000:BRB

[153] Norma E. Emerton. Book review: Wissenschaft und Mystik bei J. B. van Helmont (1579–1644) by Berthold Heinecke. Isis, 91(2):355, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236966.

Sepper:2000:BRB

[154] Dennis L. Sepper. Book review: Descartes’ Meditations: Background Source Materials by Roger Ariew; John Cottingham; Tom Sorell. Isis, 91(2):355–356, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236967.

Klein:2000:BRB

[155] Julie R. Klein. Book review: The World and Other Writings by Rene Descartes; Stephen Gaukroger. Isis, 91(2):356–357, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236968.

Belgioioso:2000:BRB

[156] Giulia Belgioioso. Book review: Letture cartesiane by Alberto Pala. Isis, 91(2):357–358, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236969.

Dear:2000:BRB

[157] Peter Dear. Book review: Descartes and the Last Scholastics by Roger Ariew. Isis, 91(2):358–359, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236970.

Bennett:2000:BRBb

[158] Jim Bennett. Book review: Wren’s “Tracts” on Architecture and Other Writings by Lydia M. Soo. Isis, 91(2):359–360, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236971.

Turner:2000:BRB

[159] G. L’E. Turner. Book review: Patrons, Artisans, and Instruments of Science, 1600–1750 by Silvio A. Bedini. Isis, 91(2):360–361, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236972.

Carson:2000:BRB

[160] Emily Carson. Book review: Critique of Pure Reason by Immanuel Kant; Paul Guyer; Allen W. Wood. Isis, 91(2):361–362, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236973.

Bork:2000:BRB

[161] Kennard B. Bork. Book review: Les plis du temps: Mythe, science et H.-B. de Saussure by Albert V. Carozzi; Bernard Crettaz; David Ripoll. Isis, 91(2):362–363, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236974.

Ranea:2000:BRBa

[162] Alberto Guillermo Ranea. Book review: Kontinuitat und Mechanismus: Eine Studie zur Philosophie des jungen Leibniz in ihrem ideengeschichtlichen Kontext by Philip Beeley. Isis, 91(2):363–364, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236975.

Wagenitz:2000:BRB

[163] Gerhard Wagenitz. Book review: Qualitates vegetabilium, Vires medicamentorum, und Oeconomicus usus plantarum bei Carl von Linne (1707– 1778): Erste Versuche einer zielgerichteten Forschung nach Arznei- und Nutzpflanzen auf wissenschaftlicher Grundlage by Gerlinde Hovel. Isis, 91(2):364–365, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236976.

Cooper:2000:BRB

[164] Alix Cooper. Book review: Medizin im Brief: Der Stadtische Arzt des 18. Jahrhunderts im Spiegel seiner Korrespondenz by Thomas Schnalke. Isis, 91(2):365–366, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236977.

deRomo:2000:BRB

[165] Ana Cecilia Rodriguez de Romo. Book review: La introduccion de la clinica en Valencia: Felix Miquel y Mico, 1754–1824 by Jorge Perez Navarro. Isis, 91(2):366, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236978.

Kevles:2000:BRB

[166] Bettyann Holtzmann Kevles. Book review: Making the Body Beautiful: A Cultural History of Aesthetic Surgery by Sander L. Gilman. Isis, 91(2):

366–367, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236979.

Baldwin:2000:BRB

[167] Martha Baldwin. Book review: The Feejee Mermaid: And Other Essays in Natural and Unnatural History by Jan Bondeson; The Platypus and the Mermaid and Other Figments of the Classifying Imagination by Harriet Ritvo. Isis, 91(2):367–369, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236980.

Feffer:2000:BRB

[168] Loren Butler Feffer. Book review: James Joseph Sylvester: Life and Work in Letters by Karen Hunger Parshall. Isis, 91(2):369–370, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236981.

Hunt:2000:BRBa

[169] Bruce J. Hunt. Book review: The Natural Philosophy of James Clerk Maxwell by P. M. Harman. Isis, 91(2):370–371, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236982.

Mulligan:2000:BRB

[170] Joseph F. Mulligan. Book review: Brieftagebuch zwischen Max Planck, Carl Runge, Bernhard Karsten, und Adolf Leopold by Max Planck; Carl Runge; Bernhard Karsten; Adolf Leopold; Klaus Hentschel; Renate Tobies. Isis, 91(2):371–372, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236983.

Rocke:2000:BRBa

[171] Alan J. Rocke. Book review: Chemischer Atomismus: Anwendung, Veranderung, Alternativen im deutschsprachigen Raum in der zweiten Halfte des 19. Jahrhunderts by Britta Gors. Isis, 91(2):372–373, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236984.

Mayr:2000:BRB

[172] Ernst Mayr. Book review: Darwin’s Spectre: Evolutionary Biology in the Modern World by Michael R. Rose. Isis, 91(2):373–374, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236985.

Ghiselin:2000:BRB

[173] Michael T. Ghiselin. Book review: Mystery of Mysteries: Is Evolution a Social Construction? by Michael Ruse. Isis, 91(2):374–375, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236986.

Pritchard:2000:BRBa

[174] James Pritchard. Book review: Yellowstone and the Great West: Journals, Letters, and Images from the 1871 Hayden Expedition by Marlene Deahl Merrill. Isis, 91(2):375–376, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236987.

Benson:2000:BRBb

[175] Maxine Benson. Book review: William Henry Jackson: Farming the Frontier by Douglas Waitley. Isis, 91(2):376, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236988.

Marchand:2000:BRB

[176] Suzanne Marchand. Book review: Finding the Walls of Troy: Frank Calvert and Heinrich Schliemann at Hissarlik by Susan Heuck Allen; Naissance de l’archeologie moderne, 1798–1945 by Eve Gran-Aymerich. Isis, 91(2):376–379, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236989.

Roldan:2000:BRB

[177] Arturo Alvarez Roldan. Book review: Cambridge and the Torres Strait: Centenary Essays on the 1898 Anthropological Expedition by Anita Herle; Sandra Rouse. Isis, 91(2):379–380, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236990.

Macdonald:2000:BRB

[178] Sharon Macdonald. Book review: Museums and American Intellectual Life, 1876–1926 by Steven Conn. Isis, 91(2):380–381, June 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236991.

Chrimes:2000:BRB

[179] Mike Chrimes. Book review: The Welland Canals and Their Communities: Engineering, Industrial, and Urban Transformation by John N.

Jackson. Isis, 91(2):381, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236992.

Post:2000:BRB

[180] Robert C. Post. Book review: The Philosophy of Railways: The Transcontinental Railway Idea in British North America by A. A. den Otter. Isis, 91(2):381–382, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236993.

Connor:2000:BRB

[181] J. T. H. Connor. Book review: Kindly Medicine: Physio-Medicalism in America, 1836–1911 by John S. Haller, Jr. Isis, 91(2):382–383, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236994.

Blustein:2000:BRB

[182] Bonnie Ellen Blustein. Book review: Gangrene and Glory: Medical Care during the American Civil War by Frank R. Freemon. Isis, 91(2):383– 385, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236995.

Ackerman:2000:BRB

[183] Evelyn Bernette Ackerman. Book review: Medecine et colonisation: L’aventure indochinoise, 1860–1939 by Laurence Monnais-Rousselot. Isis, 91(2):385, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236996.

Bynum:2000:BRB

[184] W. F. Bynum. Book review: Public Health and Social Justice in the Age of Chadwick: Britain, 1800–1854 by Christopher Hamlin. Isis, 91(2): 385–386, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236997.

Milar:2000:BRB

[185] Katharine S. Milar. Book review: An Illustrated History of American Psychology by John A. Popplestone; Marion White McPherson. Isis, 91(2):386–387, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236998.

Hale:2000:BRB

[186] Nathan G. Hale, Jr. Book review: Dispatches from the Freud Wars: Psychoanalysis and Its Passions by John Forrester; Freud among the Philosophers: The Psychoanalytic Unconscious and Its Philosophical Critics by Donald Levy; Questions for Freud: The Secret History of Psychoanalysis by Nicholas Rand; Maria Torok; Heresy: Sandor Rado and the Psychoanalytic Movement by Paul Roazen; Bluma Swerdloff. Isis, 91(2):387–389, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236999.

Marks:2000:BRB

[187] Lawrence E. Marks. Book review: Bright Colors, Falsely Seen: Synaesthesia and the Search for Transcendental Knowledge by Kevin T. Dann. Isis, 91(2):389–390, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237000.

Marx:2000:BRB

[188] Otto M. Marx. Book review: Psychiatrie im Widerstreit der Konzepte: Zur Entstehungsgeschichte der Tubinger Nervenklinik by Ioannis Pilavas. Isis, 91(2):390, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237001.

Watkins:2000:BRB

[189] Sallie A. Watkins. Book review: Nobel Prize Women in Science: Their Lives, Struggles, and Momentous Discoveries by Sharon Bertsch McGrayne. Isis, 91(2):390–391, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237002.

deLorenzo:2000:BRB

[190] Javier de Lorenzo. Book review: Alta tension: Historia, filosofia, y sociologia de la ciencia: Ensayos en memoria de Thomas Kuhn by Carlos Solis Santos. Isis, 91(2):391–392, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237003.

Berendzen:2000:BRB

[191] Richard Berendzen. Book review: The American Astronomical Society’s First Century by David H. DeVorkin. Isis, 91(2):392–394, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237004.

Malley:2000:BRB

[192] Marjorie Malley. Book review: Flash of the Cathode Rays: A History of J. J. Thomson’s Electron by Per F. Dahl; J. J. Thomson and the Discovery of the Electron by E. A. Davis; I. J. Falconer; Electron: A Centenary Volume by Michael Springford. Isis, 91(2):394–395, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237005.

Crowe:2000:BRB

[193] Michael J. Crowe. Book review: Life on Other Worlds: The Twentieth– Century Extraterrestrial Life Debate by Steven J. Dick; UFOs and Alien Contact: Two Centuries of Mystery by Robert E. Bartholomew; George S. Howard; Alien Life: The Search for Extraterrestrials and Beyond by Barry Parker. Isis, 91(2):395–397, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237006.

Deichmann:2000:BRB

[194] Ute Deichmann. Book review: Wissenschaft in turbulenter Zeit: Erinnerungen eines Chemikers an die Technische Hochschule Munchen 1933– 1953 by Helmut Behrens. Isis, 91(2):397–398, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237007.

Herzog:2000:BRB

[195] Hillary Hope Herzog. Book review: Heinrich Schirmbeck and the Two Cultures: A Post-War German Writer’s Approach to Science and Literature by Cynthia L. Appl. Isis, 91(2):398, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237008.

Morange:2000:BRB

[196] Michel Morange. Book review: Of Flies, Mice, and Men by François Jacob; Giselle Weiss. Isis, 91(2):398–399, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237009.

Farley:2000:BRB

[197] John Farley. Book review: Felix d’Herelle and the Origins of Molecular Biology by William C. Summers. Isis, 91(2):399–400, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237010.

deChadarevian:2000:BRB

[198] Soraya de Chadarevian. Book review: Picture Control: The Electron Microscope and the Transformation of Biology in America, 1940– 1960 by Nicolas Rasmussen. Isis, 91(2):400–401, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237011.

Harwood:2000:BRB

[199] Jonathan Harwood. Book review: Die Entstehung der Synthetischen Theorie: Beitrage zur Geschichte der Evolutionsbiologie in Deutschland 1930–1950 by Thomas Junker; Eve-Marie Engels. Isis, 91(2):401–402, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237012.

McCook:2000:BRB

[200] Stuart McCook. Book review: La modernizacion dificil: Henri Pittier en Venezuela, 1920–1950 by Yolanda Texera Arnal. Isis, 91(2):402– 403, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237013.

Lowy:2000:BRBa

[201] Ilana Lowy. Book review: Governing Molecules: The Discursive Politics of Genetic Engineering in Europe and the United States by Herbert Gottweis. Isis, 91(2):403–404, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237014.

Bunzl:2000:BRB

[202] Matti Bunzl. Book review: Culture: The Anthropologists’ Account by Adam Kuper. Isis, 91(2):404–405, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237015.

Kidwell:2000:BRBb

[203] Peggy Aldrich Kidwell. Book review: Makin’ Numbers: Howard Aiken and the Computer by I. Bernard Cohen; Gregory W. Welch; Robert V. D. Campbell; Howard Aiken: Portrait of a Computer Pioneer by I. Bernard Cohen. Isis, 91(2):405–406, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237016.

Norberg:2000:BRB

[204] Arthur L. Norberg. Book review: HAL’s Legacy: 2001’s Computer as Dream and Reality by David G. Stork; Arthur C. Clarke. Isis, 91(2): 406–407, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237017.

Ceruzzi:2000:BRB

[205] Paul Ceruzzi. Book review: Architects of the Information Society: Thirty-Five Years of the Laboratory for Computer Science at MIT by Simson L. Garfinkel. Isis, 91(2):407–409, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237018.

Smith:2000:BRBb

[206] John K. Smith, Jr. Book review: Paths of Innovation: Technological Change in Twentieth–Century America by David C. Mowery; Nathan Rosenberg. Isis, 91(2):409, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237019.

Coopersmith:2000:BRB

[207] Jonathan Coopersmith. Book review: What Have We Learned about Science and Technology from the Russian Experience? by Loren R. Graham. Isis, 91(2):409–410, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237020.

Lindee:2000:BRBa

[208] M. Susan Lindee. Book review: Hiroshima Traces: Time, Space, and the Dialectics of Memory by Lisa Yoneyama. Isis, 91(2):410–411, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237021.

Wajcman:2000:BRB

[209] Judy Wajcman. Book review: Situated Lives: Gender and Culture in Everyday Life by Louise Lamphere; Helena Ragone; Patricia Zavella. Isis, 91(2):411–412, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237022.

Brown:2000:BRBa

[210] Janet Welsh Brown. Book review: Women’s Science: Learning and Succeeding from the Margins by Margaret A. Eisenhart; Elizabeth Finkel; Linda Behm; Nancy Laurence; Karen Tonso. Isis, 91(2):412–413, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237023.

Yochelson:2000:BRB

[211] Ellis L. Yochelson. Book review: Women in the Antarctic by Esther D. Rothblum; Jacqueline S. Weinstock; Jessica F. Morris. Isis, 91(2):413– 414, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237024.

Weller:2000:BRB

[212] Gunter Weller. Book review: To the Pole: The Diary and Notebook of Richard E. Byrd, 1925–1927 by Richard E. Byrd; Raimund E. Goerler. Isis, 91(2):414–415, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237025.

Hansen:2000:BRB

[213] James Hansen. Book review: The Pioneers of Flight: A Documentary History by Phil Scott. Isis, 91(2):415, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237026.

Tatarewicz:2000:BRBb

[214] Joseph N. Tatarewicz. Book review: Flights of Discovery: Fifty Years at the NASA Dryden Flight Research Center by Lane E. Wallace. Isis, 91(2):415–416, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237027.

Smith:2000:BRBc

[215] George David Smith. Book review: Manufacturing the Future: A History of Western Electric by Stephen B. Adams; Orville R. Butler. Isis, 91(2): 416–417, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237028.

Bilstein:2000:BRB

[216] Roger E. Bilstein. Book review: From Engineering Science to Big Science: The NACA and NASA Collier Trophy Research Project Winners by Pamela E. Mack. Isis, 91(2):417–418, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237029.

Ziporyn:2000:BRB

[217] Terra Ziporyn. Book review: How Scientists Explain Disease by Paul Thagard. Isis, 91(2):418–419, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237030.

Iliffe:2000:BRB

[218] John Iliffe. Book review: Histories of Sexually Transmitted Diseases and HIV/AIDS in Sub-Saharan Africa by Philip W. Setel; Milton Lewis; Maryinez Lyons. Isis, 91(2):419–420, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237031.

Hartouni:2000:BRB

[219] Valerie Hartouni. Book review: On the Pill: A Social History of Oral Contraceptives, 1950–1970 by Elizabeth Siegel Watkins. Isis, 91(2):420– 421, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237032.

Lindee:2000:BRBb

[220] M. Susan Lindee. Book review: Acres of Skin: Human Experiments at Holmesburg Prison by Allen M. Hornblum. Isis, 91(2):421–422, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237033.

Brown:2000:BRBb

[221] Theodore M. Brown. Book review: “The Battle for Health”: A Political History of the Socialist Medical Association, 1930–1951 by John Stewart. Isis, 91(2):422–423, June 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237034.

Epstein:2000:BRB

[222] Steven Epstein. Book review: RePlacing Citizenship: AIDS Activism and Radical Democracy by Michael P. Brown. Isis, 91(2):423–424, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237035.

Casas:2000:BRB

[223] Martha Casas. Book review: Control: A History of Behavioral Psychology by John A. Mills. Isis, 91(2):424, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237036.

Pratten:2000:BRB

[224] Stephen Pratten. Book review: The Evolving Rationality of Rational Expectations: An Assessment of Thomas Sargent’s Achievements by Esther-Mirjam Sent. Isis, 91(2):424–425, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237037.

Green:2000:BRB

[225] Philip Green. Book review: Beyond Right and Left: Democratic Elitism in Mosca and Gramsci by Maurice Finocchiaro. Isis, 91(2):425–426, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237038.

Grosholz:2000:BRB

[226] Emily R. Grosholz. Book review: Le ragioni della logica by Carlo Cellucci. Isis, 91(2):426–427, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237039.

Meijsing:2000:BRB

[227] Monica Meijsing. Book review: Philosophy and Memory Traces: Descartes to Connectionism by John Sutton. Isis, 91(2):427–428, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237040.

MacKinnon:2000:BRB

[228] Edward MacKinnon. Book review: Conceptual Foundations of Quantum Mechanics by Bernard d’Espagnat. Isis, 91(2):428–429, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237041.

Lofts:2000:BRB

[229] Steve G. Lofts. Book review: Ernst Cassirer: Scientific Knowledge and the Concept of Man by Seymour W. Itzkoff. Isis, 91(2):429–430, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237042.

Unguru:2000:BRB

[230] Sabetai Unguru. Book review: Abstraction and Representation: Essays on the Cultural Evolution of Thinking by Peter Damerow; Renate Hanauer. Isis, 91(2):430–431, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237043.

Stanley:2000:BRB

[231] Autumn Stanley. Book review: Biographical Dictionary of the History of Technology by Lance Day; Ian McNeil. Isis, 91(2):431–432, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237044.

Anonymous:2000:Cb

[232] Anonymous. Collections. Isis, 91(2):432–434, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237045.

Anonymous:2000:BMc

[233] Anonymous. Back matter. Isis, 91(2):435–442, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237046.

Anonymous:2000:FMc

[234] Anonymous. Front matter. Isis, 91(2):??, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236914.

Luthy:2000:FDS

[235] Christoph Luthy. The fourfold Democritus on the stage of early modern science. Isis, 91(3):443–479, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237904.

Portolano:2000:JQA

[236] Marlana Portolano. John Quincy Adams’s rhetorical crusade for astronomy. Isis, 91(3):480–503, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237905.

Jones:2000:VCV

[237] David S. Jones. Visions of a cure: Visualization, clinical trials, and controversies in cardiac therapeutics, 1968–1998. Isis, 91(3):504–541, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237906.

Courtenay:2000:ECN

[238] William J. Courtenay. The early career of Nicole oresme. Isis, 91(3): 542–548, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237907.

Burnham:2000:EJZ

[239] John C. Burnham. Eloge: June Zimmerman Fullmer, 12 December 1920–31 January 2000. Isis, 91(3):549–551, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237908.

Rocke:2000:EAJ

[240] Alan J. Rocke. Eloge: Aaron J. Ihde, 31 December 1909–23 February 2000. Isis, 91(3):551–553, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237909.

Hatch:2000:LE

[241] Robert A. Hatch and G. A. Russell. Letters to the Editor. Isis, 91(3): 554–560, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237910.

Barrow-Green:2000:BRB

[242] June Barrow-Green. Book review: The History of Counting by Denise Schmandt-Besserat. Isis, 91(3):561, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237911.

Aveni:2000:BRB

[243] Anthony F. Aveni. Book review: Mapping Time: The Calendar and Its History by E. G. Richards. Isis, 91(3):561–562, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237912.

Dick:2000:BRB

[244] Steven J. Dick. Book review: Planetary Astronomy: From Ancient Times to the Third Millennium by Ronald A. Schorn. Isis, 91(3):562–563, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237913.

Sonne:2000:BRB

[245] Birgitte Sonne. Book review: The Arctic Sky: Inuit Astronomy, Star Lore, and Legend by John MacDonald. Isis, 91(3):563, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237914.

Sur:2000:BRB

[246] Abha Sur. Book review: Another Reason: Science and the Imagination of Modern India by Gyan Prakash. Isis, 91(3):563–565, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237915.

Barrow:2000:BRB

[247] Mark V. Barrow, Jr. Book review: Drawn from Life: Science and Art in the Portrayal of the New World by Victoria Dickenson; Empire’s Nature: Mark Catesby’s New World Vision by Amy R. W. Meyers; Margaret Beck Pritchard. Isis, 91(3):565–567, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237916.

Monmonier:2000:BRB

[248] Mark Monmonier. Book review: Historical Perspectives on Climate Change by James Rodger Fleming. Isis, 91(3):567–568, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237917.

Proctor:2000:BRB

[249] Robert N. Proctor. Book review: The Nature of Diamonds by George E. Harlow. Isis, 91(3):568–569, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237918.

Noble:2000:BRB

[250] David W. Noble. Book review: Meaning in Technology by Arnold Pacey. Isis, 91(3):569, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237919.

Kline:2000:BRB

[251] Ronald R. Kline. Book review: A Social History of American Technology by Ruth Schwartz Cowan. Isis, 91(3):569–570, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237920.

Willis:2000:BRB

[252] Carol Willis. Book review: Building Lives: Constructing Rites and Passages by Neil Harris. Isis, 91(3):570–572, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237921.

Hunter:2000:BRB

[253] Dianne Hunter. Book review: The Knotted Subject: Hysteria and Its Discontents by Elisabeth Bronfen. Isis, 91(3):572–573, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237922.

Caton:2000:BRB

[254] Donald Caton. Book review: Reading Birth and Death: A History of Obstetric Thinking by Jo Murphy-Lawless. Isis, 91(3):573–574, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237923.

Cueto:2000:BRB

[255] Marcos Cueto. Book review: De la responsabilidad individual a la culpabilizacion de la victima: El papel del paciente en la prevencion de la enfermedad by Luis Montiel; Isabel Porras. Isis, 91(3):574–575, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237924.

Gilman:2000:BRB

[256] Sander L. Gilman. Book review: Die Herkunft des Schonen: Grundzuge der evolutionaren Asthetik by Klaus Richter. Isis, 91(3):575–576, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237925.

Fournier:2000:BRB

[257] Marian Fournier. Book review: The Birth of the Cell by Henry Harris. Isis, 91(3):576–577, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237926.

Severino:2000:BRB

[258] Sally K. Severino. Book review: Care of the Psyche: A History of Psychological Healing by Stanley W. Jackson. Isis, 91(3):577, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237927.

Montgomery:2000:BRBa

[259] Scott L. Montgomery. Book review: The Advancement of Science, and Its Burdens by Gerald Holton; The Scientific Imagination by Gerald Holton. Isis, 91(3):577–578, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237928.

Reppy:2000:BRB

[260] Judith Reppy. Book review: The Economic Laws of Scientific Research by Terence Kealey. Isis, 91(3):578–579, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237929.

Schmandt-Besserat:2000:BRB

[261] Denise Schmandt-Besserat. Book review: The Lost Civilizations of the Stone Age by Richard Rudgley. Isis, 91(3):579–580, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237930.

vanDalen:2000:BRB

[262] Benno van Dalen. Book review: The History and Practice of Ancient Astronomy by James Evans. Isis, 91(3):580–581, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237931.

Hicks:2000:BRB

[263] Ronald Hicks. Book review: Stonehenge Revealed by David Souden. Isis, 91(3):581–582, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237932.

Savage-Smith:2000:BRB

[264] Emilie Savage-Smith. Book review: Naturwissenschaft bei den Arabern im 10. Jahrhundert n. Chr.: Briefe des Abu l-Fadl Ibn al-Amid (gest. 360/970) an Adudaddaula by Hans Daiber; Science in Medieval Islam: An Illustrated Introduction by Howard R. Turner. Isis, 91(3):582–585, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237933.

Moreton:2000:BRB

[265] Jennifer Moreton. Book review: The Kalendarium of John Somer by Linne R. Mooney. Isis, 91(3):585–586, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237934.

Gatti:2000:BRB

[266] Hilary Gatti. Book review: Giordano Bruno: Cause, Principle, and Unity and Essays on Magic. Cause, Principle, and Unity by Giordano Bruno; Robert de Lucca; Richard J. Blackwell. Isis, 91(3):586–587, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237935.

Gingerich:2000:BRB

[267] Owen Gingerich. Book review: Kepler’s Tübingen: Stimulus to a Theological Mathematics by Charlotte Methuen. Isis, 91(3):587–588, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237936.

Bates:2000:BRB

[268] Don Bates. Book review: William Harvey and the Use of Purpose in the Scientific Revolution: Cosmos by Chance or Universe by Design? by Emerson Thomas McMullen. Isis, 91(3):588, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237937.

Long:2000:BRBb

[269] Pamela O. Long. Book review: The Garden, the Ark, the Tower, the Temple: Biblical Metaphors of Knowledge in Early Modern Europe by Jim Bennett; Scott Mandelbrote. Isis, 91(3):588–589, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237938.

Wojcik:2000:BRB

[270] Jan W. Wojcik. Book review: Catholicism in the English Protestant Imagination: Nationalism, Religion, and Literature, 1660–1745 by Raymond D. Tumbleson. Isis, 91(3):589–590, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237939.

Feingold:2000:BRB

[271] Mordechai Feingold. Book review: Isaac Newton: Eighteenth-Century Perspectives by A. Rupert Hall. Isis, 91(3):590–591, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237940.

Tweedale:2000:BRB

[272] Geoffrey Tweedale. Book review: Industrial Espionage and Technology Transfer: Britain and France in the Eighteenth Century by J. R. Harris. Isis, 91(3):591–592, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237941.

Porter:2000:BRBb

[273] Dale H. Porter. Book review: The Civil Engineering of Canals and Railways before 1850 by Mike Chrimes. Isis, 91(3):592–593, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237942.

Beisswanger:2000:BRB

[274] Gabriele Beisswanger. Book review: Zum Verhaltnis Apotheker/Arzt in Hessen: Bemuhungen in Giessen um eine Novellierung der rechtlichen Grundlagen (um 1700) by Ute Fischer-Mauch. Isis, 91(3):593–594, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237943.

Porter:2000:BRBc

[275] Charlotte M. Porter. Book review: The Selected Papers of Charles Willson Peale and His Family. Volume 4: Charles Willson Peale: His Last Years, 1821–1827 by Lillian B. Miller. Isis, 91(3):594–595, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237944.

deRegt:2000:BRB

[276] Henk W. de Regt. Book review: Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906. Book 2: The Philosopher by John Blackmore. Isis, 91(3):595, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237945.

Dorfman:2000:BRB

[277] J. R. Dorfman. Book review: Ludwig Boltzmann: The Man Who Trusted Atoms by Carlo Cercignani. Isis, 91(3):595–596, September 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237946.

Kohlstedt:2000:BRB

[278] Sally Gregory Kohlstedt. Book review: Ladies in the Laboratory? American and British Women in Science, 1800–1900 by Mary R. S. Creese; Thomas M. Creese. Isis, 91(3):596–598, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237947.

Tatarewicz:2000:BRBc

[279] Joseph N. Tatarewicz. Book review: American Astronomy: Community, Careers, and Power, 1859–1940 by John Lankford; Rickey L. Slavings. Isis, 91(3):598–599, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237948.

Bailar:2000:BRB

[280] Barbara A. Bailar. Book review: The Dominion Bureau of Statistics: A History of Canada’s Central Statistical Office and Its Antecedents, 1841–1972 by David A. Worton. Isis, 91(3):599–600, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237949.

Scott:2000:BRB

[281] Eugenie C. Scott. Book review: The Creation/Evolution Controversy: A Battle for Cultural Power by Kary Doyle Smout. Isis, 91(3):600–601, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237950.

Montgomery:2000:BRBb

[282] William Montgomery. Book review: Weltratsel und Lebenswunder: Ernst Haeckel–Werk, Wirkung, und Folgen. Isis, 91(3):601–602, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237951.

Breidbach:2000:BRB

[283] Olaf Breidbach. Book review: Haeckel’s Monism and the Birth of Fascist Ideology by Daniel Gasman. Isis, 91(3):602–603, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237952.

Schickore:2000:BRB

[284] Jutta Schickore. Book review: Membranes: Metaphors of Invasion in Nineteenth-Century Literature, Science, and Politics by Laura Otis. Isis, 91(3):603–604, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237953.

Norwood:2000:BRB

[285] Vera Norwood. Book review: Kindred Nature: Victorian and Edwardian Women Embrace the Living World by Barbara T. Gates. Isis, 91(3):604, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237954.

Sawyer:2000:BRB

[286] Richard C. Sawyer. Book review: True Gardens of the Gods: Californian-Australian Environmental Reform, 1860–1930 by Ian Tyrrell. Isis, 91(3):604–605, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237955.

Champlin:2000:BRB

[287] Peggy Champlin. Book review: The Young John Muir: An Environmental Biography by Steven J. Holmes. Isis, 91(3):605–606, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237956.

Jacknis:2000:BRB

[288] Ira Jacknis. Book review: Franz Boas: Among the Inuit of Baffin Island, 1883–1884: Journals and Letters by Franz Boas; Ludger Muller-Wille; William Barr. Isis, 91(3):606–608, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237957.

Ruse:2000:BRB

[289] Michael Ruse. Book review: Sudden Origins: Fossils, Genes, and the Emergence of Species by Jeffrey H. Schwartz. Isis, 91(3):608–609, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237958.

Dean:2000:BRBb

[290] Dennis R. Dean. Book review: Santorini and Its Eruptions by Ferdinand A. Fouque; Alexander R. McBirney. Isis, 91(3):609–610, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237959.

Stern:2000:BRB

[291] Marc J. Stern. Book review: Endless Novelty: Specialty Production and American Industrialization, 1865–1925 by Philip Scranton. Isis, 91(3): 610–611, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237960.

Reynolds:2000:BRBa

[292] Terry S. Reynolds. Book review: Gentlemen Engineers: The Working Lives of Frank and Walter Shanly by Richard White. Isis, 91(3):611–612, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237961.

Tobey:2000:BRB

[293] Ronald Tobey. Book review: His and Hers: Gender, Consumption, and Technology by Roger Horowitz; Arwen Mohum. Isis, 91(3):612, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237962.

Kozloski:2000:BRB

[294] Lillian D. Kozloski. Book review: The Bends: Compressed Air in the History of Science, Diving, and Engineering by John L. Phillips. Isis, 91 (3):612–614, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237963.

Davis:2000:BRB

[295] Audrey B. Davis. Book review: Manifesting Medicine: Bodies and Machines by Robert Bud; Bernard Finn; Helmuth Trischler. Isis, 91(3): 614–615, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237964.

Sutphen:2000:BRB

[296] Mary P. Sutphen. Book review: Deciphering Global Epidemics: Analytical Approaches to the Disease Records of World Cities, 1888–1912 by Andrew Cliff; Peter Haggett; Matthew Smallman-Raynor. Isis, 91(3): 615, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237965.

Park:2000:BRB

[297] Katharine Park. Book review: Hermaphrodites and the Medical Invention of Sex by Alice Domurat Dreger. Isis, 91(3):615–616, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237966.

Gelfand:2000:BRB

[298] Toby Gelfand. Book review: Histoire des myopathies by François Delaporte; Patrice Pinell. Isis, 91(3):616–617, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237967.

Dwyer:2000:BRB

[299] Ellen Dwyer. Book review: “They’re in the Trade… of Lunacy: They ’Cannot Interfere’–They Say”: The Scottish Lunacy Commissioners and Lunacy Reform in Nineteenth-Century Scotland by Jonathan Andrews. Isis, 91(3):617–618, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237968.

Esguerra:2000:BRB

[300] Jorge Canizares Esguerra. Book review: Un siglo de ciencia en España by Jose Manuel Sanchez Ron. Isis, 91(3):618–619, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237969.

Brown:2000:BRBc

[301] Laurie M. Brown. Book review: Niels Bohr: Collected Works. Volume 10: Complementarity beyond Physics (1928–1962), by Niels Bohr and David Favrholdt. Isis, 91(3):619–620, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237970.

Pyenson:2000:BRB

[302] Lewis Pyenson. Book review: Einstein’s German World by Fritz Stern. Isis, 91(3):620–621, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237971.

Rechenberg:2000:BRB

[303] Helmut Rechenberg. Book review: The Defining Years of Nuclear Physics: 1932–1960s by M. Mladjenovic. Isis, 91(3):621, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237972.

Doel:2000:BRB

[304] Ronald E. Doel. Book review: Looking for Earth: The Race to Find New Solar Systems by Alan Boss. Isis, 91(3):621–622, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237973.

Johnson:2000:BRB

[305] Jeffrey Allan Johnson. Book review: Determinants in the Evolution of the European Chemical Industry, 1900–1939: New Technologies, Political Frameworks, Markets, and Companies by Anthony S. Travis; Harm G. Schroter; Ernst Homburg; Peter J. T. Morris. Isis, 91(3):622–623, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237974.

Slater:2000:BRB

[306] Leo B. Slater. Book review: A Fifty-Year Love Affair with Organic Chemistry by William S. Johnson. Isis, 91(3):623–624, September 2000.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237975.

Schneider:2000:BRB

[307] Laurence Schneider. Book review: Ichuanxue yu baijia zhengming: 1956 Qingdao zuotanhui yanjiu (Genetics and the “Let a Hundred Schools Contend”) [Policy]: A Study of the Consequences of the 1956 Qingdao Genetics Symposium. Isis, 91(3):624–625, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237976.

Mark:2000:BRB

[308] Joan Mark. Book review: Grit-Tempered: Early Women Archaeologists in the Southeastern United States by Nancy Marie White; Lynne P. Sullivan; Rochelle A. Marrinan. Isis, 91(3):625–626, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237977.

Murphy:2000:BRB

[309] David T. Murphy. Book review: Das Atlantropa-Projekt: Die Geschichte einer gescheiterten Vision: Hermann Sorgel und die Absenkung des Mittelmeers by Alexander Gall. Isis, 91(3):626–627, September 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237978.

Inkster:2000:BRB

[310] Ian Inkster. Book review: Manufacturing Ideology: Scientific Management in Twentieth–Century Japan by William M. Tsutsui. Isis, 91(3): 627–628, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237979.

Wang:2000:BRBb

[311] Zuoyue Wang. Book review: Thread of the Silkworm by Iris Chang. Isis, 91(3):628–629, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237980.

Rusnak:2000:BRB

[312] Kevin M. Rusnak. Book review: Way Station to Space: A History of the John C. Stennis Space Center by Mack R. Herring. Isis, 91(3):629–630, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237981.

Hacker:2000:BRB

[313] Barton C. Hacker. Book review: Fast Tanks and Heavy Bombers: Innovation in the U.S. Army, 1917–1945 by David E. Johnson. Isis, 91 (3):630–631, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237982.

Teich:2000:BRBa

[314] Albert H. Teich. Book review: Limited by Design: R&D Laboratories in the U.S. National Innovation System by Michael Crow; Barry Bozeman. Isis, 91(3):631–632, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237983.

Home:2000:BRB

[315] R. W. Home. Book review: DSIR: Making Science Work for New Zealand: Themes from the History of the Department of Scientific and Industrial Research, 1926–1992 by Ross Galbreath. Isis, 91(3):632–633, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237984.

Madison:2000:BRB

[316] Mark Madison. Book review: New Deal Medicine: The Rural Health Programs of the Farm Security Administration by Michael R. Grey. Isis, 91 (3):633–634, September 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237985.

Junod:2000:BRB

[317] Suzanne White Junod. Book review: The Politics of International Health: The Children’s Vaccine Initiative and the Struggle to Develop Vaccines for the Third World by William Muraskin. Isis, 91(3):634–635, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237986.

Jefferson:2000:BRB

[318] James W. Jefferson. Book review: The Antidepressant Era by David Healy. Isis, 91(3):635–636, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237987.

Joralemon:2000:BRB

[319] Donald Joralemon. Book review: Xenotransplantation: Scientific Frontiers and Public Policy by Jay Fishman; David Sachs; Rashid Shaikh.

Isis, 91(3):636–637, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237988.

Napoli:2000:BRB

[320] Donald S. Napoli. Book review: Psychologists on the March: Science, Practice, and Professional Identity in America, 1929–1969 by James Capshew. Isis, 91(3):637–638, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237989.

Mills:2000:BRB

[321] John A. Mills. Book review: Taming the Troublesome Child: American Families, Child Guidance, and the Limits of Psychiatric Authority by Kathleen W. Jones. Isis, 91(3):638–639, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237990.

Metaxas:2000:BRB

[322] Virginia Metaxas. Book review: A Doctor of Their Own: The History of Adolescent Medicine by Heather Munro Prescott. Isis, 91(3):639, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237991.

Mirowski:2000:BRB

[323] Philip Mirowski. Book review: How We Became Posthuman: Virtual Bodies in Cybernetics, Literature, and Informatics by Katherine Hayles. Isis, 91(3):639–640, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237992.

Capeilleres:2000:BRB

[324] Fabien Capeilleres. Book review: Ernst Cassirer: Dalla scuola di Marburgo alla filosofia della cultura by Massimo Ferrari. Isis, 91(3):640–641, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237993.

Travis:2000:BRB

[325] Anthony S. Travis. Book review: Nobel Laureates in Chemistry, 1901– 1992 by Laylin K. James. Isis, 91(3):641–642, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237994.

Rosenberg:2000:BRB

[326] Robert Rosenberg. Book review: The Text in the Machine: Electronic Texts in the Humanities by Toby Burrows. Isis, 91(3):642–643, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237995.

Anonymous:2000:Cc

[327] Anonymous. Collections. Isis, 91(3):643–644, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237996.

Anonymous:2000:BMd

[328] Anonymous. Back matter. Isis, 91(3):645–650, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237997.

Anonymous:2000:FMd

[329] Anonymous. Front matter. Isis, 91(3):??, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237903.

Lightman:2000:VTV

[330] Bernard Lightman. The visual theology of Victorian popularizers of science: From reverent eye to chemical retina. Isis, 91(4):651–680, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236819.

Schneider:2000:LKE

[331] Daniel W. Schneider. Local knowledge, environmental politics, and the founding of ecology in the United States: Stephen Forbes and “The Lake as a Microcosm” (1887). Isis, 91(4):681–705, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236820.

Santesmases:2000:SOB

[332] Maria Jesus Santesmases. Severo ochoa and the biomedical sciences in Spain under franco, 1959–1975. Isis, 91(4):706–734, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236821.

Holmes:2000:RCP

[333] Frederic L. Holmes. The “revolution in chemistry and physics”: Overthrow of a reigning paradigm or competition between contemporary research programs? Isis, 91(4):735–753, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236822.

Levinson:2000:LE

[334] Mark Levinson. Letter to the Editor. Isis, 91(4):754, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236823.

Weiner:2000:BRB

[335] Dora B. Weiner. Book review: L’Egypte, une aventure savante: Avec Bonaparte, Kleber, Menou (1798–1801). Isis, 91(4):755–757, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236824.

Wilson:2000:BRBc

[336] Curtis Wilson. Book review: The Sun in the Church: Cathedrals as Solar Observatories. Isis, 91(4):758–760, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236825.

Nyhart:2000:BRB

[337] Lynn K. Nyhart. Book review: Wissenschaftspopularisierung im 19. Jahrhundert: Burgerliche Kultur, naturwissenschaftliche Bildung und die deutsche Offentlichkeit, 1848–1914 by Andreas W. Daum. Isis, 91(4): 760–761, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236826.

Sobol:2000:BRB

[338] Peter G. Sobol. Book review: Scientific Development and Misconceptions through the Ages: A Reference Guide by Robert E. Krebs. Isis, 91(4): 761–762, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236827.

Stamhuis:2000:BRB

[339] Ida H. Stamhuis. Book review: Statistics on the Table: The History of Statistical Concepts and Methods by Stephen M. Stigler. Isis, 91(4): 762–763, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236828.

Thompson:2000:BRBa

[340] Emily Thompson. Book review: Sounds of Our Times: Two Hundred Years of Acoustics by Robert T. Beyer. Isis, 91(4):763–764, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236829.

Rocke:2000:BRBb

[341] A. J. Rocke. Book review: Communicating Chemistry: Textbooks and Their Audiences, 1789–1939 by Anders Lundgren; Bernadette Bensaude-Vincent. Isis, 91(4):764–765, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236830.

vanEe:2000:BRB

[342] Patricia Molen van Ee. Book review: Cartography in the Traditional African, American, Arctic, Australian, and Pacific Societies by David Woodward; G. Malcolm Lewis. Isis, 91(4):765–766, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236831.

Warren:2000:BRB

[343] Leonard Warren. Book review: Natural History Investigations in South Carolina: From Colonial Times to the Present by Albert Sanders; William D. Anderson. Isis, 91(4):766–767, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236832.

Reynolds:2000:BRBb

[344] Terry S. Reynolds. Book review: The Sanitary City: Urban Infrastructure in America from Colonial Times to the Present by Martin V. Melosi. Isis, 91(4):767–768, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236833.

Cooke:2000:BRB

[345] Kathy J. Cooke. Book review: Clean Living Movements: American Cycles of Health Reform by Ruth Clifford Engs. Isis, 91(4):768–769, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236834.

Craig:2000:BRB

[346] Charlotte M. Craig. Book review: Das Vorurteil im Leibe: Eine Einfuhrung in die Physiognomik by Claudia Schmolders. Isis, 91(4):769–770, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236835.

Junker:2000:BRB

[347] Thomas Junker. Book review: Wissenschaftsaustausch zwischen Russland und Westeuropa, insbesondere Deutschland, in der Botanik und Pharmazie vom 18. bis zum fruhen 20. Jahrhundert by Natalie Kunkel. Isis, 91(4):770–771, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236836.

Arnold:2000:BRB

[348] David Arnold. Book review: Climates and Constitutions: Health, Race, Environment, and British Imperialism in India, 1600–1850 by Mark Harrison. Isis, 91(4):771–772, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236837.

Guralnick:2000:BRB

[349] Stanley M. Guralnick. Book review: The Sacred and the Secular University by Jon H. Roberts; James Turner. Isis, 91(4):772–773, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236838.

Snodgrass:2000:BRB

[350] Mary Ellen Snodgrass. Book review: The Star of Bethlehem: The Legacy of the Magi by Michael R. Molnar. Isis, 91(4):773, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236839.

Lozovsky:2000:BRB

[351] Natalia Lozovsky. Book review: Mapping Time and Space: How Medieval Mapmakers Viewed Their World by Evelyn Edson. Isis, 91(4):773–774, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236840.

Schipperges:2000:BRB

[352] Heinrich Schipperges. Book review: Medical Classroom Practice: Petrus Hispanus’ Questions on Isagoge, Tegni, Regimen Acutorum, and Prognostica (c. 1245–50) by Fernando Salmon. Isis, 91(4):774, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236841.

Reeves:2000:BRB

[353] Eileen Reeves. Book review: The Triumph of Vulcan: Sculptors’ Tools, Porphyry, and the Prince in Ducal Florence by Suzanne B. Butters. Isis, 91(4):774–775, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236842.

Harney:2000:BRB

[354] Michael Harney. Book review: Le recueil des plus celebres astrologues de Simon de Phares. Volume 1 by Jean-Patrice Boudet. Isis, 91(4):775–776, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236843.

Shapiro:2000:BRB

[355] Barbara Shapiro. Book review: 1543 and All That: Image and Word, Change and Continuity in the Proto-Scientific Revolution by Guy Freeland; Anthony Corones. Isis, 91(4):776–777, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236844.

Gosselin:2000:BRB

[356] Edward Gosselin. Book review: Giordano Bruno and Renaissance Science by Hilary Gatti. Isis, 91(4):777–778, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236845.

Malet:2000:BRBb

[357] Antoni Malet. Book review: Reasoning with the Infinite: From the Closed World to the Mathematical Universe by Michel Blay; M. B. DeBevoise. Isis, 91(4):778–779, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236846.

Meli:2000:BRBb

[358] Domenico Bertoloni Meli. Book review: The Cambridge Companion to Galileo by Peter Machamer. Isis, 91(4):779–780, December 2000. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236847.

Cohen:2000:BRBb

[359] I. Bernard Cohen. Book review: Edmond Halley: Charting the Heavens and the Seas by Alan Cook. Isis, 91(4):780–781, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236848.

Ranea:2000:BRBb

[360] Alberto Guillermo Ranea. Book review: Ariadnefaden im Wissenschaftslabyrinth: Studien zu Galilei: Historiographie-Mathematik-Wirkung by Volker R. Remmert. Isis, 91(4):781, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236849.

Principe:2000:BRB

[361] Lawrence M. Principe. Book review: Aspects de la tradition alchimique au XVIIe siècle: Actes du colloque international de l’Université de Reims-Champagne-Ardenne (Reims, 28 et 29 novembre 1996) by Frank Greiner. Isis, 91(4):782, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236850.

Yost:2000:BRB

[362] Robinson M. Yost. Book review: Did the Royal Society Matter in the Eighteenth Century? by Richard Sorrenson. Isis, 91(4):782–783, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236851.

Winthrop-Young:2000:BRB

[363] Geoffrey Winthrop-Young. Book review: Observationes: Die lateinischen Schriften by Georg Christoph Lichtenberg; Dag Nikolaus Hasse. Isis, 91 (4):783–784, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236852.

Dettloff:2000:BRB

[364] John Dettloff. Book review: Lavoisier et l’Encyclopédie méthodique: Le manuscrit des régisseurs des poudres et salpêtres pour le Dictionnaire de l’artillerie (1787) by Patrice Bret. Isis, 91(4):784–785, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236853.

Evans:2000:BRB

[365] James Evans. Book review: Lectures on Natural Philosophy by Richard Helsham; D. A. Attis; P. Kelly; D. Weire. Isis, 91(4):785–786, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236854.

Dettelbach:2000:BRB

[366] Michael Dettelbach. Book review: Geography and Enlightenment by David N. Livingstone; Charles W. J. Withers. Isis, 91(4):786–788, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236855.

Rhees:2000:BRB

[367] David Rhees. Book review: Frankenstein’s Children: Electricity, Exhibition, and Experiment in Early Nineteenth-Century London by Iwan Rhys Morus. Isis, 91(4):788–789, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236856.

Crawford:2000:BRB

[368] T. Hugh Crawford. Book review: Melville’s Anatomies by Samuel Otter; Herman Melville: Stargazer by Brett Zimmerman. Isis, 91(4):789–791, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236857.

Leonard:2000:BRB

[369] Robert Leonard. Book review: The Secret Origins of Modern Microeconomics: Dupuit and the Engineers by Robert B. Ekelund,; Robert F. Hebert. Isis, 91(4):791–792, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236858.

Hunt:2000:BRBb

[370] Bruce J. Hunt. Book review: The Correspondence of Michael Faraday. Volume 4: January 1849–October 1855, Letters 2146–3032 by Michael Faraday; Frank A. J. L. James. Isis, 91(4):792–794, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236859.

Moyer:2000:BRB

[371] Albert E. Moyer. Book review: The Papers of Joseph Henry. Volume 7: January 1847–December 1849: The Smithsonian Years by Joseph Henry; Marc Rothenberg; Paul H. Theerman; Kathleen W. Dorman; John C. Rumm; Deborah Y. Jeffries; The Papers of Joseph Henry. Volume 8: January 1850–December 1853: The Smithsonian Years by Joseph Henry; Marc Rothenberg; Kathleen W. Dorman; Deborah Y. Jeffries; Frank R. Millikan. Isis, 91(4):794–795, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236860.

Herbert:2000:BRB

[372] Sandra Herbert. Book review: Lyell in America: Transatlantic Geology, 1841–1853 by Leonard G. Wilson. Isis, 91(4):795–797, December 2000.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236861.

Pritchard:2000:BRBb

[373] James Pritchard. Book review: Yellowstone: The Creation and Selling of an American Landscape, 1870–1903 by Chris J. Magoc. Isis, 91(4): 797–798, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236862.

Hinsley:2000:BRB

[374] Curtis M. Hinsley. Book review: Ancient Monuments of the Mississippi Valley by Ephraim G. Squier; Edwin H. Davis; David J. Meltzer. Isis, 91(4):798, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236863.

Castonguay:2000:BRB

[375] Stephane Castonguay. Book review: The Formative Years of Plant Pathology in the United States by C. Lee Campbell; Paul D. Peterson; Clay S. Griffith. Isis, 91(4):798–799, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236864.

Razzell:2000:BRB

[376] Peter Razzell. Book review: Sick, Not Dead: The Health of British Workingmen during the Mortality Decline by James C. Riley. Isis, 91 (4):799–800, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236865.

Greenberg:2000:BRB

[377] Valerie D. Greenberg. Book review: Freud and the Neurosciences: From Brain Research to the Unconscious by Giselher Guttmann; Inge Scholz-Strasser. Isis, 91(4):800–801, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236866.

Crabtree:2000:BRB

[378] Adam Crabtree. Book review: From Soul to Mind: The Emergence of Psychology from Erasmus Darwin to William James by Edward S. Reed; Putting Psychology in Its Place: An Introduction from a Critical Historical Perspective by Graham Richards. Isis, 91(4):801–802, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236867.

Porter:2000:BRBd

[379] Dorothy Porter. Book review: The Problem of Mental Deficiency: Eugenics, Democracy, and Social Policy in Britain, c. 1870–1959 by Mathew Thomson. Isis, 91(4):802–804, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236868.

Grabiner:2000:BRB

[380] Judith V. Grabiner. Book review: The Man Who Loved Only Numbers: The Story of Paul Erdos and the Search for Mathematical Truth by Paul Hoffman. Isis, 91(4):804–805, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236869.

Hestmark:2000:BRB

[381] Geir Hestmark. Book review: The Centennial of S. A. Andree’s North Pole Expedition by Urban Wrakberg. Isis, 91(4):805–806, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236870.

Fodera:2000:BRB

[382] Giorgia Fodera. Book review: Guido Horn d’Arturo e lo specchio a tasselli by Marina Zuccoli; Fabrizio Bonoli. Isis, 91(4):806–807, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236871.

Kragh:2000:BRB

[383] Helge Kragh. Book review: Heavy Water and the Wartime Race for Nuclear Energy by Per F. Dahl. Isis, 91(4):807–808, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236872.

Ito:2000:BRB

[384] Kenji Ito. Book review: Peace and War: Reminiscences of a Life on the Frontiers of Science by Robert Serber; Robert P. Crease. Isis, 91 (4):808–809, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236873.

Rothenberg:2000:BRB

[385] Marc Rothenberg. Book review: History of the IAU: The Birth and First Half-Century of the International Astronomical Union by Adriaan Blaauw. Isis, 91(4):809, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236874.

Sime:2000:BRB

[386] Ruth Lewin Sime. Book review: Dorothy Hodgkin: A Life by Georgina Ferry. Isis, 91(4):810, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236875.

Sigurdsson:2000:BRB

[387] Skuli Sigurdsson. Book review: Hermann Weyl und die Mathematik an der ETH Zurich, 1913–1930 by Gunther Frei; Urs Stammbach. Isis, 91 (4):810–811, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236876.

Walter:2000:BRB

[388] Maila Walter. Book review: The Diamond Makers by Robert M. Hazen. Isis, 91(4):811–812, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236877.

Weir:2000:BRB

[389] Gary E. Weir. Book review: Under Ice: Waldo Lyon and the Development of the Arctic Submarine by William M. Leary. Isis, 91(4):812–814, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236878.

Barlow:2000:BRB

[390] Jeffrey G. Barlow. Book review: Destined for Glory: Dive Bombing, Midway, and the Evolution of Carrier Air Power by Thomas Wildenberg. Isis, 91(4):814–815, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236879.

Tighe:2000:BRB

[391] Janet A. Tighe. Book review: Time to Heal: American Medical Education from the Turn of the Century to the Managed Care Era by Kenneth M. Ludmerer. Isis, 91(4):815, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236880.

Thompson:2000:BRBb

[392] Charis Thompson. Book review: Cultural Conceptions: On Reproductive Technologies and the Remaking of Life by Valerie Hartouni. Isis, 91 (4):815–816, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236881.

Gallagher:2000:BRB

[393] Nancy Gallagher. Book review: Frontiers of Medicine in the Anglo-Egyptian Sudan, 1899–1940 by Heather Bell. Isis, 91(4):816–817, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236882.

Jones:2000:BRB

[394] David S. Jones. Book review: King of Hearts: The True Story of the Maverick Who Pioneered Open-Heart Surgery by G. Wayne Miller. Isis, 91(4):817–818, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236883.

Usselman:2000:BRB

[395] Steven W. Usselman. Book review: The One Best Way: Frederick Winslow Taylor and the Enigma of Efficiency by Robert Kanigel. Isis, 91 (4):818–819, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236884.

Ericson:2000:BRB

[396] Steven J. Ericson. Book review: From Steam to Diesel: Managerial Customs and Organizational Capabilities in the Twentieth–Century American Locomotive Industry by Albert J. Churella. Isis, 91(4):819–820, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236885.

Olson:2000:BRBb

[397] Janet C. Olson. Book review: Breeding Better Vermonters: The Eugenics Project in the Green Mountain State by Nancy L. Gallagher. Isis, 91(4): 820+822, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236887.

Buhle:2000:BRBb

[398] Mari Jo Buhle. Book review: Queen Silver: The Godless Girl by Wendy McElroy. Isis, 91(4):820–821, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236886.

Peterson:2000:BRB

[399] Paul D. Peterson. Book review: Tobacco Mosaic Virus: One Hundred Years of Contributions to Virology by Karen-Beth Scholthof; John G.

Shaw; Milton Zaitlin. Isis, 91(4):822–823, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236888.

Kuklick:2000:BRB

[400] Henrika Kuklick. Book review: Malinowski’s Kiriwina: Fieldwork Photography, 1915–1918 by Michael W. Young. Isis, 91(4):823–824, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236889.

Bruce:2000:BRB

[401] Darryl Bruce. Book review: Constructing Scientific Psychology: Karl Lashley’s Mind-Brain Debates by Nadine M. Weidman. Isis, 91(4):824– 825, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236890.

Bannister:2000:BRB

[402] Robert C. Bannister. Book review: Social Control and Public Intellect: The Legacy of Edward A. Ross by Sean H. McMahon. Isis, 91(4):825– 826, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236891.

Speaker:2000:BRB

[403] Susan L. Speaker. Book review: A Social History of the Minor Tranquilizers: The Quest for Small Comfort in the Age of Anxiety by Mickey C. Smith. Isis, 91(4):826–827, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236892.

Braslow:2000:BRB

[404] Joel Braslow. Book review: The Rise of Psychopharmacology and the Story of CINP by Thomas A. Ban; David Healy; Edward Shorter. Isis, 91(4):827, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236893.

Pols:2000:BRB

[405] Hans Pols. Book review: Gestalt Psychology in German Culture, 1890– 1967: Holism and the Quest for Objectivity by Mitchell G. Ash. Isis, 91 (4):827–828, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236894.

Edwards:2000:BRB

[406] Paul Edwards. Book review: Planet Management: Limits to Growth, Computer Simulation, and the Emergence of Global Spaces by Fernando Elichirigoity. Isis, 91(4):828–829, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236895.

Gottweis:2000:BRB

[407] Herbert Gottweis. Book review: Biotechnology Backstage by Paul Aebersold. Isis, 91(4):829–830, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236896.

Nelkin:2000:BRB

[408] Dorothy Nelkin. Book review: Controlling Our Destinies: Historical, Philosophical, Ethical, and Theological Perspectives on the Human Genome Project by Phillip R. Sloan. Isis, 91(4):830–831, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236897.

Geschwind:2000:BRB

[409] Carl-Henry Geschwind. Book review: Promoting Risk: Constructing the Earthquake Threat by Robert A. Stallings. Isis, 91(4):831–832, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236898.

Teich:2000:BRBb

[410] Albert H. Teich. Book review: The Politics of Pure Science by Daniel S. Greenberg. Isis, 91(4):832, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236899.

Manier:2000:BRB

[411] Edward Manier. Book review: Introduction to the Philosophy of Science: Cutting Nature at Its Seams by Robert Klee. Isis, 91(4):832–833, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236900.

Lederer:2000:BRB

[412] Susan E. Lederer. Book review: Images of Animals: Anthropomorphism and Animal Mind by Eileen Crist; Humans and Other Animals by Arien Mack. Isis, 91(4):834–835, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236901.

Finlayson:2000:BRB

[413] Chris Finlayson. Book review: Fishy Business: Salmon, Biology, and the Social Construction of Nature by Rik Scarce. Isis, 91(4):835, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236902.

Lowy:2000:BRBb

[414] Ilana Lowy. Book review: French DNA: Trouble in Purgatory by Paul Rabinow. Isis, 91(4):835–837, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236903.

Sismondo:2000:BRB

[415] Sergio Sismondo. Book review: Reconfiguring Truth: Post-modernism, Science Studies, and the Search for a New Model of Knowledge by Steven C. Ward. Isis, 91(4):837–838, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236904.

Marcum:2000:BRB

[416] James A. Marcum. Book review: Philosophy of Science: The Historical Background by Joseph J. Kockelmans. Isis, 91(4):838, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236905.

Cat:2000:BRB

[417] Jordi Cat. Book review: Science without Laws by Ronald N. Giere. Isis, 91(4):838–839, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236906.

Barrett:2000:BRBb

[418] Jeffrey A. Barrett. Book review: Philosophical Concepts in Physics: The Historical Relation between Philosophy and Scientific Theories by James T. Cushing. Isis, 91(4):839–840, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236907.

Grattan-Guinness:2000:BRB

[419] I. Grattan-Guinness. Book review: Popper by Frederic Raphael. Isis, 91(4):840, December 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236908.

Rudwick:2000:BRB

[420] Martin Rudwick. Book review: The Papers of H. T. De la Beche (1796– 1855) in the National Museum of Wales by Tom Sharpe; Paul J. Mc-Cartney. Isis, 91(4):840–841, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236909.

Kidwell:2000:BRBc

[421] Peggy Aldrich Kidwell. Book review: A Station Favorable to the Pursuits of Science: Primary Materials in the History of Mathematics at the United States Military Academy by Joe Albree; David C. Arney; V. Frederick Rickey. Isis, 91(4):841–842, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236910.

Taylor:2000:BRB

[422] Kenneth L. Taylor. Book review: Sciences of the Earth: An Encyclopedia of Events, People, and Phenomena by Gregory A. Good. Isis, 91(4):842– 843, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236911.

Anonymous:2000:Cd

[423] Anonymous. Collections. Isis, 91(4):843–845, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236912.

Anonymous:2000:BMe

[424] Anonymous. Back matter. Isis, 91(4):846–851, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236913.

Anonymous:2000:VI

[425] Anonymous. Volume information. Isis, 91(4):852–870, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236817.

Anonymous:2000:FMe

[426] Anonymous. Front matter. Isis, 91(4):??, December 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236818.

Anonymous:2001:CBH

[427] Anonymous. Current bibliography of the history of science and its cultural influences, 2001. Isis, 92(??):1–232, ???? 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080915.

Anonymous:2001:VIa

[428] Anonymous. Volume information. Isis, 92(??):233–263, ???? 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080913.

Anonymous:2001:BMa

[429] Anonymous. Back matter. Isis, 92(??):??, ???? 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080916.

Anonymous:2001:FMa

[430] Anonymous. Front matter. Isis, 92(??):i–xxii, ???? 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080914.

Knowles:2001:IVE

[431] Scott G. Knowles and Stuart W. Leslie. “Industrial Versailles”: Eero Saarinen’s corporate campuses for GM, IBM, and AT&T. Isis, 92(1): 1–33, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237325.

Lynn:2001:DEP

[432] Michael R. Lynn. Divining the enlightenment: Public opinion and popular science in old regime France. Isis, 92(1):34–54, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237326.

Richmond:2001:WEH

[433] Marsha L. Richmond. Women in the early history of genetics: William Bateson and the Newnham College Mendelians, 1900–1910. Isis, 92(1): 55–90, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237327.

Steen:2001:PPS

[434] Kathryn Steen. Patents, patriotism, and “skilled in the art” USA v. The Chemical Foundation, Inc., 1923–1926. Isis, 92(1):91–122, March 2001.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237328.

Buchwald:2001:ECT

[435] Jed Z. Buchwald and I. Bernard Cohen. Eloge: Clifford Truesdell, 1919– 2000. Isis, 92(1):123–125, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237329.

Carlino:2001:BRB

[436] Andrea Carlino. Book review: On the Fabric of the Human Body, by Andreas Vesalius; William Franck Richardson; John Burd Carman; On the Fabric of the Human Body. Book 2: The Ligaments and Muscles by Andreas Vesalius; William Franck Richardson; John Burd Carman. Book 1: The Bones and Cartilages. Isis, 92(1):126–127, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237330.

Servos:2001:BRB

[437] John W. Servos. Book review: Science and Technology in World History: An Introduction by James E. McClellan; Harold Dorn. Isis, 92(1):127– 128, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237331.

Stoll:2001:BRBa

[438] Mark Stoll. Book review: Nature: Western Attitudes since Ancient Times by Peter Coates. Isis, 92(1):128–129, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237332.

Zebrowski:2001:BRBa

[439] Ernest Zebrowski, Jr. Book review: An Imaginary Tale: The Story of the Square Root of minus One by Paul J. Nahin. Isis, 92(1):129–130, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237333.

Nahin:2001:BRB

[440] Paul J. Nahin. Book review: A History of the Circle: Mathematical Reasoning and the Physical Universe by Ernest Zebrowski. Isis, 92(1): 130, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237334.

McAllister:2001:BRB

[441] James W. McAllister. Book review: Fearful Symmetry: The Search for Beauty in Modern Physics by A. Zee. Isis, 92(1):130–131, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237335.

McCormmach:2001:BRB

[442] Russell McCormmach. Book review: The Language of Physics: The Calculus and the Development of Theoretical Physics in Europe, 1750– 1914 by Elizabeth Garber. Isis, 92(1):131–132, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237336.

Withers:2001:BRB

[443] Charles W. J. Withers. Book review: English Maps: A History by Catherine Delano-Smith; Roger J. P. Kain. Isis, 92(1):132–133, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237337.

Slowinski:2001:BRB

[444] Joseph Slowinski. Book review: In Search of Deep Time: Beyond the Fossil Record to a New History of Life by Henry Gee. Isis, 92(1):133– 134, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237338.

McMenamin:2001:BRB

[445] Mark A. S. McMenamin. Book review: The Connecticut Valley in the Age of Dinosaurs: A Guide to the Geologic Literature, 1681–1995 by Nicholas G. McDonald. Isis, 92(1):134–135, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237339.

Blackwell:2001:BRB

[446] Richard J. Blackwell. Book review: The Jesuits: Cultures, Sciences, and the Arts, 1540–1773 by John W. O’Malley; Gauvin Alexander Bailey; Steven J. Harris; T. Frank Kennedy. Isis, 92(1):136, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237340.

Bresadola:2001:BRB

[447] Marco Bresadola. Book review: The Matter of Motion and Galvani’s Frogs by B. Innes Williams. Isis, 92(1):136–138, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237341.

Payne:2001:BRBa

[448] Lynda Stephenson Payne. Book review: Medicine and Society in Early Modern Europe by Mary Lindemann. Isis, 92(1):138, March 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237342.

Canizares-Esguerra:2001:BRB

[449] Jorge Cañizares-Esguerra. Book review: La imagen científica de la vida: La contribución Valenciana a la ilustración médica y biológica (siglos XVI–XIX) by José Maria López Piñero; Felipe Jerez Moliner. Isis, 92 (1):138–139, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237343.

Maulitz:2001:BRB

[450] Russell C. Maulitz. Book review: The Greatest Benefit to Mankind: A Medical History of Humanity by Roy Porter. Isis, 92(1):139–140, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237344.

Evans:2001:BRBa

[451] Hughes Evans. Book review: History of Medicine: A Scandalously Short Introduction by Jacalyn Duffin. Isis, 92(1):140–141, March 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237345.

Cueto:2001:BRB

[452] Marcos Cueto. Book review: High Life: A History of High-Altitude Physiology and Medicine by John B. West. Isis, 92(1):141–142, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237346.

Desgupta:2001:BRB

[453] Subrata Desgupta. Book review: Uncharted Terrains: Essays on Science Popularisation in Pre-Independence India by Narender K. Sehgal; Satpal Sangwan; Subodh Mahanti. Isis, 92(1):142, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237347.

Gates:2001:BRB

[454] Barbara T. Gates. Book review: Farther Afield: In the Study of Nature-Oriented Literature by Patrick D. Murphy. Isis, 92(1):142–143, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237348.

Marcum:2001:BRB

[455] James A. Marcum. Book review: Great Feuds in Science: Ten of the Liveliest Disputes Ever by Hal Hellman. Isis, 92(1):143–144, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237349.

Gordon:2001:BRBa

[456] Robert B. Gordon. Book review: It’s about Time: A History of Archaeological Dating in North America by Stephen E. Nash. Isis, 92(1):144–145, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237350.

Shapiro:2001:BRBa

[457] Alan E. Shapiro. Book review: A History of Color: The Evolution of Theories of Lights and Color by Robert A. Crone. Isis, 92(1):145, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237351.

Hendershot:2001:BRB

[458] Cyndy Hendershot. Book review: Chemistry and Science Fiction by Jack H. Stocker. Isis, 92(1):145–146, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237352.

Koblitz:2001:BRB

[459] Ann Hibner Koblitz. Book review: Feminism, Science, and the Philosophy of Science by Lynn Hankinson Nelson; Jack Nelson. Isis, 92(1): 146–147, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237353.

Cordle:2001:BRB

[460] Daniel Cordle. Book review: Prometheus Bedeviled: Science and the Contradictions of Contemporary Culture by Norman Levitt. Isis, 92(1): 147–148, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237354.

Robson:2001:BRB

[461] Eleanor Robson. Book review: Ancient Astronomy and Celestial Divination by N. M. Swerdlow. Isis, 92(1):148–149, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237355.

vonPlato:2001:BRB

[462] Jan von Plato. Book review: Ptolemy’s Almagest by Ptolemy; G. J. Toomer; Preceptum Canonis Ptolomei by David Pingree. Isis, 92(1): 149–150, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237356.

Lehoux:2001:BRB

[463] Daryn Lehoux. Book review: Ptolemy and the Foundations of Ancient Mathematical Optics: A Source-Based Guided Study by A. Mark Smith. Isis, 92(1):150, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237357.

Bricker:2001:BRB

[464] Harvey M. Bricker. Book review: Star Gods of the Maya: Astronomy in Art, Folklore, and Calendars by Susan Milbrath. Isis, 92(1):150– 151, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237358.

Allen:2001:BRBa

[465] James P. Allen. Book review: Ancient Egyptian Science: A Source Book. Volume 3: Ancient Egyptian Mathematics by Marshall Clagett. Isis, 92 (1):151–152, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237359.

Riddle:2001:BRB

[466] John M. Riddle. Book review: Greek Medicine: From the Heroic to the Hellenistic Age: A Source Book by James Longrigg. Isis, 92(1):152– 153, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237360.

Ganson:2001:BRB

[467] Todd Stuart Ganson. Book review: Aristotle’s Metaphysics by Aristotle; Joe Sachs. Isis, 92(1):153–154, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237361.

Craik:2001:BRB

[468] Elizabeth Craik. Book review: Hippocrates’ Woman: Reading the Female Body in Ancient Greece by Helen King. Isis, 92(1):154–155, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237362.

Ivry:2001:BRB

[469] Alfred L. Ivry. Book review: Oeuvres philosophiques et scientifiques d’al-Kindî. Volume 2: Métaphysique et cosmologie by Roshdi Rashed; Jean Jolivet. Isis, 92(1):155–156, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237363.

Crossgrove:2001:BRB

[470] William Crossgrove. Book review: Medieval Embryology in the Vernacular: The Case of De Spermate by Päivi Pahta. Isis, 92(1):156, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237364.

Colon:2001:BRB

[471] Angel R. Colón. Book review: The Rescue of the Innocents: Endangered Children in Medieval Miracles by Ronald C. Finucane. Isis, 92(1):156– 157, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237365.

Oggins:2001:BRB

[472] Robin S. Oggins. Book review: Le Liber accipitrum de Grimaldus: Un traité d’autourserie du haut Moyen Âge by Grimaldus; An Smets. Isis, 92(1):157–158, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237366.

Lemay:2001:BRB

[473] H. R. Lemay. Book review: Glosae super Boethium by Guillelmus de Conchis; L. Nauta. Isis, 92(1):158–159, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237367.

Lang:2001:BRB

[474] Helen S. Lang. Book review: Zeichen und Wissen: Das Verhältnis der Zeichentheorie zur Theorie des Wissens und der Wissenschafter im dreizehnten Jahrhundert by Michael Fuchs. Isis, 92(1):159–160, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237368.

Freudenthal:2001:BRB

[475] Gad Freudenthal. Book review: Time Matters: Time, Creation, and Cosmology in Medieval Jewish Philosophy by T. M. Rudavsky. Isis, 92 (1):160–161, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237369.

Shank:2001:BRB

[476] Michael H. Shank. Book review: Die Universität Wien im Mittelalter: Beiträge und Forschungen by Paul Uiblein; Kurt Mühlberger; Karl Kadletz. Isis, 92(1):161, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237370.

Ritvo:2001:BRB

[477] Harriet Ritvo. Book review: The Pope’s Elephant: An Elephant’s Journey from Deep in India to the Heart of Rome by Silvio A. Bedini. Isis, 92(1):161–162, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237371.

Lattis:2001:BRB

[478] James M. Lattis. Book review: The Unmaking of the Medieval Christian Cosmos, 1500–1760: From Solid Heavens to Boundless Aether by W. G. L. Randles. Isis, 92(1):163, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237372.

Phillips:2001:BRB

[479] Carla Rahn Phillips. Book review: Diálogo entre un vizcayno y un montañés sobre la fábrica de navíos by María Isabel Vicente Maroto. Isis, 92(1):163–164, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237373.

McMahon:2001:BRB

[480] Susan McMahon. Book review: Libellus de re herbaria novus (1538) by William Turner; Mats Ryden; Hans Helander; Kerstin Olsson. Isis, 92(1):164–165, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237374.

Nutton:2001:BRB

[481] Vivian Nutton. Book review: Thomas Phaer and the Boke of Chyldren (1544) by Thomas Phaer; Rick Bowers. Isis, 92(1):165–166, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237375.

Lund:2001:BRB

[482] Eric Lund. Book review: Valentin Weigel (1533–1588): German Religious Dissenter, Speculative Theorist, and Advocate of Tolerance by Andrew Weeks. Isis, 92(1):166–167, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237376.

Fraser:2001:BRB

[483] Craig Fraser. Book review: Die Werke von Jakob Bernoulli: Die Differentialgeometrie by Jakob Bernoulli; André Weil; Martin Mattmüller. Isis, 92(1):167–168, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237377.

Pourciau:2001:BRB

[484] Bruce Pourciau. Book review: Reading the Principia: The Debate on Newton’s Mathematical Philosophy from 1687 to 1736 by Niccolò Guicciardini. Isis, 92(1):168–169, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237378.

Christianson:2001:BRB

[485] J. R. Christianson. Book review: Johannes Kepler and the New Astronomy by James R. Voelkel. Isis, 92(1):169–170, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237379.

Shklar:2001:BRB

[486] Gerald Shklar. Book review: The Making of the “Dentiste,” c. 1650–1760 by Roger King. Isis, 92(1):170–171, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237380.

Willson:2001:BRB

[487] Robert F. Willson, Jr. Book review: Bodies and Selves in Early Modern England: Physiology and Inwardness in Spenser, Shakespeare, Herbert, and Milton by Michael C. Schoenfeldt. Isis, 92(1):171–172, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237381.

Hall:2001:BRB

[488] Michael G. Hall. Book review: A Fever in Salem: A New Interpretation of the New England Witch Trials by Laurie Winn Carlson. Isis, 92(1):

172–173, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237382.

Allen:2001:BRBb

[489] Peter Lewis Allen. Book review: Surgery, Skin, and Syphilis: Daniel Turner’s London (1667–1741) by Philip K. Wilson. Isis, 92(1):173–174, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237383.

Ariew:2001:BRB

[490] Roger Ariew. Book review: Descartes: Écrits physiologiques et médicaux by Vincent Aucante. Isis, 92(1):174, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237384.

Gaukroger:2001:BRB

[491] Stephen Gaukroger. Book review: Descartes: Philosophical Essays and Correspondence by René Descartes. Isis, 92(1):175, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237385.

Shapiro:2001:BRBb

[492] Barbara Shapiro. Book review: Science, Religion, and Politics in Restoration England: Richard Cumberland’s De legibus naturae by Jon Parkin. Isis, 92(1):175–176, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237386.

Rappaport:2001:BRB

[493] Rhoda Rappaport. Book review: Cosmo e apocalisse: Teorie del millennio e storia della terra nell ’inghilterra del seicento by Amalia Bettini. Isis, 92(1):176–177, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237387.

Fara:2001:BRB

[494] Patricia Fara. Book review: Représentations scientifiques et images poétiques en Angleterre au XVIIe siècle: À la recherche de l’invisible by Margaret Llasera. Isis, 92(1):177, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237388.

Findlen:2001:BRB

[495] Paula Findlen. Book review: Un intellettuale europeo e il suo universo: Vincenzo Coronelli (1650–1718) by Maria Gioia Tavoni. Isis, 92(1):177– 179, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237389.

Jacob:2001:BRB

[496] Margaret C. Jacob. Book review: Ingenious Pursuits: Building the Scientific Revolution by Lisa Jardine. Isis, 92(1):179–180, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237390.

Cover:2001:BRB

[497] J. A. Cover. Book review: Leibniz’s Science of the Rational by Emily Grosholz; Elhanan Yakira. Isis, 92(1):180–181, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237391.

Guedj:2001:BRB

[498] Denis Guedj. Book review: Medir el metro: La historia de la prolongación del arco de meridiano Dunkerque–Barcelona, base del sistema métrico decimal by Antonio E. Ten. Isis, 92(1):181, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237392.

Dijksterhuis:2001:BRB

[499] Fokko Jan Dijksterhuis. Book review: Van “Konstgenoten” en hemelse fenomenen: Nederlandse sterrenkunde in de achttiende eeuw by Huibert Jan Zuidervaart. Isis, 92(1):181–182, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237393.

Daum:2001:BRB

[500] Andreas W. Daum. Book review: Unterwegs mit wilden Tieren: Wandermenagerien zwischen Belehrung und Kommerz, 1750–1850 by Annelore Rieke–Müller; Lothar Dittrich. Isis, 92(1):182–183, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237394.

Schlote:2001:BRB

[501] Karl-Heinz Schlote. Book review: Bernard Bolzano: Miscellanea Mathematica 14 by Bernard Bolzano; Bob van Rootselaar; Anna van der Lugt. Isis, 92(1):183–184, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237395.

Larson:2001:BRBa

[502] Bruce Larson. Book review: Économie politique et économie naturelle chez Antoine-Augustin Cournot by François Vatin. Isis, 92(1):184, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237396.

Sheynin:2001:BRB

[503] Oscar Sheynin. Book review: The Politics of Large Numbers: A History of Statistical Reasoning by Alain Desrosières; Camille Naish. Isis, 92(1): 184–185, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237397.

Elzinga:2001:BRB

[504] Aant Elzinga. Book review: Vetenskapens vikingatåg: Perspektiv på svensk polarforskning, 1860–1930 by Urban Wråkberg. Isis, 92(1):185– 186, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237398.

Lagueux:2001:BRB

[505] Olivier Lagueux. Book review: Zarafa: A Giraffe’s True Story, from Deep in Africa to the Heart of Paris by Michael Allin. Isis, 92(1):186– 187, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237399.

Nicolson:2001:BRB

[506] Malcolm Nicolson. Book review: The European Origins of Scientific Ecology (1800–1901) by Pascal Acot; B. P. Hamm. Isis, 92(1):187–188, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237400.

Gregory:2001:BRB

[507] Frederick Gregory. Book review: The Life of Goethe: A Critical Biography by John R. Williams. Isis, 92(1):188–189, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237401.

Wilson:2001:BRBa

[508] Eric Wilson. Book review: Romantic Science and the Experience of the Self: Transatlantic Crosscurrents from William James to Oliver Sacks by Martin Halliwell. Isis, 92(1):189–190, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237402.

Nyhart:2001:BRB

[509] Lynn K. Nyhart. Book review: Nature’s Museums: Victorian Science and the Architecture of Display by Carla Yanni. Isis, 92(1):190–191, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237403.

Dean:2001:BRB

[510] Dennis R. Dean. Book review: The Culture of English Geology, 1815– 1851: A Science Revealed through Its Collecting by Simon J. Knell. Isis, 92(1):191, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237404.

Cassidy:2001:BRB

[511] James G. Cassidy. Book review: Surveying the Record: North American Scientific Exploration to 1930 by Edward C. Carter. Isis, 92(1):191– 193, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237405.

Gollard:2001:BRB

[512] Gareth Gollard. Book review: Proust, the Body, and Literary Form by Michael R. Finn. Isis, 92(1):193–194, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237406.

Barrow:2001:BRB

[513] Mark V. Barrow, Jr. Book review: The Emperor of Nature: Charles-Lucien Bonaparte and His World by Patricia Tyson Stroud. Isis, 92(1): 194–195, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237407.

Wellman:2001:BRB

[514] Kathleen Wellman. Book review: Constructing Paris Medicine by Caroline Hannaway; Ann La Berge. Isis, 92(1):195–196, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237408.

Lella:2001:BRB

[515] Joseph W. Lella. Book review: William Osler: A Life in Medicine by Michael Bliss. Isis, 92(1):196–197, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237409.

Lynch:2001:BRB

[516] John M. Lynch. Book review: Darwinism and the Linguistic Image: Language, Race, and Natural Theology in the Nineteenth Century by Stephen G. Alter. Isis, 92(1):197, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237410.

Gambino:2001:BRB

[517] Matthew Gambino. Book review: Insanity, Institutions, and Society, 1800–1914: A Social History of Madness in Comparative Perspective by Joseph Melling; Bill Forsythe. Isis, 92(1):197–199, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237411.

McClellan:2001:BRB

[518] Chris McClellan. Book review: Synthèse subjective; ou, Système universel des conceptions propres á l’état normal de l’humanité by Auguste Comte; Juliette Grange. Isis, 92(1):199–200, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237412.

Darnell:2001:BRB

[519] Regna Darnell. Book review: Thinking with Both Hands: Sir Daniel Wilson in the Old World and the New by Marinell Ash; Elizabeth Hulse. Isis, 92(1):200–201, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237413.

Hansen:2001:BRB

[520] LeeAnn Hansen. Book review: Rahel Levin Varnhagen: The Life and Work of a German Jewish Intellectual by Heidi Thomann Tewarson. Isis, 92(1):201, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237414.

Dror:2001:BRB

[521] Otniel E. Dror. Book review: Suspensions of Perception: Attention, Spectacle, and Modern Culture by Jonathan Crary. Isis, 92(1):201–203, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237415.

Dawson:2001:BRB

[522] John Dawson. Book review: Gödel Meets Einstein: Time Travel in the Gödel Universe by Palle Yourgrau. Isis, 92(1):203, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237416.

Grier:2001:BRB

[523] David Alan Grier. Book review: Inventing the Internet by Janet Abbate. Isis, 92(1):203–204, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237417.

Pang:2001:BRB

[524] Alex Soojung-Kim Pang. Book review: On Line and on Paper: Visual Representations, Visual Culture, and Computer Graphics in Design Engineering by Kathryn Henderson. Isis, 92(1):204–205, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237418.

Johnson:2001:BRBa

[525] Karen E. Johnson. Book review: Wiederaufbau und Kernenergie: Zur Haltung deutscher Emigranten in Amerika by Dirk Lehrach. Isis, 92(1): 205–206, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237419.

Goodstein:2001:BRB

[526] Judith Goodstein. Book review: Twentieth Century Physics: Essays and Recollections: A Selection of Historical Writings by Edoardo Amaldi by Giovanni Battimelli; Giovanni Paoloni. Isis, 92(1):206–207, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237420.

Kaiser:2001:BRBa

[527] David Kaiser. Book review: Richard Feynman: A Life in Science by John Gribbin; Mary Gribbin. Isis, 92(1):207–208, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237421.

Seidel:2001:BRBa

[528] Robert W. Seidel. Book review: Historical Encyclopedia of Atomic Energy by Stephen E. Atkins. Isis, 92(1):208–209, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237422.

MacKinnon:2001:BRBa

[529] Edward MacKinnon. Book review: Interpreting Bodies: Classical and Quantum Objects in Modern Physics by Elena Castellani. Isis, 92(1): 209–210, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237423.

Kragh:2001:BRB

[530] Helge Kragh. Book review: Quantum Dialogue: The Making of a Revolution by Mara Beller. Isis, 92(1):210–211, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237424.

Fitzpatrick:2001:BRB

[531] Anne Fitzpatrick. Book review: Soviet Scientists in the Nuclear Age: A Memoir by Veniamin Tsukerman; Zinaida Azarkh; Michael Pursglove; Timothy Sergey. Isis, 92(1):211–212, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237425.

Wittner:2001:BRB

[532] Lawrence S. Wittner. Book review: Silencing the Opposition: Antinuclear Movements and the Media in the Cold War by Andrew Rojecki. Isis, 92(1):212, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237426.

Dahl:2001:BRB

[533] Per F. Dahl. Book review: The Radiance of France: Nuclear Power and National Identity after World War II by Gabrielle Hecht. Isis, 92(1):212– 213, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237427.

Kumar:2001:BRB

[534] Deepak Kumar. Book review: The Making of the Indian Atomic Bomb: Science, Secrecy, and the Postcolonial State by Itty Abraham. Isis, 92 (1):213–214, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237428.

Schorn:2001:BRB

[535] Ronald A. Schorn. Book review: Planetary Dreams: The Quest to Discover Life beyond Earth by Robert Shapiro. Isis, 92(1):214–215, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237429.

Sassower:2001:BRB

[536] Raphael Sassower. Book review: The Intellectual Appropriation of Technology: Discourses on Modernity, 1900–1939 by Mikael Hård; Andrew Jamison. Isis, 92(1):215–216, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237430.

Kline:2001:BRBa

[537] Ronald Kline. Book review: Power Loss: The Origins of Deregulation and Restructuring in the American Electric Utility Industry by Richard F. Hirsh. Isis, 92(1):216–217, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237431.

Stranges:2001:BRB

[538] Anthony N. Stranges. Book review: Space to Earth: The Story of Solar Electricity by John Perlin. Isis, 92(1):217–218, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237432.

Gitelman:2001:BRB

[539] Lisa Gitelman. Book review: Off the Record: The Technology and Culture of Sound Recording in America by David Morton. Isis, 92(1):218– 219, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237433.

Rainger:2001:BRB

[540] Ronald Rainger. Book review: Science and the Pacific War: Science and Survival in the Pacific, 1939–1945 by Roy M. MacLeod. Isis, 92(1):219– 220, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237434.

Deacon:2001:BRBa

[541] Margaret Deacon. Book review: Of Ships and Stars: Maritime Heritage and the Founding of the National Maritime Museum, Greenwich by Kevin Littlewood; Beverley Butler. Isis, 92(1):220–221, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237435.

Hestmark:2001:BRB

[542] Geir Hestmark. Book review: By Airship to the North Pole: An Archaeology of Human Exploration by P. J. Capelotti. Isis, 92(1):221–222, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237436.

Sapolsky:2001:BRB

[543] Harvey M. Sapolsky. Book review: Sidewinder: Creative Missile Development at China Lake by Ron Westrum. Isis, 92(1):222–223, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237437.

Westwick:2001:BRB

[544] Peter J. Westwick. Book review: The Gaither Committee, Eisenhower, and the Cold War by David L. Snead. Isis, 92(1):223–224, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237438.

Chubin:2001:BRB

[545] Daryl E. Chubin. Book review: Between Politics and Science: Assuring the Integrity and Productivity of Research by David H. Guston. Isis, 92(1):224–225, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237439.

Hurt:2001:BRB

[546] R. Douglas Hurt. Book review: Dream Reaper: The Story of an Old-Fashioned Inventor in the High-Tech, High-Stakes World of Modern Agriculture. 1995 by Craig Canine. Isis, 92(1):225–226, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237440.

Miller:2001:BRB

[547] Jane Miller. Book review: The Chemists’ Club: One Hundred Years in the Chemical Community. Isis, 92(1):226–227, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237441.

Cesarani:2001:BRB

[548] David Cesarani. Book review: The Nazi War on Cancer by Robert N. Proctor. Isis, 92(1):227, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237442.

Bastos:2001:BRB

[549] Cristiana Bastos. Book review: Blood Saga: Hemophilia, AIDS, and the Survival of a Community by Susan Resnik. Isis, 92(1):227–228, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237443.

Morman:2001:BRB

[550] Edward T. Morman. Book review: No One Was Turned Away: The Role of Public Hospitals in New York City since 1900 by Sandra Opdycke. Isis, 92(1):228–229, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237444.

Burnham:2001:BRB

[551] John C. Burnham. Book review: Medicating Schizophrenia: A History by Sheldon Gelman. Isis, 92(1):229–230, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237445.

Richards:2001:BRB

[552] Graham Richards. Book review: Knowledge and Power: Perspectives in the History of Psychiatry by Eric J. Engstrom; Matthias M. Weber; Paul Hoff. Isis, 92(1):230–231, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237446.

Parascandola:2001:BRB

[553] John Parascandola. Book review: Pharmaceutical Innovation: Revolutionizing Human Health by Ralph Landau; Basil Achilladelis; Alexander Scriabine. Isis, 92(1):231–232, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237447.

Dijstelberge:2001:BRB

[554] Paul Dijstelberge. Book review: Out of Otherness: Characters and Narrators in the Dutch Venereal Disease Debates, 1850–1990 by Annet Mooij; Beverley Jackson. Isis, 92(1):232–233, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237448.

Gaudilliere:2001:BRB

[555] Jean-Paul Gaudillière. Book review: The Generation of Diversity: Clonal Election Theory and the Rise of Molecular Immunology by Scott H. Podolsky; Alfred I. Tauber. Isis, 92(1):233–234, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237449.

Winnewisser:2001:BRB

[556] Brenda P. Winnewisser. Book review: Elsa Neumann: Berlins erstes Fräulein Doktor by Annette Vogt. Isis, 92(1):234–235, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237450.

Farber:2001:BRB

[557] Paul Lawrence Farber. Book review: Flight Maps: Adventures with Nature in Modern America by Jennifer Price. Isis, 92(1):235, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237451.

Burlingame:2001:BRB

[558] Leslie J. Burlingame. Book review: Life in the Treetops: Adventures of a Woman in Field Biology by Margaret D. Lowman. Isis, 92(1):235– 236, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237452.

Firebaugh:2001:BRB

[559] Francille M. Firebaugh. Book review: The Academic Kitchen: A Social History of Gender Stratification at the University of California, Berkeley by Maresi Nerad. Isis, 92(1):236–237, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237453.

Yochelson:2001:BRBa

[560] Ellis L. Yochelson. Book review: The Great Dinosaur Extinction Controversy by Charles Officer; Jake Page. Isis, 92(1):237–238, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237454.

Westfahl:2001:BRB

[561] Gary Westfahl. Book review: Science Is Fiction: The Films of Jean Painlevé by Andy Masaki Bellows; Marina McDougall; Brigitte Berg. Isis, 92(1):238–239, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237455.

Benson:2001:BRB

[562] Maxine Benson. Book review: National Parks and the Woman’s Voice: A History by Polly Welts Kaufman; Women and Nature: Saving the “Wild” West by Glenda Riley. Isis, 92(1):239–240, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237456.

Mark:2001:BRB

[563] Joan Mark. Book review: Collecting Native America, 1870–1960 by Shepard Krech; Barbara A. Hail. Isis, 92(1):240–241, March 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237457.

Usner:2001:BRB

[564] Daniel H. Usner, Jr. Book review: The Allegany Senecas and Kinzua Dam: Forced Relocation through Two Generations by Joy Ann Bilharz. Isis, 92(1):241–242, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237458.

Buhle:2001:BRB

[565] Mari Jo Buhle. Book review: The Unknown Karen Horney: Essays on Gender, Culture, and Psychoanalysis by Bernard J. Paris. Isis, 92(1): 242–243, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237459.

Lovett:2001:BRB

[566] Laura Lovett. Book review: Inheriting Shame: The Story of Eugenics and Racism in America by Steven Selden. Isis, 92(1):243–244, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237460.

Olson:2001:BRBa

[567] Richard Olson. Book review: The Rise of Social Theory by Johan Heilbron; Sheila Gogol. Isis, 92(1):244–245, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237461.

Hackett:2001:BRB

[568] Jeremiah Hackett. Book review: Between Demonstration and Imagination: Essays in the History of Science and Philosophy by Lodi Nauta; Arjo Vanderjagt. Isis, 92(1):245–246, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237462.

Brown:2001:BRBa

[569] Janet Welsh Brown. Book review: Has Feminism Changed Science? by Londa Schiebinger. Isis, 92(1):246–247, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237463.

Philip:2001:BRB

[570] Kavita Philip. Book review: Toward a Global Science: Mining Civilizational Knowledge by Susantha Goonatilake. Isis, 92(1):247–248, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237464.

Ogilvie:2001:BRB

[571] Marilyn Bailey Ogilvie. Book review: American Women in Science: 1950 to the Present: A Biographical Dictionary by Martha J. Bailey. Isis, 92(1):249, March 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237465.

Anonymous:2001:Ca

[572] Anonymous. Collections. Isis, 92(1):249–254, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237466.

Anonymous:2001:BMb

[573] Anonymous. Back matter. Isis, 92(1):255–264, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237467.

Anonymous:2001:FMb

[574] Anonymous. Front matter. Isis, 92(1):??, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237324.

Harrison:2001:CFK

[575] Peter Harrison. Curiosity, forbidden knowledge, and the reformation of natural philosophy in early modern England. Isis, 92(2):265–290, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080629.

Rasmussen:2001:PHW

[576] Nicolas Rasmussen. Plant hormones in war and peace: Science, industry, and government in the development of herbicides in 1940s America. Isis, 92(2):291–316, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080630.

Knight:2001:CCR

[577] Mary Knight. Curing cut or ritual mutilation?: Some remarks on the practice of female and male circumcision in graeco-roman Egypt. Isis, 92(2):317–338, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080631.

Mendell:2001:EWK

[578] Henry R. Mendell. Eloge: Wilbur Knorr, 29 August 1945–18 March 1997. Isis, 92(2):339–343, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080632.

Anonymous:2001:NSA

[579] Anonymous. News of the Society: Annual meeting of the History of Science Society, Vancouver, 2–5 November 2000. Isis, 92(2):344–350, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080633.

Osler:2001:BRB

[580] Margaret J. Osler. Book review: The Works of Robert Boyle. Isis, 92(2): 351–353, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080634.

Smith:2001:BRB

[581] Crosbie Smith. Book review: The Architecture of Science by Peter Galison; Emily Thompson. Isis, 92(2):354–355, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080635.

Westman:2001:BRB

[582] Robert S. Westman. Book review: Le monde des sphères. Volume 1: Genèse et triomphe d’une représentation cosmique by Michel-Pierre Lerner; Le monde des sphères. Volume 2: La fin du cosmos classique by Michel-Pierre Lerner. Isis, 92(2):355–357, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080636.

King:2001:BRB

[583] David A. King. Book review: Osmanli astronomi literatürü tarihi (History of Astronomy Literature during the Ottoman Period) by Ekmeleddin Ihsanoglu; Osmanli matematik literatürü tarihi (History of Mathematics Literature during the Ottoman Period) by Ekmeleddin Ilhsanoglu. Isis, 92(2):357–359, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080637.

Thiele:2001:BRB

[584] Rüdiger Thiele. Book review: Number: From Ahmes to Cantor by Midhat Gazalé. Isis, 92(2):359–360, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080638.

Puig:2001:BRB

[585] Roser Puig. Book review: World-Maps for Finding the Direction and Distance to Mecca: Innovation and Tradition in Islamic Science by David A. King. Isis, 92(2):360–362, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080639.

Malet:2001:BRB

[586] Antoni Malet. Book review: Bibliographia physico-mathematica hispanica (1475–1900). Volume 1: Libros y folletos, 1475–1600 by Víctor Navarro Brotóns; Vicente L. Salavert Fabiani; Victòria Rosselló Botey; Víctor Darás Román. Isis, 92(2):362–363, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080640.

Stieb:2001:BRB

[587] Ernst W. Stieb. Book review: Über die Anfänge der Pharmaziegeschichtsschreibung: Von Johannes Ruellius (1529) bis David Peter Hermann Schmidt (1835) by Evemarie Wolf. Isis, 92(2):363–364, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080641.

Leedham-Green:2001:BRB

[588] E. S. Leedham-Green. Book review: Sir Thomas Gresham and Gresham College: Studies in the Intellectual History of London in the Sixteenth and Seventeenth Centuries by Francis Ames-Lewis. Isis, 92(2):364, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080642.

Olson:2001:BRBb

[589] Richard Olson. Book review: Adam Smith across Nations: Translations and Receptions of the Wealth of Nations by Cheng-Chung Lai. Isis, 92(2): 364–365, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080643.

Sigurdsson:2001:BRB

[590] Haraldur Sigurdsson. Book review: Fire in the Sea: The Santorini Volcano: Natural History and the Legend of Atlantis by Walter L. Friedrich; Alexander R. McBirney. Isis, 92(2):365–366, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080644.

Secord:2001:BRB

[591] James A. Secord. Book review: The Lamp of Learning: Two Centuries of Publishing at Taylor & Francis by W. H. Brock; A. J. Meadows. Isis, 92(2):366–367, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080645.

Cook:2001:BRB

[592] Harold J. Cook. Book review: A History of Science in The Netherlands: Survey, Themes, and Reference by Klaas van Berkel; Albert Van Helden; Lodewijk Palm. Isis, 92(2):367–369, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080646.

Dusek:2001:BRB

[593] Val Dusek. Book review: Physicists in Conflict by Neil A. Porter. Isis, 92(2):369–370, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080647.

Warner:2001:BRBa

[594] Deborah Jean Warner. Book review: The Quick and the Dead: Artists and Anatomy by Deanna Petherbridge; Ludmilla Jordanova. Isis, 92(2): 370, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080648.

Bruce:2001:BRB

[595] Robert V. Bruce. Book review: The Establishment of Science in America: 150 Years of the American Association for the Advancement of Science by Sally Gregory Kohlstedt; Michael M. Sokal; Bruce V. Lewenstein. Isis, 92(2):370–372, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080649.

Berendzen:2001:BRB

[596] Richard Berendzen. Book review: The “Astrophysical Journal”, American Astronomical Society Centennial Issue: Selected Fundamental Papers Published This Century in the “American Astrophysical Journal” and the “Astronomical Journal” by Helmut A. Abt. Isis, 92(2):372–373, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080650.

Worboys:2001:BRB

[597] Michael Worboys. Book review: Beriberi, White Rice, and Vitamin B: A Disease, a Cause, and a Cure by Kenneth J. Carpenter. Isis, 92(2): 373, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080651.

Ludmerer:2001:BRB

[598] Kenneth M. Ludmerer. Book review: American Medicine and the Public Interest: A History of Specialization by Rosemary Stevens. Isis, 92(2): 373–374, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080652.

Wolf:2001:BRB

[599] Jacqueline H. Wolf. Book review: Nurturing Children: A History of Pediatrics by A. R. Colón. Isis, 92(2):374–375, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080653.

Stoll:2001:BRBb

[600] Mark Stoll. Book review: Reconstructing Nature: The Engagement of Science and Religion by John Brooke; Geoffrey Cantor. Isis, 92(2):375– 376, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080654.

Fleming:2001:BRB

[601] James Rodger Fleming. Book review: Greenhouse: The Two-Hundred-Year Story of Global Warming by Gale E. Christianson. Isis, 92(2):376– 377, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080655.

Anderson:2001:BRB

[602] Katharine Anderson. Book review: The Pearly Gates of Cyberspace: A History of Space from Dante to the Internet by Margaret Wertheim. Isis, 92(2):377–378, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080656.

Jasanoff:2001:BRB

[603] Sheila Jasanoff. Book review: Legal Alchemy: The Use and Misuse of Science in the Law by David L. Faigman. Isis, 92(2):378–379, June 2001.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080657.

Rudge:2001:BRB

[604] David Wÿss Rudge. Book review: Unto Others: The Evolution and Psychology of Unselfish Behavior by Elliott Sober; David Sloan Wilson. Isis, 92(2):379–380, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080658.

Brooke:2001:BRB

[605] John Hedley Brooke. Book review: Stages of Thought: The Co-Evolution of Religious Thought and Science by Michael Horace Barnes. Isis, 92(2): 380–381, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080659.

Muroi:2001:BRB

[606] K. Muroi. Book review: Mesopotamian Mathematics, 2100–1600 B.C.: Technical Constants in Bureaucracy and Education by Eleanor Robson. Isis, 92(2):381, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080660.

McGinnis:2001:BRB

[607] Jon McGinnis. Book review: Aristoteles’ “De Anima”: Eine verlorene spätantike Paraphrase in arabischer und persischer Überlieferung by Rüdiger Arnzen. Isis, 92(2):381–382, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080661.

Gordon:2001:BRBb

[608] Robert B. Gordon. Book review: King Croesus’ Gold: Excavations at Sardis and the History of Gold Refining by Andrew Ramage; Paul Craddock. Isis, 92(2):382–383, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080662.

Cosans:2001:BRB

[609] Christopher E. Cosans. Book review: Text and Tradition: Studies in Ancient Medicine and Its Transmission by Klaus-Metrich Fischer; Diethard Nickel; Paul Potter. Isis, 92(2):383–384, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080663.

Ongaro:2001:BRB

[610] Giuseppe Ongaro. Book review: Areteo di Cappadocia: Interpretazioni e aspetti della formazione anatomo-patologica del Morgagni by Giorgio Weber. Isis, 92(2):384–385, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080664.

Campion:2001:BRB

[611] Nick Campion. Book review: Conversations with Angels: Cabala, Alchemy, and the End of Nature by Deborah E. Harkness. Isis, 92(2): 385, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080665.

Webster:2001:BRB

[612] Charles Webster. Book review: Natural Particulars: Nature and the Disciplines in Renaissance Europe by Anthony Grafton; Nancy Siraisi. Isis, 92(2):385–386, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080666.

Jesseph:2001:BRB

[613] Douglas M. Jesseph. Book review: La Contre–Réforme mathématique: Constitution et diffusion d’une culture mathématique Jésuite à la Renaissance (1540–1640) by Antonella Romano. Isis, 92(2):386–387, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080667.

White:2001:BRB

[614] Kevin White. Book review: La météorologie en français au Moyen Âge (XIIIe-XIVe siècles) by Joëlle Ducos. Isis, 92(2):387–388, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080668.

Thussen:2001:BRB

[615] J. M. M. H. Thussen. Book review: Theories of Cognition in the Later Middle Ages by Robert Pasnau. Isis, 92(2):388–390, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080669.

Givens:2001:BRBa

[616] Jean A. Givens. Book review: The Great Herbal of Leonhart Fuchs: “De historia stirpium commentarii insignes”, 1542 by Frederick G. Meyer; Emily Emmart Trueblood; John L. Heller. Isis, 92(2):390–391, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080670.

Schleiner:2001:BRB

[617] Winfried Schleiner. Book review: Geheimnisse der Alchemie by Manuel Bachmann; Thomas Hofmeier; Alchemie: Lexikon einer hermetischen Wissenschaft by Claus Priesner; Karin Figala. Isis, 92(2):391–392, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080671.

Lodge:2001:BRB

[618] Paul Lodge. Book review: Leibniz and the “Vinculum Substantiale” by Brandon Look. Isis, 92(2):392–393, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080672.

Fournier:2001:BRB

[619] Marian Fournier. Book review: Augustanus Opticus: Johann Wiesel (1583–1662) und 200 Jahre optisches Handwerk in Augsburg by Inge Keil. Isis, 92(2):393–394, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080673.

Gentilcore:2001:BRB

[620] David Gentilcore. Book review: The Clinical Consultations of Francesco Torti by Saul Jarcho. Isis, 92(2):394–395, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080674.

Barnaby:2001:BRB

[621] Andrew Barnaby. Book review: A Thinker for All Seasons: Sir Francis Bacon and His Significance Today by David Burnett. Isis, 92(2): 395, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080675.

Solomon:2001:BRB

[622] Julie Robin Solomon. Book review: Francis Bacon by Perez Zagorin. Isis, 92(2):395–398, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080676.

Zagorin:2001:BRB

[623] Perez Zagorin. Book review: Hostage to Fortune: The Troubled Life of Francis Bacon by Lisa Jardine; Alan Stewart. Isis, 92(2):398–399, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080677.

Wilson:2001:BRBb

[624] Fred Wilson. Book review: Hume by Anthony Quinton. Isis, 92(2):399– 400, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080678.

Lafuente:2001:BRB

[625] Antonio Lafuente. Book review: Navegación astronómica en la España del siglo XVIII by Manuel A. Sellés García. Isis, 92(2):400, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080679.

Gillispie:2001:BRB

[626] Charles C. Gillispie. Book review: Geography Unbound: French Geographic Science from Cassini to Humboldt by Anne Marie Claire Godlewska. Isis, 92(2):400–402, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080680.

Gohau:2001:BRB

[627] Gabriel Gohau. Book review: Louis Jurine: Chirurgien et naturaliste (1751–1819) by René Sigrist; Vincent Barras; Marc Ratcliff. Isis, 92(2): 402–403, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080681.

Gordon:2001:BRBc

[628] Robert B. Gordon. Book review: Making Iron on the Bald Eagle: Roland Curtin’s Ironworks and Workers’ Community by Gerald G. Eggert. Isis, 92(2):403–404, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080682.

Bishop:2001:BRB

[629] Paul Bishop. Book review: Goethes Naturmodell: Die Scala Naturae and ihre Transformationen by Margrit Wyder. Isis, 92(2):404, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080683.

Kim:2001:BRBa

[630] Mi Gyung Kim. Book review: Arthur Rudolf Hantzsch: Im Briefwechsel mit Wihelm Ostwald by Joachim Stocklöv. Isis, 92(2):404–405, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080684.

Cercignani:2001:BRB

[631] Carlo Cercignani. Book review: Pioneering Ideas for the Physical and Chemical Sciences: Josef Loschmidt’s Contributions and Modern Developments in Structural Organic Chemistry, Atomistics, and Statistical Mechanics by W. Fleischhacker; T. Schönfeld. Isis, 92(2):405–406, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080685.

Breeden:2001:BRB

[632] James O. Breeden. Book review: Science, Race, and Religion in the American South: John Bachman and the Charleston Circle of Naturalists, 1815–1895 by Lester D. Stephens. Isis, 92(2):406–407, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080686.

Walls:2001:BRBa

[633] Laura Dassow Walls. Book review: Romantic Turbulence: Chaos, Ecology, and American Space by Eric Wilson. Isis, 92(2):407–408, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080687.

Champlin:2001:BRB

[634] Peggy Champlin. Book review: Exploring the Highest Sierra by James G. Moore. Isis, 92(2):408–409, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080688.

Weindling:2001:BRB

[635] Paul Weindling. Book review: Krankheit und Kulturkritik: Psychiatrische Gesellschaftsdeutungen im bürgerlichen Zeitalter (1790–1914) by Volker Roelcke. Isis, 92(2):409, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080689.

Gispen:2001:BRB

[636] Kees Gispen. Book review: Der Optimismus der Ingenieure: Triumph der Technik in der Krise der Moderne um 1900 by Hans-Liudger Dienel. Isis, 92(2):409–410, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080690.

Baatz:2001:BRB

[637] Simon Baatz. Book review: The Gospel of Germs: Men, Women, and the Microbe in American Life by Nancy Tomes. Isis, 92(2):410–411, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080691.

Decker:2001:BRB

[638] Hannah S. Decker. Book review: Otto Weininger: Sex, Science, and Self in Imperial Vienna by Chandak Sengoopta. Isis, 92(2):411–412, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080692.

Fairman:2001:BRB

[639] Julie Fairman. Book review: Florence Nightingale: Avenging Angel by Hugh Small. Isis, 92(2):412–413, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080693.

Risse:2001:BRB

[640] Guenter B. Risse. Book review: City of Plagues: Disease, Poverty, and Deviance in San Francisco by Susan Craddock. Isis, 92(2):413–414, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080694.

Shapiro-Shapin:2001:BRB

[641] Carolyn G. Shapiro-Shapin. Book review: “A Time to Heal”: The Diffusion of Listerism in Victorian Britain by Jerry L. Gaw. Isis, 92(2): 414–415, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080695.

Corry:2001:BRB

[642] Leo Corry. Book review: Mathematiker auf der Flucht vor Hitler: Quellen and Studien zur Emigration einer Wissenschaft by Reinhard Siegmund-Schultze. Isis, 92(2):415–416, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080696.

Clark:2001:BRBa

[643] Eloise Clark. Book review: Funding Science in America: Congress, Universities, and the Politics of the Academic Pork Barrel by James D. Savage. Isis, 92(2):416–417, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080697.

Bocking:2001:BRB

[644] Stephen Bocking. Book review: Life Stories: World-Renowned Scientists Reflect on Their Lives and the Future of Life on Earth by Heather Newbold. Isis, 92(2):417–418, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080698.

Raina:2001:BRB

[645] Dhruv Raina. Book review: Jagadis Chandra Bose and the Indian Response to Western Science by Subrata Dasgupta. Isis, 92(2):418–419, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080699.

Hacker:2001:BRB

[646] Barton C. Hacker. Book review: A Radar History of World War II: Technical and Military Imperatives by Louis Brown. Isis, 92(2):419– 420, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080700.

Canaday:2001:BRB

[647] John Canaday. Book review: Atomic Fragments: A Daughter’s Questions by Mary Palevsky. Isis, 92(2):420–421, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080701.

Crosby:2001:BRB

[648] Donald A. Crosby. Book review: Einstein and Religion: Physics and Theology by Max Jammer. Isis, 92(2):421–422, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080702.

Seidel:2001:BRBb

[649] Robert W. Seidel. Book review: In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist by S. S. Schweber. Isis, 92(2):422–423, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080703.

Abraham:2001:BRB

[650] Itty Abraham. Book review: India’s Nuclear Bomb: The Impact on Global Proliferation by George Perkovich. Isis, 92(2):423–424, June 2001.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080704.

Morton:2001:BRB

[651] David Morton. Book review: History of International Broadcasting. Volume 1 by James Wood; History of International Broadcasting. Volume 2 by James Wood. Isis, 92(2):424–425, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080705.

Perkins:2001:BRBa

[652] John Perkins. Book review: The United States and Biological Warfare: Secrets from the Early Cold War and Korea by Stephen Endicott; Edward Hagerman. Isis, 92(2):425–426, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080706.

Hashimoto:2001:BRB

[653] Takehiko Hashimoto. Book review: Tiger Technology: The Creation of a Semiconductor Industry in East Asia by John A. Mathews; Dong-sung Cho. Isis, 92(2):426–427, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080707.

Clark:2001:BRBb

[654] Mark Henry Clark. Book review: City of Light: The Story of Fiber Optics by Jeff Hecht. Isis, 92(2):427–428, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080708.

Gifford:2001:BRB

[655] Sanford Gifford. Book review: Mind Games: American Culture and the Birth of Psychotherapy by Eric Caplan. Isis, 92(2):428–429, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080709.

Purdy:2001:BRB

[656] Laura Purdy. Book review: From Chance to Choice: Genetics and Justice by Allen Buchanan; Dan W. Brock; Norman Daniels; Daniel Wikler. Isis, 92(2):429–430, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080710.

Moser:2001:BRB

[657] Charles Moser. Book review: The Mastery of Submission: Inventions of Masochism by John K. Noyes. Isis, 92(2):430–431, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080711.

Caton:2001:BRBa

[658] Donald Caton. Book review: Pain and Suffering by William K. Livingston; Howard L. Fields. Isis, 92(2):431, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080712.

Chris:2001:BRB

[659] Cynthia Chris. Book review: Reel Nature: America’s Romance with Wildlife on Film by Gregg Mitman. Isis, 92(2):431–432, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080713.

Isern:2001:BRB

[660] Thomas D. Isern. Book review: William I. Myers and the Modernization of American Agriculture by Douglas Slaybaugh. Isis, 92(2):432–433, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080714.

Kohlstedt:2001:BRB

[661] Sally Gregory Kohlstedt. Book review: Across the Great Border Fault: The Naturalist Myth in America by Kevin Dann. Isis, 92(2):433–434, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080715.

Tjossem:2001:BRB

[662] Sara Tjossem. Book review: Lake Erie Rehabilitated: Controlling Cultural Eutrophication, 1960s-1990s by William McGucken. Isis, 92(2): 435, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080716.

Haard:2001:BRB

[663] Mikael Hård. Book review: Technoscientific Angst: Ethics and Responsibility by Raphael Sassower. Isis, 92(2):435–436, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080717.

Wilson:2001:BRBc

[664] Catherine Wilson. Book review: Thomas Kuhn: A Philosophical History for Our Times by Steve Fuller. Isis, 92(2):436–437, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080718.

Burney:2001:BRB

[665] Ian Burney. Book review: Science and Law by Tal Golan; Snait Gissis. Isis, 92(2):437–438, June 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080719.

Katz:2001:BRB

[666] Steven B. Katz. Book review: Rhetorical Figures in Science by Jeanne Fahnestock. Isis, 92(2):438–439, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080720.

Tiles:2001:BRB

[667] Mary Tiles. Book review: Reason’s Nearest Kin: Philosophies of Arithmetic from Kant to Carnap by Michael Potter. Isis, 92(2):439–440, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080721.

Cushing:2001:BRB

[668] James T. Cushing. Book review: The Philosophy of Physics by Roberto Torretti. Isis, 92(2):440–441, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080722.

Newcomb:2001:BRB

[669] Sally Newcomb. Book review: Encyclopedia of Volcanoes by Haraldur Sigurdsson; Bruce F. Houghton; Stephen R. McNutt; Hazel Rymer; John Stix. Isis, 92(2):441–442, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080723.

Anonymous:2001:Cb

[670] Anonymous. Collections. Isis, 92(2):442–443, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080724.

Anonymous:2001:BMc

[671] Anonymous. Back matter. Isis, 92(2):444–449, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080725.

Anonymous:2001:FMc

[672] Anonymous. Front matter. Isis, 92(2):??, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080628.

Weisz:2001:SMW

[673] George Weisz. Spas, mineral waters, and hydrological science in Twentieth–Century France. Isis, 92(3):451–483, September 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080730.

Pandora:2001:KHC

[674] Katherine Pandora. Knowledge held in common: Tales of Luther Burbank and science in the American vernacular. Isis, 92(3):484–516, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080731.

Krige:2001:DDC

[675] John Krige. Distrust and discovery: The case of the heavy bosons at cern. Isis, 92(3):517–540, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080732.

Hughes:2001:MDD

[676] Sally Smith Hughes. Making dollars out of DNA: The first major patent in biotechnology and the commercialization of molecular biology, 1974– 1980. Isis, 92(3):541–575, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080733.

Rothenberg:2001:BRB

[677] Marc Rothenberg. Book review: Museums of Modern Science by Svante Lindqvist; Marika Hedin; Ulf Larsson. Isis, 92(3):576, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080734.

Gross:2001:BRB

[678] Alan Gross. Book review: Scientific Discourse in Sociohistorical Context: The Philosophical Transactions of the Royal Society of London, 1675– 1975 by Dwight Atkinson. Isis, 92(3):576–577, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080735.

Caneva:2001:BRB

[679] Kenneth L. Caneva. Book review: Scientific Controversies: Philosophical and Historical Perspectives by Peter Machamer; Marcello Pera; Aristides Baltas. Isis, 92(3):577–578, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080736.

Rydell:2001:BRB

[680] Robert W. Rydell. Book review: Commemorative Practices in Science: Historical Perspectives on the Politics of Collective Memory. (Osiris, 2nd Series, 14) by Pnina G. Abir-Am; Clark A. Elliott. Isis, 92(3):578–579, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080737.

Liston:2001:BRB

[681] Michael Liston. Book review: Unifying Scientific Theories: Physical Concepts and Mathematical Structures by Margaret Morrison. Isis, 92 (3):579–580, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080738.

Swerdlow:2001:BRB

[682] N. M. Swerdlow. Book review: Galileo’s Planet: Observing Jupiter before Photography by Thomas A. Hockey. Isis, 92(3):580–581, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080739.

Garber:2001:BRBa

[683] Elizabeth Garber. Book review: The Problem of the Earth’s Shape from Newton to Clairaut: The Rise of Mathematical Science in Eighteenth-Century Paris and the Fall of “Normal” Science by John L. Greenberg. Isis, 92(3):581–582, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080740.

Barrow-Green:2001:BRB

[684] June Barrow-Green. Book review: The Nothing That Is: A Natural History of Zero by Robert Kaplan. Isis, 92(3):582–584, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080741.

Zebrowski:2001:BRBb

[685] Ernest Zebrowski, Jr. Book review: The Universal History of Numbers: From Prehistory to the Invention of the Computer by Georges Ifrah; David Bellos; E. F. Harding; Sophie Wood; Ian Mark. Isis, 92(3):584– 585, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/3080742.

Warner:2001:BRBb

[686] Deborah Jean Warner. Book review: The Universe Unveiled: Instruments and Images through History by Bruce Stephenson; Marvin Bolt; Anna Felicity Friedman. Isis, 92(3):585, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080743.

Evans:2001:BRBb

[687] James Evans. Book review: June 8, 2004: Venus in Transit by Eli Maor. Isis, 92(3):585–586, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080744.

Wilson:2001:BRBd

[688] Leonard G. Wilson. Book review: Sciences of the Earth: Studies in the History of Mineralogy and Geology by David Oldroyd. Isis, 92(3):587, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080745.

Ruse:2001:BRB

[689] Michael Ruse. Book review: Origins of Genius: Darwinian Perspectives on Creativity by Dean Keith Simonton. Isis, 92(3):587–589, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080746.

Culick:2001:BRB

[690] F. E. C. Culick. Book review: A History of Aerodynamics and Its Impact on Flying Machines by John D. Anderson. Isis, 92(3):589–590, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080747.

Agnoletti:2001:BRB

[691] Mauro Agnoletti. Book review: Navi e cantieri della repubblica di Genova: Secoli XVI-XVIII by Luciana Gatti. Isis, 92(3):590–591, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080748.

Pavon:2001:BRB

[692] José M. Cano Pavón. Book review: ‘!Todes a Madrid! La Escuela General Preparatoria de Ingenieros y Arquitectos (1886–1892) by Guillermo Lusa Monforte. Isis, 92(3):591–592, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080749.

deRenzi:2001:BRB

[693] Silvia de Renzi. Book review: Contracting a Cure: Patients, Healers, and the Law in Early Modern Bologna by Gianna Pomata; Rosemarie Foy; Anna Taraboletti-Segre. Isis, 92(3):592–593, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080750.

Rinker:2001:BRB

[694] Sylvia Rinker. Book review: “What a Blessing She Had Chloroform”: The Medical and Social Response to the Pain of Childbirth from 1800 to the Present by Donald Caton. Isis, 92(3):593–594, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080751.

Kim:2001:BRBb

[695] Yung Sik Kim. Book review: A Cultural History of Civil Examinations in Late Imperial China by Benjamin A. Elman. Isis, 92(3):594–595, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080752.

Hales:2001:BRB

[696] Peter Bacon Hales. Book review: The Tainted Desert: Environmental Ruin in the American West by Valerie L. Kuletz. Isis, 92(3):595–596, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080753.

Williams:2001:BRBa

[697] L. Pearce Williams. Book review: The Great Principles of Science by Robert M. Hazen. Isis, 92(3):596–597, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080754.

Oberheim:2001:BRB

[698] Eric Oberheim. Book review: Conquest of Abundance: A Tale of Abstraction versus the Richness of Being by Paul Feyerabend. Isis, 92(3): 597–598, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080755.

Levi:2001:BRB

[699] Joseph Levi. Book review: Jewish Thought and Scientific Discovery in Early Modern Europe by David B. Ruderman. Isis, 92(3):598–599, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080756.

Weldon:2001:BRBa

[700] Stephen P. Weldon. Book review: Belief in God in an Age of Science by John Polkinghorne. Isis, 92(3):599–600, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080757.

Giglioni:2001:BRB

[701] Guido Giglioni. Book review: L’alchimie à la cour de Côme Ier de Médicis: Savoirs, culture et politique by Alfredo Perifano. Isis, 92(3): 600–601, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080758.

Barnes:2001:BRB

[702] Bernadine Barnes. Book review: The Invention of the Italian Renaissance Printmaker by Evelyn Lincoln. Isis, 92(3):601–602, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080759.

Hellyer:2001:BRB

[703] Marcus Hellyer. Book review: On Tycho’s Island: Tycho Brahe and His Assistants, 1570–1601 by John Robert Christianson. Isis, 92(3):602–603, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080760.

Long:2001:BRB

[704] Pamela O. Long. Book review: Die Mechanik des Niccolò Tartaglia: Im Kontext der zeitgenössischen Erkenntnis- und Wissenschaftstheorie by Gerhard Arend. Isis, 92(3):603–604, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080761.

Unat:2001:BRB

[705] Yavuz Unat. Book review: The Melon-Shaped Astrolabe in Arabic Astronomy by E. S. Kennedy; P. Kunitzsch; R. P. Lorch; E. S. Kennedy; P. Kunitzsch; R. P. Lorch. Isis, 92(3):604–605, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080762.

Iliffe:2001:BRB

[706] Rob Iliffe. Book review: Generall Learning: A Seventeenth-Century Treatise on the Formation of the General Scholar by Meric Casaubon; Richard Serjeantson. Isis, 92(3):605–606, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080763.

Daston:2001:BRB

[707] Lorraine Daston. Book review: Hauptschriften zur Versicherungs- und Finanzmathematik by Gottfried Wilhelm Leibniz; Eberhard Knobloch; J.-Matthias Graf von der Schulenberg. Isis, 92(3):606–607, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080764.

Martens:2001:BRB

[708] Rhonda Martens. Book review: Optics: Paralipomena to Witelo, and Optical Part of Astronomy by Johannes Kepler; William H. Donahue. Isis, 92(3):607–608, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080765.

Steele:2001:BRB

[709] Brett Steele. Book review: War for the Every Day: Generals, Knowledge, and the Warfare in Early Modern Europe, 1680–1740 by Erik A. Lund. Isis, 92(3):608–609, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080766.

Sarasohn:2001:BRB

[710] Lisa T. Sarasohn. Book review: “The Convent of Pleasure” and Other Plays by Margaret Cavendish; Anne Shaver. Isis, 92(3):609, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080767.

Yochelson:2001:BRBb

[711] Ellis L. Yochelson. Book review: The Selected Papers of Charles Willson Peale and His Family. Volume 5: The Autobiography of Charles Willson Peale by Charles Willson Peale; Sidney Hart; David D. Ward; Lillian B. Miller. Isis, 92(3):609–611, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080768.

Bork:2001:BRB

[712] Kennard B. Bork. Book review: Manuscripts and Publications of Horace– Bénédict de Saussure on the Origin of Basalt (1772–1797) by Albert V. Carozzi. Isis, 92(3):611, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080769.

Lawrence:2001:BRB

[713] Christopher Lawrence. Book review: Obesity and Depression in the Enlightenment: The Life and Times of George Cheyne by Anita Guerrini. Isis, 92(3):611–612, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080770.

Cutter:2001:BRB

[714] Donald C. Cutter. Book review: La física de la Monarquía: Ciencia y política en el pensamiento colonial de Alejandro Malaspina (1754– 1810) by Juan Pimentel. Isis, 92(3):612–613, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080771.

Lightman:2001:BRB

[715] Bernard Lightman. Book review: The “Creed of Science” in Victorian England by Roy M. MacLeod. Isis, 92(3):613–614, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080772.

Osborne:2001:BRB

[716] Michael A. Osborne. Book review: Laboratoires du nouveau siècle: La nébuleuse réformatrice et ses réseaux en France, 1880–1914 by Christian Topalov. Isis, 92(3):614–615, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080773.

Garber:2001:BRBb

[717] Elizabeth Garber. Book review: The Science of Energy: A Cultural History of Energy Physics in Victorian Britain by Crosbie Smith. Isis, 92 (3):615–616, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080774.

Wilson:2001:BRBe

[718] David B. Wilson. Book review: From Galaxies to Turbines: Science, Technology, and the Parsons Family by Garrett W. Scaife. Isis, 92(3): 616–618, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080775.

Kurrer:2001:BRB

[719] Karl-Eugen Kurrer. Book review: Karl Culmann und die graphische Statik: Anhang mit umfangreichen Culmann-Texten by Bertram Maurer. Isis, 92(3):618, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080776.

Kline:2001:BRBb

[720] Ronald Kline. Book review: The Papers of Thomas A. Edison: The Wizard of Menlo Park, 1878 by Thomas A. Edison; Paul B. Israel; Keith A. Nier; Louis Carlat. Isis, 92(3):618–619, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080777.

Magoun:2001:BRBa

[721] Alexander Magoun. Book review: Scripts, Grooves, and Writing Machines: Representing Technology in the Edison Era by Lisa Gitelman. Isis, 92(3):619–620, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080778.

Montgomery:2001:BRB

[722] William Montgomery. Book review: Disseminating Darwinism: The Role of Place, Race, Religion, and Gender by Ronald L. Numbers; John Stenhouse. Isis, 92(3):620–621, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080779.

Dorman:2001:BRB

[723] Robert L. Dorman. Book review: George Perkins Marsh: Prophet of Conservation by David Lowenthal. Isis, 92(3):621–622, September 2001.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080780.

Larson:2001:BRBb

[724] Edward J. Larson. Book review: Charles Darwin’s Zoology Notes and Specimen Lists from H.M.S. by Richard Keynes. Isis, 92(3):622–623, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080781.

Richmond:2001:BRB

[725] Marsha L. Richmond. Book review: Der Rücktritt Richard Willstätters 1924–1925 und seine Hintergründe: Ein Münchener Universitätsskandal? by Freddy Litten. Isis, 92(3):623–624, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080782.

McLay:2001:BRL

[726] David B. McLay. Book review: Lise Meitner and the Dawn of the Nuclear Age by Patricia Rife. Isis, 92(3):624–625, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080783.

Eisberg:2001:BRB

[727] Joann Eisberg. Book review: Henry Norris Russell: Dean of American Astronomers by David H. DeVorkin. Isis, 92(3):625–626, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080784.

Johnson:2001:BRBb

[728] Jeffrey Allan Johnson. Book review: Walther Nernst and the Transition to Modern Physical Science by Diana Kormos Barkan. Isis, 92(3):626– 627, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/3080785.

Kroll:2001:BRB

[729] Gary Kroll. Book review: Wildlife Films by Derek Bousé. Isis, 92(3): 627–628, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080786.

Weir:2001:BRB

[730] Gary E. Weir. Book review: Argonaut: The Submarine Legacy of Simon Lake by John J. Poluhowich. Isis, 92(3):628–629, September 2001.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080787.

Blustein:2001:BRB

[731] Bonnie Ellen Blustein. Book review: The Woman Who Knew Too Much: Alice Stewart and the Secrets of Radiation by Gayle Greene; Dr. America: The Lives of Thomas A. Dooley, 1927–1961 by James T. Fisher. Isis, 92 (3):629–630, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080788.

Roque:2001:BRB

[732] Xavier Roqué. Book review: Pioneras españolas en las ciencias: Las mujeres del Instituto Nacional de Física y Química by Carmen Magallón Portolés. Isis, 92(3):630–631, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080789.

Williams:2001:BRBb

[733] Karen Patricia Williams. Book review: Black Women Scientists in the United States by Wini Warren. Isis, 92(3):631–632, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080790.

Crease:2001:BRBa

[734] Robert P. Crease. Book review: From Hiroshima to the Iceman: The Development and Applications of Accelerator Mass Spectrometry by Harry E. Gove. Isis, 92(3):632–633, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080791.

Maines:2001:BRB

[735] Rachel Maines. Book review: Cleavage: Technology, Controversy, and the Ironies of the Man-Made Breast by Nora Jacobson. Isis, 92(3):633– 634, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/3080792.

Abel:2001:BRB

[736] Emily K. Abel. Book review: Contagion and Confinement: Controlling Tuberculosis along the Skid Road by Barron H. Lerner. Isis, 92(3):634– 635, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/3080793.

Heitmann:2001:BRB

[737] John A. Heitmann. Book review: Brush with Death: A Social History of Lead Poisoning by Christian Warren. Isis, 92(3):635, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080794.

Ash:2001:BRB

[738] Mitchell G. Ash. Book review: Death of a “Jewish Science”: Psychoanalysis in the Third Reich by James E. Goggin; Eileen N. Brockman Goggin. Isis, 92(3):636–637, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080795.

Perkins:2001:BRBb

[739] John H. Perkins. Book review: The Agricultural Revolution of the Twentieth Century by Don Paarlberg; Philip Paarlberg. Isis, 92(3):637–638, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080796.

Berenbaum:2001:BRB

[740] May Berenbaum. Book review: Entomology, Ecology, and Agriculture: The Making of Scientific Careers in North America, 1885–1995 by Paolo Palladino. Isis, 92(3):638–639, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080797.

Steen:2001:BRB

[741] Kathryn Steen. Book review: Sources of Industrial Leadership: Studies of Seven Industries by David C. Mowery; Richard R. Nelson. Isis, 92 (3):639–640, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080798.

Chaisson:2001:BRB

[742] Eric J. Chaisson. Book review: The Rocket and the Reich: Peenemunde and the Coming of the Ballistic Missile Era by Michael J. Neufeld. Isis, 92 (3):640–641, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080799.

Bird:2001:BRB

[743] Alexander Bird. Book review: Philosophie des sciences de la nature by Angèle Kremer-Marietti. Isis, 92(3):641–642, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080800.

MacKinnon:2001:BRBb

[744] Edward MacKinnon. Book review: Models as Mediators: Perspectives on Natural and Social Science by Mary S. Morgan; Margaret Morrison. Isis, 92(3):642–643, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080801.

Brown:2001:BRBb

[745] Richard Brown. Book review: For and against Method: Including Lakatos’s Lectures on Scientific Method and the Lakatos-Feyerabend Correspondence by Imre Lakatos; Paul Feyerabend; Matteo Motterlini. Isis, 92(3):643, September 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080802.

Porter:2001:BRB

[746] Theodore M. Porter. Book review: The Norton History of the Human Sciences by Roger Smith. Isis, 92(3):644, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080803.

Johnson:2001:BRBc

[747] Dale M. Johnson. Book review: Handbook of the History of General Topology, Volume 1 by C. E. Aull; R. Lowen. Isis, 92(3):645, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080804.

Anonymous:2001:Cc

[748] Anonymous. Collections. Isis, 92(3):646–651, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080805.

Anonymous:2001:BMd

[749] Anonymous. Back matter. Isis, 92(3):652–655, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080806.

Anonymous:2001:FMd

[750] Anonymous. Front matter. Isis, 92(3):??, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080729.

Mazzotti:2001:MGA

[751] Massimo Mazzotti. Maria Gaetana Agnesi: Mathematics and the making of the Catholic Enlightenment. Isis, 92(4):657–683, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080337.

Herman:2001:FMS

[752] Ellen Herman. Families made by science: Arnold gesell and the technologies of modern child adoption. Isis, 92(4):684–715, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080338.

Quinlan-McGrath:2001:FHS

[753] Mary Quinlan-McGrath. The foundation horoscope(s) for St. Peter’s Basilica, Rome, 1506: Choosing a time, changing the storia. Isis, 92 (4):716–741, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080339.

deRomo:2001:EMD

[754] Ana Cecilia Rodríguez de Romo. Eloge: Mirko D. Grmek, 9 January 1924–6 March 2000. Isis, 92(4):742–744, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080340.

Rochberg:2001:BRB

[755] Francesca Rochberg. Book review: Astronomical Papyri from Oxyrhynchus (P Oxy. 4133–4300a), by Alexander Jones. Isis, 92(4): 745–748, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080341.

Gooday:2001:BRB

[756] Graeme Gooday. Book review: 125 Years: The Physical Society and the Institute of Physics by John L. Lewis. Isis, 92(4):749, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080342.

Hamerla:2001:BRB

[757] Rich Hamerla. Book review: Science at the American Frontier: A Biography of DeWitt Bristol Brace by David Cahan; M. Eugene Rudd. Isis, 92 (4):749–750, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080343.

Gorman:2001:BRB

[758] Michael John Gorman. Book review: Jesuitische Mathematik: In Prag im 16. und 17. Jahrhundert by Georg Schuppener. Isis, 92(4):750–751, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080344.

Goldberg:2001:BRB

[759] Michael Goldberg. Book review: Flatterland: Like Flatland Only More so by Ian Stewart. Isis, 92(4):751–752, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080345.

Sengor:2001:BRB

[760] A. M. C. Sengör. Book review: Melting the Earth: The History of Ideas on Volcanic Eruptions by Haraldur Sigurdsson. Isis, 92(4):752–753, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080346.

Alder:2001:BRB

[761] Ken Alder. Book review: Une histoire de la Méridienne: Textes, enjeux, débats et passions autour de Méridien de Paris, 1666–1827 by Jean–Pierre Martin. Isis, 92(4):753–754, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3080347.

Fisher:2001:BRB

[762] David E. Fisher. Book review: Exposing Electronics by Bernard Finn; Robert Bud; Helmuth Trischler. Isis, 92(4):754–755, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080348.

Crease:2001:BRBb

[763] Robert P. Crease. Book review: Imagined Worlds by Freeman Dyson. Isis, 92(4):755, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080349.

Magoun:2001:BRBb

[764] Alexander B. Magoun. Book review: A Spiral Way: How the Phonograph Changed Ethnography by Erika Brady. Isis, 92(4):755–756, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080350.

Marchand:2001:BRB

[765] Suzanne Marchand. Book review: Minotaur: Sir Arthur Evans and the Archaeology of the Minoan Myth by Joseph Alexander MacGillivray. Isis, 92(4):756–758, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080351.

Mathew:2001:BRB

[766] William M. Mathew. Book review: A Kind of Fate: Agricultural Change in Virginia, 1861–1920 by G. Terry Sharrer. Isis, 92(4):758–759, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080352.

Park:2001:BRB

[767] Katharine Park. Book review: At the Borders of the Human: Beasts, Bodies, and Natural Philosophy in the Early Modern Period by Erica Fudge; Ruth Gilbert; Susan Wiseman. Isis, 92(4):759–760, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080353.

Barahona:2001:BRB

[768] Ana Barahona. Book review: Plantas americanas para la España Ilustrada: Génesis, desarrollo y ocaso del proyecto español de expediciones botánicas by Antonio González Bueno; Raúl Rodríguez Nozal. Isis, 92(4):760, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080354.

Gunnoe:2001:BRB

[769] Charles D. Gunnoe, Jr. Book review: Iliaster: Literatur and Naturkunde in der frühen Neuzeit: Festgabe für Joachim Telle zum 60. Geburtstag by Wilhelm Kühlmann; Wolf-Dieter Müller–Jahncke. Isis, 92(4):760–761, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080355.

Laubichler:2001:BRB

[770] Manfred D. Laubichler. Book review: Biology Integrating Scientific Fundamentals: Contributions to the History of Interrelations between Biology, Chemistry, and Physics from the Eighteenth to the Twentieth Centuries by Brigitte Hoppe. Isis, 92(4):761–762, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080356.

Caton:2001:BRBb

[771] Donald Caton. Book review: Birth Chairs, Midwives, and Medicine by Amanda Carson Banks. Isis, 92(4):762–763, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080357.

Saliba:2001:BRB

[772] George Saliba. Book review: The Diffusion of Greco-Roman Medicine into the Middle East and the Caucasus by John A. C. Greppin; Emilie Savage-Smith; John L. Gueriguian. Isis, 92(4):763–764, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080358.

Delacy:2001:BRB

[773] Margaret Delacy. Book review: The Tragedy of Childbed Fever by Irvine Loudon. Isis, 92(4):764–765, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080359.

Givens:2001:BRBb

[774] Douglas R. Givens. Book review: Bringing Back the Past: Historical Perspectives on Canadian Archaeology by Pamela Jane Smith; Donald Mitchell. Isis, 92(4):765–766, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080360.

Lucier:2001:BRBa

[775] Paul Lucier. Book review: Gold: The California Story by Mary Hill. Isis, 92(4):766–767, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080361.

Deacon:2001:BRBb

[776] Harriet Deacon. Book review: Science and Society in Southern Africa by Saul Dubow. Isis, 92(4):767–768, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080362.

Pinna:2001:BRB

[777] Giovanni Pinna. Book review: Nel nome della razza: Il razzismo nella storia d’Italia 1870–1945 by Alberto Burgio. Isis, 92(4):768–769, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080363.

Kassler:2001:BRB

[778] Jamie C. Kassler. Book review: Number to Sound: The Musical Way to the Scientific Revolution by Paolo Gozza. Isis, 92(4):769–770, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080364.

Meadows:2001:BRB

[779] Jack Meadows. Book review: Journals and History of Science by Marco Beretta; Claudio Pogliano; Pietro Redondi. Isis, 92(4):770–771, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080365.

Seager:2001:BRB

[780] William Seager. Book review: The Taboo of Subjectivity: Toward a New Science of Consciousness by B. Alan Wallace. Isis, 92(4):771–772, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080366.

Ziporyn:2001:BRB

[781] Terra Ziporyn. Book review: Sorting Things out: Classification and Its Consequences by Geoffrey C. Bowker; Susan Leigh Star. Isis, 92 (4):772–773, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080367.

Machamer:2001:BRB

[782] Peter Machamer and Abel Franco. Book review: Biographies of Scientific Objects by Lorraine Daston. Isis, 92(4):773–774, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080368.

Brann:2001:BRB

[783] Eva Brann. Book review: Time, Conflict, and Human Values by J. T. Fraser. Isis, 92(4):774–775, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080369.

Nickles:2001:BRB

[784] Thomas Nickles. Book review: The Logic and Methodology of Science in Early Modern Thought: Seven Studies by Fred Wilson. Isis, 92(4): 775–776, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080370.

Jones:2001:BRB

[785] Greta Jones. Book review: The Contest for Social Science: Relations and Representations of Gender and Class by Eileen Janes Yeo. Isis, 92 (4):776–777, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080371.

Kelley:2001:BRB

[786] Donald R. Kelley. Book review: The Waning of the Renaissance, 1550– 1640 by William J. Bouwsma. Isis, 92(4):777–778, December 2001. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080372.

Efron:2001:BRB

[787] Noah J. Efron. Book review: The Jews and the Sciences in the Middle Ages by Y. Tzvi Langermann. Isis, 92(4):778–779, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080373.

Payne:2001:BRBb

[788] Lynda Stephenson Payne. Book review: The Nobility and Excellence of Women and the Defects and Vices of Men by Lucrezia Marinella; Anne Dunhill. Isis, 92(4):779–780, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080374.

Arrizabalaga:2001:BRB

[789] Jon Arrizabalaga. Book review: Dissection and Vivisection in the European Renaissance by Roger French. Isis, 92(4):780–781, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080375.

Getz:2001:BRB

[790] Faye Getz. Book review: Bibliographie des textes médicaux latins: Antiquité et haute moyen âge, Premier supplément 1986–1999 by Klaus-Dietrich Fischer. Isis, 92(4):781, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080376.

Walton:2001:BRB

[791] Steven A. Walton. Book review: Thomas Harriot: An Elizabethan Man of Science by Robert Fox. Isis, 92(4):781–782, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080377.

Snobelen:2001:BRB

[792] Stephen D. Snobelen. Book review: The Foundations of Newtonian Scholarship by Richard H. Dalitz; Michael Nauenberg. Isis, 92(4):782– 783, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080378.

Vickers:2001:BRB

[793] Brian Vickers. Book review: Selected Philosophical Works by Francis Bacon. Isis, 92(4):783–784, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080379.

Beyer:2001:BRB

[794] Robert T. Beyer. Book review: The Acoustic World of Early Modern England: Attending to the O-Factor by Bruce R. Smith. Isis, 92(4): 784–785, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080380.

Meli:2001:BRB

[795] Domenico Bertoloni Meli. Book review: On the Reason of the Movement of the Muscles by William Croone; Paul Maquet. Isis, 92(4):785–786, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080381.

Riskin:2001:BRB

[796] Jessica Riskin. Book review: The Sciences in Enlightened Europe by William Clark; Jan Golinski; Simon Schaffer. Isis, 92(4):786–788, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080382.

Hessenbruch:2001:BRB

[797] Arne Hessenbruch. Book review: Brandy, Balloons, and Lamps: Ami Argand, 1750–1803 by John J. Wolfe. Isis, 92(4):788–789, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080383.

Nye:2001:BRB

[798] Mary Jo Nye. Book review: Pierre Curie, 1859–1906: Le rêve scientifique by Loïc Barbo. Isis, 92(4):789–790, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080384.

Rowe:2001:BRB

[799] David E. Rowe. Book review: Bernhard Riemann, 1826–1866: Turning Points in the Conception of Mathematics by Detlef Laugwitz; Abe Shenitzer; Hardy Grant; Sarah Shenitzer. Isis, 92(4):790–791, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080385.

Atz:2001:BRB

[800] James W. Atz. Book review: The History of Steinhart Aquarium: A Very Fishy Tale by John E. McCosker. Isis, 92(4):791–792, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080386.

Smout:2001:BRB

[801] T. C. Smout. Book review: Secure from Rash Assault: Sustaining the Victorian Environment by James Winter. Isis, 92(4):792–794, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080387.

Rybczynski:2001:BRB

[802] Witold Rybczynski. Book review: Park Maker: A Life of Frederick Law Olmsted by Elizabeth Stevenson. Isis, 92(4):794–795, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080388.

Goetzmann:2001:BRB

[803] William H. Goetzmann. Book review: A River Running West: The Life of John Wesley Powell by Donald Worster. Isis, 92(4):795–796, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080389.

Lucier:2001:BRBb

[804] Paul Lucier. Book review: New York State Natural History Survey: 1836–1842 by Michele L. Aldrich. Isis, 92(4):796–798, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080390.

Lyons:2001:BRB

[805] Sherrie Lyons. Book review: The Correspondence of Charles Darwin. Volume 11: 1863 by Charles Darwin; Frederick Burkhardt; Duncan M. Porter; Sheila Ann Dean; Jonathan R. Topham; Sarah Wilmot. Isis, 92 (4):798–799, December 2001. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/3080391.

Noelke:2001:BRB

[806] Virginia Noelke. Book review: Susan La Flesche Picotte, M.D.: Omaha Indian Leader and Reformer by Benson Tong. Isis, 92(4):799–800, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080392.

Walls:2001:BRBb

[807] Laura Dassow Walls. Book review: Thoreau’s Late Career and “The Dispersion of Seeds”: The Saunterer’s Synoptic Vision by Michael Benjamin Berger. Isis, 92(4):800–801, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080393.

Herron:2001:BRB

[808] Timothy J. Herron. Book review: His Glassy Essence: An Autobiography of Charles Sanders Peirce by Charles Sanders Peirce. Isis, 92(4):801–802, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080394.

Kaiser:2001:BRBb

[809] David Kaiser. Book review: The Particle Century by Gordon Fraser. Isis, 92(4):802–803, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080395.

Larabee:2001:BRB

[810] Ann Larabee. Book review: Making a Real Killing: Rocky Flats and the Nuclear West by Len Ackland. Isis, 92(4):803–804, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080396.

Strick:2001:BRB

[811] James Strick. Book review: Carl Sagan’s Universe by Yervant Terzian; Elizabeth Bilson. Isis, 92(4):804–805, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080397.

Baird:2001:BRB

[812] Davis Baird. Book review: Arnold O. Beckman: One Hundred Years of Excellence by Arnold Thackray; Minor Myers. Isis, 92(4):805–806, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080398.

Otis:2001:BRB

[813] Laura Otis. Book review: Freud and His Aphasia Book: Language and the Sources of Psychoanalysis by Valerie D. Greenberg. Isis, 92(4):806– 807, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080399.

Todd:2001:BRB

[814] Edmund N. Todd. Book review: Die “Tönenden Funken”: Geschichte eines frühen drahtlosen Kommunikationssystems, 1905–1914 by Michael Friedewald. Isis, 92(4):807–808, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080400.

Syrett:2001:BRB

[815] David Syrett. Book review: Planned Invasion of Japan, 1945: The Siberian Weather Advantage by H. S. Yoder. Isis, 92(4):808–809, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080401.

Slotten:2001:BRB

[816] Hugh R. Slotten. Book review: The Beginning of Broadcast Regulation in the Twentieth Century by Marvin R. Bensman. Isis, 92(4):809, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/3080402.

Slaybaugh:2001:BRB

[817] Douglas Slaybaugh. Book review: Willard Cochrane and the American Family Farm by Richard A. Levins. Isis, 92(4):809–810, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080403.

Chianese:2001:BRB

[818] Robert Chianese. Book review: Back from the Far Field: American Nature Poetry in the Late Twentieth Century by Bernard W. Quetchenbach. Isis, 92(4):810–811, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080404.

Howell:2001:BRB

[819] Joel Howell. Book review: Teaching Hospitals and the Urban Poor by Eli Ginzberg. Isis, 92(4):811–812, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080405.

Wear:2001:BRB

[820] Andrew Wear. Book review: Health and Society in Britain since 1939 by Virginia Berridge. Isis, 92(4):812–813, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080406.

Young:2001:BRB

[821] Allan Young. Book review: Walter B. Cannon: Science and Society by Elin L. Wolfe; A. Clifford Barger; Saul Benison. Isis, 92(4):813–814, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080407.

Noll:2001:BRB

[822] Steven Noll. Book review: The Kennedy Family and the Story of Mental Retardation by Edward Shorter. Isis, 92(4):814–815, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080408.

Weldon:2001:BRBb

[823] Stephen P. Weldon. Book review: Science or Pseudoscience: Magnetic Healing, Psychic Phenomena, and Other Heterodoxies by Henry H. Bauer. Isis, 92(4):815–816, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/3080409.

Kincaid:2001:BRB

[824] Harold Kincaid. Book review: The Uncertain Sciences by Bruce Mazlish. Isis, 92(4):816–817, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

3080410.

Godin:2001:BRB

[825] Benoit Godin. Book review: L’opinion publique et la science: A chacun son ignorance by Bernadette Bensaude-Vincent. Isis, 92(4):817, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/3080411.

Mauskopf:2001:BRB

[826] Seymour Mauskopf. Book review: Eloge du mixte: Materiaux nouveau et philosophie ancienne by Bernadette Bensaude-Vincent. Isis, 92(4):818, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080412.

Anonymous:2001:FMe

[827] Anonymous. Front matter. Isis, 92(4):819–823, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080336.

Anonymous:2001:VIb

[828] Anonymous. Volume information. Isis, 92(4):824–844, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080335.

Anonymous:2001:BMe

[829] Anonymous. Back matter. Isis, 92(4):??, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3080413.

Weldon:2002:CBH

[830] Stephen P. Weldon. Current bibliography of the history of science and its cultural influences 2002. Isis, 93(??):1–237, ???? 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3648116.

Anonymous:2002:BM

[831] Anonymous. Back matter. Isis, 93(??):??, ???? 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3648117.

Anonymous:2002:FM

[832] Anonymous. Front matter. Isis, 93(??):i–xv, ???? 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3648115.

Hamblin:2002:NSP

[833] Jacob Darwin Hamblin. The Navy’s “sophisticated” pursuit of science: Undersea warfare, the limits of internationalism, and the utility of basic research, 1945–1956. Isis, 93(1):1–27, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343244.

Secord:2002:BPP

[834] Anne Secord. Botany on a plate: Pleasure and the power of pictures in promoting early Nineteenth-Century scientific knowledge. Isis, 93(1): 28–57, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343245.

Thurston:2002:GMA

[835] Hugh Thurston. Greek mathematical astronomy reconsidered. Isis, 93(1):58–69, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343242.

Gingerich:2002:TP

[836] Owen Gingerich. The trouble with Ptolemy. Isis, 93(1):70–74, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343243.

Cohen:2002:ECH

[837] Floris Cohen. Eloge: Casper Hakfoort, 6 January 1955–44 March 1999. Isis, 93(1):75–77, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343246.

Speiser:2002:E

[838] Ambros P. Speiser. To the editor. Isis, 93(1):78–79, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343248.

Knowles:2002:R

[839] Scott G. Knowles and Stuart W. Leslie. In reply. Isis, 93(1):79–80, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343249.

Knell:2002:E

[840] Simon Knell. To the editor. Isis, 93(1):80–81, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343250.

Dean:2002:R

[841] Dennis R. Dean. In reply. Isis, 93(1):81–82, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343251.

Spence:2002:BRR

[842] Kate Spence. Book review: Roger Herz-Fischler: The Shape of the Great Pyramid. Isis, 93(1):83–84, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343252.

Brentjes:2002:BRB

[843] Sonja Brentjes. Book review: Die Schrift des Ibrahim b, by Paul Luckey; Jan P. Hogendijk. Sinan b. Tabit über die Schatteninstrumente. Isis, 93(1):84–85, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343253.

Donnelly:2002:BRB

[844] James Donnelly. Book review: Chemistry, Society, and Environment: A New History of the British Chemical Industry. Isis, 93(1):85–86, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343254.

Cahan:2002:BRR

[845] David Cahan. Book review: Robert Fox; Anna Guagnini: Laboratories, Workshops, and Sites: Concepts and Practices of Research in Industrial Europe, 1800–1914. Isis, 93(1):86–87, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343255.

Schulten:2002:BRF

[846] Susan Schulten. Book review: Felix Driver: Geography Militant: Cultures of Exploration and Empire. Isis, 93(1):87–88, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343256.

Dean:2002:BRC

[847] Dennis R. Dean. Book review: Christopher McGowan: The Dragon Seekers: How an Extraordinary Circle of Fossilists Discovered the Dinosaurs and Paved the Way for Darwin. Isis, 93(1):88–89, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343257.

Webb:2002:BRM

[848] George E. Webb. Book review: Michael Ruse: The Evolution Wars: A Guide to the Debates. Isis, 93(1):89–90, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/343258.

Baldwin:2002:BRE

[849] Carole Baldwin. Book review: Edward J. Larson: Evolution’s Workshop: God and Science on the Galápagos Islands. Isis, 93(1):90–91, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343259.

Fleming:2002:BRBa

[850] James Rodger Fleming. Book review: Eduard Brückner: The Sources and Consequences of Climate Change and Climate Variability in Historical Times. Isis, 93(1):91, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343260.

Kidwell:2002:BRS

[851] Clara Sue Kidwell. Book review: Sherry L. Smith: Reimagining Indians: Native Americans through Anglo Eyes, 1880–1940. Isis, 93(1):92, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343261.

Abbate:2002:BRT

[852] Janet Abbate. Book review: Thomas C. Jepsen: My Sisters Telegraphic: Women in the Telegraph Office, 1846–1950. Isis, 93(1):92–93, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343262.

Broman:2002:BRD

[853] Thomas Broman. Book review: Daniel R. Headrick: When Information Came of Age: Technologies of Knowledge in the Age of Reason and Revolution, 1700–1850. Isis, 93(1):93–94, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343263.

Vermeer:2002:BRR

[854] Eduard B. Vermeer. Book review: Randall A. Dodgen: Controlling the Dragon: Confucian Engineers and the Yellow River in Late Imperial China. Isis, 93(1):94–95, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343264.

Apple:2002:BRE

[855] Rima D. Apple. Book review: Emily Abel: Hearts of Wisdom: American Women Caring for Kin, 1840–1940. Isis, 93(1):95–96, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343265.

Wilson:2002:BRP

[856] Philip K. Wilson. Book review: Peter Lewis Allen: The Wages of Sin: Sex and Disease, Past and Present. Isis, 93(1):96–97, March 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343266.

Warren:2002:BRR

[857] Christian Warren. Book review: R. Ian McCallum: Antimony in Medical History: An Account of the Medical Uses of Antimony and Its Compounds Since Early Times to the Present. Isis, 93(1):97–98, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343267.

King:2002:BRW

[858] Bill King. Book review: W. Michael Byrd; Linda A. Clayton: An American Health Dilemma: A Medical History of African Americans and the Problem of Race: Beginnings to 1900. Isis, 93(1):98–99, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343268.

Haug:2002:BRR

[859] Ruth J. Haug. Book review: Randal L. Hall: William Louis Poteat: A Leader of the Progressive-Era South. Isis, 93(1):99–100, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343269.

Wakefield:2002:BRR

[860] Andre Wakefield. Book review: Richard F. Wetzell: Inventing the Criminal: A History of German Criminology, 1880–1945. Isis, 93(1):100–101, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/343270.

Crowe:2002:BRR

[861] Michael Crowe. Book review: Roger Hennessey: Worlds without End: The Historic Search for Extraterrestrial Life. Isis, 93(1):101–102, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343271.

Behboud:2002:BRS

[862] Ali Behboud. Book review: Serafina Cuomo: Pappus of Alexandria and the Mathematics of Late Antiquity. Isis, 93(1):102–103, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343272.

French:2002:BRJ

[863] Roger French. Book review: John F. Healy: Pliny the Elder on Science and Technology. Isis, 93(1):103, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343273.

Lang:2002:BRM

[864] Helen Lang. Book review: Malcolm Wilson: Aristotle’s Theory of the Unity of Science. Isis, 93(1):104, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343274.

Lennox:2002:BRD

[865] James G. Lennox. Book review: Dennis Des Chene: Life’s Form: Late Aristotelian Conceptions of the Soul. Isis, 93(1):104–105, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343275.

Moreton:2002:BRG

[866] Jennifer Moreton. Book review: Georges Declercq: Anno Domini: The Origins of the Christian Era. Isis, 93(1):105–106, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343276.

Allen:2002:BRW

[867] James P. Allen. Book review: Wolfhart Westendorf: Handbuch der altägyptischen Medizin. Isis, 93(1):106–107, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/343277.

Whitmore:2002:BRB

[868] Thomas M. Whitmore. Book review: Imperfect Balance: Landscape Transformations in the Precolumbian Americas. Isis, 93(1):107–108, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/343278.

Egginton:2002:BRS

[869] William Egginton. Book review: Simon A. Gilson: Medieval Optics and Theories of Light in the Works of Dante. Isis, 93(1):108–109, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343279.

Elman:2002:BRY

[870] Benjamin A. Elman. Book review: Yung Sik Kim: The Natural Philosophy of Chu Hsi, 1130–1200. Isis, 93(1):109–110, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343280.

Davenport:2002:BRS

[871] Anne A. Davenport. Book review: Steven P. Marrone: The Light of Thy Countenance: Science and Knowledge of God in the Thirteenth Century. Volume 1: A Doctrine of Divine Illumination; Volume 2: God at the Core of Cognition. Isis, 93(1):110–111, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343281. See erratum [1183].

Heilbron:2002:BRN

[872] John L. Heilbron. Book review: Nigel Hiscock: The Wise Master Builder: Platonic Geometry in Plans of Medieval Abbeys and Cathedrals. Isis, 93(1):111–112, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343282.

Aiken:2002:BRA

[873] Jane Andrews Aiken. Book review: Anthony Grafton: Leon Battista Alberti: Master Builder of the Italian Renaissance. Isis, 93(1):112–113, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/343283.

Principe:2002:BRG

[874] Lawrence M. Principe. Book review: George Ripley: George Ripley’s Compound of Alchymy (1591). Isis, 93(1):113, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343284.

Pereira:2002:BRD

[875] Michela Pereira. Book review: D. Zecaire: Opuscule tres-eccellent de la vraye philosophie naturelle des metaulx. Isis, 93(1):113–114, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343285.

Meli:2002:BRB

[876] Domenico Bertoloni Meli. Book review: Isaac Newton’s Natural Philosophy. Isis, 93(1):114–115, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343286.

Finocchiaro:2002:BRG

[877] Maurice A. Finocchiaro. Book review: Giancarlo Nonnoi: Saggi Galileiani: Atomi, immagini e ideologia. Isis, 93(1):116, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343287.

Osler:2002:BRD

[878] Margaret J. Osler. Book review: Dennis Des Chene: Spirits and Clocks: Machine and Organism in Descartes. Isis, 93(1):116–117, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343288.

Shackelford:2002:BRA

[879] Jole Shackelford. Book review: Antonio Clericuzio: Elements, Principles, and Corpuscles: A Study of Atomism and Chemistry in the Seventeenth Century. Isis, 93(1):117–118, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343289.

Page:2002:BRZ

[880] Sophie Page. Book review: Zakiya Hanafi: The Monster in the Machine: Magic, Medicine, and the Marvelous in the Time of the Scientific Revolution. Isis, 93(1):118–119, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343290.

Fan:2002:BRG

[881] Fa ti Fan. Book review: Georgius Everhardus Rumphius: The Ambonese Curiosity Cabinet. Isis, 93(1):119–120, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343291.

Harrison:2002:BRBa

[882] Peter Harrison. Book review: Barbara M. Benedict: Curiosity: A Cultural History of Early Modern Inquiry. Isis, 93(1):120–121, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343292.

Williams:2002:BRC

[883] Richard S. Williams. Book review: Claire Richter Sherman: Writing on Hands: Memory and Knowledge in Early Modern Europe. Isis, 93 (1):121–122, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343293.

Canizares-Esguerra:2002:BRB

[884] Jorge Cañizares-Esguerra. Book review: The Mexican Treasury: The Writings of Dr, by Francisco Hernández; Simon Varey; Rafael Chabrán; Cynthia L. Chamberlain; Simon Varey; Simon Varey; Rafael Chabrán; Dora B. Weiner; Mauricio Nieto Olarte. Francisco Hernández Searching for the Secrets of Nature: The Life and Works of Dr. Francisco Hernández Remedios para el imperio: Historia natural y la apropriación del Nuevo Mondo. Isis, 93(1):122–123, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343294.

Sarasohn:2002:BRM

[885] Lisa T. Sarasohn. Book review: Molly McClain: Beaufort: The Duke and His Duchess, 1657–1715. Isis, 93(1):123–124, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343295.

Sarasohn:2002:BRP

[886] Lisa T. Sarasohn. Book review: Peter N. Miller: Peiresc’s Europe: Learning and Virtue in the Seventeenth Century. Isis, 93(1):124–125, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343296.

Olson:2002:BRBa

[887] Richard Olson. Book review: The Scottish Enlightenment: Essays in Reinterpretation. Isis, 93(1):125–126, March 2002. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/343297.

Corry:2002:BRC

[888] Leo Corry. Book review: Charles W. Curtis: Pioneers of Representation Theory: Frobenius, Burnside, Schur, and Brauer. Isis, 93(1):126, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343298.

Hankins:2002:BRW

[889] Thomas L. Hankins. Book review: William Rowan Hamilton: Mathematical Papers of Sir William Rowan Hamilton. Volume 4: Geometry, Analysis, Astronomy, Probability and Finite Differences, Miscellaneous. Isis, 93(1):126–127, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343299.

Ashworth:2002:BRB

[890] William J. Ashworth. Book review: Bruce Collier; James MacLachlan: Charles Babbage and the Engines of Perfection. Isis, 93(1):127– 128, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343300;
 http://www.jstor.org/stable/pdfplus/10.1086/343300.pdf.

Lanouette:2002:BRS

[891] William Lanouette. Book review: Septimus H. Paul: Nuclear Rivals: Anglo-American Atomic Relations, 1941–1952. Isis, 93(1):128–129, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/343301.

Fleck:2002:BRR

[892] George Fleck. Book review: Ruth H. Howes; Caroline L. Herzenberg: Their Day in the Sun: Women of the Manhattan Project. Isis, 93(1):129– 130, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343302.

Rothenberg:2002:BRP

[893] Marc Rothenberg. Book review: Patricia S. Whitesell: A Creation of His Own: Tappan’s Detroit Observatory. Isis, 93(1):130, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343303.

Crowe:2002:BRT

[894] Michael J. Crowe. Book review: Tom Standage: The Neptune File: A Story of Astronomical Rivalry and the Pioneers of Planet Hunting. Isis, 93(1):130–131, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343304.

Travis:2002:BRB

[895] Anthony S. Travis. Book review: The German Chemical Industry in the Twentieth Century. Isis, 93(1):131–132, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/343305.

DiGregorio:2002:BRC

[896] Mario A. Di Gregorio. Book review: Christine Hertler: Morphologische Methoden in der Evolutionsforschung. Isis, 93(1):132–133, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343306.

Wolfe:2002:BRB

[897] Audra J. Wolfe. Book review: We Can Sleep Later: Alfred D, by Franklin W. Stahl. Hershey and the Origins of Molecular Biology. Isis, 93(1):133– 134, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343307.

Herman:2002:BRJ

[898] Ellen Herman. Book review: Jonathan Gathorne-Hardy: Sex the Measure of All Things: A Life of Alfred C. Kinsey. Isis, 93(1):134–135, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343308.

Kinney:2002:BRC

[899] James Kinney. Book review: Cathy Boeckmann: A Question of Character: Scientific Racism and the Genres of American Fiction, 1892– 1912. Isis, 93(1):135–136, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343309.

Cosnett:2002:BRG

[900] J. E. Cosnett. Book review: Goldie Morgentaler: Dickens and Heredity: When Like Begets Like. Isis, 93(1):136–137, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343310.

Knight:2002:BRE

[901] David Knight. Book review: Erin O’Connor: Raw Material: Producing Pathology in Victorian Culture. (Body, Commodity, Text: Studies of Objectifying Practice.). Isis, 93(1):137–138, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343311.

ODriscoll:2002:BRM

[902] Kieran O’Driscoll. Book review: Malcolm Macmillan: An Odd Kind of Fame: Stories of Phineas Gage. Isis, 93(1):138, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343312.

Figueiroa:2002:BRJ

[903] Silvia Figueirôa. Book review: Julyan G. Peard: Race, Place, and Medicine: The Idea of the Tropics in Nineteenth-Century Brazilian Medicine. Isis, 93(1):138–139, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343313.

Lawrence:2002:BRM

[904] Christopher Lawrence. Book review: Michael Dunnill: The Plato of Praed Street: The Life and Times of Almroth Wright. Isis, 93(1):139– 140, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343314.

Brock:2002:BRM

[905] William H. Brock. Book review: Michael Worboys: Spreading Germs: Disease Theories and Medical Practice in Britain, 1865–1900. Isis, 93 (1):140–141, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343315.

Fournier:2002:BRBa

[906] Marian Fournier. Book review: Bert Theunissen: “Nut en nog eens nut”: Wetenschapsbeelden van Nederlandse natuuronderzoekers, 1800– 1900. Isis, 93(1):141–142, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343316.

Bocking:2002:BRJ

[907] Stephen Bocking. Book review: John Bellamy Foster: Marx’s Ecology: Materialism and Nature. Isis, 93(1):142–143, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343317.

Adler:2002:BRC

[908] Kraig Adler. Book review: Charles Boewe: Mantissa: A Supplement to Fitzpatrick’s. Isis, 93(1):143–144, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343318.

Zimmerman:2002:BRB

[909] Andrew Zimmerman. Book review: Bettina von Briskorn: Zur Sammlungsgeschichte afrikanischer Ethnographica im Übersee-Museum Bremen 1841–1945. Isis, 93(1):144, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343319.

Black:2002:BRL

[910] Barbara J. Black. Book review: Lynda Nead: Victorian Babylon: People, Streets, and Images in Nineteenth-Century London. Isis, 93(1):144–146, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/343320.

Burchfield:2002:BRC

[911] Joe D. Burchfield. Book review: Cherry Lewis: The Dating Game: One Man’s Search for the Age of the Earth. Isis, 93(1):146–147, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343321.

Schubert:2002:BRP

[912] Frank N. Schubert. Book review: Paula Rebert: La Gran Linea: Mapping the United States–Mexico Boundary, 1849–1857. Isis, 93(1):147– 148, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343322.

Tjossem:2002:BRR

[913] Sara Tjossem. Book review: Randal S. Beeman; James A. Pritchard: A Green and Permanent Land: Ecology and Agriculture in the Twentieth Century. Isis, 93(1):148–149, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343323.

Munday:2002:BRD

[914] Pat Munday. Book review: Donald MacMillan: Smoke Wars: Anaconda Copper, Montana Air Pollution, and the Courts, 1890–1924. Isis, 93 (1):149–150, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343324.

Haigh:2002:BRG

[915] Elizabeth Haigh. Book review: George S. Levit: Biogeochemistry– Biosphere–Noosphere: The Growth of the Theoretical System of Vladimir Ivanovich Vernadsky. Isis, 93(1):150–151, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/343325.

Lucier:2002:BRB

[916] Paul Lucier. Book review: Brian Black: Petrolia: The Landscape of America’s First Oil Boom. Isis, 93(1):151–152, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343326.

Walker:2002:BRH

[917] Mark Walker. Book review: Harry Von Kroge: Gema: Birthplace of German Radar and Sonar. Isis, 93(1):152, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343327.

Fisher:2002:BRH

[918] David E. Fisher. Book review: Hugh R. Slotten: Radio and Television Regulation: Broadcast Technology in the United States, 1920–1960. Isis, 93(1):152–153, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343328.

Seely:2002:BRD

[919] Bruce E. Seely. Book review: David A. Kirsch: The Electric Vehicle and the Burden of History. Isis, 93(1):153–154, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343329.

Vucinich:2002:BRA

[920] Alexander Vucinich. Book review: Ann Hibner Koblitz: Science, Women, and Revolution in Russia. Isis, 93(1):154–155, March 2002.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343330.

Bannister:2002:BRC

[921] Robert C. Bannister. Book review: C. Wright Mills: Letters and Autobiographical Writings. Isis, 93(1):155–156, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/343331.

Lewis:2002:BRC

[922] Albert C. Lewis. Book review: C. J. Mozzochi: The Fermat Diary. Isis, 93(1):156, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343332.

Marvin:2002:BRD

[923] Ursula B. Marvin. Book review: David H. Levy: Shoemaker by Levy: The Man Who Made an Impact. Isis, 93(1):156–157, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343333.

Bartky:2002:BRT

[924] Ian R. Bartky. Book review: Tony Jones: Splitting the Second: The Story of Atomic Time. Isis, 93(1):157–158, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343334.

Teich:2002:BRA

[925] Albert H. Teich. Book review: Alfred K. Mann: For Better or for Worse: The Marriage of Science and Government in the United States. Isis, 93(1):158–159, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343335.

Roland:2002:BRG

[926] Alex Roland. Book review: Gordon R. Mitchell: Strategic Deception: Rhetoric, Science, and Politics in Missile Defense Advocacy. Isis, 93 (1):159–160, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343336.

Guntau:2002:BRR

[927] Martin Guntau. Book review: Raymond G. Stokes: Constructing Socialism: Technology and Change in East Germany, 1945–1990. Isis, 93(1):160–161, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343337.

Olwell:2002:BRJ

[928] Russell Olwell. Book review: J. Samuel Walker: Permissible Dose: A History of Radiation Protection in the Twentieth Century. Isis, 93 (1):161–162, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343338.

Comfort:2002:BRE

[929] Nathaniel C. Comfort. Book review: Evelyn Fox Keller: The Century of the Gene. Isis, 93(1):162–163, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343339.

Strick:2002:BRC

[930] James Strick. Book review: Christopher Wills; Jeffrey Bada: The Spark of Life: Darwin and the Primeval Soup. Isis, 93(1):163–164, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343340.

Cain:2002:BRB

[931] Joe Cain. Book review: Institute of Biology: The First Fifty Years. Isis, 93(1):164, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343341.

Lederman:2002:BRD

[932] Muriel Lederman. Book review: Donna J. Haraway; Thyrza Nichols Goodeve: How Like a Leaf: An Interview with Donna J. Haraway. Isis, 93(1):164–165, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343342.

Long:2002:BRD

[933] Pamela O. Long. Book review: David Turnbull: Masons, Tricksters, and Cartographers: Comparative Studies in the Sociology of Scientific and Indigenous Knowledge. Isis, 93(1):165–166, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343343.

Harwood:2002:BRF

[934] Jonathan Harwood. Book review: Fritz Ringer: Toward a Social History of Knowledge: Collected Essays. Isis, 93(1):166–167, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343344.

Kalthoff:2002:BRM

[935] Mark A. Kalthoff. Book review: Michael Ruse: Can a Darwinian Be a Christian? The Relationship between Science and Religion. Isis, 93 (1):167–168, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343345.

Jablonski:2002:BRB

[936] Nina Jablonski. Book review: Primate Encounters: Models of Science, Gender, and Society. Isis, 93(1):168–169, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343346.

OConnor:2002:BRA

[937] Alice O’Connor. Book review: Andrew Abbott: Department and Discipline: Chicago Sociology at One Hundred. Isis, 93(1):169–170, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343347.

Grinstein:2002:BRB

[938] Louise S. Grinstein. Book review: The Biographical Dictionary of Women in Science: Pioneering Lives from Ancient Times to the Mid-Twentieth Century. Isis, 93(1):170, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343348.

Weiner:2002:BRS

[939] Dora B. Weiner. Book review: Sophie Riché; Sylvain Riquier: Des hôpitaux à Paris: Etat des fonds des Archives de I’AP-HP XIIème– XXème siècles. Isis, 93(1):170–171, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343349.

Black:2002:BRB

[940] Lois Fischer Black. Book review: Plague and Print in The Netherlands: A Short-Title Catalogue of Publications in the University Library of Amsterdam. Isis, 93(1):171–172, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343350.

Nickles:2002:BRR

[941] Thomas Nickles. Book review: Robert E. Krebs: Scientific Laws, Principles, and Theories: A Reference Guide. Isis, 93(1):172–173, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343351.

Anonymous:2002:BRBa

[942] Anonymous. Book review: Little Tools of Knowledge: Historical Essays on Academic and Bureaucratic Practices. Isis, 93(1):173, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343352.

Anonymous:2002:BRBb

[943] Anonymous. Book review: Exploring the Unknown: Selected Documents in the History of the U, by John M. Logsdon.S. Civil Space Program. Isis, 93(1):173, March 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
343353.

Anonymous:2002:BRBc

[944] Anonymous. Book review: Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century. Isis, 93(1):174, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/343354.

Anonymous:2002:BRBd

[945] Anonymous. Book review: Church, Censorship, and Culture in Early Modern Italy. Isis, 93(1):174, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343355.

Anonymous:2002:BRBe

[946] Anonymous. Book review: Images of Disease: Science, Public Policy, and Health in Post-war Europe. Isis, 93(1):174, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343356.

Anonymous:2002:BRBf

[947] Anonymous. Book review: Metaphysics, Epistemology, and Technology. Isis, 93(1):174–175, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/343357.

Anonymous:2002:BRBg

[948] Anonymous. Book review: American Philosophy of Technology: The Empirical Turn. Isis, 93(1):175, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/343358.

Anonymous:2002:BRBh

[949] Anonymous. Book review: Learning to Manage Global Environmental Risks Social Learning Group. Isis, 93(1):175, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343359.

Anonymous:2002:BRBi

[950] Anonymous. Book review: Learning to Manage Global Environmental Risks Social Learning Group. Isis, 93(1):175–176, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/343360.

Anonymous:2002:BRBj

[951] Anonymous. Book review: Elegance, Beauty, and Truth. Isis, 93(1): 176, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343361.

Anonymous:2002:NCa

[952] Anonymous. Notes on contributors. Isis, 93(1):177–182, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/343247.

Anonymous:2002:Fa

[953] Anonymous. Frontispiece. Isis, 93(1):x, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/343362.

Rossiter:2002:BOA

[954] Margaret W. Rossiter. Isis online—access and “angels”. Isis, 93(1): ix, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/343241.

Wolfe:2002:GSJ

[955] Audra J. Wolfe. Germs in space: Joshua Lederberg, exobiology, and the public imagination, 1958–1964. Isis, 93(2):183–205, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/344962.

Li:2002:NHP

[956] Shang-Jen Li. Natural history of parasitic disease: Patrick Manson’s philosophical method. Isis, 93(2):206–228, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/344961.

Kaiser:2002:NDP

[957] David Kaiser. Nuclear democracy: Political engagement, pedagogical reform, and particle physics in postwar America. Isis, 93(2):229–268, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/344960.

Rothenberg:2002:EAE

[958] Richard F. Hirsh Arthur L. Norberg Marc Rothenberg. Eloge: Albert Earl Moyer, 5 March 1945—19 December 2000. Isis, 93(2):269–271, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/344964.

Croce:2002:UEW

[959] Paul Jerome Croce. A useful eccentricity: William James’s engagement with science: William James: The correspondence of William James. Isis, 93(2):272–276, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344963.

Baldwin:2002:BRM

[960] Martha Baldwin. Book review: Michael Hunter: Robert Boyle (1627– 91): Scrupulosity and Science. Isis, 93(2):277–279, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/344966.

Weldon:2002:BRP

[961] Stephen P. Weldon. Book review: Peter Whitfield: Landmarks in Western Science: From Prehistory to the Atomic Age. Isis, 93(2):279–280, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344967.

Solovey:2002:BRT

[962] Mark Solovey. Book review: Toby A. Appel: Shaping Biology: The National Science Foundation and American Biological Research, 1945– 1975. Isis, 93(2):280–281, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/344968.

Shteir:2002:BRB

[963] Ann B. Shteir. Book review: Books and the Sciences in History. Isis, 93(2):281–282, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344969.

Chianese:2002:BRM

[964] Robert Chianese. Book review: Mary Midgley: Science and Poetry. Isis, 93(2):282–283, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344970.

Strauss:2002:BRE

[965] David Strauss. Book review: Ewen A. Whitaker: Mapping and Naming the Moon: A History of Lunar Cartography and Nomenclature. Isis, 93(2):283–284, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344971.

Burbidge:2002:BRJ

[966] Margaret Burbidge. Book review: John Gribbin: The Birth of Time: How Astronomers Measured the Age of the Universe. Isis, 93(2):284– 285, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/344972.

Pankenier:2002:BRB

[967] David W. Pankenier. Book review: Astronomy across Cultures: The History of Non-Western Astronomy. Isis, 93(2):285–286, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/344973.

Heilbron:2002:BRBa

[968] J. L. Heilbron. Book review: Sur les traces des Cassini: Astronomes et observatoires du sud de la France. Isis, 93(2):286–287, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/344974.

Bork:2002:BRBa

[969] Kennard B. Bork. Book review: Un cordée originale: Histoire des relations entre science et montagne. Isis, 93(2):287–288, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/344975.

Christianson:2002:BRB

[970] Gale E. Christianson. Book review: Brian Fagan: The Little Ice Age: How Climate Made History, 1300–1850. Isis, 93(2):288–289, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/344976.

Jankovic:2002:BRM

[971] Vladimir Janković. Book review: Mark Monmonier: Air Apparent: How Meteorologists Learned to Map, Predict, and Dramatize Weather. Isis, 93(2):289–290, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344977.

Yochelson:2002:BRP

[972] Ellis L. Yochelson. Book review: Peter Douglas Ward: Rivers in Time: The Search for Clues to Earth’s Mass Extinctions. Isis, 93(2):290, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/344978.

Abraham:2002:BRL

[973] Tara H. Abraham. Book review: Louise H. Marshall; Horace W. Magoun: Discoveries in the Human Brain: Neuroscience Prehistory, Brain Structure, and Function. Isis, 93(2):290–291, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/344979.

Anker:2002:BRJ

[974] Peder Anker. Book review: Johannes Fabian: Out of Our Minds: Reason and Madness in the Exploration of Central Africa. Isis, 93(2):291–292, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344980.

Critchlow:2002:BRB

[975] Donald T. Critchlow. Book review: Coping with Sickness: Medicine, Law, and Human Rights-Historical Perspectives. Isis, 93(2):292–293, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344981.

Gouk:2002:BRB

[976] Penelope Gouk. Book review: Une histoire de l’acoustique musicale Music Theory and Natural Order from the Renaissance to the Early Twentieth Century. Isis, 93(2):293–294, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/344982.

Donohue:2002:BRS

[977] A. A. Donohue. Book review: Stephen L. Dyson: Ancient Marbles to American Shores: Classical Archaeology in the United States. Isis, 93(2): 294–295, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/344983.

Saito:2002:BRS

[978] Ken Saito. Book review: Serafina Cuomo: Ancient Mathematics. Isis, 93(2):295–296, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344984.

Renshaw:2002:BRZ

[979] Steven Renshaw. Book review: Zhentao Xu; David W. Pankenier; Yaotiao Jiang: East Asian Archaeoastronomy: Historical Records of Astronomical Observations of China, Japan, and Korea. Isis, 93(2):296–297, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344985.

Declercq:2002:BRL

[980] Georges Declercq. Book review: Lynn E. Rose: Sun, Moon, and Sothis: A Study of Calendars and Calendar Reforms in Ancient Egypt. Isis, 93(2):297–298, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344986.

Houston:2002:BRM

[981] George W. Houston. Book review: M. J. T. Lewis: Surveying Instruments of Greece and Rome. Isis, 93(2):298–299, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/344987.

Cross:2002:BRO

[982] Richard Cross. Book review: Olivier Boulnois: Être et représentation: Une généalogie de la métaphysique moderne à l’époque de Duns Scot (XIIIe–XIVe siècle). Isis, 93(2):299–300, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/344988.

Laird:2002:BRB

[983] W. R. Laird. Book review: Renaissance Readings of the. Isis, 93(2):300– 301, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/344989.

Schildgen:2002:BRA

[984] Brenda Deen Schildgen. Book review: Alison Cornish: Reading Dante’s Stars. Isis, 93(2):301–302, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/344990.

Burns:2002:BRJ

[985] William E. Burns. Book review: Joan-Pau Rubiés: Travel and Ethnology in the Renaissance: South India through European Eyes, 1250–1625. Isis, 93(2):302–303, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
344991.

Pettigrew:2002:BRJ

[986] Todd Pettigrew. Book review: John Crawford Adams: Shakespeare’s Physic. Isis, 93(2):303, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/344992.

Jacob:2002:BRP

[987] James R. Jacob. Book review: Peter Dear: Revolutionizing the Sciences: European Knowledge and Its Ambitions, 1500–1700. Isis, 93(2):303–304, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344993.

Stewart:2002:BRM

[988] Larry Stewart. Book review: Margaret C. Jacob: Scientific Culture and the Making of the Industrial West. Isis, 93(2):304–305, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/344994.

Fara:2002:BRV

[989] Patricia Fara. Book review: Vladimir Janković: Reading the Skies: A Cultural History of English Weather, 1650–1820. Isis, 93(2):305–306, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/344995.

Newell:2002:BRE

[990] Julie R. Newell. Book review: Edward J. Cashin: William Bartram and the American Revolution on the Southern Frontier. Isis, 93(2):306, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/344996.

Morus:2002:BRJ

[991] Iwan Rhys Morus. Book review: June Z. Fullmer: Young Humphry Davy: The Making of an Experimental Chemist. Isis, 93(2):306–307, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/344997.

Parmenter:2002:BRA

[992] Jon Parmenter. Book review: Anthony F. C. Wallace: Jefferson and the Indians: The Tragic Fate of the First Americans. Isis, 93(2):307– 308, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/344998.

Evenden:2002:BRJ

[993] Doreen Evenden. Book review: Jane Sharp: The Midwives Book; or, The Whole Art of Midwifery Discovered. Isis, 93(2):308, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/344999.

Abbri:2002:BRF

[994] Ferdinando Abbri. Book review: Francesco Paolo de Ceglia: Introduzione alla fisiologia di Georg Ernst Stahl. Isis, 93(2):309, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345000.

Harrison:2002:BRBb

[995] Mark Harrison. Book review: Barbara Howard Traister: Notorious Astrological Physician of London: Works and Days of Simon Forman. Isis, 93(2):309–310, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
345001.

Sheffield:2002:BRA

[996] Suzanne Le-May Sheffield. Book review: Alan Rauch: Useful Knowledge: The Victorians, Morality, and the March of Intellect. Isis, 93(2):310– 311, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345002.

Porter:2002:BRC

[997] Theodore M. Porter. Book review: Christopher Herbert: Victorian Relativity: Radical Thought and Scientific Discovery. Isis, 93(2):311–312, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/345003.

Jonsson:2002:BRF

[998] Inge Jonsson. Book review: Francesca Maria Crasta: La filosofia della natura di Emanuel Swedenborg. Isis, 93(2):312–313, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345004.

Finn:2002:BRN

[999] Michael R. Finn. Book review: Nicola Luckhurst: Science and Structure in Proust’s. Isis, 93(2):313–314, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345005.

Churchill:2002:BRJ

[1000] Frederick B. Churchill. Book review: James A. Secord: Victorian Sensation: The Extraordinary Publication, Reception, and Secret Authorship of Vestiges of the Natural History of Creation. Isis, 93(2):314–315, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345006.

Rocke:2002:BRS

[1001] Alan J. Rocke. Book review: Sharon McGrayne: Prometheans in the Lab: Chemistry and the Making of the Modern World. Isis, 93(2):315– 316, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345007.

Silliman:2002:BRJ

[1002] Robert H. Silliman. Book review: James G. Cassidy: Ferdinand V. Hayden: Entrepreneur of Science. Isis, 93(2):316–317, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345008.

Frawley:2002:BRL

[1003] Maria H. Frawley. Book review: Lila Marz Harper: Solitary Travelers: Nineteenth-Century Women’s Travel Narratives and the Scientific Vocation. Isis, 93(2):317–318, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345009.

Podgorny:2002:BRB

[1004] Irina Podgorny. Book review: Beatrice Häsler; Thomas W. Baumann: Henri Pittier, 1857–1950: Leben und Werk eines Schweizer Naturforschers in den Neotropen. Isis, 93(2):318–319, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345010.

VanRiper:2002:BRP

[1005] A. Bowdoin Van Riper. Book review: Pat Shipman: The Man Who Found the Missing Link: Eugène Dubois and His Lifelong Quest to Prove Darwin Right. Isis, 93(2):319–320, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345011.

Faber:2002:BRB

[1006] Diana P. Faber. Book review: Benoit Marpeau: Gustave Le Bon: Parcours d’un intellectuel 1841–1931. Isis, 93(2):320–321, June 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345012.

Turner:2002:BRT

[1007] R. Steven Turner. Book review: Tracey L. Adams: A Dentist and a Gentleman: Gender and the Rise of Dentistry in Ontario. Isis, 93(2): 321, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345013.

Grob:2002:BRB

[1008] Gerald N. Grob. Book review: Traumatic Pasts: History, Psychiatry, and Trauma in the Modern Age, 1870–1930. Isis, 93(2):322, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345014.

Connor:2002:BRJ

[1009] Jennifer J. Connor. Book review: John S. Haller, Jr. The People’s Doctors: Samuel Thomson and the American Botanical Movement, 1790– 1860. Isis, 93(2):322–323, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345015.

Constant:2002:BRD

[1010] Ed Constant. Book review: David A. Mindell: War, Technology, and Experience Aboard the USS. Isis, 93(2):323–324, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345016.

Mirowski:2002:BRM

[1011] Philip Mirowski. Book review: Malachi Haim Hacohen: Karl Popper: The Formative Years, 1902–45: Politics and Philosophy in Interwar Vienna. Isis, 93(2):324–325, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345017.

Cassidy:2002:BRE

[1012] James G. Cassidy. Book review: Ellis L. Yochelson: Smithsonian Institution Secretary, Charles Doolittle Walcott. Isis, 93(2):325–326, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345018.

Rowe:2002:BRB

[1013] David E. Rowe. Book review: Hermann Weyl’s Raum–Zeit–Materie and a General Introduction to His Scientific Work. Isis, 93(2):326–327, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345019.

McLarty:2002:BRA

[1014] Colin McLarty. Book review: Alain Herreman: La topologie et ses signes: Éléments pour une histoire sémiotique des mathématiques. Isis, 93(2): 328, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345020.

Mazzotti:2002:BRJ

[1015] Massimo Mazzotti. Book review: Javier de Lorenzo: La matematica: De sus fundamentos y crisis. Isis, 93(2):328–329, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345021.

Stranges:2002:BRV

[1016] Anthony N. Stranges. Book review: Vaclav Smil: Enriching the Earth: Fritz Haber, Carl Bosch, and the Transformation of World Food Production. Isis, 93(2):329–330, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345022.

Ramberg:2002:BRU

[1017] Peter J. Ramberg. Book review: Ulrike Fell: Disziplin, Profession und Nation: Die Ideologie der Chemie in Frankreich vom Zweiten Kaiserreich bis in die Zwischenkriegszeit. Isis, 93(2):330–331, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345023.

Benson:2002:BRG

[1018] Maxine Benson. Book review: Gwyneth Hoyle: Flowers in the Snow: The Life of Isobel Wylie Hutchison. Isis, 93(2):331, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345024.

Mills:2002:BRG

[1019] Eric L. Mills. Book review: Gary E. Weir: An Ocean in Common: American Naval Officers, Scientists, and the Ocean Environment. Isis, 93 (2):332, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345025.

Fleming:2002:BRW

[1020] James Rodger Fleming. Book review: William B. Meyer: Americans and Their Weather. Isis, 93(2):332–333, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345026.

Benson:2002:BRJ

[1021] Keith R. Benson. Book review: Jan Witkowski: Illuminating Life: Selected Papers from Cold Spring Harbor (1903–1969). Isis, 93(2):333–334, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/345027.

Graebner:2002:BRR

[1022] William Graebner. Book review: Rita Caccamo: Back to Middletown: Three Generations of Sociological Reflections. Isis, 93(2):334–335, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345028.

Golan:2002:BRS

[1023] Tal Golan. Book review: Simon A. Cole: Suspect Identities: A History of Fingerprinting and Criminal Identification. Isis, 93(2):335–336, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345029.

Barnes:2002:BRL

[1024] J. A. Barnes. Book review: Lyn Schumaker: Africanizing Anthropology: Fieldwork, Networks, and the Making of Cultural Knowledge in Central Africa. Isis, 93(2):336–337, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345030.

Acker:2002:BRD

[1025] Caroline Jean Acker. Book review: David Patrick Keys; John F. Galliher: Confronting the Drug Control Establishment: Alfred Lindesmith as a Public Intellectual. Isis, 93(2):337–338, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/345031.

Eghigian:2002:BRP

[1026] Greg Eghigian. Book review: Paul Julian Weindling: Epidemics and Genocide in Eastern Europe, 1890–1945. Isis, 93(2):338–339, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345032.

Kline:2002:BRJ

[1027] Ronald R. Kline. Book review: Joy Elizabeth Hayes: Radio Nation: Communication, Popular Culture, and Nationalism in Mexico, 1920– 1945. Isis, 93(2):339–340, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345033.

Gerber:2002:BREa

[1028] Michele S. Gerber. Book review: Edmund Russell: War and Nature: Fighting Humans and Insects with Chemicals from World War I to “Silent Spring.”. Isis, 93(2):340–341, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345034.

Abbate:2002:BRB

[1029] Janet Abbate. Book review: The First Computers: History and Architectures. Isis, 93(2):341, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345035.

Belton:2002:BRS

[1030] John Belton. Book review: Semi Joseph Begun: Magnetic Recording: The Ups and Downs of a Pioneer: The Memoirs of Semi Joseph Begun. Isis, 93(2):341–343, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
345036.

Hacker:2002:BRT

[1031] Barton C. Hacker. Book review: Timothy Moy: War Machines: Transforming Technologies in the U.S. Military, 1920–1940. Isis, 93(2):343, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/345037.

Wang:2002:BRA

[1032] Zuoyue Wang. Book review: Allan A. Needell: Science, Cold War, and the American State: Lloyd V. Berkner and the Balance of Professional Ideals. Isis, 93(2):343–345, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345038.

Krige:2002:BRD

[1033] John Krige. Book review: David Leverington: New Cosmic Horizons: Space Astronomy from the V2 to the Hubble Space Telescope. Isis, 93 (2):345, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345039.

Oldroyd:2002:BRC

[1034] David Oldroyd. Book review: Constantin Roman: Continental Drift: Colliding Continents, Converging Cultures. Isis, 93(2):345–346, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345040.

Pyne:2002:BRS

[1035] Stephen J. Pyne. Book review: Seth R. Reice: The Silver Lining: The Benefits of Natural Disasters. Isis, 93(2):346–347, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345041.

Madison:2002:BRR

[1036] Mark Madison. Book review: Richard A. Rajala: Clearcutting the Pacific Rain Forest: Production, Science, and Regulation. Isis, 93(2):347–348, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/345042.

Larson:2002:BRJ

[1037] Allan Larson. Book review: John Alcock: The Triumph of Sociobiology. Isis, 93(2):348–349, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
345043.

Crumpton:2002:BRL

[1038] Amy C. Crumpton. Book review: Leslie R. Alm: Crossing Borders, Crossing Boundaries: The Role of Scientists in the U.S. Acid Rain Debate. Isis, 93(2):349–350, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345044.

Vickers:2002:BRD

[1039] John M. Vickers. Book review: D. Wade Hands: Reflection without Rules: Economic Methodology and Contemporary Science Theory. Isis, 93(2):350, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
345045.

OConnor:2002:BRJ

[1040] Erin O’Connor. Book review: Jessica Snyder Sachs: Corpse: Nature, Forensics, and the Struggle to Pinpoint Time of Death: An Exploration of the Haunting Science of Forensic Ecology. Isis, 93(2):350–351, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345046.

Schull:2002:BRR

[1041] William J. Schull. Book review: Richard F. Mould: Chernobyl Record: The Definitive History of the Chernobyl Catastrophe. Isis, 93(2):351– 352, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345047.

Hapke:2002:BRM

[1042] Thomas Hapke. Book review: Michael Friedewald: Der Computer als Werkzeug und Medium: Die geistigen und technischen Wurzeln des Personalcomputers. Isis, 93(2):352–353, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345048.

Grier:2002:BRB

[1043] David Alan Grier. Book review: Systems, Experts, and Computers: The Systems Approach in Management and Engineering, World War II and After. Isis, 93(2):353–354, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345049.

Bauer:2002:BRT

[1044] Henry H. Bauer. Book review: Toby Smith: Little Gray Men: Roswell and the Rise of a Popular Culture. Isis, 93(2):354–355, June 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345050.

Schorn:2002:BRK

[1045] Ronald A. Schorn. Book review: Karl T. Pflock: Roswell: Inconvenient Facts and the Will to Believe. Isis, 93(2):355–356, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345051.

Christenson:2002:BRB

[1046] Andrew L. Christenson. Book review: Assembling the Past: Studies in the Professionalization of Archaeology. Isis, 93(2):356, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345052.

Lang:2002:BRR

[1047] Harry G. Lang. Book review: Ruth O’Brien: Crippled Justice: The History of Modern Disability Policy in the Workplace. Isis, 93(2):356– 357, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345053.

Matchett:2002:BRR

[1048] Karin Matchett. Book review: Roberto J. González: Zapotec Science: Farming and Food in the Northern Sierra of Oaxaca. Isis, 93(2):357– 358, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345054.

Guala:2002:BRG

[1049] Francesco Guala. Book review: Giuseppe Giordano: Tra paradigmi e rivoluzioni: Thomas Kuhn. Isis, 93(2):358–359, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345055.

Fuller:2002:BRI

[1050] Steve Fuller. Book review: Isabelle Stengers: The Invention of Modern Science. Isis, 93(2):359–361, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/345056.

Paris:2002:BRR

[1051] Elizabeth Paris. Book review: Robert P. Crease: Making Physics: A Biography of Brookhaven National Laboratory, 1946–1972. Isis, 93(2): 361–362, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345057.

Cahan:2002:BRB

[1052] David Cahan. Book review: Institut im Bild, Volume 2: Bauten der Max-Planck-Gesellschaft zur Förderung der Wissenschaften. Isis, 93(2): 362, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345058.

Teich:2002:BRS

[1053] Albert H. Teich. Book review: Stephen Hilgartner: Science on Stage: Expert Advice as Public Drama. Isis, 93(2):362–363, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345059.

Rattansi:2002:BRR

[1054] P. M. Rattansi. Book review: R. W. Hunt; A. G. Watson; W. D. Macray: Bodleian Library Quarto Catalogues IX: Digby Manuscripts. Isis, 93(2): 363–364, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345060.

Horden:2002:BRB

[1055] Peregrine Horden. Book review: Scientific and Medical Writings in Old and Middle English: An Electronic Reference. Isis, 93(2):364, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345061.

Beaver:2002:BRBa

[1056] Donald deB. Beaver. Book review: American Science Leaders. Isis, 93 (2):365, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345062.

Singer:2002:BRF

[1057] Bayla Singer. Book review: Friends of the United States Air Force Academy Library: The Genesis of Flight: The Aeronautical History Collection of Colonel Richard Gimbel. Isis, 93(2):365–366, June 2002.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345063.

Anonymous:2002:BRBk

[1058] Anonymous. Book review: Technologies of Power: Essays in Honor of Thomas Parke Hughes and Agatha Chipley Hughes. Isis, 93(2):366– 367, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/345064.

Anonymous:2002:BRBl

[1059] Anonymous. Book review: Historical Aspects of Unconventional Medicine: Approaches, Concepts, Case Studies. Isis, 93(2):367, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/345065.

Anonymous:2002:BRBm

[1060] Anonymous. Book review: Millenarianism and Messianism in Early Modern European Culture. Isis, 93(2):367, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/345066.

Anonymous:2002:BRBn

[1061] Anonymous. Book review: Selection Theory and Social Construction: The Evolutionary Naturalistic Epistemology of Donald T, by Cecilia Heyes; David L. Hull. Campbell. Isis, 93(2):367, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345067.

Anonymous:2002:BRBo

[1062] Anonymous. Book review: Technology and the Rest of Culture. Isis, 93(2):367–368, June 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
345068.

Anonymous:2002:BRBp

[1063] Anonymous. Book review: Why Animal Experimentation Matters: The Use of Animals in Medical Research. Isis, 93(2):368, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345069.

Anonymous:2002:BRBq

[1064] Anonymous. Book review: Changing the Atmosphere: Expert Knowledge and Environmental Governance. Isis, 93(2):368, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/345070.

Anonymous:2002:BRBr

[1065] Anonymous. Book review: Histories of the Electron: The Birth of Microphysics. Isis, 93(2):368, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/345071.

Anonymous:2002:NCb

[1066] Anonymous. Notes on contributors. Isis, 93(2):369–375, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/345072.

Anonymous:2002:Fb

[1067] Anonymous. Frontispiece. Isis, 93(2):x, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/345073.

Buhs:2002:FAW

[1068] Joshua Blu Buhs. The fire ant wars: Nature and science in the pesticide controversies of the Late Twentieth Century. Isis, 93(3):377–400, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/512523.

Fontaine:2002:BPS

[1069] Philippe Fontaine. Blood, politics, and social science: Richard Titmuss and the Institute of Economic Affairs, 1957–1973. Isis, 93(3):401–434, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374061.

Neal:2002:MEN

[1070] Katherine Neal. Mathematics and empire, navigation and exploration: Henry Briggs and the Northwest Passage voyages of 1631. Isis, 93(3): 435–453, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374062.

Finn:2002:EBH

[1071] Bernard Finn. Eloge: Brooke Hindle, 1918–2001. Isis, 93(3):454–456, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374063.

Rowe:2002:EDJ

[1072] David E. Rowe. Eloge: Dirk Jan Struik, 1894–2000. Isis, 93(3):456–459, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374064.

Singh:2002:LE

[1073] Rajinder Singh. Letter to the Editor. Isis, 93(3):460–461, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374065.

Smocovitis:2002:BRB

[1074] Vassiliki Betty Smocovitis. Book review: Debating Darwin: Adventures of a Scholar. Isis, 93(3):462–464, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374066.

Newcomb:2002:BRA

[1075] Sally Newcomb. Book review: Alwyn Scarth: Vulcan’s Fury: Man Against the Volcano. Isis, 93(3):465–466, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374067.

Jacyna:2002:BRH

[1076] L. S. Jacyna. Book review: Howard I. Kushner: A Cursing Brain? The History of Tourette Syndrome. Isis, 93(3):466–467, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374068.

Good:2002:BRM

[1077] Gregory A. Good. Book review: Margaret Hindle Hazen; James Trefil: Good Seeing: A Century of Science at the Carnegie Institution of Washington. Isis, 93(3):467, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374069.

Glick:2002:BRD

[1078] Thomas F. Glick. Book review: Dale J. Pratt: Signs of Science: Literature, Science, and Spanish Modernity since 1868. Isis, 93(3):467–468, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374070.

Rigden:2002:BRW

[1079] John S. Rigden. Book review: William H. Cropper: Great Physicists: The Life and Times of Leading Physicists from Galileo to Hawking. Isis, 93(3):468–469, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374071.

Siegel:2002:BRP

[1080] Daniel Siegel. Book review: Prabhakar Gondhalekar: The Grip of Gravity: The Quest to Understand the Laws of Motion and Gravitation. Isis, 93(3):469–470, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374072.

Brice:2002:BRB

[1081] William R. Brice. Book review: The Age of the Earth: From 4004 B.C. to A.D. 2002, edited by S. J. Knell and C. L. E. Lewis. Isis, 93(3):470, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374073.

Burns:2002:BRB

[1082] William E. Burns. Book review: Nature and Empire: Science and the Colonial Enterprise. Isis, 93(3):470–471, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/374074.

McMahon:2002:BRS

[1083] Susan McMahon. Book review: Stanley H. Johnston, Jr. Cleveland’s Treasures from the World of Botanical Literature. Isis, 93(3):471–473, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374075.

Butrica:2002:BRJ

[1084] Andrew J. Butrica. Book review: Jean Boulaine; Jean-Paul Legros: D’Olivier de Serres à René Dumont: Portraits d’agronomes. Isis, 93 (3):473, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374076.

Rose:2002:BRB

[1085] Martha Rose. Book review: Points of Contact: Disability, Art, and Culture. Isis, 93(3):473–475, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374077.

Simonton:2002:BRM

[1086] Dean Keith Simonton. Book review: Michael J. A. Howe: Genius Explained. Isis, 93(3):475, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374078.

Maines:2002:BRJ

[1087] Rachel Maines. Book review: Jean Stengers; Anne Van Neck: Masturbation: The History of a Great Terror. Isis, 93(3):475–476, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374079.

Beaver:2002:BRBb

[1088] Donald deB. Beaver. Book review: The Social Shaping of Technology. Isis, 93(3):476–477, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374080.

Knight:2002:BRK

[1089] Mary Knight. Book review: Kamal Sabri Kolta; Doris Schwarzmann-Schafhauser: Die Hielkunde im alten Ägypten: Magie und Ratio in der Kranksheitsvorstellung und therapeutischen Praxis. Isis, 93(3):477–478, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374081.

Cleary:2002:BRP

[1090] John J. Cleary. Book review: Plato: Plato’s Timaeus. Isis, 93(3): 478–479, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374082.

Selin:2002:BRG

[1091] Helaine Selin. Book review: G. Jan Meulenbeld: A History of Indian Medical Literature. Isis, 93(3):479, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/374083.

Raphals:2002:BRC

[1092] Lisa Raphals. Book review: Charlotte Furth: A Flourishing Yin: Gender in China’s Medical History, 960–1665. Isis, 93(3):479–480, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374084.

Barnes:2002:BRH

[1093] Robin B. Barnes. Book review: Howard Hotson: Paradise Postponed: Johann Heinrich Alsted and the Birth of Calvinist Millenarianism. Isis, 93 (3):480–481, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374085.

Long:2002:BRP

[1094] Pamela O. Long. Book review: Paolo Rossi: The Birth of Modern Science. Isis, 93(3):481–482, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374086.

Harrison:2002:BRK

[1095] Peter Harrison. Book review: Kenneth James Howell: God’s Two Books: Copernican Cosmology and Biblical Interpretation in Early Modern Science. Isis, 93(3):482–483, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374087.

Steinhardt:2002:BRL

[1096] Nancy Shatzman Steinhardt. Book review: Laura Hostetler: Qing Colonial Enterprise: Ethnography and Cartography in Early Modern China. Isis, 93(3):483–484, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374088.

Zagorin:2002:BRS

[1097] Perez Zagorin. Book review: Stephen Gaukroger: Francis Bacon and the Transformation of Early-Modern Philosophy. Isis, 93(3):484–485, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374089.

Palmer:2002:BRP

[1098] Eric Palmer. Book review: Peter A. Schouls: Descartes and the Possibility of Science. Isis, 93(3):485–486, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/374090.

Lustig:2002:BRR

[1099] A. J. Lustig. Book review: Roger L. Williams: Botanophilia in Eighteenth-Century France: The Spirit of the Enlightenment. Isis, 93 (3):486–487, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374091.

Guerrini:2002:BRJ

[1100] Anita Guerrini. Book review: Jeff Loveland: Rhetoric and Natural History: Buffon in Polemical and Literary Context. Isis, 93(3):487, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374092.

Wilson:2002:BRR

[1101] Jon Wilson. Book review: Richard Drayton: Nature’s Government: Science, Imperial Britain, and the “Improvement” of the World. Isis, 93 (3):487–489, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374093.

Klein:2002:BRM

[1102] Lawrence E. Klein. Book review: Miles Ogborn: Spaces of Modernity: London’s Geographies, 1680–1780. Isis, 93(3):489–490, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374094.

Duchesneau:2002:BRB

[1103] François Duchesneau. Book review: Albrecht von Haller: Commentarius de formatione cordis in ovo incubato. Isis, 93(3):490–491, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374095.

Porter:2002:BRN

[1104] Theodore M. Porter. Book review: Nicholas Wright Gillham: A Life of Sir Francis Galton: From African Exploration to the Birth of Eugenics. Isis, 93(3):491–492, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374096.

Cosslett:2002:BRC

[1105] Tess Cosslett. Book review: Christine Kenyon-Jones: Kindred Brutes: Animals in Romantic-Period Writing. Isis, 93(3):492–493, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374097.

Marti:2002:BRB

[1106] Oscar R. Marti. Book review: Writings of Charles S, by Charles S. Peirce. Peirce: A Chronological Edition. Isis, 93(3):493, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374098.

Patton:2002:BRP

[1107] Cynthia Ellen Patton. Book review: Phyllis Weliver: Women Musicians in Victorian Fiction, 1860–1900: Representations of Music, Science, and Gender in the Leisured Home. Isis, 93(3):494, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374099.

Smith:2002:BRS

[1108] Sherry L. Smith. Book review: Shawn Michelle Smith: American Archives: Gender, Race, and Class in Visual Culture. Isis, 93(3): 494–495, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374100.

Beyler:2002:BRF

[1109] Richard H. Beyler. Book review: Freddy Litten: Mechanik und Antisemitismus: Wilhelm Müller (1880–1968). Isis, 93(3):495–496, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374101.

Mehigan:2002:BRB

[1110] Tim Mehigan. Book review: Ernst Mach’s Vienna, 1895–1930; or, Phenomenalism as Philosophy of Science. Isis, 93(3):496–497, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374102.

Guillemin:2002:BRS

[1111] Jeanne Guillemin. Book review: Soraya Boudia: Marie Curie et son laboratoire: Sciences et industrie de la radioactivité en France. Isis, 93 (3):497–498, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374103.

Marche:2002:BRD

[1112] Jordan D. Marché, II. Book review: David Strauss: Percival Lowell: The Culture and Science of a Boston Brahmin. Isis, 93(3):498–499, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374104.

Morrison:2002:BRW

[1113] Margaret Morrison. Book review: William Whewell: Of the Plurality of Worlds. Isis, 93(3):499–500, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374105.

McCook:2002:BRD

[1114] Stuart McCook. Book review: D. Graham Burnett: Masters of All They Surveyed: Exploration, Geography, and a British El Dorado. Isis, 93 (3):500, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374106.

Farber:2002:BRJ

[1115] Paul Lawrence Farber. Book review: Jeremy Gaskell: Who Killed the Great Auk? Isis, 93(3):500–501, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374107.

Gelfand:2002:BRJ

[1116] Toby Gelfand. Book review: L. S. Jacyna: Lost Words: Narratives of Language and the Brain, 1825–1926. Isis, 93(3):501–502, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374108.

Hoyt:2002:BRA

[1117] David Hoyt. Book review: Andrew Zimmerman: Anthropology and Antihumanism in Imperial Germany. Isis, 93(3):502–503, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374109.

Kingsland:2002:BRP

[1118] Sharon Kingsland. Book review: Philip Pauly: Biologists and the Promise of American Life: From Meriwether Lewis to Alfred Kinsey. Isis, 93(3):503–504, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374110.

Montgomery:2002:BRM

[1119] William Montgomery. Book review: S. M. Walters; E. A. Stow: Darwin’s Mentor: John Stevens Henslow, 1796–1861. Isis, 93(3):504–505, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374111.

Hayward:2002:BRH

[1120] Rhodri Hayward. Book review: Harry Oosterhuis: Stepchildren of Nature: Krafft-Ebing, Psychiatry, and the Making of Sexual Identity. Isis, 93(3):505–506, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374112.

Barton:2002:BRA

[1121] Ruth Barton. Book review: Alan Richardson: British Romanticism and the Science of the Mind. Isis, 93(3):506, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/374113.

Bonner:2002:BRS

[1122] Thomas Neville Bonner. Book review: Susan Wells: Out of the Dead House: Nineteenth-Century Women Physicians and the Writing of Medicine. Isis, 93(3):506–507, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374114.

Marcus:2002:BRB

[1123] Alan I. Marcus. Book review: Barbara Clow: Negotiating Disease: Power and Cancer Care, 1900–1950. Isis, 93(3):507–508, September 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374115.

Kushner:2002:BRW

[1124] Howard I. Kushner. Book review: Walter J. Friedlander: The History of Modern Epilepsy: The Beginnings, 1865–1914. Isis, 93(3):508–509, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374116.

Sturdy:2002:BRA

[1125] Steve Sturdy. Book review: Anne Hardy: Health and Medicine in Britain since 1860. Isis, 93(3):509, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374117.

Li:2002:BRD

[1126] Shang-Jen Li. Book review: Douglas M. Haynes: Imperial Medicine: Patrick Manson and the Conquest of Tropical Disease. Isis, 93(3): 510–511, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374118.

Hepler:2002:BRBa

[1127] Allison L. Hepler. Book review: Beatrix Hoffman: The Wages of Sickness: The Politics of Health Insurance in Progressive America. Isis, 93 (3):511–512, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374119.

Hansen:2002:BRL

[1128] LeeAnn Hansen. Book review: Lucy Hartley: Physiognomy and the Meaning of Expression in Nineteenth-Century Culture. Isis, 93(3): 512–513, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374120.

Fiege:2002:BRK

[1129] Mark Fiege. Book review: Kendrick A. Clements: Hoover, Conservation, and Consumerism: Engineering the Good Life. Isis, 93(3):513–514, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374121.

Mohun:2002:BRB

[1130] Arwen P. Mohun. Book review: From Fireplace to Cookstove: Technology and the Domestic Ideal in America Can She Bake a Cherry Pie? American Women and the Kitchen in the Twentieth Century. Isis, 93 (3):514–516, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374122.

Kidwell:2002:BRK

[1131] Peggy Aldrich Kidwell. Book review: Kathleen Broome Williams: Improbable Warriors: Women Scientists and the U.S. Navy in World War II. Isis, 93(3):516, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374123.

Krinsky:2002:BRJ

[1132] Carol Herselle Krinsky. Book review: Jason Goodwin: Otis: Giving Rise to the Modern City. Isis, 93(3):516–517, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374124.

Golland:2002:BRB

[1133] Louise Golland. Book review: Collected Works Collected Works Collected Works. Isis, 93(3):517–518, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374125.

Janiewski:2002:BRH

[1134] Dolores E. Janiewski. Book review: Hilary Lapsley: Margaret Mead and Ruth Benedict: The Kinship of Women. Isis, 93(3):518, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374126.

Hoffman:2002:BRA

[1135] Beatrix Hoffman. Book review: Allison L. Hepler: Women in Labor: Mothers, Medicine, and Occupational Health in the United States, 1890– 1980. Isis, 93(3):518–519, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374127.

Parlee:2002:BRB

[1136] Mary Brown Parlee. Book review: Ben Shephard: A War of Nerves: Soldiers and Psychiatrists in the Twentieth Century. Isis, 93(3):519–520, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374128.

Dawson:2002:BRT

[1137] Virginia P. Dawson. Book review: Tom D. Crouch: Aiming for the Stars: The Dreamers and Doers of the Space Age. Isis, 93(3):520–521, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374129.

Kirby:2002:BRE

[1138] David A. Kirby. Book review: Errol Vieth: Screening Science: Contexts, Texts, and Science in Fifties Science Fiction Film. Isis, 93(3):521–522, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374130.

Forman:2002:BRC

[1139] Paul Forman. Book review: Claude Audoin; Bernard Guinot: The Measurement of Time: Time, Frequency, and the Atomic Clock. Isis, 93 (3):522–523, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374131.

Hirshfeld:2002:BRW

[1140] Alan W. Hirshfeld. Book review: Wallace Tucker; Karen Tucker: Revealing the Universe: The Making of the Chandra X-Ray Observatory. Isis, 93(3):523–524, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374132.

Cittadino:2002:BRB

[1141] Eugene Cittadino. Book review: Betty Jean Craige: Eugene Odum: Ecosystem Ecologist and Environmentalist. Isis, 93(3):524–525, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374133.

Friedel:2002:BRM

[1142] Robert Friedel. Book review: Maynard Frank Wolfe: Rube Goldberg: Inventions. Isis, 93(3):525–526, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374134.

Goldman:2002:BRM

[1143] Joanne Abel Goldman. Book review: Martin V. Melosi: Effluent America: Cities, Industry, Energy, and the Environment. Isis, 93(3):526–527, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374135.

Gerber:2002:BREb

[1144] Michele S. Gerber. Book review: Eric W. Mogren: Warm Sands: Uranium Mill Tailings Policy in the American West. Isis, 93(3):527, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374136.

Morton:2002:BRB

[1145] David L. Morton, Jr. Book review: Restoring Baird’s Image John Logie Baird: Television Pioneer. Isis, 93(3):527–528, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374137.

Neufeld:2002:BRB

[1146] Michael J. Neufeld. Book review: A History of the European Space Agency, 1958–1987. Isis, 93(3):528–529, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/374138.

Rawson:2002:BRB

[1147] Timothy Rawson. Book review: Preserving Yellowstone’s Natural Conditions: Science and the Perception of Nature On Her Own Terms: Annie Montague Alexander and the Rise of Science in the American West. Isis, 93(3):529–530, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374139.

Hughes:2002:BRD

[1148] Sally Smith Hughes. Book review: Donald S. Fredrickson: The Recombinant DNA Controversy: A Memoir: Science, Politics, and the Public Interest, 1974–1981. Isis, 93(3):530–531, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374140.

Strauss:2002:BRD

[1149] Michael Strauss. Book review: Daniel Charles: Lords of the Harvest: Biotech, Big Money, and the Future of Food. Isis, 93(3):531–532, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374141.

Kenen:2002:BRA

[1150] Stephanie H. Kenen. Book review: Anne Fausto-Sterling: Sexing the Body: Gender Politics and the Construction of Sexuality. Isis, 93(3): 532–534, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374142.

Valone:2002:BRR

[1151] David A. Valone. Book review: Richard Lynn: Eugenics: A Reassessment. Isis, 93(3):534, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374143.

Fausto-Sterling:2002:BRD

[1152] Anne Fausto-Sterling. Book review: Deirdre McCloskey: Crossing: A Memoir. Isis, 93(3):534–535, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374144.

Howell:2002:BRC

[1153] Joel D. Howell. Book review: Carl Djerassi: This Man’s Pill: Reflections on the Fiftieth Birthday of the Pill. Isis, 93(3):535–536, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374145.

Hunt:2002:BRJ

[1154] Nancy Rose Hunt. Book review: John Iliffe: East African Doctors: A History of the Modern Profession. Isis, 93(3):536–537, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374146.

Jones:2002:BRK

[1155] David Jones. Book review: Kirk Jeffrey: Machines in Our Hearts: The Cardiac Pacemaker, the Implantable Defibrillator, and American Health Care. Isis, 93(3):537, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374147.

vanKeuren:2002:BRF

[1156] David van Keuren. Book review: Francis Duncan: Rickover: The Struggle for Excellence. Isis, 93(3):537–538, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374148.

Castaneda:2002:BRD

[1157] Christopher J. Castaneda. Book review: Davis Dyer; Daniel Gross: The Generations of Corning: The Life and Times of a Global Corporation. Isis, 93(3):538–539, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374149.

Badash:2002:BRB

[1158] Lawrence Badash. Book review: The Struggle Against the Bomb. Isis, 93 (3):539–540, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374150.

Kaiser:2002:BRD

[1159] David Kaiser. Book review: David Cesarani: Arthur Koestler: The Homeless Mind. Isis, 93(3):540–541, September 2002. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/374151.

Aiken:2002:BRR

[1160] Jane Andrews Aiken. Book review: Richard Padovan: Proportion: Science, Philosophy, Architecture. Isis, 93(3):541–542, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374152.

Solomon:2002:BRJ

[1161] Miriam Solomon. Book review: James Brown: Who Rules in Science? An Opinionated Guide to the Wars. Isis, 93(3):542–543, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374153.

Gal:2002:BRJ

[1162] Ofer Gal. Book review: J. E. McGuire; Barbara Tuchańska: Science Unfettered: A Philosophical Study in Sociohistorical Ontology. Isis, 93 (3):543–544, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374154.

Dorn:2002:BRJ

[1163] Harold Dorn. Book review: John V. Pickstone: Ways of Knowing: A New History of Science, Technology, and Medicine. Isis, 93(3):544, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374155.

Gieryn:2002:BRJ

[1164] Thomas F. Gieryn. Book review: John Ziman: Real Science: What It Is and What It Means. Isis, 93(3):544–545, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374156.

Brammall:2002:BRR

[1165] Kathryn M. Brammall. Book review: Richard S. Brooks; David K. Himrod: Science and Religion in the English-Speaking World, 1600–1727: A Bibliographic Guide to the Secondary Literature. Isis, 93(3):545–546, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374157.

Haller:2002:BRB

[1166] John S. Haller, Jr. Book review: An Annotated Catalogue of the Edward C, by Christopher Hoolihan. Atwater Collection of American Popular Medicine and Health Reform. Isis, 93(3):546–547, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374158.

Kaplan:2002:BRJ

[1167] Robert Kaplan. Book review: Joseph W. Dauben: The History of Mathematics from Antiquity to the Present: A Selective Annotated Bibliography. Isis, 93(3):547, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374159.

Fyfe:2002:BRB

[1168] Aileen Fyfe. Book review: The History of Science and Religion in the Western Tradition: An Encyclopedia. Isis, 93(3):547–548, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/374160.

Anonymous:2002:BRBs

[1169] Anonymous. Book review: “Defects”: Engendering the Modern Body. Isis, 93(3):548–549, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374161.

Anonymous:2002:BRBt

[1170] Anonymous. Book review: Luigi Galvani International Workshop. Isis, 93(3):549, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374162.

Anonymous:2002:BRBu

[1171] Anonymous. Book review: Cultures of Control. Isis, 93(3):549, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/374163.

Anonymous:2002:BRBv

[1172] Anonymous. Book review: Science in Islamic Civilization. Isis, 93 (3):549, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374164.

Anonymous:2002:BRBw

[1173] Anonymous. Book review: Medicine in the Italian Universities, 1250– 1600. Isis, 93(3):549–550, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374165.

Anonymous:2002:BRBaa

[1174] Anonymous. Book review: Science, Space, and Hermeneutics. Isis, 93 (3):550, September 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
374169.

Anonymous:2002:BRBx

[1175] Anonymous. Book review: Profiles in the History of the U, by Patricia J. Durana; Anne B. Effland; Douglas Helms.S. Soil Survey. Isis, 93(3):550, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374166.

Anonymous:2002:BRBy

[1176] Anonymous. Book review: New Perspectives on the Origins of Americanist Archaeology. Isis, 93(3):550, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374167.

Anonymous:2002:BRBz

[1177] Anonymous. Book review: The Environmental Tradition in English Literature. Isis, 93(3):550, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374168.

Anonymous:2002:BRBab

[1178] Anonymous. Book review: Berichte zur Geschichte der Hydro- und Meeresbiologie: Und weitere Beiträge zur 8, by Ekkehard Höxtermann; Joachim Kaasch; Michael Kaasch; Ragnar K. Kinzelbach. Jahrestagung der DGGTB in Rostock 1999. Isis, 93(3):551, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/374170.

Anonymous:2002:BRBac

[1179] Anonymous. Book review: The Blackwell Guide to the Philosophy of Science. Isis, 93(3):551, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/374171.

Anonymous:2002:BRBad

[1180] Anonymous. Book review: Learning to Manage Global Environmental Risks, by Social Learning Group; Bert Bolin. Volume 1: A Comparative History of Social Responses to Climate Change, Ozone Depletion, and Acid Rain; Volume 2: A Functional Analysis of Social Responses to Climate Change, Ozone Depletion, and Acid Rain. Isis, 93(3):551–552, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/374172.

Anonymous:2002:BRBae

[1181] Anonymous. Book review: American Philosophy of Technology: The Empirical Turn. Isis, 93(3):552, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/374173.

Anonymous:2002:NCc

[1182] Anonymous. Notes on contributors. Isis, 93(3):553–558, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/374174.

Anonymous:2002:E

[1183] Anonymous. Erratum. Isis, 93(3):viii, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/526486. In Anne A. Davenport’s re-

view of Steven P. Marrone’s book The Light of Thy Countenance in the March 2002 issue of Isis, the phrase “Late Thirteenth and Early Fourteenth Centuries” appeared incorrectly as “Late Eighteenth and Early Nineteenth Centuries.” Isis regrets the error.

Anonymous:2002:Fc

[1184] Anonymous. Frontispiece. Isis, 93(3):viii, September 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/367916.

Parascandola:2002:USF

[1185] Mark Parascandola. Uncertain science and a failure of trust: The NIH radioepidemiologic tables and compensation for radiation-induced cancer. Isis, 93(4):559–584, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/375952.

Canales:2002:PVC

[1186] Jimena Canales. Photogenic venus: The “cinematographic turn” and its alternatives in Nineteenth-Century France. Isis, 93(4):585–613, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/375953.

Schloegel:2002:GPE

[1187] Judy Johns Schloegel and Henning Schmidgen. General physiology, experimental psychology, and evolutionism: Unicellular organisms as objects of psychophysiological research, 1877–1918. Isis, 93(4):614–645, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375954.

Kleinman:2002:EJE

[1188] Kim Kleinman. Eloge: Joseph Ewan, 1909–1999, and Nesta Dunn Ewan, 1908–2000. Isis, 93(4):646–648, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375955.

Duke:2002:LE

[1189] Dennis W. Duke. Letter to the Editor. Isis, 93(4):649, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375956.

Gayon:2002:BRB

[1190] Jean Gayon. Book review: Philosophie de l’inconnu: Le vivant et la recherche. Isis, 93(4):650–654, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375957.

Tymoczko:2002:BRB

[1191] Maria Tymoczko. Book review: Science in Translation: Movements of Knowledge Through Cultures and Time. Isis, 93(4):655–657, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/375958.

Hepler:2002:BRBb

[1192] Allison L. Hepler. Book review: History of Women in the Sciences: Readings from. Isis, 93(4):658–659, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375959.

Grunden:2002:BRB

[1193] Walter E. Grunden. Book review: Current Perspectives in the History of Science in East Asia. Isis, 93(4):659–660, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375960.

Furth:2002:BRB

[1194] Charlotte Furth. Book review: Beyond Joseph Needham: Science, Technology, and Medicine in East and Southeast Asia. Isis, 93(4):660–662, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375961.

King:2002:BRB

[1195] William King. Book review: Race, Science, and Medicine, 1700–1960. Isis, 93(4):662, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
375962.

Kidwell:2002:BRBa

[1196] Peggy Aldrich Kidwell. Book review: Instruments in Print: Books from the Whipple Collection. Isis, 93(4):662–663, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375963.

Hoolihan:2002:BRB

[1197] Christopher Hoolihan. Book review: Biomedicine and Alternative Healing Systems in America: Issues of Class, Race, Ethnicity, and Gender. Isis, 93(4):663–664, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/375964.

Whitney:2002:BRB

[1198] Gordon Whitney. Book review: M. Agnoletti; S. Anderson (Editors): Methods and Approaches in Forest History. Isis, 93(4):664–665, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/375965.

Lynch:2002:BRB

[1199] William T. Lynch. Book review: Thomas Kuhn. Isis, 93(4):665–666, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375966.

Castle:2002:BRB

[1200] David Castle. Book review: Defenders of the Truth: The Battle for Science in the Sociobiology Debate and Beyond. Isis, 93(4):666–667, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/375967.

Artmann:2002:BRB

[1201] Benno Artmann. Book review: Geometry Civilized: History, Culture, and Technique. Isis, 93(4):667–668, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/375968.

Weiner:2002:BRB

[1202] Douglas R. Weiner. Book review: Academia in Upheaval: Origins, Transfers, and Transformations of the Communist Academic Regime in Russia and East Central Europe. Isis, 93(4):668–669, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375969.

Micale:2002:BRB

[1203] Mark S. Micale. Book review: Revels in Madness: Insanity in Medicine and Literature. Isis, 93(4):669–670, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/375970.

Bain:2002:BRB

[1204] Jonathan Bain. Book review: Seeing Double: Shared Identities in Physics, Philosophy, and Literature. Isis, 93(4):670–671, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375971.

Newcomb:2002:BRB

[1205] Sally Newcomb. Book review: Volcanoes in Human History: The Far-Reaching Effects of Major Eruptions. Isis, 93(4):671, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375972.

Bolt:2002:BRB

[1206] Marvin Bolt. Book review: Parallax: The Race to Measure the Cosmos. Isis, 93(4):671–672, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/375973.

Ouboter:2002:BRB

[1207] Rudolf de Bruyn Ouboter. Book review: Absolute Zero and the Conquest of Cold. Isis, 93(4):672–673, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375974.

Smith:2002:BRB

[1208] David C. Smith. Book review: A Good and Wise Measure: The Struggle for the Canadian-American Border, 1783–1842. Isis, 93(4):673–674, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375975.

Oslund:2002:BRB

[1209] Karen Oslund. Book review: A Fabulous Kingdom: The Exploration of the Arctic. Isis, 93(4):674–675, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375976.

Magner:2002:BRB

[1210] Lois N. Magner. Book review: The Rise of Experimental Biology: An Illustrated History. Isis, 93(4):675–676, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375977.

Kuklick:2002:BRB

[1211] Henrika Kuklick. Book review: History and Theory in Anthropology: Changing Perspectives. Isis, 93(4):676, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375978.

Lovelace:2002:BRB

[1212] Robert E. Lovelace. Book review: Minds Behind the Brain: A History of the Pioneers and Their Discoveries. Isis, 93(4):676–678, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375979.

Milne:2002:BRB

[1213] Antony Milne. Book review: The End of the Dinosaurs: Chicxulub Crater and Mass Extinctions. Isis, 93(4):678–679, December 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375980.

Brotherston:2002:BRB

[1214] Gordon Brotherston. Book review: Skywatchers: A Revised and Updated Version of Skywatchers of Ancient Mexico. Isis, 93(4):679–680, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375981.

Chippindale:2002:BRB

[1215] Christopher Chippindale. Book review: Stonehenge: A New Interpretation of Prehistoric Man and the Cosmos. Isis, 93(4):680–681, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/375982.

Dillon:2002:BRB

[1216] John Dillon. Book review: Timaeus. Isis, 93(4):681–682, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375983.

Gingerich:2002:BRB

[1217] Owen Gingerich. Book review: Planetary Systems from the Ancient Greeks to Kepler. Isis, 93(4):682, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/375984.

Bodson:2002:BRB

[1218] Liliane Bodson. Book review: The First Fossil Hunters: Paleontology in Greek and Roman Times. Isis, 93(4):682–683, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375985.

Nutton:2002:BRB

[1219] Vivian Nutton. Book review: Der Begriff der Physis bei Galen vor dem Hintergrund seiner Vorgänger. Isis, 93(4):683–684, December 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/375986.

Rutten:2002:BRB

[1220] Thomas Rütten. Book review: “On Second Thought” and Other Essays in the History of Medicine and Science. Isis, 93(4):684–685, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/375987.

Hanson:2002:BRB

[1221] Ann Ellis Hanson. Book review: On Head Wounds. Isis, 93(4):685–686, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375988.

Molland:2002:BRB

[1222] George Molland. Book review: Opera: Epistula de magnete; Nova compositio astrolabii particularis. Isis, 93(4):686, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375989.

Hackett:2002:BRB

[1223] Jeremiah Hackett. Book review: On Animals: A Medieval. Isis, 93 (4):686–688, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
375990.

Lemay:2002:BRB

[1224] Helen Rodnite Lemay. Book review: Women’s Healthcare in the Medieval West: Texts and Contexts. Isis, 93(4):688, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375991.

Morrison:2002:BRB

[1225] Robert Morrison. Book review: La science dans le monde iranien: À l’époque islamique. Isis, 93(4):688–690, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/375992.

Dold-Samplonius:2002:BRB

[1226] Yvonne Dold-Samplonius. Book review: Ibrahim ibn Sinan: Logique et géométrie au Xe siècle. Isis, 93(4):690–691, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375993.

Osler:2002:BRB

[1227] Margaret J. Osler. Book review: Religious Confessions and the Sciences in the Sixteenth Century. Isis, 93(4):691–692, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375994.

Gosselin:2002:BRB

[1228] Edward A. Gosselin. Book review: Oeuvres complètes. Isis, 93(4):692– 693, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/375995.

Heilbron:2002:BRBb

[1229] J. L. Heilbron. Book review: Solomon’s Child: Method in the Early Royal Society of London. Isis, 93(4):693, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/375996.

Hatfield:2002:BRB

[1230] Gary Hatfield. Book review: Vorlesungsverzeichnisse der Universität Königsberg (1720–1804). Isis, 93(4):693–694, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375997.

Malet:2002:BRB

[1231] Antoni Malet. Book review: Squaring the Circle: The War between Hobbes and Wallis. Isis, 93(4):694–695, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/375998.

Gross:2002:BRB

[1232] Alan Gross. Book review: Imagining the Unimaginable: The Poetics of Early Modern Astronomy. Isis, 93(4):695–696, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/375999.

Moran:2002:BRB

[1233] Bruce T. Moran. Book review: Chemistry and Medical Debate: Van Helmont to Boerhaave. Isis, 93(4):696–697, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376000.

Machamer:2002:BRB

[1234] Peter Machamer. Book review: Essays on Galileo and the History and Philosophy of Science. Isis, 93(4):697, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376001.

Gale:2002:BRB

[1235] George Gale. Book review: What Was Mechanical about Mechanics: The Concept of Force between Metaphysics and Mechanics from Newton to Lagrange. Isis, 93(4):697–698, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376002.

Kleinert:2002:BRB

[1236] Andreas Kleinert. Book review: Pushing Gravity: New Perspectives on Le Sage’s Theory of Gravitation. Isis, 93(4):698–699, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376003.

Hickel:2002:BRB

[1237] Erika Hickel. Book review: Pious Traders in Medicine: A German Pharmaceutical Network in Eighteenth-Century North America. Isis, 93(4): 699–700, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376004.

Panizza:2002:BRB

[1238] Letizia Panizza. Book review: Cymbalum mundi sive symbolum sapientiae: Edizione critica. Isis, 93(4):700–701, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376005.

Outram:2002:BRB

[1239] Dorinda Outram. Book review: How to Write the History of the New World: Historiographies, Epistemologies, and Identities in the Eighteenth-Century Atlantic World. Isis, 93(4):701–702, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376006.

Hart:2002:BRB

[1240] Sidney Hart. Book review: Philadelphia’s Enlightenment, 1740–1800: Kingdom of Christ, Empire of Reason. Isis, 93(4):702–703, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376007.

Guerrini:2002:BRB

[1241] Anita Guerrini. Book review: Fictions of Disease in Early Modern England: Bodies, Plagues, and Politics. Isis, 93(4):703–704, December 2002.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376008.

Pimentel:2002:BRB

[1242] Juan Pimentel. Book review: La actividad científica valenciana de la Ilustración. Isis, 93(4):704–705, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376009.

Bork:2002:BRBb

[1243] Kennard B. Bork. Book review: H, by René Sigrist.-B. de Saussure (1740–1799): Un regard sur la terre. Isis, 93(4):705–706, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376010.

Kidwell:2002:BRBb

[1244] Clara Sue Kidwell. Book review: The Myth of the Noble Savage. Isis, 93 (4):706–707, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376011.

MacMillan:2002:BRB

[1245] Ken MacMillan. Book review: Innocence Abroad: The Dutch Imagination and the New World, 1570–1670. Isis, 93(4):707–708, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376012.

Wilson:2002:BRB

[1246] Philip K. Wilson. Book review: Undertaker of the Mind: John Monro and Mad-Doctoring in Eighteenth-Century England. Isis, 93(4):708–709, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376013.

Olson:2002:BRBb

[1247] Richard Olson. Book review: Kant, Herder, and the Birth of Anthropology. Isis, 93(4):709–710, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376014.

Parezo:2002:BRB

[1248] Nancy J. Parezo. Book review: The Apache Diaries: A Father-Son Journey. Isis, 93(4):710–711, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376015.

Kaplan:2002:BRB

[1249] Robert Kaplan. Book review: The Annotated Flatland: A Romance of Many Dimensions. Isis, 93(4):711, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376016.

Kox:2002:BRB

[1250] A. J. Kox. Book review: On the Theory of the Reflection and Refraction of Light: Academic Dissertation for Acquiring the Degree of Doctor of Mathematical and Physical Sciences at Leyden University. Isis, 93(4): 711–712, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376017.

Yochelson:2002:BRB

[1251] Ellis L. Yochelson. Book review: The Coldest March: Scott’s Fatal Antarctic Expedition. Isis, 93(4):712–713, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376018.

Zeller:2002:BRB

[1252] Suzanne Zeller. Book review: The Politics of Population: State Formation, Statistics, and the Census of Canada, 1840–1875. Isis, 93(4):713– 714, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376019.

OConnor:2002:BRB

[1253] Erin O’Connor. Book review: A Traffic of Dead Bodies: Anatomy and Embodied Social Identity in Nineteenth-Century America. Isis, 93(4): 714–715, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376020.

Allen:2002:BRB

[1254] Ann T. Allen. Book review: Ärztinnen aus dem Kaiserreich: Lebensläufe einer Generation. Isis, 93(4):715–716, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376021.

Alberti:2002:BRB

[1255] Johanna Alberti Samuel J. M. M. Alberti. Book review: Mary Somerville: Science, Illumination, and the Female Mind. Isis, 93(4):716– 717, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376022.

Kenschaft:2002:BRB

[1256] Patricia Clark Kenschaft. Book review: Women Becoming Mathematicians: Creating a Professional Identity in Post-World War II America. Isis, 93(4):717–718, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376023.

Goldstein:2002:BRB

[1257] Catherine Goldstein. Book review: Lectures on Number Theory. Isis, 93(4):718, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376024.

Hunter:2002:BRB

[1258] Patti W. Hunter. Book review: Essays in the History of Lie Groups and Algebraic Groups. Isis, 93(4):719, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376025.

McKnight:2002:BRB

[1259] John L. McKnight. Book review: Instrumental Traditions and Theories of Light: The Uses of Instruments in the Optical Revolution. Isis, 93 (4):719–720, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376026.

Cassidy:2002:BRB

[1260] David C. Cassidy. Book review: Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. Isis, 93(4):720–721, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376027.

Friedman:2002:BRB

[1261] Robert Marc Friedman. Book review: Weathering the Storm: Sverre Petterssen, the D-Day Forecast, and the Rise of Modern Meteorology. Isis, 93(4):721–722, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376028.

Dean:2002:BRB

[1262] Dennis R. Dean. Book review: Bone Wars: The Excavation and Celebrity of Andrew Carnegie’s Dinosaur. Isis, 93(4):722–723, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376029.

Anderson:2002:BRB

[1263] William D. Anderson, Jr. Book review: The Poetics of Natural History: From John Bartram to William James. Isis, 93(4):723–724, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376030.

Boomgaard:2002:BRB

[1264] Peter Boomgaard. Book review: The Last Tasmanian Tiger: The History and Extinction of the Thylacine. Isis, 93(4):724–725, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376031.

Cech:2002:BRB

[1265] Paul J. Cech. Book review: The Scopes Trial: A Photographic History. Isis, 93(4):725, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376032.

Croce:2002:BRB

[1266] Paul Jerome Croce. Book review: The Transformation of Psychology: Influences of Nineteenth-Century Philosophy, Technology, and Natural Science. Isis, 93(4):725–726, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376033.

Haigh:2002:BRBa

[1267] Elizabeth V. Haigh. Book review: V poiskakh sovetskogo “soiaeuza” filosofii i biologii: Diskussii i repressii v 20-kh-nachale 30-kh gg. Isis, 93 (4):726–727, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376034.

Daum:2002:BRB

[1268] Andreas W. Daum. Book review: Socialist Darwinism: Evolution in German Socialist Thought from Marx to Bernstein. Isis, 93(4):727–738, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376035.

Vucinich:2002:BRB

[1269] Alexander Vucinich. Book review: The Evolutionary Biology Papers of Elie Metchnikoff. Isis, 93(4):728–729, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376036.

Kass:2002:BRB

[1270] L. B. Kass. Book review: The Tangled Field: Barbara McClintock’s Search for the Patterns of Genetic Control. Isis, 93(4):729–730, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376037.

Jackson:2002:BRBa

[1271] John P. Jackson, Jr. Book review: The Color of Race in America, 1900–1940. Isis, 93(4):730–731, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376038.

Viney:2002:BRB

[1272] Wayne Viney. Book review: The Divided Self of William James. Isis, 93 (4):731–732, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376039.

Harkness:2002:BRB

[1273] Jon M. Harkness. Book review: The American Medical Ethics Revolution: How the AMA’s Code of Ethics Has Transformed Physicians’ Relationships to Patients, Professionals, and Society. Isis, 93(4):732–733, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376040.

Moynahan:2002:BRB

[1274] Gregory Moynahan. Book review: Psychotherapy in the Third Reich: The Göring Institute. Isis, 93(4):733, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376041.

Gilman:2002:BRB

[1275] Sander L. Gilman. Book review: The Legend of Nietzsche’s Syphilis. Isis, 93(4):733–734, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376042.

Johnstone:2002:BRB

[1276] Megan-Jane Johnstone. Book review: Nurses in Nazi Germany: Moral Choice in History. Isis, 93(4):734–735, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376043.

Dupont:2002:BRB

[1277] Judith Dupont. Book review: Trauma: A Genealogy. Isis, 93(4):735– 737, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376044.

Darnell:2002:BRB

[1278] Regna Darnell. Book review: “Medicine That Walks”: Medicine, Disease, and Canadian Plains Aboriginal People, 1880–1945. Isis, 93(4): 737–738, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376045.

Hessenbruch:2002:BRB

[1279] Arne Hessenbruch. Book review: Die Pathologie der Person: Friedrich Kraus’ Neubestimmung des Organismus am Beginn des 20, by Martin Lindner. Jahrhunderts. Isis, 93(4):738, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376046.

Hyman:2002:BRB

[1280] Anthony Hyman. Book review: Networking: Communicating with Bodies and Machines in the Nineteenth Century. Isis, 93(4):738–739, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376047.

Levitt:2002:BRB

[1281] Theresa Levitt. Book review: Spectrum of Belief: Joseph von Fraunhofer and the Craft of Precision Optics. Isis, 93(4):739–740, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376048.

Lipartito:2002:BRB

[1282] Kenneth Lipartito. Book review: A Phone of Our Own: The Deaf Insurrection against Ma Bell. Isis, 93(4):740–741, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376049.

Mendelsohn:2002:BRB

[1283] Betsy Mendelsohn. Book review: The Thames Embankment: Environment, Technology, and Society in Victorian London. Isis, 93(4):741, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376050.

Fournier:2002:BRBb

[1284] Marian Fournier. Book review: Unsichtbar-Sichtbar-Durchschaut: Das Mikroskop als Werkzeug des Lebenswissenschaftlers. Isis, 93(4):741–742, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376051.

Mauskopf:2002:BRB

[1285] Seymour Mauskopf. Book review: Literature, Technology, and Magical Thinking, 1880–1920. Isis, 93(4):742–743, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376052.

Ruppert:2002:BRB

[1286] Peter Ruppert. Book review: Science Fiction before 1900: Imagination Discovers Technology. Isis, 93(4):743–744, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376053.

Runte:2002:BRB

[1287] Alfred Runte. Book review: The Anatomy of Nature: Geology and American Landscape Painting, 1825–1875. Isis, 93(4):744–745, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376054.

Weldon:2002:BRB

[1288] Stephen P. Weldon. Book review: Reconciling Science and Religion: The Debate in Early Twentieth–Century Britain. Isis, 93(4):745–747, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376055.

Haigh:2002:BRBb

[1289] John Haigh. Book review: Doomsday: The Science of Catastrophic Events. Isis, 93(4):747–748, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376056.

Holton:2002:BRB

[1290] Gerald Holton. Book review: Blueprints for Reform: Science, Mathematics, and Technology Education American Association for the Advancement of Science. Isis, 93(4):748–749, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376057.

Vogt:2002:BRB

[1291] Annette Vogt. Book review: Mathematikegeschichte ohne Grenzen: Die Korrespondenz zwischen K, by Kurt Vogel; A. P. Iuskevich; M. Folkerts; M. M. Rozanskaja; I. Luther; A. P. Youschkevich; K. Vogel; M. M. Rozanskaja; I. Luther; M. Folkerts. Vogel und A. P. Juschkewitsch A. P. Youshkevich-K. Vogel: Istorija matematiki bez granic. Isis, 93(4):749– 750, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376058.

Russell:2002:BRB

[1292] Colin A. Russell. Book review: The Virgin and the Dynamo: Use and Abuse of Religion in Environmental Debates. Isis, 93(4):750–751, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376059.

Harper:2002:BRB

[1293] Kristine C. Harper. Book review: Eyewitness: Evolution of the Atmospheric Sciences. Isis, 93(4):751–752, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376060.

Mackowski:2002:BRB

[1294] Maura Phillips Mackowski. Book review: Taking Science to the Moon: Lunar Experiments and the Apollo Program. Isis, 93(4):752–753, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376061.

Ericson:2002:BRB

[1295] Steven J. Ericson. Book review: Dropping the Fire: The Decline and Fall of the Steam Locomotive. Isis, 93(4):753–754, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376062.

Marvin:2002:BRB

[1296] Ursula B. Marvin. Book review: Plate Tectonics: An Insider’s History of the Modern Theory of the Earth. Isis, 93(4):754–755, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376063.

Fleming:2002:BRBb

[1297] James Rodger Fleming. Book review: The Ice Chronicles: The Quest to Understand Global Climate Change. Isis, 93(4):755–756, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376064.

Millett:2002:BRB

[1298] David Millett. Book review: The Mechanization of the Mind: On the Origins of Cognitive Science. Isis, 93(4):756–757, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376065.

Caplan:2002:BRB

[1299] Arthur L. Caplan. Book review: Bioethics in America: Origins and Cultural Politics. Isis, 93(4):757–758, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376066.

Charvolin:2002:BRB

[1300] Florian Charvolin. Book review: La nouvelle sociologie des sciences. Isis, 93(4):758–759, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376067.

Dufour:2002:BRB

[1301] Paul Dufour. Book review: Canada’s National System of Innovation. Isis, 93(4):759–760, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376068.

Jackson:2002:BRBb

[1302] Walter A. Jackson. Book review: Social Scientists for Social Justice: Making the Case against Segregation. Isis, 93(4):760–761, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376069.

Muraskin:2002:BRB

[1303] William Muraskin. Book review: Hepatitis B: The Hunt for a Killer Virus. Isis, 93(4):761–762, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376070.

Hogle:2002:BRB

[1304] Linda F. Hogle. Book review: The Story of Taxol: Nature and Politics in the Pursuit of an Anti-Cancer Drug. Isis, 93(4):762–763, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376071.

Hauger:2002:BRB

[1305] J. Scott Hauger. Book review: Inviting Disaster: Lessons from the Edge of Technology. Isis, 93(4):763–764, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376072.

Mohler:2002:BRB

[1306] Stanley R. Mohler. Book review: The Origins and Technology of the Advanced Extravehicular Space Suit. Isis, 93(4):764, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376073.

Crowe:2002:BRB

[1307] Michael J. Crowe. Book review: The Lure of the Edge: Scientific Passions, Religious Beliefs, and the Pursuit of UFOs. Isis, 93(4):764–765, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376074.

Anonymous:2002:BRBaf

[1308] Anonymous. Book review: Metaphysics, Epistemology, and Technology. Isis, 93(4):765–766, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376075.

Anonymous:2002:BRBag

[1309] Anonymous. Book review: Exploring the Unknown: Selected Documents in the History of the U, by John M. Logsdon; Amy Paige Snyder; Roger D. Launius; Stephen J. Garber; Reagan Newport.S. Civil Space Program. Isis, 93(4):766, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376076.

Anonymous:2002:BRBah

[1310] Anonymous. Book review: Documents oubliés sur l’alchimie, la kabbale et Guilaume Postel: À l’occasion de son 90e anniversaire, à François Secret par ses élèves at amis. Isis, 93(4):766, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376077.

Anonymous:2002:BRBai

[1311] Anonymous. Book review: The Structure of Knowledge: Classifications of Science and Learning since the Renaissance. Isis, 93(4):766, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376078.

Anonymous:2002:BRBaj

[1312] Anonymous. Book review: Excluded Ancestors, Inventible Traditions: Essays toward a More Inclusive History of Anthropology. Isis, 93(4):766, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376079.

Anonymous:2002:BRBak

[1313] Anonymous. Book review: Reinventing Hippocrates. Isis, 93(4):767, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376080.

Anonymous:2002:BRBal

[1314] Anonymous. Book review: Russko-nemetskie sviazi v biologii i meditsine [Russian-German Links in Biology and Medicine]. Isis, 93(4):767, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376081.

Anonymous:2002:BRBam

[1315] Anonymous. Book review: Secrets of Nature: Astrology and Alchemy in Early Modern Europe. Isis, 93(4):767, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376082.

Anonymous:2002:BRBan

[1316] Anonymous. Book review: Science Education and Culture: The Contribution of History and Philosophy of Science. Isis, 93(4):767–768, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376083.

Anonymous:2002:BRBao

[1317] Anonymous. Book review: Darwinismus und/als Ideologie. Isis, 93(4): 768, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376084.

Anonymous:2002:BRBap

[1318] Anonymous. Book review: Explanation: Theoretical Approaches and Applications. Isis, 93(4):768, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376085.

Anonymous:2002:BRBaq

[1319] Anonymous. Book review: Der ungebändigte Galilei: Beiträge zu einem Symposion. Isis, 93(4):768, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376086.

Anonymous:2002:BRBar

[1320] Anonymous. Book review: Alexander von Humboldt-Aufbruch in die Moderne. Isis, 93(4):768–769, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376087.

Anonymous:2002:BRBas

[1321] Anonymous. Book review: The Material Culture of Sex, Procreation, and Marriage in Premodern Europe. Isis, 93(4):769, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376088.

Anonymous:2002:BRBat

[1322] Anonymous. Book review: Des savants face a l’occulte, 1870–1940. Isis, 93(4):769, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376089.

Anonymous:2002:BRBau

[1323] Anonymous. Book review: Infanticide: Historical Perspectives on Child Murder and Concealment, 1550–2000. Isis, 93(4):769, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376090.

Anonymous:2002:BRBav

[1324] Anonymous. Book review: Berichte zur Geschichte und Theorie der Ökologie und weitere Beiträge. Isis, 93(4):770, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376091.

Anonymous:2002:BRBaw

[1325] Anonymous. Book review: Narrating the Arctic: A Cultural History of Nordic Scientific Practices. Isis, 93(4):770, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376092.

Anonymous:2002:BRBax

[1326] Anonymous. Book review: English Philosophy in the Age of Locke. Isis, 93(4):770, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376093.

Anonymous:2002:BRBay

[1327] Anonymous. Book review: Geography: Discipline, Profession, and Subject since 1870: An International Survey. Isis, 93(4):770–771, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376094.

Anonymous:2002:BRBaz

[1328] Anonymous. Book review: Secession, History, and the Social Sciences. Isis, 93(4):771, December 2002. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376095.

Anonymous:2002:NCd

[1329] Anonymous. Notes on contributors. Isis, 93(4):772–779, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376096.

Anonymous:2002:IBV

[1330] Anonymous. Index to ISIS, volume 93, 2002. Isis, 93(4):781–799, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/512524.

Anonymous:2002:CBV

[1331] Anonymous. Contents of ISIS, volume 93. Isis, 93(4):803–804, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/512525.

Anonymous:2002:Fd

[1332] Anonymous. Frontispiece. Isis, 93(4):x, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/375012.

Weldon:2003:CBH

[1333] Stephen P. Weldon. Current bibliography of the history of science and its cultural influences 2003. Isis, 94(??):1–193+195–275, ???? 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3652367.

Anonymous:2003:BM

[1334] Anonymous. Back matter. Isis, 94(??):??, ???? 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3652368.

Anonymous:2003:FM

[1335] Anonymous. Front matter. Isis, 94(??):i–xxii, ???? 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3652366.

Elliott:2003:EDH

[1336] Paul Elliott. Erasmus Darwin, Herbert Spencer, and the origins of the evolutionary worldview in British provincial scientific culture, 1770–1850. Isis, 94(1):1–29, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376097.

Westfall:2003:RBS

[1337] Catherine Westfall. Rethinking big science: Modest, mezzo, grand science and the development of the Bevalac, 1971–1993. Isis, 94(1):30–56, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376098.

Stanley:2003:EHW

[1338] Matthew Stanley. “An expedition to heal the wounds of war”: The 1919 eclipse and Eddington as Quaker adventurer. Isis, 94(1):57–89, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376099.

deChadarevian:2003:PDW

[1339] Soraya de Chadarevian. Portrait of a discovery: Watson, crick, and the double helix. Isis, 94(1):90–105, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376100.

Fruton:2003:EJT

[1340] Joseph S. Fruton. Eloge: John Tileston Edsall, 1902–2002. Isis, 94 (1):106–108, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376101.

Sengoopta:2003:ERP

[1341] Chundak Sengoopta. Eloge: Roy Porter, 1946–2002. Isis, 94(1):109– 111, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376102.

Lindgren:2003:EHD

[1342] Uta Lindgren. Eloge: Hertha von Dechend, 1915–2001. Isis, 94(1):112– 113, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376103.

MacLeod:2003:BRB

[1343] Roy MacLeod. Book review: The Chemical Industry in Europe, 1850– 1914: Industrial Growth, Pollution, and Professionalization. Isis, 94 (1):114–116, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376104.

Warner:2003:BRM

[1344] Deborah Jean Warner. Book review: Marc Lachieze-Rey; Jean-Pierre Luminet: Celestial Treasury: From the Music of the Spheres to the Conquest of Space. Isis, 94(1):117, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376105.

Rabkin:2003:BRB

[1345] Yakov M. Rabkin. Book review: Aleph: Historical Studies in Science and Judaism. Isis, 94(1):117–118, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376106.

Yochelson:2003:BRS

[1346] Ellis L. Yochelson. Book review: Stephen J. Pyne: Fire: A Brief History. Isis, 94(1):118, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376107.

DeYoung:2003:BRC

[1347] James DeYoung. Book review: Curtis Peebles: Asteroids: A History. Isis, 94(1):118–119, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376108.

Paddle:2003:BRP

[1348] Robert Paddle. Book review: Peter Boomgaard: Frontiers of Fear: Tigers and People in the Malay World, 1600–1950. Isis, 94(1):119–120, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376109.

Blustein:2003:BRL

[1349] Bonnie Ellen Blustein. Book review: Louise E. Robbins: Louis Pasteur and the Hidden World of Microbes Edward Edelson: Gregor Mendel and the Roots of Genetics; James R. Voelkel: Johannes Kepler and the New Astronomy; John L. Casti; Werner DePauli: Gödel: A Life in Logic. Isis, 94(1):120–121, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376110.

Stearns:2003:BRB

[1350] Peter N. Stearns. Book review: L’Europe des sciences: Constitution d’un éspace scientifique. Isis, 94(1):121–122, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376111.

McOuat:2003:BRD

[1351] Gordon McOuat. Book review: David Hancocks: A Different Nature: The Paradoxical World of Zoos and Their Uncertain Future. Isis, 94 (1):122–123, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376112.

Seidel:2003:BRBa

[1352] Robert W. Seidel. Book review: Chimie et industrie en Europe: L’apport des sociétés savantes industrielles du XIXe siècle à nos jours. Isis, 94 (1):123–124, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376113.

Richmond:2003:BRB

[1353] Marsha L. Richmond. Book review: Cambridge Scientific Minds. Isis, 94(1):124–125, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376114.

Gerovitch:2003:BRB

[1354] Slava Gerovitch. Book review: Vo glave pervenstvuiushchego uchenogo sosloviia Rossii: Ocherki zhizni I deiatel’nosti prezidentov Imperatorskoi Sankt-Peterburgskoi Akademii nauk, 1725–1917. Isis, 94(1):125–126, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376115.

Greenberg:2003:BRJ

[1355] Stephen J. Greenberg. Book review: Jennifer Connor: Guardians of Medical Knowledge: The Genesis of the Medical Library Association. Isis, 94(1):126–127, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376116.

Pyenson:2003:BRD

[1356] Lewis Pyenson. Book review: Denis Guedj: The Measure of the World: A Novel. Isis, 94(1):127–128, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376117.

Ascher:2003:BRB

[1357] Marcia Ascher. Book review: Mathematics across Cultures: The History of Non-Western Mathematics. Isis, 94(1):128–130, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376118.

Bensaude-Vincent:2003:BRT

[1358] Bernadette Bensaude-Vincent. Book review: Trevor H. Levere: Transforming Matter: A History of Chemistry from Alchemy to the Buckyball. Isis, 94(1):130–131, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376119.

Cormack:2003:BRC

[1359] Lesley B. Cormack. Book review: Charles W. J. Withers: Geography, Science, and National Identity: Scotland since 1520. Isis, 94(1):131– 132, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376120.

Reeds:2003:BRJ

[1360] Karen Reeds. Book review: Jane Eliot Sewall: Medicine in Maryland: The Practice and Profession, 1799–1999. Isis, 94(1):132, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376121.

Monforte:2003:BRB

[1361] Guillermo Lusa Monforte. Book review: Historia de la tecnología en España. Isis, 94(1):132–134, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376122.

Berggren:2003:BRR

[1362] J. L. Berggren. Book review: Reviel Netz: The Shaping of Deduction in Greek Mathematics: A Study in Cognitive History; David Fowler: The Mathematics of Plato’s Academy: A New Reconstruction. Isis, 94(1):134–136, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376123.

Goldstein:2003:BRJ

[1363] Bernard R. Goldstein. Book review: John M. Steele: Observations and Predictions of Eclipse Times by Early Astronomers. Isis, 94(1): 136, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376124.

Sabra:2003:BRM

[1364] A. I. Sabra. Book review: A. Mark Smith: Alhacen’s Theory of Visual Perception: A Critical Edition, with English Translation and Commentary, of the First Three Books of Alhacen’s De aspectibus, the Medieval Latin Version of Ibn al-Haytham’s Kitāb al-Manāzir. Isis, 94(1):136– 138, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376125.

Kalin:2003:BRD

[1365] Ibrahim Kalin. Book review: Dimitri Gutas: Greek Thought, Arabic Culture: The Graeco-Arabic Translation Movement in Baghdad and Early Ábbasid Society (Second–Fourth/Eighth–Tenth Centuries). Isis, 94(1):138–140, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376126.

Glasner:2003:BRB

[1366] Ruth Glasner. Book review: Late Medieval and Early Modern Corpuscular Matter Theories. Isis, 94(1):140, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376127.

Robinson:2003:BRJ

[1367] Francis Robinson. Book review: Jonathan M. Bloom: The History and Impact of Paper in the Islamic World. Isis, 94(1):140–141, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376128.

Ziegler:2003:BRR

[1368] Joseph Ziegler. Book review: Roger French: Canonical Medicine: Gentile da Foligno and Scholasticism. Isis, 94(1):141–143, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376129.

Sasaki:2003:BRB

[1369] Chikara Sasaki. Book review: Statecraft and Intellectual Renewal in Late Ming China: The Cross-Cultural Synthesis of Xu Guangqi (1562–1633). Isis, 94(1):143, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376130.

Sengor:2003:BRB

[1370] A. M. C. Sengör. Book review: Osmanli Cografya Literatürü Tarihi (History of Geographical Literature During the Ottoman Period). Isis, 94(1):143–144, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376131.

Bocken:2003:BRH

[1371] Inigo Bocken. Book review: Hubert Benz: Individualität und Subjektivität: Interpretationstendenzen in der Cusanus-Forschung und das Selbstverständnis des Nikolaus von Kues. Isis, 94(1):144–145, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376132.

Sturdy:2003:BRB

[1372] David J. Sturdy. Book review: Les publications de l’académie royale des sciences de Paris (1666–1793). Isis, 94(1):145–146, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376133.

Schneider:2003:BRI

[1373] Ivo Schneider. Book review: Isaac Newton: Die mathematischen Prinzipien der Physik. Isis, 94(1):146–147, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376134.

Principe:2003:BRB

[1374] Lawrence M. Principe. Book review: Thomas and Rebecca Vaughan’s Aqua Vitae: Non Vitis. Isis, 94(1):147–148, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376135.

Sarasohn:2003:BRM

[1375] Lisa T. Sarasohn. Book review: Margaret Cavendish: Observations upon Experimental Philosophy. Isis, 94(1):148, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376136.

Poirier:2003:BRM

[1376] Jean-Pierre Poirier. Book review: Marco Beretta: Imaging A Career in Science: The Iconography of Antoine-Laurent Lavoisier. Isis, 94(1):148– 149, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376137.

Nakayama:2003:BRB

[1377] Shigeru Nakayama. Book review: Dodonaeus in Japan: Translation and the Scientific Mind in the Tokugawa Period. Isis, 94(1):149–150, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376138.

Molland:2003:BRB

[1378] George Molland. Book review: The Occult Laboratory: Magic, Science, and Second Sight in Late Seventeenth-Century Scotland. Isis, 94(1):150– 151, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376139.

Maerker:2003:BRB

[1379] Anna Maerker. Book review: Science and the Visual Image in the Enlightenment. Isis, 94(1):151–152, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376140.

Doel:2003:BRP

[1380] Ronald E. Doel. Book review: Patrick J. McGrath: Scientists, Business, and the State, 1890–1960. Isis, 94(1):152–153, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376141.

Govoni:2003:BRB

[1381] Paola Govoni. Book review: Una difficile modernità: Tradizioni di ricerca e comunità scientifiche in Italia, 1890–1914. Isis, 94(1):153– 154, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376142.

Hentschel:2003:BRG

[1382] Klaus Hentschel. Book review: George Kean Sweetnam: The Command of Light: Rowland’s School of Physics and the Spectrum. (Memoirs of the American Philosophical Society, 238). Isis, 94(1):154–155, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376143.

Markkanen:2003:BRG

[1383] Tapio Markkanen. Book review: Gustav Holmberg: Reaching for the Stars: Studies in the History of Swedish Stellar and Nebular Astronomy, 1860–1940. Isis, 94(1):155–156, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376144.

Uchida:2003:BRM

[1384] Masao Uchida. Book review: Masanori Kaji: Mendeleev’s Discovery of the Periodic Law of the Chemical Elements: The Scientific and Social Context of His Discovery. Isis, 94(1):156–157, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376145.

Udias:2003:BRM

[1385] Agustín Udías. Book review: Marielle Cremer: Seismik zu Beginn des 20. Jahrhunderts: Internationalität und Disziplinbildung. Isis, 94(1):157– 158, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376146.

Stuewer:2003:BRK

[1386] Roger H. Stuewer. Book review: Klaus Hentschel: Zum Zusammenspiel von Instrument, Experiment und Theorie: Rotverschiebung im Sonnenspektrum und verwandte spektrale Verschiebungseffekte von 1880 bis 1960. Isis, 94(1):158–159, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376147.

Pyenson:2003:BRB

[1387] Lewis Pyenson. Book review: Einstein’s Daughter: The Search for Lieserl Einstein: The Formative Years, 1879–1909. Isis, 94(1):159–161, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376148.

Dean:2003:BRH

[1388] Dennis R. Dean. Book review: Hugh Torrens: The Practice of British Geology, 1750–1850. Isis, 94(1):161–162, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376149.

Gallagher:2003:BRD

[1389] Nancy Gallagher. Book review: Donald Malcolm Reid: Whose Pharaohs? Archeology, Museums, and Egyptian National Identity from Napoleon to World War I. Isis, 94(1):162–163, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376150.

Suarez:2003:BRB

[1390] Edna Suárez. Book review: The Reception of Darwinism in the Iberian World. Isis, 94(1):163–164, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376151.

Slack:2003:BRR

[1391] Nancy G. Slack. Book review: Robert A. Croker: Stephen Forbes and the Rise of American Ecology. Isis, 94(1):164–165, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376152.

Fichman:2003:BRP

[1392] Martin Fichman. Book review: Peter Raby: Alfred Russel Wallace: A Life. Isis, 94(1):165–167, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376153.

Gale:2003:BRH

[1393] George Gale. Book review: Harry W. Paul: Bacchic Medicine: Wine and Alcohol Therapies from Napoleon to the French Paradox. Isis, 94(1):167– 168, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376154.

Canizares-Esguerra:2003:BRBa

[1394] Jorge Cañizares-Esguerra. Book review: Marcos Jiménez de la Espada (1831–1898): Tras la senda de un explorador. Isis, 94(1):168, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376155.

Smocovitis:2003:BRJ

[1395] Vassiliki Betty Smocovitis. Book review: James Moore: Good Breeding: Science and Society in a Darwinian Age. Isis, 94(1):168–169, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376156.

Scholthof:2003:BRA

[1396] Karen-Beth G. Scholthof. Book review: Angela N. H. Creager: The Life of a Virus: Tobacco Mosaic Virus as an Experimental Model, 1930– 1965. Isis, 94(1):169–170, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376157.

Darnell:2003:BRB

[1397] Regna Darnell. Book review: O’Odham Creation and Related Events as Told to Ruth Benedict. Isis, 94(1):170–171, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376158.

Strick:2003:BRH

[1398] James Strick. Book review: Henry Harris: Things Come to Life: Spontaneous Generation Revisited. Isis, 94(1):171–172, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376159.

Terrada:2003:BRR

[1399] María-Luz López Terrada. Book review: Ramón Castejón Bolea: Moral sexual y enfermedad: La medicina española frente al peligro venereo (1868–1936). Isis, 94(1):172–173, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376160.

LaBerge:2003:BRD

[1400] Ann La Berge. Book review: Dora B. Weiner: Comprendre et soigner: Philippe Pinel (1745–1826) et la médecine de l’esprit. Isis, 94(1):173– 174, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376161.

Figueiroa:2003:BRH

[1401] Silvia Figueirôa. Book review: Helmut Waszkis: Dr. Moritz (Don Mauricio) Hochschild, 1881–1965: The Man and His Companies: A German Jewish Mining Entrepreneur in South America. Isis, 94(1):174–175, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376162.

Hoppe:2003:BRM

[1402] Brigitte Hoppe. Book review: Mikuláš Teich: Bier, Wissenschaft und Wirtschaft in Deutschland 1800–1914: Ein Beitrag zur deutschen Industrialisierungsgeschichte. Isis, 94(1):175–176, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376163.

Rebert:2003:BRB

[1403] Paula Rebert. Book review: The Indian Territory Journals of Colonel Richard Irving Dodge. Isis, 94(1):176–177, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376164.

Graffi:2003:BRBa

[1404] Giorgio Graffi. Book review: The Prix Volney: Its History and Significance for the Development of Linguistic Research; The Prix Volney: Its History and Significance for the Development of Linguistic Research; Early Nineteenth-Century Contributions to General and Amerindian Linguistics: Du Ponceau and Rafinesque; Contributions to Comparative Indo-European, African, and Chinese Linguistics. Isis, 94(1):177–178, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376165.

Baym:2003:BRR

[1405] Nina Baym. Book review: Robert Bernard Hass: Going by Contraries: Robert Frost’s Conflict with Science. Isis, 94(1):178–179, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376166.

Asquith:2003:BRM

[1406] Pamela J. Asquith. Book review: Morris Low; Shigeru Nakayama; Hitoshi Yoshioka: Science, Technology, and Society in Contemporary Japan. Isis, 94(1):179–181, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376167.

Greenberger:2003:BRH

[1407] Daniel M. Greenberger. Book review: Helge Kragh: Quantum Generations. Isis, 94(1):181–182, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376168.

Reinhardt:2003:BRB

[1408] Carsten Reinhardt. Book review: Robert Burns Woodward: Architect and Artist in the World of Molecules. Isis, 94(1):182–183, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376169.

Vieth:2003:BRI

[1409] Errol Vieth. Book review: Isaac Asimov: It’s Been a Good Life. Isis, 94(1):183–184, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376170.

Samios:2003:BRC

[1410] N. P. Samios. Book review: Charles H. Townes: How the Laser Happened: Adventures of a Scientist. Isis, 94(1):184–185, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376171.

Fitzpatrick:2003:BRY

[1411] Anne Fitzpatrick. Book review: Yakov Alpert: Making Waves: Stories from My Life. Isis, 94(1):185, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376172.

Cumbler:2003:BRK

[1412] John Cumbler. Book review: Kate Foss-Mollan: Hard Water: Politics and Water Supply in Milwaukee, 1870–1995. Isis, 94(1):185–186, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376173.

Sengor:2003:BRN

[1413] A. M. C. Sengör. Book review: Naomi Oreskes: The Rejection of Continental Drift: Theory and Method in American Earth Science. Isis, 94(1):186–188, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376174.

Grandin:2003:BRC

[1414] Karl Grandin. Book review: Cornelia F. Mutel: Flowing through Time: A History of the Iowa Institute of Hydraulic Research. Isis, 94(1):188– 189, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376175.

Becklake:2003:BRD

[1415] John Becklake. Book review: Dennis Piszkiewicz: Wernher von Braun: The Man Who Sold the Moon. Isis, 94(1):189–190, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376176.

VanRiper:2003:BRC

[1416] A. Bowdoin Van Riper. Book review: Carl C. Swisher III; Garniss H. Curtis; Roger Lewin: Java Man: How Two Geologists Changed Our Understanding of Human Evolution. Isis, 94(1):190–191, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376177.

Hughes:2003:BRI

[1417] Sally Smith Hughes. Book review: István Hargittai: Candid Science II: Conversations with Famous Biomedical Scientists. Isis, 94(1):191– 192, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376178.

Magner:2003:BRS

[1418] Lois N. Magner. Book review: Shane Crotty: Ahead of the Curve: David Baltimore’s Life in Science. Isis, 94(1):192–193, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376179.

Fortun:2003:BRS

[1419] Mike Fortun. Book review: Stephen Hall: Invisible Frontiers: The Race to Synthesize a Human Gene. Isis, 94(1):193–194, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376180.

Burlingame:2003:BRS

[1420] Leslie J. Burlingame. Book review: Sterling Evans: The Green Republic: A Conservation History of Costa Rica. Isis, 94(1):194–195, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376181.

Kohler:2003:BRS

[1421] Robert E. Kohler. Book review: Soraya de Chadarevian: Designs for Life: Molecular Biology after World War II. Isis, 94(1):195–196, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376182.

Snedegar:2003:BRD

[1422] Keith Snedegar. Book review: Diana Wylie: Starving on a Full Stomach: Hunger and the Triumph of Cultural Racism in Modern South Africa. Isis, 94(1):196–197, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376183.

Crowther-Heyck:2003:BRD

[1423] Hunter Crowther-Heyck. Book review: Donald T. Critchlow: Intended Consequences: Birth Control, Abortion, and the Federal Government in Modern America. Isis, 94(1):197–198, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376184.

vanKeuren:2003:BRD

[1424] David van Keuren. Book review: David W. Norton: Fifty Years Below Zero: Tributes and Meditations for the Naval Arctic Research Laboratory’s First Half Century at Barrow, Alaska. Isis, 94(1):198–199, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376185.

Mather:2003:BRC

[1425] Patricia Mather. Book review: Carolyn Rasmussen et al.: A Museum for the People: A History of Museum Victoria and Its Predecessors, 1854– 2000. Isis, 94(1):199–201, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376186.

Nader:2003:BRD

[1426] Laura Nader. Book review: Diana E. Forsythe: Studying Those Who Study Us: An Anthropologist in the World of Artificial Intelligence. Isis, 94(1):201–202, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376187.

Pitt:2003:BRL

[1427] Joseph C. Pitt. Book review: Larry A. Hickman: Philosophical Tools for Technological Culture: Putting Pragmatism to Work. Isis, 94(1): 202, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376188.

Leroux:2003:BRS

[1428] Jean Leroux. Book review: Stefania Ruzsits Jha: Reconsidering Michael Polanyi’s Philosophy. Isis, 94(1):202–203, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376189.

Roush:2003:BRP

[1429] Sherrilyn Roush. Book review: Peter Achinstein: The Book of Evidence. Isis, 94(1):203–204, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/376190.

Sanders:2003:BRA

[1430] George Sanders. Book review: Ann Larabee: Decade of Disaster. Isis, 94(1):204–205, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376191.

Ogilvie:2003:BRBa

[1431] Marilyn Ogilvie. Book review: Notable Women Scientists International Women in Science: A Biographical Dictionary to 1950 American Women in Technology: An Encyclopedia. Isis, 94(1):205–207, March 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376192.

Anonymous:2003:BRBa

[1432] Anonymous. Book review: The Return of Science: Evolution, History, and Theory. Isis, 94(1):207, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376193.

Anonymous:2003:BRBb

[1433] Anonymous. Book review: Health Care and Poor Relief in Eighteenth and Nineteenth Century Northern Europe. Isis, 94(1):207, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376194.

Anonymous:2003:BRBc

[1434] Anonymous. Book review: Historical and Philosophical Perspectives on Biomedical Ethics: From Paternalism to Autonomy? Isis, 94(1):207, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376195.

Anonymous:2003:BRBd

[1435] Anonymous. Book review: Patients in the History of Homoeopathy. Isis, 94(1):207–208, March 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
376196.

Anonymous:2003:BRBe

[1436] Anonymous. Book review: Experimenting in Tongues: Studies in Science and Language. Isis, 94(1):208, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376197.

Anonymous:2003:BRBf

[1437] Anonymous. Book review: Inventing Medieval Landscapes: Senses of Place in Western Europe. Isis, 94(1):208, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/376198.

Anonymous:2003:BRBg

[1438] Anonymous. Book review: From Energy to Information: Representation in Science and Technology, Art, and Literature. Isis, 94(1):208, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376199.

Anonymous:2003:BRBh

[1439] Anonymous. Book review: Science and Medicine in the Scottish Enlightenment. Isis, 94(1):208–209, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376200.

Anonymous:2003:BRBi

[1440] Anonymous. Book review: Word, Image, Number: Communication in the Middle Ages. Isis, 94(1):209, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376201.

Anonymous:2003:BRBj

[1441] Anonymous. Book review: From 0 to 1: An Authoritative History of Modern Computing. Isis, 94(1):209, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/376202.

Anonymous:2003:BRBk

[1442] Anonymous. Book review: Time. Isis, 94(1):209, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/376203.

Anonymous:2003:BRBl

[1443] Anonymous. Book review: Ancient and Medieval Traditions in the Exact Sciences: Essays in Memory of Wilbur Knorr. Isis, 94(1):209, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376204.

Anonymous:2003:BRBm

[1444] Anonymous. Book review: Science in Culture. Isis, 94(1):209–210, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376205.

Anonymous:2003:BRBn

[1445] Anonymous. Book review: Who Owns Life? Isis, 94(1):210, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376206.

Anonymous:2003:BRBo

[1446] Anonymous. Book review: Instruments, Travel, and Science: Itineraries of Precision from the Seventeenth to the Twentieth Century. Isis, 94(1): 210, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/376207.

Anonymous:2003:NCa

[1447] Anonymous. Notes on contributors. Isis, 94(1):211–216, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/376208.

Anonymous:2003:F

[1448] Anonymous. Frontispiece. Isis, 94(1):viii, March 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/376209.

Sturkenboom:2003:WSS

[1449] Margaret C. Jacob Dorothée Sturkenboom. A Women’s scientific society in the west: The late Eighteenth-Century assimilation of science. Isis, 94(2):217–252, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379385.

Crowther-Heyck:2003:WSN

[1450] Kathleen Crowther-Heyck. Wonderful secrets of nature: Natural knowledge and religious piety in reformation Germany. Isis, 94(2):253–273, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379386.

Stolberg:2003:WHB

[1451] Michael Stolberg. A woman down to her bones: The anatomy of sexual difference in the Sixteenth and Early Seventeenth Centuries. Isis, 94(2): 274–299, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379387.

Laqueur:2003:SF

[1452] Thomas W. Laqueur. Sex in the flesh. Isis, 94(2):300–306, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379388.

Schiebinger:2003:S

[1453] Londa Schiebinger. Skelettestreit. Isis, 94(2):307–313, June 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379389.

Graham:2003:EAV

[1454] Loren Graham. Eloge: Alexander Vucinich, 1914–2002. Isis, 94(2):314– 315, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379390.

Rhodes:2003:ESJ

[1455] Frank H. T. Rhodes. Eloge: Stephen Jay Gould, 1941–2002. Isis, 94(2): 315–316, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379391.

Robson:2003:E

[1456] Eleanor Robson. To the editor. Isis, 94(2):317–319, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379392.

Cuomo:2003:E

[1457] Serafina Cuomo. To the editor. Isis, 94(2):319–320, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379393.

Imhausen:2003:E

[1458] Annette Imhausen. To the editor. Isis, 94(2):320–321, June 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379394.

Muroi:2003:E

[1459] K. Muroi. To the editor. Isis, 94(2):320–322, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/379395.

McClellan:2003:BRBa

[1460] James E. McClellan, III. Book review: Utopia’s Garden: French Natural History From old Regime to Revolution. Isis, 94(2):324–329, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379396.

Kevles:2003:BRB

[1461] Daniel J. Kevles. Book review: American National Biography. Isis, 94(2): 330–333, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379397.

Pairo:2003:BRB

[1462] Nuria Solsona Pairó. Book review: Ciencia i tecnica als països catalans: Una aproximació biogràfica. Isis, 94(2):334–335, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379398.

Friedman:2003:BRB

[1463] Judith Friedman. Book review: Achievements: A Century of Science for the Benefit of Humankind. Isis, 94(2):335–336, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379399.

Rosner:2003:BRB

[1464] David Rosner. Book review: Mending Bodies, Saving Souls: A History of Hospitals. Isis, 94(2):336–337, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379400.

Goodstein:2003:BRB

[1465] Judith R. Goodstein. Book review: Storia materiale della scienza: Dal libro ai laboratori. Isis, 94(2):337–339, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379401.

Fox:2003:BRBa

[1466] Mary Frank Fox. Book review: The Stars Are Not Enough: Scientists-Their Passions and Professions. Isis, 94(2):339–340, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379402.

Brown:2003:BRB

[1467] Harold I. Brown. Book review: Architectures of Time: Towards a Theory of the Event in Modernist Culture. Isis, 94(2):340, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379403.

Kronick:2003:BRB

[1468] David A. Kronick. Book review: The Book of the Heart. Isis, 94(2):340– 341, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379404.

Grattan-Guinness:2003:BRB

[1469] I. Grattan-Guinness. Book review: Euclid’s Window: The Story of Geometry from Parallel Lines to Hyperspace. Isis, 94(2):341, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379405.

Hahn:2003:BRB

[1470] Robert Hahn. Book review: Praktische Geometrie und Kosmologie am Beispiel der Architektur. Isis, 94(2):341–342, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379406.

Dieks:2003:BRB

[1471] Dennis Dieks. Book review: La physique face à la probabilité. Isis, 94 (2):343, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379407.

Tymoczko:2003:BRB

[1472] Dmitri Tymoczko. Book review: The Cambridge History of Western Music Theory. Isis, 94(2):343–345, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379408.

Kosinski:2003:BRB

[1473] Antoni A. Kosinski. Book review: From Kant to Hilbert: A Source Book in the Foundations of Mathematics. Isis, 94(2):345–347, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379409.

Gingerich:2003:BRB

[1474] Owen Gingerich. Book review: Astrology: A History. Isis, 94(2):347– 348, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379410.

Gove:2003:BRB

[1475] Harry E. Gove. Book review: Measuring Mass: From Positive Rays to Proteins. Isis, 94(2):348–349, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379411.

McKnight:2003:BRB

[1476] John L. McKnight. Book review: Our Cosmic Habitat. Isis, 94(2):349– 350, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379412.

Judd:2003:BRB

[1477] Robin Judd. Book review: Medicine and the German Jews: A History. Isis, 94(2):350–351, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379413.

Scarborough:2003:BRBa

[1478] John Scarborough. Book review: Geschichte der Pharmazie. Isis, 94(2): 351–352, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379414.

Pols:2003:BRB

[1479] Hans Pols. Book review: Mad in America: Bad Science, Bad Medicine, and the Enduring Mistreatment of the Mentally Ill. Isis, 94(2):352–353, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379415.

Toman:2003:BRB

[1480] Cynthia Toman. Book review: Devices and Desires: Gender, Technology, and American Nursing. Isis, 94(2):353–355, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379416.

Pelis:2003:BRB

[1481] Kim Pelis. Book review: Blood: An Epic History of Medicine and Commerce. Isis, 94(2):355–356, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379417.

Garcia:2003:BRB

[1482] Chris Nunn Garcia. Book review: Dividing New Mexico’s Waters, 1700– 1912. Isis, 94(2):356, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379418.

DeRenzi:2003:BRB

[1483] Silvia De Renzi. Book review: Natura-Cultura: L’interpretazione del mondo fisico nei testi e nelle immagini. Isis, 94(2):356–357, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379419.

Andrews:2003:BRB

[1484] Bridie Andrews. Book review: Sex, Law, and Society in Late Imperial China. Isis, 94(2):357–358, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379420.

Steele:2003:BRB

[1485] J. M. Steele. Book review: Astral Sciences in Mesopotamia. Isis, 94(2): 358–359, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379421.

Farquhar:2003:BRB

[1486] Judith Farquhar. Book review: The Expressiveness of the Body and the Divergence of Greek and Chinese Medicine. Isis, 94(2):359–360, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379422.

Evans:2003:BRBa

[1487] Harry B. Evans. Book review: The Water Supply of Ancient Rome: City Area, Water, and Population. Isis, 94(2):360–361, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379423.

Bernard:2003:BRB

[1488] Alain Bernard. Book review: L’aube des mathématiques grecques. Isis, 94(2):361–362, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379424.

Sluiter:2003:BRB

[1489] Ineke Sluiter. Book review: Clavis Commentariorum der antiken medizinischen Texte. Isis, 94(2):362, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379425.

Ogilvie:2003:BRBb

[1490] Brian W. Ogilvie. Book review: Pristina Medicamenta: Ancient and Medieval Medical Botany Herbs and Herbalism in the Middle Ages and Renaissance. Isis, 94(2):362–364, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379426.

Livesey:2003:BRB

[1491] Steven J. Livesey. Book review: God and Reason in the Middle Ages. Isis, 94(2):364, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379427.

Brotons:2003:BRB

[1492] Víctor Navarro Brotóns. Book review: El umbral de la modernidad: Estudios sobre filosofía, religión y ciencia entre Petrarca y Descartes. Isis, 94(2):365–366, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379428.

Poulle:2003:BRB

[1493] Emmanuel Poulle. Book review: Astronomy in the Iberian Peninsula: Abraham Zacut and the Transition from Manuscript to Print. Isis, 94(2): 366–367, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379429.

Burns:2003:BRB

[1494] William E. Burns. Book review: Broader Horizons: A Study of Johannes Witte de Hese’s Itinerarius and Medieval Travel Narratives. Isis, 94(2): 367–368, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379430.

Hall:2003:BRB

[1495] Bert Hall. Book review: The Martial Arts of Renaissance Europe. Isis, 94(2):368–369, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379431.

Cifoletti:2003:BRB

[1496] Giovanna Cifoletti. Book review: Pedro Nunes (1502–1578): His Lost Algebra and Other Discoveries. Isis, 94(2):369–371, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379432.

Charette:2003:BRB

[1497] François Charette. Book review: La Alquibla en al-Andalus y al-Magrib al-Aqs.à. Isis, 94(2):371, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379433.

Mills:2003:BRB

[1498] Allan Mills. Book review: Hartmann’s Practika: A Manual for Making Sundials and Astrolabes with the Compass and Rule. Isis, 94(2):372– 373, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379434.

Reeds:2003:BRB

[1499] Karen Reeds. Book review: The Clock and the Mirror: Girolamo Cardano and Renaissance Medicine. Isis, 94(2):373, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379435.

PresasiPuig:2003:BRB

[1500] Albert Presas i Puig. Book review: Los veintiún libros de los ingenios y máquinas de Juanelo, atribuidos a Pedro Juan de Lastanosa. Isis, 94(2): 373–374, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379436.

Applebaum:2003:BRB

[1501] Wilbur Applebaum. Book review: Groundbreaking Scientific Experiments, Inventions, and Discoveries of the Seventeenth Century. Isis, 94(2):374–375, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379437.

Clericuzio:2003:BRB

[1502] Antonio Clericuzio. Book review: The Philosophy of Robert Boyle. Isis, 94(2):375–376, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379438.

Rabin:2003:BRB

[1503] Sheila J. Rabin. Book review: Kepler’s Philosophy and the New Astronomy. Isis, 94(2):376–377, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379439.

Blackwell:2003:BRB

[1504] Richard J. Blackwell. Book review: Morandi’s Last Prophecy and the End of Renaissance Politics. Isis, 94(2):377, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/379440.

Orland:2003:BRB

[1505] Barbara Orland. Book review: Zur Geschichte der Naturzerstörung: Frau und Tier in der wissenschaftlichen Revolution. Isis, 94(2):377– 378, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379441.

Cody:2003:BRB

[1506] Lisa Forman Cody. Book review: The Midwives of Seventeenth-Century London. Isis, 94(2):378–379, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379442.

Efron:2003:BRB

[1507] Noah J. Efron. Book review: Judaism in the Theology of Sir Isaac Newton. Isis, 94(2):379–380, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379443.

Fell:2003:BRB

[1508] Ulrike Fell. Book review: Lavoisier: Die Revolution in der Chemie. Isis, 94(2):380–381, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379444.

Franceschetti:2003:BRB

[1509] Donald R. Franceschetti. Book review: The Art of Teaching Physics: The Eighteenth-Century Demonstration Apparatus of Jean Antoine Nollet. Isis, 94(2):381, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379445.

Muller-Wille:2003:BRB

[1510] Staffan Müller-Wille. Book review: Linnaeus: Nature and Nation. Isis, 94(2):382–383, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379446.

Rothfels:2003:BRB

[1511] Nigel Rothfels. Book review: Elephant Slaves and Pampered Parrots: Exotic Animals in Eighteenth-Century Paris. Isis, 94(2):383–384, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379447.

Nieto-Galan:2003:BRB

[1512] Augustí Nieto-Galan. Book review: Papermaking in Eighteenth-Century France: Management, Labor, and Revolution at the Montgolfier Mill, 1761–1805. Isis, 94(2):384–385, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379448.

Menke:2003:BRB

[1513] Richard Menke. Book review: American Women of Letters and the Nineteenth-Century Sciences: Styles of Affiliation. Isis, 94(2):385, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379449.

Schubring:2003:BRB

[1514] Gert Schubring. Book review: Extension Theory. Isis, 94(2):386–387, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379450.

VanEvra:2003:BRB

[1515] James W. Van Evra. Book review: The Search for Mathematical Roots, 1870–1940: Logics, Set Theories, and the Foundations of Mathematics from Cantor through Russell to Gödel. Isis, 94(2):387–388, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379451.

Swenson:2003:BRB

[1516] Loyd S. Swenson, Jr. Book review: Ferdinand Braun, 1850–1918: Ein wilhelminischer Physiker. Isis, 94(2):388–389, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/379452.

Schroder:2003:BRB

[1517] Wilfried Schröder. Book review: L’espistemologia di Planck: Nel suo contesto stori. Isis, 94(2):389, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379453.

Seifer:2003:BRB

[1518] Marc J. Seifer. Book review: Tesla: Master of Lightning. Isis, 94(2): 389–390, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379454.

Cassidy:2003:BRB

[1519] David C. Cassidy. Book review: Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). Isis, 94(2):390–391, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379455.

Kaiser:2003:BRB

[1520] David Kaiser. Book review: Mapping the Spectrum: Techniques of Visual Representation in Research and Teaching. Isis, 94(2):391–392, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379456.

Metailie:2003:BRB

[1521] Georges Métailié. Book review: Translating Science: The Transmission of Western Chemistry into Late Imperial China, 1840–1900. Isis, 94(2): 392–393, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379457.

Davies:2003:BRB

[1522] Gordon H. Herries Davies. Book review: Gideon Mantell and the Discovery of Dinosaurs. Isis, 94(2):393–394, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379458.

Browne:2003:BRB

[1523] Janet Browne. Book review: Sparks of Life: Darwinism and the Victorian Debates over Spontaneous Generation. Isis, 94(2):394–396, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379459.

Mehigan:2003:BRB

[1524] Tim Mehigan. Book review: Fundamentals of the Theory of Movement Perception. Isis, 94(2):396–397, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379460.

Ruse:2003:BRB

[1525] Michael Ruse. Book review: The Structure of Evolutionary Theory. Isis, 94(2):397–398, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379461.

Yans-McLaughlin:2003:BRB

[1526] Virginia Yans-McLaughlin. Book review: Margaret Mead, Gregory Bateson, and Highland Bali: Fieldwork Photographs of Bayung Gedé, 1936– 1939. Isis, 94(2):398, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379462.

Golden:2003:BRB

[1527] Janet Golden. Book review: Bill W. and Mr. Wilson: The Legend and Life of A.A.’s Cofounder, by Matthew J. Raphael. Isis, 94(2):398–399, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379463.

Edelson:2003:BRBa

[1528] Paul J. Edelson. Book review: Malaria: Poverty, Race, and Public Health in the United States. Isis, 94(2):399–400, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379464.

Hamlin:2003:BRB

[1529] Christopher Hamlin. Book review: Cheated Not Poisoned? Food Regulation in the United Kingdom, 1875–1938. Isis, 94(2):400–401, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379465.

Simon:2003:BRB

[1530] Jonathan Simon. Book review: Embryos in Wax: Models from the Ziegler Studio. Isis, 94(2):401–402, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379466.

Peitzman:2003:BRB

[1531] Steven J. Peitzman. Book review: Making the Modern Medical School: The Wisconsin Stories. Isis, 94(2):402–404, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/379467.

Finlay:2003:BRB

[1532] Mark Finlay. Book review: Poems in Steel: National Socialism and the Politics of Inventing from Weimar to Bonn. Isis, 94(2):404–405, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379468.

Jepsen:2003:BRB

[1533] Thomas Jepsen. Book review: Telegraph Messenger Boys: Labor, Technology, and Geography, 1850–1950. Isis, 94(2):405–406, June 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379469.

Hanson:2003:BRB

[1534] Elizabeth Hanson. Book review: Savages and Beasts: The Birth of the Modern Zoo. Isis, 94(2):406–407, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379470.

Nickles:2003:BRB

[1535] Thomas Nickles. Book review: At the End of an Age. Isis, 94(2):407– 408, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379471.

Kojevnikov:2003:BRB

[1536] Alexei Kojevnikov. Book review: Andrei Sakharov: Nauka i Svoboda Sakharov: A Biography. Isis, 94(2):408–409, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/379472.

Saller:2003:BRB

[1537] Heinrich Saller. Book review: Conceptual Foundations of Quantum Field Theory. Isis, 94(2):409–410, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379473.

Fastovsky:2003:BRB

[1538] David E. Fastovsky. Book review: Time Traveler: In Search of Dinosaurs and Ancient Mammals from Montana to Mongolia. Isis, 94(2): 410–411, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379474.

Rainger:2003:BRBa

[1539] Ronald Rainger. Book review: The Eternal Darkness: A Personal History of Deep-Sea Exploration. Isis, 94(2):411–412, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379475.

Perkins:2003:BRB

[1540] John H. Perkins. Book review: Travels in the Genetically Modified Zone. Isis, 94(2):412–413, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379476.

Sloan:2003:BRB

[1541] Phillip R. Sloan. Book review: Evolution of Sameness and Difference: Perspectives on the Human Genome Project. Isis, 94(2):413–414, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379477.

Steen:2003:BRB

[1542] Tomoko Y. Steen. Book review: Perspectives on Genetics: Anecdotal, Historical, and Critical Commentaries, 1987–1998. Isis, 94(2):414–415, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379478.

Nunn:2003:BRB

[1543] Chris Nunn. Book review: The Creation of Psychopharmacology. Isis, 94(2):415–416, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
379479.

Lambeth:2003:BRB

[1544] Benjamin S. Lambeth. Book review: MiG Aircraft since 1937. Isis, 94(2): 416–417, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379480.

Lankford:2003:BRB

[1545] John Lankford. Book review: NASA and the Space Industry. Isis, 94(2): 417–418, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379481.

Light:2003:BRB

[1546] Jennifer S. Light. Book review: Proceedings of the 1998 Conference on the History and Heritage of Science Information Systems. Isis, 94(2): 418, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/379482.

Mitcham:2003:BRB

[1547] Carl Mitcham. Book review: Technological Innovation as an Evolutionary Process. Isis, 94(2):418–419, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379483.

Wang:2003:BRE

[1548] Zuoyue Wang. Book review: Edward Teller and Judith L. Shoolery: Memoirs: A Twentieth-Century Journey in Science and Politics. Isis, 94(2):419–420, June 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.
1086/379484;   http://www.jstor.org/stable/pdfplus/10.1086/
379484.pdf.

McNeill:2003:BRB

[1549] J. R. McNeill. Book review: A History of Environmental Politics since 1945. Isis, 94(2):420–421, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379485.

Pyne:2003:BRB

[1550] Stephen J. Pyne. Book review: In the Absence of Predators: Conservation and Controversy on the Kaibab Plateau. Isis, 94(2):421–422, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/379486.

Gatensby:2003:BRB

[1551] Anne Gatensby. Book review: Social Empiricism. Isis, 94(2):422–423, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/379487.

Gingras:2003:BRB

[1552] Yves Gingras. Book review: Introduction à la sociologie des sciences et des connaissances scientifiques. Isis, 94(2):423–424, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379488.

Barcaro:2003:BRB

[1553] Umberto Barcaro. Book review: L’università italiana: Bibliografia 1848– 1914. Isis, 94(2):424–425, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/379489.

Greene:2003:BRB

[1554] Mott Greene. Book review: Records and History of the United States Geological Survey. Isis, 94(2):425–426, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/379490.

Anonymous:2003:NCb

[1555] Anonymous. Notes on contributors. Isis, 94(2):427–433, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/379491.

Wigelsworth:2003:CPN

[1556] Jeffrey R. Wigelsworth. Competing to popularize Newtonian philosophy: John Theophilus Desaguliers and the preservation of reputation. Isis, 94 (3):435–455, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380653.

Levitt:2003:BPA

[1557] Theresa Levitt. Biot’s paper and Arago’s plates: Photographic practice and the transparency of representation. Isis, 94(3):456–476, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380654.

Chimisso:2003:MHO

[1558] Cristina Chimisso and Gad Freudenthal. A mind of her own: Hélène Metzger to Émile Meyerson, 1933. Isis, 94(3):477–491, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380655.

Kargon:2003:EHW

[1559] Robert Kargon. Eloge: Harry Woolf, 1923–2003. Isis, 94(3):492–493, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380656.

Fischer:2003:ELK

[1560] Michael Fischer, Sarah Jansen, and Charles Weiner. Eloge: Lily E. Kay, 1947–2000. Isis, 94(3):493–495, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380657.

MacLeod:2003:EPF

[1561] Roy MacLeod. Eloge: Philip F. Rehbock, 1942–2001. Isis, 94(3):495–498, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380658.

MacKenzie:2003:EDO

[1562] Donald MacKenzie. Eloge: David Owen Edge, 1932–2003. Isis, 94(3): 498–499, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380659.

Rawlins:2003:E

[1563] Dennis Rawlins. To the editor. Isis, 94(3):500–502, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380660.

Hallyn:2003:BRF

[1564] Fernand Hallyn. Book review: Frederika Oosterhoff: Ideas Have a History: Perspectives on the Western Search for Truth. Isis, 94(3):503, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380661.

Runte:2003:BRR

[1565] Alfred Runte. Book review: Raymond F. Dasmann: Called by the Wind: The Autobiography of a Conservationist. Isis, 94(3):503–504, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380662.

Rothenberg:2003:BRT

[1566] Marc Rothenberg. Book review: Timothy W. Luke: Museum Politics: Power Plays at the Exhibition. Isis, 94(3):504–505, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380663.

Whitworth:2003:BRS

[1567] Michael H. Whitworth. Book review: Steven Meyer: Irresistible Dictation: Gertrude Stein and the Correlations of Writing and Science. Isis, 94(3):505–506, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380664.

Freedman:2003:BRS

[1568] Carl Freedman. Book review: Susan E. Lederer: Frankenstein: Penetrating the Secrets of Nature. Isis, 94(3):506–507, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380665.

Mirowski:2003:BRR

[1569] Philip Mirowski. Book review: E. Roy Weintraub: How Economics Became a Mathematical Science. Isis, 94(3):507–508, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380666.

Rainger:2003:BRBb

[1570] Ronald Rainger. Book review: Oceanographic History: The Pacific and Beyond. Isis, 94(3):508–509, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380667.

Newcomb:2003:BRB

[1571] Sally Newcomb. Book review: Salt: Grain of Life. Isis, 94(3):509–510, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380668.

Rudge:2003:BRB

[1572] David Wÿss Rudge. Book review: Biology and Epistemology. Isis, 94 (3):510–511, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380669.

Pemberton:2003:BRB

[1573] Stephen Pemberton. Book review: A History of Disability. Isis, 94 (3):511, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380670.

vanderGeest:2003:BRF

[1574] Sjaak van der Geest. Book review: Fritz Krafft: Christus als Apotheker: Ursprung, Aussage und Geschichte eines christlichen Sinnbildes. Isis, 94 (3):511–513, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380671.

Mazumdar:2003:BRJ

[1575] Pauline M. H. Mazumdar. Book review: J. T. H. Connor: Doing Good: The Life of Toronto’s General Hospital. Isis, 94(3):513–514, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380672.

Adams:2003:BRB

[1576] Tracey L. Adams. Book review: A Sourcebook of Dental Medicine. Isis, 94(3):514–515, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380673.

Bennett:2003:BRR

[1577] James T. Bennett. Book review: S. Robert Lichter; Stanley Rothman: Environmental Cancer: A Political Disease? Isis, 94(3):515–516, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380674.

Connor:2003:BRW

[1578] Jennifer J. Connor. Book review: William H. Helfand: Quack, Quack, Quack: The Sellers of Nostrums in Prints, Posters, Ephemera, and Books. Isis, 94(3):516–517, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380675.

Grunden:2003:BRB

[1579] Walter E. Grunden. Book review: Culture and Technology in Modern Japan. Isis, 94(3):517–518, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380676.

Riddle:2003:BRM

[1580] John M. Riddle. Book review: Mark Grant: Galen on Food and Diet. Isis, 94(3):518, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380677.

Howard:2003:BRB

[1581] Robert A. Howard. Book review: The Fireworks Book: Gunpowder in Medieval Germany. Isis, 94(3):518–519, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380678.

Pironet:2003:BRB

[1582] Fabienne Pironet. Book review: On the Purity of the Art of Logic: The Shorter and the Longer Treatises. Isis, 94(3):519–520, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380679.

Freudenthal:2003:BRR

[1583] Gad Freudenthal. Book review: Ruth Glasner: A Fourteenth-Century Scientific Philosophic Controversy: Jedaiah ha-Penini’s Treatise on Opposite Motions and Book of Confutation: Critical Edition, Introduction, and Commentary. Isis, 94(3):520–521, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380680.

Hirtler:2003:BRB

[1584] Eva Hirtler. Book review: Sinnlichkeit und Vernunft in der mittelalterlichen Musiktheorie: Strategien der Konsonanzwertung und der Gegenstand der. Isis, 94(3):521–522, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380681.

Chabas:2003:BRB

[1585] José Chabás. Book review: Reportorio de los tiempos. Isis, 94(3):522, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380682.

MacLeod:2003:BRW

[1586] Christine MacLeod. Book review: W. Patrick McCray: Glassmaking in Renaissance Venice: The Fragile Craft. Isis, 94(3):522–523, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380683.

Finocchiaro:2003:BRB

[1587] Maurice A. Finocchiaro. Book review: Galileo in Context. Isis, 94(3): 523–524, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380684.

Jenkin:2003:BRJ

[1588] John Jenkin. Book review: John Gascoigne: The Enlightenment and the Origins of European Australia. Isis, 94(3):524–525, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380685.

Neal:2003:BRA

[1589] Katherine Neal. Book review: Amir R. Alexander: Geometrical Landscapes: The Voyages of Discovery and the Transformation of Mathematical Practice. Isis, 94(3):525–526, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380686.

Hedrick:2003:BRE

[1590] Elizabeth Hedrick. Book review: Elizabeth Potter: Gender and Boyle’s Law of Gases. Isis, 94(3):526–527, September 2003. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/380687.

Kemp:2003:BRL

[1591] Martin Kemp. Book review: Lawrence M. Principe; Lloyd DeWitt: Transmutations: Alchemy in Art: Selected Works from the Eddleman and Fisher Collections at the Chemical Heritage Foundation. Isis, 94 (3):527–528, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380688.

Eddy:2003:BRB

[1592] M. D. Eddy. Book review: Brian Nance: Turquet de Mayerne as Baroque Physician: The Art of Medical Portraiture. Isis, 94(3):528–529, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380689.

Dhombres:2003:BRI

[1593] Jean Dhombres. Book review: Isabelle Stengers: La guerre de sciences aura-t-elle lieu? Isis, 94(3):529, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380690.

Salmon:2003:BRR

[1594] Marylynn Salmon. Book review: Rebecca J. Tannenbaum: The Healer’s Calling: Women and Medicine in Early New England. Isis, 94(3): 529–530, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380691.

Raj:2003:BRD

[1595] Kapil Raj. Book review: David Arnold: Science, Technology, and Medicine in Colonial India. Isis, 94(3):530–532, September 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380692.

Williams:2003:BRO

[1596] L. Pearce Williams. Book review: Olivier Darrigol: Electrodynamics from Ampère to Einstein. Isis, 94(3):532, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380693.

Wachelder:2003:BRP

[1597] Joseph Wachelder. Book review: Paul F. H. Lauxtermann: Schopenhauer’s Broken World-View: Colours and Ethics between Kant and Goethe. Isis, 94(3):532–533, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380694.

Hentschel:2003:BRM

[1598] Klaus Hentschel. Book review: Mike Ware: Cyanotype: The History, Science, and Art of Photographic Printing in Prussian Blue. Isis, 94 (3):533–534, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380695.

Pinna:2003:BRB

[1599] Giovanni Pinna. Book review: Giovanni Canestrini: Zoologist and Darwinist. Isis, 94(3):534–535, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380696.

Callery:2003:BRB

[1600] Bernadette G. Callery. Book review: American Garden Literature in the Dumbarton Oaks Collection (1785–1900): From the Newengland Farmer to Italian Gardens: An Annotated Bibliography. Isis, 94(3):535, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380697.

Turner:2003:BRR

[1601] Frank M. Turner. Book review: Randal Keynes: Darwin, His Daughter, and Human Evolution. Isis, 94(3):535–536, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380698.

Estow:2003:BRF

[1602] Clara Estow. Book review: Felipe Jerez Moliner: Los artistas valencianos de la Ilustración y el grabado biológico y médico (1759–1814). Isis, 94 (3):536–538, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380699.

Trombold:2003:BRI

[1603] Charles D. Trombold. Book review: Ian Graham: Alfred Maudslay and the Maya: A Biography. Isis, 94(3):538–539, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380700.

Matless:2003:BRV

[1604] David Matless. Book review: Volker M. Welter: Biopolis: Patrick Geddes and the City of Life. Isis, 94(3):539–540, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380701.

Rothstein:2003:BRJ

[1605] William G. Rothstein. Book review: John S. Haller, Jr.: A Profile in Alternative Medicine: The Eclectic Medical College of Cincinnati, 1845– 1942. Isis, 94(3):540–541, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380702.

Weiss:2003:BRB

[1606] Sheila Faith Weiss and Thomas M. Berez. Book review: Medicine and Medical Ethics in Nazi Germany: Origins, Practices, Legacies. Isis, 94 (3):541–543, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380703.

Chinard:2003:BRJ

[1607] Francis P. Chinard. Book review: Jean Adrian: Louis C. Maillard: De la médecine à l’alimentation. Isis, 94(3):543–544, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380704.

Paton:2003:BRC

[1608] Miranda Paton. Book review: Charles Musser: Edison Motion Pictures, 1890–1900: An Annotated Filmography. Isis, 94(3):544–545, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380705.

Platt:2003:BRR

[1609] Harold L. Platt. Book review: Robert W. Jackson: Rails across the Mississippi: A History of the St. Louis Bridge. Isis, 94(3):545–546, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380706.

Seidel:2003:BRM

[1610] Robert W. Seidel. Book review: Michel Cotte: Le fonds d’archives Seguin: Aux origines de la révolution industrielle en France 1790–1860.

Isis, 94(3):546–547, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380707.

Oslund:2003:BRD

[1611] Karen Oslund. Book review: David Thomas Murphy: German Exploration of the Polar World: A History, 1870–1940. Isis, 94(3):547–549, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380708.

Elliott:2003:BRBa

[1612] Clark A. Elliott. Book review: The Papers of Joseph Henry. Isis, 94 (3):549, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380709.

Doctorow:2003:BRB

[1613] Osher Doctorow. Book review: The Theory of Probability. Isis, 94(3): 549–550, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380710.

Kidwell:2003:BRB

[1614] Peggy Aldrich Kidwell. Book review: Maria Mitchell: A Life in Journals and Letters. Isis, 94(3):550–551, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380711.

Kline:2003:BRE

[1615] Ronald Kline. Book review: Emily Thompson: The Soundscape of Modernity: Architectural Acoustics and the Culture of Listening in America, 1900–1933. Isis, 94(3):551–552, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/380712.

Chalk:2003:BRG

[1616] Rosemary Chalk. Book review: Gordon Moran: Silencing Scientists and Scholars in Other Fields: Power, Paradigm Controls, Peer Review, and Scholarly Communication. Isis, 94(3):552–553, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380713.

Doob:2003:BRB

[1617] J. L. Doob. Book review: Kolmogorov in Perspective. Isis, 94(3):553–554, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380714.

Gough:2003:BRG

[1618] Jerry B. Gough. Book review: Glenn T. Seaborg; Eric Seaborg: Adventures in the Atomic Age: From Watts to Washington. Isis, 94(3):554–555, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380715.

Pessoa:2003:BRO

[1619] Osvaldo Pessoa, Jr. Book review: Olival Freire, Jr.: David Bohm e a controvérsia dos quanta. Isis, 94(3):555, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380716.

Haigh:2003:BRD

[1620] Elizabeth Haigh. Book review: Douglas R. Weiner: A Little Corner of Freedom: Russian Nature Protection from Stalin to Gorbachëv. Isis, 94 (3):555–556, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380717.

Pickford:2003:BRJ

[1621] Martin Pickford. Book review: Jon Kalb: Adventures in the Bone Trade: The Race to Discover Human Ancestors in Ethiopia’s Afar Depression. Isis, 94(3):556–557, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380718.

Weidman:2003:BRD

[1622] Nadine Weidman. Book review: Deborah Blum: Love at Goon Park: Harry Harlow and the Science of Affection. Isis, 94(3):557–558, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380719.

Strasser:2003:BRF

[1623] Bruno J. Strasser. Book review: Frederic Lawrence Holmes: Meselson, Stahl, and the Replication of DNA: A History of “The Most Beautiful Experiment in Biology”. Isis, 94(3):558–560, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380720.

Hamblin:2003:BRH

[1624] Jacob Darwin Hamblin. Book review: Helen Rozwadowski: The Sea Knows No Boundaries: A Century of Marine Science under ICES. Isis, 94(3):560–561, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380721.

Jeffrey:2003:BRW

[1625] Kirk Jeffrey. Book review: Wilson Greatbatch: The Making of the Pacemaker: Celebrating a Lifesaving Invention. Isis, 94(3):561, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/380722.

Cantor:2003:BRJ

[1626] David Cantor. Book review: John O’Donnell: Coriell: The Coriell Institute for Medical Research and a Half Century of Science. Isis, 94 (3):561–562, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380723.

Eyler:2003:BRP

[1627] John M. Eyler. Book review: Peter C. English: Rheumatic Fever in America and Britain: A Biological, Epidemiological, and Medical History. Isis, 94(3):562–563, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380724.

Yost:2003:BRM

[1628] Jeffrey Yost. Book review: M. Mitchell Waldrop: The Dream Machine: J. C. R. Licklider and the Revolution That Made Computing Personal. Isis, 94(3):563–564, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380725.

Griffin:2003:BRP

[1629] Nicholas Griffin. Book review: Peter H. Denton: The ABC of Armageddon: Bertrand Russell on Science, Religion, and the Next War, 1919– 1938. Isis, 94(3):564–565, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380726.

Riordan:2003:BRA

[1630] Michael Riordan. Book review: Allan Franklin: Selectivity and Discord: Two Problems of Experiment. Isis, 94(3):565–566, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380727.

Secord:2003:BRB

[1631] James A. Secord. Book review: Thornton and Tully’s Scientific Books, Libraries, and Collectors: A Study of Bibliography and the Book Trade in Relation to the History of Science. Isis, 94(3):566–567, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380728.

Anonymous:2003:BRBp

[1632] Anonymous. Book review: The Arts of Seventeenth-Century Science: Representations of the Natural World in European and North American Culture. Isis, 94(3):567, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/380729.

Anonymous:2003:BRBq

[1633] Anonymous. Book review: The New Science and Jesuit Science: Seventeenth-Century Perspectives, by Mordechai Feingold. (Archimedes: New Studies in the History and Philosophy of Science and Technology, 6.). Isis, 94(3):567, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380730.

Anonymous:2003:BRBr

[1634] Anonymous. Book review: Ad Quadratum: The Practical Application of Geometry in Medieval Architecture. Isis, 94(3):567, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380731.

Anonymous:2003:BRBs

[1635] Anonymous. Book review: Thomas Kuhn. Isis, 94(3):567–568, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380732.

Anonymous:2003:BRBt

[1636] Anonymous. Book review: “The Common Purposes of Life”: Science and Society at the Royal Institution of Great Britain. Isis, 94(3):568, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380733.

Anonymous:2003:BRBu

[1637] Anonymous. Book review: The Changing Image of the Sciences. Isis, 94(3):568, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380734.

Anonymous:2003:BRBv

[1638] Anonymous. Book review: Galen and Galenism. Isis, 94(3):568, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/380735.

Anonymous:2003:BRBw

[1639] Anonymous. Book review: Über Schall: Ernst Machs und Peter Salchers Geschossfotografien. Isis, 94(3):568–569, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/380736.

Anonymous:2003:BRBaa

[1640] Anonymous. Book review: Jesuit Science and the Republic of Letters. Isis, 94(3):569, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/380740.

Anonymous:2003:BRBx

[1641] Anonymous. Book review: The Path to Justice: Following in the Footsteps of Henry George. Isis, 94(3):569, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/380737.

Anonymous:2003:BRBy

[1642] Anonymous. Book review: The Road to Medical Statistics. Isis, 94 (3):569, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380738.

Anonymous:2003:BRBz

[1643] Anonymous. Book review: Dahlemer Archivgespräche. Isis, 94(3):569, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/380739.

Anonymous:2003:BRBab

[1644] Anonymous. Book review: Jahrbuch für Universitätsgeschichte. Isis, 94 (3):569–570, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380741.

Anonymous:2003:BRBac

[1645] Anonymous. Book review: History of Rocketry and Astronautics. Isis, 94(3):570, September 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
380742.

Anonymous:2003:NCc

[1646] Anonymous. Notes on contributors. Isis, 94(3):571–575, September 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/380743.

Guerrini:2003:DS

[1647] Anita Guerrini. Duverney’s skeletons. Isis, 94(4):577–603, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386383.

DeBont:2003:CPM

[1648] Raf De Bont. The creation of Prehistoric Man: Aimé Rutot and the Eolith Controversy, 1900–1920. Isis, 94(4):604–630, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386384.

Kirsh:2003:PGI

[1649] Nurit Kirsh. Population genetics in Israel in the 1950s: The unconscious internalization of ideology. Isis, 94(4):631–655, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386385.

Batten:2003:WED

[1650] Alan H. Batten. What Eddington did not say. Isis, 94(4):656–659, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/386386.

Hiebert:2003:RB

[1651] Erwin N. Hiebert. Response to batten. Isis, 94(4):660, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386387.

Rocke:2003:EFL

[1652] Alan Rocke and John Harley Warner. Eloge: Frederic Lawrence Holmes, 1932–2003. Isis, 94(4):661–665, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386388.

Lindee:2003:EDN

[1653] M. Susan Lindee. Eloge: Dorothy Nelkin, 1933–2003. Isis, 94(4):666– 669, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386389.

Malet:2003:ELG

[1654] Antoni Malet. Eloge: Luis García Ballester, 1936–2000. Isis, 94(4):669– 671, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386390.

Murdoch:2003:EGM

[1655] John E. Murdoch. Eloge: A. George Molland, 1941–2002. Isis, 94(4):671– 674, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386391.

Bouse:2003:E

[1656] Derek Bousé. To the editor. Isis, 94(4):675–676, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386392.

Kroll:2003:R

[1657] Gary Kroll. In reply. Isis, 94(4):676–677, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386393.

Hentschel:2003:E

[1658] Klaus Hentschel. To the editor. Isis, 94(4):677–678, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386394.

Stanley:2003:R

[1659] Matthew Stanley. In reply. Isis, 94(4):678, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386395.

Caneva:2003:BRB

[1660] Kenneth Caneva. Book review: The Romantic Conception of Life: Science and Philosophy in the Age of Goethe. Isis, 94(4):679–683, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386396.

Turner:2003:BRB

[1661] Frank M. Turner. Book review: Science in Theistic Contexts: Cognitive Dimensions. Isis, 94(4):684, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386397.

Lee:2003:BRB

[1662] Paula Lee. Book review: Making Space for Science: Territorial Themes in the Shaping of Knowledge. Isis, 94(4):684–685, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386398.

Tobey:2003:BRB

[1663] Ronald Tobey. Book review: History by Numbers: An Introduction to Quantitative Approaches. Isis, 94(4):685–686, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386399.

Brush:2003:BRB

[1664] Stephen G. Brush. Book review: The Cambridge History of Science. Isis, 94(4):687–688, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386400.

Benson:2003:BRB

[1665] Keith R. Benson. Book review: Stuffed Animals and Pickled Heads: The Culture and Evolution of Natural History Museums. Isis, 94(4):688–689, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386401.

Cravens:2003:BRBa

[1666] Hamilton Cravens. Book review: An Elusive Science: The Troubling History of Education Research. Isis, 94(4):689–690, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386402.

Park:2003:BRBa

[1667] Katharine Park. Book review: Spectacular Bodies: The Art and Science of the Human Body from Leonardo da Vinci to Now. Isis, 94(4):690, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386403.

Elliott:2003:BRBb

[1668] Clark A. Elliott. Book review: The Web of Knowledge: A Festschrift in Honor of Eugene Garfield. Isis, 94(4):690–691, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386404.

Hudson:2003:BRB

[1669] Robert G. Hudson. Book review: The Emergence of Sexuality: Historical Epistemology and the Formation of Concepts. Isis, 94(4):691–692, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386405.

Graffi:2003:BRBb

[1670] Giorgio Graffi. Book review: Limiting the Arbitrary: Linguistic Naturalism and Its Opposites in Plato’s Cratylus and Modern Theories of Language. Isis, 94(4):692–693, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386406.

Martinez:2003:BRB

[1671] Alberto A. Martínez. Book review: Knots: Mathematics with a Twist. Isis, 94(4):693–694, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386407.

Sengers:2003:BRB

[1672] Johanna Levelt Sengers. Book review: Cohesion: A Scientific History of Intermolecular Forces. Isis, 94(4):694–695, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386408.

Swan:2003:BRB

[1673] Patricia B. Swan. Book review: The Backbone of History: Health and Nutrition in the Western Hemisphere. Isis, 94(4):695–696, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386409.

Erlen:2003:BRB

[1674] Jonathon Erlen. Book review: Rising Life Expectancy: A Global History. Isis, 94(4):696, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386410.

Kenen:2003:BRB

[1675] Stephanie H. Kenen. Book review: An American Obsession: Science, Medicine, and Homosexuality in Modern Society. Isis, 94(4):696–697, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386411.

Rasmussen:2003:BRB

[1676] Margaret Rasmussen. Book review: Ceramic Innovations in the Twentieth Century. Isis, 94(4):697–698, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386413.

Johnson:2003:BRB

[1677] Stephen B. Johnson. Book review: Between Human and Machine: Feedback, Control, and Computing before Cybernetics. Isis, 94(4):698–699, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386412.

Valone:2003:BRB

[1678] David A. Valone. Book review: The Kingdom of Science: Literary Utopianism and British Education, 1612–1870. Isis, 94(4):700, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386414.

Early:2003:BRB

[1679] Julie Early. Book review: The Aurelian Legacy: British Butterflies and Their Collectors. Isis, 94(4):700–701, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386415.

Noble:2003:BRB

[1680] David W. Noble. Book review: Inventing Ourselves Out of Jobs? America’s Debate over Technological Unemployment, 1929–1981. Isis, 94(4): 701–702, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386416.

Shinn:2003:BRB

[1681] Terry Shinn. Book review: Instrument-Experiment: Historische Studien. Isis, 94(4):702–703, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386417.

Smith:2003:BRB

[1682] Merritt Roe Smith. Book review: The Gifts of Athena: Historical Origins of the Knowledge Economy. Isis, 94(4):703–705, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386418.

Evans:2003:BRBb

[1683] James Evans. Book review: Ptolemy’s Geography: An Annotated Translation of the Theoretical Chapters. Isis, 94(4):705, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386419.

Tybjerg:2003:BRB

[1684] Karin Tybjerg. Book review: Prolegomena Mathematica: From Apollonius of Perga to Late Neoplatonism. Isis, 94(4):705–706, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386420.

Bezza:2003:BRB

[1685] Giuseppe Bezza. Book review: Sesto Empirico: Contro gli astrologi. Isis, 94(4):706–707, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386421.

Scarborough:2003:BRBb

[1686] John Scarborough. Book review: Des causes et des signes des maladies aiguës et chroniques. Isis, 94(4):707–708, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386422.

Wickkiser:2003:BRB

[1687] Bronwen Wickkiser. Book review: The Art Is Long: On the Sacred Disease and the Scientific Tradition. Isis, 94(4):708–709, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386423.

Carpenter:2003:BRB

[1688] Kenneth J. Carpenter. Book review: On the Properties of Foodstuffs (De alimentorum facultatibus). Isis, 94(4):709, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386424.

Brentjes:2003:BRB

[1689] Sonja Brentjes. Book review: Suhayl: Journal for the History of the Exact and Natural Sciences in Islamic Civilisation. Isis, 94(4):709–710, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386425.

Barmore:2003:BRB

[1690] Frank E. Barmore. Book review: The Hereford Map: A Transcription and Translation of the Legends with Commentary. Isis, 94(4):710–711, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386426.

OBoyle:2003:BRB

[1691] Cornelius O’Boyle. Book review: Les “images astrologiques” au Moyen Âge et à la Renaissance: Spéculations intellectuelles et pratiques magiques (XIIe-XVe siècle). Isis, 94(4):711–712, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386427.

Jones:2003:BRBa

[1692] Matthew L. Jones. Book review: A Culture of Fact: England, 1550–1720. Isis, 94(4):712–713, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386428.

Rutkin:2003:BRB

[1693] H. Darrel Rutkin. Book review: An Annotated Census of Copernicus’. Isis, 94(4):713–714, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386429.

Savoia:2003:BRB

[1694] Andrea Ubrizsy Savoia. Book review: Osservazione di cose straordinarie: Il. Isis, 94(4):714–715, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386430.

Lines:2003:BRB

[1695] David A. Lines. Book review: The Universities of the Italian Renaissance. Isis, 94(4):715–716, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386431.

Perez:2003:BRB

[1696] Jorge Navarro Pérez. Book review: Tesoros de la biblioteca histórica Doctor Nicolás León: Libros de medicina de los siglos XVI, XVII y XVIII. Isis, 94(4):716–717, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386432.

Jacquart:2003:BRB

[1697] Danielle Jacquart. Book review: Thirteenth- and Fourteenth-Century Copies of the Ars medicine: A Checklist and Contents Description of the Manuscripts The Articella in the Early Press, c, by Cornelius O’Boyle; Jon Arrizabalaga; Roger French. 1476–1534 Papers of the Articella Project Meeting: Cambridge, December 1995. Isis, 94(4):717–718, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386433.

Martens:2003:BRB

[1698] Rhonda Martens. Book review: Luminaries in the Natural World: The Sun and the Moon in England, 1400–1720. Isis, 94(4):718–719, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386434.

Gabbey:2003:BRB

[1699] Alan Gabbey. Book review: The Principia: Mathematical Principles of Natural Philosophy. Isis, 94(4):719–721, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386435.

Marchitello:2003:BRB

[1700] Howard Marchitello. Book review: Charting an Empire: Geography at the English Universities, 1580–1620. Isis, 94(4):721–722, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386436.

Boschiero:2003:BRB

[1701] Luciano Boschiero. Book review: Isaac Beeckman: Atomista e lettore critico di Lucrezio. Isis, 94(4):722–723, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386437.

Ariew:2003:BRB

[1702] Roger Ariew. Book review: La recherche de la vérité par la lumière naturelle de René Descartes. Isis, 94(4):723, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386438.

Meli:2003:BRB

[1703] Domenico Bertoloni Meli. Book review: The Eye of the Lynx: Galileo, His Friends, and the Beginnings of Modern Natural History. Isis, 94 (4):723–725, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386439.

Bird:2003:BRB

[1704] Alexander Bird. Book review: The Mechanisation of Aristotelianism: The Late Aristotelian Setting of Thomas Hobbes’ Natural Philosophy. Isis, 94(4):725–726, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386440.

McMullin:2003:BRB

[1705] Ernan McMullin. Book review: Largo campo di filosofare. Isis, 94(4):726– 727, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386441.

Rosenberg:2003:BRB

[1706] Gary D. Rosenberg. Book review: Fossil Woods and Other Geological Specimens. Isis, 94(4):727–728, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386442.

Fara:2003:BRB

[1707] Patricia Fara. Book review: An Age of Wonders: Prodigies, Politics, and Providence in England, 1657–1727. Isis, 94(4):728, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386443.

Snobelen:2003:BRB

[1708] Stephen D. Snobelen. Book review: The Cambridge Companion to Newton. Isis, 94(4):728–729, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386444.

LeCoat:2003:BRB

[1709] Nanette Le Coat. Book review: The Wild Girl, Natural Man, and the Monster: Dangerous Experiments in the Age of Enlightenment. Isis, 94 (4):729–731, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386445.

Fox:2003:BRBb

[1710] Robert Fox. Book review: L’État, l’armée, la science: L’invention de la recherche publique en France, 1763–1830. Isis, 94(4):731–732, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386446.

Mandelbrote:2003:BRB

[1711] Scott Mandelbrote. Book review: The Cure of the Passions and the Origins of the English Novel. Isis, 94(4):732–733, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386447.

Shea:2003:BRB

[1712] William R. Shea. Book review: Volta: Science and Culture in the Age of Enlightenment. Isis, 94(4):733, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386448.

Canizares-Esguerra:2003:BRBb

[1713] Jorge Cañizares-Esguerra. Book review: Los naturalistas españoles en el África Hispana (1860–1936). Isis, 94(4):733–734, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386449.

Mathis:2003:BRB

[1714] Wolfgang Mathis. Book review: Wissenschaftlicher Briefwechsel, by Arnold Sommerfeld; Michael Eckert; Karl Märker. Volume 1: 1892–1918. Isis, 94(4):734–735, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386450.

Tjossem:2003:BRB

[1715] Sara Tjossem. Book review: A Dame Full of Vim and Vigor: A Biography of Alice Middleton Boring, Biologist in China. Isis, 94(4):735–736, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386451.

Bocking:2003:BRB

[1716] Stephen Bocking. Book review: Fishing the Great Lakes: An Environmental History, 1783–1933. Isis, 94(4):736–737, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386452.

Kleinman:2003:BRB

[1717] Kim Kleinman. Book review: Journal of a Georgia Woman, 1870–1872. Isis, 94(4):737, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386453.

Haule:2003:BRB

[1718] John Ryan Haule. Book review: Un voyant prodigieux: Alexis Didier, 1826–1886. Isis, 94(4):737–738, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386454.

Brooke:2003:BRB

[1719] John Hedley Brooke. Book review: Darwin’s Cathedral: Evolution, Religion, and the Nature of Society. Isis, 94(4):738–739, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386455.

Montgomery:2003:BRB

[1720] William Montgomery. Book review: Charles Darwin, by Janet Browne. Volume 2: The Power of Place. Isis, 94(4):739–740, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386456.

Pandora:2003:BRB

[1721] Katherine Pandora. Book review: Hugo De Vries: Travels of a Dutch Botanist in America, 1904–1912. Isis, 94(4):740–741, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386457.

Jones:2003:BRBb

[1722] Greta Jones. Book review: Science, Reform, and Politics in Victorian Britain: The Social Science Association, 1857–1886. Isis, 94(4):741, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386458.

Neeley:2003:BRB

[1723] Kathryn A. Neeley. Book review: Dying to Know: Scientific Epistemology and Narrative in Victorian England. Isis, 94(4):741–742, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386459.

Rossi:2003:BRB

[1724] William Rossi. Book review: Henry David Thoreau and the Moral Agency of Knowing. Isis, 94(4):742–743, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386460.

Brice:2003:BRB

[1725] William R. Brice. Book review: Agassiz’s Legacy: Scientists’ Reflections on the Value of Field Experience. Isis, 94(4):743–744, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386461.

Larson:2003:BRB

[1726] Edward J. Larson. Book review: The Scopes Trial: A Brief History with Documents. Isis, 94(4):744–745, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386462.

Cravens:2003:BRBb

[1727] Hamilton Cravens. Book review: Max Weber and the Idea of Economic Sociology. Isis, 94(4):745–746, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386463.

Golan:2003:BRB

[1728] Tal Golan. Book review: Bodies of Evidence: Medicine and the Politics of the English Inquest, 1830–1926. Isis, 94(4):746–747, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386464.

Sangwan:2003:BRB

[1729] Satpal Sangwan. Book review: Romanticism and Colonial Disease. Isis, 94(4):747–748, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386465.

Green:2003:BRB

[1730] Christopher D. Green. Book review: Insights of Genius: Imagery and Creativity in Science and Art. Isis, 94(4):748–749, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386466.

Marchand:2003:BRB

[1731] Suzanne Marchand. Book review: Max Ernst and Alchemy: A Magician in Search of Myth. Isis, 94(4):749–750, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386467.

Hinsley:2003:BRB

[1732] Curtis M. Hinsley. Book review: The Limits of Multiculturalism: Interrogating the Origins of American Anthropology. Isis, 94(4):750–751, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386468.

Slaton:2003:BRB

[1733] Amy Slaton. Book review: The Works: The Industrial Architecture of the United States. Isis, 94(4):751–752, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/386469.

Killen:2003:BRB

[1734] Andreas Killen. Book review: Träume nach Freud: Die “Traumdeutung” und die Geschichte der psychoanalytischen Bewegung. Isis, 94(4):752– 753, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386470.

Weisz:2003:BRB

[1735] George Weisz. Book review: Instruire, guérir, servir: Formation, recherche et pratique médicales en France et en Allemagne pendant la deuxième moitié du XIXe siècle. Isis, 94(4):753–754, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386471.

Birn:2003:BRB

[1736] Anne-Emanuelle Birn. Book review: Virus, moustiques et modernité: La fièvre jaune au Brésil entre science et politique. Isis, 94(4):754–755, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386472.

Cueto:2003:BRB

[1737] Marcos Cueto. Book review: Rockefeller Philanthropy and Modern Biomedicine: International Initiatives from World War I to the Cold War. Isis, 94(4):755–756, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386473.

Dawson:2003:BRB

[1738] John W. Dawson, Jr. Book review: Rockefeller and the Internationalization of Mathematics between the Two World Wars: Documents and Studies for the Social History of Mathematics in the Twentieth Century. Isis, 94(4):756–757, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386474.

McFarland-Icke:2003:BRB

[1739] Bronwyn McFarland-Icke. Book review: Medical Films, Ethics, and Euthanasia in Nazi Germany: The History of Medical Research and Teaching Films of the Reich Office for Educational Films/Reich Institute for Films in Science and Education, 1933–1945. Isis, 94(4):757–758, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/386475.

Weisel:2003:BRB

[1740] Gary J. Weisel. Book review: Einstein: The Passions of a Scientist. Isis, 94(4):758, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386476.

Haigh:2003:BRB

[1741] Elizabeth Haigh. Book review: The Earth’s Biosphere: Evolution, Dynamics, and Change. Isis, 94(4):758–759, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/386477.

Goss:2003:BRB

[1742] David A. Goss. Book review: History of Light and Colour Measurements: Science in the Shadows. Isis, 94(4):759–760, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386478.

Hahner-Rombach:2003:BRB

[1743] Sylvelyn Hähner-Rombach. Book review: “Captain of All These Men of Death”: The History of Tuberculosis in Nineteenth and Twentieth Century Ireland. Isis, 94(4):760–761, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386479.

Carlson:2003:BRB

[1744] Elof Axel Carlson. Book review: Building a Better Race: Gender, Sexuality, and Eugenics from the Turn of the Century to the Baby Boom. Isis, 94(4):761, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386480.

Park:2003:BRBb

[1745] Buhm Soon Park. Book review: Noble Conspirator: Florence S, by Judith Robinson. Mahoney and the Rise of the National Institutes of Health. Isis, 94(4):761–763, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386481.

Dando:2003:BRB

[1746] Malcolm Dando. Book review: Britain and Biological Warfare: Expert Advice and Science Policy, 1930–65. Isis, 94(4):763, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386482.

Creager:2003:BRB

[1747] Angela N. H. Creager. Book review: Edwin J, by Douglas M. Surgenor. Cohn and the Development of Protein Chemistry: With a Detailed Account of His Work on the Fractionation of Blood During and After World War II. Isis, 94(4):763–765, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386483.

Sapp:2003:BRB

[1748] Jan Sapp. Book review: Linus Pauling: Scientist and Peacemaker. Isis, 94(4):765–766, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386484.

Wang:2003:BRBc

[1749] Zuoyue Wang. Book review: Sputnik: The Shock of the Century. Isis, 94 (4):766–767, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386485.

Wolfe:2003:BRB

[1750] Audra J. Wolfe. Book review: Scientists in the Classroom: The Cold War Reconstruction of American Science Education. Isis, 94(4):767–768, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386486.

Grabiner:2003:BRB

[1751] Judith V. Grabiner. Book review: California Dreaming: Reforming Mathematics Education. Isis, 94(4):768, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/386487.

Bassett:2003:BRB

[1752] Ross Bassett. Book review: MIT and the Rise of Entrepreneurial Science. Isis, 94(4):768–769, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386488.

Backhouse:2003:BRB

[1753] Roger E. Backhouse. Book review: Machine Dreams: Economics Becomes a Cyborg Science. Isis, 94(4):769–771, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386489.

Keller:2003:BRB

[1754] Evelyn Fox Keller. Book review: The Misunderstood Gene. Isis, 94(4): 771, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386490.

Michel:2003:BRB

[1755] Sonya Michel. Book review: Poverty Knowledge: Social Science, Social Policy, and the Poor in Twentieth–Century U, by Alice O’Connor.S. History. Isis, 94(4):771–772, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386491.

Edelson:2003:BRBb

[1756] Paul J. Edelson. Book review: The Cure: A Story of Cancer and Politics from the Annals of the Cold War. Isis, 94(4):772–773, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386492.

Murphy:2003:BRB

[1757] Michelle Murphy. Book review: Into Our Own Hands: The Women’s Health Movement in the United States, 1969–1990. Isis, 94(4):773–774, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386493.

Markowitz:2003:BRB

[1758] Gerald Markowitz. Book review: Old Paint: A Medical History of Childhood Lead-Paint Poisoning in the United States to 1980. Isis, 94(4):774– 775, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386494.

Kline:2003:BRB

[1759] Ronald Kline. Book review: Boys and Their Toys? Masculinity, Class, and Technology in America Making Technology Masculine: Men, Women, and Modern Machines in America, 1870–1945. Isis, 94(4):775– 776, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386495.

Clark:2003:BRB

[1760] Mark H. Clark. Book review: Hacker Culture. Isis, 94(4):776–777, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/386496.

Coopersmith:2003:BRB

[1761] Jonathan Coopersmith. Book review: Inventing Modern America: From the Microwave to the Mouse. Isis, 94(4):777–778, December 2003. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386497.

Beyler:2003:BRB

[1762] Richard H. Beyler. Book review: Analog Days: The Invention and Impact of the Moog Synthesizer. Isis, 94(4):778–779, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386498.

Rozwadowski:2003:BRB

[1763] Helen M. Rozwadowski. Book review: Sea Legs: Tales of a Woman Oceanographer. Isis, 94(4):779, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386499.

Meyer:2003:BRB

[1764] William B. Meyer. Book review: Scanning the Skies: A History of Tornado Forecasting. Isis, 94(4):779–780, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/386500.

Olby:2003:BRB

[1765] Robert Olby. Book review: A Life in Science. Isis, 94(4):780–781, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/386501.

Mackowski:2003:BRB

[1766] Maura Phillips Mackowski. Book review: Uplink-Downlink: A History of the Deep Space Network, 1957–1997. Isis, 94(4):781–782, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386502.

Tatarewicz:2003:BRB

[1767] Joseph N. Tatarewicz. Book review: “Before This Decade Is Out…, by Glen E. Swanson… ”: Personal Reflections on the Apollo Program. Isis, 94(4):782–783, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/386503.

Ghamari-Tabrizi:2003:BRB

[1768] Sharon Ghamari-Tabrizi. Book review: Nuclear Reactions: The Politics of Opening a Radioactive Waste Disposal Site. Isis, 94(4):783–784, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386504.

Seidel:2003:BRBb

[1769] Robert W. Seidel. Book review: From Newspeak to Cyberspeak: A History of Soviet Cybernetics. Isis, 94(4):784–785, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386505.

Fedunkiw:2003:BRB

[1770] Marianne Fedunkiw. Book review: Federal Bodysnatchers and the New Guinea Virus: Parasites, People, and Politics. Isis, 94(4):785, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386506.

Toon:2003:BRB

[1771] Elizabeth Toon. Book review: Critical Care Nursing: A History. Isis, 94 (4):785–786, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386507.

Ceruzzi:2003:BRB

[1772] Paul E. Ceruzzi. Book review: Retooling: A Historian Confronts Technological Change. Isis, 94(4):786–787, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/386508.

Marti:2003:BRB

[1773] Oscar R. Martí. Book review: Peirce’s Scientific Metaphysics: The Philosophy of Chance, Law, and Evolution. Isis, 94(4):787–788, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/386509.

McClellan:2003:BRBb

[1774] James McClellan, III. Book review: The Social Origins of Modern Science. Isis, 94(4):788–789, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386510.

Jasanoff:2003:BRB

[1775] Sheila Jasanoff. Book review: The Social Construction of What? Isis, 94 (4):789–790, December 2003. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
386511.

Moore:2003:BRB

[1776] Kelly Moore. Book review: Ivory Bridges: Connecting Science and Society. Isis, 94(4):790–791, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386512.

Kornblith:2003:BRB

[1777] Hilary Kornblith. Book review: How Scientific Practices Matter: Reclaiming Philosophical Naturalism. Isis, 94(4):791–792, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386513.

Anonymous:2003:BRBad

[1778] Anonymous. Book review: The Politics of the Healthy Life: An International Perspective. Isis, 94(4):792, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/386514.

Anonymous:2003:BRBae

[1779] Anonymous. Book review: Worldly Provincialism: German Anthropology in the Age of Empire. Isis, 94(4):792, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386515.

Anonymous:2003:BRBaf

[1780] Anonymous. Book review: New Media, 1740–1915. Isis, 94(4):792, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/386516.

Anonymous:2003:BRBag

[1781] Anonymous. Book review: Electric Bodies: Episodes in the History of Medical Electricity. Isis, 94(4):792–793, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386517.

Anonymous:2003:BRBah

[1782] Anonymous. Book review: Paracelsian Moments: Science, Medicine, and Astrology in Early Modern Europe. Isis, 94(4):793, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386518.

Anonymous:2003:BRBai

[1783] Anonymous. Book review: From China to Paris: 2000 Years’ Transmission of Mathematical Ideas. Isis, 94(4):793, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386519.

Anonymous:2003:BRBaj

[1784] Anonymous. Book review: On the Edge: Soviet Science in the First Half of the Twentieth Century. Isis, 94(4):793, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/386520.

Anonymous:2003:BRBak

[1785] Anonymous. Book review: Bodily Extremities: Preoccupations with the Human Body in Early Modern European Culture. Isis, 94(4):793–794, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386521.

Anonymous:2003:BRBal

[1786] Anonymous. Book review: The Enterprise of Science in Islam: New Perspectives. Isis, 94(4):794, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386522.

Anonymous:2003:BRBam

[1787] Anonymous. Book review: History of Mathematical Sciences: Portugal and East Asia II: Scientific Practices and the Portuguese Expansion in Asia, 1498–1759. Isis, 94(4):794, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/386523.

Anonymous:2003:NCd

[1788] Anonymous. Notes on contributors. Isis, 94(4):795–802, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/386524.

Anonymous:2003:IBV

[1789] Anonymous. Index to ISIS, volume 94, 2003. Isis, 94(4):803–820, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/512527.

Anonymous:2003:CV

[1790] Anonymous. Contents, volume 94. Isis, 94(4):821–822, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/512526.

Rossiter:2003:FIB

[1791] Margaret W. Rossiter. Forty issues on; or, Isis midwifery. Isis, 94(4):viii– ix, December 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/386382.

Weldon:2004:CBH

[1792] Stephen P. Weldon. Current bibliography of the history of science and its cultural influences 2004. Isis, 95(??):1–205+207–249, ???? 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3652370.

Anonymous:2004:BM

[1793] Anonymous. Back matter. Isis, 95(??):??, ???? 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3652371.

Anonymous:2004:FM

[1794] Anonymous. Front matter. Isis, 95(??):i–xxiii, ???? 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/3652369.

Zuidervaart:2004:BOL

[1795] Huib J. Zuidervaart and Rob H. van Gent. “A bare outpost of learned European culture on the edge of the jungles of Java”: Johan Maurits Mohr (1716–1775) and the emergence of instrumental and institutional science in Dutch Colonial Indonesia. Isis, 95(1):1–33, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423509.

Tuchman:2004:SGM

[1796] Arleen Marcia Tuchman. Situating gender: Marie E. Zakrzewska and the place of science in Women’s medical education. Isis, 95(1):34–57, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423510.

Jenkin:2004:WHB

[1797] John Jenkin. William Henry Bragg in adelaide: Beginning research at a colonial locality. Isis, 95(1):58–90, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423511.

Gieryn:2004:ERK

[1798] Thomas F. Gieryn. Eloge: Robert K. Merton, 1910–2003. Isis, 95(1): 91–94, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423512.

Dauben:2004:EKR

[1799] Joseph W. Dauben and Christoph J. Scriba. Eloge: Kurt-Reinhard Biermann, 1919–2002. Isis, 95(1):94–95, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423513.

Lurie:2004:ETN

[1800] Edward Lurie. Eloge: Thomas Neville Bonner, 1923–2003. Isis, 95(1): 96, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423514.

Cadden:2004:E

[1801] Joan Cadden, Rebecca Flemming, Monica H. Green, and Helen King. To the editor. Isis, 95(1):97–98, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423515.

Rossiter:2004:R

[1802] Margaret W. Rossiter. In reply. Isis, 95(1):98, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423516.

Chabas:2004:E

[1803] José Chabás and Bernard R. Goldstein. To the editor. Isis, 95(1):98– 100, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423517.

Poulle:2004:R

[1804] Emmanuel Poulle. In reply. Isis, 95(1):100, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423518.

Walter:2004:BRP

[1805] Maila Walter. Book review: Peter Atkins: Galileo’s Finger: The Ten Great Ideas of Science. Isis, 95(1):101, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423519.

Lynn:2004:BRN

[1806] Michael R. Lynn. Book review: Nicolas Witkowski: Une histoire sentimentale des sciences. Isis, 95(1):101–102, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423520.

Bolt:2004:BRH

[1807] Marvin Bolt. Book review: Hester Higton: Sundials at Greenwich: A Catalogue of the Sundials, Nocturnals, and Horary Quadrants in the National Maritime Museum, Greenwich. Isis, 95(1):102–103, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423521.

Cohen:2004:BRC

[1808] Seymour S. Cohen. Book review: Cathy Cobb: Magick, Mayhem, and Mavericks: The Spirited History of Physical Chemistry. Isis, 95(1):103– 105, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423522.

Farber:2004:BRM

[1809] Paul Lawrence Farber. Book review: Michael A. Walters: A Concise History of Ornithology. Isis, 95(1):105, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423523.

OConnor:2004:BRM

[1810] Erin O’Connor. Book review: Mary Roach: Stiff: The Curious Lives of Human Cadavers. Isis, 95(1):105–106, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/423524.

Shapiro:2004:BRR

[1811] Alan E. Shapiro. Book review: Raymond L. Lee, Jr.; Alistair B. Fraser: The Rainbow Bridge: Rainbows in Art, Myth, and Science. Isis, 95 (1):106–107, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423525.

James:2004:BRR

[1812] Kathyrn James. Book review: Roger French: Medicine before Science: The Business of Medicine from the Middle Ages to the Enlightenment. Isis, 95(1):107–108, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423526.

Leung:2004:BRB

[1813] Angela Ki Che Leung. Book review: Innovation in Chinese Medicine. Isis, 95(1):108–109, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423527.

Grob:2004:BRB

[1814] Gerald N. Grob. Book review: The Cambridge Historical Dictionary of Disease. Isis, 95(1):109–110, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423528.

Broman:2004:BRB

[1815] Thomas Broman. Book review: Geschichte der Medizinischen Fakultät Greifswald: Geschichte der Medizinischen Fakultät von 1456 bis 1713 von Christoph Helwig D, by Hans Georg Thümmel.J. und das Dekanatsbuch der Medizinischen Fakultät von 1714 bis 1823. Isis, 95(1):110–111, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423529.

Touwaide:2004:BRB

[1816] Alain Touwaide. Book review: On My Own Opinions. Isis, 95(1):111, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423530.

Livesey:2004:BRJ

[1817] Steven J. Livesey. Book review: Jack Zupko: John Buridan: Portrait of a Fourteenth-Century Arts Master. Isis, 95(1):111–112, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423531.

Ochs:2004:BRB

[1818] Sidney Ochs. Book review: Maimonides and the Sciences. Isis, 95 (1):112–113, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423532.

Gosselin:2004:BRB

[1819] Edward A. Gosselin. Book review: The Cabala of Pegasus. Isis, 95 (1):113–114, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423533.

Esguerra:2004:BRB

[1820] Jorge Cañizares-Esguerra. Book review: El Códice de Jaume Honorat Pomar (c. 1550–1606), by Jaume Honorat Pomar; José María López Piñero; Rita Barberá Nolla; María José Alcón Miquel: Plantas y animales del Viejo Mundo y de América. Isis, 95(1):114–115, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423534.

Hall:2004:BRR

[1821] Bert Hall. Book review: Roberta J. Magnusson: Water Technology in the Middle Ages: Cities, Monasteries, and Waterworks after the Roman Empire. Isis, 95(1):115–116, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423535.

Hedrick:2004:BRB

[1822] Elizabeth Hedrick. Book review: Encyclopedia of the Scientific Revolution: From Copernicus to Newton. Isis, 95(1):116, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423536.

Palmerino:2004:BRI

[1823] Carla Rita Palmerino. Book review: Ingrid D. Rowland: The Ecstatic Journey: Athanasius Kircher in Baroque Rome. Isis, 95(1):116–117, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423537.

Yost:2004:BRR

[1824] Robinson M. Yost. Book review: A. R. T. Jonkers: Earth’s Magnetism in the Age of Sail. Isis, 95(1):117–118, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423538.

Tarver:2004:BRW

[1825] John Tarver. Book review: Walter Kingsley Taylor; Eliane M. Norman: André Michaux in Florida: An Eighteenth-Century Botanical Journey. Isis, 95(1):118–119, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423539.

Larson:2004:BRA

[1826] Jim Larson. Book review: Andreas Mettenleiter: Adam Christian Thebesius (1686–1732) und die Entdeckung der Vasa Cordis Minima: Biographie, Medizinhistorische Würdigung und Rezeptionsgeschichte. Isis, 95 (1):119–120, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423540.

Duchesneau:2004:BRB

[1827] François Duchesneau. Book review: Repertorium zu Albrecht von Hallers Korrespondenz, 1724–1777. Isis, 95(1):120–121, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423541.

Burns:2004:BRJ

[1828] William E. Burns. Book review: Jan Bondeson: The Two-Headed Boy and Other Medical Marvels. Isis, 95(1):121, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423542.

Wilson:2004:BRAa

[1829] Philip K. Wilson. Book review: Anne Borsay: Medicine and Charity in Georgian Bath: A Social History of the General Infirmary, c. 1739–1830. Isis, 95(1):122, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423543.

Jones:2004:BRE

[1830] David Jones. Book review: Elizabeth A. Fenn: Pox Americana: The Great Smallpox Epidemic of 1775–82. Isis, 95(1):122–123, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423544.

Lawrence:2004:BRK

[1831] Christopher Lawrence. Book review: Kay K. Moss: Southern Folk Medicine, 1750–1820. Isis, 95(1):123–124, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/423545.

Ariew:2004:BRK

[1832] Roger Ariew. Book review: Kathleen Wellman: Making Science Social: The Conferences of Théophraste Renaudot, 1633–1642. Isis, 95(1):124– 125, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423546.

Cumo:2004:BRP

[1833] Christopher Cumo. Book review: Patrick Brantlinger: Dark Vanishings: Discourse on the Extinction of Primitive Races, 1800–1930. Isis, 95 (1):125–126, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423547.

Madison:2004:BRM

[1834] Mark Madison. Book review: Michael A. Bryson: Visions of the Land: Science, Literature, and the American Environment from the Era of Exploration to the Age of Ecology. Isis, 95(1):126–127, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423548.

Kohlstedt:2004:BRK

[1835] Sally Gregory Kohlstedt. Book review: Kim Tolley: The Science Education of American Girls: A Historical Perspective. Isis, 95(1):127–128, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423549.

Weldon:2004:BRJ

[1836] Stephen P. Weldon. Book review: James Turner: Language, Religion, Knowledge: Past and Present. Isis, 95(1):128, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423550.

Haule:2004:BRA

[1837] John Ryan Haule. Book review: Alison Winter: Mesmerized: Powers of Mind in Victorian Britain. Isis, 95(1):128–129, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423551.

Koblitz:2004:BRR

[1838] Ann Hibner Koblitz. Book review: Richard Pipes: The Degaev Affair: Terror and Treason in Tsarist Russia. Isis, 95(1):129–130, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423552.

Wilson:2004:BRAb

[1839] David B. Wilson. Book review: Andrew Warwick: Masters of Theory: Cambridge and the Rise of Mathematical Physics. Isis, 95(1):130–131, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423553.

Badash:2004:BRJ

[1840] Lawrence Badash. Book review: J. L. Heilbron: Ernest Rutherford and the Explosion of Atoms. Isis, 95(1):131–132, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423554.

Levitt:2004:BRT

[1841] Martin L. Levitt. Book review: Tom Tucker: Bolt of Fate: Benjamin Franklin and His Electric Kite Hoax. Isis, 95(1):132–133, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423555.

Donnelly:2004:BRJ

[1842] Russell Donnelly. Book review: Johanna Levelt Sengers: How Fluids Unmix: Discoveries by the School of Van der Waals and Kamerlingh Onnes. Isis, 95(1):133–134, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423556.

Ash:2004:BRK

[1843] Eric H. Ash. Book review: Karen Piper: Cartographic Fictions: Maps, Race, and Identity. Isis, 95(1):134–135, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/423557.

Faragher:2004:BRC

[1844] John Mack Faragher. Book review: Conevery Bolton Valenius: The Health of the Country: How American Settlers Understood Themselves and Their Land. Isis, 95(1):135, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423558.

Smocovitis:2004:BRB

[1845] Vassiliki Betty Smocovitis. Book review: On the Origin of Species. Isis, 95(1):135–137, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423559.

Moran:2004:BRK

[1846] Jeffrey P. Moran. Book review: Karl W. Giberson; Donald A. Yerxa: Species of Origins: America’s Search for a Creation Story. Isis, 95(1): 137, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423560.

Montgomery:2004:BRBa

[1847] William Montgomery. Book review: Ernst Haeckel-Wilhelm Bölsche: Briefwechsel 1887–1919. Isis, 95(1):137–138, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423561.

Camerini:2004:BRM

[1848] Jane R. Camerini. Book review: Michael Shermer: In Darwin’s Shadow: The Life and Science of Alfred Russel Wallace: A Biographical Study on the Psychology of History. Isis, 95(1):138–139, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423562.

Baumann:2004:BRN

[1849] Fred Baumann. Book review: Nasser Behnegar: Leo Strauss, Max Weber, and the Scientific Study of Politics. Isis, 95(1):139–140, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423563.

Yans:2004:BRD

[1850] Virginia Yans. Book review: Derek Freeman: The Fateful Hoaxing of Margaret Mead: A Historical Analysis of Her Samoan Research. Isis, 95(1):140–141, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423564.

Barnhart:2004:BRB

[1851] Terry A. Barnhart. Book review: John D, by Charles Boewe. Clifford’s “Indian Antiquities”: With Related Material by C. S. Rafinesque. Isis, 95(1):141–142, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423565.

Helfand:2004:BRR

[1852] William H. Helfand. Book review: R. Alton Lee: The Bizarre Careers of John R. Brinkley. Isis, 95(1):142–143, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423566.

Edelson:2004:BRT

[1853] Paul J. Edelson. Book review: Terrie M. Romano: Making Medicine Scientific: John Burdon Sanderson and the Culure of Victorian Science. Isis, 95(1):143–144, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423567.

Kneeland:2004:BRB

[1854] Timothy W. Kneeland. Book review: Gracefully Insane: The Rise and Fall of America’s Premier Mental Hospital Creating Mental Illness. Isis, 95(1):144–146, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423568.

Crozier:2004:BRP

[1855] Ivan Crozier. Book review: Paul Lerner: Hysterical Men: War, Psychiatry, and the Politics of Trauma, 1890–1930. Isis, 95(1):146, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423569.

Arens:2004:BRA

[1856] Katherine Arens. Book review: Andreas Mayer: Mikroskopie der Psyche: Die Anfänge der Psychoanalyse im Hypnose-Labor. Isis, 95(1):146–147, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423570.

Lehrer:2004:BRP

[1857] Ronald Lehrer. Book review: Paul Roazen: The Historiography of Psychoanalysis. Isis, 95(1):147–148, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/423571.

Elliott:2004:BRT

[1858] Clark A. Elliott. Book review: Thomas Neville Bonner: Iconoclast: Abraham Flexner and a Life in Learning. Isis, 95(1):148–149, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423572.

Carpenter:2004:BRA

[1859] Kenneth J. Carpenter. Book review: Alan M. Kraut: Goldberger’s War: The Life and Work of a Public Health Crusader. Isis, 95(1):149–150, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423573.

Wolf:2004:BRW

[1860] Jacqueline H. Wolf. Book review: Wendy Mitchinson: Giving Birth in Canada, 1900–1950. Isis, 95(1):150–151, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/423574.

Hawes:2004:BRB

[1861] Joseph Hawes. Book review: Formative Years: Children’s Health in the United States, 1880–2000. Isis, 95(1):151–152, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423575.

Withuhn:2004:BRB

[1862] William L. Withuhn. Book review: The Engineering and History of Rocket: A Survey Report. Isis, 95(1):152–153, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423576.

Israel:2004:BRK

[1863] Paul Israel. Book review: Ken Beauchamp: A History of Telegraphy. Isis, 95(1):153–154, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423577.

Weightman:2004:BRS

[1864] Gavin Weightman. Book review: Sungook Hong: Wireless: From Marconi’s Black-Box to the Audion. Isis, 95(1):154, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423578.

Hunt:2004:BRG

[1865] Bruce J. Hunt. Book review: Gavin Weightman: Signor Marconi’s Magic Box: The Most Remarkable Invention of the Nineteenth Century and the Amateur Inventor Whose Genius Sparked a Revolution. Isis, 95(1):154– 155, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423579.

Teich:2004:BRD

[1866] Albert H. Teich. Book review: Daniel S. Greenberg: Science, Money, and Politics: Political Triumph and Ethical Erosion. Isis, 95(1):155– 156, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423580.

Crowe:2004:BRJ

[1867] Michael J. Crowe. Book review: John F. Moffitt: Picturing Extraterrestrials: Alien Images in Modern Mass Culture. Isis, 95(1):156–157, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423581.

Kaesler:2004:BRB

[1868] Roger L. Kaesler. Book review: Starring T, by José Luis Sanz; Philip Mason. rex! Dinosaur Mythology and Popular Culture. Isis, 95(1):157– 158, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423582.

Pandora:2004:BRM

[1869] Katherine Pandora. Book review: Michael Whitworth: Einstein’s Wake: Relativity, Metaphor, and Modernist Literature. Isis, 95(1):158–159, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423583.

Allen:2004:BRB

[1870] Garland E. Allen. Book review: The Funding of Scientific Racism: Wickliffe Draper and the Pioneer Fund Intelligence, Race, and Genetics: Conversations with Arthur Jensen. Isis, 95(1):159–161, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423584.

Hauger:2004:BRB

[1871] J. Scott Hauger. Book review: Bart Simon: Undead Science: Science Studies and the Afterlife of Cold Fusion. Isis, 95(1):161–162, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423585.

Lewis:2004:BRB

[1872] Albert C. Lewis. Book review: Collected Works Correspondence, A-G Collected Works Correspondence, H-Z. Isis, 95(1):162–163, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423586.

DeVorkin:2004:BRM

[1873] David DeVorkin. Book review: Michael W. Friedlander: A Thin Cosmic Rain: Particles from Outer Space. Isis, 95(1):163, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423587.

Kidwell:2004:BRD

[1874] Peggy Aldrich Kidwell. Book review: Dorrit Hoffleit: Misfortunes as Blessings in Disguise: The Story of My Life. Isis, 95(1):163, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423588.

Stranges:2004:BRA

[1875] Anthony N. Stranges. Book review: Ahmed Zewail: Voyage through Time: Walks of Life to the Nobel Prize. Isis, 95(1):164–165, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423589.

Figueiroa:2004:BRBa

[1876] Silvia Figueirôa. Book review: A Passion for Gold: An Autobiography. Isis, 95(1):165, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423590.

Newcomb:2004:BRD

[1877] Sally Newcomb. Book review: Davis A. Young: Mind over Magma: The Story of Igneous Petrology. Isis, 95(1):165–166, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423591.

Schnare:2004:BRK

[1878] Susan E. Schnare. Book review: Karl Sabbagh: A Rum Affair: A True Story of Botanical Fraud. Isis, 95(1):166, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/423592.

Smith:2004:BRAa

[1879] Adrian Smith. Book review: Andrew Jamison: The Making of Green Knowledge: Environmental Politics and Cultural Transformation. Isis, 95(1):166–167, March 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
423593.

McElheny:2004:BRC

[1880] Victor McElheny. Book review: Celeste Michelle Condit: The Meanings of the Gene: Public Debates about Human Heredity. Isis, 95(1):167– 168, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423594.

Guerrini:2004:BRD

[1881] Anita Guerrini. Book review: Deborah Rudacille: The Scalpel and the Butterfly: The Conflict between Animal Research and Animal Protection. Isis, 95(1):168–169, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423595.

Matthews:2004:BRD

[1882] J. Rosser Matthews. Book review: N. D. A. Kemp: “Merciful Release”: The History of the British Euthanasia Movement. Isis, 95(1):169–170, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423596.

Humphreys:2004:BRJ

[1883] Margaret Humphreys. Book review: James C. Whorton: Nature Cures: The History of Alternative Medicine in America. Isis, 95(1):170, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423597.

Smith:2004:BRAb

[1884] Robert W. Smith. Book review: Andrew J. Dunar; Stephen P. Waring: Power to Explore: A History of Marshall Space Flight Center, 1960– 1990. Isis, 95(1):171, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423598.

Low:2004:BRB

[1885] Morris Low. Book review: A Social History of Science and Technology in Contemporary Japan The Occupation Period, 1945–1952. Isis, 95(1):171– 172, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423599.

Stotz:2004:BRG

[1886] Karola Stotz. Book review: Gordon Graham: Genes: A Philosophical Inquiry. Isis, 95(1):172–173, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/423600.

Magnello:2004:BRB

[1887] Eileen Magnello. Book review: Mathematics in a Postmodern Age: A Christian Perspective. Isis, 95(1):173, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/423601.

Bowler:2004:BRD

[1888] Peter J. Bowler. Book review: David L. Hull: Science and Selection: Essays on Biological Evolution and the Philosophy of Science. Isis, 95 (1):174, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/423602.

Anonymous:2004:BRBa

[1889] Anonymous. Book review: The Medieval Heritage in Early Modern Metaphysics and Modal Theory, 1400–1700. Isis, 95(1):174–175, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423603.

Anonymous:2004:BRBb

[1890] Anonymous. Book review: Classics on Fractals. Isis, 95(1):175, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/423604.

Anonymous:2004:NCa

[1891] Anonymous. Notes on contributors. Isis, 95(1):176–181, March 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/423605.

Cittadino:2004:BSE

[1892] Eugene Cittadino. Borderline science: Expert testimony and the red river boundary dispute. Isis, 95(2):183–219, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426194.

Stothers:2004:ASB

[1893] Richard B. Stothers. Ancient scientific basis of the “Great Serpent” from historical evidence. Isis, 95(2):220–238, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/426195.

Tweedale:2004:CVC

[1894] Geoffrey Tweedale and Jock McCulloch. Chrysophiles versus chrysophobes: The white asbestos controversy, 1950s–2004. Isis, 95(2):239– 259, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426196.

Nauenberg:2004:EBB

[1895] Michael Nauenberg and Bruce Pourciau. Eloge: Bruce Brackenridge, 1927–2003. Isis, 95(2):260–262, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426197.

Blair:2004:HPC

[1896] Ann Blair. 2003 HSS Prize Citations: Watson Davis and Helen Miles Davis Prize, Joseph H. Hazen Education Prize, History of Women in Science Prize, Pfizer Prize, Derek Price/Rod Webster Prize, Henry and Ida Schuman Prize, Sarton Medal. Isis, 95(2):263–269, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426198.

Chalk:2004:R

[1897] Rosemary Chalk. In reply. Isis, 95(2):270, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/426200.

Moran:2004:E

[1898] Gordon Moran. To the editor. Isis, 95(2):270, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426199.

Pasachoff:2004:E

[1899] Jay M. Pasachoff. To the editor. Isis, 95(2):270, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426201.

Richards:2004:E

[1900] Robert J. Richards. To the editor. Isis, 95(2):270–271, June 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426202.

Ihsanoglu:2004:E

[1901] Ekmeleddîn Ihsanoglu. To the editor. Isis, 95(2):272–274, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426203.

Sengor:2004:R

[1902] A. M. C. Sengör. In reply. Isis, 95(2):273–274, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426204.

McCartney:2004:BRB

[1903] Mark McCartney. Book review: Irish Innovators in Science and Technology. Isis, 95(2):275, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426205.

Carron:2004:BRE

[1904] Jean-Claude Carron. Book review: Emmanuel Faye: Philosophie et perfection de l’homme: De la Renaissance à Descartes. Isis, 95(2):275–276, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426206.

Taub:2004:BRB

[1905] Liba Taub. Book review: Treasures of the Science Museum. Isis, 95(2): 277, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426207.

Montgomery:2004:BRBb

[1906] Scott L. Montgomery. Book review: Visualizations. Isis, 95(2):277– 279, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426208.

Gray:2004:BRC

[1907] Jeremy Gray. Book review: Christian Houzel: La géométrie algébrique: Recherches historiques. Isis, 95(2):279, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426209.

Barrow-Green:2004:BRC

[1908] June Barrow-Green. Book review: Charles Seife: Zero: The Biography of a Dangerous Idea. Isis, 95(2):279–280, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426210.

Farber:2004:BRT

[1909] Paul Lawrence Farber. Book review: Tim Birkhead: A Brand-New Bird: How Two Amateur Scientists Created the First Genetically Engineered Animal. Isis, 95(2):280–281, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426211.

Nyhart:2004:BRB

[1910] Lynn K. Nyhart. Book review: Die Kulturgeschichte des Zoos. Isis, 95(2): 281–282, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426212.

Sweet:2004:BRB

[1911] Victoria Sweet. Book review: Western Medical Thought from Antiquity to the Middle Ages. Isis, 95(2):282–283, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/426213.

Lohff:2004:BRG

[1912] Brigitte Lohff. Book review: Gabriele Beisswanger; Gudrun Hahn; Evelyn Seibert; Ildikó Szász; Christl Trischler: Frauen in der Pharmazie: Die Geschichte eines Frauenberufs. Isis, 95(2):283, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426214.

Lautner:2004:BRB

[1913] Peter Lautner. Book review: Eudemus of Rhodes. Isis, 95(2):283–285, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426215.

Rutten:2004:BRB

[1914] Thomas Rütten. Book review: Hippocrates. Isis, 95(2):285–286, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426216.

Rihll:2004:BRR

[1915] T. E. Rihll. Book review: Rosalind Thomas: Herodotus in Context: Ethnography, Science, and the Art of Persuasion. Isis, 95(2):286–287, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426217.

Nauta:2004:BRB

[1916] Lodi Nauta. Book review: Letters. Isis, 95(2):287–288, June 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426218.

Jones:2004:BRW

[1917] Matthew L. Jones. Book review: W. R. Laird: The Unfinished Mechanics of Giuseppe Moletti: An Edition and English Translation of His Dialogue on Mechanics. Isis, 95(2):288–289, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426219.

Lynch:2004:BRM

[1918] William T. Lynch. Book review: Marie Boas Hall: Henry Oldenburg: Shaping the Royal Society. Isis, 95(2):289–290, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426220.

Wilson:2004:BRR

[1919] Philip K. Wilson. Book review: Richard C. Allen: David Hartley on Human Nature. Isis, 95(2):290, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426221.

Noske:2004:BRI

[1920] Rainer Noske. Book review: Ingrid Hupp: Arithmetik- und Algebralehrbücher Würzburger Mathematiker des 18. Jahrhunderts. Isis, 95 (2):290–291, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426222.

Gozza:2004:BRB

[1921] Paolo Gozza. Book review: Roger North’s Writings on Music, c, by Mary Chan; Jamie C. Kassler; Janet D. Hine; Mary Chan; Jamie C. Kassler; Janet D. Hine. 1704–c. 1709 Roger North’s “Of Sounds” and Prendcourt Tracts, c. 1710–c. 1716. Isis, 95(2):291–293, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426223.

Graziano:2004:BRP

[1922] Amy Graziano. Book review: Penelope Gouk: Music, Science, and Natural Magic in Seventeenth-Century England. Isis, 95(2):293, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426224.

Belcastro:2004:BRB

[1923] Alessandro Belcastro. Book review: Die Werke von Daniel Bernoulli. Isis, 95(2):293–294, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426225.

Shank:2004:BRS

[1924] J. B. Shank. Book review: Scott Mandelbrote: Footprints of the Lion: Isaac Newton at Work. Isis, 95(2):294–296, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/426226.

Dijksterhuis:2004:BRB

[1925] Fokko Jan Dijksterhuis. Book review: Optics: Paralipomena to Witelo and Optical Part of Astronomy. Isis, 95(2):296–297, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426227.

Ogilvie:2004:BRB

[1926] Brian W. Ogilvie. Book review: The World in a Box: The Story of an Eighteenth-Century Picture Encyclopedia. Isis, 95(2):297–298, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426228.

Turner:2004:BRJ

[1927] Steven Turner. Book review: Joyce E. Chaplin: Subject Matter: Technology, the Body, and Science on the Anglo-American Frontier, 1500–1676. Isis, 95(2):298–299, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426229.

Yearley:2004:BRN

[1928] Steven Yearley. Book review: Nicholas Whyte: Science, Colonialism, and Ireland. Isis, 95(2):299–300, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426230.

Greene:2004:BRS

[1929] John Greene. Book review: Silvio A. Bedini: Jefferson and Science. Isis, 95(2):300–301, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426231.

Wyatt:2004:BRP

[1930] Antony Wyatt. Book review: Peter Allen: A Geological Survey in Transition. Isis, 95(2):301–302, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426232.

Branch:2004:BRL

[1931] Michael P. Branch. Book review: Laura Dassow Walls: Emerson’s Life in Science: The Culture of Truth. Isis, 95(2):302, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426233.

Schoepflin:2004:BRC

[1932] Rennie B. Schoepflin. Book review: Craig James Hazen: The Village Enlightenment in America: Popular Religion and Science in the Nineteenth Century. Isis, 95(2):303–304, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426234.

Robey:2004:BRB

[1933] Ethan Robey. Book review: Bruno Giberti: Designing the Centennial: A History of the 1876 International Exhibition in Philadelphia. Isis, 95 (2):304, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426235.

Schubring:2004:BRB

[1934] Gert Schubring. Book review: Naturwissenschaften um 1800: Wissenschaftskultur in Jena-Weimar. Isis, 95(2):305–306, June 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426236.

Hume:2004:BRB

[1935] Brad D. Hume. Book review: Barry Alan Joyce: The Shaping of American Ethnography: The Wilkes Exploring Expedition, 1838–1842. Isis, 95(2):306–307, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426237.

Kubbinga:2004:BRA

[1936] Henk Kubbinga. Book review: Ad J. P. Maas: Atomisme en individualisme: De Amsterdamse Natuurkunde tussen 1877 en 1940. Isis, 95(2): 307–308, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426238.

Worboys:2004:BRM

[1937] Michael Worboys. Book review: Mark W. Weatherall: Gentlemen, Scientists, and Doctors: Medicine at Cambridge, 1800–1940. Isis, 95(2):

308–309, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426239.

Quinn:2004:BRB

[1938] Philip L. Quinn. Book review: Bill Cooke: A Rebel to His Last Breath: Joseph McCabe and Rationalism. Isis, 95(2):309–310, June 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426240.

Benson:2004:BRE

[1939] Maxine Benson. Book review: Ellen E. Wohl: Virtual Rivers: Lessons from the Mountain Rivers of the Colorado Front Range. Isis, 95(2): 310, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426241.

Coodley:2004:BRC

[1940] Lauren Coodley. Book review: Carolyn Merchant: Green versus Gold: Sources in California’s Environmental History. Isis, 95(2):310–311, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426242.

Dean:2004:BRF

[1941] Sheila Ann Dean. Book review: Frank N. Egerton: Hewett Cottrell Watson: Victorian Plant Ecologist and Evolutionist. Isis, 95(2):311– 312, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426243.

Mooney:2004:BRE

[1942] Graham Mooney. Book review: Evelynn Maxine Hammonds: Childhood’s Deadly Scourge: The Campaign to Control Diphtheria in New York City, 1880–1930. Isis, 95(2):312–313, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426244.

Olson:2004:BRM

[1943] Richard Olson. Book review: Martin S. Staum: Labeling People: French Scholars on Society, Race, and Empire, 1815–1848. Isis, 95(2):313–314, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426245.

Fritscher:2004:BRT

[1944] Bernhard Fritscher. Book review: Thomas Junker; Uwe Hoßfeld: Die Entdeckung der Evolution: Eine revolutionäre Theorie und ihre Geschichte. Isis, 95(2):314–315, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426246.

Benbow:2004:BRE

[1945] Mary Benbow. Book review: Elizabeth Hanson: Animal Attractions: Nature on Display in American Zoos. Isis, 95(2):315–316, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426247.

Flesher:2004:BRG

[1946] Mary Mosher Flesher. Book review: Gerald Sweeney: “Fighting for the Good Cause”: Reflections on Francis Galton’s Legacy to American Hereditarian Psychology. Isis, 95(2):316–317, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426248.

Gruenbaum:2004:BRN

[1947] Ellen Gruenbaum. Book review: Nancy Rose Hunt: A Colonial Lexicon of Birth Ritual, Medicalization, and Mobility in the Congo. Isis, 95(2): 317–318, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426249.

Harrison:2004:BRL

[1948] Mark Harrison. Book review: Lise Wilkinson; Anne Hardy: Prevention and Cure: The London School of Hygiene and Tropical Medicine: A Twentieth-Century Quest for Global Public Health. Isis, 95(2):318– 319, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426250.

Killen:2004:BRC

[1949] Andreas Killen. Book review: Carolyn Thomas de la Peña: The Body Electric: How Strange Machines Built the Modern American. Isis, 95 (2):319, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426251.

Channell:2004:BRG

[1950] David F. Channell. Book review: Gillian Cookson; Colin Hempstead: A Victorian Scientist and Engineer: Fleeming Jenkin and the Birth of Electrical Engineering. Isis, 95(2):319–320, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426252.

Blaszczyk:2004:BRA

[1951] Regina Lee Blaszczyk. Book review: Amy K. Glasmeier: Manufacturing Time: Global Competition in the Watch Industry, 1795–2000. Isis, 95(2): 320–321, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426253.

Aracil:2004:BRA

[1952] Javier Aracil. Book review: Agustí Nieto-Galan: La seducción de la máquina: Santponç, Monturiol, Isaac Peral: Vapores, submarinos e inventores. Isis, 95(2):321–322, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426254.

Tucker:2004:BRB

[1953] Jennifer Tucker. Book review: African-American Pioneers in Anthropology. Isis, 95(2):322–324, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426255.

Needell:2004:BRP

[1954] Allan A. Needell. Book review: Peter J. Westwick: The National Labs: Science in an American System, 1947–1974. Isis, 95(2):324–325, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426256.

James:2004:BRC

[1955] Kathryn James. Book review: Chaomei Chen: Mapping Scientific Frontiers: The Quest for Knowledge Visualization. Isis, 95(2):325, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426257.

Trenn:2004:BRJ

[1956] Thaddeus J. Trenn. Book review: Johannes Abele: “Wachhund des Atomzeitalters”: Geigerzähler in der Geschichte des Strahlenschutzes. Isis, 95(2):325–327, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426258.

Brown:2004:BRB

[1957] Laurie M. Brown. Book review: Brian Greene: The Elegant Universe: Superstrings, Hidden Dimensions, and the Quest for the Ultimate Theory. Isis, 95(2):327, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426259.

Jennings:2004:BRJ

[1958] Mark R. Jennings. Book review: Joseph E. Taylor III: Making Salmon: An Environmental History of the Northwest Fisheries Crisis. Isis, 95(2): 327–328, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426260.

McCalman:2004:BRJ

[1959] Janet McCalman. Book review: Julie Marcus: The Indomitable Miss Pink: A Life in Anthropology. Isis, 95(2):328–329, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426261.

Lowy:2004:BRBa

[1960] Ilana Löwy. Book review: Science as Autobiography: The Troubled Life of Niels Jerne. Isis, 95(2):329–330, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426262.

Brooks:2004:BRJa

[1961] Nathan M. Brooks. Book review: Jeanne Guillemin: Anthrax: The Investigation of a Deadly Outbreak. Isis, 95(2):330–331, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426263.

Cole:2004:BRD

[1962] Leonard A. Cole. Book review: David Koplow: Smallpox: The Fight to Eradicate a Global Scourge. Isis, 95(2):331, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/426264.

Kay:2004:BRR

[1963] Gwen Kay. Book review: Robyn L. Rosen: Reproductive Health, Reproductive Rights: Reformers and the Politics of Maternal Welfare, 1917– 1940. Isis, 95(2):331–332, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426265.

Weiss:2004:BRH

[1964] Sheila Faith Weiss. Book review: Heiner Fangerau: Etablierung eines rassenhygienischen Standardwerkes 1921–1941: Der Baur-Fischer-Lenz im Spiegel der zeitgenössischen Rezensionsliteratur. Isis, 95(2):332–333, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426266.

Williams:2004:BRC

[1965] Michael R. Williams. Book review: Christine A. Finn: Artifacts: An Archaeologist’s Year in Silicon Valley. Isis, 95(2):333, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426267.

Kidwell:2004:BRJ

[1966] Peggy Aldrich Kidwell. Book review: Jeffrey R. Yost: A Bibliographic Guide to Resources in Scientific Computing, 1945–1975. Isis, 95(2):334, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426268.

Bishop:2004:BRS

[1967] Robert C. Bishop. Book review: Steen Brock: Niels Bohr’s Philosophy of Quantum Physics in the Light of the Helmholtzian Tradition of Theoretical Physics. Isis, 95(2):334–335, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426269.

Brock:2004:BRS

[1968] Steen Brock. Book review: Shimon Malin: Nature Loves to Hide: Quantum Physics and the Nature of Reality: A Western Perspective. Isis, 95(2):335–336, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426270.

Ragep:2004:BRB

[1969] F. Jamil Ragep. Book review: History of Islamic Philosophy. Isis, 95(2): 336–337, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426271.

Kimler:2004:BRJ

[1970] William Kimler. Book review: John A. Moore: From Genesis to Genetics: The Case of Evolution and Creationism. Isis, 95(2):337–338, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426272.

Darden:2004:BRM

[1971] Lindley Darden. Book review: Michael Ruse: Darwin and Design: Does Evolution Have a Purpose? Isis, 95(2):338–339, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426273.

Kruytbosch:2004:BRB

[1972] Carlos Kruytbosch. Book review: Science Bought and Sold: Essays in the Economics of Science. Isis, 95(2):339–341, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/426274.

Lewenstein:2004:BRR

[1973] Bruce V. Lewenstein. Book review: Robert L. Park: Voodoo Science: The Road from Foolishness to Fraud. Isis, 95(2):341, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426275.

Bazerman:2004:BRA

[1974] Charles Bazerman. Book review: Alan G. Gross; Joseph E. Harmon; Michael Reidy: Communicating Science: The Scientific Article from the Seventeenth Century to the Present. Isis, 95(2):341–342, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426276.

Anonymous:2004:BRBc

[1975] Anonymous. Book review: Revisiting the Foundations of Relativistic Physics: Festschrift in Honor of John Stachel. Isis, 95(2):342–343, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426277.

Anonymous:2004:BRBd

[1976] Anonymous. Book review: La ilustración y las ciencias: Para una historia de la objetividad. Isis, 95(2):343, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426278.

Anonymous:2004:BRBe

[1977] Anonymous. Book review: Mathematics and War. Isis, 95(2):343, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426279.

Anonymous:2004:BRBf

[1978] Anonymous. Book review: Celebrating the Life and Times of Hugh Miller: Scotland in the Early Nineteenth Century. Isis, 95(2):343–344, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426280.

Anonymous:2004:BRBg

[1979] Anonymous. Book review: Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. Isis, 95(2):344, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426281.

Anonymous:2004:BRBh

[1980] Anonymous. Book review: Writing the History of Mathematics: Its Historical Development. Isis, 95(2):344, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426282.

Anonymous:2004:BRBi

[1981] Anonymous. Book review: Ceteris Paribus Laws. Isis, 95(2):344, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/426283.

Anonymous:2004:BRBj

[1982] Anonymous. Book review: Essays on Early Medieval Mathematics: The Latin Tradition. Isis, 95(2):344–345, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426284.

Anonymous:2004:BRBk

[1983] Anonymous. Book review: Observation and Experiment in the Natural and Social Sciences. Isis, 95(2):345, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426285.

Anonymous:2004:BRBl

[1984] Anonymous. Book review: Nature in the Global South: Environmental Projects in South and Southeast Asia. Isis, 95(2):345, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426286.

Anonymous:2004:BRBm

[1985] Anonymous. Book review: Die “Wissenschaft Hinter dem eisernen Vorhang”: Realitaten und Mythen der Sowjetischen Wissenschaft. Isis, 95(2):345–346, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426287.

Anonymous:2004:BRBn

[1986] Anonymous. Book review: The Scientific Revolution: The Essential Readings. Isis, 95(2):346, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426288.

Anonymous:2004:BRBo

[1987] Anonymous. Book review: The Cambridge Companion to Darwin. Isis, 95(2):346, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426289.

Anonymous:2004:BRBp

[1988] Anonymous. Book review: The Enterprise of Science in Islam: New Perspectives. Isis, 95(2):346, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426290.

Anonymous:2004:BRBq

[1989] Anonymous. Book review: A History of Analysis. Isis, 95(2):346–347, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/426291.

Anonymous:2004:BRBr

[1990] Anonymous. Book review: Reconsidering a Century of Flight. Isis, 95 (2):347, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426292.

Anonymous:2004:BRBs

[1991] Anonymous. Book review: The Power of Images in Early Modern Science. Isis, 95(2):347, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426293.

Anonymous:2004:BRBt

[1992] Anonymous. Book review: When Science and Christianity Meet. Isis, 95 (2):347, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/426294.

Anonymous:2004:BRBu

[1993] Anonymous. Book review: The Treatise on Perspective: Published and Unpublished. Isis, 95(2):347–348, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426295.

Anonymous:2004:BRBv

[1994] Anonymous. Book review: Opioids and Pain Relief: A Historical Perspective. Isis, 95(2):348, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/426296.

Anonymous:2004:BRBw

[1995] Anonymous. Book review: Johannes Kepler, Collection of Articles, by N. I. Nevskaia. Vol. 2: The Works about Kepler in Russia and Germany. Isis, 95(2):348, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426297.

Anonymous:2004:BRBx

[1996] Anonymous. Book review: Magic in History: Prayer, Magic, and the Stars in the Ancient and Late Antique World. Isis, 95(2):348, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426298.

Anonymous:2004:BRBy

[1997] Anonymous. Book review: The Confinement of the Insane: International Perspectives, 1800–1965. Isis, 95(2):348–349, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/426299.

Anonymous:2004:BRBaa

[1998] Anonymous. Book review: Renaissance Astrolabes and Their Makers. Isis, 95(2):349, June 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
426301.

Anonymous:2004:BRBz

[1999] Anonymous. Book review: Niccolò Stenone (1638–1686): Anatomista, geologo, vescovo. Isis, 95(2):349, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/426300.

Anonymous:2004:NCb

[2000] Anonymous. Notes on contributors. Isis, 95(2):350–355, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/426302.

Lightman:2004:E

[2001] Bernard Lightman. Editorial. Isis, 95(3):357–358, September 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428958.

Rentetzi:2004:GPR

[2002] Maria Rentetzi. Gender, politics, and radioactivity research in Interwar Vienna: The case of the Institute for Radium Research. Isis, 95(3): 359–393, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428960.

Kenny:2004:RSS

[2003] Michael G. Kenny. Racial science in social context: John r. Baker on eugenics, race, and the public role of the scientist. Isis, 95(3):394–419, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/428959.

Blair:2004:EMP

[2004] Ann Blair. An early Modernist’s perspective. Isis, 95(3):420–430, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/428961.

Topham:2004:VIA

[2005] Jonathan R. Topham. A view from the industrial age. Isis, 95(3):431–442, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/428962.

Daston:2004:TN

[2006] Lorraine Daston. Taking note(s). Isis, 95(3):443–448, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428963.

Lesch:2004:EGL

[2007] John E. Lesch. Eloge: Gerald Lynn Geison, 1943–2001. Isis, 95(3): 449–451, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428965.

Rosenberg:2004:EOT

[2008] Charles E. Rosenberg. Eloge: Owsei Temkin, 1902–2002. Isis, 95(3): 451–453, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428966.

Richards:2004:ESE

[2009] Robert J. Richards. Eloge: Susan Elizabeth Abrams, 1945–2003. Isis, 95 (3):453–454, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428964.

Barnhart:2004:R

[2010] Terry A. Barnhart. In reply. Isis, 95(3):455, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428971.

Boewe:2004:E

[2011] Charles Boewe. To the Editor. Isis, 95(3):455, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428970.

Hunt:2004:R

[2012] Bruce J. Hunt. In reply. Isis, 95(3):457, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428969.

Weightman:2004:E

[2013] Gavin Weightman. To the Editor. Isis, 95(3):457, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428968.

Baumann:2004:E

[2014] Fred Baumann. To the Editor. Isis, 95(3):459, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428967.

Hacking:2004:BRB

[2015] Ian Hacking. Book review: The Science of Conjecture: Evidence and Probability before Pascal. Isis, 95(3):460–464, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428972.

Pantalony:2004:BRB

[2016] David Pantalony. Book review: Science in the Provinces: A Descriptive Catalogue of the “Deventer” Collection. Isis, 95(3):465, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428973.

Kleiner:2004:BRB

[2017] Israel Kleiner. Book review: Sherlock Holmes in Babylon and Other Tales of Mathematical History. Isis, 95(3):465–466, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428974.

Sobol:2004:BRA

[2018] Peter G. Sobol. Book review: Anthony Aveni: Behind the Crystal Ball: Magic, Science, and the Occult from Antiquity through the New Age. Isis, 95(3):466–467, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428975.

Rigden:2004:BRP

[2019] John S. Rigden. Book review: Philip Ball: The Ingredients: A Guided Tour of the Elements. Isis, 95(3):467–468, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428976.

Nieto-Galan:2004:BRP

[2020] Agustí Nieto-Galan. Book review: Philip Ball: Bright Earth: Art and the Invention of Color. Isis, 95(3):468–469, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428977.

Calinger:2004:BRW

[2021] Ronald Calinger. Book review: William P. Berlinghoff; Fernando Q. Gouvêa: Math through the Ages: A Gentle History for Teachers and Others. Isis, 95(3):469, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428978.

DiBono:2004:BRB

[2022] Mario Di Bono. Book review: I lettori di astronomia: Presso lo Studio di Bologna dal XII al XX secolo Seventh Centenary of the Teaching of Astronomy in Bologna, 1297–1997. Isis, 95(3):469–470, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428979.

Dietrich:2004:BRL

[2023] Michael R. Dietrich. Book review: Leah Ceccarelli: Shaping Science with Rhetoric: The Cases of Dobzhansky, Schrödinger, and Wilson. Isis, 95 (3):470–471, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428980.

Roland:2004:BRA

[2024] Alex Roland. Book review: Alfred W. Crosby: Throwing Fire: Projectile Technology through History. Isis, 95(3):471–472, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428981.

Bedini:2004:BRB

[2025] Silvio A. Bedini. Book review: Scientific Instruments: Originals and Imitations. Isis, 95(3):472–473, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/428982.

Forsyth:2004:BRJ

[2026] Patrick Forsyth. Book review: John Gage: Color and Meaning: Art, Science, and Symbolism. Isis, 95(3):473–474, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428983.

Bray:2004:BRB

[2027] Francesca Bray. Book review: Science and Civilisation in China, by Peter J. Golas. Volume 5: Chemistry and Chemical Technology; Part 13: Mining. Isis, 95(3):474–475, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/428984.

Pulliam:2004:BRG

[2028] Gregory Pulliam and John Snapper. Book review: Giorgio Graffi: Two Hundred Years of Syntax: A Critical Survey. Isis, 95(3):475–476, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/428985.

Wise:2004:BRA

[2029] Steven Wise. Book review: Anita Guerrini: Experimenting with Humans and Animals: From Galen to Animal Rights. Isis, 95(3):476–477, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/428986.

Leary:2004:BRR

[2030] William M. Leary. Book review: Richard P. Hallion: Taking Flight: Inventing the Aerial Age from Antiquity through the First World War. Isis, 95(3):477, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428987.

Greene:2004:BRBa

[2031] John C. Greene. Book review: The Oxford Companion to the History of Modern Science. Isis, 95(3):477–478, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428988.

Smith:2004:BRW

[2032] David C. Smith. Book review: De Witt Douglas Kilgore: Astrofuturism: Science, Race, and Visions of Utopia in Space. Isis, 95(3):478–479, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/428989.

Whitney:2004:BRP

[2033] Elspeth Whitney. Book review: Pamela O. Long: Openness, Secrecy, Authorship: Technical Arts and the Culture of Knowledge from Antiquity to the Renaissance. Isis, 95(3):479–480, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428990.

Runte:2004:BRC

[2034] Alfred Runte. Book review: Carolyn Merchant: The Columbia Guide to American Environmental History. Isis, 95(3):480–481, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/428991.

Figueiroa:2004:BRH

[2035] Silvia F. de M. Figueirôa. Book review: Hugh Raffles: In Amazonia: A Natural History. Isis, 95(3):481–482, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428992.

Lucier:2004:BRG

[2036] Paul Lucier. Book review: George E. Webb: Science in the American Southwest: A Topical History. Isis, 95(3):482, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/428993.

Leonard:2004:BRB

[2037] Robert Leonard. Book review: The Future of the History of Economics. Isis, 95(3):482–483, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428994.

Bhayro:2004:BRB

[2038] Siam Bhayro. Book review: The Science of Man in Ancient Greece. Isis, 95(3):483–484, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428995.

Goodrum:2004:BRG

[2039] Matthew R. Goodrum. Book review: Gary Urton: Signs of the Inka Khipu: Binary Coding in the Andean Knotted-String Records. Isis, 95 (3):484–485, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
428996.

Payne:2004:BRB

[2040] Lynda Payne. Book review: Historia corporis humani sive Anatomice Paper Bodies: A Catalogue of Anatomical Fugitive Sheets, 1538–1687. Isis, 95(3):485–486, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428997.

Eastwood:2004:BRI

[2041] Bruce Eastwood. Book review: Irene Caiazzo: Lectures médiévales de Macrobe: Les Glosae Colonienses super Macrobium: Étude et édition. Isis, 95(3):486–487, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/428998.

Dupre:2004:BRJ

[2042] Sven Dupré. Book review: John North: The Ambassadors’ Secret: Holbein and the World of the Renaissance. Isis, 95(3):487–488, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/428999.

Nauert:2004:BRU

[2043] Charles Nauert. Book review: Urszula Szulakowska: The Alchemy of Light: Geometry and Optics in Late Renaissance Alchemical Illustration. Isis, 95(3):488, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429000.

Findlen:2004:BRS

[2044] Paula Findlen. Book review: Simone Contardi: La casa di Salomone a Firenze: L’Imperiale e Reale Museo di Fisico e Storia Naturale (1755– 1801). Isis, 95(3):488–489, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429001.

Davies:2004:BRA

[2045] Gordon L. Herries Davies. Book review: Alan Cutler: The Seashell on the Mountaintop: A Story of Science, Sainthood, and the Humble Genius Who Discovered a New History of the Earth. Isis, 95(3):489–490, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429002.

Laird:2004:BRB

[2046] W. R. Laird. Book review: Le mecaniche. Isis, 95(3):490–492, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429003.

Tannenbaum:2004:BRG

[2047] Rebecca Tannenbaum. Book review: Gerald N. Grob: The Deadly Truth: A History of Disease in America. Isis, 95(3):492, September 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429004.

Cook:2004:BRM

[2048] Harold J. Cook. Book review: Margaret Pelling: Medical Conflicts in Early Modern London: Patronage, Physicians, and Irregular Practitioners, 1550–1640. Isis, 95(3):492–493, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429005.

Lawrence:2004:BRR

[2049] Christopher Lawrence. Book review: Roy Porter: Flesh in the Age of Reason: The Modern Foundations of Body and Soul. Isis, 95(3):493–494, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429006.

Finocchiaro:2004:BRW

[2050] Maurice A. Finocchiaro. Book review: William R. Shea; Mariano Artigas: Galileo in Rome: The Rise and Fall of a Troublesome Genius. Isis, 95(3):494–495, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429007.

Canizares-Esguerra:2004:BRBa

[2051] Jorge Cañizares-Esguerra. Book review: The Torner Collections of Sessé and Mociño: Biological Illustrations. Isis, 95(3):495–496, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429008.

Gelbart:2004:BRH

[2052] Nina Rattner Gelbart. Book review: Holly Tucker: Pregnant Fictions: Childbirth and the Fairy Tale in Early-Modern France. Isis, 95(3): 496–497, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429009.

Musselman:2004:BRW

[2053] Elizabeth Green Musselman. Book review: Warwick Anderson: The Cultivation of Whiteness: Science, Health, and Racial Destiny in Australia. Isis, 95(3):497–498, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429010.

Brooks:2004:BRJb

[2054] Nathan M. Brooks. Book review: James T. Andrews: Science for the Masses: The Bolshevik State, Public Science, and the Popular Imagination in Soviet Russia, 1917–1934. Isis, 95(3):498–499, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429011.

Schmidt:2004:BRS

[2055] Benjamin Schmidt. Book review: Stephen Dow Beckham; Doug Erickson; Jeremy Skinner; Paul Merchant: The Literature of the Lewis and Clark Expedition: A Bibliography and Essays. Isis, 95(3):499–500, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429012.

Scholl:2004:BRJ

[2056] Lars U. Scholl. Book review: Jobst Broelmann: Intuition und Wissenschaft in der Kreiseltechnik, 1750 bis 1930. Isis, 95(3):500, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429013.

Rice:2004:BRB

[2057] Adrian Rice. Book review: The Political Pamphlets and Letters of Charles Lutwidge Dodgson and Related Pieces: A Mathematical Approach. Isis, 95(3):500–501, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429014.

Castonguay:2004:BRT

[2058] Stéphane Castonguay and Darin S. Kinsey. Book review: Thomas R. Dunlap: Nature and the English Diaspora: Environment and History in the United States, Canada, Australia, and New Zealand. Isis, 95(3): 501–502, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429015.

Sorensen:2004:BRT

[2059] Conner Sorensen. Book review: Thomas Eisner: For Love of Insects. Isis, 95(3):502–503, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429016.

Blustein:2004:BRA

[2060] Bonnie Ellen Blustein. Book review: Amy L. Fairchild: Science at the Borders: Immigrant Medical Inspection and the Shaping of the Modern Industrial Labor Force. Isis, 95(3):503–504, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429017.

Ertter:2004:BRP

[2061] Barbara Ertter. Book review: Paul Lawrence Farber: Finding Order in Nature: The Naturalist Tradition from Linnaeus to E. O. Wilson. Isis, 95 (3):504–505, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429018.

Warner:2004:BRB

[2062] Deborah Warner. Book review: Caroline Herschel’s Autobiographies The Herschel Partnership as Viewed by Caroline. Isis, 95(3):505, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429019.

Barnes:2004:BRB

[2063] Michael Horace Barnes. Book review: The Lost Itinerary of Frank Hamilton Cushing. Isis, 95(3):505–506, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/429020.

Stranges:2004:BRB

[2064] Anthony N. Stranges. Book review: Instruments and Experimentation in the History of Chemistry. Isis, 95(3):506–507, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429021.

Herbert:2004:BRB

[2065] Sandra Herbert. Book review: Fossils, Finches, and Fuegians: Darwin’s Adventures and Discoveries on the Beagle. Isis, 95(3):507–508, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429022.

Rusnock:2004:BRB

[2066] Andrea Rusnock. Book review: The Statistical Mind in a Pre-Statistical Era: The Netherlands, 1750–1850. Isis, 95(3):508–509, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429023.

Kingsland:2004:BRR

[2067] Sharon Kingsland. Book review: Robert E. Kohler: Landscapes and Labscapes: Exploring the Lab–Field Border in Biology. Isis, 95(3):509–510, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429024.

Cole:2004:BRL

[2068] Simon A. Cole. Book review: Lawrence Frank: Victorian Detective Fiction and the Nature of Evidence: The Scientific Investigations of Poe, Dickens, and Doyle. Isis, 95(3):510–511, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/429025.

Stam:2004:BRE

[2069] Henderikus J. Stam. Book review: Elizabeth Lunbeck; Bennett Simon: Family Romance, Family Secrets: Case Notes from an American Psychoanalysis, 1912. Isis, 95(3):511–512, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429026.

Kaiser:2004:BRB

[2070] David Kaiser. Book review: Selected Correspondence (1911–1946) of Karl Mannheim, Scientist, Philosopher, and Sociologist. Isis, 95(3): 512–513, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429027.

Walls:2004:BRD

[2071] Laura Dassow Walls. Book review: David E. Nye: America as Second Creation: Technology and Narratives of New Beginnings. Isis, 95(3): 513–514, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429028.

Regal:2004:BRN

[2072] Brian Regal. Book review: Noel Keith Roberts: From Piltdown Man to Point Omega: The Evolutionary Theory of Teilhard De Chardin. Isis, 95(3):514, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429029.

Hufbauer:2004:BRA

[2073] Karl Hufbauer. Book review: Alan J. Rocke: Nationalizing Science: Adolphe Wurtz and the Battle for French Chemistry. Isis, 95(3):514–515, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429030.

Lawrence:2004:BRL

[2074] Christopher Lawrence. Book review: Lisa Rosner: The Most Beautiful Man in Existence: The Scandalous Life of Alexander Lesassier. Isis, 95 (3):515–516, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429031.

Bailey:2004:BRP

[2075] Steve Bailey. Book review: Peter L. Rudnytsky: Reading Psychoanalysis: Freud, Rank, Ferenczi, Groddeck. Isis, 95(3):516–517, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429032.

Backhouse:2004:BRC

[2076] Roger E. Backhouse. Book review: Christopher K. Ryan: Harry Gunnison Brown: An Orthodox Economist and His Contributions. Isis, 95 (3):517–518, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429033.

Cooke:2004:BRR

[2077] Kathy J. Cooke. Book review: Rennie B. Schoepflin: Christian Science on Trial: Religious Healing in America. Isis, 95(3):518, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429034.

Cassidy:2004:BRP

[2078] James G. Cassidy. Book review: Paul Schullery; Lee Whittlesey: Myth and History in the Creation of Yellowstone National Park. Isis, 95(3): 518–519, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429035.

Wood:2004:BRS

[2079] Patricia K. Wood. Book review: Susan Schulten: The Geographical Imagination in America, 1880–1950. Isis, 95(3):519–520, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429036.

Sorlin:2004:BRB

[2080] Sverker Sörlin. Book review: Vetenskapsbärarna: Naturvetenskapen i det svenska samhället, 1880–1950 [The Carriers of Science: Science in Swedish Society, 1880–1950]. Isis, 95(3):520–522, September 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429037.

Cohn:2004:BRC

[2081] Lawrence H. Cohn. Book review: Charles F. Wooley: The Irritable Heart of Soldiers and the Origins of Anglo-American Cardiology: The U.S. Civil War (1861) to World War I (1918). Isis, 95(3):522, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429038.

Maeir:2004:BRN

[2082] Aren M. Maeir. Book review: Nadia Abu el-Haj: Facts on the Ground: Archaeological Practice and Territorial Self-Fashioning in Israeli Society. Isis, 95(3):523–524, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429039.

Fontaine:2004:BRM

[2083] Philippe Fontaine. Book review: S. M. Amadae: Rationalizing Capitalist Democracy: The Cold War Origins of Rational Choice Liberalism. Isis, 95(3):524–525, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429040.

Mifflin:2004:BRR

[2084] Jeffrey Mifflin. Book review: Roger E. Bilstein: Testing Aircraft, Exploring Space: An Illustrated History of NACA and NASA. Isis, 95 (3):525–526, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429041.

Channell:2004:BRJ

[2085] David F. Channell. Book review: John M. Charap: Explaining the Universe: The New Age of Physics. Isis, 95(3):526–527, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429042.

Cassedy:2004:BRC

[2086] James H. Cassedy. Book review: Claude E. Dolman; Richard J. Wolfe: Suppressing the Diseases of Animals and Man: Theobald Smith, Microbiologist. Isis, 95(3):527, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429043.

Guillemin:2004:BRI

[2087] Jeanne Guillemin. Book review: Igor V. Domaradskij; Wendy Orent: Biowarrior: Inside the Soviet/Russian Biological War Machine. Isis, 95 (3):527–528, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429044.

Challey:2004:BRG

[2088] James Challey. Book review: Gérard Emptoz; Valérie Marchal: Aux sources de la propriété industrielle: Guide des archives de l’INPI. Isis, 95 (3):528–529, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429045.

Good:2004:BRC

[2089] Gregory A. Good. Book review: Carl-Henry Geschwind: California Earthquakes: Science, Risk, and the Politics of Hazard Mitigation. Isis, 95(3):529–530, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429046.

Roland:2004:BRM

[2090] Alex Roland. Book review: Michael H. Gorn: Expanding the Envelope: Flight Research at NACA and NASA. Isis, 95(3):530–531, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429047.

Drouard:2004:BRG

[2091] Alain Drouard. Book review: Geneviève Heller; Gilles Jeanmonod; Jacques Gasser: Rejetées, rebelles, mal adaptées: Débats sur l’eugénisme: Pratiques de la sterilization non volontaire en Suisse romande au XXe siècle. Isis, 95(3):531–532, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429048.

Lowen:2004:BRD

[2092] Rebecca Lowen. Book review: Daniel Lee Kleinman: Impure Cultures: University Biology and the World of Commerce. Isis, 95(3):532, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/429049.

Elliott:2004:BRB

[2093] Clark A. Elliott. Book review: Portraits of Great American Scientists. Isis, 95(3):532–533, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429050.

Lunbeck:2004:BRJ

[2094] Elizabeth Lunbeck. Book review: Jonathan Michel Metzl: Prozac on the Couch: Prescribing Gender in the Era of Wonder Drugs. Isis, 95 (3):533–534, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429051.

Mark:2004:BRC

[2095] Joan Mark. Book review: Char Solomon: Tatiana Proskouriakoff: Interpreting the Ancient Maya. Isis, 95(3):534–535, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429052.

Staley:2004:BRJ

[2096] Richard Staley. Book review: John Stachel: Einstein from “B” to “Z,”. Isis, 95(3):535–536, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429053.

Canaday:2004:BRD

[2097] John Canaday. Book review: David J. Tietge: Flash Effect: Science and the Rhetorical Origins of Cold War America. Isis, 95(3):536–537, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429054.

Broome:2004:BRC

[2098] Taft H. Broome. Book review: C. G. Williams: Technology and the Dream: Reflections on the Black Experience at MIT, 1941–1999. Isis, 95(3):537, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429055.

Finocchiaro:2004:BRA

[2099] Maurice A. Finocchiaro. Book review: Armando Brissoni: Saggio su Galileo Galilei: Discorsi di dottrina. Isis, 95(3):537–538, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429056.

Stump:2004:BRA

[2100] David J. Stump. Book review: André Kukla: Social Constructivism and the Philosophy of Science. Isis, 95(3):538–539, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429057.

Barberis:2004:BRH

[2101] Daniela Barberis. Book review: Helen E. Longino: The Fate of Knowledge. Isis, 95(3):539–540, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429058.

Sismondo:2004:BRR

[2102] Sergio Sismondo. Book review: Robert K. Merton; Elinor Barber: The Travels and Adventures of Serendipity: A Study in Sociological Semantics and the Sociology of Science. Isis, 95(3):540, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429059.

Junker:2004:BRT

[2103] Thomas Junker. Book review: Thomas Potthast: Die Evolution und der Naturschutz: Zum Verhältnis von Evolutionsbiologie, Ökologie und Naturethik. Isis, 95(3):540–541, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429060.

Fuller:2004:BRS

[2104] Steve Fuller. Book review: Stephen Toulmin: Return to Reason. Isis, 95 (3):541–542, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429061.

Anonymous:2004:BRBab

[2105] Anonymous. Book review: Disease in the History of Modern Latin America: From Malaria to AIDS. Isis, 95(3):542, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/429062.

Anonymous:2004:BRBac

[2106] Anonymous. Book review: Robert Stephenson: The Eminent Engineer. Isis, 95(3):542, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429063.

Anonymous:2004:BRBad

[2107] Anonymous. Book review: The Victorian Supernatural. Isis, 95(3):542, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429064.

Anonymous:2004:BRBae

[2108] Anonymous. Book review: Useful Bodies: Humans in the Service of Medical Science in the Twentieth Century. Isis, 95(3):542–543, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429065.

Anonymous:2004:BRBaf

[2109] Anonymous. Book review: Early Modern Natural Law Theories: Contexts and Strategies in Early Enlightenment. Isis, 95(3):543, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429066.

Anonymous:2004:BRBag

[2110] Anonymous. Book review: The Challenges of Ivan Illich: A Collective Reflection. Isis, 95(3):543, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429067.

Anonymous:2004:BRBah

[2111] Anonymous. Book review: The Politics of Healing: Histories of Alternative Medicine in Twentieth–Century North America. Isis, 95(3):543–544, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/429068.

Anonymous:2004:BRBai

[2112] Anonymous. Book review: Herbert Spencer: The Intellectual Legacy. Isis, 95(3):544, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429069.

Anonymous:2004:BRBaj

[2113] Anonymous. Book review: From Newton to Hawking: A History of Cambridge University’s Lucasian Professors of Mathematics. Isis, 95 (3):544, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429070.

Anonymous:2004:BRBak

[2114] Anonymous. Book review: “Our Lords, the Sick”: McGovern Lectures in the History of Medicine and Medical Humanism. Isis, 95(3):544, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/429071.

Anonymous:2004:BRBal

[2115] Anonymous. Book review: Micrologus: Natura, scienze e società medieval, Vol. 12: Il sole e la luna. Isis, 95(3):544–545, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429072.

Anonymous:2004:BRBam

[2116] Anonymous. Book review: Die Verbindung nach Auschwitz: Biowissenschaften und Menschenversuche an Kaiser-Wilhelm-Instituten. Isis, 95(3):545, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429074.

Anonymous:2004:BRBan

[2117] Anonymous. Book review: Adolf Butenandt und die Kaiser-Wilhelm-Gesellschaft-Wissenschaft, Industrie und Politik im “Dritten Reich”, by Wolfgang Schieder; Achim Trunk. Isis, 95(3):545, September 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429075.

Anonymous:2004:TJM

[2118] Anonymous. Thomas J. Misa: Leonardo to the Internet: Technology and Culture from the Renaissance to the Present. Isis, 95(3):545, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/429073.

Anonymous:2004:BRBao

[2119] Anonymous. Book review: In Search of the Rainforest. Isis, 95(3): 545–546, September 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
429076.

Anonymous:2004:BRBap

[2120] Anonymous. Book review: Stem Cell Research: New Frontiers in Science and Ethics. Isis, 95(3):546, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/429077.

Anonymous:2004:BRBaq

[2121] Anonymous. Book review: Science and Dissent in England, 1688–1945. Isis, 95(3):546, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/429078.

Anonymous:2004:NCc

[2122] Anonymous. Notes on contributors. Isis, 95(3):547–553, September 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/429079.

Ratcliff:2004:ATS

[2123] Marc J. Ratcliff. Abraham Trembley’s strategy of generosity and the Scope of celebrity in the Mid-Eighteenth century. Isis, 95(4):555–575, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/430649.

Lykknes:2004:AAD

[2124] Annette Lykknes, Lise Kvittingen, and Anne Kristine Børresen. Appreciated abroad, depreciated at home. The career of a radiochemist in Norway: Ellen Gleditsch (1879–1968). Isis, 95(4):576–609, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/430650.

Galison:2004:I

[2125] Peter Galison. Introduction. Isis, 95(4):610–613, December 2004. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/430651.

Darrigol:2004:MEP

[2126] Olivier Darrigol. The mystery of the Einstein–Poincaré connection. Isis, 95(4):614–626, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
430652.

Jerome:2004:ERM

[2127] Fred Jerome. Einstein, race, and the myth of the cultural icon. Isis, 95 (4):627–639, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
430653.

Renn:2004:RRP

[2128] Jürgen Renn. The Relativity revolution from the perspective of historical epistemology. Isis, 95(4):640–648, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/430654.

Gleason:2004:ECE

[2129] Mary Louise Gleason and Joseph W. Dauben. Eloge: Carolyn Eisele, 1902–2000. Isis, 95(4):649–652, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/430655.

Benson:2004:EDK

[2130] Keith R. Benson and Helen Rozwadowski. Eloge: David K. van Keuren, 1950–2004. Isis, 95(4):652–653, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/430656.

Secord:2004:KT

[2131] James A. Secord. Knowledge in transit. Isis, 95(4):654–672, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/430657.

Cross:2004:BRB

[2132] Jim Cross. Book review: Essays on the History of Mechanics: In Memory of Clifford Ambrose Truesdell and Edoardo Benvenuto. Isis, 95(4):673, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432273.

Canizares-Esguerra:2004:BRBb

[2133] Jorge Cañizares-Esguerra. Book review: Abriendo las cajas negras: Colección de instrumentos científicos de la Universitat de València. Isis, 95(4):673–674, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432274.

Bergandi:2004:BRB

[2134] Donato Bergandi. Book review: Earth, Air, Fire, Water: Humanistic Studies of the Environment. Isis, 95(4):674–675, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432275.

Smith:2004:BRB

[2135] Pamela H. Smith. Book review: The Moral Authority of Nature. Isis, 95 (4):675–677, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432276.

Usselman:2004:BRJ

[2136] Mel Usselman. Book review: Joseph S. Fruton: Methods and Styles in the Development of Chemistry. Isis, 95(4):677–678, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432277.

Blustein:2004:BRB

[2137] Bonnie Ellen Blustein. Book review: Science, Race, and Ethnicity: Readings from Isis and Osiris. Isis, 95(4):678–679, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432278.

Woodward:2004:BRM

[2138] William R. Woodward. Book review: Martin Kusch: Psychological Knowledge: A Social History and Philosophy. Isis, 95(4):679–680, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432279.

Rogers:2004:BRK

[2139] Catherine Rogers. Book review: Keith J. Laidler: Energy and the Unexpected. Isis, 95(4):680–681, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432280.

Marvin:2004:BRD

[2140] Ursula B. Marvin. Book review: David Leverington: Babylon to Voyager and Beyond: A History of Planetary Astronomy. Isis, 95(4):681–682, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432281.

DeVorkin:2004:BRD

[2141] David DeVorkin. Book review: David Levy; Wendee Wallach-Levy: Cosmic Discoveries: The Wonders of Astronomy. Isis, 95(4):682–683, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432282.

deRomo:2004:BRJ

[2142] Ana Cecilia Rodríguez de Romo. Book review: José María López Piñero: La medicina en la historia. Isis, 95(4):683–684, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432283.

Phillips:2004:BRB

[2143] Denise Phillips. Book review: Beiträge zur Geschichte der Martin– Luther–Universität Halle-Wittenberg 1502–2002: Im Auftrag der Rektoratskommission für das Universitätsjubiläum. Isis, 95(4):684–685, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432284.

Scholz:2004:BRG

[2144] Albrecht Scholz. Book review: Gerhard Saalmann; Heinrich Schlange-Schöningen: Die Psoriasis: Geschichte und Therapie vom Altertum bis zur Gegenwart. Isis, 95(4):685, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432285.

Jones:2004:BRM

[2145] Karen Jones. Book review: Michael Williams: Deforesting the Earth: From Prehistory to Global Crisis. Isis, 95(4):685–686, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432286.

Bolt:2004:BRA

[2146] Marvin Bolt. Book review: Anthony Aveni: Empires of Time: Calendars, Clocks, and Cultures. Isis, 95(4):686–687, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432287.

Johnson:2004:BRH

[2147] Monte Johnson. Book review: Helen S. Lang: The Order of Nature in Aristotle’s Physics: Place and the Elements. Isis, 95(4):687–688, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432288.

Torijano:2004:BRB

[2148] Pablo A. Torijano. Book review: Revista Internacional de Investigación sobre Magia y Astrología Antiguas. Isis, 95(4):688–689, December 2004.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432289.

Higginbottom:2004:BRC

[2149] Gail Higginbottom. Book review: Clive Ruggles: Astronomy in Prehistoric Britain and Ireland. Isis, 95(4):689–690, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432290.

Martin:2004:BRL

[2150] Craig Martin. Book review: Liba Taub: Ancient Meteorology. Isis, 95 (4):690–691, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432291.

Lejbowicz:2004:BRB

[2151] Max Lejbowicz. Book review: Picatrix: Un traité de magie médiéval. Isis, 95(4):691–692, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432292.

Moran:2004:BRD

[2152] Bruce T. Moran. Book review: Douglas Biow: Doctors, Ambassadors, Secretaries: Humanism and Professions in Renaissance Italy. Isis, 95 (4):692–693, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432293.

Withers:2004:BRD

[2153] Charles W. J. Withers. Book review: David Buisseret: The Mapmakers’ Quest: Depicting New Worlds in Renaissance Europe. Isis, 95(4):693– 694, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432294.

Snedegar:2004:BRB

[2154] Keith Snedegar. Book review: A Treatise on the Astrolabe Time and the Astrolabe in The Canterbury Tales, by Geoffrey Chaucer; Sigmund Eisner; Marijane Osborn. (Science and Culture.). Isis, 95(4):694–695, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432295.

Touwaide:2004:BRM

[2155] Alain Touwaide. Book review: Minta Collins: Medieval Herbals: The Illustrative Traditions. Isis, 95(4):695–697, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432296.

Morse:2004:BRN

[2156] Victoria Morse. Book review: Natalia Lozovsky: “The Earth Is Our Book”: Geographical Knowledge in the Latin West ca. 400–1000. Isis, 95(4):697, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432297.

Sepkoski:2004:BRA

[2157] David Sepkoski. Book review: Ann E. Moyer: The Philosopher’s Game: Rithmomachia in Medieval and Renaissance Europe. Isis, 95(4):697–699, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432298.

Hu:2004:BRB

[2158] Minghui Hu. Book review: The Europeanization of Astronomical Instruments in Ming and Qing China [in Chinese]. Isis, 95(4):699–700, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432299.

Gal:2004:BRJ

[2159] Ofer Gal. Book review: Jim Bennett; Michael Cooper; Michael Hunter; Lisa Jardine: London’s Leonardo: The Life and Work of Robert Hooke. Isis, 95(4):700–701, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432300.

Steigerwald:2004:BRD

[2160] Joan Steigerwald. Book review: David Bindman: Ape to Apollo: Aesthetics and the Idea of Race in the Eighteenth Century. Isis, 95(4):701–702, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432301.

Dritsas:2004:BRJ

[2161] Lawrence S. Dritsas. Book review: Judith A. Carney: Black Rice: The African Origins of White Rice Cultivation. Isis, 95(4):702–703, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432302.

Eddy:2004:BRT

[2162] M. D. Eddy. Book review: Ted Dadswell: The Selborne Pioneer: Gilbert White as Naturalist and Scientist: A Re-examination. Isis, 95(4):703, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432303.

Jacob:2004:BRB

[2163] Margaret C. Jacob. Book review: Newton: The Making of a Genius Isaac Newton. Isis, 95(4):703–704, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432304.

Crosland:2004:BRJ

[2164] Maurice Crosland. Book review: James E. McClellan III: Specialist Control: The Publications Committee of the Académie Royale des Sciences (Paris), 1700–1793. Isis, 95(4):704–705, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432305.

Pimentel:2004:BRS

[2165] Juan Pimentel. Book review: Stuart McCook: States of Nature: Science, Agriculture, and Environment in the Spanish Caribbean, 1760–1940. Isis, 95(4):705–706, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432306.

Figueiroa:2004:BRBb

[2166] Silvia Figueirôa. Book review: Itinerários histórico-naturais. Isis, 95 (4):706–707, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432307.

Prandi:2004:BRA

[2167] Julie D. Prandi. Book review: Astrida Orle Tantillo: The Will to Create: Goethe’s Philosophy of Nature. Isis, 95(4):707–708, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432308.

Balis:2004:BRC

[2168] Andrea Balis. Book review: Caroline Jean Acker: Creating the American Junkie: Addiction Research in the Classic Era of Narcotic Control. Isis, 95(4):708, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432309.

Harper:2004:BRB

[2169] Kristine C. Harper. Book review: Brian Austin: Schonland, Scientist and Soldier: From Lightning on the Veld to Nuclear Power at Harwell: The Life of Field Marshal Montgomery’s Scientific Advisor. Isis, 95(4):709, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432310.

Gariepy:2004:BRG

[2170] Thomas P. Gariepy. Book review: Gerard N. Burrow: A History of Yale’s School of Medicine: Passing Torches to Others. Isis, 95(4):709– 710, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432311.

Carpenter:2004:BRG

[2171] Kenneth J. Carpenter. Book review: George W. Bruyn; Charles M. Poser: The History of Tropical Neurology: Nutritional Disorders. Isis, 95(4):710–711, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432312.

Harvey:2004:BRL

[2172] Joy Harvey. Book review: Linda L. Clark: The Rise of Professional Women in France: Gender and Public Administration since 1830. Isis, 95(4):711–712, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432313.

Abbri:2004:BRL

[2173] Ferdinando Abbri. Book review: Luigi Cerruti: Bella e potente: La chimica del Novecento fra scienza e società. Isis, 95(4):712–713, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432314.

Iguiniz:2004:BRB

[2174] Ruth Iguiniz. Book review: The Return of Epidemics: Health and Society in Peru during the Twentieth Century Culpa y coraje: Historia de las políticas sobre el VIH/Sida en el Perú. Isis, 95(4):713–714, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432315.

Badash:2004:BRP

[2175] Lawrence Badash. Book review: Per F. Dahl: From Nuclear Transmutation to Nuclear Fission, 1932–1939. Isis, 95(4):714–715, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432316.

Rothenberg:2004:BRS

[2176] Marc Rothenberg. Book review: Steven J. Dick: Sky and Ocean Joined: The U.S. Naval Observatory, 1830–2000. Isis, 95(4):715–716, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432317.

Lowy:2004:BRBb

[2177] Ilana Löwy. Book review: Louis Pasteur. Isis, 95(4):716, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432318.

Gilman:2004:BRG

[2178] Sander L. Gilman. Book review: Georges Didi-Huberman: Invention of Hysteria: Charcot and the Photographic Iconography of the Salpêtrière, translated by Alisa Hartz. Isis, 95(4):716–717, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432319.

Shuttleworth:2004:BRJ

[2179] Sally Shuttleworth. Book review: Joel Peter Eigen: Unconscious Crime: Mental Absence and Criminal Responsibility in Victorian London. Isis, 95(4):717–718, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432320.

Newell:2004:BRG

[2180] Julie R. Newell. Book review: Gail Fishman: Journeys through Paradise: Pioneering Naturalists in the Southeast. Isis, 95(4):718, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432321.

Strauss:2004:BRD

[2181] Michael Strauss. Book review: Deborah Fitzgerald: Every Farm a Factory: The Industrial Ideal in American Agriculture. Isis, 95(4):719, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432322.

Lagercrantz:2004:BRB

[2182] Hugo Lagercrantz. Book review: Benedikt Föger; Klaus Taschwer: Die andere Seite des Spiegels: Konrad Lorenz und der Nationalsozialismus. Isis, 95(4):719–720, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432323.

Teichmann:2004:BRJ

[2183] Jürgen Teichmann. Book review: Jan Frercks: Die Forschungspraxis Hippolyte Fizeaus: Eine Charakterisierung ausgehend von der Replikation seines Ätherwindexperiments von 1852. Isis, 95(4):720–721, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432324.

Feldmann:2004:BRD

[2184] Hans-Uli Feldmann. Book review: David Gugerli; Daniel Speich: Topografien der Nation: Politik, kartografische Ordnung und Landschaften im 19. Jahrhundert. Isis, 95(4):721–722, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/432325.

Andrews:2004:BRR

[2185] James T. Andrews. Book review: Rolf Hellebust: Flesh to Metal: Soviet Literature and the Alchemy of Revolution. Isis, 95(4):722–723, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432326.

Weindling:2004:BRL

[2186] Paul Weindling. Book review: Lane Ryo Hirabayashi: The Politics of Fieldwork: Research in an American Concentration Camp. Isis, 95(4): 723, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432327.

Mahoney:2004:BRD

[2187] Michael S. Mahoney. Book review: Diana Hook; Jeremy Norman: Origins of Cyberspace: A Library on the History of Computing, Networking, and Telecommunications. Isis, 95(4):723–724, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432328.

Houpt:2004:BRS

[2188] Katherine Albro Houpt. Book review: Susan D. Jones: Valuing Animals: Veterinarians and Their Patients in Modern America. Isis, 95(4):724– 725, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432329.

Bettin:2004:BRM

[2189] Hartmut Bettin. Book review: Martin Kirschke: Liebigs Lehrer Karl W. G. Kastner (1783–1857): Eine Professorenkarriere in Zeiten naturwissenschaftlichen Umbruchs. Isis, 95(4):725–726, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432330.

Morus:2004:BRT

[2190] Iwan Rhys Morus. Book review: Timothy W. Kneeland; Carol A. B. Warren: Pushbutton Psychiatry: A History of Electroshock in America. Isis, 95(4):726, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432331.

Krementsov:2004:BRB

[2191] Nikolai Krementsov. Book review: U istokov akademicheskoi genetiki v Sankt-Peterburge. Isis, 95(4):726–728, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/432332.

Laszlo:2004:BRB

[2192] Pierre Laszlo. Book review: L’histoire du concept de molécule. Isis, 95(4): 728, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432333.

Parthasarathy:2004:BRB

[2193] Shobita Parthasarathy. Book review: A Darker Ribbon: Breast Cancer, Women, and Their Doctors in the Twentieth Century The Breast Cancer Wars: Hope, Fear, and the Pursuit of a Cure in Twentieth–Century America. Isis, 95(4):728–729, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432334.

Rowe:2004:BRF

[2194] David E. Rowe. Book review: Frank Linhard: Historische Elemente einer Prinzipienphysik. Isis, 95(4):730–731, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432335.

Brain:2004:BRB

[2195] Robert M. Brain. Book review: The Mind of Modernism: Medicine, Psychology, and the Cultural Arts in Europe and America, 1880–1940. Isis, 95(4):731–732, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432336.

Crozier:2004:BRF

[2196] Ivan Crozier. Book review: Frank Mort: Dangerous Sexualities: Medico-Moral Politics in England since 1830. Isis, 95(4):732–733, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432337.

Hewitt:2004:BRS

[2197] Hermi Hyacinth Hewitt. Book review: Sioban Nelson: Say Little, Do Much: Nursing, Nuns, and Hospitals in the Nineteenth Century. Isis, 95(4):733, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432338.

Siegmund-Schultze:2004:BRP

[2198] Reinhard Siegmund-Schultze. Book review: Peter Pesic: Abel’s Proof: An Essay on the Sources and Meaning of Mathematical Unsolvability. Isis, 95(4):733–734, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432339.

Levine:2004:BRT

[2199] George Levine. Book review: Theodore Porter: Karl Pearson: The Scientific Life in a Statistical Age. Isis, 95(4):734–736, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432340.

Williams:2004:BRD

[2200] Richard S. Williams, Jr. Book review: Dick Ringler: Bard of Iceland: Jónas Hallgrímsson, Poet and Scientist. Isis, 95(4):736, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432341.

Tobies:2004:BRS

[2201] Renate Tobies. Book review: Sandra L. Singer: Adventures Abroad: North American Women at German-Speaking Universities, 1868–1915.

Isis, 95(4):736–737, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432342.

Chimisso:2004:BRJ

[2202] Cristina Chimisso. Book review: Jean-François Stoffel: Bibliographie d’Alexandre Koyré. Isis, 95(4):737–738, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/432343.

Pearl:2004:BRJ

[2203] Sharrona Pearl. Book review: John van Wyhe: Phrenology and the Origins of Victorian Scientific Naturalism. Isis, 95(4):738–739, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432344.

Gerovitch:2004:BRA

[2204] Slava Gerovitch. Book review: Alexander Vucinich: Einstein and Soviet Ideology. Isis, 95(4):739–740, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432345.

Richmond:2004:BRH

[2205] Marsha L. Richmond. Book review: Heiko Weber: Monistische und antimonistische Weltanschauung: Eine Auswahlbibliographie. Isis, 95 (4):740–741, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432346.

Henderson:2004:BRD

[2206] C. Michael B. Henderson. Book review: Davis A. Young: N. L. Bowen and Crystallization-Differentiation: The Evolution of a Theory. Isis, 95 (4):741–742, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432347.

Weldon:2004:BRS

[2207] Stephen P. Weldon. Book review: Stephen M. Barr: Modern Physics and Ancient Faith. Isis, 95(4):742–743, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/432348.

Anker:2004:BRM

[2208] Peder Anker. Book review: Michel Bess: The Light-Green Society: Ecology and Technological Modernity in France, 1960–2000. Isis, 95(4):743, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432349.

Dawson:2004:BRB

[2209] Virginia P. Dawson. Book review: The Century of Space Science. Isis, 95 (4):743–744, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432350.

Flesher:2004:BRR

[2210] Mary Mosher Flesher. Book review: Richard Fardon: Mary Douglas: An Intellectual Biography. Isis, 95(4):744–745, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432351.

Woodward:2004:BRB

[2211] William R. Woodward. Book review: Wissenschaft und Wiedervereinigung: Disziplinen im Umbruch. Isis, 95(4):745–746, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432352.

Dasher:2004:BRD

[2212] Douglas Dasher. Book review: Dean W. Kohlhoff: Amchitka and the Bomb: Nuclear Testing in Alaska. Isis, 95(4):746–747, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432353.

Swan:2004:BRB

[2213] Patricia B. Swan. Book review: Collected Writings Brain, Nerves, Muscles, and Electricity: My Life in Science. Isis, 95(4):747–748, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432354.

Muraskin:2004:BRC

[2214] William Muraskin. Book review: Cynthia A. Needham; Richard Canning: Global Disease Eradication: The Race for the Last Child. Isis, 95 (4):748–749, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432355.

Bassett:2004:BRF

[2215] Ross Bassett. Book review: T. F. Peterson: Nightwork: A History of Hacks and Pranks at MIT. Isis, 95(4):749–750, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432356.

Wang:2004:BRR

[2216] Zuoyue Wang. Book review: Richard Polenberg (Editor): In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing. Isis, 95 (4):750–751, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432357.

Doel:2004:BRJ

[2217] Ronald E. Doel. Book review: Jeffrey T. Richelson: The Wizards of Langley: Inside the CIA’s Directorate of Science and Technology. Isis, 95(4):751–752, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432358.

Sarmiento:2004:BRT

[2218] Augusto Sarmiento. Book review: Thomas Schlich: Surgery, Science, and Industry: A Revolution in Fracture Care, 1950s–1990s. Isis, 95 (4):752–753, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432359.

Pickren:2004:BRB

[2219] Wade Pickren. Book review: Ben Shephard: A War of Nerves: Soldiers and Psychiatrists in the Twentieth Century. Isis, 95(4):753–754, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432360.

Cravens:2004:BRA

[2220] Hamilton Cravens. Book review: Andrew Abbott: Chaos of Disciplines. Isis, 95(4):754, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432361.

Finocchiaro:2004:BRD

[2221] Maurice A. Finocchiaro. Book review: Dušan I. Bjelić: Galileo’s Pendulum: Science, Sexuality, and the Body–Instrument Link. Isis, 95(4):754– 755, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432362.

Marti:2004:BRC

[2222] Oscar R. Martí. Book review: Cristina Chimisso: Gaston Bachelard: Critic of Science and the Imagination. Isis, 95(4):755–756, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432363.

Bowker:2004:BRS

[2223] Geoffrey C. Bowker. Book review: Steven Darian: Understanding the Language of Science. Isis, 95(4):756–757, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/432364.

Greene:2004:BRBb

[2224] Mott Greene. Book review: Prematurity in Scientific Discovery: On Resistance and Neglect. Isis, 95(4):757–758, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432365.

Houpt:2004:BRJ

[2225] Katherine Albro Houpt. Book review: Janet V. Dohner: The Encyclopedia of Historic and Endangered Livestock and Poultry Breeds. Isis, 95 (4):758–759, December 2004. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432366.

Anonymous:2004:NCd

[2226] Anonymous. Notes on contributors. Isis, 95(4):760–766, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/430659.

Anonymous:2004:IBV

[2227] Anonymous. Index to ISIS, volume 95, 2004. Isis, 95(4):767–783, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/512528.

Anonymous:2004:CV

[2228] Anonymous. Contents, volume 95. Isis, 95(4):787–788, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/512529.

Anonymous:2004:F

[2229] Anonymous. Frontispiece. Isis, 95(4):viii, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432272.

Weldon:2005:CBH

[2230] Stephen P. Weldon. Current bibliography of the history of science and its cultural influences 2005. Isis, 96(??):1–199+201–242, ???? 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3652364.

Anonymous:2005:BM

[2231] Anonymous. Back matter. Isis, 96(??):??, ???? 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3652365.

Anonymous:2005:FM

[2232] Anonymous. Front matter. Isis, 96(??):i–xxi, ???? 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/3652363.

Warwick:2005:XRE

[2233] Andrew Warwick. X-rays as evidence in German orthopaedic surgery, 1895–1900. Isis, 96(1):1–24, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491543.

Lassman:2005:GSP

[2234] Thomas C. Lassman. Government science in postwar America: Henry A. Wallace, Edward U. Condon, and the transformation of the National Bureau of Standards, 1945–1951. Isis, 96(1):25–51, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/430676.

Schiebinger:2005:I

[2235] Londa Schiebinger. Introduction. Isis, 96(1):52–55, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/430677.

Harrison:2005:SBE

[2236] Mark Harrison. Science and the British empire. Isis, 96(1):56–63, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/430678.

Canizares-Esguerra:2005:ICS

[2237] Jorge Cañizares-Esguerra. Iberian colonial science. Isis, 96(1):64–70, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/430679.

Harris:2005:JSA

[2238] Steven J. Harris. Jesuit scientific activity in the overseas missions, 1540– 1773. Isis, 96(1):71–79, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/430680.

Osborne:2005:SFE

[2239] Michael A. Osborne. Science and the French empire. Isis, 96(1):80– 87, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/430681.

Feingold:2005:EFM

[2240] Mordechi Feingold. Eloge: Frank E. Manuel, 1910–2003. Isis, 96(1): 88–90, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/430682.

Hankins:2005:BRB

[2241] Thomas Hankins. Book review: Things That Talk: Object Lessons from Art and Science Models: The Third Dimension of Science. Isis, 96(1): 91–94, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/430683.

Gruber:2005:BRR

[2242] Jacob W. Gruber. Book review: Robert Carneiro: Evolutionism and Cultural Anthropology: A Critical History. Isis, 96(1):95, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432978.

Jacquart:2005:BRB

[2243] Danielle Jacquart. Book review: The Trotula: A Medieval Compendium of Women’s Medicine Born to Procreate: Women and Childbirth in France from the Middle Ages to the Eighteenth Century. Isis, 96(1): 95–97, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432979.

Cormack:2005:BRB

[2244] Lesley B. Cormack. Book review: The New Nature of Maps: Essays in the History of Cartography The Apollo’s Eye: A Cartographic Genealogy of the Earth in the Western Imagination. Isis, 96(1):97–98, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432980.

Saliba:2005:BRB

[2245] George Saliba. Book review: Arabic Astronomy in Sanskrit: Al-Birjandi on Tadhkira II, Chapter 11 and its Sanskrit Translation. Isis, 96(1):98– 99, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432981.

Weikart:2005:BRJ

[2246] Richard Weikart. Book review: Jan Sapp: Genesis: the Evolution of Biology. Isis, 96(1):99, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432982.

Parr:2005:BRJ

[2247] Joy Parr. Book review: Julie Wosk: Women and the Machine: Representations from the Spinning Wheel to the Electronic Age. Isis, 96(1): 99–100, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432983.

Shank:2005:BRG

[2248] Michael H. Shank. Book review: Geoffrey Lloyd: The Ambitions of Curiosity: Understanding the World in Ancient Greece and China. Isis, 96(1):100–101, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432984.

Riegel:2005:BRG

[2249] Jeffrey Riegel. Book review: Geoffrey Lloyd; Nathan Sivin: The Way and the Word: Science and Medicine in Early China and Greece. Isis, 96(1):101–102, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432985.

Rocca:2005:BRT

[2250] Julius Rocca. Book review: Tullio Manzoni: Il cervello secondo Galeno. Isis, 96(1):102–103, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432986.

Sherry:2005:BRP

[2251] David Sherry. Book review: Patricia F. O’Grady: Thales of Miletus: The Beginnings of Western Science and Philosophy. Isis, 96(1):103, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/432987.

Pingree:2005:BRE

[2252] David Pingree. Book review: Enrico G. Rafaelli: L’oroscopo del mundo. Il teme di nascita del mundo e del primo uomo seconda Pastrologla zoroastriana. Isis, 96(1):104, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432988.

Harkness:2005:BRA

[2253] Deborah E. Harkness. Book review: Anthony Grafton: Cardano’s Cosmos: The World and Works of a Renaissance Astrologer. Isis, 96(1):104– 105, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/432989.

Burnett:2005:BRP

[2254] D. Graham Burnett. Book review: Peter Russel: Prince Henry “the Navigator”: A Life. Isis, 96(1):105–106, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432990.

Burnett:2005:BRK

[2255] D. Graham Burnett. Book review: Kirsten A. Seaver: Maps, Myths, and Men: The Story of the Vínland Map. Isis, 96(1):106–107, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432991.

Reeds:2005:BRR

[2256] Karen Reeds. Book review: Rebecca Bushnell: Green Desire: Imagining Early Modern English Gardens. Isis, 96(1):107–108, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432992.

Dacome:2005:BRB

[2257] Lucia Dacome. Book review: Lana Caprina: Une controverse médicale sur l’Utérus pensant à l’Université de Bologne en 1771–1772. Isis, 96 (1):108–109, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432993.

Hochadel:2005:BRA

[2258] Oliver Hochadel. Book review: Andreas Gipper: Wunderbare Wissenschaft: Literarische Strategien naturwissenschaftlicher Vulgarisierung in Frankreich. Von Cyrano de Bergerac bis zur Encyclopédie. Isis, 96 (1):109–110, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
432994.

Bray:2005:BRR

[2259] Francesca Bray. Book review: Rania Huntington: Alien Kind: Foxes and Late Imperial Chinese Narrative. Isis, 96(1):110–111, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/432995.

Bleichmar:2005:BRB

[2260] Daniela Bleichmar. Book review: Bernardino de Sahagún: First Anthropologist. Isis, 96(1):111–112, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432996.

Sylla:2005:BRK

[2261] Edith Dudley Sylla. Book review: Katherine Neal: From Discrete to Continuous: The Broadening of Number Concepts in Early Modern England. Isis, 96(1):112, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432997.

Gordin:2005:BRB

[2262] Michael D. Gordin. Book review: Istochniki po istorii astronomii Rossii XVIII v. Isis, 96(1):113, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/432998.

Padron:2005:BRB

[2263] Ricardo Padrón. Book review: El Atlas del Rey Planeta: La “Descripcion de España y de las costas y puertos de sus reinos” de Pedro Texeira (1634). Isis, 96(1):113–114, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/432999.

Canizares-Esguerra:2005:BRJ

[2264] Jorge Cañizares-Esguerra. Book review: Juan Pimentel: Testigos del mundo: Ciencia, literatura y viajes la Ilustración. Isis, 96(1):114–115, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433000.

Duchesneau:2005:BRN

[2265] François Duchesneau. Book review: Nicholas Rescher: On Leibniz. Isis, 96(1):115–116, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433001.

Lynch:2005:BRA

[2266] Katherine A. Lynch. Book review: Andrea Rusnock: Vital Accounts: Quantifying Health and Population in Eighteenth-Century England and France. Isis, 96(1):116–117, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433002.

Ciardi:2005:BRR

[2267] Marco Ciardi. Book review: Raffaella Seligardi: Lavoisier in Italia: la comunità scientifica italiana e la rivoluzione chimica. Isis, 96(1):117– 118, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/433003.

Heaman:2005:BRK

[2268] E. A. Heaman. Book review: Kevin P. Siena: Venereal Disease, Hospitals, and the Urban Poor: London’s “Foul Wards,” 1600–1800. Isis, 96(1):118–119, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433004.

Powers:2005:BRR

[2269] John C. Powers. Book review: Robert Siegfried: From Elements to Atoms: A History of Chemical Composition. Isis, 96(1):119–120, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/433005.

Cavazza:2005:BRM

[2270] Marta Cavazza. Book review: Marta Stefani: Corruzione e generazione: John T. Needham e l’origine del vivente. Isis, 96(1):120–121, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433006.

Stewart:2005:BRM

[2271] Larry Stewart. Book review: Mary Terrall: The Man Who Flattened the Earth: Maupertuis and the Sciences in the Enlightenment. Isis, 96 (1):121–122, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433007.

Chianese:2005:BRR

[2272] Robert Louis Chianese. Book review: Robert D. Tobin: Doctor’s Orders: Goethe and Enlightenment Thought. Isis, 96(1):122–123, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433008.

Roberts:2005:BRR

[2273] Lissa Roberts. Book review: Rienk Vermij: The Calvinist Copernicans: The Reception of the New Astronomy in the Dutch Republic, 1575–1750. Isis, 96(1):123–124, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433009.

Turner:2005:BRN

[2274] Steven Turner. Book review: Nicholas J. Wade: Destined for Distinguished Oblivion: The Scientific Vision of William Charles Wells (1757–1817). Isis, 96(1):124–125, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433010.

Frost:2005:BRR

[2275] Alan Frost. Book review: Roger L. Williams: French Botany in the Enlightenment: The Ill-fated Voyages of La Pérouse and His Rescuers. Isis, 96(1):125, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433011.

Richmond:2005:BRH

[2276] Marsha L. Richmond. Book review: Heng-an Chen: Die Sexualitätstheorie und “Theoretische Biologie” von Max Hartmann in der ersten Hälfte des zwanzigsten Jahrhunderts. Isis, 96(1):126, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433012.

Evans:2005:BRI

[2277] James Evans. Book review: Ileana Chinnici (Editor): La carte du ciel: Correspondence inédite conservée dans les Archives de l’Observatoire de Paris. Isis, 96(1):126–127, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433013.

Sigurdson:2005:BRB

[2278] Jon Sigurdson. Book review: Traduire, transposer, naturaliser: La formation d’une langue scientifique moderne hors des frontières de l’Europe au XIXe siècle. Isis, 96(1):127–128, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433014.

Krebs:2005:BRL

[2279] Edgardo Krebs. Book review: Luiz de Castro Faria: Another Look: A Diary of the Serra do Norte Expedition. Isis, 96(1):128–129, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433015.

White:2005:BRM

[2280] Paul White. Book review: Martin Fichman: An Elusive Victorian: The Evolution of Alfred Russel Wallace. Isis, 96(1):129–130, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433016.

Sengor:2005:BRB

[2281] A. M. Celâl Sengör. Book review: Mojsisovics on the Triassic Reefs of the Dolomites, 1879. Isis, 96(1):130–131, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433017.

Delpiano:2005:BRP

[2282] Patrizia Delpiano. Book review: Paola Govoni: Un pubblico per la scienza. La divulgazione scientifica nell’Italia in formazione. Isis, 96 (1):131–132, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433018.

Neeley:2005:BRH

[2283] Kathryn A. Neeley. Book review: Holly Henry: Virginia Woolf and the Discourse of Science: The Aesthetics of Astronomy. Isis, 96(1):132– 133, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/433019.

Arens:2005:BRH

[2284] Katherine Arens. Book review: Hans-Georg Hofer: Nervenschwäche und Krieg: Modernitätskritik und Krisenbewältigung in der österreichischen Psychiatrie (1880–1920). Isis, 96(1):133, March 2005. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/433020.

Cocks:2005:BRI

[2285] Geoffrey Cocks. Book review: Ian F. McNeely: “Medicine on a Grand Scale”: Rudolf Virchow, Liberalism, and the Public Health. Isis, 96 (1):133–134, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433021.

Noakes:2005:BRA

[2286] Richard Noakes. Book review: Alex Owen: The Place of Enchantment: British Occultism and the Culture of the Modern. Isis, 96(1):134–135, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433022.

Holmberg:2005:BRA

[2287] Gustav Holmberg. Book review: Alex Soojung-Kim Pang: Empire and the Sun: Victorian Solar Eclipse Expeditions. Isis, 96(1):135–136, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/433023.

Chapman:2005:BRH

[2288] William Chapman. Book review: H. Glenn Penny: Objects of Culture: Ethnology and Ethnographic Museums in Imperial Germany. Isis, 96 (1):136–137, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433024.

VanRiper:2005:BRB

[2289] A. Bowdoin Van Riper. Book review: Brian Regal: Henry Fairfield Osborn: Race and the Search for the Origins of Man. Isis, 96(1):137– 138, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/433025.

Robinson:2005:BRD

[2290] David K. Robinson. Book review: Daniel P. Todes: Pavlov’s Physiology Factory: Experiment, Interpretation, Laboratory Enterprise. Isis, 96(1):138–139, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433026.

Wakefield:2005:BRB

[2291] Andre Wakefield. Book review: Zwischen Aufklärung, Policey, und Verwaltung Zur Genese des Medizinalwesens, 1750–1850. Isis, 96(1):139– 140, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/433027.

Bynum:2005:BRP

[2292] W. F. Bynum. Book review: Paul C. Winther: Anglo-European Science and the Rhetoric of Empire: Malaria, Opium and British Rule in India. Isis, 96(1):140–141, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433028.

Sarkar:2005:BRJ

[2293] Sahotra Sarkar. Book review: Jon Beckwith: Making Genes, Making Waves: A Social Activist in Science. Isis, 96(1):141, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433029.

Leonard:2005:BRB

[2294] Robert Leonard. Book review: Bruce Caldwell: Hayek’s Challenge: An Intellectual Biography of F. A. Hayek. Isis, 96(1):141–142, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433030.

Handler:2005:BRS

[2295] Richard Handler. Book review: Sally Cole: Ruth Landes: A Life in Anthropology. Isis, 96(1):143, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433031.

Leslie:2005:BRH

[2296] Stuart W. Leslie. Book review: Harold E. Edgerton: Exploring the Art and Science of Stopping Time: A CD-ROM Based on the Life and Work of Harold E. Edgerton. Isis, 96(1):143–144, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/433032.

LeMahieu:2005:BRM

[2297] D. L. Le Mahieu. Book review: Mark Frankland: Radio Man: The Remarkable Rise and Fall of C. O. Stanley. Isis, 96(1):144–145, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/433033.

Harris:2005:BRG

[2298] Martha Harris. Book review: Graeme K. Hunter: Light Is a Messenger: The Life and Science of William Lawrence Bragg. Isis, 96(1):145–146, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433034.

Paul:2005:BRP

[2299] Diane B. Paul. Book review: Peter Keating; Alberto Cambrosio: Biomedical Platforms: Realigning the Normal and the Pathological in Late-Twentieth-Century Medicine. Isis, 96(1):146–147, March 2005. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433035.

Lowy:2005:BRJ

[2300] Ilana Löwy. Book review: Jane Maienschein: Whose View of Life? Embryos, Cloning, and Stem Cells. Isis, 96(1):147–148, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433036.

Pesic:2005:BRC

[2301] Peter Pesic. Book review: C. A. Meier (Editor): Atom and Archetype: The Pauli/Jung Letters, 1932–1958. Isis, 96(1):148–149, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/433037.

Ruse:2005:BRH

[2302] Michael Ruse. Book review: Henry L. Minton: Departing from Deviance: A History of Homosexual Rights and Emancipatory Science in America. Isis, 96(1):149–150, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433038.

Lovett:2005:BRN

[2303] Laura Lovett. Book review: Nancy Ordover: American Eugenics: Race, Queer Anatomy, and the Science of Nationalism. Isis, 96(1):150–151, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433039.

Gorelik:2005:BRP

[2304] Gennady Gorelik. Book review: Pavel Podvig: Russian Strategic Nuclear Forces. Isis, 96(1):151, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433040.

Kohler:2005:BRK

[2305] Robert E. Kohler. Book review: Karen Rader: Making Mice: Standardizing Animals for American Biomedical Research, 1900–1955. Isis, 96(1):151–153, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433041.

Dick:2005:BRB

[2306] Steven J. Dick. Book review: The Machine in Neptune’s Garden: Historical Perspectives on Technology and the Marine Environment. Isis, 96(1): 153, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/433042.

Auger:2005:BRJ

[2307] Jean-François Auger. Book review: John N. Vardalas: The Computer Revolution in Canada: Building National Technological Competence. Isis, 96(1):153–154, March 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
433043.

Gottwaldt:2005:BRT

[2308] Alfred Gottwaldt. Book review: Thomas Zeller: Strasse, Bahn, Panorama: Verkehrswege und Landschaftsveraenderung in Deutschland von 1930 bis 1990. Isis, 96(1):154–155, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/433044.

Weidman:2005:BRT

[2309] Nadine Weidman. Book review: Thomas C. Dalton: Becoming John Dewey: Dilemmas of a Philosopher and Naturalist. Isis, 96(1):155–156, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433045.

Porter:2005:BRG

[2310] Theodore M. Porter. Book review: Graeme J. N. Gooday: The Morals of Measurement: Accuracy, Irony, and Trust in Late Victorian Electrical Practice. Isis, 96(1):156–157, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433046.

Miller:2005:BRA

[2311] James B. Miller. Book review: Alister E. McGrath: A Scientific Theology: Volume 1 (Nature). Isis, 96(1):157–158, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/433047.

Burbidge:2005:BRE

[2312] John W. Burbidge. Book review: Emmanuel Renault: Philosophie Chimique: Hegel et la Science Dynamiste de son Temps. Isis, 96(1):158–159, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/433048.

Stump:2005:BRJ

[2313] David J. Stump. Book review: John H. Zammito: A Nice Derangement of Epistemes: Postpositivism in the Study of Science from Quine to Latour. Isis, 96(1):159–160, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/433049.

Higby:2005:BRB

[2314] Gregory J. Higby. Book review: The Pharmaceutical Industry: A Guide to Historical Records. Isis, 96(1):160, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/433050.

Saliba:2005:BRF

[2315] George Saliba. Book review: Francis Maddison; Emile Savage-Smith: Science, Tools and Magic, Volume I and II: The Nasser D. Khalili Collection of Islamic Art. Isis, 96(1):160–161, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/433051.

Anonymous:2005:NCa

[2316] Anonymous. Notes on contributors. Isis, 96(1):162–166, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/430685.

Anonymous:2005:Fa

[2317] Anonymous. Frontispiece. Isis, 96(1):viii, March 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/430675.

Colgrove:2005:SDC

[2318] James Colgrove. “Science in a democracy”: The contested status of vaccination in the progressive era and the 1920s. Isis, 96(2):167–191, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/431531.

Fyfe:2005:CWB

[2319] Aileen Fyfe. Conscientious workmen or booksellers’ hacks? The professional identities of science writers in the Mid-Nineteenth Century. Isis, 96(2):192–223, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
431532.

Kohler:2005:GV

[2320] Robert E. Kohler. A Generalist’s vision. Isis, 96(2):224–229, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/431533.

Findlen:2005:TCS

[2321] Paula Findlen. The two cultures of scholarship? Isis, 96(2):230–237, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/431534.

Shapin:2005:HCR

[2322] Steven Shapin. Hyperprofessionalism and the crisis of readership in the history of science. Isis, 96(2):238–243, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/431535.

Kaiser:2005:TGV

[2323] David Kaiser. Training and the Generalist’s vision in the history of science. Isis, 96(2):244–251, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/431536.

DeVorkin:2005:ERP

[2324] David DeVorkin. Eloge: Robert Phillip Multhauf, 1919–2004. Isis, 96(2): 252–257, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/431537.

Zabell:2005:BRB

[2325] S. L. Zabell. Book review: The History of Mathematical Tables: From Sumer to Spreadsheets. Isis, 96(2):258, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491481.

Roll-Hansen:2005:BRI

[2326] Nils Roll-Hansen. Book review: Iris Fry: The Emergence of Life on Earth: A Historical and Scientific Overview. Isis, 96(2):258–260, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491482.

McEvoy:2005:BRF

[2327] John G. McEvoy. Book review: Frederic Lawrence Holmes: Investigative Pathways: Patterns and Stages in the Careers of Experimental Scientists. Isis, 96(2):260–261, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491483.

Bowler:2005:BRB

[2328] Peter J. Bowler. Book review: When Science and Christianity Meet. Isis, 96(2):261, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491484.

Williams:2005:BRS

[2329] Elizabeth A. Williams. Book review: Sidney Ochs: A History of Nerve Functions: From Animal Spirits to Molecular Mechanisms. Isis, 96(2): 261–262, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491485.

Douglas:2005:BRM

[2330] Rob-Roy Douglas. Book review: A. M. C. Şengör: The Large-Wavelength Deformations of the Lithosphere: Materials for a History of the Evolution of Thought from the Earliest Times to Plate Tectonics. Isis, 96(2):262– 263, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491486.

Toepell:2005:BRB

[2331] Michael Toepell. Book review: Die geometrische Revolution. Isis, 96(2): 263–264, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491487.

Forgan:2005:BRD

[2332] Sophie Forgan. Book review: Dalibor Vesely: Architecture in the Age of Divided Representation: The Question of Creativity in the Shadow of Production. Isis, 96(2):264–266, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491488.

Balmer:2005:BRA

[2333] Brian Balmer. Book review: Adrienne Mayor: Greek Fire, Poison Arrows, and Scorpion Bombs: Biological and Chemical Warfare in the Ancient World. Isis, 96(2):266, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/491489.

Dickson:2005:BRJ

[2334] Keith Dickson. Book review: Julius Rocca: Galen on the Brain: Anatomical Knowledge and Physiological Speculation in the Second Century A.D. Isis, 96(2):267, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491490.

Steele:2005:BRF

[2335] J. M. Steele. Book review: Francesca Rochberg: The Heavenly Writing: Divination, Horoscopy, and Astronomy in Mesopotamian Culture. Isis, 96(2):267–268, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491491.

Robson:2005:BRC

[2336] Eleanor Robson. Book review: Corinna Rossi: Architecture and Mathematics in Ancient Egypt. Isis, 96(2):268–270, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491492.

Carusi:2005:BRB

[2337] Paola Carusi. Book review: L’art sacré et divin de la chrysopée par un anonyme. Isis, 96(2):270–271, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491493.

Beretta:2005:BRB

[2338] Marco Beretta. Book review: Arte e scienza delle acque nel Rinascimento. Isis, 96(2):271, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/491494.

Efron:2005:BRB

[2339] Noah J. Efron. Book review: The Wars of the Lord. Isis, 96(2):271– 273, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491495.

Plofker:2005:BRB

[2340] Kim Plofker. Book review: The Suryapraka’sa of Suryadasa. Isis, 96 (2):273, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491496.

Howard:2005:BRB

[2341] Nicole Howard. Book review: “Devant le Deluge” and Other Essays on Early Modern Scientific Communication Science, Reading, and Renaissance Literature: The Art of Making Knowledge, 1580–1670. Isis, 96(2): 273–275, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491497.

Rubies:2005:BRB

[2342] Joan-Pau Rubiés. Book review: Merchants and Marvels: Commerce and the Representation of Nature in Early Modern Europe. Isis, 96(2):275– 276, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491498.

Cadden:2005:BRM

[2343] Joan Cadden. Book review: Maaike van der Lugt: Le ver, le démon et la Vierge: Les théories médiévales de la génération extraordinaire: Une étude sur les rapports entre théologie, philosophie naturelle et medicine. Isis, 96(2):276–277, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491499.

Jacobs:2005:BRJ

[2344] Frederika H. Jacobs. Book review: Jessica Wolfe: Humanism, Machinery, and Renaissance Literature. Isis, 96(2):277–278, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491500.

Kollerstrom:2005:BRD

[2345] Nick Kollerstrom. Book review: David Clark; Stephen P. H. Clark: Newton’s Tyranny: The Suppressed Scientific Discoveries of Stephen Gray and John Flamsteed. Isis, 96(2):278–279, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491501.

Bedini:2005:BRE

[2346] Silvio A. Bedini. Book review: Edwin Danson: Drawing the Line: How Mason and Dixon Surveyed the Most Famous Border in America. Isis, 96 (2):279, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491502.

Muller-Sievers:2005:BRS

[2347] Helmut Müller-Sievers. Book review: Simone De Angelis: Von Newton zu Haller: Studien zum Naturbegriff zwischen Empirismus und deduktiver Methode in der Schweizer Frühaufklärung. Isis, 96(2):280, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/491503.

Jones:2005:BRT

[2348] Carol Jones. Book review: Tal Golan: Laws of Men and Laws of Nature: The History of Scientific Expert Testimony in England and America. Isis, 96(2):280–281, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491504.

Debus:2005:BRM

[2349] Allen G. Debus. Book review: Mi Gyung Kim: Affinity, That Elusive Dream: A Genealogy of the Chemical Revolution. Isis, 96(2):281–282, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/491505.

Buisseret:2005:BRJ

[2350] David Buisseret. Book review: Janis Langins: Conserving the Enlightenment: French Military Engineering from Vauban to the Revolution. Isis, 96(2):282–283, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491506.

Cutter:2005:BRF

[2351] Donald C. Cutter. Book review: Fernando Monge: En la costa de la niebla: El paisaje y el discurso etnográfico ilustrado de la expedición Malaspina en el Pacífico. Isis, 96(2):283–284, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/491507.

Bertucci:2005:BRM

[2352] Paola Bertucci. Book review: Marco Piccolino; Marco Bresadola: Rane, torpedini e scintille: Galvani, Volta e l’elettricità animale. Isis, 96(2):

284–285, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491508.

Adler:2005:BRB

[2353] Kraig Adler. Book review: Profiles of Rafinesque. Isis, 96(2):285–286, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/491509.

Zupko:2005:BRB

[2354] Ronald Edward Zupko. Book review: Weights and Measures in Scotland: A European Perspective. Isis, 96(2):286–287, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491510.

Ogilvie:2005:BRM

[2355] Marilyn Bailey Ogilvie. Book review: Mary R. S. Creese: Ladies in the Laboratory II: West European Women in Science, 1800–1900: A Survey of Their Contributions to Research. Isis, 96(2):287–288, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491511.

Marsden:2005:BRB

[2356] Ben Marsden. Book review: Technology, Industrial Conflict, and the Development of Technical Education in Nineteenth-Century England. Isis, 96(2):288–289, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491512.

Rocke:2005:BRM

[2357] A. J. Rocke. Book review: Michael D. Gordin: A Well-Ordered Thing: Dmitrii Mendeleev and the Shadow of the Periodic Table. Isis, 96(2): 289–291, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491513.

Harvey:2005:BRJ

[2358] Joy Harvey. Book review: Jennifer Michael Hecht: The End of the Soul: Scientific Modernity, Atheism, and Anthropology in France. Isis, 96(2): 291, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491514.

Ramos:2005:BRD

[2359] Sérgio de Paula Ramos. Book review: Darian Leader: Freud’s Footnotes. Isis, 96(2):292, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491515.

Sauer:2005:BRB

[2360] Bev Sauer. Book review: Bernadette Longo: Spurious Coin: A History of Science, Management, and Technical Writing. Isis, 96(2):292–293, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491516.

Figueiroa:2005:BRD

[2361] Silvia Figueirôa. Book review: David West: Fritz Müller: A Naturalist in Brazil. Isis, 96(2):293–294, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/491517.

Ebeling:2005:BRE

[2362] Werner Ebeling. Book review: Eri Yagi: A Historical Approach to Entropy: Collected Papers of Eri Yagi and Her Coworkers. Isis, 96(2):294, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/491518.

Salvini:2005:BRB

[2363] Giorgio Salvini. Book review: Da via Panisperna all’America: I fisici italiani e la seconda guerra mondial. Isis, 96(2):295, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491519.

Bergandi:2005:BRR

[2364] Donato Bergandi. Book review: Richard N. L. Andrews: Managing the Environment, Managing Ourselves: A History of American Environmental Policy. Isis, 96(2):295–296, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491520.

Rojecki:2005:BRJ

[2365] Andrew Rojecki. Book review: John Canaday: The Nuclear Muse: Literature, Physics, and the First Atomic Bomb. Isis, 96(2):296–297, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491521.

Roque:2005:BRB

[2366] Xavier Roqué. Book review: La física en España a través de los Anales de la Sociedad Española de Física y Química, 1903–1965. Isis, 96(2):297– 298, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491522.

Matchett:2005:BRJ

[2367] Karin Matchett. Book review: Joseph Cotter: Troubled Harvest: Agronomy and Revolution in Mexico, 1880–2002. Isis, 96(2):298–299, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491523.

Weitze:2005:BRB

[2368] Marc-Denis Weitze. Book review: Geschichte des Deutschen Museums: Akteure, Artefakte, Ausstellungen. Isis, 96(2):299–300, June 2005. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/491524.

Smith:2005:BRM

[2369] John K. Smith. Book review: Margaret B. W. Graham; Alec T. Shuldiner: Corning and the Craft of Innovation. Isis, 96(2):300–301, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/491525.

Harwood:2005:BRS

[2370] Jonathan Harwood. Book review: Sarah Jansen: “Schädlinge”: Geschichte eines wissenschaftlichen und politischen Konstrukts, 1840– 1920. Isis, 96(2):301–302, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/491526.

Kimmelman:2005:BRR

[2371] Barbara Kimmelman. Book review: Ronald R. Kline: Consumers in the Country: Technology and Social Change in Rural America. Isis, 96(2): 302–303, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491527.

Hopwood:2005:BRC

[2372] Nick Hopwood. Book review: Cay-Rüdiger Prüll: Medizin am Toten oder am Lebenden? Pathologie in Berlin und in London, 1900–1945. Isis, 96(2):303–304, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491528.

Palmer:2005:BRP

[2373] Bill Palmer. Book review: Peter J. Ramberg: Chemical Structure, Spatial Arrangement: The Early History of Stereochemistry, 1874–1914. Isis, 96 (2):304, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491529.

Markowitz:2005:BRW

[2374] Gerald Markowitz. Book review: William G. Rothstein: Public Health and the Risk Factor: A History of an Uneven Medical Revolution. Isis, 96 (2):305, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/491530.

Schmalzer:2005:BRL

[2375] Sigrid Schmalzer. Book review: Laurence Schneider: Biology and Revolution in Twentieth-Century China. Isis, 96(2):305–306, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/491531.

Rowe:2005:BRS

[2376] David E. Rowe. Book review: Sanford L. Segal: Mathematicians under the Nazis. Isis, 96(2):306–308, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491532.

Sutphen:2005:BRN

[2377] Molly P. Sutphen. Book review: Nayan Shah: Contagious Divides: Epidemics and Race in San Francisco’s Chinatown. Isis, 96(2):308, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491533.

DeVorkin:2005:BRA

[2378] David DeVorkin. Book review: Asif A. Siddiqi: Challenge to Apollo: The Soviet Union and the Space Race, 1945–1974. Isis, 96(2):308–309, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491534.

Marti:2005:BRJ

[2379] Oscar R. Martí. Book review: Jeffrey Sklansky: The Soul’s Economy: Market Society and Selfhood in American Thought, 1820–1920. Isis, 96(2):309–310, June 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
491535.

Smil:2005:BRD

[2380] Vaclav Smil. Book review: Dietrich Stoltzenberg: Fritz Haber: Chemist, Nobel Laureate, German, Jew. Isis, 96(2):310–312, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491536.

Couvalis:2005:BRR

[2381] George Couvalis. Book review: Robert P. Farrell: Feyerabend and Scientific Values: Tightrope-Walking Rationality. Isis, 96(2):312–313, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491537.

Zammito:2005:BRD

[2382] John H. Zammito. Book review: Duncan F. Kennedy: Rethinking Reality: Lucretius and the Textualization of Nature. Isis, 96(2):313–314, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/491538.

Parshall:2005:BRB

[2383] Karen Hunger Parshall. Book review: New Trends in the History and Philosophy of Mathematics. Isis, 96(2):314–315, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491539.

Pratt:2005:BRB

[2384] Vernon Pratt. Book review: Action and Reaction: The Life and Adventures of a Couple. Isis, 96(2):315–316, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/491540.

DeLacy:2005:BRM

[2385] Margaret DeLacy. Book review: Martin Wallen: City of Health, Fields of Disease: Revolutions in the Poetry, Medicine, and Philosophy of Romanticism. Isis, 96(2):316, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/491541.

Pyle:2005:BRJ

[2386] Cynthia M. Pyle. Book review: José María López Piñero et alia: Los animales en la ciencia y la vida humana: Illustraciones zoológicas de un milenio (siglos XI–XX). Isis, 96(2):317, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/491542.

Anonymous:2005:NCb

[2387] Anonymous. Notes on contributors. Isis, 96(2):318–322, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/431539.

Anonymous:2005:Fb

[2388] Anonymous. Frontispiece. Isis, 96(2):viii, June 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/491480.

Lightman:2005:E

[2389] Bernard Lightman. Editorial. Isis, 96(3):323, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/447744.

Kohlstedt:2005:NBS

[2390] Sally Gregory Kohlstedt. Nature, not books: Scientists and the origins of the Nature-Study movement in the 1890s. Isis, 96(3):324–352, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/447745.

Rudolph:2005:TSA

[2391] John L. Rudolph. Turning science to account: Chicago and the general science movement in secondary education, 1905–1920. Isis, 96(3): 353–389, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
447746.

Dear:2005:WHS

[2392] Peter Dear. What is the history of science the history of? Early modern roots of the ideology of modern science. Isis, 96(3):390–406, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/447747.

Creager:2005:HPR

[2393] Angels N. H. Creager and Adrian Johns. HSS 2004 program report. Isis, 96(3):407–408, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/447748.

Anonymous:2005:HPC

[2394] Anonymous. HSS 2004 Prize citations. Isis, 96(3):409, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/447749.

Kohlstedt:2005:ENR

[2395] Sally Gregory Kohlstedt. Eloge: Nathan Reingold, 1927–2004. Isis, 96 (3):410–412, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
447750.

Watson:2005:ERH

[2396] Richard A. Watson. Eloge: Richard H. Popkin, 1923–2005. Isis, 96 (3):412–415, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
447751.

Winsor:2005:EEM

[2397] Mary P. Winsor. Eloge: Ernst Mayr, 1904–2005. Isis, 96(3):415–418, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/447752.

Kohler:2005:BRB

[2398] Robert E. Kohler. Book review: Science and the City. Isis, 96(3):419, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498759.

Lindee:2005:BRB

[2399] Susan Lindee. Book review: Science, History, and Social Activism: A Tribute to Everett Mendelsohn. Isis, 96(3):420–421, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498760.

Fayter:2005:BRA

[2400] Paul Fayter. Book review: Aaron Parrett: The Translunar Narrative in the Western Tradition. Isis, 96(3):421–422, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498761.

Miraldi:2005:BRA

[2401] Elisabetta Miraldi. Book review: Andrea Ubrizsy Savoia: Rapporti italo-ungheresi nella nascita della botanica in Ungheria. Isis, 96(3):422, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498762.

Finkelberg:2005:BRD

[2402] Aryeh Finkelberg. Book review: Dirk L. Couprie; Robert Hahn; Gerard Naddaf: Anaximander in Context: New Studies in the Origins of Greek Philosophy. Isis, 96(3):422–424, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498763.

Ernst:2005:BRB

[2403] Germana Ernst. Book review: Cardano e la tradizione dei saperi. Isis, 96 (3):424–425, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498764.

Kremer:2005:BRJ

[2404] Richard L. Kremer. Book review: José Chabás; Bernard R. Goldstein: The Alfonsine Tables of Toledo. Isis, 96(3):425–426, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498765.

Alfonso-Goldfarb:2005:BRB

[2405] Ana Maria Alfonso-Goldfarb. Book review: Alchimia e medicina nel medioevo. Isis, 96(3):426–427, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498766.

Gingerich:2005:BRB

[2406] Owen Gingerich. Book review: Texte zur Aufnahme der Copernicanischen Theorie. Isis, 96(3):427, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498767.

Anzulewicz:2005:BRJ

[2407] Henryk Anzulewicz. Book review: Jörn Müller: Natürliche Moral und philosophische Ethik bei Albertus Magnus. Isis, 96(3):428–429, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/498768.

Sylla:2005:BRB

[2408] Edith Dudley Sylla. Book review: Une arithmétique commerciale du XVe siècle: Le “Compendy de la praticque des nombres” de Barthélemy de Romans. Isis, 96(3):429–430, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498769.

Reiss:2005:BRA

[2409] Timothy J. Reiss. Book review: Andrew Barnaby; Lisa J. Schnell: Literate Experience: The Work of Knowing in Seventeenth-Century English Writing. Isis, 96(3):430–431, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498770.

Knobloch:2005:BRH

[2410] Eberhard Knobloch. Book review: Henk J. M. Bos: Redefining Geometrical Exactness: Descartes’ Transformation of the Early Modern Concept of Construction. Isis, 96(3):431–432, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498771.

Meli:2005:BRM

[2411] Domenico Bertoloni Meli. Book review: Michele Camerota: Galileo Galilei e la cultura scientifica nell’età della Controriforma. Isis, 96(3): 432–433, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498772.

Johnson:2005:BRF

[2412] Nuala C. Johnson. Book review: Fa-ti Fan: British Naturalists in Qing China: Science, Empire, and Cultural Encounter. Isis, 96(3):433–434, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498773.

Henry:2005:BRB

[2413] John Henry. Book review: Newton and Newtonianism: New Studies. Isis, 96(3):434–435, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498774.

Warner:2005:BRM

[2414] Deborah Warner. Book review: Marian Fournier: Early Microscopes: A Descriptive Catalogue. Isis, 96(3):435–436, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498775.

Nauenberg:2005:BRO

[2415] Michael Nauenberg. Book review: Ofer Gal: Meanest Foundations and Nobler Superstructures: Hooke, Newton, and the “Compounding of the Celestiall Motions of the Planets”. Isis, 96(3):436, September 2005. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498776.

DesChene:2005:BRS

[2416] Dennis Des Chene. Book review: Stephen Gaukroger: Descartes’ System of Natural Philosophy. Isis, 96(3):436–437, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498777.

Smith:2005:BRN

[2417] Kurt Smith. Book review: Nicholas Kollerstrom: Newton’s Forgotten Lunar Theory: His Contribution to the Quest for Longitude. Isis, 96 (3):437–438, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498778.

Delbourgo:2005:BRM

[2418] James Delbourgo. Book review: Marco Piccolino: The Taming of the Ray: Electric Fish Research in the Enlightenment from John Walsh to Alessandro Volta. Isis, 96(3):438–439, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/498779.

Bensaude-Vincent:2005:BRR

[2419] Bernadette Bensaude-Vincent. Book review: Robert E. Schofield: The Enlightened Joseph Priestley: A Study of His Life and Work from 1773 to 1804. Isis, 96(3):439–440, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498780.

Lolck:2005:BRB

[2420] Maiken Lolck. Book review: Draft of a Physical, Geographical, and Historical Description of Icelandic Ice Mountains on the Basis of a Journey to the Most Prominent of Them in 1792–1794, with Four Maps and Eight Perspective Drawings, by Sveinn Pálsson; Willi Dansgaard. Frozen Annals: Greenland Ice Sheet Research. Isis, 96(3):440–441, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498781.

Sterling:2005:BRL

[2421] Keir B. Sterling. Book review: Leonard Warren: Constantine Samuel Rafinesque: A Voice in the American Wilderness. Isis, 96(3):441–442, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498782.

Rauch:2005:BRR

[2422] Alan Rauch. Book review: Richard Yeo: Encyclopaedic Visions: Scientific Dictionaries and Enlightenment Culture. Isis, 96(3):442–443, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498783.

Potthast:2005:BRP

[2423] Thomas Potthast. Book review: Pederf Anker: Imperial Ecology: Environmental Order in the British Empire, 1895–1945. Isis, 96(3):443–444, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498784.

Wallen:2005:BRM

[2424] Martin Wallen. Book review: Maria H. Frawley: Invalidism and Identity in Nineteenth-Century Britain. Isis, 96(3):444–445, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498785.

Appel:2005:BRB

[2425] Toby Anita Appel. Book review: Étienne Geoffroy Saint-Hilaire, 1772– 1844: A Visionary Naturalist. Isis, 96(3):445–446, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498786.

Grattan-Guinness:2005:BRB

[2426] I. Grattan-Guinness. Book review: Le système du monde: Pierre Simon Laplace: Un itinéraire dans la science. Isis, 96(3):446–447, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498787.

Crouch:2005:BRP

[2427] Tom D. Crouch. Book review: Paul R. Josephson: Industrialized Nature: Brute Force Technology and the Transformation of the Natural World. Isis, 96(3):447–448, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498788.

Lundgren:2005:BRU

[2428] Anders Lundgren. Book review: Ursula Klein: Experiments, Models, Paper Tools: Cultures of Organic Chemistry in the Nineteenth Century. Isis, 96(3):448–449, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498789.

Bergandi:2005:BRJ

[2429] Donato Bergandi. Book review: John R. McNeill: Something New under the Sun: An Environmental History of the Twentieth-Century World. Isis, 96(3):449–450, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498790.

Garaizar:2005:BRB

[2430] Isabel Garaizar. Book review: La creación de la Escuela Industrial Barcelonesa (1851), by Guillermo Lusa Monforte; Guillermo Lusa Monforte; Guillermo Lusa Monforte. Inquietudes y reformas de cambio de siglo: El proyecto de nueva Escuela Industrial (1899–1910) Quaderns d’Història de l’Enginyería: 150 anys d’enginyeria industrial. Isis, 96 (3):450–451, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498791.

Atzema:2005:BRB

[2431] Eisso J. Atzema. Book review: Mathematics Unbound: The Evolution of an International Mathematical Research Community, 1800–1945. Isis, 96 (3):451–452, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498792.

Needell:2005:BRB

[2432] Allan A. Needell. Book review: Jerry Wiesner: Scientist, Statesman, Humanist: Memories and Memoirs. Isis, 96(3):452, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498793.

Buhle:2005:BRS

[2433] Mari Jo Buhle. Book review: Sonu Shamdasani: Jung and the Making of Modern Psychology: The Dream of a Science. Isis, 96(3):452–453, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498794.

Biggs:2005:BRA

[2434] Lindy Biggs. Book review: Amy Slaton: Reinforced Concrete and the Modernization of American Building, 1900–1930. Isis, 96(3):453–454, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498795.

Fichman:2005:BRR

[2435] Martin Fichman. Book review: Ross A. Slotten: The Heretic in Darwin’s Court: The Life of Alfred Russel Wallace. Isis, 96(3):454–455, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498796.

Checkovich:2005:BRN

[2436] Alex Checkovich. Book review: Neil Smith: American Empire: Roosevelt’s Geographer and the Prelude to Globalization. Isis, 96(3):455–456, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498797.

Holmes:2005:BRD

[2437] Steven Pavlos Holmes. Book review: Dennis C. Williams: God’s Wilds: John Muir’s Vision of Nature. Isis, 96(3):456–457, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498798.

Johnson:2005:BRJ

[2438] Ann Johnson. Book review: Jon Agar: The Government Machine: A Revolutionary History of the Computer. Isis, 96(3):457–458, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498799.

Jones-Imhotep:2005:BRH

[2439] Edward Jones-Imhotep. Book review: Harry Collins: Gravity’s Shadow: The Search for Gravitational Waves. Isis, 96(3):458–459, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498800.

Spinardi:2005:BRL

[2440] Graham Spinardi. Book review: Lynn Eden: Whole World on Fire: Organizations, Knowledge, and Nuclear Weapons Devastation. Isis, 96 (3):459–460, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498801.

Mighetto:2005:BRM

[2441] Lisa Mighetto. Book review: Matthew D. Evenden: Fish versus Power: An Environmental History of the Fraser River. Isis, 96(3):460–461, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498802.

Rasmussen:2005:BRA

[2442] Nicholas Rasmussen. Book review: Alison Li: J. B. Collip and the Development of Medical Research in Canada: Extracts and Enterprise. Isis, 96 (3):461–462, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498803.

Kenen:2005:BRJ

[2443] Stephanie H. Kenen. Book review: Joanne Meyerowitz: How Sex Changed: A History of Transsexuality in the United States. Isis, 96 (3):462–463, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498804.

Lopez-Ocon:2005:BRM

[2444] Leoncio López-Ocón. Book review: María Jesús Santesmases: Entre Cajal y Ochoa: Ciencias biomédicas en la España de Franco, 1939–1975. Isis, 96(3):463–464, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498805.

Freire:2005:BRK

[2445] Olival Freire, Jr. Book review: Kent W. Staley: The Evidence for the Top Quark: Objectivity and Bias in Collaborative Experimentation. Isis, 96 (3):464–465, September 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498806.

Mazumdar:2005:BRK

[2446] Pauline M. H. Mazumdar. Book review: Keith Wailoo: Dying in the City of the Blues: Sickle Cell Anemia and the Politics of Race and Health. Isis, 96(3):465–466, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498807.

Levinton:2005:BRG

[2447] Jeffrey S. Levinton. Book review: Garland E. Allen; Jeffrey Baker: Biology: Scientific Process and Social Issues. Isis, 96(3):466, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498808.

Gooday:2005:BRD

[2448] Graeme Gooday. Book review: Davis Baird: Thing Knowledge: A Philosophy of Scientific Instruments. Isis, 96(3):466–467, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498809.

Cech:2005:BRJ

[2449] Paul J. Cech. Book review: John F. Haught: Deeper Than Darwin: The Prospect for Religion in the Age of Evolution. Isis, 96(3):467–468, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498810.

Zeller:2005:BRD

[2450] Suzanne Zeller. Book review: David N. Livingstone: Putting Science in Its Place: Geographies of Scientific Knowledge. Isis, 96(3):468–469, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498811.

Rudge:2005:BRE

[2451] David Wÿss Rudge. Book review: Ernst Mayr: What Makes Biology Unique? Considerations on the Autonomy of a Scientific Discipline. Isis, 96(3):469–470, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498812.

Weirich:2005:BRB

[2452] Paul Weirich. Book review: Auguste Comte: Trajectoires positivistes 1798–1998. Isis, 96(3):470–471, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498813.

Hawkins:2005:BRW

[2453] Mike Hawkins. Book review: Warren Schmaus: Rethinking Durkheim and His Tradition. Isis, 96(3):471–472, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/498814.

Ronzitti:2005:BRB

[2454] Giuseppina Ronzitti. Book review: The Nature of Scientific Evidence: Statistical, Philosophical, and Empirical Considerations. Isis, 96(3):472, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498815.

Anonymous:2005:NCc

[2455] Anonymous. Notes on contributors. Isis, 96(3):473–476, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/447754.

Anonymous:2005:Fc

[2456] Anonymous. Frontispiece. Isis, 96(3):viii, September 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/447743.

Lane:2005:GMC

[2457] K. Maria D. Lane. Geographers of mars: Cartographic inscription and exploration narrative in late Victorian representations of the red Planet. Isis, 96(4):477–506, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498590.

McKaughan:2005:INB

[2458] Daniel J. McKaughan. The influence of niels Bohr on max Delbrück: Revisiting the hopes inspired by “light and life”. Isis, 96(4):507–529, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498591.

Staley:2005:CCC

[2459] Richard Staley. On the co-creation of classical and modern physics. Isis, 96(4):530–558, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
498592.

Alberti:2005:OM

[2460] Samuel J. M. M. Alberti. Objects and the museum. Isis, 96(4):559–571, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498593.

Forgan:2005:BMK

[2461] Sophie Forgan. Building the museum: Knowledge, conflict, and the power of place. Isis, 96(4):572–585, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/498594.

Kohlstedt:2005:TTM

[2462] Sally Gregory Kohlstedt. “Thoughts in things” modernity, history, and North American museums. Isis, 96(4):586–601, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498595.

Bennett:2005:MHS

[2463] Jim Bennett. Museums and the history of science: Practitioner’s postscript. Isis, 96(4):602–608, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498596.

Grattan-Guinness:2005:EJG

[2464] I. Grattan-Guinness and Daniel Siegel. Eloge: John L. Greenberg, 1945– 2004. Isis, 96(4):609–611, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498597.

Taylor:2005:ETM

[2465] Kenneth L. Taylor. Eloge: Thomas Malcolm Smith, Jr., 1921–2005. Isis, 96(4):612–614, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/498598.

Hamm:2005:ESS

[2466] Ernst Hamm, Alan Richardson, and Catherine Crawford. Eloge: Stephen Straker, 1942–2004. Isis, 96(4):615–617, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/498599.

Kubbinga:2005:ERT

[2467] Henk Kubbinga. Eloge: René Taton, 1915–2004. Isis, 96(4):618–620, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498600.

Sweeney:2005:E

[2468] Gerald Sweeney. To the Editor. Isis, 96(4):621–622, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498601.

Flesher:2005:R

[2469] Mary Mosher Flesher. In reply. Isis, 96(4):622, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498602.

Kubbinga:2005:E

[2470] Henk Kubbinga. To the Editor. Isis, 96(4):622–623, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498603.

Laszlo:2005:R

[2471] Pierre Laszlo. In reply. Isis, 96(4):623–624, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/498604.

Bartholomew:2005:ERO

[2472] James R. Bartholomew. Essay review: One hundred years of the Nobel Science Prizes: Elisabeth Crawford (Editor). Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. viii + 161 pp., Tokyo: Universal Academy Press, 2002. Elisabeth Crawford. The Nobel Population, 1901–1950: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry. vi + 420 pp., Tokyo: Universal Academy Press, 2002. Mauro Dardo. Nobel Laureates and Twentieth-Century Physics. x + 515 pp., Cambridge: Cambridge University Press, 2004. Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. xv + 400 pp., notes, index. New York: W. H. Freeman, 2001. István Hargittai. The Road to Stockholm: Nobel Prizes, Science, and Scientists. xvii + 342 pp., Oxford/New York: Oxford University Press, 2002. George Thomas Kurian. The Nobel Scientists: A Biographical Encyclopedia. 675 pp., Amherst, N.Y.: Prometheus Books, 2002. Isis, 96(4):625–632, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/498605;
 http://www.jstor.org/stable/3652242.

Hamlin:2005:BRB

[2473] Christopher Hamlin. Book review: Science Serialized: Representation of the Sciences in Nineteenth-Century Periodicals Culture and Science in the Nineteenth-Century Media Science in the Nineteenth-Century Periodical: Reading the Magazine of Nature. Isis, 96(4):633–642, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/498606.

Brescia:2005:BRM

[2474] Pablo Brescia. Book review: Miguel de Asúa: Ciencia y literatura: Un relato histórico. Isis, 96(4):643–644, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/501378.

Shea:2005:BRM

[2475] William R. Shea. Book review: Maurice A. Finocchiaro: Retrying Galileo, 1633–1992. Isis, 96(4):644, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/501379.

Mosley:2005:BRO

[2476] Adam Mosley. Book review: Owen Gingerich: The Book Nobody Read: Chasing the Revolutions of Nicolaus Copernicus. Isis, 96(4):644–645, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501380.

Harris:2005:BRA

[2477] Steve Harris. Book review: Augustín Udías: Searching the Heavens and the Earth: The History of Jesuit Observatories. Isis, 96(4):645–646, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501381.

Hubbard:2005:BRV

[2478] Jennifer Hubbard. Book review: Virginia M. Walsh: Global Institutions and Social Knowledge: Generating Research at the Scripps Institution and the Inter-American Tropical Tuna Commission, 1900s–1990s. Isis, 96(4):646–647, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501382.

Hankinson:2005:BRR

[2479] R. J. Hankinson. Book review: Robert Mayhew: The Female in Aristotle’s Biology: Reason or Rationalization. Isis, 96(4):647–649, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/501383.

Bertolacci:2005:BRG

[2480] Amos Bertolacci. Book review: Gotthard Strohmaier: Avicenna. Isis, 96(4):649, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501384.

Vermeir:2005:BRS

[2481] Koen Vermeir. Book review: Steven Vanden Broecke: The Limits of Influence: Pico, Louvain, and the Crisis of Renaissance Astrology. Isis, 96(4):650–651, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501385.

Wood:2005:BRB

[2482] Paul Wood. Book review: The Letters of Sir Joseph Banks: A Selection, 1768–1820. Isis, 96(4):651, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501386.

Vienne:2005:BRR

[2483] Jean-Michel Vienne. Book review: Robert Crocker: Henry more, 1614– 1687: A Biography of the Cambridge Platonist. Isis, 96(4):651–652, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501387.

Hahn:2005:BRC

[2484] Roger Hahn. Book review: Charles Coulston Gillispie: Science and Polity in France: The Revolutionary and Napoleonic Years. Isis, 96(4):652–653, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501388.

Pranghofer:2005:BRB

[2485] Sebastian Pranghofer. Book review: Anatomie: Sektionen einer medizinischen Wissenschaft im 18, by Jürgen Helm; Karin Stukenbrock. Jahrhundert. Isis, 96(4):653–654, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501389.

Marr:2005:BRR

[2486] Alexander Marr. Book review: Robert D. Huerta: Giants of Delft: Johannes Vermeer and the Natural Philosophers: The Parallel Search for Knowledge during the Age of Discovery. Isis, 96(4):654–655, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/501390.

Snobelen:2005:BRR

[2487] Stephen D. Snobelen. Book review: Rina Knoeff: Herman Boerhaave (1668–1738): Calvinist Chemist and Physician. Isis, 96(4):655–656, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501391.

Markley:2005:BRA

[2488] Robert Markley. Book review: Ayval Leshem: Newton on Mathematics and Spiritual Purity. Isis, 96(4):656–657, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501392.

Serfati:2005:BRC

[2489] Michel Serfati. Book review: Chikara Sasaki: Descartes’s Mathematical Thought. Isis, 96(4):657–658, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501393.

Daniel:2005:BRH

[2490] Dane T. Daniel. Book review: Hereward Tilton: The Quest for the Phoenix: Spiritual Alchemy and Rosicrucianism in the Work of Count Michael Maier (1569–1622). Isis, 96(4):658–659, December 2005. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501394.

Dingwall:2005:BRA

[2491] Helen Dingwall. Book review: Andrew Wear: Knowledge and Practice in English Medicine, 1550–1680. Isis, 96(4):659–661, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501395.

Evenden:2005:BRS

[2492] Matthew Evenden. Book review: Stéphane Castonguay: Protection des cultures, construction de la nature: Agriculture, foresterie et entomologie au Canada, 1884–1959. Isis, 96(4):661–662, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501396.

Anker:2005:BRB

[2493] Peder Anker. Book review: Norsk polarhistorie. Isis, 96(4):662, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501397.

Tuchman:2005:BRE

[2494] Arleen Marcia Tuchman. Book review: Eric J. Engstrom: Clinical Psychiatry in Imperial Germany: A History of Psychiatric Practice. Isis, 96 (4):662–663, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501398.

Olea-Franco:2005:BRS

[2495] Adolfo Olea-Franco. Book review: Silvia Fernanda de Mendonça Figueirôa: As ciências geológicas no Brasil: Uma história social e institucional, 1875–1934. Isis, 96(4):663–664, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501399.

Pestre:2005:BRP

[2496] Dominique Pestre. Book review: Peter Galison: Einstein’s Clocks, Poincaré’s Maps: Empires of Time. Isis, 96(4):664–665, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501400.

Morawski:2005:BRJ

[2497] Jill G. Morawski. Book review: John D. Greenwood: The Disappearance of the Social in American Social Psychology. Isis, 96(4):665–666, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501401.

Allen:2005:BRB

[2498] Garland E. Allen. Book review: Criminal Woman, the Prostitute, and the Normal Woman. Isis, 96(4):666–667, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/501402.

Rabkin:2005:BRV

[2499] Yakov M. Rabkin. Book review: Vadim Menzhulin: Drugoi Sikorskii: Neudobnye starnitsy istorii psikhiatrii. Isis, 96(4):667–668, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/501403.

Conlin:2005:BRW

[2500] Michael F. Conlin. Book review: William Tobin: The Life and Science of Léon Foucault: The Man Who Proved the Earth Rotates. Isis, 96 (4):668–669, December 2005. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501404.

Roll-Hansen:2005:BRR

[2501] Nils Roll-Hansen. Book review: Richard Weikart: From Darwin to Hitler: Evolutionary Ethics, Eugenics, and Racism in Germany. Isis, 96(4):669– 671, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501405.

Balis:2005:BRB

[2502] Andrea Balis. Book review: Shots in the Dark: The Wayward Search for an AIDS Vaccine Big Shot: Passion, Politics, and the Struggle for an AIDS Vaccine. Isis, 96(4):671–672, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501406.

Henriksen:2005:BRP

[2503] Paul Henriksen. Book review: Peter Goodchild: Edward Teller: The Real Dr. Strangelove. Isis, 96(4):672–673, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501407.

Leone:2005:BRB

[2504] Matteo Leone. Book review: Per una storia della fisica italiana, 1945– 1965. Isis, 96(4):673–674, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501408.

Klappenbach:2005:BRM

[2505] Hugo Klappenbach. Book review: Mariana Ben Plotkin: Freud in the Pampas: The Emergence and Development of a Psychoanalytic Culture in Argentina. Isis, 96(4):674–675, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501409.

Weiner:2005:BRB

[2506] Douglas R. Weiner. Book review: History of the Russian Zapovedniks, 1895–1995. Isis, 96(4):675–676, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501410.

Weidman:2005:BRS

[2507] Nadine Weidman. Book review: Steven C. Ward: Modernizing the Mind: Psychological Knowledge and the Remaking of Society. Isis, 96(4):676– 677, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501411.

Singer:2005:BRM

[2508] Bayla Singer. Book review: Margaret A. Weitekamp: Right Stuff, Wrong Sex: America’s First Women in Space Program. Isis, 96(4):677–678, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501412.

Westfall:2005:BRB

[2509] Catherine Westfall. Book review: Growing Explanations: Historical Perspectives on Recent Science. Isis, 96(4):678–679, December 2005. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501413.

Bunge:2005:BRP

[2510] Mario Bunge. Book review: Philip Kitcher: Science, Truth, and Deomocracy. Isis, 96(4):679–680, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501414.

Trenn:2005:BRJ

[2511] Thaddeus J. Trenn. Book review: John Polkinghorne: Science and the Trinity: The Christian Encounter with Reality. Isis, 96(4):680–681, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501415.

Anonymous:2005:NCd

[2512] Anonymous. Notes on contributors. Isis, 96(4):682–684, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/498608.

Anonymous:2005:IBV

[2513] Anonymous. Index to Isis, volume 96, 2005. Isis, 96(4):685–696, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501416.

Anonymous:2005:CV

[2514] Anonymous. Contents, volume 96. Isis, 96(4):699–700, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501417.

Anonymous:2005:Fd

[2515] Anonymous. Frontispiece. Isis, 96(4):vi, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/498589.

Anonymous:2006:BM

[2516] Anonymous. Back matter. Isis, 97(??):??, ???? 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/4125061.

Anonymous:2006:CBH

[2517] Anonymous. Current bibliography of the history of science and its cultural influences, 2006. Isis, 97(??):i–v + vii–xxiii + 1–252, ???? 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/4125060.

Anonymous:2006:FM

[2518] Anonymous. Front matter. Isis, 97(??):??, ???? 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/4125059.

Turchetti:2006:SNS

[2519] Simone Turchetti. “For slow neutrons, slow pay”: Enrico Fermi’s patent and the U.S. Atomic Energy Program, 1938–1953. Isis, 97(1):1–27, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501097;
 http://www.jstor.org/stable/pdfplus/10.1086/501097.pdf.

Clark:2006:HGB

[2520] J. F. M. Clark. History from the ground up: Bugs, political economy, and God in Kirby and Spence’s Introduction to Entomology (1815–1856). Isis, 97(1):28–55, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501099.

Graham:2006:CTC

[2521] Loren Graham and Jean-Michel Kantor. A comparison of two cultural approaches to mathematics: France and Russia, 1890–1930. Isis, 97(1): 56–74, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/501100.

Wise:2006:MV

[2522] M. Norton Wise. Making visible. Isis, 97(1):75–82, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/501101.

Smith:2006:ASV

[2523] Pamela H. Smith. Art, science, and visual culture in early modern Europe. Isis, 97(1):83–100, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/501102.

Morus:2006:SBS

[2524] Iwan Rhys Morus. Seeing and believing science. Isis, 97(1):101–110, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/501103.

Tucker:2006:HPA

[2525] Jennifer Tucker. The historian, the picture, and the archive. Isis, 97 (1):111–120, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501104.

Landecker:2006:MHS

[2526] Hannah Landecker. Microcinematography and the history of science and film. Isis, 97(1):121–132, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/501105.

Anonymous:2006:HPC

[2527] Anonymous. HSS 2004 Prize citations. Isis, 97(1):133–139, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501106.

Cravens:2006:BRB

[2528] Hamilton Cravens. Book review: The Modern Social Sciences. Isis, 97(1):140–143, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
501107.

Wang:2006:BRB

[2529] Jessica Wang. Book review: Blackett: Physics, War, and Politics in the Twentieth Century J, by Mary Jo Nye; David C. Cassidy. Robert Oppenheimer and the American Century. Isis, 97(1):144–147, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/501108.

vanWyhe:2006:BRM

[2530] John van Wyhe. Book review: Michael Hagner: Geniale Gehirne: Zur Geschichte der Elitegehirnforschung. Isis, 97(1):148, March 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504529.

Evans:2006:BRC

[2531] James Evans. Book review: C. M. Linton: From Eudoxus to Einstein: A History of Mathematical Astronomy. Isis, 97(1):148–149, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504530.

Canales:2006:BRA

[2532] Jimena Canales. Book review: Arnaud Maillet: The Claude Glass: Use and Meaning of the Black Mirror in Western Art. Isis, 97(1):149–150, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504531.

Wu:2006:BRP

[2533] Yi-Li Wu. Book review: Paul U. Unschuld: Huang Di nei jing su wen: Nature, Knowledge, Imagery in an Ancient Chinese Medical Text. Isis, 97(1):150–151, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504532.

Anstey:2006:BRH

[2534] Peter Anstey. Book review: Hiro Hirai: Le concept de semence dans les théories de la matière à la Renaissance: De Marsile Ficin à Pierre Gassendi. Isis, 97(1):151–152, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504533.

Hamm:2006:BRN

[2535] Ernst Hamm. Book review: Noah Heringman: Romantic Rocks, Aesthetic Geology. Isis, 97(1):152–153, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504534.

Applebaum:2006:BRM

[2536] Wilbur Applebaum. Book review: Michał Kokowski: Copernicus’s Originality: Towards Integration of Contemporary Copernican Studies. Isis, 97(1):153–154, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504535.

Zammito:2006:BRR

[2537] John H. Zammito. Book review: Raphaël Lagier: Les races humaines selon Kant. Isis, 97(1):154–155, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504536.

Walton:2006:BRW

[2538] Steven A. Walton. Book review: Wolfgang Lefèvre: Picturing Machines, 1400–1700. Isis, 97(1):155–156, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504537.

Withers:2006:BRM

[2539] Charles Withers. Book review: Mary Sponberg Pedley: The Commerce of Cartography: Making and Marketing Maps in Eighteenth-Century France and England. Isis, 97(1):156–157, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/504538.

Shteir:2006:BRL

[2540] Ann B. Shteir. Book review: Londa Schiebinger: Plants and Empire: Colonial Bioprospecting in the Atlantic World. Isis, 97(1):157–158, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504539.

Webster:2006:BRJ

[2541] Charles Webster. Book review: Jole Shackelford: A Philosophical Path for Paracelsian Medicine: The Ideas, Intellectual Context, and Influence of Petrus Severinus (1540/2–1602). Isis, 97(1):158–159, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504540.

Eamon:2006:BRP

[2542] William Eamon. Book review: Pamela H. Smith: The Body of the Artisan: Art and Experience in the Scientific Revolution. Isis, 97(1):159–161, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504541.

Roe:2006:BRC

[2543] Shirley A. Roe. Book review: Caspar Friedrich Wolff: De formatione intestinorum/La formation des intestins (1768–1769). Isis, 97(1):161, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504542.

Nyhart:2006:BRC

[2544] Lynn K. Nyhart. Book review: Carol Armstrong and Catherine de Zegher: Ocean Flowers: Impressions from Nature. Isis, 97(1):161–162, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504543.

Blackmore:2006:BRE

[2545] John T. Blackmore. Book review: Erik C. Banks: Ernst Mach’s World Elements: A Study in Natural Philosophy. Isis, 97(1):163–164, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/504544.

Goodrum:2006:BRT

[2546] Matthew R. Goodrum. Book review: Terry A. Barnhart: Ephraim George Squier and the Development of American Anthropology. Isis, 97(1):164–165, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504545.

Lustig:2006:BRA

[2547] A. J. Lustig. Book review: Augustin Pyramus de Candolle: Mémoires et souvenirs (1778–1841). Isis, 97(1):165–166, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/504546.

Haller:2006:BRC

[2548] John S. Haller, Jr. Book review: Christopher Crenner: Private Practice: In the Early Twentieth-Century Medical Office of Dr. Richard Cabot. Isis, 97(1):166–167, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504547.

Staum:2006:BRN

[2549] Martin S. Staum. Book review: Nélia Dias: La mesure des sens: Les anthropologues et le corps humain au XIXe siècle. Isis, 97(1):167–168, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504548.

Lyon:2006:BRA

[2550] Eileen L. G. Lyon. Book review: Aileen Fyfe: Science and Salvation: Evangelical Popular Science Publishing in Victorian Britain. Isis, 97(1):

169, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504549.

Westfall:2006:BRC

[2551] Catherine Westfall. Book review: C. Stewart Gillmor: Fred Terman at Stanford: Building a Discipline, a University, and Silicon Valley. Isis, 97(1):169–170, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504550.

Sommer:2006:BRB

[2552] Marianne Sommer. Book review: First in Line: Tracing Our Ape Ancestry The Metaphysics of Apes: Negotiating the Animal-Human Boundary. Isis, 97(1):170–172, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/504551.

Laubichler:2006:BRT

[2553] Manfred D. Laubichler. Book review: Thomas Junker: Die zweite Darwinsche Revolution: Geschichte des Synthetischen Darwinismus in Deutschland 1924 bis 1950. Isis, 97(1):172, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/504552.

Kroker:2006:BRH

[2554] Kenton Kroker. Book review: Horace Winchell Magoun: American Neuroscience in the Twentieth Century: Confluence of the Neural, Behavioral, and Communicative Streams. Isis, 97(1):172–174, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504553.

Bellon:2006:BRJ

[2555] Richard Bellon. Book review: Jack Meadows: The Victorian Scientist: The Growth of a Profession. Isis, 97(1):174–175, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/504554.

Ashworth:2006:BRS

[2556] William Ashworth. Book review: Steven Ruskin: John Herschel’s Cape Voyage: Private Science, Public Imagination, and the Ambitions of Empire. Isis, 97(1):175–176, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/504555.

Connor:2006:BRB

[2557] J. T. H. Connor. Book review: Creating a Tradition of Biomedical Research: Contributions to the History of the Rockefeller University The Science We Have Loved and Taught: Dartmouth Medical School’s First Two Centuries. Isis, 97(1):176–178, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504556.

Owen:2006:BRC

[2558] Alex Owen. Book review: Corinna Treitel: A Science for the Soul: Occultism and the Genesis of the German Modern. Isis, 97(1):178, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/504557.

Ensmenger:2006:BRR

[2559] Nathan Ensmenger. Book review: Ross Knox Bassett: To the Digital Age: Research Labs, Start-Up Companies, and the Rise of MOS Technology. Isis, 97(1):178–179, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/504558.

Krech:2006:BRN

[2560] Hartmut Krech. Book review: Natalie Janine Chaoui: Mit Messzirkel und Schrotflinte: Das primatologische Werk von Adolph Hans Schultz (1891–1976). Isis, 97(1):179–180, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504559.

Figueiroa:2006:BRM

[2561] Silvia Figueirôa. Book review: Marcos Chor Maio: Ciência, política e relações internacionais: Ensaios sobre Paulo Carneiro. Isis, 97(1): 181, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504560.

Herzig:2006:BRR

[2562] Rebecca Herzig. Book review: Ron Eglash, Jennifer L. Croissant, Giovanna Di Chiro and Rayvon Fouché: Appropriating Technology: Vernacular Science and Social Power. Isis, 97(1):181–182, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/504561.

Plafcan:2006:BRU

[2563] Dan Plafcan. Book review: Ulf Hashagen, Reinhard Keil-Slawik and Arthur L. Norberg: History of Computing: Software Issues. Isis, 97 (1):182–183, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504562.

Teigen:2006:BRF

[2564] Philip M. Teigen. Book review: Frank Huisman and John Harley Warner: Locating Medical History: The Stories and Their Meanings. Isis, 97 (1):183–185, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504563.

French:2006:BRD

[2565] Steven French. Book review: David Kaiser: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Isis, 97 (1):185–186, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504564.

Smith:2006:BRB

[2566] Robert W. Smith. Book review: Bettyann Holtzmann Kevles: Almost Heaven: The Story of Women in Space. Isis, 97(1):186, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504565.

Millbrooke:2006:BRH

[2567] Anne Millbrooke. Book review: Howard E. McCurdy: Faster, Better, Cheaper: Low-Cost Innovation in the U.S. Space Program. Isis, 97(1):186–188, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504566.

Gray:2006:BRA

[2568] Chris Hables Gray. Book review: Alex Roland: Strategic Computing: DARPA and the Quest for Machine Intelligence, 1983–1993. Isis, 97 (1):188–189, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504567.

Reynolds:2006:BRM

[2569] Terry S. Reynolds. Book review: Mikko Saikku: This Delta, This Land: An Environmental History of the Yazoo-Mississippi Floodplain. Isis, 97 (1):189–190, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504568.

Orland:2006:BRS

[2570] Barbara Orland. Book review: Susan Merrill Squier: Liminal Lives: Imagining the Human at the Frontiers of Biomedicine. Isis, 97(1):190– 191, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504569.

French:2006:BRA

[2571] Steven French. Book review: Andrew Watson: The Quantum Quark. Isis, 97(1):191–192, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504570.

Hull:2006:BRR

[2572] David L. Hull. Book review: Richard M. Burian: The Epistemology of Development, Evolution, and Genetics: Selected Essays. Isis, 97(1):192– 193, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504571.

Summers:2006:BRR

[2573] William C. Summers. Book review: Rowland H. Davis: The Microbial Models of Molecular Biology: From Genes to Genomes. Isis, 97(1):193– 194, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504572.

Shanks:2006:BRB

[2574] Niall Shanks. Book review: Animal Rights and Moral Philosophy Thinking with Animals: New Perspectives on Anthropomorphism. Isis, 97 (1):194–195, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504573.

Barnes:2006:BRO

[2575] Jonathan Barnes. Book review: Orna Harari: Knowledge and Demonstration: Aristotle’s Posterior Analytics. Isis, 97(1):195–196, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/504574.

Richardson:2006:BRM

[2576] Alan Richardson. Book review: Martin Jay: Songs of Experience: Modern American and European Variations on a Universal Theme. Isis, 97(1):196–197, March 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504575.

Rolin:2006:BRC

[2577] Kristina Rolin. Book review: Cassandra L. Pinnick, Noretta Koertge and Robert F. Almeder: Scrutinizing Feminist Epistemology: An Examination of Gender in Science. Isis, 97(1):197–198, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/504576.

Darden:2006:BRM

[2578] Lindley Darden. Book review: Marcel Weber: Philosophy of Experimental Biology. Isis, 97(1):198–199, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/504577.

Anonymous:2006:NCa

[2579] Anonymous. Notes on contributors. Isis, 97(1):200–203, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501110.

Anonymous:2006:Fa

[2580] Anonymous. Frontispiece. Isis, 97(1):vi, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/501098.

Reinhardt:2006:LUI

[2581] Carsten Reinhardt. A lead user of instruments in science: John D. Roberts and the adaptation of nuclear magnetic resonance to organic chemistry, 1955–1975. Isis, 97(2):205–236, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/504732.

Bromberg:2006:DPV

[2582] Joan Lisa Bromberg. Device physics vis-à-vis fundamental physics in Cold War America: The case of quantum optics. Isis, 97(2):237–259, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504733.

Hopwood:2006:PEC

[2583] Nick Hopwood. Pictures of evolution and charges of fraud: Ernst Haeckel’s embryological illustrations. Isis, 97(2):260–301, June 2006.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504734.

Richards:2006:IFL

[2584] Joan L. Richards. Introduction: Fragmented lives. Isis, 97(2):302–305, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/504735.

Terrall:2006:BCH

[2585] Mary Terrall. Biography as cultural history of science. Isis, 97(2):306– 313, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504736.

Porter:2006:LSS

[2586] Theodore M. Porter. Is the life of the scientist a scientific unit? Isis, 97(2):314–321, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504737.

Nye:2006:SBH

[2587] Mary Jo Nye. Scientific biography: History of science by another means? Isis, 97(2):322–329, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
504738.

Grant:2006:EMC

[2588] Edward Grant. Eloge: Marshall Clagett, 1916–2005. Isis, 97(2):330– 333, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/504739.

Finkelberg:2006:R

[2589] Aryeh Finkelberg. In reply. Isis, 97(2):334, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/507348.

Naddaf:2006:E

[2590] Gerard Naddaf. To the Editor. Isis, 97(2):334, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507347.

Piccolino:2006:E

[2591] Marco Piccolino. To the Editor. Isis, 97(2):334–336, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507349.

Delbourgo:2006:R

[2592] James Delbourgo. In reply. Isis, 97(2):336, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/507350.

Knight:2006:BRB

[2593] David Knight. Book review: The 1702 Chair of Chemistry at Cambridge: Transformation and Change. Isis, 97(2):337, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507351.

Morawski:2006:BRB

[2594] Jill G. Morawski. Book review: From Séance to Science: A History of the Profession of Psychology in America A Social History of Psychology. Isis, 97(2):337–339, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507352.

Myers:2006:BRF

[2595] Greg Myers. Book review: Fernand Hallyn: Les structures rhétoriques de la science: De Kepler à Maxwell. Isis, 97(2):339–340, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507353.

Lawrence:2006:BRBa

[2596] Christopher Lawrence. Book review: Body Counts: Medical Quantification in Historical and Sociological Perspective. Isis, 97(2):340–341, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507354.

Schramm:2006:BRS

[2597] Jan-Melissa Schramm. Book review: Stephen Kern: A Cultural History of Causality: Science, Murder Novels, and Systems of Thought. Isis, 97(2):341–342, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507355.

Tuana:2006:BRE

[2598] Nancy Tuana. Book review: Elisabeth A. Lloyd: The Case of the Female Orgasm: Bias in the Science of Evolution. Isis, 97(2):342–343, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507356.

Klaassen:2006:BRW

[2599] Frank Klaassen. Book review: William R. Newman: Promethean Ambitions: Alchemy and the Quest to Perfect Nature. Isis, 97(2):343–344, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/507357.

Hahm:2006:BRG

[2600] David Hahm. Book review: Gerard Naddaf: The Greek Concept of Nature. Isis, 97(2):344–345, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/507358.

McKirahan:2006:BRB

[2601] Richard McKirahan. Book review: Pythagoras: His Life, Teaching, and Influence. Isis, 97(2):345–346, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/507359.

Mayhew:2006:BRB

[2602] Robert Mayhew. Book review: Hippocrates on Ancient Medicine. Isis, 97(2):346–347, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507360.

OBoyle:2006:BRI

[2603] Cornelius O’Boyle. Book review: Ian Maclean: Logic, Signs, and Nature in the Renaissance: The Case of Learned Medicine. Isis, 97(2):347– 348, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507361.

Johnston:2006:BRE

[2604] Stephen Johnston. Book review: Eric H. Ash: Power, Knowledge, and Expertise in Elizabethan England. Isis, 97(2):348–349, June 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507362.

OMalley:2006:BRM

[2605] John W. O’Malley. Book review: Marcus Hellyer: Catholic Physics: Jesuit Natural Philosophy in Early Modern Germany. Isis, 97(2):349– 351, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507363.

Linhard:2006:BRB

[2606] Frank Linhard. Book review: The Beginnings of the Modern Philosophy of Music in England: Francis North’s A Philosophical Essay of Musick (1677) with Comments of Isaac Newton, Roger North, and in the Philosophical Transactions. Isis, 97(2):351–352, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507364.

Jensen:2006:BRB

[2607] Derek Jensen. Book review: Biographia Copernicana: Die Copernicus-Biographien des 16, by Andreas Kühne; Stefan Kirschner. bis 18. Jahrhunderts: Texte und Übersetzungen. Isis, 97(2):352, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507365.

Lagerlund:2006:BRJ

[2608] Henrik Lagerlund. Book review: Jaap Maat: Philosophical Languages in the Seventeenth Century: Delgarno, Wilkins, Leibniz. Isis, 97(2):352– 353, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507366.

Finocchiaro:2006:BRE

[2609] Maurice A. Finocchiaro. Book review: Ernan McMullin (Editor): The Church and Galileo. Isis, 97(2):353–355, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/507367.

Dettelbach:2006:BRB

[2610] Michael Dettelbach. Book review: Colonial Botany: Science, Commerce, and Politics in the Early Modern World. Isis, 97(2):355, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507368.

Spary:2006:BRP

[2611] E. C. Spary. Book review: Peter Hanns Reill: Vitalizing Nature in the Enlightenment. Isis, 97(2):355–356, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/507369.

Feldhay:2006:BRW

[2612] Rivka Feldhay. Book review: Wade Rowland: Galileo’s Mistake: A New Look at the Epic Confrontation between Galileo and the Church. Isis, 97(2):356–358, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507370.

Darnell:2006:BRS

[2613] Regna Darnell. Book review: Stephen G. Alter: William Dwight Whitney and the Science of Language. Isis, 97(2):358–359, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507371.

Bellon:2006:BRB

[2614] Richard Bellon. Book review: Bert Bender: Evolution and “the Sex Problem”: American Narratives during the Eclipse of Darwinism. Isis, 97(2):359–360, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507372.

Crist:2006:BRR

[2615] Eileen Crist. Book review: Richard W. Burkhardt, Jr.: Patterns of Behavior: Konrad Lorenz, Niko Tinbergen, and the Founding of Ethology. Isis, 97(2):360–361, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507373.

Lindemann:2006:BRF

[2616] Mary Lindemann. Book review: Fritz Dross: Krankenhaus and lokale Politik 1770–1850: Das Beispiel Düsseldorf. Isis, 97(2):361–362, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507374.

Oberdan:2006:BRA

[2617] Thomas Oberdan. Book review: Anita Burdman Feferman; Solomon Feferman: Alfred Tarski: Life and Logic. Isis, 97(2):362–363, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507375.

Richards:2006:BRJ

[2618] Joan Richards. Book review: Jeremy J. Gray: János Bolyai, Non-Euclidean Geometry, and the Nature of Space. Isis, 97(2):363–364, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507376.

Glick:2006:BRC

[2619] Thomas F. Glick. Book review: Cédric Grimoult: Histoire de l’histoire des sciences: Historiographie de l’évolutionnisme dans le monde francophone. Isis, 97(2):364, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/507377.

Chazan:2006:BRM

[2620] Michael Chazan. Book review: Marc-Antoine Kaeser: L’univers du préhistorien: Science, foi et politique dans l’œuvre et la vie d’Éduard Desor (1811–1882). Isis, 97(2):365, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/507378.

Crowe:2006:BRH

[2621] Michael J. Crowe. Book review: Helge Kragh: Matter and Spirit in the Universe: Scientific and Religious Preludes to Modern Cosmology. Isis, 97(2):365–366, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507379.

Apple:2006:BRM

[2622] Rima D. Apple. Book review: Miriam R. Levin: Defining Women’s Scientific Enterprise: Mount Holyoke Faculty and the Rise of American Science. Isis, 97(2):366–367, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/507380.

Parezo:2006:BRF

[2623] Nancy J. Parezo. Book review: Florence C. Lister: Troweling through Time: The First Century of Mesa Verdean Archaeology. Isis, 97(2): 368, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507381.

Knell:2006:BRJ

[2624] Simon J. Knell. Book review: Jack Morrell: John Phillips and the Business of Victorian Science. Isis, 97(2):368–369, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507382.

Romero:2006:BRS

[2625] Aldemaro Romero. Book review: Steven Palmer: From Popular Medicine to Medical Populism: Doctors, Healers, and Public Power in Costa Rica, 1800–1940. Isis, 97(2):369–370, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/507383.

Paul:2006:BRK

[2626] John J. Paul. Book review: Kavita Philip: Civilizing Natures: Race, Resources, and Modernity in Colonial South India. Isis, 97(2):370–371, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/507384.

Caneva:2006:BRI

[2627] Kenneth L. Caneva. Book review: Iwan Rhys Morus: When Physics Became King. Isis, 97(2):371–373, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/507385.

Pearl:2006:BRS

[2628] Sharrona Pearl. Book review: Sharon Ruston: Shelley and Vitality. Isis, 97(2):373, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507386.

Simon:2006:BRA

[2629] Linda Simon. Book review: Allen Thiher: Fiction Refracts Science: Modernist Writers from Proust to Borges. Isis, 97(2):374, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507387.

Churchill:2006:BRR

[2630] Frederick B. Churchill. Book review: Ron Amundson: The Changing Role of the Embryo in Evolutionary Thought: Roots of Evo-Devo. Isis, 97 (2):375, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507388.

Kelman:2006:BRJ

[2631] Ari Kelman. Book review: Joshua Blu Buhs: The Fire Ant Wars: Nature, Science, and Public Policy in Twentieth-Century America. Isis, 97(2):375–376, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507389.

Everson:2006:BRI

[2632] Ted Everson. Book review: Ira H. Carmen: Politics in the Laboratory: The Constitution of Human Genomics. Isis, 97(2):376–377, June 2006.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507390.

Tweedale:2006:BRA

[2633] Geoffrey Tweedale. Book review: Alfred D. Chandler, Jr.: Shaping the Industrial Century: The Remarkable Story of the Evolution of the Modern Chemical and Pharmaceutical Industries. Isis, 97(2):377–378, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507391.

Krimsky:2006:BRT

[2634] Sheldon Krimsky. Book review: Thomas R. DeGregori: Origins of the Organic Agriculture Debate. Isis, 97(2):378–379, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507392.

Kaiser:2006:BRS

[2635] David Kaiser. Book review: Sharon Ghamari-Tabrizi: The Worlds of Herman Kahn: The Intuitive Science of Thermonuclear War. Isis, 97(2): 379–382, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/507393.

Sanchez-Ron:2006:BRB

[2636] José M. Sánchez-Ron. Book review: The Scientific Legacy of Fred Hoyle. Isis, 97(2):382, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507394.

Moon:2006:BRP

[2637] Suzanne Moon. Book review: Paul R. Josephson: Resources under Regimes: Technology, Environment, and the State. Isis, 97(2):382–383, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/507395.

Barnes:2006:BRJ

[2638] Barry Barnes. Book review: Jenny Reardon: Race to the Finish: Identity and Governance in an Age of Genomics. Isis, 97(2):383–384, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/507396.

Herman:2006:BRF

[2639] Ellen Herman. Book review: F. Robert Rodman: Winnicott: Life and Work. Isis, 97(2):384–386, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/507397.

Krementsov:2006:BRN

[2640] Nikolai Krementsov. Book review: Nils Roll-Hansen: The Lysenko Effect: The Politics of Science. Isis, 97(2):386–387, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/507398.

Marks:2006:BRL

[2641] Jonathan Marks. Book review: Linda Stone; Paul F. Lurquin: A Genetic and Cultural Odyssey: The Life and Work of L. Luca Cavalli-Sforza. Isis, 97(2):387, June 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
507399.

Ruse:2006:BRP

[2642] Michael Ruse. Book review: Phil Dowe: Galileo, Darwin, and Hawking: The Interplay of Science, Reason, and Religion. Isis, 97(2):387–388, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507400.

Porter:2006:BRB

[2643] Theodore M. Porter. Book review: Oxford Dictionary of Scientific Quotations. Isis, 97(2):388–389, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/507401.

Anonymous:2006:NCb

[2644] Anonymous. Notes on contributors. Isis, 97(2):390–393, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/504741.

Anonymous:2006:Fb

[2645] Anonymous. Frontispiece. Isis, 97(2):vi, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/507346.

Coen:2006:LMF

[2646] Deborah R. Coen. A lens of many facets: Science through a Family’s eyes. Isis, 97(3):395–419, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/508074.

Crowther-Heyck:2006:PRI

[2647] Hunter Crowther-Heyck. Patrons of the revolution: Ideals and institutions in postwar behavioral science. Isis, 97(3):420–446, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/508075.

Richmond:2006:DCR

[2648] Marsha L. Richmond. The 1909 Darwin celebration: Reexamining evolution in the light of mendel, mutation, and meiosis. Isis, 97(3):447–484, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/508076.

Cadden:2006:I

[2649] Joan Cadden. Introduction. Isis, 97(3):485–486, September 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/508077.

Park:2006:WGU

[2650] Katharine Park. Women, gender, and utopia: The death of nature and the historiography of early modern science. Isis, 97(3):487–495, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/508078.

Mitman:2006:WEN

[2651] Gregg Mitman. Where ecology, nature, and politics meet: Reclaiming the death of nature. Isis, 97(3):496–504, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/508079.

Thompson:2006:BNR

[2652] Charis Thompson. Back to nature? Resurrecting ecofeminism after poststructuralist and Third-Wave feminisms. Isis, 97(3):505–512, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/508080.

Merchant:2006:SRD

[2653] Carolyn Merchant. The scientific revolution and the death of nature. Isis, 97(3):513–533, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/508090.

Blackmore:2006:E

[2654] John T. Blackmore. To the Editor. Isis, 97(3):534, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/508094.

Fan:2006:SLI

[2655] Fa ti Fan. Science in Late Imperial China, Benjamin A. Elman: On Their Own Terms: Science in China, 1550–1900. Isis, 97(3):535–538, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/508095.

Lindee:2006:BBJ

[2656] Susan Lindee. Intimate Biotechnology Joan Rothschild: The Dream of the Perfect Child Charis Thompson: Making Parents: The Ontological Choreography of Reproductive Technologies. Isis, 97(3):539–542, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/508096.

Nye:2006:BRS

[2657] David E. Nye. Book review: From Science to Industry? Flaws in the Linear Model. Karl Grandin; Nina Wormbs; Sven Widmalm (Editors): The Science-Industry Nexus: History, Policy, Implications. Isis, 97(3): 543–545, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
508097.

Malley:2006:BRB

[2658] Marjorie C. Malley. Book review: An Entertainment for Angels: Electricity in the Enlightenment, by Patricia Fara; Jeff Hughes; John Waller. (Revolutions in Science.) The Manhattan Project: Big Science and the Atom Bomb The Discovery of the Germ: Twenty Years That Transformed the Way We Think about Disease. Isis, 97(3):546–547, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509959.

Ramberg:2006:BRA

[2659] Peter Ramberg. Book review: Arthur Greenberg: The Art of Chemistry: Myths, Medicines, and Materials. Isis, 97(3):547–548, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509960.

Murphy:2006:BRG

[2660] Michelle Murphy. Book review: Gerald N. Grob: The Deadly Truth: A History of Disease in America. Isis, 97(3):548–549, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509961.

Benson:2006:BRB

[2661] Keith R. Benson. Book review: Stätten biologischer Forschung/Places of Biological Research. Isis, 97(3):549–550, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/509962.

Boschiero:2006:BRS

[2662] Luciano Boschiero. Book review: Santiago Huerta: Arcos, bóvedas y cúpulas: Geometría y equilibrio en el cálculo tradicional de estructuras de fábrica. Isis, 97(3):550–551, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509963.

Checkovich:2006:BRG

[2663] Alex Checkovich. Book review: Geoffrey J. Martin: All Possible Worlds: A History of Geographical Ideas. Isis, 97(3):551–552, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509964.

Fara:2006:BRD

[2664] Patricia Fara. Book review: David Park: The Grand Contraption: The World as Myth, Number, and Chance. Isis, 97(3):552–553, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509965.

Maxwell:2006:BRC

[2665] Keely Maxwell. Book review: Christian C. Young: The Environment and Science: Social Impact and Interaction. Isis, 97(3):553, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509966.

Debarbat:2006:BRK

[2666] Suzanne Débarbat. Book review: Ken Alder: The Measure of All Things: The Seven-Year Odyssey and Hidden Error That Transformed the World. Isis, 97(3):553–554, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/509967.

Cohen:2006:BRS

[2667] Shana Cohen. Book review: Sylvia Bowerbank: Speaking for Nature: Women and Ecologies of Early Modern England. Isis, 97(3):554–555, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509968.

Jordanova:2006:BRL

[2668] Ludmilla Jordanova. Book review: Lisa Forman Cody: Birthing the Nation: Sex, Science, and the Conception of Eighteenth-Century Britons. Isis, 97(3):555–556, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/509969.

Goodrum:2006:BRD

[2669] Matthew R. Goodrum. Book review: David Boyd Haycock: William Stukeley: Science, Religion, and Archaeology in Eighteenth-Century England. Isis, 97(3):556–557, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509970.

Hunter:2006:BRC

[2670] Michael Hunter. Book review: Claire Preston: Thomas Browne and the Writing of Early Modern Science. Isis, 97(3):557–558, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509971.

Duchesneau:2006:BRJ

[2671] François Duchesneau. Book review: Jean-Marc Rohrbasser; Jacques Véron: Leibniz et les raisonnements sur la vie humaine. Isis, 97(3): 558–559, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509972.

Iliffe:2006:BRB

[2672] Rob Iliffe. Book review: Oeconomies in the Age of Newton. Isis, 97 (3):559–561, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509973.

Jones:2006:BRW

[2673] Matthew L. Jones. Book review: William R. Shea: Designing Experiments and Games of Chance: The Unconventional Science of Blaise Pascal. Isis, 97(3):561–562, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509974.

McNeil:2006:BRB

[2674] Maureen McNeil. Book review: The Genius of Erasmus Darwin. Isis, 97 (3):562–563, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509975.

Schmaus:2006:BRC

[2675] Warren Schmaus. Book review: Claude Blanckaert: La nature de la société: Organicisme et sciences sociales au XIXe siècle. Isis, 97(3): 563–564, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509976.

Stamhuis:2006:BRM

[2676] Ida H. Stamhuis. Book review: Mineke Bosch: Aletta Jacobs 1854– 1929: Een onwrikbaar geloof in rechtvaardigheid. Isis, 97(3):564–565, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509977.

Sangwan:2006:BRP

[2677] Satpal Sangwan. Book review: Pratik Chakrabarti: Western Science in Modern India: Metropolitan Methods, Colonial Practices. Isis, 97 (3):565–566, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509978.

Agar:2006:BRH

[2678] Jon Agar. Book review: Hunter Crowther-Heyck: Herbert A. Simon: The Bounds of Reason in Modern America. Isis, 97(3):566, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509979.

Miron:2006:BRR

[2679] Janet Miron. Book review: Robert Darby: A Surgical Temptation: The Demonization of the Foreskin and the Rise of Circumcision in Britain. Isis, 97(3):567–568, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/509980.

Capelotti:2006:BRB

[2680] P. J. Capelotti. Book review: Antarctic Challenges: Historical and Current Perspectives on Otto Nordenskjöld’s Antarctic Expedition, 1901– 1903. Isis, 97(3):568–569, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509981.

Beyler:2006:BRN

[2681] Richard H. Beyler. Book review: Nancy Thorndike Greenspan: The End of the Certain World: The Life and Science of Max Born: The Nobel Physicist Who Ignited the Quantum Revolution. Isis, 97(3):569–570, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509982;
 http://www.jstor.org/stable/pdfplus/10.1086/509982.pdf.

Keating:2006:BRC

[2682] Peter Keating. Book review: Charles Hayter: An Element of Hope: Radium and the Response to Cancer in Canada, 1900–1940. Isis, 97 (3):570–571, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509983.

Cao:2006:BRD

[2683] Tian Yu Cao. Book review: Danian Hu: China and Albert Einstein: The Reception of the Physicist and His Theory in China, 1917–1979. Isis, 97 (3):571–573, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509984.

Murray:2006:BRD

[2684] David J. Murray. Book review: D. Brett King; Michael Wertheimer: Max Wertheimer and Gestalt Theory. Isis, 97(3):573–574, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509985.

Palladino:2006:BRN

[2685] Paolo Palladino. Book review: Nikolai Krementsov: International Science between the World Wars: The Case of Genetics. Isis, 97(3):574–575, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509986.

Kohler:2006:BRC

[2686] Robert E. Kohler. Book review: Christopher Lawrence: Rockefeller Money, the Laboratory, and Medicine in Edinburgh, 1919–1930: New Science in an Old Country. Isis, 97(3):575–576, September 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509987.

Rozwadowski:2006:BRD

[2687] Helen M. Rozwadowski. Book review: David Lebrun: Proteus: A Nineteenth-Century Vision. Isis, 97(3):576–577, September 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509988.

Lawrence:2006:BRBb

[2688] Susan C. Lawrence. Book review: Walking the Paris Hospitals: Diary of an Edinburgh Medical Student, 1834–1835. Isis, 97(3):577–578, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509989.

Radick:2006:BRD

[2689] Gregory Radick. Book review: Daniel Pauly: Darwin’s Fishes: An Encyclopedia of Ichthyology, Ecology, and Evolution. Isis, 97(3):578–579, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/509990.

Stine:2006:BRH

[2690] Jeffrey K. Stine. Book review: Harold L. Platt: Shock Cities: The Environmental Transformation and Reform of Manchester and Chicago. Isis, 97(3):579–581, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/509991.

Davis:2006:BRM

[2691] Edward B. Davis. Book review: Michael Ruse: The Evolution–Creation Struggle. Isis, 97(3):581–582, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509992.

Keiner:2006:BRR

[2692] Christine Keiner. Book review: Robert J. Spear: The Great Gypsy Moth War: The History of the First Campaign in Massachusetts to Eradicate the Gypsy Moth, 1890–1901. Isis, 97(3):582–583, September 2006.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509993.

Burnham:2006:BRS

[2693] John C. Burnham. Book review: Sarah W. Tracy: Alcoholism in America: From Reconstruction to Prohibition. Isis, 97(3):583–584, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509994.

Thurston:2006:BRB

[2694] Floyd E. Thurston. Book review: Labour in the Laboratory: Medical Laboratory Workers in the Maritimes, 1900–1950. Isis, 97(3):584, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509995.

Ash:2006:BRA

[2695] Mitchell G. Ash. Book review: Alexander von Schwerin: Experimentalisierung des Menschen: Der Genetiker Hans Nachtsheim und die vergleichende Erbpathologie 1920–1945. Isis, 97(3):584–586, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509996.

McCray:2006:BRD

[2696] W. Patrick McCray. Book review: David M. Berube: Nano-Hype: The Truth Behind the Nanotechnology Buzz. Isis, 97(3):586–587, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/509997.

Agnew:2006:BRI

[2697] Duncan Agnew. Book review: Irene K. Fischer: Geodesy? What’s That? My Personal Involvement in the Age-Old Quest for the Size and Shape of the Earth, with a Running Commentary on Life in a Government Research Office. Isis, 97(3):587, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509998.

Burstyn:2006:BRJ

[2698] Harold L. Burstyn. Book review: Jacob Darwin Hamblin: Oceanographers and the Cold War: Disciples of Marine Science. Isis, 97(3): 587–588, September 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
509999.

Low:2006:BRB

[2699] Morris Low. Book review: Reasonable Men, Powerful Words: Political Culture and Expertise in Twentieth–Century Japan. Isis, 97(3):588–589, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/510000.

Dawson:2006:BRW

[2700] Virginia P. Dawson. Book review: W. D. Kay: Defining NASA: The Historical Debate over the Agency’s Mission. Isis, 97(3):589–590, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/510001.

Sepkoski:2006:BRC

[2701] David Sepkoski. Book review: Chris Mooney: The Republican War on Science. Isis, 97(3):590–591, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/510002.

Seidel:2006:BRB

[2702] Robert W. Seidel. Book review: Three Mile Island: A Nuclear Crisis in Historical Perspective. Isis, 97(3):591–592, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/510003.

Gillespie:2006:BRS

[2703] Tarleton Gillespie. Book review: Steven Weber: The Success of Open Source. Isis, 97(3):592–593, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/510004.

Borgerson:2006:BRJ

[2704] Kirstin Borgerson. Book review: Jeanne Daly: Evidence-Based Medicine and the Search for a Science of Clinical Care. Isis, 97(3):593–594, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/510005.

McOuat:2006:BRJ

[2705] Gordon McOuat. Book review: Joseph LaPorte: Natural Kinds and Conceptual Change. Isis, 97(3):594–595, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/510006.

Stengers:2006:BRB

[2706] Isabelle Stengers. Book review: Science, argent et politique: Un essai d’interprétation. Isis, 97(3):595–596, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/510007.

Wall:2006:BRB

[2707] Byron E. Wall. Book review: Symmetry and Its Discontents: Essays on the History of Inductive Probability. Isis, 97(3):596–597, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/510008.

Musselman:2006:BRJ

[2708] Elizabeth Green Musselman. Book review: John P. Jackson, Jr.; Nadine M. Weidman: Race, Racism, and Science: Social Impact and Interaction. Isis, 97(3):597–598, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/510009.

Slaton:2006:BRB

[2709] Amy Slaton. Book review: Science, Technology, and Society: An Encyclopedia. Isis, 97(3):598–599, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/510010.

Anonymous:2006:NCc

[2710] Anonymous. Notes on contributors. Isis, 97(3):600–603, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/509958.

Anonymous:2006:Fc

[2711] Anonymous. Frontispiece. Isis, 97(3):vi, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/509957.

Hankins:2006:LGS

[2712] Thomas L. Hankins. A “large and graceful sinuosity” John Herschel’s graphical method. Isis, 97(4):605–633, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/509946.pdf.

McCray:2006:ASI

[2713] W. Patrick McCray. Amateur scientists, the International Geophysical Year, and the ambitions of Fred Whipple. Isis, 97(4):634–658, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/509947.pdf.

Pettit:2006:JBC

[2714] Michael Pettit. “The joy in believing”: The Cardiff Giant, commercial deceptions, and styles of observation in Gilded Age America. Isis, 97 (4):659–677, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/509948.pdf.

Alexander:2006:II

[2715] Amir R. Alexander. Introduction. Isis, 97(4):678–682, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/509949.pdf.

Terrall:2006:MNG

[2716] Mary Terrall. Mathematics in narratives of geodetic expeditions. Isis, 97 (4):683–699, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/509950.pdf.

Richards:2006:HMF

[2717] Joan L. Richards. Historical mathematics in the French Eighteenth Century. Isis, 97(4):700–713, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/509951.pdf.

Alexander:2006:TMR

[2718] Amir R. Alexander. Tragic mathematics: Romantic narratives and the refounding of mathematics in the Early Nineteenth Century. Isis, 97 (4):714–726, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/509952.pdf.

Gischler:2006:LE

[2719] Eberhard Gischler, A. M. C. Sengör, and David C. Cassidy. Letters to the Editor. Isis, 97(4):727–730, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/509953.pdf.

Crowther-Heyck:2006:BRP

[2720] Hunter Crowther-Heyck. Book review: Peter J. Bowler and Iwan Rhys Morus: Making Modern Science: A Historical Survey A History of Science in Society: From Philosophy to Utility. Isis, 97(4):731–732, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512866.pdf.

Stewart:2006:BRY

[2721] Ian G. Stewart. Book review: Yves Gingras, Peter Keating and Camille Limoges: Du scribe au savant: Les porteurs du savoir de l’Antiquité à la révolution industrielle. Isis, 97(4):732–733, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512867.pdf.

Fenster:2006:BRI

[2722] Della Fenster. Book review: Ivor Grattan-Guinness: Landmark Writings in Western Mathematics, 1640-1940. Isis, 97(4):733–734, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512868.pdf.

Doel:2006:BRD

[2723] Ronald E. Doel. Book review: David Kaiser: Pedagogy and the Practice of Science: Historical and Contemporary Perspectives. Isis, 97(4):734– 735, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512869.pdf.

Zuidervaart:2006:BRC

[2724] Huib J. Zuidervaart. Book review: Chunglin Kwa: De ontdekking van het weten: Een andere geschiedenis van de wetenschap. Isis, 97(4):736, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512870.pdf.

Hessenbruch:2006:BRJ

[2725] Arne Hessenbruch. Book review: Julius A. Stratton and Loretta H. Mannix: Mind and Hand: The Birth of MIT Pursuing the Endless Frontier: Essays on MIT and the Role of Research Universities. Isis, 97 (4):736–737, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512871.pdf.

Ogilvie:2006:BRG

[2726] Brian W. Ogilvie. Book review: Gian Battista Vai and William Cavazza: Four Centuries of the Word Geology: Ulisse Aldrovandi 1603 in Bologna. Isis, 97(4):738–739, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/512872.pdf.

Hamblin:2006:BRS

[2727] Jacob Darwin Hamblin. Book review: Steven A. Walton: Instrumental in War: Science, Research, and Instruments between Knowledge and the World. Isis, 97(4):739–740, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512873.pdf.

Cosans:2006:BRP

[2728] Christopher Cosans. Book review: Philip J. van der Eijk: Medicine and Philosophy in Classical Antiquity: Doctors and Philosophers on Nature, Soul, Health, and Disease. Isis, 97(4):740–741, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512874.pdf.

Wood:2006:BRC

[2729] Rega Wood. Book review: Cecilia Trifogli: Oxford Physics in the Thirteenth Century (ca. 1250-1270): Motion, Infinity, Place, and Time. Cecilia Trifogli: Liber Tertius Physicorum Aristotelis: Repertorio delle Questioni com menti inglesi ca. 1250–1270. Isis, 97(4):741–743, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512875.pdf.

Schechner:2006:BRG

[2730] Sara Schechner. Book review: Gerard L’E. Turner: Elizabethan Instrument Makers: The Origins of the London Trade in Precision Instrument Making. Isis, 97(4):743, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512876.pdf.

Osler:2006:BRP

[2731] Margaret J. Osler. Book review: Philip Beeley and Christoph J. Scriba (Editors): The Correspondence of John Wallis (1616-1703). Volume 1: 1641–1659. The Correspondence of John Wallis (1616–1703). Volume 2: 1660–September 1668. Isis, 97(4):743–745, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512877.pdf.

Clark:2006:BRR

[2732] William Clark. Book review: Roger Berkowitz: The Gift of Science: Leibniz and the Modern Legal Tradition. Isis, 97(4):745–746, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512878.pdf.

Dupre:2006:BRJ

[2733] Sven Dupré. Book review: Jean-Vincent Blanchard: L’optique du discours au XVIIe siècle: De la rhétorique des jésuites au style de la raison moderne (Descartes, Pascal). Isis, 97(4):746–747, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512879.pdf.

Greene:2006:BRH

[2734] Mott Greene. Book review: Hasok Chang: Inventing Temperature: Measurement and Scientific Progress. Isis, 97(4):747–748, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512880.pdf.

Ash:2006:BRR

[2735] Eric H. Ash. Book review: Ronald W. Cooley: “Full of All Knowledg”: George Gerbert’s Country Parson and Early Modern Social Discourse. Isis, 97(4):748–749, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/512881.pdf.

Endersby:2006:BRF

[2736] Jim Endersby. Book review: Felix Driver and Luciana Martens: Tropical Visions in an Age of Empire. Isis, 97(4):749, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512882.pdf.

Ducheyne:2006:BRT

[2737] Steffen Ducheyne. Book review: Thierry Gontier: Descartes et la causa sui: Autoproduction divine, autodétermination humaine. Isis, 97(4):749– 750, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512883.pdf.

Finocchiaro:2006:BRP

[2738] Maurice A. Finocchiaro. Book review: Pierre-Noël Mayaud: Le conflit entre l’astronomie nouvelle et l’Écriture sainte aux XVIe et XVIIe siècles: Un moment de l’histoire des idées: Autour de l’affaire Galilée. Isis, 97(4):750–752, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/512884.pdf.

Blay:2006:BRM

[2739] Michel Blay. Book review: Marco Panza: Newton et les origines de l’analyse: 1664–1666. Isis, 97(4):752, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/512885.pdf.

Leach:2006:BRS

[2740] William Leach. Book review: Susan Scott Parrish: American Curiosity: Cultures of Natural History in the Colonial British Atlantic World. Isis, 97(4):752–753, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512886.pdf.

Emerson:2006:BRG

[2741] Roger L. Emerson. Book review: Guenter B. Risse: New Medical Challenges during the Scottish Enlightenment. Isis, 97(4):753–754, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512887.pdf.

Alves:2006:BRR

[2742] Abel Alves. Book review: Rafael Ángel Rodríguez Sánchez: La introducción de la medicina moderna en España: Una imagen de nuestra renovación científica (1687-1727) desde la Teoría de la Ciencia de Thomas S. Kuhn. Isis, 97(4):754–755, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512888.pdf.

Rappaport:2006:BRM

[2743] Rhoda Rappaport. Book review: Martin J. S. Rudwick: Bursting the Limits of Time: The Reconstruction of Geohistory in the Age of Revolution. Isis, 97(4):755–757, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512889.pdf.

Bertomeu-Sanchez:2006:BRJ

[2744] José Ramón Bertomeu-Sánchez. Book review: Jonathan Simon: Chemistry, Pharmacy, and Revolution in France, 1777-1809. Isis, 97(4):757– 758, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512890.pdf.

Sargent:2006:BRJ

[2745] Rose-Mary Sargent. Book review: Julie Robin Solomon and Catherine Gimelli Martin: Francis Bacon and the Refiguring of Early Modern Thought: Essays to Commemorate “The Advancement of Learning” (1605-2005). Isis, 97(4):758–759, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512891.pdf.

Knott:2006:BRH

[2746] Sarah Knott. Book review: Hubert Steinke: Irritating Experiments: Haller’s Concept and the European Controversy on Irritability and Sensibility, 1750-90. Isis, 97(4):759–760, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/512892.pdf.

Nye:2006:BRJ

[2747] Robert A. Nye. Book review: Judith Zinsser: Men, Women, and the Birthing of Modern Science. Isis, 97(4):760–761, December 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512893.pdf.

Good:2006:BRK

[2748] Gregory Good. Book review: Katharine Anderson: Predicting the Weather: Victorians and the Science of Meteorology. Isis, 97(4):761–763, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512894.pdf.

Nickles:2006:BRT

[2749] Thomas Nickles. Book review: Theodore Arabatzis: Representing Electrons: A Biographical Approach to Theoretical Entities. Isis, 97(4):763– 764, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512895.pdf.

Pelis:2006:BRM

[2750] Kim Pelis. Book review: Michael Bliss: Harvey Cushing: A Life in Surgery. Isis, 97(4):764–766, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512896.pdf.

Pyenson:2006:BRA

[2751] Lewis Pyenson. Book review: Albert Einstein: The Collected Papers of Albert Einstein, edited by Michel Janssen, Robert Schulmann, József Illy, Christoph Lehner, Diana Kormos Buchwald, Daniel Kennefick, A. J. Kox, David Rowe, R. Hirschmann, O. Moses, A. Mynttinen, A. Pringle, and R. Fountain. Albert Einstein. The Collected Papers of Albert Einstein: Volume 7: The Berlin Years: Writings, 1918–1921. Isis, 97(4): 766–767, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512897.pdf.

Hunt:2006:BRR

[2752] Bruce J. Hunt. Book review: Robert Fox and Graeme Gooday: Physics in Oxford, 1839-1939: Laboratories, Learning, and College Life. Isis, 97 (4):767–768, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512898.pdf.

Coen:2006:BRW

[2753] Deborah R. Coen. Book review: Wilhelm Füßl: Oskar von Miller 1855-1934: Eine Biographie. Isis, 97(4):768–769, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512899.pdf.

Dean:2006:BRS

[2754] Dennis R. Dean. Book review: Sandra Herbert: Charles Darwin, Geologist. Isis, 97(4):769–770, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512900.pdf.

Hagen:2006:BRS

[2755] Joel B. Hagen. Book review: Sharon E. Kingsland: The Evolution of American Ecology, 1890-2000. Isis, 97(4):770, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512901.pdf.

Marsden:2006:BRH

[2756] Ben Marsden. Book review: Harro Maas: William Stanley Jevons and the Making of Modern Economics. Isis, 97(4):770–772, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512902.pdf.

Lane:2006:BRR

[2757] K. Maria D. Lane. Book review: Robert Markley: Dying Planet: Mars in Science and the Imagination. Isis, 97(4):772–773, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512903.pdf.

MacKinnon:2006:BRM

[2758] Edward MacKinnon. Book review: Michela Massimi: Pauli’s Exclusion Principle: The Origin and Validation of a Scientific Principle. Isis, 97 (4):773–774, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512904.pdf.

Roberts:2006:BRJ

[2759] David Lindsay Roberts. Book review: John Parker: R. L. Moore: Mathematician and Teacher. Isis, 97(4):774–775, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512905.pdf.

Robinson:2006:BRS

[2760] Michael F. Robinson. Book review: Sumathi Ramaswamy: The Lost Land of Lemuria: Fabulous Geographies, Catastrophic Histories. Isis, 97 (4):775–776, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512906.pdf.

Blaisdell:2006:BRH

[2761] Muriel L. Blaisdell. Book review: Helen M. Rozwadowski: Fathoming the Ocean: The Discovery and Exploration of the Deep Sea. Isis, 97 (4):776–778, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512907.pdf.

Nahin:2006:BRT

[2762] Paul J. Nahin. Book review: Tapan K. Sarkar, Robert J. Mailloux, Arthur A. Oliner, Magdalena Salazar-Palma and Dipak L. Sengupta:

History of Wireless. Isis, 97(4):778–779, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/pdfplus/10.1086/512908.pdf.

Weiss:2006:BRH

[2763] Sheila Faith Weiss. Book review: Hans-Walter Schmuhl: Grenzüberschreitungen: Das Kaiser-Wilhelm-Institut für Anthropologie, menschliche Erblehre und Eugenik 1927-1945. Isis, 97(4):779–780, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512909.pdf.

Harrington:2006:BRJ

[2764] Anne Harrington. Book review: Jay Schulkin: Curt Richter: A Life in the Laboratory. Isis, 97(4):780–781, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/512910.pdf.

Turner:2006:BRB

[2765] Frank M. Turner. Book review: Bart Schultz: Henry Sidgwick-Eye of the Universe: An Intellectual Biography. Isis, 97(4):781–782, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512911.pdf.

Smith:2006:BRH

[2766] Merritt Roe Smith. Book review: Howard P. Segal: Recasting the Machine Age: Henry Ford’s Village Industries. Isis, 97(4):782–783, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512912.pdf.

Biggs:2006:BRK

[2767] David Biggs. Book review: Kenneth E. Wilkening: Acid Rain Science and Politics in Japan: A History of Knowledge and Action toward Sustainability. Isis, 97(4):783–784, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512913.pdf.

Hooker:2006:BRD

[2768] Claire Hooker. Book review: Diana Wyndham: Eugenics in Australia: Striving for National Fitness. Isis, 97(4):784–786, December 2006. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512914.pdf.

Strenski:2006:BRM

[2769] Ivan Strenski. Book review: Michael W. Young: Malinowski: Odyssey of an Anthropologist, 1884-1920. Isis, 97(4):786–787, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512915.pdf.

Craft:2006:BRE

[2770] Stephen G. Craft. Book review: Erik M. Conway: High-Speed Dreams: NASA and the Technopolitics of Supersonic Transportation, 1945-1999. Isis, 97(4):787–788, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/512916.pdf.

Creager:2006:BRA

[2771] Angela N. H. Creager. Book review: Alan H. Goodman, Deborah Heath and M. Susan Lindee: Genetic Nature/Culture: Anthropology and Science beyond the Two-Culture Divide. Isis, 97(4):788–790, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512917.pdf.

Gordin:2006:BRG

[2772] Michael D. Gordin. Book review: Gennady Gorelik: The World of Andrei Sakharov: A Russian Physicist’s Path to Freedom. Isis, 97(4):790–791, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512918.pdf.

Vetter:2006:BRM

[2773] Jeremy Vetter. Book review: Marcus Hall: Earth Repair: A Transatlantic History of Environmental Restoration Natural Visions: The Power of Images in American Environmental Reform. Isis, 97(4):791–793, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512919.pdf.

Siddiqi:2006:BRJ

[2774] Asif A. Siddiqi. Book review: James R. Hansen: First Man: The Life of Neil A. Armstrong. Isis, 97(4):793–794, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/512920.pdf.

Rushton:2006:BRO

[2775] Alan H. Rushton. Book review: Oren Solomon Harman: The Man Who Invented the Chromosome: A Life of Cyril Darlington. Isis, 97(4):794– 795, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512921.pdf.

Otis:2006:BRK

[2776] Laura Otis. Book review: N. Katherine Hayles: My Mother Was a Computer. Isis, 97(4):795–796, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512922.pdf.

Lassman:2006:BRJ

[2777] Thomas Lassman. Book review: Jeff Hecht: Beam: The Race to Make the Laser. Isis, 97(4):796–797, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512923.pdf.

Kline:2006:BRJ

[2778] Wendy Kline. Book review: Judith A. Houck: Hot and Bothered: Women, Medicine, and Menopause in Modern America. Isis, 97(4):797– 798, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512924.pdf.

Hall:2006:BRA

[2779] Karl Hall. Book review: Alexei B. Kojevnikov: Stalin’s Great Science: The Times and Adventures of Soviet Physicists. Isis, 97(4):798–799, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512925.pdf.

Kenney:2006:BRC

[2780] Martin Kenney. Book review: Christophe Lécuyer: Making Silicon Valley: Innovation and the Growth of High Tech, 1930-1970. Isis, 97(4): 799–800, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512926.pdf.

Foster:2006:BRM

[2781] Amy E. Foster. Book review: Maura Phillips Mackowski: Testing the Limits: Aviation Medicine and the Origins of Manned Space Flight. Isis, 97(4):800–801, December 2006. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/512927.pdf.

Yood:2006:BRA

[2782] Charles Yood. Book review: Arthur L. Norberg: Computers and Commerce: A Study of Technology and Management at Eckert-Mauchly Computer Company, Engineering Research Associates, and Remington Rand, 1946-1957. Isis, 97(4):801–802, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512928.pdf.

McKenzie:2006:BRS

[2783] Matthew McKenzie. Book review: Sara Tjossem: The Journey to PICES: Scientific Cooperation in the North Pacific. Isis, 97(4):802–803, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512929.pdf.

Craver:2006:BRE

[2784] Carl Craver. Book review: Elliot S. Valenstein: The War of the Soups and the Sparks: The Discovery of Neurotransmitters and the Dispute over How Nerves Communicate. Isis, 97(4):803–804, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512930.pdf.

Serlin:2006:BRJ

[2785] David Serlin. Book review: José van Dijck: The Transparent Body: A Cultural Analysis of Medical Imaging. Isis, 97(4):804, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512931.pdf.

Young:2006:BRS

[2786] Chris Young. Book review: Spencer R. Weart: The Discovery of Global Warming. Isis, 97(4):804–806, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/512932.pdf.

Schwach:2006:BRG

[2787] Vera Schwach. Book review: Gary Wright: Ocean Sciences Bridging the Millennia: A Spectrum of Historical Accounts. Isis, 97(4):806–807, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512933.pdf.

Sismondo:2006:BRH

[2788] Sergio Sismondo. Book review: Harry Collins and Trevor Pinch: Dr. Golem: How to Think about Medicine. Isis, 97(4):807–808, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512934.pdf.

Anker:2006:BRK

[2789] Peder Anker. Book review: Kim Cuddington and Beatrix E. Beisner: Ecological Paradigms Lost: Routes of Theory Change (theoretical ecology). Isis, 97(4):808, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/512935.pdf.

Lipton:2006:BRS

[2790] Peter Lipton. Book review: Susan Haack: Defending Science-Within Reason: Between Science and Cynicism. Isis, 97(4):808–809, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/512936.pdf.

Cat:2006:BRG

[2791] Jordi Cat. Book review: George Reisch: How the Cold War Transformed Philosophy of Science: To the Icy Slopes of Logic. Isis, 97(4):809–811, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512937.pdf.

Anonymous:2006:NC

[2792] Anonymous. Notes on contributors. Isis, 97(4):812–816, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/509956.pdf.

Anonymous:2006:IIV

[2793] Anonymous. Index to ISIS, volume 97, 2006. Isis, 97(4):817–828, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
512938.pdf.

Anonymous:2006:CV

[2794] Anonymous. Contents, volume 97. Isis, 97(4):831–832, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/512939.pdf.

Anonymous:2006:F

[2795] Anonymous. Frontispiece. Isis, 97(4):vi, December 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/pdfplus/10.1086/512940.pdf.

Anonymous:2007:CBH

[2796] Anonymous. Current bibliography of the history of science and its cultural influences, 2007. Isis, 98(S1):i–305, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/529412.

Naylor:2007:GTT

[2797] Ron Naylor. Galileo’s tidal theory. Isis, 98(1):1–22, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/512829.

Rankin:2007:BEP

[2798] Alisha Rankin. Becoming an expert practitioner: Court experimentalism and the medical skills of Anna of Saxony (1532–1585). Isis, 98(1):23– 53, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/512830.

Williams:2007:NSE

[2799] Elizabeth A. Williams. Neuroses of the stomach: Eating, gender, and psychopathology in French medicine, 1800–1870. Isis, 98(1):54–79, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/512831.

Alder:2007:I

[2800] Ken Alder. Introduction. Isis, 98(1):80–83, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/512832.

Tresch:2007:TWP

[2801] John Tresch. Technological World-Pictures: Cosmic things and cosmograms. Isis, 98(1):84–99, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/512833.

Hecht:2007:CNT

[2802] Gabrielle Hecht. A cosmogram for nuclear things. Isis, 98(1):100–108, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/512834.

Bijker:2007:DDT

[2803] Wiebe E. Bijker. Dikes and dams, thick with politics. Isis, 98(1):109– 123, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/512835.

Alder:2007:ATG

[2804] Ken Alder. America’s two gadgets: Of bombs and polygraphs. Isis, 98(1):124–137, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
512836.

Latour:2007:CWG

[2805] Bruno Latour. Can we get our materialism back, please? Isis, 98(1):138– 142, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/512837.

Smith:2007:EOH

[2806] Pamela H. Smith. Eloge: Owen Hannaway, 8 October 1939–21 January 2006. Isis, 98(1):143–148, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/512838.

Anonymous:2007:BE

[2807] Anonymous. Isis editorial. Isis, 98(1):149, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/512839.

Hatch:2007:CEP

[2808] Robert Alan Hatch. Clio Electric: Primary texts and digital research in pre-1750 history of science. Isis, 98(1):150–160, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/512840.

Furth:2007:BRL

[2809] Charlotte Furth. Book review: Linda L. Barnes: Needles, Herbs, Gods, and Ghosts: China, Healing, and the West to 1848. Isis, 98(1):161– 162, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519083.

Farber:2007:BRM

[2810] Paul Lawrence Farber. Book review: Marco Beretta: From Private to Public: Natural Collections and Museums. Isis, 98(1):162–163, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519084.

Austin:2007:BRP

[2811] Suzanne Austin. Book review: Patrice Bourdelais and Bart K. Holland: Epidemics Laid Low: A History of What Happened in Rich Countries. Isis, 98(1):163–164, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/519085.

Paradis:2007:BRJ

[2812] Jim Paradis. Book review: John A. Cartwright and Brian Baker: Literature and Science: Social Impact and Interaction. Isis, 98(1):164–165, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/519086.

Pauly:2007:BRP

[2813] Philip J. Pauly. Book review: Patricia Craig: The Department of Plant Biology. Isis, 98(1):165–167, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519087.

Staley:2007:BRO

[2814] Richard Staley. Book review: Olivier Darrigol: Worlds of Flow: A History of Hydrodynamics from the Bernoullis to Prandtl. Isis, 98(1):167– 168, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519088.

Hannaway:2007:BRJ

[2815] Caroline Hannaway. Book review: Jacalyn Duffin: Lovers and Livers: Disease Concepts in History. Isis, 98(1):168–169, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519089.

McConnell:2007:BRC

[2816] Craig Sean McConnell. Book review: Christopher Frayling: Mad, Bad, and Dangerous? The Scientist and the Cinema. Isis, 98(1):169–170, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/519090.

Kassell:2007:BRW

[2817] Lauren Kassell. Book review: Wouter J. Hanegraaff: Dictionary of Gnosis and Western Esotericism. Isis, 98(1):170–171, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519091.

Walker:2007:BRU

[2818] Mark Walker. Book review: Uwe Hoßfeld: Geschichte der biologischen Anthropologie in Deutschland: Von den Anfängen bis in die Nachkriegszeit. Isis, 98(1):171–172, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519092.

Rabin:2007:BRG

[2819] Sheila J. Rabin. Book review: Günther Oestmann, H. Darrel Rutkin and Kocku von Stuckrad: Horoscopes and Public Spheres: Essays on the History of Astrology. Isis, 98(1):172–173, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519093.

Knobloch:2007:BRM

[2820] Eberhard Knobloch. Book review: Michel Serfati and Jacques Bouveresse: La révolution symbolique: La constitution de l’écriture symbolique mathématique. Isis, 98(1):173–174, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519094.

Hollister-Short:2007:BRB

[2821] Graham Hollister-Short. Book review: Brett D. Steele and Tamera Dorland: The Heirs of Archimedes: Science and the Art of War through the Age of Enlightenment. Isis, 98(1):174–175, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519095.

Raphals:2007:BRK

[2822] Lisa Raphals. Book review: Karine Chemla, Guo Shuchun and Geoffrey Lloyd: Les neuf chapitres: Le classique mathématique de la Chine ancienne et ses commentaires. Isis, 98(1):175–176, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519096.

Albala:2007:BRB

[2823] Ken Albala. Book review: Bruno Laurioux: Une histoire culinaire du Moyen Âge. Isis, 98(1):176–177, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519097.

Edson:2007:BRB

[2824] Evelyn Edson. Book review: Barbara Obrist: La cosmologie médiévale: Textes et images. Isis, 98(1):177–178, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/519098.

Lefevre:2007:BRC

[2825] Wolfgang Lefèvre. Book review: Charles van den Heuvel: “De Huysbou”: A Reconstruction of an Unfinished Treatise on Architecture, Town Planning, and Civil Engineering by Simon Stevin. Isis, 98(1):178–179, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519099.

Reeves:2007:BRM

[2826] Eileen Reeves. Book review: Mario Biagioli: Galileo’s Instruments of Credit: Telescopes, Images, Secrecy. Isis, 98(1):179–180, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/519100.

Cormack:2007:BRM

[2827] Lesley Cormack. Book review: Martin Brückner: The Geographic Revolution in Early America: Maps, Literacy, and National Identity. Isis, 98(1):180–181, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519101.

DeLacy:2007:BRS

[2828] Margaret DeLacy. Book review: Stanley Finger: Doctor Franklin’s Medicine. Isis, 98(1):181–182, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519102.

Cushman:2007:BRA

[2829] Gregory T. Cushman. Book review: Antonio González Bueno: Antonio José Cavanilles (1745–1804): La pasión por la ciencia. Isis, 98(1):182– 183, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519103.

Moran:2007:BRM

[2830] Bruce T. Moran. Book review: Margaret C. Jacob: Strangers Nowhere in the World: The Rise of Cosmopolitanism in Early Modern Europe. Isis, 98(1):183–184, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519104.

Zammito:2007:BRE

[2831] John H. Zammito. Book review: Ernst Müller: Ästhetische Religiosität und Kunstreligion in den Philosophien von der Aufklärung bis zum Ausgang des deutschen Idealismus. Isis, 98(1):185–186, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519106.

Rodriguez:2007:BRC

[2832] Julia Rodriguez. Book review: Claudia Agostoni: Monuments of Progress: Modernization and Public Health in Mexico City, 1876–1910. Isis, 98(1):186–187, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/519107.

Ogawa:2007:BRP

[2833] Mariko Ogawa. Book review: Peter Ayres: Harry Marshall Ward and the Fungal Thread of Death. Isis, 98(1):187–188, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519108.

Burba:2007:BRF

[2834] Juliet Burba. Book review: Fredrik Barth, Andre Gingrich, Robert Parkin and Sydel Silverman: One Discipline, Four Ways: British, German, French, and American Anthropology. Isis, 98(1):188–189, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519109.

Nash:2007:BRD

[2835] Stephen Nash. Book review: David L. Caffey: Frank Springer and New Mexico: From the Colfax County War to the Emergence of Modern Santa Fe. Isis, 98(1):189–190, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/519110.

Geschwind:2007:BRG

[2836] Carl-Henry Geschwind. Book review: Gregory Clancey: Earthquake Nation: The Cultural Politics of Japanese Seismicity, 1868–1930. Isis, 98 (1):190–191, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519111.

Newcomb:2007:BRH

[2837] Sally Newcomb. Book review: Helmut W. Flügel: Der Abgrund der Zeit: Die Entwicklung der Geohistorik 1670–1830. Isis, 98(1):191–192, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519112.

Harwood:2007:BRB

[2838] Jonathan Harwood. Book review: Bernd Gausemeier: Natürliche Ordnungen und politische Allianzen: Biologische und biochemische Forschung an Kaiser-Wilhelm-Instituten, 1933–1945. Isis, 98(1):192– 193, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519113.

Hong:2007:BRW

[2839] Sungook Hong. Book review: Walter E. Grunden: Secret Weapons and World War II: Japan in the Shadow of Big Science. Isis, 98(1):193– 194, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519114.

Carson:2007:BRK

[2840] Cathryn Carson. Book review: Klaus Hentschel: Die Mentalität deutscher Physiker in der frühen Nachkriegszeit (1945–1949). Isis, 98 (1):194–195, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519115.

Hoppe:2007:BRR

[2841] Brigitte Hoppe. Book review: Robin Keen: The Life and Work of Friedrich Wöhler (1800–1882). Isis, 98(1):195–196, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519116.

Paoloni:2007:BRS

[2842] Giovanni Paoloni. Book review: Sandra Linguerri: Vito Volterra e il Comitato Talassografico Italiano: Imprese per aria e per mare nell’Italia Unita (1883–1930). Isis, 98(1):196–197, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/519117.

Linhard:2007:BRJ

[2843] Frank Linhard. Book review: Jesper Lützen: Mechanistic Images in Geometric Form: Heinrich Hertz’s Principles of Mechanics. Isis, 98(1):199– 200, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519119.

Bullough:2007:BRB

[2844] Vern L. Bullough. Book review: Bernadette McCauley: Who Shall Take Care of Our Sick? Roman Catholic Sisters and the Development of Catholic Hospitals in New York City. Isis, 98(1):200–201, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/519120.

Nash:2007:BRN

[2845] Linda Nash. Book review: Natalia Molina: Fit to Be Citizens? Public Health and Race in Los Angeles, 1879–1939. Isis, 98(1):201–202, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519121.

Mylott:2007:BRG

[2846] Anne Mylott. Book review: Gerhard Müller-Strahl: Zell-Theorien organischer Materie: Ein Beitrag zu Bildung und Wandel einer Elementarlehre der Lebenswissenschaften. Isis, 98(1):202–203, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519122.

Soderqvist:2007:BRN

[2847] Thomas Söderqvist. Book review: Nicolaas A. Rupke: Alexander von Humboldt: A Metabiography. Isis, 98(1):203–204, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519123.

Webb:2007:BRA

[2848] George E. Webb. Book review: Allan Sandage: The Mount Wilson Observatory: Breaking the Code of Cosmic Evolution. Isis, 98(1):204– 205, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519124.

Fortun:2007:BRA

[2849] Mike Fortun. Book review: Alexandra Minna Stern: Eugenic Nation: Faults and Frontiers of Better Breeding in Modern America. Isis, 98 (1):205–206, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519125.

Guntau:2007:BRJ

[2850] Martin Guntau. Book review: Jacques L. R. Touret and Robert P. W. Visser: Dutch Pioneers of the Earth Sciences. Isis, 98(1):206–207, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519126.

Morus:2007:BRT

[2851] Iwan Rhys Morus. Book review: Ted Underwood: The Work of the Sun: Literature, Science, and Political Economy, 1760–1860. Isis, 98(1):207– 208, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/519127.

Allchin:2007:BRW

[2852] Douglas Allchin. Book review: William Bechtel: Discovering Cell Mechanisms: The Creation of Modern Cell Biology. Isis, 98(1):208–209, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519128.

Rees:2007:BRD

[2853] Amanda Rees. Book review: Donald A. Dewsbury: Monkey Farm: A History of the Yerkes Laboratories of Primate Biology, Orange Park, Florida, 1930–1965. Isis, 98(1):209–210, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/519129.

Garber:2007:BRD

[2854] Elizabeth Garber. Book review: David Edgerton: Warfare State: Britain, 1920–1970. Isis, 98(1):210–211, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/519130.

Hentschel:2007:BRD

[2855] Klaus Hentschel. Book review: David Gugerli, Patrick Kupper and Daniel Speich: Die Zukunftsmaschine: Konjunkturen der ETH Zürich 1855–2005. Isis, 98(1):211–212, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519131.

Maienschein:2007:BRF

[2856] Jane Maienschein. Book review: Frederic Lawrence Holmes and William C. Summers: Reconceiving the Gene: Seymour Benzer’s Adventures in Phage Genetics. Isis, 98(1):212–213, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/519132.

deChadarevian:2007:BRS

[2857] Soraya de Chadarevian. Book review: Susan Lindee: Moments of Truth in Genetic Medicine. Isis, 98(1):213–214, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519133.

Jarrell:2007:BRJ

[2858] Richard A. Jarrell. Book review: Jordan D. Marché II: Theaters of Time and Space: American Planetaria, 1930–1970. Isis, 98(1):214–215, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/519134.

Marche:2007:BRW

[2859] Jordan D. Marché, II. Book review: W. Patrick McCray: Giant Telescopes: Astronomical Ambition and the Promise of Technology. Isis, 98(1):215–216, March 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
519135.

Quirke:2007:BRT

[2860] Viviane Quirke. Book review: Toine Pieters: Interferon: The Science and Selling of a Miracle Drug. Isis, 98(1):216–217, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519136.

Kohlstedt:2007:BRS

[2861] Sally Gregory Kohlstedt. Book review: Sandra Harding: Science and Social Inequality: Feminist and Postcolonial Issues. Isis, 98(1):217–218, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/519137.

Shumelda:2007:BRN

[2862] Mark Shumelda. Book review: Nicholas Rescher: What If? Thought Experimentation in Philosophy. Isis, 98(1):218–219, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/519138.

Anonymous:2007:NCa

[2863] Anonymous. Notes on contributors. Isis, 98(1):220–223, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/512843.

Anonymous:2007:Fa

[2864] Anonymous. Frontispiece. Isis, 98(1):vi, March 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/519082.

Carroll:2007:BPO

[2865] Victoria Carroll. “Beyond the pale of ordinary criticism” eccentricity and the fossil books of Thomas Hawkins. Isis, 98(2):225–265, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/518187.

Keene:2007:EBG

[2866] Melanie Keene. “Every boy & girl a scientist” instruments for children in interwar Britain. Isis, 98(2):266–289, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/518188.

Proctor:2007:LII

[2867] Robert N. Proctor. “-Logos,” “-ismos,” and “-ikos” the political iconicity of denominative suffixes in science (or, phonesthemic tints and taints in the coining of science domain names). Isis, 98(2):290–309, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/518189.

Burnett:2007:ICE

[2868] D. Graham Burnett. Introduction: Cross-Examination? Isis, 98(2):310– 314, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/518190.

DeRenzi:2007:MEB

[2869] Silvia De Renzi. Medical expertise, bodies, and the law in early modern courts. Isis, 98(2):315–322, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/518191.

Kevles:2007:PPP

[2870] Daniel J. Kevles. Patents, protections, and privileges: The establishment of intellectual property in animals and plants. Isis, 98(2):323–331, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/518192.

Winter:2007:FM

[2871] Alison Winter. A forensics of the mind. Isis, 98(2):332–343, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/518193.

Jasanoff:2007:BTK

[2872] Sheila Jasanoff. Bhopal’s trials of knowledge and ignorance. Isis, 98(2): 344–350, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/518194.

Cirillo:2007:EDC

[2873] Vincent J. Cirillo and Karen Reeds. Eloge: David L. Cowen, 1909–2006. Isis, 98(2):351–353, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
518195.

Anonymous:2007:LEa

[2874] Anonymous. Letters to the Editor. Isis, 98(2):354, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/518200.

Dolan:2007:CEO

[2875] Brian Dolan. Clio Electric: Online resources in post-1750 history of science, medicine, and technology. Isis, 98(2):355–361, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/518196.

Jacob:2007:BRB

[2876] Margaret C. Jacob. Book review: The Cambridge History of Science. Isis, 98(2):361–365, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
518197.

Gingras:2007:BRB

[2877] Yves Gingras. Book review: Tout est relatif, Monsieur Poincaré! Isis, 98(2):366–372, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521515.

Hull:2007:BRL

[2878] David L. Hull. Book review: Lee Alan Dugatkin: The Altruism Equation: Seven Scientists Search for the Origins of Goodness. Isis, 98(2):373, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521445.

Griffel:2007:BRA

[2879] Frank Griffel. Book review: Averroès (Ibn Rušd): Commentaire moyen à la Rhétorique d’Aristote: Édition critique du texte arabe et traduction française. Volume 1: Introduction générale. Volume 2: Édition et traduction. Volume 3: Commentaire du commentaire. Isis, 98(2):373–375, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521446.

Lang:2007:BRM

[2880] Helen S. Lang. Book review: Monte Ransome Johnson: Aristotle on Teleology. Isis, 98(2):375, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521447.

Celenza:2007:BRC

[2881] Christopher S. Celenza. Book review: Christiane L. Joost-Gaugier: Measuring Heaven: Pythagoras and His Influence on Thought and Art in Antiquity and the Middle Ages. Isis, 98(2):376–377, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521448.

Akbari:2007:BRD

[2882] Suzanne Conklin Akbari. Book review: Dallas G. Denery II: Seeing and Being Seen in the Later Medieval World: Optics, Theology, and Religious Life. Isis, 98(2):377–378, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521449.

Berggren:2007:BRB

[2883] J. L. Berggren. Book review: In Synchrony with the Heavens: Studies in Astronomical Timekeeping and Instrumentation in Medieval Islamic Civilization, by David A. King; David A. King. Volume 1: The Call of the Muezzin. In Synchrony with the Heavens: Studies in Astronomical Timekeeping and Instrumentation in Medieval Islamic Civilization. Volume 2: Instruments of Mass Calculation. Isis, 98(2):378–379, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521450.

McClellan:2007:BRE

[2884] James E. McClellan, III. Book review: Edward G. Andrew: Patrons of Enlightenment. Isis, 98(2):380, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521451.

Bleichmar:2007:BRB

[2885] Daniela Bleichmar. Book review: Experiencing Nature: The Spanish American Empire and the Early Scientific Revolution A New World of Animals: Early Modern Europeans on the Creatures of Iberian America. Isis, 98(2):380–383, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521452.

Fritscher:2007:BRA

[2886] Bernhard Fritscher. Book review: Albert V. Carozzi: Horace-Bénédict de Saussure (1740–1799): Un pionnier des sciences de la Terre. Isis, 98(2): 383–384, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521453.

Driver:2007:BRP

[2887] Felix Driver. Book review: Philippe Despoix: Le monde mesuré: Dispositifs de l’exploration à l’âge des Lumières. Isis, 98(2):384–385, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521454.

Castelao-Lawless:2007:BRS

[2888] Teresa Castelão-Lawless. Book review: Saul Fisher: Pierre Gassendi’s Philosophy and Science: Atomism for Empiricists. Isis, 98(2):385–386, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/521455.

Kim:2007:BRR

[2889] Mi Gyung Kim. Book review: Russell McCormmach: Speculative Truth: Henry Cavendish, Natural Philosophy, and the Rise of Modern Theoretical Science. Isis, 98(2):386–387, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521456.

Jones:2007:BRS

[2890] Matthew L. Jones. Book review: Stephen A. McKnight: The Religious Foundations of Francis Bacon’s Thought. Isis, 98(2):387–388, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521457.

Richards:2007:BRE

[2891] Joan L. Richards. Book review: Elizabeth Green Musselman: Nervous Conditions: Science and the Body Politic in Early Industrial Britain. Isis, 98(2):388–389, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521458.

Garber:2007:BRW

[2892] Margaret Garber. Book review: William R. Newman: Atoms and Alchemy: Chymistry and the Experimental Origins of the Scientific Revolution. Isis, 98(2):389–390, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521459.

Henry:2007:BRG

[2893] John Henry. Book review: Gianna Pomata and Nancy G. Siraisi: Historia: Empiricism and Erudition in Early Modern Europe. Isis, 98(2): 390–391, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521460.

Charles:2007:BRS

[2894] Loïc Charles. Book review: Sabine Reungoat: William Petty: Observateur des îles Britanniques. Isis, 98(2):391–392, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521461.

Crowther-Heyck:2007:BRG

[2895] Kathleen Crowther-Heyck. Book review: Gerhild Scholz Williams: Ways of Knowing in Early Modern Germany: Johannes Praetorius as a Witness to His Time. Isis, 98(2):392–393, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521462.

Davenport:2007:BRJ

[2896] Anne Davenport. Book review: Jean Seidengart: Dieu, l’univers et la sphère infinie: Penser l’infinité cosmique à l’aube de la science classique. Isis, 98(2):393–394, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521463.

Steigerwald:2007:BRM

[2897] Joan Steigerwald. Book review: Martin Stuber, Stefan Hächler and Luc Lienhard: Hallers Netz: Ein europäischer Gelehrtenbriefwechsel zur Zeit der Aufklärung. Isis, 98(2):394–395, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521464.

Anderson:2007:BRD

[2898] Katharine Anderson. Book review: David Arnold: The Tropics and the Traveling Gaze: India, Landscape, and Science, 1800–1856. Isis, 98(2): 395–396, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521465.

Herman:2007:BRJ

[2899] Ellen Herman. Book review: Jesse F. Ballenger: Self, Senility, and Alzheimer’s Disease in Modern America: A History. Isis, 98(2):396– 397, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521466.

Brock:2007:BRR

[2900] Emily Brock. Book review: Richard G. Beidleman: California’s Frontier Naturalists. Isis, 98(2):397–398, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521467.

Smocovitis:2007:BRM

[2901] Vassiliki Betty Smocovitis. Book review: Michael Bulmer: Francis Galton: Pioneer of Heredity and Biometry. Isis, 98(2):398–399, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521468.

Illy:2007:BRJ

[2902] József Illy. Book review: Jeffrey Crelinsten: Einstein’s Jury: The Race to Test Relativity. Isis, 98(2):399–400, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521469.

Levit:2007:BRM

[2903] Georgy S. Levit. Book review: Mario A. Di Gregorio: From Here to Eternity: Ernst Haeckel and Scientific Faith. Isis, 98(2):400–401, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521470.

Laubichler:2007:BRM

[2904] Manfred D. Laubichler. Book review: Matthias Dörries: Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Isis, 98(2):401– 402, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521471.

Lanzoni:2007:BRD

[2905] Susan Lanzoni. Book review: Denis Forest: Histoire des aphasies: Une anatomie de l’expression. Isis, 98(2):402–403, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/521472.

Bowker:2007:BRB

[2906] Geoffrey C. Bowker. Book review: Benoît Godin: Measurement and Statistics on Science and Technology: 1920 to the Present. Isis, 98(2): 403–404, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521473.

Maat:2007:BRJ

[2907] Harro Maat. Book review: Jonathan Harwood: Technology’s Dilemma: Agricultural Colleges between Science and Practice in Germany, 1860– 1934. Isis, 98(2):404–405, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521474.

Pols:2007:BRL

[2908] Hans Pols. Book review: Lynette A. Jackson: Surfacing Up: Psychiatry and Social Order in Colonial Zimbabwe, 1908–1968. Isis, 98(2):405– 406, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521475.

Mills:2007:BRM

[2909] James H. Mills. Book review: Margaret Jones: Health Policy in Britain’s Model Colony: Ceylon (1900–1948). Isis, 98(2):406–407, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521476.

Morat:2007:BRA

[2910] Daniel Morat. Book review: Andreas Killen: Berlin Electropolis: Shock, Nerves, and German Modernity. Isis, 98(2):407–408, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521477.

Brain:2007:BRB

[2911] Robert M. Brain. Book review: Barbara Larson: The Dark Side of Nature: Science, Society, and the Fantastic in the Work of Odilon Redon. Isis, 98(2):408–409, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521478.

Besse:2007:BRD

[2912] Jean-Marc Besse and Marie-Claire Robic. Book review: David N. Livingstone and Charles W. J. Withers: Geography and Revolution. Isis, 98(2): 409–411, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521479.

Landecker:2007:BRK

[2913] Hannah Landecker. Book review: Kirsten Ostherr: Cinematic Prophylaxis: Globalization and Contagion in the Discourse of World Health. Isis, 98(2):411–412, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521480.

Hardy:2007:BRS

[2914] Anne Hardy. Book review: Scott H. Podolsky: Pneumonia before Antibiotics: Therapeutic Evolution and Evaluation in Twentieth-Century America. Isis, 98(2):412–413, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521481.

Boudia:2007:BRJ

[2915] Soraya Boudia. Book review: Jean-Pierre Poirier: Marie Curie et les conquérants de l’atome, 1896–2006. Isis, 98(2):413, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521482.

Kragh:2007:BRC

[2916] Helge Kragh. Book review: Carsten Reinhardt: Shifting and Rearranging: Physical Methods and the Transformation of Modern Chemistry. Isis, 98(2):413–414, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521483.

Mandelbrote:2007:BRM

[2917] Scott Mandelbrote. Book review: Margaret Schabas: The Natural Origins of Economics. Isis, 98(2):414–416, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/521484.

Goodfellow:2007:BRC

[2918] Sarah Goodfellow. Book review: Chandak Sengoopta: The Most Secret Quintessence of Life: Sex, Glands, and Hormones, 1850–1950. Isis, 98(2):416–417, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521485.

Josefowicz:2007:BRT

[2919] Diane Greco Josefowicz. Book review: Thomas K. Simpson: Figures of Thought: A Literary Appreciation of Maxwell’s. Isis, 98(2):417, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521486.

Rusnock:2007:BRF

[2920] Andrea Rusnock. Book review: Frank M. Snowden: The Conquest of Malaria: Italy, 1900–1962. Isis, 98(2):417–418, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521487.

Fontaine:2007:BRP

[2921] Philippe Fontaine. Book review: Philippe Steiner: L’école durkheimienne et l’économie: Sociologie, religion et connaissance. Isis, 98(2):418–420, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/521488.

Heffernan:2007:BRL

[2922] Michael Heffernan. Book review: Leon I. Yacher: The Role of Geographer and Natural Scientist Henri François Pittier (1857–1950) in the Evolution of Geography as a Science in Costa Rica. Isis, 98(2):420, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521489.

Rothenberg:2007:BRL

[2923] Marc Rothenberg. Book review: Louis Brown: The Department of Terrestrial Magnetism The Geophysical Laboratory. Hatten S. Yoder, Jr.: The Geophysical Laboratory. Isis, 98(2):420–422, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521490.

Durbach:2007:BRJ

[2924] Nadja Durbach. Book review: James Colgrove: State of Immunity: The Politics of Vaccination in Twentieth-Century America. Isis, 98(2):422– 423, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521491.

Rothschild:2007:BRS

[2925] Lynn Rothschild. Book review: Stephen J. Dick and James E. Strick: The Living Universe: NASA and the Development of Astrobiology. Isis, 98(2): 423–424, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521492.

Lederer:2007:BRB

[2926] Susan E. Lederer. Book review: The Nuremberg Medical Trial: The Holocaust and the Origin of the Nuremberg Medical Code, by Horst H. Freyhofer; Paul Julian Weindling. Nazi Medicine and the Nuremberg Trials: From Medical War Crimes to Informed Consent. Isis, 98(2):424– 425, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521493.

Burnett:2007:BRM

[2927] D. Graham Burnett. Book review: Michael Heazle: Scientific Uncertainty and the Politics of Whaling. Isis, 98(2):425–427, June 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521494.

Kroker:2007:BRM

[2928] Kenton Kroker. Book review: Mark Jackson: Allergy: The History of a Modern Malady. Isis, 98(2):427–428, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521495.

Davis:2007:BRT

[2929] Frederick R. Davis. Book review: Tina Loo: States of Nature: Conserving Canada’s Wildlife in the Twentieth Century. Isis, 98(2):428, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521496.

Rainger:2007:BRJ

[2930] Ronald Rainger. Book review: Jane Maienschein, Marie Glitz and Garland E. Allen: The Department of Embryology. Isis, 98(2):429–430, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521497.

Sze:2007:BRM

[2931] Julie Sze. Book review: Michelle Murphy: Sick Building Syndrome and the Problem of Uncertainty: Environmental Politics, Technoscience, and Women Workers. Isis, 98(2):430–431, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521498.

Linker:2007:BRN

[2932] Beth Linker. Book review: Nicholas L. Tilney: A Perfectly Striking Departure: Surgeons and Surgery at the Peter Bent Brigham Hospital, 1912–1980. Isis, 98(2):431–432, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521499.

Ankeny:2007:BRC

[2933] Rachel A. Ankeny. Book review: Catherine Waldby and Robert Mitchell: Tissue Economies: Blood, Organs, and Cell Lines in Late Capitalism. Isis, 98(2):432–433, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521500.

Wynn:2007:BRA

[2934] James Wynn. Book review: Alan G. Gross: Starring the Text: The Place of Rhetoric in Science Studies. Isis, 98(2):433–434, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521501.

Restivo:2007:BRB

[2935] Sal Restivo. Book review: Barbara Herrnstein Smith: Scandalous Knowledge: Science, Truth, and the Human. Isis, 98(2):434–435, June 2007.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521502.

Brigandt:2007:BRP

[2936] Ingo Brigandt. Book review: P. Kyle Stanford: Exceeding Our Grasp: Science, History, and the Problem of Unconceived Alternatives. Isis, 98(2):435–436, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521503.

Anonymous:2007:ACB

[2937] Anonymous. Adrian C. Brock: Internationalizing the History of Psychology. Isis, 98(2):436, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521505.

Anonymous:2007:JRB

[2938] Anonymous. José Ramón Bertomeu-Sánchez and Agustí Nieto-Galan: Chemistry, Medicine, and Crime: Mateu J. B. Orfila (1787–1853) and His Times. Isis, 98(2):436, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521504.

Anonymous:2007:BRN

[2939] Anonymous. Book review: Nina Byers and Gary Williams: Out of the Shadows: Contributions of Twentieth-Century Women to Physics. Isis, 98(2):436–437, June 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521506.

Anonymous:2007:BRMa

[2940] Anonymous. Book review: Marco Mamone Capria: Physics before and after Einstein. Isis, 98(2):437, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521507.

Anonymous:2007:MFF

[2941] Anonymous. Mary Frank fox, doborah g. Johnson and sue v. rosser: Women, Gender, and Technology. Isis, 98(2):437–438, June 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521508.

Anonymous:2007:HGH

[2942] Anonymous. Hart grob and Hans hooijmaijers: Who Needs Scientific Instruments. Isis, 98(2):438, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521509.

Anonymous:2007:BRK

[2943] Anonymous. Book review: E. I. Kolchinsky and M. B. Konashev: Na perelome: Otechestvennaja nauka v konce XIX–XX veke: istochniki, issledovanija, istoriografija [On the Edge: Science in Russia in the End of XIX–XX Century: Researches, Sources, Historiography]. Isis, 98(2): 438–439, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521510.

Anonymous:2007:BRG

[2944] Anonymous. Book review: G. J. H. McCall, A. J. Bowden and R. J. Howarth: The History of Meteoritics and Key Meteorite Collections: Fireballs, Falls, and Finds. Isis, 98(2):439, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521511.

Anonymous:2007:BRMb

[2945] Anonymous. Book review: M. Teresa Tavormina: Sex, Aging, and Death in a Medieval Medical Compendium: Trinity College Cambridge MS R. 14.52, Its Texts, Language, and Scribe. Isis, 98(2):439, June 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521513.

Anonymous:2007:TRJ

[2946] Anonymous. Tom Rockmore and Joseph Margolis: History, Historicity, and Science. Isis, 98(2):439, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521512.

Anonymous:2007:CWB

[2947] Anonymous. Charles T. Wolfe: Monsters and Philosophy. Isis, 98(2): 439–440, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521514.

Anonymous:2007:NCb

[2948] Anonymous. Notes on contributors. Isis, 98(2):441–445, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/518199.

Anonymous:2007:Fb

[2949] Anonymous. Frontispiece. Isis, 98(2):vi, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/521516.

Lehoux:2007:OOE

[2950] Daryn Lehoux. Observers, objects, and the embedded eye; or, Seeing and knowing in Ptolemy and Galen. Isis, 98(3):447–467, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521152.

Kingsland:2007:MCT

[2951] Sharon E. Kingsland. Maintaining continuity through a scientific revolution: A rereading of E. B. Wilson and T. H. Morgan on sex determination and Mendelism. Isis, 98(3):468–488, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/521153.

Pickstone:2007:WKB

[2952] John V. Pickstone. Working knowledges before and after circa 1800: Practices and disciplines in the history of science, technology, and medicine. Isis, 98(3):489–516, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/521154.

Elman:2007:NDH

[2953] Benjamin A. Elman. New directions in the history of modern science in China: Global science and comparative history. Isis, 98(3):517–523, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521155.

Fan:2007:RMS

[2954] Fa ti Fan. Redrawing the map: Science in Twentieth-Century China. Isis, 98(3):524–538, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/521156.

Hu:2007:RRC

[2955] Danian Hu. The reception of Relativity in China. Isis, 98(3):539–557, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521157.

Wang:2007:SSM

[2956] Zuoyue Wang. Science and the state in modern China. Isis, 98(3): 558–570, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
521158.

Schmalzer:2007:AUR

[2957] Sigrid Schmalzer. On the appropriate use of Rose-Colored glasses: Reflections on science in socialist China. Isis, 98(3):571–583, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521159.

Shen:2007:MWT

[2958] Grace Shen. Murky waters: Thoughts on desire, utility, and the “sea of modern science”. Isis, 98(3):584–596, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/521160.

Anonymous:2007:NP

[2959] Anonymous. News of the profession. Isis, 98(3):597–601, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/521161.

Musselman:2007:ESF

[2960] Elizabeth Green Musselman and Karen A. Rader. Eloge: Sylvia Freeman Wallace Mcgrath, 1937–2006. Isis, 98(3):602–604, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524227.

Plofker:2007:EDP

[2961] Kim Plofker. Eloge: David Pingree, 1933–2005. Isis, 98(3):605–607, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524228.

Anonymous:2007:LEb

[2962] Anonymous. Letters to the Editor. Isis, 98(3):608–612, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521162.

Hu:2007:BRB

[2963] Danian Hu. Book review: Benjamin A. Elman: A Cultural History of Modern Science in China. (New Histories of Science, Technology, and Medicine.). Isis, 98(3):613–614, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524230.

Pitt:2007:BRM

[2964] Joseph C. Pitt. Book review: Mikael Hård and Andrew Jamison: Hubris and Hybrids: A Cultural History of Technology and Science. Isis, 98 (3):614–615, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524231.

Schulten:2007:BRC

[2965] Susan Schulten. Book review: Christian Jacob: The Sovereign Map: Theoretical Approaches in Cartography throughout History. Isis, 98(3): 615–616, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524232.

Nordlund:2007:BRD

[2966] Christer Nordlund. Book review: David R. Oldroyd: Earth, Water, Ice, and Fire: Two Hundred Years of Geological Research in the English Lake District. (Geological Society Memoirs, 25.). Isis, 98(3):616–617, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524233.

Davis:2007:BRR

[2967] Edward B. Davis. Book review: Richard G. Olson: Science and Religion, 1450–1900: From Copernicus to Darwin. Isis, 98(3):617–618, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524234.

Ede:2007:BRJ

[2968] Andrew Ede. Book review: John Waller: Leaps in the Dark: The Making of Scientific Reputations. Isis, 98(3):618–619, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524235.

Betegh:2007:BRG

[2969] Gábor Betegh. Book review: Gabriela Roxana Carone: Plato’s Cosmology and Its Ethical Dimensions. Isis, 98(3):619–620, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524236.

Schmidt:2007:BRA

[2970] Jeremy Schmidt. Book review: Angus Gowland: The Worlds of Renaissance Melancholy: Robert Burton in Context. (Ideas in Context.). Isis, 98(3):620–621, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524237.

Hoyrup:2007:BRS

[2971] Jens Høyrup. Book review: Sabine Rommevaux: Clavius: Une clé pour Euclide au XVIe siècle. (Mathesis.). Isis, 98(3):621–622, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524238.

Wallis:2007:BRV

[2972] Faith Wallis. Book review: Victoria Sweet: Rooted in the Earth, Rooted in the Sky: Hildegard of Bingen and Premodern Medicine. (Studies in Medieval History and Culture.). Isis, 98(3):622–623, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524285.

Smith:2007:BRV

[2973] Justin E. H. Smith. Book review: Vincent Aucante: La philosophie médicale de Descartes. Isis, 98(3):623–625, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524240.

Downing:2007:BRM

[2974] Lisa Downing. Book review: Michael Ben-Chaim: Experimental Philosophy and the Birth of Empirical Science: Boyle, Locke, and Newton. Isis, 98(3):625–626, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524286.

Wilson:2007:BRM

[2975] Catherine Wilson. Book review: Michael Cooper and Michael Hunter: Robert Hooke: Tercentennial Studies. Isis, 98(3):626–627, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524242.

Levere:2007:BRM

[2976] Trevor Levere. Book review: Maurice Crosland: The Language of Science: From the Vernacular to the Technical. Isis, 98(3):627–628, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/524243.

Segala:2007:BRL

[2977] Marco Segala. Book review: Lucia De Frenza: I sonnambuli delle miniere: Amoretti, Fortis, Spallanzani e il dibattito sull’elettrometria organica e minerale in Italia (1790–1816). (Biblioteca di Storia della Scienza, 50.). Isis, 98(3):628–629, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/524244.

Golinski:2007:BRJ

[2978] Jan Golinski. Book review: James Delbourgo: A Most Amazing Scene of Wonders: Electricity and Enlightenment in Early America. Isis, 98 (3):629–630, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524245.

Scharf:2007:BRL

[2979] Sara T. Scharf. Book review: Lipke B. Holthuis and Theodore W. Pietsch: Les planches inédites de poissons et autres animaux marins de l’Indo-Ouest Pacifique d’Isaac Johannes Lamotius/Isaac Johannes Lamotius (1646–c. 1718) and His Paintings of Indo–West Pacific Fishes and Other Marine Animals. Isis, 98(3):630–631, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524246.

Lynn:2007:BRM

[2980] Michael R. Lynn. Book review: Margaret C. Jacob and Larry Stewart: Practical Matter: Newton’s Science in the Service of Industry and Empire, 1687–1851. Isis, 98(3):631–632, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/524247.

Cottingham:2007:BRM

[2981] John Cottingham. Book review: Matthew L. Jones: The Good Life in the Scientific Revolution: Descartes, Pascal, Leibniz, and the Cultivation of Virtue. Isis, 98(3):632–633, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524248.

Lachapelle:2007:BRM

[2982] Sofie Lachapelle. Book review: Michael R. Lynn: Popular Science and Public Opinion in Eighteenth-Century France. Isis, 98(3):633–634, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524249.

Clulee:2007:BRB

[2983] Nicholas Clulee. Book review: Bruce T. Moran: Distilling Knowledge: Alchemy, Chemistry, and the Scientific Revolution. (New Histories of Science, Technology, and Medicine.). Isis, 98(3):634–635, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524250.

Herbert:2007:BRP

[2984] Sandra Herbert. Book review: Patrick Armstrong: Darwin’s Other Islands. Isis, 98(3):635–636, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524251.

Carroll:2007:BRW

[2985] Victoria Carroll. Book review: William J. Astore: Observing God: Thomas Dick, Evangelicalism, and Popular Science in Victorian Britain and America. Isis, 98(3):636–637, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/524252.

Beck:2007:BRV

[2986] Naomi Beck. Book review: Valeria P. Babini: Il caso Murri: Una storia italiana. Isis, 98(3):637–638, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524253.

Amsterdamska:2007:BRD

[2987] Olga Amsterdamska. Book review: David S. Barnes: The Great Stink of Paris and the Nineteenth-Century Struggle against Filth and Germs. Isis, 98(3):638–639, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524287.

Paul:2007:BRA

[2988] Harry W. Paul. Book review: Antonio Cadeddu: Les vérités de la science: Pratique, récit, histoire: Le cas Pasteur. (Biblioteca di Nuncius: Studi e Testi, 57.). Isis, 98(3):639–641, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524255.

Abeles:2007:BRB

[2989] Francine F. Abeles. Book review: Arthur Cayley: Mathematician Laureate of the Victorian Age, by Tony Crilly; Karen Hunger Parshall. James Joseph Sylvester: Jewish Mathematician in a Victorian World. Isis, 98 (3):641–642, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524256.

Hughes:2007:BRD

[2990] Thomas P. Hughes. Book review: David Edgerton: The Shock of the Old: Technology and Global History since 1900. Isis, 98(3):642–643, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524257.

Holton:2007:BRJ

[2991] Gerald Holton. Book review: Jean Eisenstaedt: The Curious History of Relativity: How Einstein’s Theory of Gravity Was Lost and Found Again. Isis, 98(3):643–644, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524258.

Widmalm:2007:BRA

[2992] Sven Widmalm. Book review: Aant Elzinga: Einstein’s Nobel Prize: A Glimpse behind Closed Doors: The Archival Evidence. Isis, 98(3): 644–645, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524259.

Lawrence:2007:BRM

[2993] Christopher Lawrence. Book review: Marianne P. Fedunkiw: Rockefeller Foundation Funding and Medical Education in Toronto, Montreal, and Halifax. (McGill–Queen’s/Associated Medical Services in the History of Medicine, Health, and Society, 24.). Isis, 98(3):645–646, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524260.

Huneman:2007:BRM

[2994] Philippe Huneman. Book review: Michael Friedman and Alfred Nordmann: (Editors). The Kantian Legacy in Nineteenth-Century Science. (Dibner Institute Studies in the History of Science and Technology.). Isis, 98(3):646–647, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524261.

Grapi:2007:BRL

[2995] Pere Grapí. Book review: Lluís Garrigós-Oltra, Carles Millán-Verdú and Georgina Blanes-Nadal: El color líquido: Instrumentos y útiles de la colorimetría en el siglo XIX. Isis, 98(3):647–648, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524262.

Mazzotti:2007:BRA

[2996] Massimo Mazzotti. Book review: Angelo Guerraggio and Pietro Nastasi: Italian Mathematics between the Two World Wars. (Science Networks Historical Studies, 29.). Isis, 98(3):648–649, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524263.

Kremer:2007:BRK

[2997] Richard L. Kremer. Book review: Klaus Hentschel: Gaußens unsichtbare Hand: Der Universitäts-Mechanicus und Maschinen-Inspector Moritz Meyerstein: Ein Instrumentenbauer im 19. Jahrhundert. (Abhandlungen der Akademie der Wissenschaften zu Göttingen, Ser. 3, 52). Isis, 98 (3):649–650, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524264.

Barrow:2007:BRR

[2998] Mark V. Barrow, Jr. Book review: Robert E. Kohler: All Creatures: Naturalists, Collectors, and Biodiversity, 1850–1950. Isis, 98(3):650–651, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524265.

Gates:2007:BRL

[2999] Barbara T. Gates. Book review: Linda Lear: Beatrix Potter: A Life in Nature. Isis, 98(3):651–652, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524266.

Nieto-Galan:2007:BRE

[3000] Agustí Nieto-Galan. Book review: Esther Leslie: Synthetic Worlds: Nature, Art, and the Chemical Industry. Isis, 98(3):652–653, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524267.

Klein:2007:BRD

[3001] Ursula Klein. Book review: David Philip Miller: Discovering Water: James Watt, Henry Cavendish, and the Nineteenth-Century “Water Controversy.” (Science, Technology, and Culture, 1700–1945.). Isis, 98 (3):653–654, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524268.

Aisenberg:2007:BRK

[3002] Andrew Aisenberg. Book review: Kim Pelis: Charles Nicolle, Pasteur’s Imperial Missionary: Typhus and Tunisia. (Rochester Studies in Medical History.). Isis, 98(3):654–655, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524269.

Reidy:2007:BRS

[3003] Michael S. Reidy. Book review: Sujit Sivasundaram: Nature and the Godly Empire: Science and Evangelical Mission in the Pacific, 1795– 1850. (Cambridge Social and Cultural Histories, 7.). Isis, 98(3):655–656, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524270.

Yeo:2007:BRL

[3004] Richard Yeo. Book review: Laura J. Snyder: Reforming Philosophy: A Victorian Debate on Science and Society. Isis, 98(3):656–657, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524271.

Gass:2007:BRL

[3005] Gillian Gass. Book review: Lester D. Stephens and Dale R. Calder: Seafaring Scientist: Alfred Goldsborough Mayor, Pioneer in Marine Biology. Isis, 98(3):658, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/524272.

Richmond:2007:BRA

[3006] Marsha L. Richmond. Book review: Arleen Marcia Tuchman: Science Has No Sex: The Life of Marie Zakrzewska, M.D. Isis, 98(3):658–659, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524273.

Thomas:2007:BRB

[3007] Julia Adeney Thomas. Book review: Brett L. Walker: The Lost Wolves of Japan. Isis, 98(3):659–660, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524274.

Faye:2007:BRG

[3008] Jan Faye. Book review: Gerald Holton: Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Isis, 98(3):660–661, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/524275.

DeVorkin:2007:BRM

[3009] David H. DeVorkin. Book review: Malcolm S. Longair: The Cosmic Century: A History of Astrophysics and Cosmology. Isis, 98(3):661–662, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524276.

Enebakk:2007:BRP

[3010] Vidar Enebakk. Book review: Peter Broks: Understanding Popular Science. Isis, 98(3):662–663, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524277.

Oberdan:2007:BRB

[3011] Thomas Oberdan. Book review: Bernard d’Espagnat: On Physics and Philosophy. Isis, 98(3):663–664, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/524278.

Mitchell:2007:BRV

[3012] Sandra D. Mitchell. Book review: Vittorio Hösle and Christian Illies: Darwinism and Philosophy. Isis, 98(3):664–665, September 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524279.

Weber:2007:BRH

[3013] Marcel Weber. Book review: Hans-Jörg Rheinberger: Epistemologie des Konkreten: Studien zur Geschichte der modernen Biologie. (Suhrkamp Taschenbuch Wissenschaft.). Isis, 98(3):665–666, September 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/524280.

Gordin:2007:BRE

[3014] Michael D. Gordin. Book review: Eric R. Scerri: The Periodic Table: Its Story and Its Significance. Isis, 98(3):666–667, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/524281.

Perini:2007:BRE

[3015] Laura Perini. Book review: Edward R. Tufte: Beautiful Evidence. Isis, 98 (3):667–668, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
524282.

Anonymous:2007:NCc

[3016] Anonymous. Notes on contributors. Isis, 98(3):669–673, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/521164.

Anonymous:2007:Fc

[3017] Anonymous. Frontispiece. Isis, 98(3):vi, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/524290.

Burkhardt:2007:LGL

[3018] Richard W. Burkhardt, Jr. The leopard in the garden: Life in close quarters at the Muséum d’Histoire Naturelle. Isis, 98(4):675–694, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529263.

Binzley:2007:ACS

[3019] Ronald A. Binzley. American Catholicism’s science crisis and the Albertus Magnus Guild, 1953–1969. Isis, 98(4):695–723, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529264.

Sutter:2007:NAA

[3020] Paul S. Sutter. Nature’s agents or agents of empire? Entomological workers and environmental change during the construction of the Panama Canal. Isis, 98(4):724–754, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529265.

Golinski:2007:I

[3021] Jan Golinski. Introduction. Isis, 98(4):755–759, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529266.

Olesko:2007:WWL

[3022] Kathryn M. Olesko. The world we have lost: History as art. Isis, 98 (4):760–768, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529267.

Gascoigne:2007:GFL

[3023] John Gascoigne. “Getting a fix”: The longitude phenomenon. Isis, 98 (4):769–778, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529268.

Principe:2007:TH

[3024] Lawrence M. Principe. Transmuting history. Isis, 98(4):779–787, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529269.

Gillispie:2007:DM

[3025] Charles C. Gillispie. The distorted Meridian. Isis, 98(4):788–795, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529270.

King:2007:EAH

[3026] David A. King. Eloge: Asger Hartvig Aaboe, 1922–2007. Isis, 98(4):796– 798, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529271.

Haubold:2007:LE

[3027] Hans J. Haubold. Letters to the Editor. Isis, 98(4):799–800, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529272.

Siegel:2007:BRC

[3028] Daniel Siegel. Book review: The collected papers of Albert Einstein. Isis, 98(4):800, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Daston:2007:BRB

[3029] Lorraine Daston. Book review: The Emergence of Probability: A Philosophical Study of Early Ideas about Probability, Induction, and Statistical Inference. Isis, 98(4):801–808, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529273.

Mijnhardt:2007:BRB

[3030] Wijnand Mijnhardt. Book review: Matters of Exchange: Commerce, Medicine, and Science in the Dutch Golden Age. Isis, 98(4):809–812, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529274.

Todes:2007:BRW

[3031] Daniel P. Todes. Book review: W. F. Bynum; Anne Hardy; Stephen Jacyna; Christopher Lawrence; E. M. Tansey: The Western Medical Tradition, 1800 to 2000. Isis, 98(4):813–814, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529279.

Ruse:2007:BRR

[3032] Michael Ruse. Book review: Richard Dawkins: The God Delusion. Isis, 98(4):814–816, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529280.

Maienschein:2007:BRS

[3033] Jane Maienschein. Book review: Sarah Franklin: Dolly Mixtures: The Remaking of Genealogy. Isis, 98(4):816–817, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529281.

Albu:2007:BRA

[3034] Emily Albu. Book review: Alessandro Scafi: Mapping Paradise: A History of Heaven on Earth. Isis, 98(4):817–818, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529282.

Hoppe:2007:BRB

[3035] Brigitte Hoppe. Book review: Geschichte der Pharmazie, by Rudolph Schmitz; Christoph Friedrich; Wolf–Dieter Müller–Jahncke. Volume 2:

Von der Frühen Neuzeit bis zur Gegenwart. Isis, 98(4):818–819, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529283.

Fara:2007:BRB

[3036] Patricia Fara. Book review: Figuring It Out: Science, Gender, and Visual Culture. Isis, 98(4):819–821, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529284.

Mastorakou:2007:BRB

[3037] Stamatina Mastorakou. Book review: Euclid’s Phaenomena: A Translation and Study of a Hellenistic Treatise in Spherical Astronomy. Isis, 98 (4):821–822, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529285.

Gere:2007:BRB

[3038] Cathy Gere. Book review: Brian M. Fagan: Return to Babylon: Travelers, Archaeologists, and Monuments in Mesopotamia. Isis, 98(4):822– 823, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529286.

Pellegrin:2007:BRA

[3039] Pierre Pellegrin. Book review: Andrea Falcon: Aristotle and the Science of Nature: Unity without Uniformity. Isis, 98(4):823–824, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529287.

Alberti:2007:BRS

[3040] Samuel J. M. M. Alberti. Book review: Stephanie Moser: Wondrous Curiosities: Ancient Egypt at the British Museum. Isis, 98(4):824–825, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529288.

Smith:2007:BRB

[3041] A. Mark Smith. Book review: Nicole Oresme’s De visione stellarum (On Seeing the Stars): A Critical Edition of Oresme’s Treatise on Optics and Atmospheric Refraction, with an Introduction, Commentary, and English Translation. Isis, 98(4):825–826, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529289.

Hayton:2007:BRL

[3042] Darin Hayton. Book review: Lauren Kassell: Medicine and Magic in Elizabethan London: Simon Forman: Astrologer, Alchemist, and Physician. Isis, 98(4):826–827, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529290.

Hehmeyer:2007:BRP

[3043] Ingrid Hehmeyer. Book review: Peter E. Pormann; Emilie Savage-Smith: Medieval Islamic Medicine. Isis, 98(4):827–828, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529291.

Beretta:2007:BRK

[3044] Marco Beretta. Book review: Ken Arnold: Cabinets for the Curious: Looking Back at Early English Museums. Isis, 98(4):828–829, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529292.

Sobol:2007:BRA

[3045] Peter G. Sobol. Book review: Arun Bala: The Dialogue of Civilizations in the Birth of Modern Science. Isis, 98(4):829–830, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529293.

Clark:2007:BRW

[3046] Stephanie Brown Clark. Book review: A. W. Bates: Emblematic Monsters: Unnatural Conceptions and Deformed Births in Early Modern Europe. Isis, 98(4):830–831, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529294.

Muller-Sievers:2007:BRO

[3047] Helmut Müller-Sievers. Book review: Olaf Breidbach: Goethes Metamorphosenlehre. Isis, 98(4):831–832, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529295.

Conlin:2007:BRJ

[3048] Michael F. Conlin. Book review: Joyce E. Chaplin: The First Scientific American: Benjamin Franklin and the Pursuit of Genius. Isis, 98(4):832– 833, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529296.

Wakefield:2007:BRA

[3049] Andre Wakefield. Book review: Alix Cooper: Inventing the Indigenous: Local Knowledge and Natural History in Early Modern Europe. Isis, 98 (4):833–834, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529297.

McEvoy:2007:BRB

[3050] John G. McEvoy. Book review: James Watt, by Richard L. Hills; Richard L. Hills; Richard L. Hills. Volume 1: His Time in Scotland, 1736–1774 James Watt. Volume 2: The Years of Toil, 1775–1785 James Watt. Volume 3: Triumph through Adversity, 1785–1819. Isis, 98(4):834–836, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529298.

Langins:2007:BRJ

[3051] Janis Langins. Book review: Jeff Horn: The Path Not Taken: French Industrialization in the Age of Revolution, 1750–1830. Isis, 98(4):836– 837, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529299.

Gaukroger:2007:BRA

[3052] Stephen Gaukroger. Book review: Antonia LoLordo: Pierre Gassendi and the Birth of Early Modern Philosophy. Isis, 98(4):837–838, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529300.

Finocchiaro:2007:BRA

[3053] Maurice A. Finocchiaro. Book review: Antonio Beltrán Marí: Talento y poder: Historia de las relaciones entre Galileo y la Iglesia católica. Isis, 98(4):838–839, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529301.

Carneiro:2007:BRJ

[3054] Ana Carneiro. Book review: José Ramón Bertomeu Sánchez; Antonio García Belmar: La revolución química: Entre la historia y la memoria. Isis, 98(4):839–840, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529302.

Daniel:2007:BRU

[3055] Dane T. Daniel. Book review: Urszula Szulakowska: The Sacrificial Body and the Day of Doom: Alchemy and Apocalyptic Discourse in the Protestant Reformation. Isis, 98(4):840–841, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529303.

Canizares-Esguerra:2007:BRS

[3056] Jorge Cañizares-Esguerra. Book review: Santiago Aragón Albillos: El Zoológico del Museo de Ciencias Naturales de Madrid: Mariano de la Paz Graells (1809–1898), la Sociedad de Aclimatación y los animales útiles. Isis, 98(4):841–842, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529304.

Faye:2007:BRL

[3057] Cathy Faye. Book review: Ludy T. Benjamin, Jr.: A Brief History of Modern Psychology. Isis, 98(4):842–843, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529305.

Gradmann:2007:BRB

[3058] Christoph Gradmann. Book review: Hirnströme: Eine Kulturgeschichte der Elektroenzephalographie. Isis, 98(4):843–844, December 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529306.

Darnell:2007:BRS

[3059] Regna Darnell. Book review: Steven Conn: History’s Shadow: Native Americans and Historical Consciousness in the Nineteenth Century. Isis, 98(4):844–845, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529307.

Scharf:2007:BRP

[3060] Sara Scharf. Book review: Pietro Corsi; Jean Gayon; Gabriel Gohau; Stéphane Tirard: Lamarck, philosophe de la nature. Isis, 98(4):846, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529308.

Golinski:2007:BRB

[3061] Jan Golinski. Book review: Intimate Universality: Local and Global Themes in the History of Weather and Climate. Isis, 98(4):846–847, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529309.

Waterman:2007:BRJ

[3062] A. M. C. Waterman. Book review: James P. Huzel: The Popularization of Malthus in Early Nineteenth-Century England: Martineau, Cobbett, and the Pauper Press. Isis, 98(4):847–849, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529310.

Bach:2007:BRB

[3063] Thomas Bach. Book review: Bernhard Kleeberg: Theophysis: Ernst Haeckels Philosophie des Naturganzen. Isis, 98(4):849–850, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/529311.

Sivasundaram:2007:BRH

[3064] Sujit Sivasundaram. Book review: Harry Liebersohn: The Travelers’ World: Europe to the Pacific. Isis, 98(4):850–851, December 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529312.

Navarro:2007:BRD

[3065] Jaume Navarro. Book review: David Lindley: Degrees Kelvin: A Tale of Genius, Invention, and Tragedy. Isis, 98(4):851–852, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529313.

Lawrence:2007:BRH

[3066] Susan C. Lawrence. Book review: Helen MacDonald: Human Remains: Dissection and Its Histories. Isis, 98(4):852–853, December 2007. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529314.

Decker:2007:BRB

[3067] Hannah S. Decker. Book review: Brenda Maddox: Freud’s Wizard: Ernest Jones and the Transformation of Psychoanalysis. Isis, 98(4):853– 854, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529315.

Kursell:2007:BRM

[3068] Julia Kursell. Book review: Matthias Rieger: Helmholtz Musicus: Die Objektivierung der Musik im 19. Jahrhundert durch Helmholtz’ Lehre von den Tonempfindungen. Isis, 98(4):854–855, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529316.

Buhs:2007:BRC

[3069] Joshua Blu Buhs. Book review: Charlotte Sleigh: Six Legs Better: A Cultural History of Myrmecology. Isis, 98(4):855–856, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529317.

Beyler:2007:BRG

[3070] Richard H. Beyler. Book review: Gerhard Sonnert: Einstein and Culture. Isis, 98(4):856–857, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529318.

Badino:2007:BRF

[3071] Massimiliano Badino. Book review: Friedrich Steinle: Explorative Experimente: Ampère, Faraday und die Ursprünge der Elektrodynamik. Isis, 98 (4):857–858, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529319.

Clark:2007:BRP

[3072] J. F. M. Clark. Book review: Peter Thorsheim: Inventing Pollution: Coal, Smoke, and Culture in Britain since 1800. Isis, 98(4):858–859, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529320.

Markowitz:2007:BRW

[3073] Gerald Markowitz. Book review: Werner Troesken: The Great Lead Water Pipe Disaster. Isis, 98(4):859–860, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/529321.

Worboys:2007:BRG

[3074] Michael Worboys. Book review: George N. Vlahakis; Isabel Maria Malquias; Nathan M. Brooks; François Regourd; Feza Gunergun; David Wright: Imperialism and Science: Social Impact and Interaction. Isis, 98 (4):860–861, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529322.

Lutzen:2007:BRJ

[3075] Jesper Lützen. Book review: Jean-Daniel Voelke: Renaissance de la géométrie non euclidienne entre 1860 et 1900. Isis, 98(4):861–862, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529323.

Rainger:2007:BRB

[3076] Ronald Rainger. Book review: Deep Freeze: The United States, the International Geophysical Year, and the Origins of Antarctica’s Age of Science Drift Station: Arctic Outposts of Superpower Science. Isis, 98 (4):862–864, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529324.

Lucas:2007:BRM

[3077] Matthew Lucas. Book review: Matthew J. Bellamy: Profiting the Crown: Canada’s Polymer Corporation, 1942–1990. Isis, 98(4):864–865, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529325.

Nicholson:2007:BRT

[3078] Ian Nicholson. Book review: Thomas Blass: The Man Who Shocked the World: The Life and Legacy of Stanley Milgram. Isis, 98(4):865–866, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529326.

Dyck:2007:BRR

[3079] Erika Dyck. Book review: Richard DeGrandpre: The Cult of Pharmacology: How America Became the World’s Most Troubled Drug Culture. Isis, 98(4):866–867, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529327.

Strasser:2007:BRB

[3080] Bruno J. Strasser. Book review: The Historiography of Contemporary Science, Technology, and Medicine: Writing Recent Science. Isis, 98 (4):867–868, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529328.

Snedegar:2007:BRS

[3081] Keith Snedegar. Book review: Saul Dubow: A Commonwealth of Knowledge: Science, Sensibility, and White South Africa, 1820–2000. Isis, 98 (4):868–869, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529329.

Segal:2007:BRB

[3082] Judy Z. Segal. Book review: SARS in Context: Memory, History, Policy. Isis, 98(4):870, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529330.

Clark:2007:BRD

[3083] William Clark. Book review: David John Frank; Jay Gabler: Reconstructing the University: Worldwide Shifts in Academia in the Twentieth Century. Isis, 98(4):870–871, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529331.

Weldon:2007:BRR

[3084] Stephen P. Weldon. Book review: Ronald L. Numbers: The Creationists: From Scientific Creationism to Intelligent Design. Isis, 98(4):871–873, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529332.

Davidovitch:2007:BRP

[3085] Nadav Davidovitch. Book review: Paul A. Offit: The Cutter Incident: How America’s First Polio Vaccine Led to the Growing Vaccine Crisis. Isis, 98(4):873–874, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529333.

Josephson:2007:BRE

[3086] Paul Josephson. Book review: Ethan Pollock: Stalin and the Soviet Science Wars. Isis, 98(4):874–875, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529334.

Gray:2007:BRS

[3087] Jeremy J. Gray. Book review: Siobhan Roberts: King of Infinite Space: Donald Coxeter, the Man Who Saved Geometry. Isis, 98(4):875–876, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529335.

Solovey:2007:BRJ

[3088] Mark Solovey. Book review: James A. Schellenberg: Searchers, Seers, and Shakers: Masters of Social Science. Isis, 98(4):876, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529336.

Sepkoski:2007:BRS

[3089] David Sepkoski. Book review: Seth Shulman: Undermining Science: Suppression and Distortion in the Bush Administration. Isis, 98(4):877– 878, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529337.

Choi:2007:BRJ

[3090] Hyungsub Choi. Book review: Joel N. Shurkin: Broken Genius: The Rise and Fall of William Shockley, Creator of the Electronic Age. Isis, 98(4):878–879, December 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
529338.

Hall:2007:BRG

[3091] Bert Hall. Book review: Giles Slade: Made to Break: Technology and Obsolescence in America. Isis, 98(4):879–880, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529339.

Teich:2007:BRW

[3092] Mikulás Teich. Book review: William Taussig Scott; Martin X. Moleski, S.J.: Michael Polanyi: Scientist and Philosopher. Isis, 98(4):880–881, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529340.

Johnson:2007:BRP

[3093] Stephen Johnson. Book review: Peter J. Westwick: Into the Black: JPL and the American Space Program, 1976–2004. Isis, 98(4):881–882, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529341.

Gregory:2007:BRB

[3094] Jane Gregory. Book review: Science, Technology, and Society: A Sociological Approach An Introduction to Science and Technology Studies Society and Knowledge: Contemporary Perspectives in the Sociology of Knowledge and Science. Isis, 98(4):882–884, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529342.

Healey:2007:BRT

[3095] Richard Healey. Book review: Talal A. Debs; Michael L. G. Redhead: Objectivity, Invariance, and Convention: Symmetry in Physical Science. Isis, 98(4):884–885, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529343.

Matthen:2007:BRH

[3096] Mohan Matthen. Book review: Hans Radder: The World Observed/ The World Conceived. Isis, 98(4):885–886, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529344.

Walsh:2007:BRA

[3097] D. M. Walsh. Book review: Alex Rosenberg: Darwinian Reductionism: Or, How to Stop Worrying and Love Molecular Biology. Isis, 98(4):886– 887, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529345.

Anonymous:2007:BRBa

[3098] Anonymous. Book review: Understanding Moore’s Law: Four Decades of Innovation. Isis, 98(4):887, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529346.

Anonymous:2007:BRBb

[3099] Anonymous. Book review: Removing Barriers: Women in Academic Science, Technology, Engineering, and Mathematics. Isis, 98(4):887–888, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529347.

Anonymous:2007:BRBc

[3100] Anonymous. Book review: History of Rocketry and Astronautics: Proceedings of the Thirty-Second History Symposium of the International Academy of Astronautics, Melbourne, Australia, 1998. Isis, 98(4):888, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/529348.

Anonymous:2007:BRBd

[3101] Anonymous. Book review: “Moderne” Anstaltspsychiatrie im 19, by Heiner Fangerau; Karen Nolte. und 20. Jahrhundert-Legitimation und Kritik. Isis, 98(4):888–889, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529349.

Anonymous:2007:BRBe

[3102] Anonymous. Book review: Jacopo Aconcio: Il pensiero scientifico e l’idea di tolleranza. Isis, 98(4):889, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529350.

Anonymous:2007:BRBf

[3103] Anonymous. Book review: Am Mittelpunkt der zwischen Hannover und Berlin vorfallenden Mitteilungen: Gottfried Wilhelm Leibniz in Hundisburg. Isis, 98(4):889, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/529351.

Anonymous:2007:BRBg

[3104] Anonymous. Book review: Science, Literature, and Film in the Hispanic World. Isis, 98(4):889–890, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529352.

Anonymous:2007:BRBh

[3105] Anonymous. Book review: Essays in the History of the Theory of Structures, in Honour of Jacques Heyman. Isis, 98(4):890, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529353.

Anonymous:2007:BRBi

[3106] Anonymous. Book review: Notions of Space and Time: Early Modern Concepts and Fundamental Theories. Isis, 98(4):890, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529354.

Anonymous:2007:BRBj

[3107] Anonymous. Book review: Plague and the End of Antiquity: The Pandemic of 541–750. Isis, 98(4):891, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/529355.

Anonymous:2007:BRBk

[3108] Anonymous. Book review: Mental Health and Canadian Society: Historical Perspectives. Isis, 98(4):891, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529356.

Anonymous:2007:BRBl

[3109] Anonymous. Book review: The Cambridge Companion to Early Modern Philosophy. Isis, 98(4):891, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529357.

Anonymous:2007:BRBm

[3110] Anonymous. Book review: Devices and Designs: Medical Technologies in Historical Perspective. Isis, 98(4):891–892, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/529358.

Anonymous:2007:BRBn

[3111] Anonymous. Book review: Medicating Modern America: Prescription Drugs in History. Isis, 98(4):892, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/529359.

Anonymous:2007:NCd

[3112] Anonymous. Notes on contributors. Isis, 98(4):893–897, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529276.

Anonymous:2007:IBV

[3113] Anonymous. Index to ISIS, volume 98, 2007. Isis, 98(4):899–911, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/529278.

Anonymous:2007:CV

[3114] Anonymous. Contents, volume 98. Isis, 98(4):913–916, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/529275.

Anonymous:2007:Fd

[3115] Anonymous. Frontispiece. Isis, 98(4):vii, December 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/529277.

Anonymous:2008:CBH

[3116] Anonymous. Current bibliography of the history of science and its cultural influences, 2008. Isis, 99(S1):i–320, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/597813.

Higgitt:2008:SSW

[3117] Rebekah Higgitt and Charles W. J. Withers. Science and sociability: Women as audience at the British Association for the Advancement of Science, 1831–1901. Isis, 99(1):1–27, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/587538.

Kohler:2008:FFC

[3118] Robert E. Kohler. From farm and family to career naturalist: The apprenticeship of Vernon Bailey. Isis, 99(1):28–56, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/587555.

Wellerstein:2008:PBN

[3119] Alex Wellerstein. Patenting the bomb: Nuclear weapons, intellectual property, and technological control. Isis, 99(1):57–87, March 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/587556.

Richardson:2008:SPT

[3120] Alan Richardson. Scientific philosophy as a topic for history of science. Isis, 99(1):88–96, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
587534.

Daston:2008:SO

[3121] Lorraine Daston. On scientific observation. Isis, 99(1):97–110, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/587535.

Galison:2008:TPH

[3122] Peter Galison. Ten problems in history and philosophy of science. Isis, 99(1):111–124, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
587536.

Friedman:2008:HPS

[3123] Michael Friedman. History and philosophy of science in a new key. Isis, 99(1):125–134, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
587537.

Kondratas:2008:EAB

[3124] Ramunas Kondratas. Eloge: Audrey Blyman Davis, 9 November 1933– 29 August 2006. Isis, 99(1):135–137, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589414.

Savage-Smith:2008:EFR

[3125] Emilie Savage-Smith. Eloge: Francis Romeril Maddison, 27 July 1927– 12 July 2006. Isis, 99(1):138–139, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589415.

Anonymous:2008:EBVa

[3126] Anonymous. Errata in Isis, volume 98, number 3. Isis, 99(1):140, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589418.

Rupke:2008:LE

[3127] Nicolaas A. Rupke. Letter to the Editor. Isis, 99(1):140, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589416.

Soderqvist:2008:LE

[3128] Thomas Söderqvist. Letter to the Editor. Isis, 99(1):140, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589417.

Anonymous:2008:EBVb

[3129] Anonymous. Erratum in Isis, volume 98, number 4. Isis, 99(1):140– 141, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589419.

Coen:2008:BRB

[3130] Deborah R. Coen. Book review: The Greening of German History. Isis, 99(1):142–148, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
587714.

Long:2008:BRB

[3131] Pamela O. Long. Book review: Fortress of the Soul: Violence, Metaphysics, and Material Life in the Huguenots’ New World, 1517–1751, by Neil Kamil. (Early America: History, Context, Culture.). Isis, 99(1):149– 153, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589406.

Akera:2008:BRB

[3132] Atsushi Akera. Book review: Dominance by Design: Technological Imperatives and America’s Civilizing Mission. Isis, 99(1):154–155, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589335.

Ede:2008:BRC

[3133] Andrew Ede. Book review: Clifford D. Conner: A People’s History of Science: Miners, Midwives, and “Low Mechanicks.”. Isis, 99(1):155– 156, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589337.

Stahnisch:2008:BRB

[3134] Frank W. Stahnisch. Book review: Der Geist bei der Arbeit: Historische Untersuchungen zur Hirnforschung. Isis, 99(1):157, March 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589338.

Susalla:2008:BRB

[3135] Peter J. Susalla. Book review: The History of Astronomy: A Very Short Introduction. Isis, 99(1):158, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589339.

Callender:2008:BRB

[3136] Craig Callender. Book review: Concepts of Simultaneity: From Antiquity to Einstein and Beyond. Isis, 99(1):158–160, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589340.

Vermij:2008:BRB

[3137] Rienk Vermij. Book review: Rhumb Lines and Map Wars: A Social History of the Mercator Projection. Isis, 99(1):160, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589341.

Palmieri:2008:BRB

[3138] Paolo Palmieri. Book review: Lo zufolo e la cicala: Divagazioni galileiane tra la scienza e la sua storia, by Marco Piccolino. (Saggi Scienze.). Isis, 99(1):160–161, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589342.

Martello:2008:BRB

[3139] Robert Martello. Book review: The Machine in America: A Social History of Technology. Isis, 99(1):161–162, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/589343.

Rodriguez:2008:BRB

[3140] Julia Rodriguez. Book review: Science in Latin America: A History, by Juan José Saldaña. Translated by Bernabé Madrigal. Isis, 99(1):162– 164, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589344.

Goodstein:2008:BRB

[3141] Judith R. Goodstein. Book review: Scienza a due voci, by Raffaella Simili; Valeria P. Babini; Raffaella Simili. (Biblioteca di Nuncius, 50.) More Than Pupils: Italian Women in Science at the Turn of the Twentieth Century. (Biblioteca di Nuncius, 63.). Isis, 99(1):164–165, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589345.

Bolt:2008:BRB

[3142] Marvin Bolt. Book review: Stargazer: The Life and Times of the Telescope. Isis, 99(1):165–166, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589346.

Lehoux:2008:BRBa

[3143] Daryn Lehoux. Book review: Geminos’s Introduction to the Phenomena: A Translation and Study of a Hellenistic Survey of Astronomy. Isis, 99(1):166–167, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589347.

Curtis:2008:BRB

[3144] Robert I. Curtis. Book review: A Handbook of Food Processing in Classical Rome: For Her Bounty No Winter, by David L. Thurmond. (Technology and Change in History, 9.). Isis, 99(1):167–168, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589348.

Unguru:2008:BRBa

[3145] Sabetai Unguru. Book review: Campanus of Novara and Euclid’s Elements, by H. L. L. Busard. 2 volumes. (Boethius, 51.). Isis, 99(1):168– 169, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589349.

Nuti:2008:BRB

[3146] Lucia Nuti. Book review: Fra Mauro’s World Map, by Piero Falchetta. (Terrarum Orbis: History of the Representation of Space in Text and Image, 5.). Isis, 99(1):169–170, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589350.

Walton:2008:BRBa

[3147] Steven A. Walton. Book review: Books, Banks, Buttons, and Other Inventions from the Middle Ages, by Chiara Frugoni. Translated by William McCuaig. Isis, 99(1):171, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589351.

Schettino:2008:BRB

[3148] Ernesto Schettino. Book review: La reivindicación de la filosofía en Giordano Bruno. Isis, 99(1):171–172, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589352.

Kusukawa:2008:BRB

[3149] Sachiko Kusukawa. Book review: Secrets of Women: Gender, Generation, and the Origins of Human Dissection. Isis, 99(1):172–174, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589353.

Wilding:2008:BRB

[3150] Nick Wilding. Book review: Behind the Scenes at Galileo’s Trial: Including the First English Translation of Melchior Inchofer’s Tractatus syllepticus Galileo Observed: Science and the Politics of Belief. Isis, 99(1):174–175, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589354.

Miller:2008:BRB

[3151] David Philip Miller. Book review: Joseph Banks and the British Museum: The World of Collecting, 1770–1830, by Neil Chambers. Foreword by Michael Dixon. Isis, 99(1):175–176, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/589355.

McClellan:2008:BRB

[3152] James E. McClellan, III. Book review: La bibliothèque de l’Académie royale des sciences au XVIIIe siècle. Isis, 99(1):176–177, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589356.

Hatfield:2008:BRB

[3153] Gary Hatfield. Book review: Descartes: A Biography Cogito, Ergo Sum: The Life of René Descartes. Isis, 99(1):177–178, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589357.

Machamer:2008:BRB

[3154] Peter Machamer. Book review: Descartes’s Theory of Action, by Anne Ashley Davenport. (Brill’s Studies in Intellectual History, 142.). Isis, 99(1):178–179, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589358.

McGee:2008:BRB

[3155] David McGee. Book review: Ships and Science: The Birth of Naval Architecture in the Scientific Revolution, 1600–1800. Isis, 99(1):179– 180, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589359.

Serpell:2008:BRB

[3156] James A. Serpell. Book review: Brutal Reasoning: Animals, Rationality, and Humanity in Early Modern England. Isis, 99(1):180–181, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589360.

Ariew:2008:BRB

[3157] Roger Ariew. Book review: Élie Diodati et Galilée: Naissance d’un réseau scientifique dans l’Europe du XVIIe siècle, by Stéphane Garcia. Preface by Isabelle Pantin. (Bibliothèque d’Histoire des Sciences, 6.). Isis, 99(1):181–182, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589361.

Anstey:2008:BRB

[3158] Peter Anstey. Book review: The Boyle Papers: Understanding the Manuscripts of Robert Boyle, by Michael Hunter. With contributions by Edward B. Davis, Harriet Knight, Charles Littleton, and Lawrence M. Principe. Isis, 99(1):182–183, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589362.

Luthy:2008:BRB

[3159] Christoph Lüthy. Book review: Philosophie naturelle et géométrie au XVIIe siècle, by Vincent Jullien. (Science, Techniques et Civilisations du Moyen Âge à l’Aube des Lumières). Isis, 99(1):183–184, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589363.

McTavish:2008:BRB

[3160] Lianne McTavish. Book review: Generating Bodies and Gendered Selves: The Rhetoric of Reproduction in Early Modern England. Isis, 99(1): 185, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589364.

Duchesneau:2008:BRBa

[3161] François Duchesneau. Book review: Spallanzani e le rigenerazioni animali: L’inchiesta, la communicazione, la rete. Isis, 99(1):185–187, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589365.

Eddy:2008:BRB

[3162] M. D. Eddy. Book review: The Enlightenment and the Book: Scottish Authors and Their Publishers in Eighteenth-Century Britain, Ireland, and America. Isis, 99(1):187–189, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589366.

Guerrini:2008:BRB

[3163] Anita Guerrini. Book review: Murder after Death: Literature and Anatomy in Early Modern England. Isis, 99(1):189, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589367.

Bellon:2008:BRB

[3164] Richard Bellon. Book review: Before Darwin: Reconciling God and Nature. Isis, 99(1):190–191, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589368.

Secord:2008:BRB

[3165] Anne Secord. Book review: Nature’s Engraver: A Life of Thomas Bewick. Isis, 99(1):191–192, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589369.

Breen:2008:BRB

[3166] Louise A. Breen. Book review: Dr, by Philip Cash. Benjamin Waterhouse: A Life in Medicine and Public Service (1754–1846). Isis, 99 (1):192–193, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589370.

Parshall:2008:BRB

[3167] Karen Hunger Parshall. Book review: Equations from God: Pure Mathematics and Victorian Faith. Isis, 99(1):193–194, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589371.

Pisano:2008:BRB

[3168] Dominick A. Pisano. Book review: Blind Landings: Low-Visibility Operations in American Aviation, 1918–1958. Isis, 99(1):194–196, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589372.

vanWyhe:2008:BRB

[3169] John van Wyhe. Book review: The Correspondence of Charles Darwin, by Charles Darwin; Frederick Burkhardt. Volume 15: 1867. Isis, 99 (1):196–197, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589373.

Martini:2008:BRB

[3170] Laura Martini. Book review: The Volterra Chronicles: The Life and Times of an Extraordinary Mathematician, 1860–1940, by Judith R. Goodstein. (History of Mathematics, 31.). Isis, 99(1):197–198, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589374.

Toepell:2008:BRB

[3171] Michael Toepell. Book review: Walther von Dyck (1856–1934): Mathematik, Technik und Wissenschaftsorganisation an der TH München, by Ulf Hashagen. (Boethius: Texte und Abhandlungen zur Geschichte der Mathematik und der Naturwissenschaften, 47.). Isis, 99(1):198–199, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/589375.

Weindling:2008:BRBa

[3172] Paul Weindling. Book review: The Cult of Health and Beauty in Germany: A Social History, 1890–1930. Isis, 99(1):199, March 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589376.

Stanley:2008:BRB

[3173] Matthew Stanley. Book review: Traveling at the Speed of Thought: Einstein and the Quest for Gravitational Waves. Isis, 99(1):199–200, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589377.

Moro-Abadia:2008:BRB

[3174] Oscar Moro-Abadía. Book review: Los americanistas del siglo XIX: La construcción de una comunidad científica internacional. Isis, 99(1):200– 201, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589378.

Burkhardt:2008:BRB

[3175] Richard W. Burkhardt, Jr. Book review: Science and Culture for Members Only: The Amsterdam Zoo Artis in the Nineteenth Century. Isis, 99(1):201–202, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589379.

Sutter:2008:BRBa

[3176] Paul S. Sutter. Book review: Inescapable Ecologies: A History of Environment, Disease, and Knowledge. Isis, 99(1):202–203, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589380.

Caneva:2008:BRB

[3177] Kenneth Caneva. Book review: Thermodynamik als kultureller Kampfplatz: Zur Faszinationsgeschichte der Entropie, 1850–1915. Isis, 99 (1):204–205, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589381.

Schickore:2008:BRB

[3178] Jutta Schickore. Book review: Müller’s Lab: The Story of Jakob Henle, Theodor Schwann, Emil du Bois-Reymond, Hermann von Helmholtz, Rudolf Virchow, Robert Remak, Ernst Haeckel, and Their Brilliant, Tormented Advisor. Isis, 99(1):205–206, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589382.

Sommer:2008:BRB

[3179] Marianne Sommer. Book review: Anthropologie nach Haeckel. Isis, 99 (1):206–207, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589383.

Blustein:2008:BRB

[3180] Bonnie Ellen Blustein. Book review: Nerve Endings: The Discovery of the Synapse. Isis, 99(1):207–208, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589384.

Kucuk:2008:BRB

[3181] B. Harun Küçük. Book review: Turkey’s Modernization: Refugees from Nazism and Atatürk’s Vision. Isis, 99(1):208–209, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589385.

McCook:2008:BRB

[3182] Stuart McCook. Book review: Civilizing Argentina: Science, Medicine, and the Modern State. Isis, 99(1):209–210, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/589386.

DiGregorio:2008:BRB

[3183] Mario A. Di Gregorio. Book review: Darwin and the Nature of Species. Isis, 99(1):210–211, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589387.

Chilvers:2008:BRB

[3184] C. A. J. Chilvers. Book review: Moscow Stories. Isis, 99(1):211–212, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/589388.

Nye:2008:BRBa

[3185] Mary Jo Nye. Book review: Eisenhower, Science Advice, and the Nuclear Test-Ban Debate, 1945–1963. Isis, 99(1):212–213, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589389.

Fan:2008:BRB

[3186] Fa ti Fan. Book review: Jingsheng sheng wu diao cha suo shi gao [Historical manuscript of Fan Memorial Institute of Biology], by Zonggang Hu. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu). Isis, 99(1): 214, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589390.

Herman:2008:BRB

[3187] Ellen Herman. Book review: The Averaged American: Surveys, Citizens, and the Making of a Mass Public. Isis, 99(1):214–216, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589391.

Tchalakov:2008:BRB

[3188] Ivan Tchalakov. Book review: Holographic Visions: A History of New Science. Isis, 99(1):216–217, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589392.

Needell:2008:BRB

[3189] Allan A. Needell. Book review: American Hegemony and the Postwar Reconstruction of Science in Europe, by John Krige. (Transformations: Studies in the History of Science and Technology.). Isis, 99(1):217–218, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/589393.

Shorter:2008:BRB

[3190] Edward Shorter. Book review: Pharmaceutical Reason: Knowledge and Value in Global Psychiatry, by Andrew Lakoff. (Cambridge Studies in Society and the Life Sciences.). Isis, 99(1):218–219, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589394.

Selcer:2008:BRB

[3191] Perrin Selcer. Book review: Sixty Years of Science at UNESCO, 1945– 2005. Isis, 99(1):219–220, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589395.

Epstein:2008:BRB

[3192] Steven Epstein. Book review: When Illness Goes Public: Celebrity Patients and How We Look at Medicine. Isis, 99(1):220–222, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589396.

Rheinberger:2008:BRB

[3193] Hans-Jörg Rheinberger. Book review: Das Labor in der Box: Technikentwicklung und Unternehmensgründung in der frühen deutschen Biotechnologie Biotech: The Countercultural Origins of an Industry. Isis, 99 (1):222–223, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589397.

Adas:2008:BRB

[3194] Michael Adas. Book review: Transforming the Twentieth Century: Technical Innovations and Their Consequences. Isis, 99(1):223–225, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589398.

November:2008:BRB

[3195] Joseph November. Book review: Biocapital: The Constitution of Postgenomic Life. Isis, 99(1):225–227, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589399.

vanDelft:2008:BRB

[3196] Dirk van Delft. Book review: The Whipple Museum of the History of Science: Instruments and Interpretations, to Celebrate the Sixtieth Anniversary of R, by Liba Taub; Frances Willmoth. S. Whipple’s Gift to the University of Cambridge. Isis, 99(1):227–228, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/589400.

Wang:2008:BRBa

[3197] Zuoyue Wang. Book review: Oppenheimer: The Tragic Intellect. Isis, 99(1):228–229, March 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
589401.

Morange:2008:BRB

[3198] Michel Morange. Book review: The Troubled Dream of Genetic Medicine: Ethnicity and Innovation in Tay-Sachs, Cystic Fibrosis, and Sickle Cell Disease. Isis, 99(1):229–230, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/589402.

Sengoopta:2008:BRB

[3199] Chandak Sengoopta. Book review: The Estrogen Elixir: A History of Hormone Replacement Therapy in America. Isis, 99(1):230–231, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/589403.

Schmalzer:2008:BRB

[3200] Sigrid Schmalzer. Book review: He cheng yi ge dan bai zhi: Jie jing niu yi dao su de ren gong quan he cheng [ Synthesize a protein: The story of total synthesis of crystalline insulin project in China ], by Weimin Xiong; Kedi Wang. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(1):231–232, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/589404.

Fisher:2008:BRB

[3201] Erik Fisher. Book review: The Philosophy of Expertise. Isis, 99(1):232– 233, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/589405.

Anonymous:2008:NCa

[3202] Anonymous. Notes on contributors. Isis, 99(1):234–238, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/589422.

Anonymous:2008:Fa

[3203] Anonymous. Frontispiece. Isis, 99(1):vii, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/589421.

Creager:2008:ADH

[3204] Angela N. H. Creager and Gregory J. Morgan. After the double helix. Isis, 99(2):239–272, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
588626.

Vetter:2008:CSF

[3205] Jeremy Vetter. Cowboys, scientists, and fossils. Isis, 99(2):273–303, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/588688.

Pesic:2008:PR

[3206] Peter Pesic. Proteus rebound. Isis, 99(2):304–317, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/588627.

Maienschein:2008:WDD

[3207] Jane Maienschein and George Smith. What difference does history of science make, anyway? Isis, 99(2):318–321, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/588689.

Gooday:2008:DSE

[3208] Graeme Gooday, John M. Lynch, Kenneth G. Wilson, and Constance K. Barsky. Does science education need the history of science? Isis, 99(2): 322–330, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/588690.

Hamilton:2008:TWH

[3209] Andrew Hamilton and Quentin D. Wheeler. Taxonomy and why history of science matters for science. Isis, 99(2):331–340, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/588691.

Maienschein:2008:HCH

[3210] Jane Maienschein, Manfred Laubichler, and Andrea Loettgers. How can history of science matter to scientists? Isis, 99(2):341–349, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/588692.

Pandora:2008:SEW

[3211] Katherine Pandora and Karen A. Rader. Science in the everyday world. Isis, 99(2):350–364, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
588693.

Wang:2008:HSA

[3212] Zuoyue Wang and Naomi Oreskes. History of science and American science policy. Isis, 99(2):365–373, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/588694.

Walter:2008:LE

[3213] Scott Walter. Letter to the Editor. Isis, 99(2):374, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/588695.

Gingras:2008:LE

[3214] Yves Gingras. Letter to the Editor. Isis, 99(2):374–375, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591388.

Appel:2008:BRB

[3215] Toby A. Appel. Book review: Dictionary of Medical Biography. Isis, 99(2):376–377, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591319.

Zenderland:2008:BRB

[3216] Leila Zenderland. Book review: The Measure of Merit: Talents, Intelligence, and Inequality in the French and American Republics, 1750–1940. Isis, 99(2):377–378, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591320.

Joy:2008:BRB

[3217] Lynn S. Joy. Book review: The Intelligibility of Nature: How Science Makes Sense of the World. Isis, 99(2):378–379, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591321.

Ogilvie:2008:BRB

[3218] Brian W. Ogilvie. Book review: Curiosity and Wonder from the Renaissance to the Enlightenment. Isis, 99(2):379–380, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591322.

Li:2008:BRB

[3219] Shang-Jen Li. Book review: Jin dai xi fang shi Hua sheng wu shi [History of Western Botanical and Zoological Studies in China], by Guihan Luo. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(2): 380–381, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591325.

Bocking:2008:BRB

[3220] Stephen Bocking. Book review: Breathing Space: How Allergies Shape Our Lives and Landscapes. Isis, 99(2):381–382, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591326.

Cushman:2008:BRB

[3221] Gregory T. Cushman. Book review: Dirt: The Erosion of Civilizations. Isis, 99(2):382–383, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591327.

Borck:2008:BRB

[3222] Cornelius Borck. Book review: Visual Cultures of Science: Rethinking Representational Practices in Knowledge Building and Science Communication, by Luc Pauwels. (Interfaces: Studies in Visual Culture.). Isis, 99(2):383–384, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591328.

Sivasundaram:2008:BRB

[3223] Sujit Sivasundaram. Book review: Relocating Modern Science: Circulation and the Construction of Knowledge in South Asia and Europe, 1650– 1900. Isis, 99(2):384–385, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/591329.

McIver:2008:BRB

[3224] Tom McIver. Book review: Evolution vs, by Eugenie C. Scott; Niles Eldredge. Creationism: An Introduction. Isis, 99(2):385–386, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591330.

Rensing:2008:BRB

[3225] Susan Rensing. Book review: Women and Science: Social Impact and Interaction, by Suzanne Le-May Sheffield. (Science and Society.). Isis, 99(2):386–387, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591331.

Henry:2008:BRB

[3226] John Henry. Book review: The History and Poetics of Scientific Biography, by Thomas Söderqvist. (Science, Technology, and Culture, 1700– 1945.). Isis, 99(2):387–388, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/591332.

Dym:2008:BRBa

[3227] Warren Alexander Dym. Book review: Chemie in Österreich: Bergbau, Alchemie, und frühe Chemie: Von den Anfängen bis zum Ende des 18, by Rudolf Werner Soukup. Jahrhunderts. (Beiträge zur Wissenschaftsgeschichte und Wissenschaftsforschung, 7.). Isis, 99(2):388–389, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/591333.

Shindell:2008:BRB

[3228] Matthew Shindell. Book review: Is Pluto a Planet? A Historical Journey through the Solar System. Isis, 99(2):389–391, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591334.

Kroker:2008:BRB

[3229] Kenton Kroker. Book review: Sleep and Society: Sociological Ventures into the (Un)known. Isis, 99(2):391–392, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/591335.

Trepanier:2008:BRB

[3230] Simon Trépanier. Book review: Explaining the Cosmos: The Ionian Tradition of Scientific Philosophy. Isis, 99(2):392–393, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591336.

Clouse:2008:BRB

[3231] Michele L. Clouse. Book review: New Philosophy of Human Nature: Neither Known to nor Attained by the Great Ancient Philosophers, Which Will Improve Human Life and Health. Isis, 99(2):393–394, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591337.

Montgomery:2008:BRB

[3232] Scott L. Montgomery. Book review: Observation and Image-Making in Gothic Art Visualizing Medieval Medicine and Natural History, 1200– 1550, by Jean A. Givens; Jean A. Givens; Karen M. Reeds; Alain Touwaide. (AVISTA Studies in the History of Medieval Technology, Science, and Art.). Isis, 99(2):394–395, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591339.

Gouk:2008:BRB

[3233] Penelope Gouk. Book review: The Jewel House: Elizabethan London and the Scientific Revolution. Isis, 99(2):395–397, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591340.

Broomhall:2008:BRB

[3234] Susan Broomhall. Book review: Hermaphrodites in Renaissance Europe, by Kathleen P. Long. (Women and Gender in the Early Modern World.). Isis, 99(2):397–398, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591341.

Barker:2008:BRB

[3235] Peter Barker. Book review: Bearing the Heavens: Tycho Brahe and the Astronomical Community of the Late Sixteenth Century. Isis, 99(2):398– 399, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591342.

Akasoy:2008:BRB

[3236] Anna Akasoy. Book review: Islamic Science and the Making of the European Renaissance. Isis, 99(2):399–400, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591343.

Ducheyne:2008:BRB

[3237] Steffen Ducheyne. Book review: Huygens: The Man behind the Principle. Isis, 99(2):400–401, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591344.

Scharf:2008:BRB

[3238] Sara Scharf. Book review: Linnaeus in Italy: The Spread of a Revolution in Science, by Marco Beretta; Alessandro Tosi. (Uppsala Studies in History of Science, 34.). Isis, 99(2):401–402, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591345.

Shea:2008:BRBa

[3239] William R. Shea. Book review: Galilei der Künstler: Der Mond-Die Sonne-Die Hand. Isis, 99(2):402–403, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591346.

Benzaquen:2008:BRB

[3240] Adriana S. Benzaquén. Book review: De l’homme. Isis, 99(2):403– 404, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591347.

Gascoigne:2008:BRB

[3241] John Gascoigne. Book review: The Scientific Correspondence of Sir Joseph Banks, 1765–1820, by Neil Chambers. Volumes 1–6. (Pickering Masters.). Isis, 99(2):404–406, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591348.

Fissell:2008:BRB

[3242] Mary E. Fissell. Book review: Hubbub: Filth, Noise, and Stench in England, 1600–1770. Isis, 99(2):406–407, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591349.

Bleichmar:2008:BRB

[3243] Daniela Bleichmar. Book review: Nature, Empire, and Nation: Explorations of the History of Science in the Iberian World. Isis, 99(2):407– 408, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591350.

Roberts:2008:BRBa

[3244] Lissa Roberts. Book review: Tulipmania: Money, Honor, and Knowledge in the Dutch Golden Age. Isis, 99(2):408–409, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591351.

Snobelen:2008:BRB

[3245] Stephen D. Snobelen. Book review: The Early Biographies of Isaac Newton, 1660–1885. Isis, 99(2):409–411, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591352.

Sylla:2008:BRB

[3246] Edith Dudley Sylla. Book review: Thinking with Objects: The Transformation of Mechanics in the Seventeenth Century. Isis, 99(2):411–412, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/591353.

Cook:2008:BRB

[3247] Harold J. Cook. Book review: Indian Ink: Script and Print in the Making of the English East India Company. Isis, 99(2):412–413, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591354.

Duchesneau:2008:BRBb

[3248] François Duchesneau. Book review: Les forces vitales et leur distribution dans la nature: Un essai de ”systématique physiologique, by Stéphane Schmitt.” (De Diversis Artibus.). Isis, 99(2):413–414, June 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591355.

Stewart:2008:BRBa

[3249] Larry Stewart. Book review: Placing the Enlightenment: Thinking Geographically about the Age of Reason. Isis, 99(2):414–415, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591356.

Brundell:2008:BRB

[3250] Barry Brundell. Book review: Negotiating Darwin: The Vatican Confronts Evolution, 1877–1902. Isis, 99(2):415–416, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591357.

Efron:2008:BRB

[3251] Noah Efron. Book review: Jewish Tradition and the Challenge of Darwinism. Isis, 99(2):416–418, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/591358.

Harrison:2008:BRB

[3252] Mark Harrison. Book review: Medical Lives in the Age of Surgical Revolution, by M. Anne Crowther; Marguerite W. Dupree. (Cambridge Studies in Population, Economy, and Society in Past Time.). Isis, 99(2):418– 419, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591359.

Rudwick:2008:BRB

[3253] Martin Rudwick. Book review: Whatever Is under the Earth: The Geological Society of London, 1807 to 2007. Isis, 99(2):419–420, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591360.

Brain:2008:BRB

[3254] Robert Michael Brain. Book review: Wissenschaft in der Machinenstadt: Emil Du Bois-Reymond und seine Laboratorien in Berlin. Isis, 99(2): 420–422, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591361.

Good:2008:BRBa

[3255] Greg Good. Book review: The Callendar Effect: The Life and Work of Guy Stewart Callendar (1898–1964), the Scientist Who Established the Carbon Dioxide Theory of Climate Change. Isis, 99(2):422–423, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/591362.

Elwick:2008:BRB

[3256] James Elwick. Book review: Herbert Spencer and the Invention of Modern Life. Isis, 99(2):423–424, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591363.

Marchisotto:2008:BRB

[3257] Elena Anne Corie Marchisotto. Book review: Episodes in the History of Modern Algebra (1800–1950), by Jeremy J. Gray; Karen Hunger Parshall. (History of Mathematics, 32.). Isis, 99(2):424–425, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591364.

Topham:2008:BRB

[3258] Jonathan R. Topham. Book review: Jeremiah Joyce: Radical, Dissenter, and Writer, by John Issitt. (Science, Technology, and Culture, 1700– 1945.). Isis, 99(2):426, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/591365.

Hentschel:2008:BRBa

[3259] Klaus Hentschel. Book review: Harmonious Triads: Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany, by Myles W. Jackson. (Transformations: Studies in the History of Science and Technology.). Isis, 99(2):426–428, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591366.

Stein:2008:BRB

[3260] Claudia Stein. Book review: Klio und Hippokrates: Eine Liaison littéraire des 18, by Hans-Uwe Lammel. Jahrhunderts und die Folgen für die Wissenschaftskultur bis 1850 in Deutschland. (Sudhoffs Archiv, 55.). Isis, 99(2):428–429, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/591367.

Chung:2008:BRB

[3261] Juliette Yuehtsen Chung. Book review: Ji shu yu di guo yi yan jiu: riben zai Zhongguo de zhi min ke yan ji gou [Researches on Technology and Imperialism: Japanese Colonial Scientific Research Institutes in China], by Bo Liang; Jianping Han; Xingsui Cao; Liwei Wu. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.) Ri wei shi qi de zhi min di ke yan ji gou: li shi yu wen xian [Colonial Scientific Institutions during the Japanese Occupation and Puppet Manchukuo Period: History and Literature]. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(2):429–430, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591369.

Lewis:2008:BRB

[3262] John W. Lewis and Xue Litai. Book review: Liang dan yi xing gong cheng yu da ke xue [The Project of “Two Bombs, One Satellite”: A Model of the Big Science], by Jifeng Liu; Yanqiong Liu; Haiyan Xie. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(2):430–431, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/591370.

Stern:2008:BRB

[3263] Alexandra Minna Stern. Book review: Conceiving the Future: Pronatalism, Reproduction, and the Family in the United States, 1890–1938. Isis, 99(2):431–432, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591371.

Pyenson:2008:BRB

[3264] Lewis Pyenson. Book review: Einstein on Politics: His Private Thoughts and Public Stands on Nationalism, Zionism, War, Peace, and the Bomb. Isis, 99(2):432–434, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591372.

Enebakk:2008:BRB

[3265] Vidar Enebakk. Book review: Perspectives on Scandinavian Science in the Early Twentieth Century. Isis, 99(2):434–435, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591373.

VanRiper:2008:BRB

[3266] A. Bowdoin Van Riper. Book review: Science Talk: Changing Notions of Science in American Popular Culture. Isis, 99(2):435–436, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591374.

Chazan:2008:BRB

[3267] Michael Chazan. Book review: L’affaire de l’abri du poisson: Patrie et préhistoire. Isis, 99(2):436–437, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591375.

Wang:2008:BRBb

[3268] Zuoyue Wang. Book review: Ke xue she tuan zai jin dai Zhongguo de ming yun: yi Zhongguo ke xue she wei zhong xin [The Science Association and the Change of Society in Modern China: A Study on the Science Society of China], by Jian Zhang. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(2):437–438, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591376.

Anderson:2008:BRB

[3269] Katharine Anderson. Book review: Extremes: Oceanography’s Adventures at the Poles. Isis, 99(2):438–439, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591377.

Young:2008:BRB

[3270] Chris Young. Book review: Barry Commoner and the Science of Survival: The Remaking of American Environmentalism. Isis, 99(2):439– 440, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591378.

Fox:2008:BRB

[3271] Robert Fox. Book review: The History of Imperial College London, 1907–2007: Higher Education and Research in Science, Technology, and Medicine. Isis, 99(2):440–441, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591379.

Macrakis:2008:BRB

[3272] Kristie Macrakis. Book review: The Star Wars Enigma: Behind the Scenes of the Cold War Race for Missile Defense. Isis, 99(2):441–442, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/591380.

Seidel:2008:BRB

[3273] Robert W. Seidel. Book review: Proving Grounds: Project Plowshare and the Unrealized Dream of Nuclear Earthmoving. Isis, 99(2):442–443, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/591381.

Hughes:2008:BRB

[3274] Jeff Hughes. Book review: Nylon and Bombs: DuPont and the March of Modern America. Isis, 99(2):443–444, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591382.

Chu:2008:BRB

[3275] Pingyi Chu. Book review: Zhongguo jin xian dai ke ji jiang li zhi du [Science and Technology Awards in Modern China], by Anjing Qu. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(2):444–445, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/591383.

Brooke:2008:BRB

[3276] John Hedley Brooke. Book review: Eminent Lives in Twentieth–Century Science and Religion. Isis, 99(2):445–446, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/591384.

Zhu:2008:BRB

[3277] Yuelin Zhu. Book review: Xin Zhongguo yu xin ke xue: gao fen zi xue zai xian dai Zhongguo de jian li [New Science for a New China: Institutionalization of Polymer Science in the P R. China], by Li Zhang. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu). Isis, 99(2): 446–447, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591385.

Richardson:2008:BRB

[3278] Alan Richardson. Book review: Value-Free Science? Ideals and Illusions. Isis, 99(2):448, June 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591386.

Anonymous:2008:NCb

[3279] Anonymous. Notes on contributors. Isis, 99(2):449–453, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/588696.

Anonymous:2008:Fb

[3280] Anonymous. Frontispiece. Isis, 99(2):vi, June 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/591387.

Jones:2008:ATP

[3281] Susan D. Jones and Philip M. Teigen. Anthrax in transit: Practical experience and intellectual exchange. Isis, 99(3):455–485, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/591709.

Krementsov:2008:HBO

[3282] Nikolai Krementsov. Hormones and the Bolsheviks: From organotherapy to experimental endocrinology, 1918–1929. Isis, 99(3):486–518, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591710.

Goldstein:2008:OSK

[3283] Daniel Goldstein. Outposts of science: The knowledge trade and the expansion of scientific community in post–civil war America. Isis, 99 (3):519–546, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591711.

Radick:2008:IWW

[3284] Gregory Radick. Introduction: Why what if? Isis, 99(3):547–551, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/591712.

Henry:2008:IIS

[3285] John Henry. Ideology, inevitability, and the scientific revolution. Isis, 99 (3):552–559, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
591713.

Bowler:2008:WDD

[3286] Peter J. Bowler. What Darwin disturbed: The biology that might have been. Isis, 99(3):560–567, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591714.

French:2008:GPS

[3287] Steven French. Genuine possibilities in the scientific past and how to spot them. Isis, 99(3):568–575, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/591715.

Fuller:2008:NTC

[3288] Steve Fuller. The normative turn: Counterfactuals and a philosophical historiography of science. Isis, 99(3):576–584, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/591716.

Davenport:2008:LE

[3289] Anne Davenport and Peter Machamer. Letter to the Editor. Isis, 99 (3):585, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593286.

Broks:2008:LE

[3290] Peter Broks. Letter to the Editor. Isis, 99(3):585–586, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593287.

Enebakk:2008:LE

[3291] Vidar Enebakk. Letter to the Editor. Isis, 99(3):586, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593288.

Hughes:2008:LE

[3292] A. M. Hughes. Letter to the Editor. Isis, 99(3):586, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593289.

Carson:2008:BRB

[3293] John Carson. Book review: The Post-Revolutionary Self: Politics and Psyche in France, 1750–1850. Isis, 99(3):587–591, September 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591717.

Pursell:2008:BRB

[3294] Carroll Pursell. Book review: The Lie Detectors: The History of an American Obsession. Isis, 99(3):592–593, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593218.

Croce:2008:BRB

[3295] Paul Jerome Croce. Book review: Monkey Trials and Gorilla Sermons: Evolution and Christianity from Darwin to Intelligent Design. Isis, 99 (3):593–594, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593219.

Lawrence:2008:BRB

[3296] Christopher Lawrence. Book review: Penicillin: Triumph and Tragedy. Isis, 99(3):594–595, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593220.

Waddell:2008:BRB

[3297] Mark A. Waddell. Book review: The Secret History of Hermes Trismegistus: Hermeticism from Ancient to Modern Times, by Florian Ebeling. Translated by David Lorton. Foreword by Jan Assmann. Isis, 99 (3):595–596, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593221.

Ramberg:2008:BRB

[3298] Peter J. Ramberg. Book review: Fermentation: Vital or Chemical Process? (History of Science and Medicine Library, 1), by Joseph S. Fruton. Isis, 99(3):596–597, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593222.

Lefevre:2008:BRBa

[3299] Wolfgang Lefèvre. Book review: The Emergence of a Scientific Culture: Science and the Shaping of Modernity, 1210–1685. Isis, 99(3):597–599, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593223.

Evans:2008:BRB

[3300] James Evans. Book review: Revolutionaries of the Cosmos: The Astro-Physicists Conceptions of Cosmos: From Myths to the Accelerating Universe: A History of Cosmology. Isis, 99(3):599–600, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593224.

Turner:2008:BRBa

[3301] Frank M. Turner. Book review: A History of Natural Philosophy from the Ancient World to the Nineteenth Century. Isis, 99(3):600–601, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/593225.

Larson:2008:BRB

[3302] Edward J. Larson. Book review: Breeding Contempt: The History of Coerced Sterilization in the United States. Isis, 99(3):601–602, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/593226.

Churchill:2008:BRB

[3303] Frederick B. Churchill. Book review: Heredity Produced: At the Crossroads of Biology, Politics, and Culture, 1500–1870. Isis, 99(3):602–604, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593227.

Dym:2008:BRBb

[3304] Warren Alexander Dym. Book review: Chymists and Chymistry: Studies in the History of Alchemy and Early Modern Chemistry. Isis, 99(3): 604–605, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593228.

Selya:2008:BRB

[3305] Rena Selya. Book review: The Poetics of DNA, by Judith Roof. (PostHumanities, 2.). Isis, 99(3):605–606, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593229.

Xu:2008:BRB

[3306] Yibao Xu. Book review: Zhongguo shuxue de xihua licheng [The Westernization of Mathematics in China] Zhongguo jindai shuxue jiaoyu shigao [A Draft History of Mathematics Education in the Late Qing Dynasty] Zhongguo jindai daishu shi jianbian [A Short History of Algebra in Modern China]. Isis, 99(3):606–608, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593230.

McCluskey:2008:BRB

[3307] Stephen McCluskey. Book review: The Apotheosis of Janaab’ Pakal: Science, History, and Religion at Classic Maya Palenque. Isis, 99(3): 608–609, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593231.

Burns:2008:BRB

[3308] Elizabeth Burns. Book review: Hypatia of Alexandria: Mathematician and Martyr. Isis, 99(3):609–610, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/593232.

Lehoux:2008:BRBb

[3309] Daryn Lehoux. Book review: Theophrastus of Eresus: On Weather Signs. Isis, 99(3):610–611, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593233.

Worthen:2008:BRB

[3310] Shana Worthen. Book review: L’ordre du temps: L’invention de la ponctualité au XVIe siècle. Isis, 99(3):611–612, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593234.

Hirai:2008:BRB

[3311] Hiro Hirai. Book review: Alchimie et paracelsisme en France à la fin de la Renaissance (1567–1625), by Didier Kahn. (Cahiers d’Humanisme et Renaissance, 80.). Isis, 99(3):612–613, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593235.

Kahana-Smilansky:2008:BRB

[3312] Hagar Kahana-Smilansky. Book review: Medical Aphorisms, Treatises 6–9: A Parallel Arabic-English Edition, by Maimonides. Translated by Gerrit Bos. (Complete Medical Works of Maimonides, 2.). Isis, 99(3): 613–614, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593236.

Pumfrey:2008:BRB

[3313] Stephen Pumfrey. Book review: The English Renaissance Stage: Geometry, Poetics, and the Practical Spatial Arts, 1580–1630. Isis, 99(3): 614–615, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593237.

Bates:2008:BRB

[3314] David Bates. Book review: L’Encyclopédie méthodique: Des Lumières au positivisme (1782–1832). Isis, 99(3):615–616, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593238.

Fraser:2008:BRB

[3315] Craig Fraser. Book review: Euler at Three Hundred: An Appreciation, by Robert E. Bradley; Lawrence A. D’Antonio; C. Edward Sandifer. (Spectrum Series.). Isis, 99(3):616–617, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593239.

Dijksterhuis:2008:BRB

[3316] Fokko Jan Dijksterhuis. Book review: La philosophie naturelle de Christiaan Huygens, by Fabien Chareix. (Mathesis.). Isis, 99(3):617–618, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593240.

Kahn:2008:BRBa

[3317] Didier Kahn. Book review: The Chemical Promise: Experiment and Mysticism in the Chemical Philosophy, 1550–1800. Isis, 99(3):618–619, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593241.

Malone:2008:BRB

[3318] Robert J. Malone. Book review: William Dunbar: Scientific Pioneer of the Old Southwest. Isis, 99(3):619–620, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593242.

Bandinelli:2008:BRB

[3319] Angela Bandinelli. Book review: Storia, scienza e società: Ricerche sulla scienza in Italia nell’età moderna e contemporanea, by Paola Govoni. (Bologna Studies in History of Science, 11.). Isis, 99(3):620–621, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593243.

Somsen:2008:BRB

[3320] Geert Somsen. Book review: Les racines sociales et économiques des Principia de Newton: Une rencontre entre Newton et Marx à Londres en 1931, by Boris Hessen. Translated with commentary by Serge Guérout. Postscript by Christopher Chilvers. Isis, 99(3):621–622, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593244.

Staum:2008:BRB

[3321] Martin S. Staum. Book review: L’avènement de la médecine clinique moderne en Europe, 1750–1815: Politiques, institutions et savoirs, by Othmar Keel. Bibiliothèque d’Histoire de la Médicine et de la Santé. Isis, 99(3):622–623, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593245.

Roberts:2008:BRBb

[3322] Lissa Roberts. Book review: Materials in Eighteenth-Century Science: A Historical Ontology, by Ursula Klein; Wolfgang Lefèvre. (Transformations: Studies in the History of Science and Technology.). Isis, 99 (3):623–624, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593246.

Gronim:2008:BRB

[3323] Sara Stidstone Gronim. Book review: The Art and Science of William Bartram. Isis, 99(3):625, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/593247.

Bennett:2008:BRB

[3324] Jim Bennett. Book review: Making Scientific Instruments in the Industrial Revolution, by A. D. Morrison-Low. Foreword by Ludmilla Jordanova. (Science, Technology, and Culture, 1700–1945.). Isis, 99(3): 625–626, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593248.

Clark:2008:BRBa

[3325] William Clark. Book review: Melancholy and the Care of the Soul: Religion, Moral Philosophy, and Madness in Early Modern England, by Jeremy Schmidt. (History of Medicine in Context.). Isis, 99(3):626–627, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593249.

Cozzoli:2008:BRB

[3326] Daniele Cozzoli. Book review: Verità matematiche e forme della natura da Galileo a Newton. Isis, 99(3):627–628, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593250.

Olshin:2008:BRB

[3327] Benjamin B. Olshin. Book review: Cidadão do mundo: Uma biografia científica do Abade Correia da Serra. Isis, 99(3):628–629, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/593251.

Choi:2008:BRB

[3328] Tina Young Choi. Book review: Poison, Detection, and the Victorian Imagination, by Ian Burney. (Encounters: Cultural Histories.). Isis, 99 (3):629–631, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593252.

Buklijas:2008:BRB

[3329] Tatjana Buklijas. Book review: Vienna in the Age of Uncertainty: Science, Liberalism, and Private Life. Isis, 99(3):631–632, September 2008.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593253.

Kursell:2008:BRB

[3330] Julia Kursell. Book review: Elektrische Klangmaschinen: Die Pionierzeit in Deutschland und Österreich. Isis, 99(3):632–633, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593254.

Hentschel:2008:BRBb

[3331] Klaus Hentschel. Book review: Die Welt als Raum und Zeit, by Alexander Friedmann. Edited and translated by Georg Singer. (Ostwalds Klassiker der exakten Wissenschaften, 287.). Isis, 99(3):633, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593255.

Sheffield:2008:BRB

[3332] Suzanne Le-May Sheffield. Book review: “Good Observers of Nature”: American Women and the Scientific Study of the Natural World, 1820– 1885. Isis, 99(3):634, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593256.

Shen:2008:BRBa

[3333] Grace Y. Shen. Book review: Di zhi xue yu Minguo she hui: 1916–1950 [Geology and Society: A Study in Chinese National Geological Survey], by Zhang Jiuchen. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(3):634–635, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593257.

Mahone:2008:BRB

[3334] Sloan Mahone. Book review: Colonial Madness: Psychiatry in French North Africa. Isis, 99(3):635–636, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/593258.

Burstyn:2008:BRB

[3335] Harold L. Burstyn. Book review: The Emergence of Numerical Weather Prediction: Richardson’s Dream. Isis, 99(3):636–637, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593259.

Elkind:2008:BRB

[3336] Sarah S. Elkind. Book review: The Landscape of Reform: Civic Pragmatism and Environmental Thought in America. Isis, 99(3):637–638, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593260.

Josephson:2008:BRB

[3337] Paul Josephson. Book review: Dictatorship of the Air: Aviation Culture and the Fate of Modern Russia. Isis, 99(3):639, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593261.

Clark:2008:BRBb

[3338] J. F. M. Clark. Book review: Science, Politics, and Business in the Work of Sir John Lubbock: A Man of Universal Mind, by Mark Patton. (Science, Technology, and Culture, 1700–1945.). Isis, 99(3):639–640, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/593262.

Friedman:2008:BRB

[3339] Robert Marc Friedman. Book review: The Coldest Crucible: Arctic Exploration and American Culture. Isis, 99(3):641, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593263.

Elliott:2008:BRB

[3340] Clark A. Elliott. Book review: The Papers of Joseph Henry, by Marc Rothenberg; Kathleen W. Dorman; Frank R. Millikan; Deborah Y. Jeffries; Sarah Schoenfeld; Marc Rothenberg; Kathleen W. Dorman; Frank R. Millikan; Deborah Y. Jeffries; Sarah Schoenfeld. Volume 10: January 1858–December 1865: The Smithsonian Years The Papers of Joseph Henry. Volume 11: January 1866–May 1878: The Smithsonian Years. Isis, 99(3):641–643, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593264.

Cravens:2008:BRB

[3341] Hamilton Cravens. Book review: Savage Perils: Racial Frontiers and Nuclear Apocalypse in American Culture. Isis, 99(3):643–644, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/593265.

Ash:2008:BRB

[3342] Mitchell G. Ash. Book review: “Blood and Homeland”: Eugenics and Racial Nationalism in Central and Southeast Europe, 1900–1940. Isis, 99 (3):644–645, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593266.

Fang:2008:BRB

[3343] Li-Zhi Fang. Book review: Zijin shan tian wen tai shi gao: Zhongguo tian wen xue xian dai hua ge an, by Jiang Xiaoyuan; Wu Yan. [History of Purplemountain Observatory.] (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 99(3):645–646, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593267.

Agar:2008:BRB

[3344] Jon Agar. Book review: Calculating a Natural World: Scientists, Engineers, and Computers during the Rise of U, by Atsushi Akera.S. Cold War Research. Isis, 99(3):646–647, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/593268.

Fedunkiw:2008:BRB

[3345] Marianne P. Fedunkiw. Book review: A Generation of Excellence: A History of the Canadian Institute for Advanced Research. Isis, 99(3): 647–649, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593269.

Cambrosio:2008:BRB

[3346] Alberto Cambrosio. Book review: Inclusion: The Politics of Difference in Medical Research. Isis, 99(3):649–650, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593270.

Meldrum:2008:BRB

[3347] Marcia Meldrum. Book review: Prescribing by Numbers: Drugs and the Definition of Disease. Isis, 99(3):650–652, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593271.

Schweber:2008:BRBa

[3348] Libby Schweber. Book review: Le point de vue du nombre, 1936: Précédé de l’avant-propos au Tome VII de l’Encyclopédie française de Lucien Febvre et suivi de trois articles de Maurice Halbwachs. Isis, 99(3): 652–653, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593272.

DSouza:2008:BRB

[3349] Rohan D’Souza. Book review: Triumph of the Expert: Agrarian Doctrines of Development and the Legacies of British Colonialism. Isis, 99 (3):653–654, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593273.

Weitekamp:2008:BRB

[3350] Margaret A. Weitekamp. Book review: A History of the Kennedy Space Center. Isis, 99(3):654–655, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/593274.

Nakayama:2008:BRB

[3351] Shigeru Nakayama. Book review: Japan on Display: Photography and the Emperor, by Morris Low. (Routledge/Asian Studies Association of Australia East Asia Series.). Isis, 99(3):655–656, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/593275.

Westwick:2008:BRB

[3352] Peter Westwick. Book review: The Nuclear Borderlands: The Manhattan Project in Post-Cold War New Mexico. Isis, 99(3):656–658, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/593276.

Doel:2008:BRB

[3353] Ronald E. Doel. Book review: The Honest Broker: Making Sense of Science in Policy and Politics. Isis, 99(3):658–659, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593277.

Nye:2008:BRBb

[3354] David E. Nye. Book review: Technology in Postwar America: A History. Isis, 99(3):659–660, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/593278.

Sloan:2008:BRB

[3355] Phillip R. Sloan. Book review: La fabrique d’une nouvelle science: La biologie moléculaire à l’âge atomique (1945–1964), by Bruno J. Strasser. (Bibliothèque d’Histoire des Sciences, 8.). Isis, 99(3):660–661, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/593279.

Gooding:2008:BRB

[3356] David Gooding. Book review: The Cognitive Structure of Scientific Revolutions. Isis, 99(3):661–662, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/593280.

Chimisso:2008:BRB

[3357] Cristina Chimisso. Book review: Time and the Shape of History. Isis, 99 (3):662–664, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593281.

Borrello:2008:BRB

[3358] Mark E. Borrello. Book review: Science without Laws: Model Systems, Cases, and Exemplary Narratives. Isis, 99(3):664–665, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593282.

Doing:2008:BRB

[3359] Park Doing. Book review: Inside the Politics of Technology: Agency and Normativity in the Co-Production of Technology and Society. Isis, 99 (3):665–666, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593283.

Maas:2008:BRB

[3360] Harro Maas. Book review: Equilibrium in Economics: Scope and Limits, by Valeria Mosini. (Routledge Frontiers of Political Economy.). Isis, 99 (3):666–667, September 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
593284.

Gingras:2008:BRB

[3361] Yves Gingras. Book review: Controverses sur la science: Pour une sociologie transversaliste de l’activité scientifique, by Terry Shinn; Pascal Ragouet. (Cours et Travaux.). Isis, 99(3):667–668, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/593285.

Anonymous:2008:NCc

[3362] Anonymous. Notes on contributors. Isis, 99(3):669–673, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/591719.

Anonymous:2008:Fc

[3363] Anonymous. Frontispiece. Isis, 99(3):vi, September 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/593290.

Adkins:2008:RPA

[3364] G. Matthew Adkins. The Renaissance of peiresc: Aubin-louis millin and the postrevolutionary Republic of letters. Isis, 99(4):675–700, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/595766.

Elshakry:2008:KMC

[3365] Marwa S. Elshakry. Knowledge in motion: The cultural politics of modern science translations in Arabic. Isis, 99(4):701–730, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/595767.

Merchant:2008:VIF

[3366] Carolyn Merchant. “The violence of impediments”: Francis Bacon and the origins of experimentation. Isis, 99(4):731–760, December 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597767.

Kohler:2008:LHR

[3367] Robert E. Kohler. Lab history: Reflections. Isis, 99(4):761–768, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/595769.

Klein:2008:LCS

[3368] Ursula Klein. The laboratory challenge: Some revisions of the standard view of early modern experimentation. Isis, 99(4):769–782, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/595771.

Gooday:2008:PRL

[3369] Graeme Gooday. Placing or replacing the laboratory in the history of science? Isis, 99(4):783–795, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/595772.

Gieryn:2008:LDP

[3370] Thomas F. Gieryn. Laboratory design for post-fordist science. Isis, 99 (4):796–802, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
595773.

Vogt:2008:EVS

[3371] Annette B. Vogt. Eloge: Vladimir Semyonovich Kirsanov, 26 December 1936–12 May 2007. Isis, 99(4):803–805, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/595774.

Ferreiro:2008:LE

[3372] Larrie D. Ferreiro. Letters to the Editor. Isis, 99(4):806, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597747.

McGee:2008:LE

[3373] David McGee. Letters to the Editor. Isis, 99(4):806–807, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597748.

Anonymous:2008:EN

[3374] Anonymous. Editorial note. Isis, 99(4):807, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597749.

Kremer:2008:BRB

[3375] Richard L. Kremer. Book review: Academic Charisma and the Origins of the Research University. Isis, 99(4):808–810, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/595776.

Cohen-Cole:2008:BRB

[3376] Jamie Cohen-Cole. Book review: Mind as Machine: A History of Cognitive Science. Isis, 99(4):811–812, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597681.

Walton:2008:BRBb

[3377] Steven A. Walton. Book review: Gunpowder, Explosives, and the State: A Technological History. Isis, 99(4):812–813, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597682.

Laird:2008:BRB

[3378] W. R. Laird. Book review: Mechanics from Aristotle to Einstein. Isis, 99 (4):813–814, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597683.

Drayton:2008:BRB

[3379] Ralph Drayton. Book review: Science and Religion, 400 B.C. to A.D. 1550: From Aristotle to Copernicus, by Edward Grant. Isis, 99(4):814– 816, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597684.

Buckley:2008:BRB

[3380] Eve Buckley. Book review: An Environmental History of Latin America. Isis, 99(4):816–817, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597685.

vanBerkel:2008:BRB

[3381] Klaas van Berkel. Book review: The Two Books: Historical Notes on Some Interactions between Natural Science and Theology. Isis, 99(4): 817–818, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597686.

Turner:2008:BRBb

[3382] R. Steven Turner. Book review: The Microscope and the Eye: A History of Reflections, 1740–1870. Isis, 99(4):818–819, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597687.

Thurs:2008:BRB

[3383] Daniel Patrick Thurs. Book review: Health and the Rhetoric of Medicine. Isis, 99(4):819–820, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597688.

Sepkoski:2008:BRB

[3384] David Sepkoski. Book review: On Deep History and the Brain. Isis, 99 (4):820–821, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597689.

Unguru:2008:BRBb

[3385] Sabetai Unguru. Book review: Amazing Traces of a Babylonian Origin in Greek Mathematics. Isis, 99(4):821–822, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597750.

Zuccato:2008:BRB

[3386] Marco Zuccato. Book review: Ordering the Heavens: Roman Astronomy and Cosmology in the Carolingian Renaissance. Isis, 99(4):823–824, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597690.

Marr:2008:BRB

[3387] Alexander Marr. Book review: Gutenberg and the Impact of Printing. Isis, 99(4):824–825, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597691.

Stevens:2008:BRB

[3388] Wesley M. Stevens. Book review: Les Alchandreana primitifs: Étude sur les plus anciens traités astrologiques Latins d’origine arabe (Xe siècle). Isis, 99(4):825–826, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597692.

Chamberland:2008:BRB

[3389] Celeste Chamberland. Book review: Books of Secrets: Natural Philosophy in England, 1550–1600. Isis, 99(4):826–827, December 2008.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597693.

Davenport:2008:BRB

[3390] Anne Davenport. Book review: Demonstration and Scientific Knowledge in William of Ockham: A Translation of Summa Logicae III-II: De Syllogismo Demonstrativo, and Selections from the Prologue to the Ordinatio. Isis, 99(4):827–828, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597694.

Crowther:2008:BRB

[3391] Kathleen Crowther. Book review: The Science of Describing: Natural History in Renaissance Europe. Isis, 99(4):828–829, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597695.

Kahn:2008:BRBb

[3392] Didier Kahn. Book review: Alchemy: A Bibliography of English Language Writings. Isis, 99(4):829–830, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597696.

Clucas:2008:BRB

[3393] Stephen Clucas. Book review: John Dee’s Occultism: Magical Exaltation through Powerful Signs. Isis, 99(4):830–831, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597697.

Caianiello:2008:BRB

[3394] Silvia Caianiello. Book review: Viaggio nel paese delle meraviglie: Scienza e curiosità nell’Italia del Settecento. Isis, 99(4):832–833, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/597698.

Finocchiaro:2008:BRB

[3395] Maurice A. Finocchiaro. Book review: Galileo e Keplero: Filosofia, cosmologia e teologia nell’Età della Controriforma. Isis, 99(4):833–834, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597699.

Spary:2008:BRB

[3396] Emma Spary. Book review: Histoire naturelle, générale et particulière, avec la description du Cabinet du roy. Isis, 99(4):834–835, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/597700.

Marchand:2008:BRB

[3397] Suzanne Marchand. Book review: The Science of Culture in Enlightenment Germany. Isis, 99(4):835–837, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/597701.

Edwards:2008:BRB

[3398] Michael Edwards. Book review: Thomas Hobbes and the Politics of Natural Philosophy. Isis, 99(4):837–838, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/597702.

Good:2008:BRBb

[3399] Greg Good. Book review: British Weather and the Climate of Enlightenment. Isis, 99(4):838–839, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597703.

Cremiere:2008:BRB

[3400] Cédric Crémière. Book review: Buffon: Histoire naturelle et philosophie. Isis, 99(4):839–840, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597704.

Jacob:2008:BRB

[3401] Margaret C. Jacob. Book review: Divided by Faith: Religious Conflict and the Practice of Toleration in Early Modern Europe. Isis, 99(4):840– 841, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597705.

Lefevre:2008:BRBb

[3402] Wolfgang Lefèvre. Book review: Theatermaschine und Festungsbau: Paradoxien der Wissensproduktion im 17, by Jan Lazardzig. Jahrhundert. Isis, 99(4):841–842, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597706.

Malet:2008:BRB

[3403] Antoni Malet. Book review: John Wallis: Una vita per un progetto. Isis, 99(4):842–843, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597707.

Shea:2008:BRBb

[3404] William R. Shea. Book review: Galileo’s Glassworks: The Telescope and the Mirror. Isis, 99(4):843–844, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597708.

DeAngelis:2008:BRB

[3405] Simone De Angelis. Book review: Widmung, Welterklärung und Wissenschaftslegitimierung: Titelbilder und ihre Funktionen in der Wissenschaftlichen Revolution. Isis, 99(4):844–846, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597709.

Vermeir:2008:BRB

[3406] Koen Vermeir. Book review: Die große kosmologische Kontroverse: Rekonstruktionsversuche anhand des Itinerarium exstaticum von Athanasius Kircher SJ (1602–1680). Isis, 99(4):846–847, December 2008. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597710.

Sutter:2008:BRBb

[3407] Paul Sutter. Book review: The Tropical World of Samuel Taylor Darling: Parasites, Pathology, and Philanthropy. Isis, 99(4):847–848, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/597711.

Gossin:2008:BRB

[3408] Pamela Gossin. Book review: The Sun Kings: The Unexpected Tragedy of Richard Carrington and the Tale of How Modern Astronomy Began. Isis, 99(4):848–849, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/597712.

Russett:2008:BRB

[3409] Cynthia Russett. Book review: Darwin, Literature, and Victorian Respectability. Isis, 99(4):849–850, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597713.

Lacki:2008:BRB

[3410] Jan Lacki. Book review: The Collected Papers of Albert Einstein, by Albert Einstein; Diana Kormos Buchwald; Tilman Sauer; Ze’ev Rosenkranz; József Illy; Virginia Iris Holmes. Volume 10: The Berlin Years: Correspondence, May-December 1920, and Supplementary Correspondence, 1909–1920. Isis, 99(4):850–851, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597714.

Reynolds:2008:BRB

[3411] Andrew Reynolds. Book review: Styles of Reasoning in the British Life Sciences: Shared Assumptions, 1820–1858. Isis, 99(4):851–852, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/597715.

Bowler:2008:BRB

[3412] Peter J. Bowler. Book review: Science in the Marketplace: Nineteenth-Century Sites and Experiences. Isis, 99(4):852–853, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597716.

Knight:2008:BRB

[3413] David Knight. Book review: Victorian Popularizers of Science: Designing Nature for New Audiences. Isis, 99(4):853–855, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597717.

Wenkel:2008:BRB

[3414] Simone Wenkel. Book review: From Berlin to New York: Life and Work of the Almost Forgotten German-Jewish Biochemist Carl Neuberg (1877– 1956). Isis, 99(4):855–856, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597718.

Fayter:2008:BRB

[3415] Paul Fayter. Book review: Science, Time, and Space in the Late Nineteenth-Century Periodical Press: Movable Types. Isis, 99(4):856– 858, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597719.

Stoltzner:2008:BRB

[3416] Michael Stöltzner. Book review: The Reign of Relativity: Philosophy in Physics, 1915–1925. Isis, 99(4):858–859, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597720.

Stewart:2008:BRBb

[3417] Mart A. Stewart. Book review: The Humboldt Current: Nineteenth-Century Exploration and the Roots of American Environmentalism. Isis, 99(4):859–860, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597721.

Weindling:2008:BRBb

[3418] Paul Weindling. Book review: “Freiheit der Naturforschung!” Carl Freiherr von Rokitansky und die Wiener Medizinische Schule: Wissenschaft und Politik im Konflikt. Isis, 99(4):860, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/597722.

Dowd:2008:BRB

[3419] Matthew F. Dowd. Book review: Practical Mystic: Religion, Science, and A, by Matthew Stanley. S. Eddington. Isis, 99(4):861, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/597723.

Beck:2008:BRB

[3420] Naomi Beck. Book review: The Philosophy of Herbert Spencer. Isis, 99(4):862, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597724.

Veneer:2008:BRB

[3421] Leucha Veneer. Book review: Excavating Victorians. Isis, 99(4):862–864, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597725.

Mazumdar:2008:BRB

[3422] Pauline M. H. Mazumdar. Book review: Expunging Variola: The Control and Eradication of Smallpox in India, 1947–1977. Isis, 99(4):864, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597726.

Shell:2008:BRB

[3423] Hanna Rose Shell. Book review: Films of Fact: A History of Science in Documentary Films and Television. Isis, 99(4):865–866, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597727.

Neushul:2008:BRB

[3424] Peter Neushul. Book review: Science in Flux: NASA’s Nuclear Program at Plum Brook Station, 1955–2005. Isis, 99(4):866–867, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597728.

Wilkins:2008:BRB

[3425] John Wilkins. Book review: The Panda’s Black Box: Opening up the Intelligent Design Controversy Living with Darwin: Evolution, Design, and the Future of Faith. Isis, 99(4):867–868, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597729.

Armus:2008:BRB

[3426] Diego Armus. Book review: Cold War, Deadly Fevers: Malaria Eradication in Mexico, 1955–1975. Isis, 99(4):868–869, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597730.

Topper:2008:BRB

[3427] David R. Topper. Book review: Six Stories from the End of Representation: Images in Painting, Photography, Astronomy, Microscopy, Particle Physics, and Quantum Mechanics, 1980–2000. Isis, 99(4):869–870, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597731.

Olley:2008:BRB

[3428] Allan Olley. Book review: When Computers Were Human. Isis, 99 (4):870–871, December 2008. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597732.

Ensmenger:2008:BRB

[3429] Nathan Ensmenger. Book review: Beyond AI: Creating the Conscience of the Machine. Isis, 99(4):872, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597733.

Meng:2008:BRB

[3430] Yue Meng. Book review: Selected Materials on the Science and Technology in the People’s Republic of China (1949–1995). Isis, 99(4):872–873, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597734.

Krige:2008:BRB

[3431] John Krige. Book review: Half-Lives and Half-Truths: Confronting the Radioactive Legacies of the Cold War. Isis, 99(4):873–874, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/597735.

Shen:2008:BRBb

[3432] Grace Y. Shen. Book review: Weng Wenhao nian pu [The Chronicle of Dr. Weng Wenhao], by Xuetong Li. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu). Isis, 99(4):874–875, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597736.

WellDone:2008:BRB

[3433] Steve “The Bookie” Well Done. Book review: The Wonderful Adventures of Nat Selleck and Eva Lou Shinn in Sci Fi Land: A Spoof on Evolution and Natural Selection. Isis, 99(4):875–876, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597737.

Malloy:2008:BRP

[3434] Sean L. Malloy. Book review: P. D. Smith: Doomsday Men: The Real Dr. Strangelove and the Dream of the Superweapon. Isis, 99(4):876–877, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597738;
 http://www.jstor.org/stable/pdfplus/10.1086/597738.pdf.

Pietikainen:2008:BRB

[3435] Petteri Pietikainen. Book review: Coercion as Cure: A Critical History of Psychiatry. Isis, 99(4):877–878, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/597739.

Bachi:2008:BRB

[3436] Beat Bächi. Book review: Plastik und politische Kultur in Westdeutschland. Isis, 99(4):878–879, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597740.

Schweber:2008:BRBb

[3437] S. S. Schweber. Book review: Meeting the Universe Halfway: Quantum Physics and the Entanglement of Matter and Meaning. Isis, 99(4):879– 882, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597741.

Endersby:2008:BRB

[3438] Jim Endersby. Book review: When Species Meet. Isis, 99(4):882–883, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597742.

Myers:2008:BRB

[3439] Natasha Myers. Book review: Marking Time: On the Anthropology of the Contemporary. Isis, 99(4):883–884, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597743.

Lynch:2008:BRB

[3440] Michael Lynch. Book review: The Politics of Life Itself: Biomedicine, Power, and Subjectivity in the Twenty-first Century. Isis, 99(4):884–885, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597744.

Pitt:2008:BRB

[3441] Joseph C. Pitt. Book review: Philosophical Instruments: Minds and Tools at Work. Isis, 99(4):885–886, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/597745.

Waller:2008:BRB

[3442] John C. Waller. Book review: Mind in Life: Biology, Phenomenology, and the Sciences of Mind. Isis, 99(4):886–887, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597746.

Anonymous:2008:NCd

[3443] Anonymous. Notes on contributors. Isis, 99(4):888–892, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/595778.

Anonymous:2008:IBV

[3444] Anonymous. Index to Isis, volume 99, 2008. Isis, 99(4):893–906, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/597752.

Anonymous:2008:CV

[3445] Anonymous. Contents, volume 99. Isis, 99(4):909–911, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597830.

Anonymous:2008:Fd

[3446] Anonymous. Frontispiece. Isis, 99(4):vi, December 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/isis.2008.99.issue-4.vi.

Anonymous:2009:CBH

[3447] Anonymous. Current bibliography of the history of science and its cultural influences, 2009. Isis, 100(S1):i–280, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652112.

Anonymous:2009:JUT

[3448] Anonymous. Journals under threat: A joint response from history of science, technology, and medicine Editors. Isis, 100(1):1–3, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597576.

Dauben:2009:SDH

[3449] Joseph W. Dauben, Mary Louise Gleason, and George E. Smith. Seven decades of history of science: I. Bernard Cohen (1914–2003), second Editor of Isis. Isis, 100(1):4–35, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597575.

Renwick:2009:PSS

[3450] Chris Renwick. The practice of Spencerian science: Patrick Geddes’s biosocial program, 1876–1889. Isis, 100(1):36–57, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597574.

Lightman:2009:I

[3451] Bernard Lightman. Introduction. Isis, 100(1):58–59, March 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597573.

Pyenson:2009:EIM

[3452] Lewis Pyenson and Christophe Verbruggen. Ego and the international: The modernist circle of George Sarton. Isis, 100(1):60–78, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597572.

Holton:2009:GSH

[3453] Gerald Holton. George Sarton, his Isis, and the aftermath. Isis, 100(1): 79–88, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/597571.

Dear:2009:HSH

[3454] Peter Dear. The history of science and the history of the sciences: George Sarton, Isis, and the two cultures. Isis, 100(1):89–93, March 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597570.

Maienschein:2009:RSI

[3455] Jane Maienschein. Rethinking Sarton’s Institute for History of Science and Civilization — virtually. Isis, 100(1):94–102, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/597569.

Numbers:2009:AHS

[3456] Ronald L. Numbers. The American History of Science Society or the International History of Science Society? The fate of cosmopolitanism since George Sarton. Isis, 100(1):103–107, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/597568.

Grant:2009:ECW

[3457] Edward Grant. Eloge: Claudia Wilson Kren, 29 October 1927–26 January 2008. Isis, 100(1):108–110, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599630.

Herbert:2009:EFH

[3458] Sandra Herbert and Alison Pearn. Eloge: Frederick H. Burkhardt, 1912–2007. Isis, 100(1):111–114, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599631.

Lynch:2009:LEa

[3459] John M. Lynch and Graeme Gooday. Letters to the Editor. Isis, 100(1): 115, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599633.

Weikart:2009:LE

[3460] Richard Weikart. Letters to the Editor. Isis, 100(1):115, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599632.

Fuller:2009:LE

[3461] Steve Fuller. Letters to the Editor. Isis, 100(1):115–116, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599634.

Lynch:2009:LEb

[3462] John M. Lynch and Graeme Gooday. Letters to the Editor. Isis, 100(1): 116, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599635.

Wilson:2009:LE

[3463] John J. Wilson, Robin Floyd, Robert H. Hanner, and David Castle. Letters to the Editor. Isis, 100(1):117, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599636.

Hamilton:2009:LE

[3464] Andrew Hamilton and Quentin D. Wheeler. Letters to the Editor. Isis, 100(1):117–118, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599637.

Kubbinga:2009:ERB

[3465] Henk Kubbinga. Essay review: Niels Bohr’s Collected Works, Volumes 11 and 12, edited by Finn Aaserud. Isis, 100(1):119–126, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597565.

Kusch:2009:BRB

[3466] Martin Kusch. Book review: Objectivity and Historiography. Isis, 100 (1):127–131, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
597564.

Belle:2009:BRB

[3467] Riccardo Bellé. Book review: The Geometry of an Art: The History of the Mathematical Theory of Perspective from Alberti to Monge, by Kirsti Andersen. (Sources and Studies in the History of Mathematics and Physical Sciences.). Isis, 100(1):132–133, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599638.

Nash:2009:BRB

[3468] Richard Nash. Book review: Encounters with Wild Children: Temptation and Disappointment in the Study of Human Nature. Isis, 100(1):133– 134, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599639.

Maienschein:2009:BRB

[3469] Jane Maienschein. Book review: Heredity and Hope: The Case for Genetic Screening Babies by Design: The Ethics of Genetic Choice The Case against Perfection: Ethics in the Age of Genetic Engineering. Isis, 100(1):134–136, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599640.

Savitt:2009:BRB

[3470] Steven Savitt. Book review: Understanding Space-Time: The Philosophical Development of Physics from Newton to Einstein. Isis, 100(1):136– 137, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599641.

Bangu:2009:BRB

[3471] Sorin Bangu. Book review: Representation and Productive Ambiguity in Mathematics and the Sciences. Isis, 100(1):137–139, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599642.

Lachmund:2009:BRB

[3472] Jens Lachmund. Book review: Urban Machinery: Inside Modern European Cities, by Mikael Hård; Thomas J. Misa. (Inside Technology.). Isis, 100(1):139–140, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599643.

Spanagel:2009:BRB

[3473] David Spanagel. Book review: Four Centuries of Geological Travel: The Search for Knowledge on Foot, Bicycle, Sledge, and Camel, by Patrick N. Wyse Jackson. (Geological Society Special Publications, 287.). Isis, 100(1):140–141, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599644.

Kimmelman:2009:BRB

[3474] Barbara Kimmelman. Book review: Fruits and Plains: The Horticultural Transformation of America. Isis, 100(1):141–142, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599645.

Golinski:2009:BRB

[3475] Jan Golinski. Book review: The Mindful Hand: Inquiry and Invention from the Late Renaissance to Early Industrialisation, by Lissa Roberts; Simon Schaffer; Peter Dear. (History of Science and Scholarship in The Netherlands, 9.). Isis, 100(1):142–144, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599646.

Cooter:2009:BRBa

[3476] Roger Cooter. Book review: Our Present Complaint: American Medicine, Then and Now. Isis, 100(1):144–145, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599647.

McClelland:2009:BRB

[3477] Charles E. McClelland. Book review: Examen, Titel, Promotionen: Akademisches und staatliches Qualifikationswesen vom 13 bis zum 21. Jahrhundert, by Rainer Christoph Schwinges. (Veröffentlichungen der Gesellschaft für Universitäts- und Wissenschaftsgeschichte, 7.). Isis, 100 (1):145–146, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599648.

Reinarz:2009:BRB

[3478] Jonathan Reinarz. Book review: Sensing the Past: Seeing, Hearing, Smelling, Tasting, and Touching in History. Isis, 100(1):146–147, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599649.

Huff:2009:BRB

[3479] Toby Huff. Book review: Naturwissenschaften im Kulturvergleich: Europa-Islam-China. Isis, 100(1):147–149, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/599650.

Olshin:2009:BRBa

[3480] Benjamin B. Olshin. Book review: Technology and Culture in Greek and Roman Antiquity, by S. Cuomo. (Key Themes in Ancient History.). Isis, 100(1):149–150, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599651.

Hahm:2009:BRB

[3481] David E. Hahm. Book review: The Origin of the History of Science in Classical Antiquity. Isis, 100(1):150–151, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/599652.

Klaassen:2009:BRB

[3482] Frank Klaassen. Book review: Entre science et nigromance: Astrologie, divination et magie dans l’Occident médiéval (XIIe-XVe siècle). Isis, 100(1):151–152, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599653.

Siraisi:2009:BRB

[3483] Nancy G. Siraisi. Book review: Le Liber de virtutibus herbarum, lapidum et animalium (Liber aggregationis): Un texte à succès attribué à Albert le Grand, by Isabelle Draelants. (Micrologus’ Library, 22.). Isis, 100(1): 152, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599654.

Minehan:2009:BRBa

[3484] Bridie Andrews Minehan. Book review: Explorations in Daoism: Medicine and Alchemy in Literature. Isis, 100(1):153–154, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599655.

Folkerts:2009:BRB

[3485] Menso Folkerts. Book review: The Archimedes Codex: How a Medieval Prayer Book Is Revealing the True Genius of Antiquity’s Greatest Scientist. Isis, 100(1):154–155, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/599656.

Rius:2009:BRB

[3486] Mònica Rius. Book review: Volkstümliche Astronomie im islamischen Mittelalter: Zur Bestimmung der Gebetszeiten und der Qibla bei al-As.bah.i, Ibn Rah.iq una al-Farisi, by Petra G. Schmidl. Volume 1: Texte und Übersetzungen; Volume 2: Erläuterungen und Zusammenfassungen. (Islamic Philosophy, Theology, and Science: Texts and Studies, 68.). Isis, 100(1):155–156, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599657.

Stolzenberg:2009:BRB

[3487] Daniel Stolzenberg. Book review: History, Medicine, and the Traditions of Renaissance Learning. Isis, 100(1):156–157, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599658.

Rammer:2009:BRB

[3488] Gerhard Rammer. Book review: Wind and Water in the Middle Ages: Fluid Technologies from Antiquity to the Renaissance, by Steven A. Walton. (Medieval and Renaissance Texts and Studies, 322; Penn State Medieval Studies, 2.). Isis, 100(1):157–158, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/599659.

Outram:2009:BRB

[3489] Dorinda Outram. Book review: Storms and Dreams: Louis de Bougainville: Soldier, Explorer, Statesman. Isis, 100(1):158–159, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599660.

Jones:2009:BRB

[3490] Matthew L. Jones. Book review: Language, Mind, and Nature: Artificial Languages in England from Bacon to Locke, by Rhodri Lewis. (Ideas in Context, 80.). Isis, 100(1):159–160, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599661.

Shapiro:2009:BRB

[3491] Alan E. Shapiro. Book review: The Creation of Color in Eighteenth-Century Europe. Isis, 100(1):160–161, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/599662.

Richards:2009:BRB

[3492] Joan L. Richards. Book review: The World of Maria Gaetana Agnesi, Mathematician of God. Isis, 100(1):161–162, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/599663.

Shackelford:2009:BRB

[3493] Jole Shackelford. Book review: Alchemy and Authority in the Holy Roman Empire. Isis, 100(1):162–164, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599664.

Taylor:2009:BRBa

[3494] Kenneth L. Taylor. Book review: Les tremblements de terre aux XVIIe et XVIIIe siècles: La naissance d’un risque, by Grégory Quenet. (Époques.). Isis, 100(1):164–165, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599665.

Barnett:2009:BRBa

[3495] Lydia Barnett. Book review: Les neiges labiles: Une histoire culturelle de l’avalanche au XVIIIe siècle, by Raphaël Rabusseau. (Travaux d’Histoire Suisse, 4.). Isis, 100(1):165–166, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599666.

Boantza:2009:BRB

[3496] Victor D. Boantza. Book review: The Salt of the Earth: Natural Philosophy, Medicine, and Chymistry in England, 1650–1750, by Anna Marie Roos. (History of Science and Medicine Library, 3.). Isis, 100(1):166– 167, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599667.

Rappaport:2009:BRBa

[3497] Rhoda Rappaport. Book review: The Earth Sciences in the Enlightenment: Studies on the Early Development of Geology, by Kenneth L.

Taylor. (Variorum Collected Studies Series.). Isis, 100(1):167–168, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599668.

Hunt:2009:BRB

[3498] Bruce J. Hunt. Book review: One Time Fits All: The Campaigns for Global Uniformity. Isis, 100(1):168–169, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/599669.

Robinson:2009:BRB

[3499] Michael Robinson. Book review: Trying Leviathan: The Nineteenth-Century New York Court Case That Put the Whale on Trial and Challenged the Order of Nature. Isis, 100(1):169–171, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599670.

Finkelstein:2009:BRB

[3500] Gabriel Finkelstein. Book review: Zwischen den Sternen: Lichtbildarchive: Was Einstein und Uexküll, Benjamin und das Kino der Astronomie des 19, by Karl Clausberg. Jahrhunderts verdanken. Isis, 100 (1):171, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599671.

Morris:2009:BRB

[3501] Peter J. T. Morris. Book review: The Rise and Decline of Colloid Science in North America, 1900–1935: The Neglected Dimension, by Andrew Ede. (Science, Technology, and Culture, 1700–1945.). Isis, 100(1):171– 172, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599672.

Gasper:2009:BRB

[3502] Michael Gasper. Book review: The Great Social Laboratory: Subjects of Knowledge in Colonial and Postcolonial Egypt. Isis, 100(1):172–173, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/599673.

Pilpel:2009:BRB

[3503] Avital Pilpel. Book review: Ending the Mendel-Fisher Controversy. Isis, 100(1):173–174, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599674.

Shelton:2009:BRB

[3504] Robert Shelton. Book review: Thomas Hardy’s Novel Universe: Astronomy, Cosmology, and Gender in the Post-Darwinian World, by Pamela Gossin. (Nineteenth Century.). Isis, 100(1):175–176, March 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599675.

Henry:2009:BRBa

[3505] John Henry. Book review: Recreating Newton: Newtonian Biography and the Making of Nineteenth-Century History of Science. Isis, 100(1):176– 177, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599676.

Good:2009:BRB

[3506] Greg Good. Book review: The Herschels of Hanover. Isis, 100(1):177– 178, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599677.

Burian:2009:BRB

[3507] Richard Burian. Book review: From Embryology to Evo-Devo: A History of Developmental Evolution, by Manfred D. Laubichler; Jane Maienschein. (Dibner Institute Studies in the History of Science and Technology.). Isis, 100(1):178–180, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599678.

Forman:2009:BRD

[3508] Paul Forman. Book review: David Lindley: Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for the Soul of Science. Isis, 100 (1):180–181, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599679.

Reeds:2009:BRBa

[3509] Karen Reeds. Book review: Britton’s Botanical Empire: The New York Botanical Garden and American Botany, 1888–1929, by Peter Mickulas. (Memoirs of the New York Botanical Garden, 94.). Isis, 100(1):181– 182, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599680.

Smith:2009:BRBa

[3510] Robert W. Smith. Book review: Von Braun: Dreamer of Space, Engineer of War. Isis, 100(1):182–183, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599681.

Brooke:2009:BRB

[3511] John Hedley Brooke. Book review: Science and Scientism in Nineteenth-Century Europe. Isis, 100(1):183–184, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/599682.

Weisel:2009:BRB

[3512] Gary J. Weisel. Book review: The Genesis of General Relativity, by Jürgen Renn. Volume 1: Einstein’s Zurich Notebook: Introduction and Source. Volume 2: Einstein’s Zurich Notebook: Commentary and Essays. Volume 3: Gravitation in the Twilight of Classical Physics: Between Mechanics, Field Theory, and Astronomy. Volume 4: Gravitation in the Twilight of Classical Physics: The Promise of Mathematics. (Boston Studies in the Philosophy of Science, 250.). Isis, 100(1):185–186, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599683.

Gordin:2009:BRB

[3513] Michael D. Gordin. Book review: Einstein and Oppenheimer: The Meaning of Genius. Isis, 100(1):186–188, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599684.

Regal:2009:BRB

[3514] Brian Regal. Book review: Bones and Ochre: The Curious Afterlife of the Red Lady of Paviland. Isis, 100(1):188–189, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599685.

Beyler:2009:BRH

[3515] Richard Beyler. Book review: Vom Hintereingang zum Hauptportal? Lise Meitner und ihre Kolleginnen an der Berliner Universität und in der Kaiser-Wilhelm-Gesellschaft. Isis, 100(1):189–190, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599686.

Comfort:2009:BRB

[3516] Nathaniel C. Comfort. Book review: Davenport’s Dream: Twenty-first Century Reflections on Heredity and Eugenics. Isis, 100(1):191–192, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/599687.

Minehan:2009:BRBb

[3517] Bridie Andrews Minehan. Book review: Zhongguo jin dai ji bing she hui shi [A Social History of Diseases in Modern China (1912–1937)], by Zhang Daqing. (Zhongguo jin xian dai ke xue ji shu shi yan jiu cong shu.). Isis, 100(1):192–193, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/599688.

Cowan:2009:BRB

[3518] Ruth Schwartz Cowan. Book review: Perfect Motherhood: Science and Childrearing in America. Isis, 100(1):193–194, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599689.

Neufeld:2009:BRB

[3519] Michael J. Neufeld. Book review: Red Prometheus: Engineering and Dictatorship in East Germany, 1945–1990. Isis, 100(1):194–195, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599690.

Greene:2009:BRBa

[3520] Mott Greene. Book review: Authors of the Storm: Meteorologists and the Culture of Prediction. Isis, 100(1):195–197, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599691.

Ceruzzi:2009:BRB

[3521] Paul E. Ceruzzi. Book review: The Internet Imaginaire. Isis, 100(1): 197, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599692.

Harper:2009:BRB

[3522] Kristine C. Harper. Book review: Disrupting Science: Social Movements, American Scientists, and the Politics of the Military, 1945–1975. Isis, 100(1):197–199, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599693.

Servos:2009:BRB

[3523] John W. Servos. Book review: The Price of Truth: How Money Affects the Norms of Science, by David B. Resnik; Daniel S. Greenberg. (Practical and Professional Ethics.) Science for Sale: The Perils, Rewards, and Delusions of Campus Capitalism. Isis, 100(1):199–200, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599694.

Simmonds:2009:BRB

[3524] Jeanette Simmonds. Book review: Structures of Scientific Collaboration, by Wesley Shrum; Joel Genuth; Ivan Chompalov. (Inside Technology.). Isis, 100(1):200–201, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/599695.

Durant:2009:BRB

[3525] Darrin Durant. Book review: Conservation Fallout: Nuclear Protest at Diablo Canyon. Isis, 100(1):201–202, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599696.

Richardson:2009:BRB

[3526] Alan Richardson. Book review: The Challenge of the Social and the Pressure of Practice: Science and Values Revisited. Isis, 100(1):202– 203, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599697.

Papayannakos:2009:BRB

[3527] Dimitris Papayannakos and Stathis Psillos. Book review: A Metaphysics for Scientific Realism: Knowing the Unobservable. Isis, 100(1):204–205, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/599698.

Lynch:2009:BRB

[3528] William T. Lynch. Book review: Rethinking Expertise. Isis, 100(1):205– 206, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599699.

Radick:2009:BRB

[3529] Gregory Radick. Book review: Scientific Perspectivism Scientific Pluralism, by Ronald N. Giere; Stephen H. Kellert; Helen E. Longino; C. Kenneth Waters. (Minnesota Studies in the Philosophy of Science.). Isis, 100(1):206–207, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
599700.

Fuller:2009:BRBa

[3530] Steve Fuller. Book review: The Handbook of Science and Technology Studies. Isis, 100(1):207–209, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599701.

Tucker:2009:BRB

[3531] Aviezer Tucker. Book review: Historical Knowledge, Historical Error: A Contemporary Guide to Practice. Isis, 100(1):209–211, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599702.

Lloyd:2009:BRB

[3532] G. E. R. Lloyd. Book review: What Science Offers the Humanities: Integrating Body and Culture. Isis, 100(1):211–212, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599703.

Anonymous:2009:NCa

[3533] Anonymous. Notes on contributors. Isis, 100(1):213–218, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597562.

Anonymous:2009:Fa

[3534] Anonymous. Frontispiece. Isis, 100(1):viii, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/599704.

Cohen-Cole:2009:CAC

[3535] Jamie Cohen-Cole. The creative American: Cold War salons, social science, and the cure for modern society. Isis, 100(2):219–262, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599554.

Vicedo:2009:FEF

[3536] Marga Vicedo. The father of ethology and the Foster mother of ducks: Konrad Lorenz as expert on motherhood. Isis, 100(2):263–291, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599553.

Porter:2009:HSB

[3537] Theodore M. Porter. How science became technical. Isis, 100(2):292– 309, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/599552.

Topham:2009:I

[3538] Jonathan R. Topham. Introduction. Isis, 100(2):310–318, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599551.

Daum:2009:VPS

[3539] Andreas W. Daum. Varieties of popular science and the transformations of public knowledge: Some historical reflections. Isis, 100(2):319–332, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/599550.

OConnor:2009:RPS

[3540] Ralph O’Connor. Reflections on popular science in Britain: Genres, categories, and historians. Isis, 100(2):333–345, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/599549.

Pandora:2009:PSN

[3541] Katherine Pandora. Popular science in national and transnational perspective: Suggestions from the American context. Isis, 100(2):346–358, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/599548.

Bensaude-Vincent:2009:HPS

[3542] Bernadette Bensaude-Vincent. A historical perspective on science and its “others”. Isis, 100(2):359–368, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/599547.

Maienschein:2009:EPJ

[3543] Jane Maienschein. Eloge: Philip J. Pauly, 3 September 1950–2 April 2008. Isis, 100(2):369–371, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/599546.

Darling:2009:LE

[3544] Karen Darling. Letters to the Editor. Isis, 100(2):372, June 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599545.

Monaldi:2009:BRB

[3545] Daniela Monaldi. Book review: Particle or Wave: The Evolution of the Concept of Matter in Modern Physics. Isis, 100(2):373–374, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605210.

Garber:2009:BRB

[3546] Daniel Garber. Book review: The Artificial and the Natural: An Evolving Polarity. Isis, 100(2):374–375, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605211.

Tredwell:2009:BRB

[3547] Katherine A. Tredwell. Book review: Mechanics and Cosmology in the Medieval and Early Modern Period. Isis, 100(2):375–376, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605212.

Eisenberg:2009:BRB

[3548] Ziv Eisenberg. Book review: Who Chooses? American Reproductive History since 1830. Isis, 100(2):376–377, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605213.

Brysse:2009:BRB

[3549] Keynyn Brysse. Book review: A History of Paleontology Illustration. Isis, 100(2):377–378, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605214.

Knight:2009:BRB

[3550] David Knight. Book review: From Alchemy to Chemistry in Picture and Story. Isis, 100(2):378–379, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/605215.

Fallwell:2009:BRB

[3551] Lynne Fallwell. Book review: The Healthy Jew: The Symbiosis of Judaism and Modern Medicine. Isis, 100(2):379–380, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605216.

vanBerkel:2009:BRB

[3552] Klaas van Berkel. Book review: Advancements of Learning: Essays in Honour of Paolo Rossi. Isis, 100(2):380–381, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605217.

Archibald:2009:BRB

[3553] Tom Archibald. Book review: The Mathematics of Egypt, Mesopotamia, China, India, and Islam: A Sourcebook. Isis, 100(2):381–382, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605218.

Sturdy:2009:BRB

[3554] Steve Sturdy. Book review: Crafting Immunity: Working Histories of Clinical Immunology. Isis, 100(2):382–383, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/605219.

Ragep:2009:BRB

[3555] F. Jamil Ragep. Book review: The Beginnings of Western Science: The European Scientific Tradition in Philosophical, Religious, and Institutional Context, Prehistory to a, by David C. Lindberg.d. 1450. Isis, 100(2):383–385, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605220.

Ruse:2009:BRBa

[3556] Michael Ruse. Book review: The Evolving World: Evolution in Everyday Life. Isis, 100(2):385, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/605221.

Hoskin:2009:BRB

[3557] Michael Hoskin. Book review: Cosmos: An Illustrated History of Astronomy and Cosmology. Isis, 100(2):386, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605222.

Mueller-Wille:2009:BRB

[3558] Staffan Mueller-Wille. Book review: In Pursuit of the Gene: From Darwin to DNA. Isis, 100(2):386–387, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605223.

Fayter:2009:BRB

[3559] Paul Fayter. Book review: Hollow Earth: The Long and Curious History of Imagining Strange Lands, Fantastical Creatures, Advanced Civilizations, and Marvelous Machines below the Earth’s Surface. Isis, 100(2): 387–388, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605224.

Pols:2009:BRB

[3560] Hans Pols. Book review: History of Psychiatry and Medical Psychology: With an Epilogue on Psychiatry and the Mind-Body Relation. Isis, 100(2):388–389, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605225.

Chen:2009:BRB

[3561] Xiang Chen. Book review: Zhongguo jin xian dai ji liang shi gao [A Draft of the History of Modern and Contemporary Metrology in China]. Isis, 100(2):389–390, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605226.

Truitt:2009:BRB

[3562] E. R. Truitt. Book review: The Natural and the Supernatural in the Middle Ages. Isis, 100(2):390–391, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605227.

Seidengart:2009:BRBa

[3563] Jean Seidengart. Book review: De la causa, principio e uno/Über die Ursache, das Prinzip und das Eine. Isis, 100(2):391–392, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605228.

Seidengart:2009:BRBb

[3564] Jean Seidengart. Book review: De l’infinito, universo et mondi/Über das Unendliche, das Universum und die Welten. Isis, 100(2):392, June 2009.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605229.

Buisseret:2009:BRB

[3565] David Buisseret. Book review: The Tropics of Empire: Why Columbus Sailed South to the Indies. Isis, 100(2):392–393, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605230.

Harvey:2009:BRB

[3566] P. D. A. Harvey. Book review: Pictura et scriptura: Textes, images, et herméneutique des Mappae mundi (XIIIe-XVIe siècle). Isis, 100(2):393– 394, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605231.

Forshaw:2009:BRB

[3567] Peter Forshaw. Book review: Andreas Libavius and the Transformation of Alchemy: Separating Chemical Cultures with Polemical Fire. Isis, 100(2):394–395, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605232.

Rocca:2009:BRB

[3568] Julius Rocca. Book review: Résumé de ses livres sur la fabrique du corps humain. Isis, 100(2):395–396, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605233.

Olshin:2009:BRBb

[3569] Benjamin B. Olshin. Book review: The History of Cartography, by David Woodward. Volume 3: Cartography in the European Renaissance. Isis, 100(2):396–398, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605234.

Sarasohn:2009:BRB

[3570] Lisa T. Sarasohn. Book review: Science, Literature, and Rhetoric in Early Modern England. Isis, 100(2):398–399, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/605235.

Scharf:2009:BRB

[3571] Sara Scharf. Book review: L’herbier des philosophes. Isis, 100(2):399– 400, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605236.

Alder:2009:BRB

[3572] Ken Alder. Book review: Un astronome des lumières: Jérôme Lalande Lettres à Madame du Pierry et au juge Honoré Flaugergues. Isis, 100(2): 400–401, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605325.

Calinger:2009:BRB

[3573] Ronald Calinger. Book review: Leonhard Euler. Isis, 100(2):401–402, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605237.

Eamon:2009:BRB

[3574] William Eamon. Book review: Medical Charlatanism in Early Modern Italy. Isis, 100(2):402–403, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/605238.

Osler:2009:BRBa

[3575] Margaret J. Osler. Book review: The Scientific Revolution and the Origins of Modern Science. Isis, 100(2):403–404, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605239.

Gaukroger:2009:BRB

[3576] Stephen Gaukroger. Book review: Natural Science and the Origins of the British Empire. Isis, 100(2):404–405, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605240.

Hine:2009:BRB

[3577] William L. Hine. Book review: Galileo in France: French Reactions to the Theories and Trial of Galileo. Isis, 100(2):405–406, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605241.

Fillafer:2009:BRB

[3578] Franz Leander Fillafer. Book review: Sinnstiftung durch Geschichte? Zur Enstehung spekulativ-universalistischer Geschichtsphilosophie zwischen Bayle und Kant. Isis, 100(2):406–407, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605242.

Shteir:2009:BRB

[3579] Ann Shteir. Book review: The Aliveness of Plants: The Darwins at the Dawn of Plant Science. Isis, 100(2):407–408, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605243.

Harris:2009:BRB

[3580] Martha Harris. Book review: Walther Nernst: Pioneer of Physics and of Chemistry. Isis, 100(2):408–409, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605244.

Schaffer:2009:BRB

[3581] Simon Schaffer. Book review: Hans Christian Ørsted and the Romantic Legacy in Science: Ideas, Disciplines, Practices. Isis, 100(2):409–411, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605245.

Schneiderman:2009:BRB

[3582] Jill Schneiderman. Book review: The Role of Women in the History of Geology. Isis, 100(2):411–412, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605246.

Ferreiros:2009:BRB

[3583] José Ferreirós. Book review: Modern Algebra and the Rise of Mathematical Structures. Isis, 100(2):412–413, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605247.

OConnor:2009:BRB

[3584] Ralph O’Connor. Book review: Romantic Landscapes: Geology and Its Cultural Influence in Britain, 1765–1835. Isis, 100(2):413–414, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/605248.

Braun:2009:BRB

[3585] Marta Braun. Book review: Anatomy of the Passions. Isis, 100(2):414– 415, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605249.

Richmond:2009:BRBa

[3586] Jesse Richmond. Book review: Debating Humankind’s Place in Nature, 1860–2000: The Nature of Paleoanthropology. Isis, 100(2):415–416, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/605250.

Sleigh:2009:BRB

[3587] Charlotte Sleigh. Book review: Thaddeus William Harris (1795–1856): Nature, Science, and Society in the Life of an American Naturalist. Isis, 100(2):416–417, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605251.

King:2009:BRB

[3588] Nicholas B. King. Book review: Searching Eyes: Privacy, the State, and Disease Surveillance in America. Isis, 100(2):417–419, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605252.

Hardcastle:2009:BRB

[3589] Gary Hardcastle. Book review: The Cambridge Companion to Carnap. Isis, 100(2):419–420, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/605253.

Hentschel:2009:BRBa

[3590] Klaus Hentschel. Book review: Einstein for the Twenty-first Century: His Legacy in Science, Art, and Modern Culture. Isis, 100(2):420–421, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605254.

Dawson:2009:BRB

[3591] Gowan Dawson. Book review: Jane Austen & Charles Darwin: Naturalists and Novelists. Isis, 100(2):421–422, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605255.

Morus:2009:BRB

[3592] Iwan Rhys Morus. Book review: Frankenstein’s Science: Experimentation and Discovery in Romantic Culture, 1780–1830. Isis, 100(2):422– 423, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605256.

Hume:2009:BRB

[3593] Brad D. Hume. Book review: Adam’s Ancestors: Race, Religion, and the Politics of Human Origins. Isis, 100(2):423–424, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605257.

Parsons:2009:BRB

[3594] Charles Parsons. Book review: Logic’s Lost Genius: The Life of Gerhard Gentzen. Isis, 100(2):424–425, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605258.

Williams:2009:BRB

[3595] Elizabeth A. Williams. Book review: Laboratories of Faith: Mesmerism, Spiritism, and Occultism in Modern France. Isis, 100(2):425–426, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/605259.

Heitmann:2009:BRB

[3596] John A. Heitmann. Book review: Fighting Traffic: The Dawn of the Motor Age in the American City. Isis, 100(2):426–427, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605260.

White:2009:BRB

[3597] Paul White. Book review: Samuel Butler, Victorian against the Grain: A Critical Overview. Isis, 100(2):427–428, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605261.

Ramsey:2009:BRB

[3598] Matthew Ramsey. Book review: The Great Nation in Decline: Sex, Modernity, and Health Crises in Revolutionary France, c, by Sean M. Quinlan. 1750–1850. Isis, 100(2):428–429, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605262.

Miller:2009:BRB

[3599] David Philip Miller. Book review: Tides of History: Ocean Science and Her Majesty’s Navy. Isis, 100(2):429–430, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605263.

Bellon:2009:BRB

[3600] Richard Bellon. Book review: Charles Darwin. Isis, 100(2):430–431, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/605264.

Ablard:2009:BRB

[3601] Jonathan D. Ablard. Book review: The Technical Imagination: Argentine Culture’s Modern Dreams. Isis, 100(2):431–432, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605265.

Veneer:2009:BRB

[3602] Leucha Veneer. Book review: Hugh Miller: Stonemason, Geologist, Writer. Isis, 100(2):432–433, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605266.

VanRiper:2009:BRB

[3603] A. Bowdoin Van Riper. Book review: The Legacy of the Mastodon: The Golden Age of Fossils in America. Isis, 100(2):433–434, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605267.

Nieto-Galan:2009:BRB

[3604] Agustí Nieto-Galan. Book review: La pasión científica de un liberal romántico: Lorenzo Gómez Pardo y Ensenyat (1801–1847). Isis, 100(2): 434–435, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/605268.

Rowe:2009:BRB

[3605] David E. Rowe. Book review: Der Erfinder der 5, by Daniela Wuensch. Dimension: Theodor Kaluza, Leben und Werk. Isis, 100(2):435–436, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/605269.

Burnham:2009:BRB

[3606] John C. Burnham. Book review: Psychedelic Psychiatry: LSD from Clinic to Campus. Isis, 100(2):436–437, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605270.

Hardy:2009:BRB

[3607] Anne Hardy. Book review: Brock Chisholm, the World Health Organization, and the Cold War. Isis, 100(2):437–438, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/605271.

Dyck:2009:BRB

[3608] Erika Dyck. Book review: Mania: A Short History of Bipolar Disorder. Isis, 100(2):438–439, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/605272.

Hentschel:2009:BRBb

[3609] Klaus Hentschel. Book review: Spying on Science: Western Intelligence in Divided Germany, 1945–1961. Isis, 100(2):439–440, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605273.

Markowitz:2009:BRB

[3610] Gerald Markowitz. Book review: Bending Science: How Special Interests Corrupt Public Health Research. Isis, 100(2):440–441, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/605274.

Conway:2009:BRB

[3611] Erik M. Conway. Book review: Digital Apollo: Human and Machine in Spaceflight. Isis, 100(2):441–442, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605275.

Schliesser:2009:BRB

[3612] Eric Schliesser. Book review: A Treatise of Human Nature: A Critical Edition. Isis, 100(2):442–444, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605276.

Bouchard:2009:BRB

[3613] Frédéric Bouchard. Book review: Science, Politics, and Evolution. Isis, 100(2):444–445, June 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
605277.

Porter:2009:BRB

[3614] Theodore M. Porter. Book review: Agnotology: The Making and Unmaking of Ignorance. Isis, 100(2):445–446, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/605278.

Stefani:2009:BRB

[3615] Marta Stefani. Book review: Dante e la filosofia della natura: Percezioni, linguaggi, cosmologie. Isis, 100(2):446–447, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605279.

Anonymous:2009:NCb

[3616] Anonymous. Notes on contributors. Isis, 100(2):448–452, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/599543.

Anonymous:2009:Fb

[3617] Anonymous. Frontispiece. Isis, 100(2):vi, June 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/605280.

Rego:2009:PBH

[3618] Brianna Rego. The polonium brief: A hidden history of cancer, radiation, and the tobacco industry. Isis, 100(3):453–484, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/644613.

Richmond:2009:DDR

[3619] Jesse Richmond. Design and dissent: Religion, authority, and the scientific spirit of Robert Broom. Isis, 100(3):485–504, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/644626.

Valverde:2009:SPC

[3620] Nuria Valverde. Small parts: Crisóstomo Martínez (1638–1694), bone histology, and the visual making of body wholeness. Isis, 100(3):505–536, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/644627.

Secord:2009:I

[3621] James A. Secord. Introduction. Isis, 100(3):537–541, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/644629.

Browne:2009:LDP

[3622] Janet Browne. Looking at Darwin: Portraits and the making of an icon. Isis, 100(3):542–570, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/644630.

Clark:2009:YHM

[3623] Constance Areson Clark. “You are here”: Missing links, chains of being, and the language of cartoons. Isis, 100(3):571–589, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/644631.

Smocovitis:2009:SHP

[3624] Vassiliki Betty Smocovitis. Singing his praises: Darwin and his theory in song and musical production. Isis, 100(3):590–614, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/644632.

Anonymous:2009:HSS

[3625] Anonymous. 2008 History of Science Society Prize citations. Isis, 100 (3):615–618, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649135.

Summers:2009:EJS

[3626] William C. Summers. Eloges: Joseph Stewart Fruton, 1912–2007. Isis, 100(3):619–622, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649136.

Buchwald:2009:EMM

[3627] Jed Z. Buchwald and D. Graham Burnett. Eloge: Michael S. Mahoney, 1939–2008. Isis, 100(3):623–626, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649137.

Ruley:2009:LE

[3628] John D. Ruley. Letters to the Editor. Isis, 100(3):627, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649138.

Pisano:2009:LE

[3629] Dominick A. Pisano. Letters to the Editor. Isis, 100(3):627–628, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/649139.

Cooter:2009:LE

[3630] Roger Cooter. Letters to the Editor. Isis, 100(3):628, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649140.

Agar:2009:BRB

[3631] Jon Agar. Book review: The Mantra of Efficiency: From Waterwheel to Social Control. Isis, 100(3):629–630, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/649141.

Topper:2009:BRB

[3632] David R. Topper. Book review: Cosmic Imagery: Key Images in the History of Science. Isis, 100(3):630–631, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/649142.

Fox:2009:BRB

[3633] Robert Fox. Book review: Aurora Torealis: Studies in the History of Science and Ideas in Honor of Tore Frängsmyr. Isis, 100(3):631–632, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649143.

Smith:2009:BRBb

[3634] Hilary A. Smith. Book review: Graphics and Text in the Production of Technical Knowledge in China: The Warp and the Weft, by Francesca Bray; Vera Dorofeeva-Lichtmann; Georges Métailié. (Sinica Leidensia, 79.). Isis, 100(3):632–633, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649144.

Cheung:2009:BRB

[3635] Tobias Cheung. Book review: Les corps vils: Expérimenter sur les êtres humains aux XVIIIe et XIXe siècles. Isis, 100(3):633–634, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649145.

Kingsland:2009:BRB

[3636] Sharon Kingsland. Book review: Rebels, Mavericks, and Heretics in Biology. Isis, 100(3):634–635, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649146.

Mauck:2009:BRB

[3637] Aaron Pascal Mauck. Book review: Mortal Coil: A Short History of Living Longer. Isis, 100(3):635–636, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649147.

Freiburger:2009:BRB

[3638] Dana A. Freiburger and Ronald L. Numbers. Book review: Catholicism and Science, by Peter M. J. Hess; Paul L. Allen; Don O’Leary. (Greenwood Guides to Science and Religion.) Roman Catholicism and Modern Science: A History. Isis, 100(3):636–638, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649148.

Lachapelle:2009:BRB

[3639] Sofie Lachapelle. Book review: L’enseignement secondaire scientifique en France d’un siècle à l’autre, 1802–1980: Évolution, permanences et décalages Les femmes, l’enseignement et les sciences: Un long cheminement (XIXe-XXe siècle). Isis, 100(3):638–639, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649149.

Boothby:2009:BRB

[3640] Thomas Boothby. Book review: The History of the Theory of Structures: From Arch Analysis to Computational Mechanics. Isis, 100(3):639–640, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649150.

Inkster:2009:BRB

[3641] Ian Inkster. Book review: The Applied-Science Problem, by James E. McClellan. (Based on papers presented at a workshop at the Stevens Institute of Technology, 6–8 May 2005). Isis, 100(3):640–641, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649151.

Fritscher:2009:BRB

[3642] Bernhard Fritscher. Book review: Myth and Geology, by L. Piccardi;; W. B. Masse. (Geological Society Special Publications, 273.). Isis, 100 (3):641–642, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649152.

Intemann:2009:BRB

[3643] Kristen Intemann. Book review: Gendered Innovations in Science and Engineering. Isis, 100(3):642–643, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/649153.

Yano:2009:BRB

[3644] Michio Yano. Book review: The Archaic and the Exotic: Studies in the History of Indian Astronomical Instruments. Isis, 100(3):643–644, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649154.

Dym:2009:BRB

[3645] Warren Alexander Dym. Book review: Ein schillerndes Kristall: Eine Wissensgeschichte des Salzes zwischen Früher Neuzeit und Moderne. Isis, 100(3):644–645, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649155.

Brentjes:2009:BRB

[3646] Sonja Brentjes. Book review: O Ye Gentlemen: Arabic Studies on Science and Literary Culture in Honour of Remke Kruk, by Arnoud Vrolijk;; Jan P. Hogendijk. (Islamic Philosophy, Theology, and Science, 74.). Isis, 100(3):645–646, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649156.

Acerbi:2009:BRB

[3647] Fabio Acerbi. Book review: Apollonius of Perga’s Conica: Text, Context, Subtext. Isis, 100(3):646–647, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649157.

Totelin:2009:BRB

[3648] Laurence Totelin. Book review: Galen and the Rhetoric of Healing. Isis, 100(3):647–648, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649158.

Livesey:2009:BRB

[3649] Steven J. Livesey. Book review: Aristote au Mont-Saint-Michel: Les racines grecques de l’Europe chrétienne, by Sylvain Gouguenheim. (L’Univers Historique.). Isis, 100(3):648–650, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649159.

Park:2009:BRB

[3650] Katharine Park. Book review: Midwifery, Obstetrics, and the Rise of Gynaecology: The Uses of a Sixteenth-Century Compendium, by Helen King. (Women and Gender in the Early Modern World.). Isis, 100(3): 650–651, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649160.

Reeds:2009:BRBb

[3651] Karen Reeds. Book review: Used Books: Marking Readers in Renaissance England, by William H. Sherman. (Material Texts.). Isis, 100 (3):651–652, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649161.

McClellan:2009:BRBa

[3652] James E. McClellan, III. Book review: Where Fate Beckons: The Life of Jean–François de la Pérouse, by John Dunmore. (Lives of Great Explorers, 2.). Isis, 100(3):652–653, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649162.

Wilson:2009:BRB

[3653] David B. Wilson. Book review: Academic Patronage in the Scottish Enlightenment: Glasgow, Edinburgh, and St, by Roger L. Emerson. Andrews Universities. Isis, 100(3):653–654, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/649163.

Fraser:2009:BRB

[3654] Craig Fraser. Book review: Infinitesimal Differences: Controversies between Leibniz and His Contemporaries. Isis, 100(3):654–655, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649164.

Henson:2009:BRB

[3655] Pamela Henson. Book review: Everyday Nature: Knowledge of the Natural World in Colonial New York. Isis, 100(3):655–656, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649165.

Eddy:2009:BRB

[3656] M. D. Eddy. Book review: Protogaea. Isis, 100(3):657–658, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649166.

Grave:2009:BRB

[3657] Patricia Radelet de Grave. Book review: Newtons Spirits und der Leibnizsche Raum, by Frank Linhard. (Studien und Materialien zur Geschichte der Philosophie, 77.). Isis, 100(3):658–659, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649167.

Wolfe:2009:BRB

[3658] Charles T. Wolfe. Book review: The Brain Takes Shape: An Early History. Isis, 100(3):659, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649168.

Raposo:2009:BRB

[3659] Pedro Raposo. Book review: Jesse Ramsden (1735- 1800): London’s Leading Scientific Instrument Maker, by Anita McConnell; Lord Martin Rees. (Science, Technology, and Culture, 1700–1945). Isis, 100(3): 659–660, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649169.

Harrison:2009:BRB

[3660] Peter Harrison. Book review: Theology and Science in the Thought of Francis Bacon. Isis, 100(3):660–661, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/649170.

Pitt:2009:BRB

[3661] Joseph C. Pitt. Book review: Reenacting Galileo’s Experiments: Rediscovering the Techniques of Seventeenth-Century Science. Isis, 100 (3):661–662, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649171.

Cohen:2009:BRB

[3662] H. Floris Cohen. Book review: Making Knowledge in Early Modern Europe: Practices, Objects, and Texts, 1400–1800. Isis, 100(3):662–663, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649172.

Mazumdar:2009:BRB

[3663] Pauline M. H. Mazumdar. Book review: Colonial Pathologies: American Tropical Medicine, Race, and Hygiene in the Philippines. Isis, 100(3): 663–664, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649173.

Todes:2009:BRB

[3664] Daniel P. Todes. Book review: Renovating Russia: The Human Sciences and the Fate of Liberal Modernity, 1880–1930. Isis, 100(3):664–665, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649174.

Noakes:2009:BRB

[3665] Richard Noakes. Book review: William Crookes (1832–1919) and the Commercialization of Science, by William H. Brock. (Science, Technology, and Culture, 1700–1945.). Isis, 100(3):665–667, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649175.

Schmaus:2009:BRB

[3666] Warren Schmaus. Book review: Writing the History of the Mind: Philosophy and Science in France, 1900 to 1960s, by Cristina Chimisso. (Science, Technology, and Culture, 1700–1945.). Isis, 100(3):667–668, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649176.

Paradowski:2009:BRB

[3667] Robert J. Paradowski. Book review: Cathedrals of Science: The Personalities and Rivalries That Made Modern Chemistry. Isis, 100(3):668–669, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649177.

Parshall:2009:BRB

[3668] Karen Hunger Parshall. Book review: Mr, by Alex D. D. Craik. Hopkins’ Men: Cambridge Reform and British Mathematics in the Nineteenth Century. Isis, 100(3):669–670, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649178.

Qureshi:2009:BRB

[3669] Sadiah Qureshi. Book review: Darwin’s Sacred Cause: How a Hatred of Slavery Shaped Darwin’s Views on Human Evolution. Isis, 100(3): 670–671, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649179.

Bowler:2009:BRB

[3670] Peter J. Bowler. Book review: H, by Sander Gliboff. G. Bronn, Ernst Haeckel, and the Origins of German Darwinism: A Study in Translation and Transformation. (Transformations: Studies in the History of Science and Technology.). Isis, 100(3):671–672, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649180.

Gearhart:2009:BRB

[3671] Clayton Gearhart. Book review: Max Planck: Die Entstehung der modernen Physik. Isis, 100(3):672–673, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/649181.

Reynolds:2009:BRB

[3672] Andrew Reynolds. Book review: Krebszelle und Zellenstaat: Zur medizinischen und politischen Metaphorik in Rudolf Virchows Zellularpathologie, by Eva Johach. (Berliner Kulturwissenschaft, 5.). Isis, 100(3): 673–674, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649182.

Sommer:2009:BRB

[3673] Marianne Sommer. Book review: Nature’s Clocks: How Scientists Measure the Age of Almost Everything. Isis, 100(3):674–675, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649183.

Paradis:2009:BRB

[3674] Jim Paradis. Book review: The Earth on Show: Fossils and the Poetics of Popular Science, 1802–1856. Isis, 100(3):675–676, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649184.

Oldroyd:2009:BRB

[3675] David Oldroyd. Book review: Finding Time for the Old Stone Age: A History of Palaeolithic Archaeology and Quaternary Geology in Britain, 1860–1960. Isis, 100(3):676–677, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649185.

Munz:2009:BRB

[3676] Tania Munz. Book review: The Simian Tongue: The Long Debate about Animal Language. Isis, 100(3):677–679, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649186.

Shulman:2009:BRB

[3677] Peter Shulman. Book review: Power Struggles: Scientific Authority and the Creation of Practical Electricity before Edison. Isis, 100(3):679–680, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649187.

Edgerton:2009:BRB

[3678] David Edgerton. Book review: Science, War, and Imperialism: India in the Second World War, by Jagdish N. Sinha. (Social Sciences in Asia, 18.). Isis, 100(3):680–681, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649188.

Gregory:2009:BRB

[3679] Frederick Gregory. Book review: Asklepios und die Philosophen: Paradigmawechsel in der Medizin im 19, by Nelly Tsouyopoulos; Claudia Wiesemann; Barbara Bröker; Sabine Rogge; Claudia Wiesemann. Jahrhundert. Isis, 100(3):681–682, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/649189.

Kroker:2009:BRB

[3680] Kenton Kroker. Book review: Biomedicine in the Twentieth Century: Practices, Policies, and Politics, by Caroline Hannaway. (Biomedical and Health Research, 72.). Isis, 100(3):682–683, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649190.

Hong:2009:BRB

[3681] Sungook Hong. Book review: Ham Radio’s Technical Culture, by Kristen Haring. (Inside Technology.). Isis, 100(3):683–684, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649191.

Dussauge:2009:BRB

[3682] Isabelle Dussauge. Book review: Magnetic Appeal: MRI and the Myth of Transparency. Isis, 100(3):684–685, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/649192.

Dick:2009:BRB

[3683] Steven J. Dick. Book review: Keep Watching the Skies! The Story of Operation Moonwatch and the Dawn of the Space Age. Isis, 100(3): 685–686, September 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
649193.

McCray:2009:BRB

[3684] W. Patrick McCray. Book review: Earthrise: How Man First Saw the Earth. Isis, 100(3):686–687, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649194.

Cottebrune:2009:BRB

[3685] Anne Cottebrune. Book review: History and Theory of Human Experimentation: The Declaration of Helsinki and Modern Medical Ethics, by Ulf Schmidt; Andreas; Frewer. (Geschichte und Philosophie der Medizin, 2.). Isis, 100(3):687–688, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/649195.

Dror:2009:BRB

[3686] Otniel E. Dror. Book review: Shock Therapy: A History of Electroconvulsive Treatment in Mental Illness. Isis, 100(3):688–690, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649196.

Foster:2009:BRB

[3687] Jay Foster. Book review: Thomas Kuhn’s Revolution: An Historical Philosophy of Science. Isis, 100(3):690–691, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649197.

Sismondo:2009:BRB

[3688] Sergio Sismondo. Book review: Knowledge as Social Order: Rethinking the Sociology of Barry Barnes. Isis, 100(3):691, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649198.

Ziegler:2009:BRB

[3689] Rafael Ziegler. Book review: Historische Epistemologie zur Einführung. Isis, 100(3):692, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/649199.

Keller:2009:BRB

[3690] Evelyn Fox Keller. Book review: Becoming Beside Ourselves: The Alphabet, Ghosts, and Distributed Human Being. Isis, 100(3):692–693, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/649200.

Anonymous:2009:NCc

[3691] Anonymous. Notes on contributors. Isis, 100(3):694–698, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/649201.

Anonymous:2009:Fc

[3692] Anonymous. Frontispiece. Isis, 100(3):vi, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/649202.

Lucier:2009:PSN

[3693] Paul Lucier. The professional and the scientist in Nineteenth-Century America. Isis, 100(4):699–732, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652016.

Manias:2009:RPC

[3694] Chris Manias. The race prussienne controversy: Scientific internationalism and the nation. Isis, 100(4):733–757, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652017.

Schlich:2009:PML

[3695] Thomas Schlich. The perfect machine: Lorenz Böhler’s rationalized fracture treatment in World War I. Isis, 100(4):758–791, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652018.

White:2009:I

[3696] Paul White. Introduction. Isis, 100(4):792–797, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652019.

Alberti:2009:BHM

[3697] Fay Bound Alberti. Bodies, hearts, and minds: Why emotions matter to historians of science and medicine. Isis, 100(4):798–810, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652020.

White:2009:DES

[3698] Paul White. Darwin’s emotions: The scientific self and the sentiment of objectivity. Isis, 100(4):811–826, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652021.

Hayward:2009:EEJ

[3699] Rhodri Hayward. Enduring emotions: James L. Halliday and the invention of the psychosocial. Isis, 100(4):827–838, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652022.

Wilson:2009:WHM

[3700] Elizabeth A. Wilson. “Would I had him with me always”: Affects of longing in early artificial intelligence. Isis, 100(4):839–847, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652023.

Dror:2009:ARF

[3701] Otniel E. Dror. Afterword: A reflection on feelings and the history of science. Isis, 100(4):848–851, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652024.

Baldwin:2009:EMM

[3702] Mary E. Baldwin. Eloge: Marianne (Marika) Gosztonyi Ainley, 1937– 2008. Isis, 100(4):852–855, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652025.

Burian:2009:EMG

[3703] Richard Burian and Roger Ariew. Eloge: Marjorie Glicksman Grene, 1910–2009. Isis, 100(4):856–859, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652026.

Eastwood:2009:LE

[3704] Bruce Eastwood. Letters to the Editor. Isis, 100(4):860, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652027.

Zuccato:2009:LE

[3705] Marco Zuccato. Letters to the Editor. Isis, 100(4):860–861, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652028.

Carey:2009:LE

[3706] Toni Vogel Carey. Letters to the Editor. Isis, 100(4):861–862, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652029.

Hopwood:2009:BRB

[3707] Nick Hopwood. Book review: The Tragic Sense of Life: Ernst Haeckel and the Struggle over Evolutionary Thought. Isis, 100(4):863–867, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/652030.

Pickering:2009:BRB

[3708] Andrew Pickering. Book review: The Scientific Life: A Moral History of a Late Modern Vocation. Isis, 100(4):868–871, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652031.

Taylor:2009:BRBb

[3709] Kenneth L. Taylor. Book review: Worlds Before Adam: The Reconstruction of Geohistory in the Age of Reform. Isis, 100(4):872–880, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652032.

Danziger:2009:BRB

[3710] Kurt Danziger. Book review: Psychology’s Territories: Historical and Contemporary Perspectives from Different Disciplines. Isis, 100(4):881– 882, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652033.

Carruthers:2009:BRB

[3711] Mary Carruthers. Book review: Marking the Mind: A History of Memory. Isis, 100(4):882–883, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652034.

Ruse:2009:BRBb

[3712] Michael Ruse. Book review: Critique of Intelligent Design: Materialism versus Creationism from Antiquity to the Present. Isis, 100(4):883–884, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652035.

Clutton-Brock:2009:BRB

[3713] Martin Clutton-Brock. Book review: How We Got Here: From Bows and Arrows to the Space Age. Isis, 100(4):884–885, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652036.

Jackson:2009:BRB

[3714] John P. Jackson, Jr. Book review: The Scientific Literature: A Guided Tour. Isis, 100(4):885–886, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652037.

Dijksterhuis:2009:BRB

[3715] Fokko Jan Dijksterhuis. Book review: Unsichtbares Licht? Dunkle Wärme? Chemische Strahlen? Eine wissenschaftshistorische und -theoretische Analyse von Argumenten für das Klassifizieren von Strahlungssorten 1650–1925 mit Schwerpunkt auf den Jahren 1770–1900. Isis, 100(4):886–887, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652038.

Fehige:2009:BRB

[3716] Yiftach J. H. Fehige. Book review: Sexualized Brains: Scientific Modeling of Emotional Intelligence from a Cultural Perspective. Isis, 100(4):887– 888, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652039.

Mayer:2009:BRB

[3717] Andreas Mayer. Book review: The Sleep of Others and the Transformations of Sleep Research. Isis, 100(4):888–889, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652040.

Kidwell:2009:BRB

[3718] Peggy Aldrich Kidwell. Book review: Reinventing Knowledge: From Alexandria to the Internet. Isis, 100(4):889–890, December 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652041.

Asaro:2009:BRB

[3719] Peter Asaro. Book review: The Robot: The Life Story of a Technology. Isis, 100(4):890–891, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652042.

Simoes:2009:BRB

[3720] Ana Simões. Book review: Beyond Borders: Fresh Perspectives in History of Science. Isis, 100(4):891–892, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/652043.

Igo:2009:BRB

[3721] Sarah E. Igo. Book review: Smoot’s Ear: The Measure of Humanity. Isis, 100(4):892–893, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652044.

Scarborough:2009:BRB

[3722] John Scarborough. Book review: The Cambridge Companion to Galen. Isis, 100(4):893–895, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652045.

Sider:2009:BRB

[3723] David Sider. Book review: The Encyclopedia of Ancient Natural Scientists: The Greek Tradition and Its Many Heirs. Isis, 100(4):895–896, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652046.

Montelle:2009:BRB

[3724] Clemency Montelle. Book review: Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Near Eastern Societies. Isis, 100(4):896–897, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652047.

Evans:2009:BRB

[3725] James Evans. Book review: Aetna and the Moon: Explaining Nature in Ancient Greece and Rome. Isis, 100(4):897–898, December 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652048.

Lo:2009:BRB

[3726] Vivienne Lo. Book review: A Dictionary of the Huang Di nei jing su wen, by Hermann Tessenow; Paul U. Unschuld. (Huang Di nei jing su wen Project). Isis, 100(4):898–900, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652049.

Hogendijk:2009:BRB

[3727] Jan P. Hogendijk. Book review: Aspects of the Astrolabe: “Architectonica Ratio” in Tenth- and Eleventh-Century Europe, by Arianna Borrelli. (Sudhoffs Archiv, 57). Isis, 100(4):900–901, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652050.

Klestinec:2009:BRB

[3728] Cynthia Klestinec. Book review: Le scienze, by Antonio Clericuzio; Germana Ernst; Maria Conforti. (Il Rinascimento Italiano e l’Europa, 5.). Isis, 100(4):901–902, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652051.

Nuti:2009:BRB

[3729] Lucia Nuti. Book review: Terra Incognita: Mapping the Antipodes before 1600. Isis, 100(4):902–903, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652052.

Bohlandt:2009:BRB

[3730] Marco Böhlandt. Book review: Astrolabes and Angels, Epigrams and Enigmas: From Regiomontanus’ Acrostic for Cardinal Bessarion to Piero della Francesca’s Flagellation of Christ. Isis, 100(4):903–904, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652053.

Rosenberg:2009:BRB

[3731] Daniel Rosenberg. Book review: Diderot: Natural Philosopher, by Kurt Ballstadt. (Studies on Voltaire and the Eighteenth Century.). Isis, 100 (4):905–906, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652054.

Domingues:2009:BRB

[3732] Beatriz Domingues. Book review: Más allá de la Leyenda Negra: España y la Revolutión Científica/Beyond the Black Legend: Spain and the Scientific Revolution. Isis, 100(4):906–907, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652055.

MacLeod:2009:BRB

[3733] Roy MacLeod. Book review: Science and Empire in the Atlantic World. Isis, 100(4):907–908, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652056.

Henry:2009:BRBb

[3734] John Henry. Book review: The Essential Galileo. Isis, 100(4):908–909, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652057.

DalPrete:2009:BRB

[3735] Ivano Dal Prete. Book review: From Makers to Users: Microscopes, Markets, and Scientific Practices in the Seventeenth and Eighteenth Centuries/Dagli artigiani ai naturalisti: Microscopi, offerta dei mercati e pratiche scientifiche nei secoli XXVII e XVIII, by Dario Generali; Marc J. Ratcliff. (Edizione Nazionale delle Opere di Antonio Vallisneri, 3.). Isis, 100(4):909–910, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652058.

Delisle:2009:BRB

[3736] Candice Delisle. Book review: Conrad Gessner’s Private Library, by Urs B. Leu; Raffael Keller; Sandra; Weidmann. (History of Science and Medicine Library, 5.). Isis, 100(4):910–911, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652059.

McClellan:2009:BRBb

[3737] James E. McClellan III. Book review: Measuring the New World: Enlightenment Science and South America. Isis, 100(4):911–912, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652060.

Wilding:2009:BRB

[3738] Nick Wilding. Book review: Galileo’s Inquisition Trial Revisited. Isis, 100 (4):912–913, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652061.

Osler:2009:BRBb

[3739] Margaret J. Osler. Book review: Mémoire de Gassendi: Vies et célébrations écrites avant 1700, by Sylvie Taussig; Anthony Turner. (Les Styles du Savoir: Défense et Illustration de la Pensée à l’âge Classique.). Isis, 100(4):913–914, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/652062.

Rappaport:2009:BRBb

[3740] Rhoda Rappaport. Book review: Science et théologie dans les débats savants de la seconde moitié du XVIIe siècle: La genèse dans les Philosophical Transactions et le Journal des savants (1665–1710). Isis, 100 (4):914–915, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652063.

Kassler:2009:BRB

[3741] Jamie C. Kassler. Book review: Music, Experiment, and Mathematics in England, 1653–1705. Isis, 100(4):915–916, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652064.

Kohlstedt:2009:BRB

[3742] Sally Gregory Kohlstedt. Book review: William Barton Rogers and the Idea of MIT. Isis, 100(4):916–917, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/652065.

Richmond:2009:BRBb

[3743] Marsha Richmond. Book review: Darwins Korallen: Die frühen Evolutionsdiagramme und die Tradition der Naturgeschichte. Isis, 100(4):917– 918, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652066.

Pearl:2009:BRB

[3744] Sharrona Pearl. Book review: Science and Eccentricity: Collecting, Writing, and Performing Science for Early Nineteenth-Century Audiences, by Victoria Carroll. (Science and Culture in the Nineteenth Century, 4.). Isis, 100(4):918–919, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652067.

Fichman:2009:BRB

[3745] Martin Fichman. Book review: Evolution: Selected Letters of Charles Darwin, 1860–1870. Isis, 100(4):919–920, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/652068.

Reidy:2009:BRB

[3746] Michael S. Reidy. Book review: Imperial Nature: Joseph Hooker and the Practices of Victorian Science. Isis, 100(4):920–921, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652069.

Schickore:2009:BRB

[3747] Jutta Schickore. Book review: Johannes Müller und die Pathologische Anatomie: Eine kommentierte Edition der Vorlesungsmitschrift von Jakob Henle (1830), by Ildikó Gágyor. (Studies on History, Theory, and Ethics of Medicine, 17.). Isis, 100(4):922, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652070.

Israel:2009:BRB

[3748] Paul Israel. Book review: Domesticating Electricity: Technology, Uncertainty, and Gender, 1880–1914, by Graeme Gooday. (Science and Culture in the Nineteenth Century, 7.). Isis, 100(4):923–924, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652071.

Brain:2009:BRB

[3749] Robert Michael Brain. Book review: Dances with Darwin, 1875–1910: Vernacular Modernity in France. Isis, 100(4):924–925, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652072.

Shell-Gellasch:2009:BRB

[3750] Amy Shell-Gellasch. Book review: Pioneering Women in American Mathematics: The Pre-1940 Ph.D.’s, by Judy Green; Jeanne LaDuke (History of Mathematics, 34). Isis, 100(4):925–926, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652073.

McOuat:2009:BRB

[3751] Gordon McOuat. Book review: What I Require from Life: Writings on Science and Life from J. Isis, 100(4):926–927, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652074.

Gordon:2009:BRB

[3752] David Gordon. Book review: Butterflies and Barbarians: Swiss Missionaries and Systems of Knowledge in South-East Africa. Isis, 100(4):927– 928, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652075.

Beyerchen:2009:BRB

[3753] Alan Beyerchen. Book review: Inside IG Farben: Hoechst during the Third Reich. Isis, 100(4):928–929, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/652076.

Cantor:2009:BRB

[3754] Geoffrey Cantor. Book review: The Victorian Eye: A Political History of Light and Vision in Britain, 1800–1910. Isis, 100(4):929–930, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652077.

Anderson:2009:BRB

[3755] Warwick Anderson. Book review: The Making of a Tropical Disease: A Short History of Malaria. Isis, 100(4):930–931, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652078.

Burba:2009:BRB

[3756] Juliet Burba. Book review: Anthropology Goes to the Fair: The 1904 Louisiana Purchase Exposition. Isis, 100(4):931–933, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652079.

Gossin:2009:BRB

[3757] Pamela Gossin. Book review: H, by John S. Partington. G. Wells in Nature, 1893–1946: A Reception Reader. (Arbeiten zur Literarischen Phantastik, 3.). Isis, 100(4):933–934, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652080.

Rothenberg:2009:BRB

[3758] Marc Rothenberg. Book review: The Passionate Empiricist: The Eloquence of John Quincy Adams in the Service of Science. Isis, 100(4): 934–935, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652081.

Mooney:2009:BRB

[3759] Graham Mooney. Book review: Typhoid in Uppingham: Analysis of a Victorian Town and School in Crisis, 1875–1877, by Nigel Richardson. (Science and Culture in the Nineteenth Century, 5.). Isis, 100(4):935– 936, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652082.

Goodrum:2009:BRB

[3760] Matthew R. Goodrum. Book review: From Genesis to Prehistory: The Archaeological Three Age System and Its Contested Reception in Denmark, Britain, and Ireland, by Peter Rowley-Conwy. (Oxford Studies in the History of Archaeology.). Isis, 100(4):936–937, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652083.

Schuring:2009:BRB

[3761] Michael Schüring. Book review: Kontinuitäten und Diskontinuitäten in der Wissenschaftsgeschichte des 20, by Rüdiger vom Bruch; Uta Gerhardt; Aleksandra Pawliczek. Jahrhunderts. (Wissenschaft, Politik und Gesellschaft, 1.). Isis, 100(4):937–938, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652084.

Greene:2009:BRBb

[3762] Mott Greene. Book review: Gesammelte Schriften. Isis, 100(4):938–939, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652085.

Stadler:2009:BRB

[3763] Friedrich Stadler and Donata Romizi. Book review: Wissenschaftslandschaften um 1900: Philosophie, die Wissenschaften und der nichtreduktive Szientismus, by Paul Ziche. (Legierungen, 3.). Isis, 100(4):939–941, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652086.

Cooter:2009:BRBb

[3764] Roger Cooter. Book review: The Collectors of Lost Souls: Turning Kuru Scientists into Whitemen. Isis, 100(4):941–942, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652087.

Kursell:2009:BRB

[3765] Julia Kursell. Book review: Mechanical Sound: Technology, Culture, and Public Problems of Noise in the Twentieth Century. Isis, 100(4):942–943, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652088.

Stamhuis:2009:BRB

[3766] Ida Stamhuis. Book review: Darwins kleinkinderen: De evolutietheorie in België, 1865–1945. Isis, 100(4):943–944, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652089.

Creager:2009:BRB

[3767] Angela N. H. Creager. Book review: Promising Genomics: Iceland and deCODE Genetics in a World of Speculation. Isis, 100(4):944–945, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/652090.

Martin:2009:BRB

[3768] Terry Martin. Book review: The Plot to Kill God: Findings from the Soviet Experiment in Secularization. Isis, 100(4):945–947, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652091.

Kojevnikov:2009:BRB

[3769] Alexei Kojevnikov. Book review: Science in the New Russia: Crisis, Aid, Reform. Isis, 100(4):947–948, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652092.

Sellers:2009:BRB

[3770] Christopher Sellers. Book review: Poison in the Well: Radioactive Waste in the Oceans at the Dawn of the Nuclear Age. Isis, 100(4):948–949, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652093.

Strick:2009:BRB

[3771] James Strick. Book review: Alien Ocean: Anthropological Voyages in Microbial Seas. Isis, 100(4):949–950, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/652094.

Weidman:2009:BRB

[3772] Nadine Weidman. Book review: Medicine and Modernism: A Biography of Sir Henry Head, by L. S. Jacyna. (Science and Culture in the Nineteenth Century, 6.). Isis, 100(4):950–951, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652095.

Nye:2009:BRB

[3773] David E. Nye. Book review: Invented Edens: Techno-Cities of the Twentieth Century, by Robert H. Kargon; Arthur P. Molella. (Lemelson Center Studies in Invention and Innovation.). Isis, 100(4):951–952, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652096.

Pandora:2009:BRB

[3774] Katherine Pandora. Book review: Science on the Air: Popularizers and Personalities on Radio and Early Television. Isis, 100(4):952–953, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/652097.

Ruswick:2009:BRB

[3775] Brent Ruswick. Book review: Three Generations, No Imbeciles: Eugenics, the Supreme Court, and Buck v, by Paul A. Lombardo. Bell. Isis, 100 (4):953–954, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652098.

Badash:2009:BRB

[3776] Lawrence Badash. Book review: The Atomic Bomb and American Society: New Perspectives. Isis, 100(4):954–955, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652099.

Kline:2009:BRB

[3777] Wendy Kline. Book review: The Fertility Doctor: John Rock and the Reproductive Revolution. Isis, 100(4):955–956, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652100.

Barnett:2009:BRBb

[3778] Richard Barnett. Book review: Racial Science and British Society, 1930–62. Isis, 100(4):956–957, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652101.

Pennock:2009:BRB

[3779] Robert T. Pennock. Book review: Rescuing Science from Politics: Regulation and the Distortion of Scientific Research. Isis, 100(4):957–958, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652102.

Tauber:2009:BRB

[3780] Alfred Tauber. Book review: Knowledge of Life. Isis, 100(4):958–959, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652103.

Verbeek:2009:BRB

[3781] Theo Verbeek. Book review: Cartesian Reflections: Essays on Descartes’s Philosophy. Isis, 100(4):959–960, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652104.

Tosh:2009:BRB

[3782] Nick Tosh. Book review: Historical Judgement: The Limits of Historiographical Choice. Isis, 100(4):960–961, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652105.

Churchland:2009:BRB

[3783] Paul Churchland. Book review: Creating Scientific Concepts. Isis, 100 (4):961–963, December 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652106.

Fuller:2009:BRBb

[3784] Steve Fuller. Book review: After the Open Society: Selected Social and Political Writings. Isis, 100(4):963–964, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652107.

Cat:2009:BRB

[3785] Jordi Cat. Book review: The Cambridge Companion to Logical Empiricism. Isis, 100(4):964–965, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652108.

Anonymous:2009:NCd

[3786] Anonymous. Notes on contributors. Isis, 100(4):966–971, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652109.

Anonymous:2009:IBV

[3787] Anonymous. Index to Isis vol 100, 2009. Isis, 100(4):973–986, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652110.

Anonymous:2009:CBV

[3788] Anonymous. Contents of Isis vol 100, 2009. Isis, 100(4):989–991, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/652192.

Anonymous:2009:Fd

[3789] Anonymous. Frontispiece. Isis, 100(4):vii, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652111.